A Complete Bibliography of Publications in SIGCSE Bulletin: 2000–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

21 April 2018 Version 1.08

Title word cross-reference

0 [Gin09b]. 0/1 [Rol07]. 1 [Gin09b]. 2 [Hu05, LM05]. 3 [CPJ07, CF02, Sei09]. st [Mar05c, Sal03a, Laz02b]. ^{th} [Laz02b]. λ [CKMW02]. μ [GDM05]. N [Rol08b, Rol06b]. O(1) [SJ05, SJ06]. O(n) [SJ05, SJ06]. τϵ́χνη [DGMW04]. τϵχνη [DGMW07]. V^{2} [DG05].

*is* [Spa04].

-queens [Rol06b, Rol08b].

.coms [Mar01a]. .Ethics [Mar01a]. .NET [AO06b, GS05a, Gui05].

/Java [Kum04, Kum05b].

0 [AFJ^{+}03]. 05 [Chr05b].

1 [AF03b, BDM01a, Che00, DLHAK08, WAC^{+}08]. 1-bit [BSV04]. 10 [HB01]. 10th [CFPL05]. 11/10 [HB01]. 11th [DGS06]. 12 [BT06, Goo08, HEB08, OM08, SLPG07]. 12th [HPT07]. 13th [ALRY08].

2000 [CDWH00, Ell00b, MvW01]. 2001 [FKCM01, RCE^{+}00, RCD^{+}02, WMGR01]. 2002 [CJGU02]. 2003 [Ano03, DSF^{+}03, GKJD03]. 2004 [BCK04, JKDN04]. 2005 [CFPL05, DNTB05, SMS^{+}06]. 2006 [ACM06, BTHR06, DGS06]. 2007 [HPT07, LB08b]. 2008 [ALRY08, DRFG08]. 21st [BDRH09, Laz02a, Lee01b, Mar01b, Sal03b]. 2nd [MAFLR05, MAFLR06]. 2nd-year [MAFLR05, MAFLR06].

3 [BN09, MMBAS04]. 32rd [WMGR01]. 34th [GKJD03]. 35th [Imp04]. 39th [BTHR06, DRFG08].

40th [ADM^{+}09]. 4th [Ell00b].

6th [FKCM01, Fin01].

7th [CJGU02, Fin02, KB08].

8 [BT06]. 8th [DSF^{+}03, Fin03].

99 [CLK00]. 9th [BCK04].

= [Jen01a, Reg08].

Aarhus [CJGU02, Fin02]. AARTIC [MR09]. abacus [Pop03b]. abbreviated [Reb05]. abduction [SA07]. abductive [Rad07]. ABET [Gor03a, Lea08, LSMB00, San09a]. ABET/CSAB [LSMB00]. abilities [CLM^{+}07]. ability [BC06a, CB00, CJ01b, DW04]. above [Yeh03]. abroad [FAL^{+}05]. Abstract [BGG^{+}00, AGE06, PRK05, PP07, Ste04]. Abstraction [BC06a, Hen08a, SD09, BLW01, Ens09, Haz02, Haz08, Kop01, MVD01, OBL04, PGK05, PK06, Sch09b, Soo00, Soo09, WSP04]. abstractions [Bru05b]. abuse [God06]. Academia [RRD^{+}09, DJS08, YH05]. Academia-academia-industry [RRD^{+}09]. Academic [Ber05, PHZ^{+}03, Wal04b, Con04, CT09, CLP^{+}01, Joy06b, MNN^{+}09, MA01, PM02b, Rag04, RMPD09, Wal01a]. academic-strength [Con04]. academically [CR02]. academically-diverse [CR02]. academics [GPP00]. Academy [CLR09, Cle04a, MTBS09]. accelerated [Reb05]. acceptance [LBA07]. accepted [BZ04a]. access [Bor07, GP02, GL05, Hil04, Kes04, Rol03a, WGS05, YH04]. Accessibility [BP07, CFGL05, RDCE06, Ros06, CHUB08, FS02, FdMFPT07]. Accessible [MMRI07, LC08, McC01b, MICM08, PJ01, Rob00c, Ros07, SBR^{+}07].

Accommodating [PRW04]. accord [Tur09]. account [GLC07]. accountability [How05]. Accreditation [Kac01b, Kac01a, CS03a, Gor00a, GLF01, Gor03a, Gor03b, Gor04a, Gor05, Gor06b, Har01b, LLP04, Lea08, LKGY00, LFGY02, Top09a, YLRZ07, ZRY05, ZRY06]. accumulation [Gin07a]. ACE [Dal01]. achievement [Gib00, LML07, Pop05a, RHY07, Rob00e, RBJ^{+}09]. achievements [AGE04, BZ04a]. achievent [EH07]. achievent-related [EH07]. ACM [Ell00b, AHL04, BT06, Cam03a, CHK03a, CHK05, CHK06, CCD^{+}08b, DCP^{+}09, GOB^{+}07, GC02, HCK07, IHL^{+}04, MMLT09, Pre01, RCE^{+}00, RBC^{+}05, RBC^{+}06, SL03]. ACM-W [GOB^{+}07, GC02]. ACM/IEEE [MMLT09, RCE^{+}00]. ACM/IEEE-CS [MMLT09, RCE^{+}00]. acquisition [App01, MGH^{+}06]. across [ABE^{+}09b, AHJ^{+}07, BBW04, CS05, Lee00b, Lee01a, LCPQR05, WSW05, WR07]. ACSE [Ell00b]. act [Pat06, Wal02b]. Action [HLL03, Edw04, FB02]. actions [Boe06]. Activating [HK01]. Active [Lau07, Lud05, McC05b, McC05a, McC06c, McC06a, McC06d, McC06b, VIPC09, AAD^{+}07, BGGR01, Bud06, Cas02, CS08, CMS06c, GM09, Gon06, GKL^{+}02, GS06b, Ham07, LDK^{+}07, LNR05, PBK04, Qua06, Rad09, STTBC07, SB07a, SFPFV08, SAHS04, Wir03]. Active-learning [Lud05, GM09]. actively [NEN00]. Activities [CHK06, Bow07a, Bur08, Cam05b, CBJL08, CMCB09, Kle01, KB01, LLP04, Lud05, LR08, Mor08, MKLH09, OBL04, SP07, SSP^{+}05a, SM05a, SM05b, Tur07a, Tur07b]. activity [AMW06, Cam04b, CCM07, Par06b, TBM09]. activity-based [TBM09]. actors [KT01]. actual [HL04b]. Ada [Fag00]. Ada-based [Fag00]. adapt [BAdC07]. Adaptation [Las09, BdCGB07, CCK01]. Adapting [PG07]. adaptive [AEPQ06, BL06b, DWT00]. added [ZJ03]. adder [BSV04]. Adding [AHL04, BMP06]. address [BU09, HILS05, RPG04, Rob05, Rob03b, Tra03]. Addressing [DSB^{+}03, JY04, Cle02b, HRBC03]. adjunct [KCF^{+}00]. administer [TT06]. administration [Bor07, Cig03, Dav04, RN03]. admissions [ACL^{+}03]. adolescent [Gor02]. adopt [To04]. Adopting [CS08]. adoption [JS08a, LBA08, Ni09]. ’ADT [BS09]. adults [GY03]. advance [KK09]. Advanced [AWS^{+}09, BZ05, UD01, Wis05, AFJ^{+}03, Ari00, ACK^{+}00, BZ07, HCHP01, Nør00, OHRL01, PY05, Pop09b, SL08b, Ska02, SHP^{+}06]. Advancing [CLR09]. advantage [BT01, BP02, CMT04a, CGT05, PMT03]. advantages [CS03a]. adventure [BL06a]. adventures [Gig02, Lab09]. Advertising [Wal08b]. advice [HMS^{+}09, KCF^{+}00, Pop03b]. affect [HY09, PR02]. affecting [GR00a, JY01b, PJ05, To04]. Affective [MD05, RBJ^{+}09, Lew07b, MD04]. affects [CMD08]. afraid [GGH03]. Africa [GM07]. African [GSyC07, LS02, PV08]. after [Cli04, Gor00b, Got09b, SK09, SKM02, Vas02]. after-the-fact [SKM02]. again [DSE02, GCD^{+}09, Kum09d, MRGS08, Rol05]. against [BCD08]. Age [NL04, BHM^{+}06b]. Agent [SSC00a, SJA04, Par02a, SSC00b]. Agent-based [SJA04]. Agents [BL00, HSM06, HA04, MLH05, Pop09b]. Agile [BKR^{+}05, McC01a, Cle03a, Cle04b, Got04, KPCM06, KD07, LCW06, MD05]. agility [HWRC08]. agonies [Edm07a]. agreement [RP06]. Aha [Gin02a]. AI [AAC^{+}05, Can06, KF03, Kum05a, KKR06, MT09, Pop09b, Rao05, RM08a, RMN06, Sie00]. aid [Ext00, Geo02, Red01]. aided [Jen08, Rol03b, SL00b, WP06b]. aim [WVMN05]. aimed [LR01]. aiming [IMRJ09]. AJAX [DV07]. Alan [Sor05]. Alfonse [Har00, Har01a]. algebra [MG00]. Algebraic [GET06]. Algorithm [And00, CPJ07, KM00, SCE07, Arm08, Boz06, Cha04b, CWHL06, CKMW02, CN07, GEVZ03, Gin08b, HAG05, HTK02, HLN07, JFH00, KMS01, KLOT05, Lau06, MH08, MK00, MLK07, NEN00, NRB^{+}05, PP03b, PGK05, PK06, PCW09, Rol02, RSF00, RF01, RN02, RKS07, SSMN04, TBAK04, VIPCUF08, WP02]. AlgorithmA [BFMC05, CLK00]. Algorithmic [Gin04a, Ast03, CWHH05, CKW04, GGAB01, Gin01c, Gin01d, KB08, KB09, MH05, RD00]. algorithmics [GT01]. Algorithms [BHLW02, Sch02b, AF03a, And00, ACSS06, Arm06, Ber09b, BN00, CPJ07, Chi05, CSM07b, CM01, Cle07b, D’A03, GMFA09, GOGPGP08, HKL^{+}02, HYN03, Kes04, KH00, KH01, KPT00, Kos00, Lan01, Lev00, LP02, Lev05, Mér09b, MT09, NM02, OG05, PP03a, Pap03c, PPRV06, RL09, STTBC07, San02, She03b, SP07, Sna06b, Sok02, Sol07b, SN02, Sur01, VI01, VIPC09, WB06, Wu05]. AlgoVista [CKW04]. Alice [Ada07, CCM07, LS08c, Pau08, PEH07, RHL^{+}09, SL09b, Sei09]. all-female [SM02a]. alliances [Pop07a]. allocation [Wal09b]. allowing [WG07]. almost [MHA04]. alouds [Ars09]. also [FLS04b, Liu06, Pow03]. alterations [Kac09b]. Alternating [Gin07a]. alternative [HS02, KC07, LB00, Mos06, RP04, Rol07, Rou04]. Alternatives [Oka05]. Aluminum [YM06]. always [GGH03]. AM [BGL^{+}07]. amateur [Sko07]. Amazon [BA06]. ambient [BC07e, Bur06]. American [Dav01a, KTH^{+}00, Kla05a, LS02]. among [AEH05, AL09, BGD03b, Lew07a, Lin05, LZL08, MGS07, MJ03, Per09]. amortised [Bla09d]. amortised-linear [Bla09d]. Analogies [NLMN06]. analogy [Mas04]. Analysis [Lea08, MAFO08, Tra03, AR00a, AEH05, Arm06, Ast03, BGNM04, Bow07b, Cha00b, Cha04b, Dic05, ESS^{+}07, EH05, Ham05, Hoo08, Jin08, KB08, Lev00, Lev05, LB08b, MDM04, Mir09, MS03, MLM^{+}08, NH05, Qui06, RD08, RKD02, RK09a, San02, SB07b, TG05, UD03, Val04, WLCL08, WW06b, dBGdMB09]. analytic [Rol03b]. Analyze [Lev05, BL05b, Fon03]. Analyzing [AK09, Han09, MLK07, Ode00, Rus09, SSP^{+}05a, AF03a, GBO08, LL00, LCS^{+}09, SSP^{+}05c]. Anchor [MSK09]. Ancient [GH01]. Andes [YM06]. Andragogy [Ell02]. ANIMAL [RSF00]. AnimalScript [RF01]. animated [DCP00]. Animating [Gri02c]. Animation [DMU02, Mos06, NRRR06, AFJ^{+}03, Boi06, CM01, JFH00, KPT00, KT01, KMS01, Kum09a, MMBAS04, Mor08, Rod02, RSF00, RF01, RKS07, TLH^{+}09, VIPCUF08, McD02]. Animations [Soj03a, ABL07, Hu03, HLN07, RF00b, UFVI05]. animator [Gri03]. anniversary [ADM^{+}09]. Annotation [Ran00, RLM08, SKM02, WC04]. annotator [RJ03]. Annual [ALRY08, BCK04, CJGU02, CFPL05, DSF^{+}03, DGS06, FKCM01, Fin01, Fin02, Fin03, HPT07, CPP07]. anonymousness [GG06b]. answer [HW08]. answering [Cle09a, RLM08]. anthropology [KBAP00]. anthropomorphic [Sav05]. Antipodal [Got01, Got05]. ANY [Ral05]. anymore [BA04b, BA05a]. anyone [Wil04a]. AP [OHRL01, DHRH05, GLR^{+}01, HHRA06, TW09b]. APEINTA [IMRJ09]. API [Gag02, MWM01, MS05b]. APOGEE [FPQ^{+}08]. Apoo [RM01]. apparent [Car06]. applet [EGST08, Hu03, Rob07b]. Applets [Hei03]. Application [Boz06, CF02, Haz03a, Ols05, Vau00, Bac03, Bri06, BBS06, CDF^{+}07, CSD^{+}07, GMGM06, Had01, Kag09, MPJC09, NRDP07, PC08, Rol08b, SGMJ09, SL06, Shi06, Tre02, Wan06, WP02, YD04, ZLL^{+}07]. application-driven [NRDP07]. applications [CCC07, CWHL06, DEJ^{+}06, DJLA07, Gui05, KHPC08, cL05, Röß06, SL07a, Sch09a, TYT^{+}09, ZB08]. Applied [MFL01, GV09, LW06, MH08, ML04, Tra07]. apply [Gor03a, RPR03]. Applying [AD01, LKPC04, SV00]. apportioning [Bec08]. appreciation [HY09]. Apprentice [KB04a]. Apprentice-Based [KB04a]. Apprenticeship [Cha00a]. apprenticeships [Lon08]. Approach [BB03a, All03, ABB^{+}03b, App06, BG08a, BGDR05, Bec03c, BC04a, BP09a, BDM01a, Bud06, BCB09, Chr09, Cli01, Cox05, DAJDRG08, DW05b, DMKN02, Ell06a, Ens09, GEVZ03, GPF^{+}01, Gon06, GH06, Gri08, GA03, Had01, HHK^{+}09, Ham02b, Ham04b, Ham06, Ham05, HEB08, Hol08, HB09b, Imp07, Jac00a, KAPG07, Kru03, cL05, LV05, LCW06, Lee02a, LE07, MD08, MGH^{+}06, MMC01, MS07, Mos06, MLC04, MPK08b, Mur07b, NM04, ÓT04, OG06, OM08, PHC09, PPP06, PHMS08, PPA08, Rag04, Raj05, RM08a, Ros06, Sha03, ST08, SSS06, Sna06a, Spo00, Str00, Ten03, TV08, VW00a, WS01a, WF08, WP07, Xu07, Zil05, Zin08]. Approaches [ABB^{+}06, PH04a, Tho06, Bow07b, BDRH09, BS05, HB09a, HTW04, HP06, LW05, PB05a, PFVI00, RMZG06, dMBM08]. appropriate [Ber09b, PHM^{+}01]. approximate [MG04]. approximation [FLS04a]. apron [Ast07a]. aptitude [Car06, CLB07, LC06]. Arabian [ASA00]. archaeological [Ast03]. architectural [Chr05b]. Architecture [Don01, Bem02, BP03, BB01, CHK^{+}01, CK02, Chu01, CGR^{+}08, HMS04, Ibb02, IDD06, JCK03, LL04, Lee03a, MHK03a, Osb02, SNA09, Ska02, WYOH02, Yeh02, YBAD05]. Architecture-dependent [Don01]. architectures [CSP09]. archiving [BGKT00b]. area [GP02, GLC07]. areas [KSRD05, SGC^{+}01]. arena [Gor02, Pra05a]. Arguments [Hen09a]. Aristotle [RC00]. arithmetic [Cha03b, War04]. arrangements [Gin07b]. Array [Bur03, Hu05, Rol03a]. ArrayLists [JT04]. arrays [Bla09d, BDM05, EB08, JT04]. arsenic [Jon01]. art [BR09, Mas05, PKR06, ZE01]. articulating [Kol08a]. Articulation [CTP^{+}04]. artificial [HBH07, HA04, Imb03, Kla02a, MF07, Mer01, SS06, Yan09a]. arts [ABB^{+}03b, BCD08, BGDR05, BBN^{+}04, Cli06, Con00, DKG06, DH06, Doo03, Gup01, HS03, KvLA09, KA09, MKBW07, PBM03]. Asian [PV08]. ask [GGH03]. asked [SM05b, VT01]. aspect [CSP09, Haz01b, Yan07]. aspect-oriented [CSP09]. aspects [Cha02, Haz03b, OG01, SSR07]. Asperger [Ega05]. assembling [Ort03]. Assembly [BD01, BBE06, Gou09, RM01, VS06, Zil05]. assertions [Gin04c]. Assess [CAMF^{+}03, Imp07, MNBK03, SCL^{+}08]. assessed [ES00]. assessible [SMS08]. Assessing [CLP06, DW04, FK08, Gha05b, JFH00, LdLM^{+}08, MIS^{+}07, VMD01, Wol07, FLK^{+}09, Imp09, Mic00, Wic07]. Assessment [BPD03, HPCZ04, PTT08, ALR04, Abu04, APR06, AFR08, AKA09, AMW06, BdMG08, BH03, BDMY05, BS05, CHB06, Chi05, Cle08b, DW04, Dic05, DLM07, DW05b, Eng04, Eng06, ER09, FMK^{+}00, FH05, GLC07, GSW08, Gov07, Haw09a, HTSW07, HB06, Hob00, HHS02, HL04c, Jac00a, Kac01b, Kac01a, KM00, KLOT05, Lap02, Lis01b, Liu06, MKS02, McC03b, MAD^{+}01a, MAD^{+}01b, MD05, MSS^{+}04, MPJC09, NLCvL09, SMK01, SM03, San09a, San09b, SDE01, SV03, Sha05, SJ04, Sol07a, TSW07, TG05, VA02, WC04, WP06b, WP06a, WM03b, pYB08, YLRZ07]. assessments [Doo09]. assignment [BP00, Cas01, Djo09, Edm07b, HW08, Hol00, HWH^{+}05, KF03, Liu08, MV01, PP03b, Par02a, PLA^{+}05, PL04a, Red01, Rol09a, Rol09b, dR04]. assignment-structured [HWH^{+}05]. Assignments [Bec03a, LDB^{+}03a, Par04a, Par09, Ros02, Bec03b, BGKT00b, CM08a, EBC^{+}08, EPQ08, Gig03, GSW08, Gre00b, HW08, Hel07b, Hui01, LWS07, Moh09, OJJ00, PFVI00, PB03a, PCR^{+}01, Par01, PER^{+}02, PPR^{+}03, PML^{+}04, PWM^{+}06, PCS^{+}07, Par08, PMS^{+}09, RW05, SD00, SWR05, SW06b, Sud08, SPW^{+}08]. assimilation [MDH^{+}00]. assist [BC04b, RTET03]. assistance [PH04c, SFM^{+}07]. assistants [DVE06, Reg03]. assisted [Eng02, TSW07]. Assisting [BAdC07, HP06]. assistive [Lif04a, Lif04b]. Associate [Cam05a, Cam06a, HKCW08, Haw08b, HKCW09, Kle02]. Associate-degree [Cam05a, Cam06a, Haw08b, Kle02]. associate-level [HKCW08, HKCW09]. associated [Haw08a]. associated-degree [Haw08a]. association [CA02]. associations [PHC08]. assurance [Bry09, Cam04a, CNP^{+}09, Gha07, GSW08, MS03, YLYX06]. Astrachan [Par03a]. asymmetric [Sok00]. asynchronous [BKW05, Bow07b]. ATDD [SN08]. ATE [CTP^{+}04]. AtoCC [HW06]. attack [HAuR04, TW05]. attempting [DSE02]. attempts [GPP00]. attendance [BGL^{+}07]. attention [Kac01b, Kac01a]. attitude [AR00a, TRR03]. Attitudes [Lew07a, BU09, BDW^{+}05, DTP03, FS09, Han06, MJ03, Per09, SPM00, TT00, Wic07]. attract [CMD^{+}07, PBWS06, PB06]. attracts [AL05b]. Attribute [Nef04]. attributes [BL01c, Imp09, RP06, SB09]. attrition [BM05]. audience [CM06a, HPC^{+}03, Rob03b]. auditing [Lin07, Yan09b]. auditory [CMS06c]. Austin [CDWH00]. Australasian [Ell00b]. Australia [Ell00b]. authentic [FP09]. authoring [AKA09, LR01, PP03a, Ros08, SKM02]. authorizes [Gor00a]. authors [DBBK02]. Automata [McD02, ARVV06, Boz06, CFR04, CCY03, CGM04, GET05, GOMSJVGP08, Gri02c, HS04, HW06, HR00, LJ08, Mer09a, Pil09, Rod06, RBFR06, RLR07, RWL^{+}09, SW06a, SBR^{+}07, Tay02a, Ver05, WD05, WW06a]. automate [Sol07a]. Automated [Eng04, FMR^{+}06, KP05, O’N05, OHZ01, BUCP02, BGNM04, Eng06, ER09, FPQ^{+}08, HTSW07, Jac00a, Kur08b, PL04a, San09b, TIKT04]. Automatic [DAH08, POMA08, pYB08, dBGdMB09, GS06a, GRLDPG^{+}01, Hel07b, KM00, MKS02, Par02a, RLM08, RS09, SMK01, TWS05, TWS06, TWS09, TG05]. Automatically [KB09, YH03, EPQ08, ES00, OB07]. Automation [OTUW07a]. automaton [Gri03]. autonomous [GP02, MH00, SS07]. available [Gor06a, Kno02, McC02]. Avatars [CF02]. AVBL [Sar08, SK09]. average [Bec03a, She08]. aversion [GMS09]. avoid [Kla00]. Avoiding [CHB06]. award [Kno01]. awarding [OHRL01]. Awards [Hen03b, Pop03b, Pop03a]. AWARE [TW05]. awareness [Gin01c, Got03, GHG^{+}08, HLL03, Joy06b, PMC^{+}04, TW05]. awe [GCD^{+}09, MRGS08]. Aztec [Gig02].

b [Reg08, FLS04b, LS03]. B-splines [FLS04b, LS03]. B.Sc. [ASA00]. Babel [PY05]. BabyOS [LCG07]. Bachelor [Per09, Spo00]. Bachelors [BGD03a]. Back [Kum09c, Man04, Reg06, Noo06]. background [EG08, GM08, MN01]. Backstop [MKKC08]. backtracking [GM06b, Noo00, RP04]. backwash [PL04b]. bad [MD05, MLM^{+}08, Rol01]. bakery [Boz06]. balance [Cle03a]. balanced [BNSW05]. balanced-sample [BNSW05]. Balancing [Wal00a, Rol02, Sok02]. BalloonSAT [AJF07]. ballot [Med04]. balls [BM07a]. Banner [Pop04]. Bantam [CL08b]. bare [HHB04]. barely [Mur07a]. Baron [Mar00a]. base [BS02, HA02]. baseball [Pop03a]. Based [KB04a, Kum04, MFL01, AK09, AFJ^{+}03, AC08, AMUFVI09, APR07, AO06a, AAS^{+}04, AAD^{+}07, BP00, Bem02, BJJ01, BBW04, Ber09b, BC06b, BLA^{+}09e, BPK^{+}01, BG08b, BDMY05, BDM01a, BS00, BP05, BB01, CS02, CTP09, Che00, CS08, CTR^{+}03, CR02, CLK00, Con04, CCMC05, CL08b, CI03, Cul00, CBH09, DWT00, Dia05, DAJDRG08, DSBK01, DKN08, EH05, ESP03, Ekm06, EA00, ELC07, Est01, Fag00, Fow01, FPQ^{+}08, GE09, GS06a, Gol01, GSVG05, GSW08, GHG^{+}08, Hab02, Had01, Ham07, Ham05, Hel07b, Hic04, HB06, Hol00, HKF00, HNC08, Imp07, JT04, JFH00, JKKL04, Kag09, KK02, KB08, KvD02, Lab09, cL05, LV03, LV05, LCB^{+}03, Las09, LFL06, Lin07, Mar07c, Mat03, MG04, Mat06, MD08, MSTS02, Mér09b, MLH05, MPJC09, MHA04, MKLH09, NEN00, NRB^{+}05, NL04]. based [Nef04, Nør00, ÓT04, OG06, O’L06, PDS07, PY05, PPRV06, Pul00, Qua06, Qui06, Rad06, Rad08, RH03, Raj05, RN03, RC04a, RSK06, RLM08, Ros01, Röß06, SM00a, SC04, Sca00, Sci07, SNH08, Sha00, SPM00, SMR03, SSC01, SAHS04, SJA04, SS00, SBZ09, Ste03, SJS^{+}07, TIKT04, Tho06, TT00, To04, Tre02, TBM09, UFVI05, Vat01, Way05, Wic05, Xu07, YLRZ07, YH03, YB01, Zil05, ZRY06]. bases [Cle05a]. basic [Bel02, CGR^{+}08, GELS04a, HK01, Jin08, KPT03, MSK09, TH06, WB09, HM02]. basics [Kay00, Reg06, Shu04]. basis [Bur06, DeP06, Hic04]. basket [WW06b]. battle [GPP00]. Bayesian [KN09, Yan07]. be [BDM05, BGSL06, Chr03, ESP03, FG08, HHRA06, Kac06, KBC01, LS02, MBD04, PP07, RWGM00, RR06, Sic09, Tra07, Wal05c]. Beacond [Cha03a]. beauty [GCD^{+}09, MRGS08]. become [KTH^{+}00, TKB01, WG07]. Becoming [PPR^{+}02, Cle04b]. bedfellows [Cle04a]. been [McC05c]. before [APGR07, BDM05, HM00, SM05b, Vas02]. beginner [Lau07]. beginners [Eng01, KK03, Wis06]. beginning [CLM^{+}06, Cun02, CLP^{+}01, HM00, Jac00b, NBF^{+}08, PRR02]. beginnings [CAK09, San08]. behaved [Lee02d]. behavior [DSE02, DDC^{+}03a, GSVG05, LBA08, Wal05d, WLCL08, WHCT09]. behavioral [RBJ^{+}09]. behaviors [FNB^{+}09, HM05]. behaviour [JDGAGMGC05]. behaviourist [PB03a]. being [BAAM06, CS03a, DABJ00, Got00b]. beliefs [Lew07a, Per09, WI09]. benchmark [CWHL06]. benders [Spe02]. beneficial [ESP03]. Benefits [BKS^{+}04, JFH00, OB05, Sar08]. Beowulf [AV02a, AB08]. Berkeley [BM07b]. Best [BAET09, Con08, DGHC03, MA01, WG07]. better [ABE^{+}02, Ast01, BL01b, MP09a, MSP^{+}06, MAR07d, RJM^{+}08, RK09b, ST08]. Between [Coy04, CF04b, CBC09, DBBK02, DFGM08, Dic07c, GPP00, Gin08c, GCF04, GCFP04, Haz03c, HL04b, HPM^{+}08, Kum01, LFT09, Pre07, RB09, SSP^{+}05a, Sud08]. Beyond [Med04, ADM^{+}09, BMK03, Gag02, GDHW06, GL08a, GL09, SLSM02, TA08]. BFS [EG08]. big [Alb09, Rou04, Wal08a]. biggest [Kac06]. Binary [PDC01, BNSW05, Hab02, Kru03, PHC08, Rol02, SW06a, SM00b, War04]. Binomial [Man02, Rol01]. Bioinformatics [BDDL04, FJ06, LD04, BBN07, BJ07, BS04, Cut07, D’A03, DD08, DRKG02, Ell06b, Khu08, Qin09, TLW^{+}04, TLW^{+}09, ZLOB07]. biology [BCL00, Coh02a, NRW04, Qin09]. bison [Kag09]. bison/flex [Kag09]. bit [BSV04, Cha04b, SWR05]. bit-level [Cha04b]. bits [Chu03]. Black [Got02, HK01, WTTT^{+}08]. blackboard [Qua07]. blackboards [RTM^{+}04]. Blackjack [Pop09b]. Blackjack-playing [Pop09b]. BLAST [Cut07]. Blended [wH09a, DFS03]. blessing [McC00b]. blind [BAB^{+}08, CCL^{+}07, Sie06]. Blogger [Mar07a]. Blogging [Mar08]. blogosphere [Got07]. bloom [LL03, SMS08]. blower [Bri09]. blowing [BVPE06]. blue [Got02]. BlueJ [FMKQ08, Hag00, KR00, PH03, PHBM05, PHN06, PKR03, XSD07, PH04b]. Board [Gin03a, Imb03, Lim07, CCD^{+}08b, DCP^{+}09]. boards [JY01a]. body [CTP^{+}04, Jon02]. BoF [Hen08b]. Bologna [DGS06, Cun07, FPA^{+}06, Por09]. bolts [CLP^{+}01]. bombardment [EA07]. Book [Lis04a, Lis09a, Gib02, Lor00, Noy02, Lis06a]. books [CI03, GE08a, MUPVPF^{+}02, WC04, Yai08]. both [FH01, LNR03]. Bottom [Knu03, Gin01e, Got00b]. Bottom-up [Knu03, Gin01e]. bounded [Rob00b]. bounding [GM06b, RP04]. bounty [McC01b]. box [Ast09a, Med04]. boxes [BMG03, HK01, Hab02]. boys [Kla02b]. Braided [Pre07]. Braille [Dic07a]. brain [Par01, Tha07, PFJC05]. Branches [MC08]. Brazil [Tur08a]. Brazilian [LdLM^{+}08]. breach [Got03]. Breadth [DAKLH07, Pow03, WW09, Bla09b, BZ04b, DLHAK08, Mur01, PSW03, RSK06, Sha03, VW00a, Vas02, Ver05]. Breadth-also [Pow03]. Breadth-first [DAKLH07, BZ04b, DLHAK08, Mur01, PSW03, RSK06, Sha03, VW00a]. Breadth-last [WW09]. breadth-second [Bla09b]. breaks [TRT05]. Bridging [BPS09, HDH06b, TRT05, Wic07, BHLW02, HL04b]. brief [BSV04, Min08]. bright [CAK09, Wol02]. Bringing [Alb09, Moo02, UCJ^{+}04, MDD06, PD06, STS09]. brings [Tea02]. Broadening [PHS^{+}07, RI08, DLM09]. broader [FG08, His09]. broken [Mil09]. brokers [Jan01]. brown [Cle05a]. browser [KB01, Lab09, NM02]. browser-based [Lab09]. BS [ND08]. Bubble [Ast03]. budding [ML07]. Buffer [Sha04, PT03, Rob00b]. bug [BA01, Lar06]. Bugs [Ric00, GAGG04]. Build [Bla09a, Cle09a, SSS07]. Building [AH08a, ABO^{+}05, BC06b, FB02, MR02, MPD^{+}05, SS07, AV02a, AL09, BRP^{+}07, KAPG07, PV08, SJH06]. builds [SLC^{+}09]. built [Dic00, KLOT05, VT01]. Bulletin [Wal01b]. burning [Lis09a]. bus [Alv05]. business [BC04d, Dic07a, GE08b, GR00b, Gor05, Top08b, Van08]. bytecode [JCOP07]. Byzantine [TBAK04].

C [CMRR04, Car03, EMM00, EPM09, Gol08, Kum01, Kum04, Kum05b, Kum09a, MD06a, NC07, Nod04, Phi04, PWH00, PNF00, War01, WSP02]. C# [ABE^{+}02, GS05a, JJOA07, Nod04, Reg02a, Sax03]. CAC [San09a]. caching [She06]. CAD [San08]. calculation [BLA^{+}09e]. calculations [Gin00b, Wil04b]. calculator [LS06, MG00]. calculus [GS03]. calibration [MT05]. Call [Lis06a, AH08b, Har01a, Fow01]. calls [BMP06, BM07b, Man05]. cam [SS07]. CAMERA [NR05]. camp [CPP07]. campaigning [Mar08]. camps [Ada07, DLM07, GHLL08]. campus [FH01, HE06, OTUW07b, PW01]. Can [BZ04a, EMM^{+}06b, EMdL^{+}07, FG08, Kla00, LBK08, Reg02a, Tra07, Abo09, BGDR05, CLM^{+}06, Har05, HB09b, Kac07b, LZ07, MAFW08, NPM08, Pat06, Pet09, SMH05, Wal04c]. cannot [Wal04c]. canon [Eis03]. Can’t [Pat06, Har05]. Canterbury [FKCM01, Fin01]. CAP [Haw09a]. CAP-space [Haw09a]. capacity [PV08]. Caparica [CFPL05]. capitalize [Gin04b]. Capstone [BC04d, ADG^{+}03, CB02, CGY^{+}01, Cle02b, Cle07a, Cle09b, Con04, Goo03, Hua03, Jon00, LFL06, MH00, O’L06, PRK05, Ram03, Sta05, Tut00]. capture [Cha07, Lew07b]. Capturing [RTET03]. cards [GGR03, KNT02]. care [Ast04]. careers [CLP^{+}01, HTBS01, PHW07, PBWS06, PB06, SBLH07]. Carnegie [FM02, FB02]. carnivorous [Kac08a]. Carolina [WMGR01, Wil07, ACM01, GE08b]. cars [Wir08]. cartography [SWR05]. Case [ACL^{+}03, MFL01, NW03, PDS07, Qui06, RCR04, AGHM02, AGHC03, AGE06, Arm08, BCD^{+}04, BDM^{+}07, BCL00, BL06b, BS01, Cal08, CCM08, Cun00, Dei08, EH07, ER09, ENO09, GE08b, GEVZ09a, Gin04a, GSS00, HA02, Har05, Haz01b, Kla02a, KPT03, KT07, KvdMvDvdV00, Kur09b, LL06a, LL02, MBD04, PB03b, PP06, RJM^{+}05, RKD02, Ruc07, She08, SS00, SH07, Sud08, Utt06, VZGE07, Dix04]. Case-Based [MFL01, PDS07, Qui06]. cases [HA02]. castle [EB08]. CAT [EPQ08]. Catalan [PPRV06]. catch [GPP00, MRM^{+}06]. cause [SB07b]. causes [KT07, MBE^{+}09]. CBIS [WD03]. CBL [Gel00]. CC2001 [Doo04, Doo07, Doo08, Gor01a, MBPB03, PSW03, RJM^{+}05]. CCLI [MR00, Pre01]. CE [SK09]. center [COR07a, EA00, GWG02, Mon02]. Centered [AF03b, AO06a, App06, BG08a, Old08, Tra07, WD03]. centering [DMRS06]. Central [GCW^{+}03, JDC00, WO07, CL02]. centralized [GL05]. centre [Ekm06]. centric [DV07, Gol01, Gup06, HKW^{+}02, San09b]. century [BDRH09, Laz02b, Laz02a, Lee01b, Mar01b, Mar05c, Sal03a, Sal03b]. certificate [ZV04]. certification [Hit05, MK01]. CFX [RJAH04]. CGTDEMO [JDC00]. Chain [Gin01a, Mau02, WSP02]. chaining [SNH08]. chair [DWY03]. challenge [Fri05, GS03, Kac06, Sie06]. Challenges [ADG^{+}03, AHGP05, GET05, GV09, Gin09a, GM07, Wen04, Car05a, CM02, FLLM06, Gin08a, Gin09b, GPL^{+}05, GAL^{+}06, GGKT04, KDH^{+}00, LBC04, Lis05a, NRRR06, PHS^{+}07, Rya01, Sut00, TMD^{+}04, WB05, WBGM02]. Champaign [CCM08]. change [LM03, Ni09, GWG02]. Changes [BDW^{+}05, DHRH05, ACK^{+}00, HHRA06, MBE^{+}09]. Changing [Kla05a, Sam02, Cun00, Har01b, Lau06, SLPG07]. Channels [CCJ^{+}02]. character [vLK09]. Characteristics [CHMG01, SB08, BGD03a, Gra03, PJ05]. characterization [GET06]. characters [RBC07]. Charles [Lis06a]. Charlotte [ACM01, WMGR01]. charming [Doo08]. chat [AC08, CS00b, MV01]. chatbots [BAB^{+}08]. cheat [DSM01]. Cheating [Par07a, SDM^{+}02, CCH^{+}03c, DSB^{+}03, SD03b]. check [Gri01]. checked [PWH00]. Checking [OB07, Hen08a, NSH^{+}09, TYT^{+}09]. Checklists [ST07]. Checkmate [PWH00]. checkpoint [Eng06]. cheesecakes [Kum08b]. Chernobyl [Mar04b]. chessboard [Gin05a]. children [BT05, God06, Pet09]. chips [Gin09a, KSE02]. Chirp [XM06]. Choice [AU05, Bec06, Fon02, FS06, MPK^{+}08a, Rad09, Soj03c]. choices [EH07, SM08]. choose [Car06]. choosing [Pat06]. CID [Boi04]. CIMEL [BPK^{+}01]. cipher [HAuR04]. ciphers [MWM01]. circuits [JMS02, Kur08b]. circularly [She04b]. CIS [AKM00, PN08]. Cisco [AM07]. CISE [CPT09a]. CIspace [AAC^{+}05]. citation [LB08b]. cite [Law05, Lis08a]. CITIDEL [Cas03, CFL^{+}04, Kno02, KCI03]. citizen [Di 06]. claim [DV04]. clandestine [Lis04b]. class [Aba08, Bec08, BAET09, BDM^{+}07, CS00a, Cas01, CFL^{+}04, DV04, Dei05, DKL^{+}07, ET05, Fek02, Fos05, HLM01, HBCW03, Kac03, Las09, LNR03, NRDP07, PHC08, PG06, RN03, RF00a, Rut01, SLC^{+}09, SL09a, SM02a, Soh04, Sud08, SSR07, Tre02, UD03, Wal07a]. Classes [Fos05, AR00b, Bor07, Bri09, CIK^{+}03, FP09, Fek08, Guz01a, JJOA07, KW07, LC08, PH07, PP07, PD06, RMT09, Rou04, SFD08, TRT04, Wol02]. classic [Eis03, Oka05, WP07]. classical [Min09, Ste04]. classical/linear [Min09]. classification [PB03b, SDR09]. Classroom [AAV^{+}03, GL03b, HSSC05, Rol06a, Wal08a, Wal09a, Wal09b, ACR03, AAD^{+}07, AU05, BGDJ02, BGW04, Bla09c, BR05b, Bow06, Bow07a, BEE07, BGL^{+}07, CP03, CTP09, CS08, CL08b, CT07, DGSW06, Deu08b, DAH08, DAHA09, Ega05, GC05, GS06b, HS04, HHB05, HP07, IMRJ09, JKDS01, JMS02, KD07, KV08, KMMS02, KTKP02, LNH09, LDK^{+}07, MHW03, MNBK03, PTT08, Pho01, Qua04, RC04b, RTM^{+}04, SV03, SB07a, TF07b, TS02, TEP06, UCJ^{+}04, VA02, Vau00, WJD03, Wal03a, Wal04b, Wal05b]. cleaning [MHW07]. clicking [OG01, PS06]. Client [JBB^{+}03, Sha02]. client/server [Sha02]. climate [BGDJ02, HHRA06]. Climbing [LSG05]. clips [Har03]. ClockIt [NBF^{+}08]. clogosphere [Got07]. Closed [Les03, SSP^{+}05b, PHMS08, SSP^{+}05a, Soh06a, Wic05]. closed-loop [PHMS08]. Closing [WP06a]. clothes [Man07]. cloud [CBKA09]. clouds [Ast07b, Kac08b]. clubhouse [FM02, FM03b]. clubs [BBC06]. clue [HBH07, Kac09a]. Cluster [KMSB08, AV02a, AB08]. CMeRun [Eth04]. CMPS [Whi03]. CMS [Kül09]. co [DGHC03, SM02c]. co-design [DGHC03]. co-operative [SM02c]. coach [Tho01]. Coached [LV03]. COBOL [Fri04]. Code [TRR03, AEPQ09, Bai05, BP00, BC09, Cle05a, Gar04, Gib09, HTSW07, HAFT09, JWC07, JB07, MP09a, PH07, RMT09, Try05, TQCPQE08, WLCL08, WW00, WSS05, Zel00]. code-a-phobes [TRR03]. code-once [Bai05]. code’n [Boo07]. codes [Chu01, JWC08, Mar01a]. Codewitz [KK03]. coding [PHC09]. coefficient [Man02, Rol01]. coffee [Kac08b]. cogito [Ast09a, Aha00]. Cognitive [Haz03b, MH08, OBL04, Aha00, BS00, BS01, MFR^{+}08, MFR^{+}09, MGH^{+}06, MMC01, PR02, RWP03, WTTT^{+}08]. collaborating [CWP09]. Collaboration [BC07a, SGKL^{+}01, AHM^{+}04, Cle08a, Gra05, JKDS01, LDA^{+}00, LDHW02, MGE^{+}05, RHE04, RC02, TCK09]. collaborations [PBM03, RRD^{+}09]. collaborative [AL09, BYS09, BPK^{+}01, BDF^{+}08, BROB00, CK08, DDS08, GCDJ08, Goe08, Guz01a, Ham04a, HR02, HYN03, JP03, LHS09, Mar07c, MDD06, MD06b, MW09, MPJC09, MLK07, O’N05, RTET03, Reb05, THH09, WHCT09]. collaboratively [FL05]. collaboratory [Pra05b]. colleagues [Kac08a]. collecting [NBF^{+}08]. collection [BKM^{+}01, Fek02, Poh05]. Collections [Wic03, BGN09, UFVI05]. collective [WHCT09]. College [CTP^{+}04, HCK07, AV02a, BS08, Bro09, Cam01, Cam03a, CHK03a, Cam04b, CHK05, Cam05b, CHK06, CM08a, CKMP09, Doo04, Doo07, Gou08, Haw08a, Haw09a, Haw09b, Hen07a, Kla02a, Kle01, LS04a, NP02a, Pop05b, TFG05, TT00, FLL^{+}01]. colleges [Cam02b, CHK03b, Cam04a, ESS^{+}07, Haw09b, MKBW07, WTTT^{+}08]. collide [ZE01]. Color [Gin01b, Tay02b]. Colorful [Gin00a, GGAB01, Gin08a, Gin09a, Gin09b]. column [Hen05b]. columnists [Imp05b, Imp06b]. columnsort [McC04a]. Combating [GG06b, WW00]. Combinatorial [RM08b]. combinatorics [RS08]. combine [HM05]. Combining [BP09a, CO00, Doo03, NM02]. comes [AGGR04, Vas02]. comfort [LM03]. command [Sol07a]. command-line [Sol07a]. commanding [Bry09]. commemorate [ADM^{+}09]. Comments [Mas03]. commerce [GS00, Wea04]. commercial [LCB^{+}03]. committee [BMW04, Cam01, Cam03a, CHK03a, CHK05, CHK06, DDLR01, FKO07, HAJ07, Mar05b, Mar06, HCK07]. committees [Wal02a]. Common [PP04, AL01, CL00, EBC^{+}08, Mar07a, Rob04b]. commons [Got07, TF07b].

Commonsense [CLM^{+}07]. communicate [Boi06]. communicating [Boi04]. communication [BU09, BBCD09, CFJ^{+}02, Cun02, DGSW06, Etl06, Gra05, HL07, Mic00, Pol01, Ree00, WB05]. communications [EA00, HL08]. communities [AL09, GM02, BGN09, She04a]. Community [Bro08, CTP^{+}04, Haw08a, Haw09a, Haw09b, How05, KP04, GCF04, GCFP04, Imp08, LFL06, Lin07, MA01, PW08, PBM07, TM03]. community-based [LFL06, Lin07]. Community-oriented [KP04]. companion [All03]. company [BFMC05, Way05]. company-based [Way05]. comparative [ASA00, Mau02, Sok02, UD03]. Comparing [PV08, WP02, BBY08, POM07]. comparison [PTT06, PTT08]. COMPASS [ALR04, Abu04]. compatibility [KWW^{+}04]. compelling [Ber09b]. compendium [Gor01a, Haw08a]. competence [SB08, Sic09]. competences [RVP08]. competencies [Rus09]. competency [Jac00b]. Competing [Man00]. Competition [Cle00a, BVPE06, KH08, Hen03b]. competitions [Bow08c]. competitive [JWC08]. Compiler [NPM08, AT04, Bal03, BC04b, CL08b, Deb02, Dem05, DLP08, dOG07, Hen05c, MP09a, Mea06, MPD^{+}05, Nef04, Ruc07, SL07b, Wai06, WJJD06, WSS05]. compilers [Aho08, Bal03, FM06, SNA09, XM06]. Complecto [Con08]. complementary [PFVI00]. complete [BA05b, Eng01, GH01, GGR03]. Complex [GD07, Gou08, NH05, ENO09, Hum04, KT07]. Complexity [BC07e, BS00, Chr05b, Hov05, Tay02a]. component [APR07, CIL^{+}03, CLK00, Gor05, Rou02, SSC01]. component-based [APR07]. Components [HTW04, PDC01, PHM^{+}01, ZLOB07]. Components-first [HTW04]. composition [Gin09c, Ham04b, LV05, dBGdMB09]. Comprehending [Cle05a, Soo01b]. comprehensibility [HCMM00]. comprehensible [Yeh03]. comprehension [Ext00, GBO08]. comprehensive [AV02a, CMS06a, GPF^{+}01, MKC09, NM02, Nyg02, PFVI00, Whi03]. compression [BP00, KH00, RL09]. computability [Haz02]. computation [BJJ01, Cor07c, Guz03, Han02, Kap04, Man02, Sav05, Sig07, Ste04]. Computational [CPST02, CPT^{+}09b, HCW07, And00, AGE04, AGE06, AHPS09, CWHL06, CS03b, DK06, HHK^{+}09, LF09, Noy02, PB03b, Qin09, SHN^{+}00, Shi02, TSC^{+}02]. computations [Gin09d, Mir09]. Computer [ACM01, ACM06, ASA00, ALRY08, Ano03, BTHR06, BCK04, BGL^{+}07, Car02b, CJGU02, CDWH00, CH07, CS03a, CFPL05, CS03b, DSF^{+}03, DNTB05, DGS06, DRFG08, Ell00b, FKCM01, Fin01, Fin02, Fin03, FS02, GESZ09, GWG02, Ges07, GLD^{+}05, GKJD03, Gup06, GS09, GS03, HKCW08, Haz03c, HV05, HMRM03, HPT07, IDD06, Imp02b, ISMS03, Imp03, Imp04, JKDN04, KH08, Kol08a, LC08, LdLM^{+}08, Laz08, LH02, MDS04, McC00a, MBE^{+}09, PBM03, Pra05a, Ran07a, Rob04c, RPH^{+}03, RHDR07, SL07a, Shi02, Shu06a, SdSK05, TSW07, WMGR01, WP06b, YBAD05, YH05, AHLY03, AEH05, ABAE^{+}09, ACL^{+}03, AKWZ07, AHGP05, AHH^{+}06, AAF^{+}06, APR06, AAS^{+}04, ACSS06, ARVV06, ALH07, ACK^{+}00, AWS^{+}09, ALO^{+}03, BC07a, BGDJ02, BGD03a, BGDR05, BMK09, BBMR03, Bea02, Bea03, BM05, BM07a, Bec07]. computer [BBN07, Bem02, BP03, BABBH04, BKM^{+}01, BKR^{+}05, BSV04, BBW04, BRP^{+}03, BAMSK05, BT06, BBY08, BAB^{+}08, BBCD09, BN00, Bot05, BEM^{+}07, Bou09, Bow08a, BDM^{+}07, Boy08, BBN^{+}04, BR02, BMR04, Bra05, BPS09, Bri09, Bro08, BO01, BS01, BKS^{+}04, BCB09, BR09, BBT07, BDDL04, BS04, BB01, CGB08, Cam06a, CL08a, Car05a, Car04, Car08, Car06, Cas09, CHK^{+}01, CK02, CB00, CM02, Cha06b, Cha04a, CO00, Cha04b, CCY03, CBJL08, Chr03, CR03, Cle00a, CHM^{+}01, CFGL05, CKMP09, CKLW05, CF04b, CDP03, CCMC05, CGR^{+}08, Cor07c, CPT02, CGI05, Cun00, Cun02, Cun07, CA02, Cur03, Cur07, CM08b, CS06, CBKM07, DKG06, DR07, DG05, DFGH04, DV04, DMRS06, DDA^{+}07, DD08, DW07, DRKG02, DTL^{+}02, DTW07, EMM^{+}06a, Ega07, EH07, Eis03, ELE^{+}02, Eng02, Eps06, Erl06]. computer [Etl06, Fag00, FM03a, FKKW00, FM09, Fon09, FG07, FH05, FB02, Fri05, Fro07, FJA^{+}07, FIG07, GT00, GEHR09, GP02, GSyC07, GKHPQ07, GS00, Geh01, GG06a, Gib00, Gol04a, GD08, GMGM06, Goo08, GR00b, Got03, Got09a, GD07, GL03a, Gra07, GHG^{+}08, GK02, GMHS01, GGKT04, GHLL08, GBO08, Gür02a, Guz01a, HAM00, Han07, HB01, HEB08, HALW01, HL07, Haw08b, HKCW09, Haz01a, Haz03a, HL04a, HL04b, HDE^{+}06, HL06, HGEB08, Hel07a, HDA^{+}02, HBEM03, Hep06, HPC^{+}03, HPCZ04, HB06, HE06, HETD09, HDH06b, HV07, HAJ07, HRB^{+}09, HB09b, Hug09, HS02, Ibb02, IK06, IHL^{+}04, Jac00b, Jam02, Jen08, JHLT06, Joe06a, Jon00, JCK03, KK09, Kac03, KK02, Kay00, KvdHR05b, KH05a, Ker04, KW07, Khu08, Kla05a, Kla05b, Kle02, KH06, Kli05, KR08, KAPG07, KSRD05]. computer [cL05, cLwW07, cLwW09, LH03, LH05, LA08, LL04, Lax01, LD04, LS04a, Lee00b, Lee01a, Lee02a, Lee02b, LJG05, LS05, Lew07a, LLP^{+}09, LZBO09, LHL03, Lis04a, LGA^{+}01, Lom07, LZL08, LC06, LLB02, Luq05, MMN^{+}09, MVC00, MVD01, Mal07, ML07, MIS^{+}07, Mar01b, Mar02, Mar03, Mar04a, MA01, MDH^{+}00, MM00, McC02, McC03a, MPQB^{+}00, MDM04, McG00, ML04, MRH09, Med04, MAFLR05, MAFLR06, Mic00, MR01, MHK03a, MK01, MJ03, MPJC09, MKC09, Mos02, MMK02, MWF^{+}02, MT05, MMW^{+}05, MT08, Mus06, NRA^{+}03, NL04, NN00, Nod03, NH05, O’L06, Ola00, Oli05, Ols06, Osb02, OHRL01, OM08, PRC06, Pan09, Pap03a, Pap03b, Pap03d, Par06b, PS06, PR02, Par05b, PDC01, PH04c, PPR^{+}02, PSH^{+}07, Per09, PTC^{+}01, PSS09, PV08, PD06, Pol01, PPA08, PMG^{+}04, Pow08]. computer [Pra05b, PM02b, PT02, PP04, Pro00, Pul06, Qui06, Rad06, Raf08, RH03, RH08a, Ral05, Ran07b, RMZG06, Ras00, RRKS06, RHY07, Red01, RBC^{+}02, RB09, RMMH01, RMPD09, Rio02, Rob00e, RKI02, Rob03b, Rob04b, Rod02, Ros01, Ros09, RDCE06, Ros07, RF00b, RJM^{+}08, RS03, Rya01, Sam02, SM00a, SM03, San05, SV00, SC04, SC07, SB08, SBG^{+}08, Sco02, SSC00a, She03a, SL00b, SLPG07, SSC00b, SHVP05, SAHS04, SJ04, Ska02, SF02, ST08, Sna06a, Sol07b, SS00, SSA08, SMS08, SBZ09, SHL^{+}00, SGKL^{+}01, SH07, SJS^{+}07, SB06, SSR07, Sur02, TW09a, TPK^{+}06, TF07b, Tha07, Tho06, TT06, TW05, TBS^{+}00, TM03, THH09, TDJ^{+}03, TMD^{+}04, TYC^{+}07, TT05, Tut00, VW00a, VMD01, VCC^{+}05, VCC^{+}07, VVV04, VZ06, WW04, WJDL04, Wal03a, Wal04a, Wal09a, Wal05d]. computer [WAS^{+}02, Wan06, WHCT09, WS02, WW09, Wil00, Wil03, WS01b, Wir03, WYOH02, WBS04, WR07, Wos06, pYB08, Yeh02, ZSW04, ZLL^{+}07, ZB08, ZV04, ZIBG05, dMBG^{+}08, vdK07, BPT^{+}04, DTL^{+}02, HKW^{+}02, YB01]. computer-aided [SL00b]. computer-assisted [Eng02]. computer-based [GHG^{+}08, SS00]. computer-centric [Gup06]. Computer-related [RPH^{+}03]. Computer-science [Kol08a, Geh01, KK02, Tho06]. computer-software [PRC06]. Computers [BB03a, All03, ADM^{+}09, BJJ01, BBW04, CTR^{+}03, Cli04, Got03, JKDS01, Joe06b, Kla02b, LML^{+}01, NL04, PF03, Sch07]. computes [BBE^{+}09]. Computing [ASA08a, ASA08b, AAO^{+}08, ABS08, AH08b, Bor02, CIL^{+}03, Cle07a, Cle07b, Hem05, Imp05a, Imp06a, LS01, MSM08, MMD^{+}09, MFL01, Mar02, Mar03, MK06, MPH09, Raf08, Rob00d, REC^{+}01, SCC^{+}04, SL00a, Wir02, ZRY05, Abo09, ANS00, AGHM02, AGHC03, ABO^{+}05, AAEI08, App06, AHJ^{+}07, BHBG07, BAK^{+}09, BBC06, Bow08b, Bow00, Bro09, BBE^{+}09, BNS08, BZ04b, Bur04, Cam02b, CBKA09, CIK^{+}03, CHM^{+}05, CSS05, CMS06a, CMGR07, CDF^{+}07, CSD^{+}07, CMD^{+}07, CLMW09, CDS09, Cha03a, CLM^{+}07, CL02, CY01, CGY^{+}01, Cle06b, CK08, Cle08b, Coh02b, Coh07a, CM06b, CBC09, CT07, CT09, CA02, DCT06, DABJ00, DPBP09, Dei08, Den04, DDRG04, DLM07, DKN08, DKR09, EB09, Edm08b, FP09, FJ06, FM03b, Fle06, FHBD06, FPQ^{+}08, FPA^{+}06, FLK^{+}09, GM02, Gab03, GOB^{+}07, Gal02, GCD^{+}09, GLA^{+}08, GCPQA05]. computing [GM06a, GM08, GGH^{+}08, GL05, GR00a, GR00b, Gor01a, Gor03a, Gor04b, GKL^{+}02, GC02, Gür02b, GF05, HM00, Har01b, Haw08a, Haw09b, HILS05, HKW^{+}02, His09, HY01, HWH^{+}05, HAH^{+}06, HH05b, HNC08, Imp01, Imp02a, ISMS03, IHL^{+}04, IL04, Imp07, ISC^{+}09, JM02, JY01a, Joy06a, Kac06, KSE02, KMSB08, Kno00b, KEK^{+}07, LC08, LNCC09, LD04, Lew07b, BGN09, LS02, LLLC06, MAFW08, Man00, Mar00b, MRGS08, MNBK03, MSS^{+}04, MBK09, Mun09, NLMN06, NSW08, OG01, OAB07, PHW07, PKB06, PSE^{+}05, PHS^{+}07, PBWS06, PB06, Phe07, PPJH09, Pop02, PBM07, Rad07, RPR03, RP06, RJM^{+}05, RBG02, RBTC03, RK09b, Sam02, SR07, Sic09, SD03c, Sko07, SBLH07, SL03, SSS^{+}09, Sok00, Tay02b, Tea02, TFIJ08, Tho03a, TT00, Top09b, Tur01, UHFT09, Vau00, Wal02a, Wal03a, WF08]. computing [Xu07, YLRZ07, ZRY06, vVML04, CLR09, RCD^{+}02, SMS^{+}06]. computing/informatics [vVML04]. concentration [CKLW05, CP00]. concentrations [BMP06]. Concept [AHH^{+}06, All03, ABA09, BDMY05, Bow07b, Cha00b, Fon01, KRTT01, KMS01, MSK09, PGK05, PK06, SBE^{+}08, SJA04]. concept-based [BDMY05]. conception [Haz03c]. conceptions [ET05, PL04a]. Concepts [Bur03, JBMP03, Kaw07, Ari00, ALH07, BL01a, BEM^{+}07, BMR04, Bra05, Bru05b, CHM^{+}05, Ciz03, DWJ04, EMM^{+}06a, FH01, Fon01, Gag02, Gol04a, GGH^{+}08, GMGM06, Gra05, HK01, Haz02, Hen07a, HRBC03, HH05b, HH05a, Jin08, LR04, LZBO09, LM05, ML04, MWF^{+}02, MMW^{+}05, MRM^{+}06, NR05, RW02, Ric02, ST07, SC04, SV03, SM05b, SWR05, Sto06, TH06, TS00, UD01, VZGE07, WB09, Yue07, ZJ03]. concepts-first [Gol04a]. conceptual [HLL03, Kac09b, PPA08, Wic07]. Conceptualizing [NM04]. concerning [SBG^{+}08]. concerns [Rou06]. Concorde [BA04b, BA05a]. concrete [Bru05b, JJ05]. Concurrency [CLMW09, HAB^{+}00, BA04a, BA07a, JJ05, Rob03a]. Concurrent [ES08, Bar01, Bar02b, CFJ^{+}02, Ext00, Har02, OB07, Rob06a]. conditions [FKMR04]. Conference [ALRY08, BCK04, CJGU02, CFPL05, DSF^{+}03, DGS06, Ell00b, FKCM01, Fin01, Fin02, Fin03, HPT07, Fra02, WMM02, Whi09]. Conferencing [SJ02, RBG02]. confidence [BHM^{+}06b, Ira04, Kum08a, PMC^{+}04]. configuration [Xiz01]. configurations [Min09]. Configuring [AL05a]. conflict [MFR^{+}08, MFR^{+}09]. conform [San09a]. conformance [Cle08b]. congestion [Bog04]. connected [Wal08a]. connecting [PD03, TM03]. connection [DCP00]. connections [Bro08, SV03, Sud08]. CONNJUR [Ell06b]. Conservatively [Reg00]. consider [Kum09b]. consideration [McG09]. considerations [GS05b]. considered [BS00, Got00a, HGH^{+}01, Mas03, MMK02]. considering [PT02]. Consolidating [DGHC03]. constant [Cha03b]. constraints [GD05]. construct [BDMY05]. constructed [Fle00]. Constructing [Gui05, PSE^{+}05, GHBG^{+}03a, Lau06, Mér09b]. Construction [HL04a, Dem05, DLP08, Geo02, dOG07, Hen05c, MPD^{+}05, MGH07, Ruc07, SV03, TMSZ07, Yue07]. constructions [DKN08]. Constructive [HYN03, IS09, BPK^{+}01]. Constructivism [Had05, LKPC04]. Constructivist [ML04, Gib00, MMP05, PB03a, Reb05, Vat01]. constructivists [Lis05b]. constructs [LS08c]. Contemplate [McC04a]. Content [Dal05, BC04c, BAdC07, BdCGB07, CP00, HB01, LH05, NBL09, PHMS08, Rao05, SSC01, ZJ03]. contents [ARVV06, dBGdMB09]. contest [AHL04, Ros09]. contests [JWC08]. context [BC04a, Cle02a, DH06, DUM^{+}05, EMM^{+}06a, KvdHR05a, KvdHR05b, KvD06, LTHN09, MK06, Mic00, MBK09, PMC05, Poh05, RC02, SBZ09, Xu08]. context-based [SBZ09]. continental [Sha00]. continued [GCD^{+}09, PSS09]. continuing [Pop07a]. continuous [LDHW02]. contracts [GHBG^{+}03a]. Contrasting [ZIBG05]. Contributing [HCJ^{+}08, WS01b, Ham06]. contributions [GSW08, wH07]. control [Bog04, CKM07, Gin01c, GSVG05, HCMM00, LBA08, RW05, WGS05]. controlled [Ste09a]. controller [XM06]. controls [Hu03]. Controversy [Bru04, Bru05a]. convergence [DW05a]. converging [Pul00]. conversation [Kla05a]. conversational [WJD03]. conversion [Gin01b]. cool [La 06, Nyg02]. cooperation [BVPE06, Cle00a, Haz03c]. Cooperative [BCM05, Mar09b, PMF08, BC08, CO00, GDHW06, Gon06, Hug09, McC05b, McC05a, McC06c, McC06a, McC06d, McC06b, PJ01, SJA04]. coordinated [CS00b]. cope [GET06]. copy [Car03]. copyright [GPP00]. Core [BMR04, Dem05, Lea01, PSE^{+}05, WW09]. corn [Sna06b]. corner [Haw08a, Haw09a, Haw09b]. cornering [PWH00]. Corporate [Har03, Bri09, BC04d]. corporate-ready [BC04d]. corpus [TMSZ07]. corrected [Sar08]. corrected-errors [Sar08]. Correcting [HMRM03, Chu03]. correction [Chu01]. correctness [GHBG^{+}03a, GHBG^{+}03b, Hen08a, Hen08b, SBR^{+}07]. cost [SC07, ZF04]. costs [WAS^{+}02]. could [ESP03, KBC01, MBD04]. countries [PV08]. counts [Hen08b, Hen09a, Hen09b]. County [CRPZ06]. Coupling [Van04]. Course [Dei08, Hei03, JCG06, LS04b, MFL01, POM^{+}08, Aba08, AGGR04, AT04, AL05a, ABB^{+}06, Aho08, AKM00, ABAA09, AA04, And07, Ari00, ARVV06, AGE06, Arm06, Arm08, BMW04, BHBG07, BMK09, BL01a, BUCP02, Bar01, Bar02b, BED^{+}03, BHY09, BZ07, BBMW03, Ber01, BLOM06, BCS^{+}07, BVPE06, Bla09a, BLA^{+}09e, BBCD09, BDMY05, BCC^{+}05, Bou03, BC04d, Bud06, CW03, CHDH02, Can06, CK02, CD08, CFR04, CB00, CM02, Chi05, Chr09, Cig03, CTR^{+}03, CK08, Cle08a, Cli06, Cli01, CKM07, Coh07b, CKMP09, CHP^{+}05, CR02, CGR^{+}08, CM06b, CI03, CN07, Cun00, Cun02, D’A03, DO07, DDL03, DR07, DV04, Dei07, DLRHP09, Dic05, DTZZ05, Doo08, DKR09, Ega07, Eis03, Ell02, ES00, Eps06, ESW09, FK04, GEVZ03, GET05, GE09, GEVZ09b, GMFA09, GV09, GV06, GM08]. course [Goe03, GL07a, Gol04b, Gou06, Guz03, GF05, HH07, HHB05, HBH07, Haz03b, HL04a, HL06, HBIG03, HPCZ04, Hol00, HWH^{+}05, HAFT09, HR00, Imb03, Ira04, Jac08, JJK03, JHLT06, Jon00, KH05a, KNT02, KBC01, KT01, KPT03, KvD02, Kos00, Kum05c, KKR06, LH03, Lar06, LL04, LAKE07, Lau07, LNCC09, LB08a, LBC04, LCPQR05, Lif04a, Lif04b, LAP08, LHS04, LS08b, LS01, LNR05, LKPC04, Mar05b, Mar06, MK06, Mat03, Mau02, MWBF02, Mea06, MSTS02, MAFLR05, MAFLR06, Mer01, Moo02, MPD^{+}05, MB05, MH05, Mus06, NLCvL09, NH05, NH07, Nør00, Noy02, O’L06, Old08, PB05a, Pap03c, PRK05, Par06a, Pas05, PHBM05, PL04a, POMA08, PQEA^{+}04, PSW03, PG07, Pil04, Pil09, Pol01, Pop09b, Pro00, RH03, Rag04, Raj05, RN03, Rao05, Rao06, RM01, RPG04]. course [RWGM00, Rob00a, RBFR06, RLR07, Roi05, Ros09, Ros00, RNH^{+}06, RRR02, RRRH04, SCHS09, SL07a, SL07b, Sch04, SNA09, Sha04, Sha06, SL00a, SSC01, Shu04, Shu05, Shu06a, Ska02, SLSM02, SSP^{+}05a, Sor05, SRR00, Sta05, Ste03, SHS^{+}08, Sur01, SK05, TYT^{+}09, TIKT04, Tra03, TMSZ07, Tre02, TN02, Tut00, TBM09, UD01, VI01, VI05, Ver05, WM03a, WSC03, WW04, Wai06, WJJD06, Wal09a, WAS^{+}02, Wan01, Way05, WS02, WWK^{+}05, WG07, WS01a, WF08, WS01b, Wis05, Wol07, WK07, Wos06, YLYX06, YD04, YA05, ZJ03, ZSW04, ZE01, ZH02, vW00b]. course-automated [TIKT04]. course-project [HBH07]. CourseMaster [HST02]. Courses [ES05b, JT04, ADG^{+}03, AFJ^{+}03, AMUFVI09, ADP01, AAS^{+}04, AGE05, AGEH06, BvdH09, BGD03b, Bay09, BKM^{+}01, Ber03b, Ber09b, BKW05, BP09a, BSL^{+}08, BCB09, BBS06, CG05, CL08a, CFJ^{+}02, CGY^{+}01, Con04, CHS05, CT07, CT09, CBH09, Dav09, DV07, ES05c, ET02, Erl06, EGMS08, FH01, FJ06, GE08b, GEVZ09a, GKHPQ07, Geh01, Gha07, GAGR05, GGH^{+}08, GD08, Goo03, GG06b, GBO08, Ham05, HB01, Hel07a, HPM^{+}08, His09, HKF00, IL04, JKS00, JS08b, JA03, JY01a, JBB^{+}03, Kla00, Kla02a, KC07, KDH^{+}00, KEK^{+}07, KH05b, KS04, LLP^{+}09, MKS02, MS05a, MDM04, MWF^{+}02, NCA^{+}02, Par06b, PS06, POM07, PBOP07, Rad07, RMZG06, Rob00e, SD03b, SJ02, SFPFV08, SAHS04, SW04, SSP^{+}05c, SHP^{+}06, Ste09a, SGKL^{+}01, TYC^{+}07, TSC^{+}02, VIPCUF08, Wal00a, WR07, YA05]. courses [vLK09, vTNC08]. courseware [DUK02, Eth04, Hel07a, wH09b, JFH00, LR01]. coursework [SJS^{+}07]. Coverage [DLRHP09]. Covington [RHDR07]. CPATH [CWP09]. crafting [Haw08a]. Crawfish [Gig02]. CRC [GGR03]. CRCD [GCW^{+}03]. create [LS08b, SL09a]. creates [BC04d]. Creating [Eis03, GDB03, HM05, HPM^{+}08, KvD06, MF07, SM02b, Aba08, ABE^{+}02, MMN^{+}09, SRGF08, WBBC09, YH03]. creation [ABL07, ABAE^{+}09, BP09b, Cha06a, DAH08, JKM07, OB05, Shu06b, WP06a]. Creative [HAM00, Gin08b, LJG05, Mas05]. Creativity [KR08, cLwW09, RB09]. credit [OHRL01]. crevasse [Dic07c]. crickets [XM06]. crisis [CMGR07]. criteria [CCMS05, CI03, Gor00a, Gor03a, Gor03b, Gor04a, LKGY00, SDR09, ZRY05, ZRY06]. criterion [LL03]. criterion-referencing [LL03]. critical [Boz06, KI08, Rao05]. Cross [LAP08, ABE^{+}09b, Bri06, FG08, Gri02c, PRW04, SC08]. Cross-cultural [LAP08, ABE^{+}09b]. cross-disciplinary [Bri06, FG08]. cross-domain [SC08]. cross-over [PRW04]. cross-platform [Gri02c]. crossroads [Lis07a]. cruelty [Cle06b]. Cryptanalysis [HAuR04]. Cryptography [AA04, DFGH04, GV09, McA08]. CS [DHRH05, SMK09, ALR04, Abu04, AF03b, AH08a, AFJ^{+}03, AL05b, AGE05, AGEH06, BP00, BLH01, BAS04, BKM^{+}01, BW06, BDW^{+}05, BKW05, BT05, BBY08, BYS09, Bla09b, BMP06, BC07d, BC04c, BPS09, BDM01a, Bru04, Bru05a, BDM05, Che00, CMS07, Cli01, Coh02a, CSC03, CP00, DBK^{+}09, Dal02, Deb04, DSBK01, DAKLH07, DLHAK08, Ega05, EMdL^{+}07, EGB05, EGB07, ES08, FHBD06, GH01, GCDJ08, GM08, Gin04b, GAGR05, GCF04, GCFP04, GL03a, GLR^{+}01, Gri02a, HK01, HHRA06, HWRC08, HEB08, Hen09b, HKF00, HFHK05, HM02, HAFT09, HLCS08, Ira04, JS06, JCG06, JA03, KKLK04, KvdHR05a, KDM^{+}00, KP04, Kül09, LA07, Lap07, Lee01b, LR04, Lis08a, Lis09a, Lis09b, MNN^{+}09, Mal09, Man07, MR00, MAD^{+}01a, MAD^{+}01b, MMLT09, MDL^{+}02, MD06b, MGFK06, Med01, MN01, MH05, Mur07b]. CS [Ni09, NKI^{+}09, NCA^{+}02, OBLH03, Par05a, PN08, Pea02, PMC^{+}04, PW08, Reb05, RMB00, Ric09, RCE^{+}00, RVWB04, Rus09, SW06a, Sha03, SW04, SLSM02, ST08, SSP^{+}05c, SGC^{+}01, SM07, SdSK05, TBAA09, TFG05, Tra03, TS02, TW09b, Val04, VA02, Vas02, Wal05b, WW09, WAC^{+}08, WP07, WMP08, WLMM09]. CS-1 [AF03b, WAC^{+}08, BLH01, Kül09]. CS-complete [GH01]. CS/ACM [IHL^{+}04, SL03]. CS/IS [HFHK05, SM07]. CS/SE [Cli01, JS06]. CS0 [BF05, Cli06, DTZZ05, EBG^{+}05, Hic04, HAB^{+}08, Kru03, PEH07, Ree01, TA08]. CS1 [All03, ABE^{+}02, AV02b, App01, BRP^{+}07, BS06a, BCM05, BC08, Bec01, BC04a, BU09, Bla09c, BCB^{+}08, BDRH09, BDM01b, CG02, CC02, Cic09, CCG05, CR02, CLP06, CH06, Dal05, Dal06, DVE06, DJC09, DTZZ05, Djo09, DW05b, EPM09, Eth04, Fit09, GEVZ03, GELS04a, GEVZ09a, GEVZ09b, Gao00, Gig03, Gin00a, GL07a, GL08b, Gon06, Gri00, GG06b, HMDK04, Han05a, HMS^{+}09, HW08, HL04c, HTW04, HP06, HAB^{+}08, Hun03, JS08a, LNH09, Law01, LDB^{+}03a, LJM05, Lis01b, LH02, LS08c, LRD09, MD06a, MSP^{+}06, MD04, MD05, MO09, MGE^{+}05, MK02, MB05, Mos06, Mur07a, Mur07b, NWF^{+}03, NSS^{+}05, PW01, Rad08, RM08a, RJAH04, Reg00, Reg02a, Reg06, RPG04, RKDM01, Rou02, Rou06, RRR02, RRRH04, SCHS09, ST07, Sch00, Sch09a, SSP^{+}05a, SSP^{+}05b, Soh06b, Soh06a, SW06b, Sto06]. CS1 [SM08, SMF09, SKO^{+}09, TA08, TBS^{+}00, TMS07, Tow09, TG05, VMD01, VR04, VZGE07, Wal07b, WVMN05, WN05, Wic01, Wic05, Yue07]. CS1-CS2 [AV02b]. CS1/CS2 [DJC09, Eth04, HTW04, LJM05, RKDM01, Rou06, Wal07b]. CS102 [PB03a]. CS2 [Ada09, ABE^{+}02, AV02b, Ars09, BP09b, CTLW03, Cor07b, CN07, CH06, Cut07, DJC09, EPM09, Eth04, Gig03, Gon06, GG06b, HTW04, Hun03, KB01, KB04b, LCPQR05, LM06, LJM05, LH02, MGE^{+}05, NM02, PB09a, RM08a, RF00a, Reg02a, Reg06, RKDM01, RKD02, Rou06, SJ05, SJ06, TBS^{+}00, TS00, Wal07b, WKB02]. CS2/CS7 [BP09b, PB09a]. CS2008 [CCD^{+}08b]. CS262 [Bla09b]. CS4HS [BC07d, BCLW08]. CS7 [BP09b, PB09a]. CS SDI [BPT^{+}04]. CSAB [Gor00a, LSMB00]. CSbots [LNH09]. CSCL [WHCT09]. CSE [BR05b, CRPZ06, CWC09, Hal08]. CSEd [Lis07a, PD03]. CSO [OJJ00]. CT [CPT^{+}09b]. cues [Kac09a]. Cultural [LZL08, ABE^{+}09b, Kol08a, LAP08, LGA^{+}01]. Culture [FHBD06, DDA^{+}07, How03, Ira04, KSE02, Mar00a, MNBK03, WJDL04]. cultures [Jon02]. CUPM [Wal03b]. CUPV [KS00]. Curb [JJOA07]. Curing [GMS09]. curious [Jon02]. current [Cas01, CMGR07, CNP^{+}09, FKMR04, HHRA06]. Curricula [RCD^{+}02, SMS^{+}06, CR03, DMI^{+}09, DDLR01, Den04, DJC09, FG07, Gor01b, Gor04b, Gra07, Haw08a, Haw09a, Hen05a, ISMS03, IHL^{+}04, Imp05a, Imp06a, ISC^{+}09, LD04, LZBO09, Mar02, Mar03, MMLT09, POM07, Ral05, RJM^{+}05, REC^{+}01, Sam02, SCC^{+}04, SL03, SS00, WBS04]. Curricular [CHK03a, CHK05, HCK07, HKCW08, HKCW09, PH04c, SBZ09]. Curriculum [CCD^{+}08b, Lie05, MvW01, SMTH00, AHJ^{+}07, BBF^{+}09, BBW04, BBN^{+}04, Bry09, BS01, BEE07, BF06, Bur04, BDDL04, CGB08, Cam02b, Cam05a, Cam06a, Cha05, CS05, Cha06b, CBJL08, Chr03, CDGS01, CKMP09, CKLW05, Con00, Cor07b, CFKS02, CCK01, Cun07, Cur03, DW05a, DGMW07, Deb04, DDLR01, DUM^{+}05, DDA^{+}07, DRKG02, DTL^{+}02, FK05, FJ06, FSAF03, FIG07, GT00, Gab03, Gol02, GCF^{+}01, GPF^{+}01, Gol01, Gor00b, Gor01a, Gor06a, GGS^{+}06b, Gou08, Gra03, HWRC08, HLM01, Haw08b, HKW^{+}02, Hic04, HT00, HHLS06, HRS06, HH05b, HFHK05, HM02, Imp02a, Imp03, JS06, Jon01, KKLK04, KTH^{+}00, Lee00b, Lee01a, Lee02b, Les04, LGA^{+}01, LC05, MNN^{+}09, Mar02, Mar03, MDH^{+}00, McG09, MD06b, MTHS00, MB05, MP09b, NRRR06, Ni09, Our04, PDC01, PM06, PSW03, Ram02, RD08, RJM^{+}05, RSK06]. curriculum [RHKS01, RMB00, RMMH01, RR03, RS03, RTH^{+}08, Sam02, SM00a, SSC00a, SF02, Sob00, SHL^{+}00, Top07a, Top07b, TKB01, TDJ^{+}03, TMD^{+}04, Tur01, VCC^{+}05, VCC^{+}07, Wai06, WSW05, WTTT^{+}08, YSLL00, YA05, ZLL^{+}07, dMBM08, vVML04, RCE^{+}00]. curriculums [Dem05, PRC06]. curse [Qua06]. Curve [FLS04a]. curves [Ras01]. custom [ESW09]. customizable [CL08b]. customized [GA03, LR01]. cut [NW06]. Cutting [Ast07a]. cyber [BBMW03, CHK03b, CD01b]. cyber-icebreaker [CD01b]. cyber-security [CHK03b]. Cybercivics [OM08]. cyberinfrastructure [NSW08]. cyberporn [Old08]. Cybersecurity [Cam03b, MR02, Shu04]. Cyberspace [CF02]. cyberTech [PB06, PBWS06, Wos06]. cyberTech-ITEST [PB06, PBWS06]. cyberwar [MR02, WW04]. cycle [Got03]. cycles [Gin06a, NCA^{+}02]. cyclic [Chu01].

D [CPJ07, CF02, Hu05, LM05, Sei09]. daily [BU09, Cle02a]. dance [Pat06]. dancing [Pat06]. Dangers [MT08]. dark [BCD08, Mar00b]. Data [EVS07, Kum09a, ND08, Ada09, Aha00, BMG03, Bud06, CCL^{+}07, CCH03a, Cam04a, CSM07b, CHJB07, DeP06, DUM^{+}05, EA00, ESW09, Fek02, GEVZ09a, GMFA09, Gig02, GSS00, HTK02, HCB04, HL08, Hoo08, JG08, KMSB08, LBM^{+}04, Lis09a, LCS^{+}09, LZL08, MS00, MR06, MHW07, MD08, MSK09, MT05, Mus06, NM04, NM02, NBF^{+}08, Oka05, PHBM05, Poh05, Rah08, Roi05, RTH^{+}08, Sha04, She06, Shi06, Ten03, TE08, WSC03, Wag05, WSP04, GKP06]. data-driven [GSS00]. Database [PH03, SGMJ09, AGGR04, AYWM08, Ber03b, BSL^{+}08, BBS06, DUM^{+}05, GV06, Gui05, KP05, MKB05, Mer00, Moo02, MICM08, PFF04, PBK04, RN03, RWGM00, Rob00a, RR03, Sci07, SYLL00, SRR00, Udo06, UD01, UD03, WM03a, WSC03, Wan01, Yan09b]. Databases [Goe08, DSCU01, GLC07, TM04]. datapath [VS07b]. daughters [Pop08b]. Day [BCS^{+}07, GL08a, GL09, Par05a, Rol02]. Day/Stout/Warren [Rol02]. days [Bla09c, FM06, Gri02b, Soh04]. DBMS [RN03]. DBMSs [Wan01]. dead. [Cle03b]. Deaf [CLR09, BAdC07, BdCGB07, KW07]. dealing [Aha00]. dean [BMP06]. Dear [Edm07a]. debate [LS05, LBC^{+}06, Pre07]. debugger [RM07]. Debugging [CL04, MLM^{+}08, SFM^{+}07, AEH05, Eth04, LH05, MKKC08, Ric00]. Debunking [Mar00a, GL03a]. decade [SRR00]. decades [SRR00]. deception [Got09b]. decide [ZTS^{+}09]. decision [Bri09, DWJ02, SSA08]. decision-making [Bri09, SSA08]. declarative [RHKS01]. declare [RHKS01]. decline [BLOM06]. decomposition [Gin02c, MGH07, Soo09]. decreasing [WAS^{+}02]. deep [Car03, Vas02]. defend [TT06]. Defense [BCD08, LW07]. Defensive [BGDJ02]. defined [BA07b]. Defining [McG00, Pea02, ISC^{+}09]. definition [Gor01b, RVP08]. definitions [CSD02, DSB^{+}03]. Degenerate [Ber09a]. degree [BGD03a, BBN^{+}04, Cam05a, Cam06a, DH06, DSBK01, GGS^{+}06b, HM00, Haw08a, Haw08b, HE06, Kle02, Top09a, TLW^{+}04, TLW^{+}09, Wan06]. degrees [ACL^{+}03, CCD^{+}08b, CDS09, Cle00b, ELE^{+}02, SM00b]. delivery [AM07, BvdH09, Cle08a, DDL03, Lea01, SSC01]. Delphi [GGH^{+}08]. delta [Gin04a]. demand [PRC06, PW01]. demand-driven [PRC06]. Democracy [Bla09c, Bor02]. demographic [Kum09b]. demonstrate [Fon01, GMGM06]. demonstrating [Rol06a]. demonstration [BBK05, BLN06, DUK02, Hen06b, Pul07, RRP02]. demonstrations [Ell00a, Liu06]. Denmark [CJGU02, Fin02]. Denormalization [BS02]. denotational [Gol08]. Department [BPT^{+}04, CP00, DABJ00, DWY03, PH04c]. departmental [MNBK03]. departments [CSC03, KDM^{+}00, LC08, MBPB03, RMPD09]. dependence [PRW04]. dependencies [MG04]. dependent [Don01]. dependent/independent [Gib00]. deployment [Bor07, LNH09]. depth [Mur01, Mur07b]. depth-first [Mur01, Mur07b]. Dereferee [AEPQ09]. descent [Sax03]. describing [POM07]. description [KYC07]. descriptions [LL02, Wal09a]. desert [PN08]. Design [AF03b, BS00, CWHL06, Cle02a, DHRH05, FK02, Fon04, GF05, Jon00, Lev00, MKB05, MCS04, Mer09a, Mos03, NW03, NW00, NW01, NW02, NRW04, NRW05, NSS^{+}05, PG06, Ree04, Rou02, Sma07, SP03, WJJD06, YD04, AV02b, ACD07, AT01, Arm06, BvdH09, BL01a, BCM05, Ber01, Boi04, BC04b, BROB00, BGG^{+}00, CCMS05, CMS06b, CC01, Chr04, Chr05a, Chr09, CC05, Cle05b, CK04, CKLW05, CKMW02, Con08, CM06b, Cox05, Dav09, Deb02, DFGM08, Dew05, DGHC03, Dic05, DSBK01, EMM^{+}06b, FLL^{+}01, Gao00, GP07, Ges07, GGAB01, Gin03c, Gin04c, GA06, Gin07d, Gin08b, Gin09d, GMJDA09, GGR03, Ham02b, Ham04b, Ham05, Han04, HTSW07, Hol05, HS03, Huf02, Ing09a, Ing09b, JMS02, KvdHR05a, KvdMvDvdV00, KA09, LV03, LBB^{+}01, LNCC09, LNH09, Laz02b, Laz02a, LRPQ04, Lom07, Mar07c, MH08, May08]. design [MSTS02, MAFLR05, MAFLR06, MWM01, MICM08, MB05, MWPB05, NKI^{+}09, ÓT04, Ort08, Pas05, PH04b, PHN06, PH07, PPP06, PSW03, PK08, Pro00, Pro09, Qua07, RD08, RJM^{+}05, RRP01, RF00a, RI08, RKD02, Ros06, Ruc07, SL09a, SL07b, SL00b, SNS07, SSP^{+}05b, SF04, Sur01, TRT06, Tra07, Tre02, Try05, TQCPQE08, Vat01, WS02, Wei05a, Wic01, WP02, WSP04, Wic05, Wol07, Wol01, vW00b]. design-by-keyword [Gin03c]. design-code [HTSW07]. design-first [MB05, MWPB05]. design-patterns-first [PPP06]. designed [Gig03, LNH09, LNR03, RPG04]. designers [Kan07, Rya01, WG07]. Designing [AEPQ06, AAC^{+}05, BL01b, CBP03, CBJL08, SSP^{+}05c, WW04, AMW06, MKS02, MH00, SHP^{+}06, Sta06]. DesignMentor [SL00b]. designs [CH09, RTET03]. Desirable [RP06]. desk [AB08]. despite [BAMSK05]. destroyed [BA01]. detecting [JWC07, JWC08, NW06]. detection [CCH^{+}03c, ELC07, EG08, GS06a, LC01, Lar06]. detective [DWJ04, DJWD05]. Determinants [To04, FHBD06]. determination [PA03]. Develop [Les03, SL06, SL07b, CFL^{+}04, CI03, FH01, Fek08, GMN02, Ham04a, HKN^{+}02, LR01]. developed [RM08b]. developer [McG09]. developers [NBF^{+}08]. Developing [BH04, CFKS02, EBC^{+}08, FP09, FJA^{+}07, HHS02, cLwW09, LR08, MD06b, MSS^{+}04, Pil03, Qua06, SDR09, SW06b, VCC^{+}05, VCC^{+}07, VCJ09, Bro08, CGI05, FLK^{+}09, HEM09, KHPC08, RKD02]. Development [CGR^{+}08, HR03, Lin05, NRB^{+}05, OTUW07b, ABA09, BS08, BCD^{+}04, BKR^{+}05, Boi06, Bor07, BDF^{+}08, BS00, BS01, COR07a, CDGS01, Ciz03, Cli01, Con08, CSP09, DJS08, DJC09, Doo08, EMdL^{+}07, FK02, FLL^{+}01, Fil01, FSAF03, GE09, GPF^{+}01, Gol01, GMGM06, Got04, GD07, Gou08, HY09, Had01, HALW01, Hau05, Haz05, HBIG03, HHF^{+}03, Hum04, Imp03, Kac09b, Kla00, KI08, KB01, Kur09a, Kur09b, LB00, Lee01b, LM06, Lie05, Mar07b, Mat06, Mau02, McC01a, McG09, MAFLR05, MAFLR06, MW09, MTHS00, MR09, MLH05, Mon02, NH07, NSS^{+}05, Ols06, Pap03d, Par06a, PMC05, PWC^{+}05, PSS09, Pol01, RMY08, Rob00c, RD00, RIP^{+}09, SN08, Sch09a, Shu06a, SMTH00, SBZ09, Tur01, TS00, Van08, Wis06, WP07, YD04, vTNC08]. device [War04]. devices [Esp08, HLN07, MNN^{+}09, RA06, RW02, Ric02, TBM09]. DFS [BL05a, EG08]. DGE [SLPG07]. diagnosis [FKMR04]. diagnostic [Cle02b, PBDWL09]. diagram [HB06, TWS09]. diagrams [Dic07c, HCMM00, HL04c, TRT04, TWS05, TWS06, TSW07, UD03]. dialog [AKWZ07, DBBK02]. dialogue [LV03, SBG^{+}08]. dialogue-based [LV03]. Dickens [Lis06a]. did [Lee02c]. Didactic [CGM04, Xin03]. didactics [ZV04]. didn’t [Got09a]. difference [Abo09, Cle00b, HV07, KMMS02, Tow02]. Differences [Per09, BRP^{+}03, BDW^{+}05, CJ01a, CJ02, LCPQR05, LW05, MNBK03, MRM^{+}06, RPH^{+}03, SPM00, VZ06, VHF07]. different [BW06, CC02, CY01, CS03a, GEVZ03, GE09, HRBC03, LM03, PR02, Son09, WW00, ZIBG05]. differentiate [CBC09]. difficult [Dal06, Fon01, GGH^{+}08, HA02, LS03, Mas05, VI00, Yue07]. Difficulties [MHK03a, WVMN05, CM01, HA02, Kol08a, LAMJ05, Pil09, SBM05, TRT06]. difficulty [SB09]. Digital [CIL^{+}03, EGST08, Rad07, Rad08, WBPW06, WBS04, AHJ^{+}07, AH08b, BR09, CIK^{+}03, DKG06, ICL02, Joe06b, Kur08b, MV05, MICM08, Rad06, RTM^{+}04, SD09, WGS05, Wol07, vLK09]. dilemma [HLNT01, RKDM01]. dimension [LR07]. Dimensional [Bur03]. Dimensions [CHM^{+}01, LBM^{+}04, MH08, SB08]. dinosaur [Lab09]. DIODE [EBG^{+}05]. direct [KH05b]. directed [Ell06a]. direction [YSLL00]. directions [Owe08]. DIS [ND08]. disabilities [Dei07, Dei08, FS02, LV08, NL07]. disability [GHG^{+}08]. disadvantaging [Bec03a]. DisASTer [OG05]. disciplinary [Bri06, FG08, LW05, LLLC06, PBM03, TF07b]. discipline [HDH06b, Lew07b, SR07, Top08b]. Disciplined [Cle05b]. disciplines [CMS06a, CWP09, CBC09, Coy04, LLLC06, Pop02]. discourse [HAG05, LL00, Lis09b]. discoveries [BAK^{+}09]. Discovering [Cha00b, Geo00, Tal06]. discovery [BF06, MG04, Ola00, Sie00]. Discrete [LL06a, MAR07d, AHH^{+}06, BMW04, DV04, Gol04b, HDH06a, Hen07b, HD07, JJK03, Kos00, Mar05b, Mar06, SWR05, TBS^{+}00, WW06b]. discrimination [Cle01a]. discussion [Bru04, Bru05a, Cla00, Dia05, HTK02, JY01a, San05, SMR03, SB06, VA02]. discussions [BKW05, Cra07, GL03b]. Disequilibration [BR02]. dishes [YM06]. dishonesty [CCH^{+}03c, PHZ^{+}03]. disk [DSSP06, Rob04a]. disorders [Gin03d]. disruptive [OGMA07]. dissemination [BF06]. Distance [Dav01a, Dou00, GE08a, GE09, GPP00, Pet09, vW00b, AJD09, CHM^{+}01, DDLFS01, Dav01b, ET02, GE06, GE07, HE06, JKS00, Kac01b, Kac01a, KV08, KvdMvDvdV00, MTBS09, MICM08, MPK08b, Pul06, PC08, RH03, RBG02, RBlR01, Ros01, Sha00, TL01, WPS01]. distance-learning [ET02]. distinct [BP09a]. distinction [Sam05]. distribute [LR01]. Distributed [CHK^{+}01, Hau05, Kos00, PC08, Sch02b, She03b, Aba08, Alb09, BDF^{+}08, CWP09, CB04, Han05a, HALW01, HA04, HKL^{+}02, HWH^{+}05, KPT00, KT01, KPT03, Las03, MW09, NB01, OG05, Phe07, SP07, VB02, WKC^{+}09, Wis05, WB05]. distribution [AR08, SM00b]. District [CRPZ06]. ditch [Kac09a]. diverse [AL09, CR02, RHL^{+}09]. Diversifying [Fri05]. diversity [BHYA05, Coh07b, Goo08, GHLL08, JM02, MBD04, Wic09]. division [Bec08]. Divisor [Gin02b]. DLSim [SD08]. Do [Boy00, BLW01, Cha03b, CBC09, Gin04b, HCMM00, MGFK06, MO09, MT05, PR02, Ral05, Wal03a, WG07, Wil00, ABH^{+}07, Bar05, BABBH04, BCS^{+}07, BEM^{+}07, CS02, CLM^{+}06, Cle06a, Cle09a, DSE02, FG07, Got09a, Kac07b, KB08, LRPQ04, MAFW08, MPH09, OPS^{+}02, PT02, RHKS01, Ric09, SSA08, Wal04c, Wal07b, WB09, WBBC09]. doctoral [CWC09]. doctorate [MSS^{+}04]. doctorates [Mos02]. Documentation [Cle03a, Gib02]. documenting [Cas03]. documents [SL09a]. Does [HM00, KMMS02, Wil04a, DGSW06, DKL^{+}07, GG06a, Mos02, PHW07, RPR03, Rol06b]. Doesn’t [MHK^{+}03b, BA04b, BA05a, FMK^{+}00]. doing [BG08a, RW05]. DoIT [Ram02]. domain [Gol08, Hen05c, Mir09, SC08]. domains [HB06]. Domino [Gin07b]. DOMOSIMCOL [BROB00]. domotic [BROB00]. Don’t [Got07, RR06, BB06, Car06, Dic07a, Edm08a, GAGG04, LBA07]. door [TF07b]. Dorothy [Edm07a]. doses [Jon01]. doubly [Mur07b]. down [ABE^{+}02, GOMSJVGP08, Gin01e, Got05, MVC00, WF08]. Draft [Gor04a]. drag [Ber06]. dramatic [Wal01c, Wal02b]. Draw [Mar04a, EMdL^{+}07]. Drawing [PHC08]. dream [Rob04b]. Drink [Kac08b]. drive [BC06b, BCB09]. driven [Ada09, BSV04, BDM01b, CC02, DJS08, DJC09, GSS00, Ing09b, JS06, JS08b, KI08, NRDP07, PRC06, PRK05, PHC09, Pro09, Sch09a, SNH08]. drizzle [EBG^{+}05]. DrJava [ACS02]. drop [Ber06]. Dropping [Man07]. DrProject [RW07]. DSW [Rol02]. Dubai [Mar05d]. Dundee [HPT07]. DUPLEX [CCH^{+}03b]. during [Dei05]. dying [Pau08]. Dynamic [CHJB07, BC09, CH06, HCB04, JDGAGMGC05, Pho01, PWH00, Ram02, Rol07, RIP^{+}09, SNH08, WPS01]. Dynamically [Kum00]. dynamics [JP03, RBA05]. Dyslexia [PMG^{+}04]. Dzver [Pan09].

E-1 [Mal07]. e-assessment [APR06, AFR08]. e-books [WC04]. e-citizen [Di 06]. E-commerce [GS00]. e-IMC [LR01]. e-learning [APR07, BAdC07, Dav06, GGÁC03, LW05, MBD04, Ols05, SKM02, Tra03, BPK^{+}01, BdCGB07, TSO05, dBGdMB09]. E-notes [Wir03]. e-readiness [GM07]. e-tailing [To04]. e-voting [PL05, Got06a]. E.H.E.A. [PG07]. Early [NW03, ACL^{+}03, ABK^{+}05, BBK05, BWC^{+}05, BS00, BHM^{+}06b, CMS06b, GAGR05, HGH^{+}01, JS08b, Lie05, LBC^{+}06, MS05a, Mar07c, MD06b, PQEA^{+}04, RM08a]. earned [Mos02]. Easy [RKS07, YLS05, AFJ^{+}03, BH03, Phe07, RH08b]. easyaccept [SN08]. Eccles [EH07]. Eclipse [Deu08a, Deu08b]. economic [Gre00a]. economical [AH08a]. ed [Dal02, DGHC03]. Ed-tech [DGHC03]. editor [Ber06, CCJ^{+}02]. editorial [Tur01]. Educating [Di 06, Joy06a, AKWZ07, CHK03b, Mar05d, Mik05]. Education [ACM01, ACM06, ALRY08, Ano03, BTHR06, BCK04, CJGU02, CDWH00, CIL^{+}03, CCD^{+}08b, CFPL05, DSF^{+}03, DNTB05, DCP^{+}09, DGS06, DRFG08, Ell00b, FKCM01, Fin01, Fin02, Fin03, GWG02, GKJD03, HCK07, HPT07, Imp02b, JKDN04, MFL01, Ran07a, RHDR07, Top09b, WMGR01, AH08a, ASA00, AGHM02, AGHC03, AHGP05, AAEI08, ABS08, APR06, AJD09, AM07, BG08a, BABBH04, BKR^{+}05, Bor02, BN00, BPS09, BBE^{+}09, BKS^{+}04, BBT07, BS04, Cam01, Cam03a, CHK03a, CHK05, CHK06, Cap02, Car05a, Car02b, CCD^{+}08a, CIK^{+}03, CDF^{+}07, CSD^{+}07, CBP03, Cle00a, Cle00b, CHM^{+}01, Cle07a, Cle08b, CCMC05, CNP^{+}09, CS03b, CS06, DDLFS01, Dav01b, DG05, DPM00, DW07, EMM^{+}06a, Edm08b, ESP03, EGB05, EGB07, FK05, FS09, FdMFPT07, GE06, GE08a, GE09, GLA^{+}08, GS00, GCDJ08, GPP00, Gha05a, GCF04, GCFP04]. education [GPF^{+}01, Got00a, GLD^{+}05, GK02, Gri02a, GKP06, GS09, Had05, Han07, HEB08, Haw09b, HL04b, HDE^{+}06, HGEB08, Hep06, HETD09, HDH06b, HY01, HMS04, Hug09, HNC08, IMRJ09, Imp04, Joe06a, Jon00, Kac01b, Kac01a, Kac06, KHPC08, Ker04, Knu03, Kol08a, KI08, KV08, KvdMvDvdV00, KAPG07, KSRD05, KPN02, LdLM^{+}08, LH05, LML^{+}01, LCW06, Lee00b, Lee01a, Lee02a, Les04, BGN09, Lis04a, MMP05, MSM08, Mar01b, MUPVPF^{+}02, McC00a, MGFK06, Med04, Med01, MTBS09, MWG08, MS03, MMRI07, MvW01, MT08, NRA^{+}03, NP02a, NSH^{+}09, OCM01, Par05a, PHM^{+}01, Pea02, PSE^{+}05, Pet09, PV08, PD06, Por09, Pul06, PC08, RF09, Ran07b, Rob00d, Rob04b, RR06, RDCE06, RBTC03, RJM^{+}08, RK09b, Sae09, Sal03a, Sal03b, SB08, SB09, SBG^{+}08, SNH08, SRGF08, SSC00b, SHN^{+}00, Sic09]. education [SBM05, SBZ09, SGC^{+}01, SM02c, SdSK05, TS08, TFIJ08, TFG05, Top08a, Top09a, Tra07, THH09, UHFT09, VS06, VKV02, Wal02a, Wir02, YM02, DTW07]. education-friendly [MWG08]. education-oriented [VS06]. Educational [OB05, AMUFVI09, AMVZ06, AS02, BPS09, Bri06, CBKA09, CIL^{+}03, CR03, CKW04, DCT06, EB09, Fil01, HKL^{+}02, IMRJ09, JDC00, KH01, Kno00b, KPT00, Kum09b, Mik05, NCK^{+}03, NvdH04, NB01, PWC^{+}05, PPJH09, PHMS08, RW07, Rob00c, Rya01, Tur07a, Tur07b, Val04, Yai08, YH05]. educationware [Pul07]. educators [ABO^{+}05, AAO^{+}08, BAK^{+}09, BAS04, BDDL04, CY01, Dal06, HBEM03, LNR03, MMD^{+}09, MM00, McC02, OAB07, Pre07]. EduComponents [APR06, APR07]. effect [ABAA09, BGD03a, BL01c, CCG05, GEVZ09a, Gib00, Kum08a, PL04b, RSK06, TBAA09]. Effective [CG05, KPN02, SSMN04, SJ04, BH03, BDMY05, CTP09, CD01b, Esh07, FB02, HAM00, HHS02, LW05, SD00, Shu04]. Effectiveness [WM03b, CW03, CB04, Cic09, FM03a, Gri03, Kum05d, MLC04, PB03a]. effects [BEW08, Gon06, LCPQR05, MWBF02, SK09]. efficacy [RLW04, WHCT09]. efficiency [AKT07, Fon03, GEVZ03, RK09a, WC04]. Efficient [Cha05, BN00, Deb04]. effort [CMT04b]. effortless [AKA09]. efforts [Poh05, WB02]. eFuzion [PWC^{+}05]. Egyptian [GH01]. EHR [CPT09a]. eight [HAM00]. eighteen [Gor00b]. Eighth [RHDR07]. eL-CID [Boi04]. Elaborating [Gin07c, Gin00a]. elaboration [Gin07d]. ElderCare [Eps01]. elearning [KK03]. electives [WW09]. Electronic [Esp08, Geh01, JY01a, MUPVPF^{+}02, She04a, Wea04, AM05, BJJ01, BEE07, CF04a, CF06, CI03, Del00, MDD06, SD00, Tho03b, Wir03]. elementary [CBJL08, KS04, PD06]. elements [DLPQ04, Gin01c, Got02, HGEB08, Wen04]. elicit [SV03]. elicitation [RVP08]. Eliciting [Yan07]. Elucidate [Ext00]. Email [Hob00, San09b]. email-centric [San09b]. embedded [BHH^{+}08, BLN06, Bry08, Cli01, MK06, McC05c, MS07, QLT09, SSP^{+}05b, XM06]. Embedding [Gin04c, LZBO09, WZ00]. Embracing [PBM07]. Embryonic [Ras03]. Emergence [PL04a]. emergent [HM05]. Emerging [KSRD05, KDP06, Pre07]. EMMA [TIKT04]. emperor [Man07]. emphasis [CIL^{+}03, Dal05, Shi02]. Emphasize [JT04]. emphasizing [KvdHR05a, MS05a, SSW08, Ver05]. Empirical [ES06, RMB00, Sob00, BW06, BMR04, Bra05, Gri03, HAFT09, OBL04, RBC^{+}02, RKD02, San02, SB09]. employing [Jac01]. emporium [Lee02a]. empowerment [Goe03]. empty [Gin04e]. Emulab [LFG08]. emulated [DEJ^{+}06, DJLA07]. emulation [GD08]. emulator [CD01a, GDM05, TW05, VVV04]. Emulators [Lee00a]. enabled [Cul00]. Enabling [Hei03, GP02]. Encapsulation [Fle01, WSP04]. Encoding [Wic03]. encounter [Kol08a]. Encouraging [AR00b, Ols06, RKI02, Way06, Ber03a, CT09, GSW08, Rob00e, WB02]. encryption [Ber09a, Gou09, HAuR04]. encryption/cipher [HAuR04]. end [BLOM06, Noo06]. ended [HW08, KMN^{+}03]. endorses [Kno00b]. enemies [Lis07b]. enforcement [Rol03a]. engage [HETD09, Lis09b, NEN00, Tow09]. Engagement [HE07, MH06, MP09b, Cox05, Lau08, MLK07, NRA^{+}03, Poh05, RWL^{+}09]. Engaging [BGGR01, BS05, CM08b, DKN08, Hug09, Joe06a, Kan07, Kur09a, RHL^{+}09, Sig07, SHW^{+}09, FM09, MdL09, SR07, Zil05]. engine [AKT07, BC06b, BSA^{+}05, CSA^{+}07, NM02, SC08]. engineer [FC00]. Engineering [CIL^{+}03, AD01, BG08a, BL01a, BHY09, Bec08, BAET09, Boe06, BKS^{+}04, CM06a, Cam05a, Cam06a, Cap02, Cha00a, CB02, Chr09, Ciz03, CC05, Cle04a, Cle05b, CLK00, Con04, CHS05, CPST02, CA02, Cur03, DPM00, Doo03, EGMS08, FC00, FK05, GP02, Gha05a, GPF^{+}01, Had01, Had05, Haz01b, Haz03b, HHLS06, HP07, HMS08, IK06, ISMS03, Imp03, IHL^{+}04, JKS00, Jam02, Jen08, JBB^{+}03, KvLA09, Kur08a, KFE^{+}07, LCW06, LBC04, Lie05, LHL03, LTHN09, Lud05, LNR05, Mai09, Mar02, Mar03, MDH^{+}00, MPQB^{+}00, MCS04, NvdH04, Our04, PTT06, PTT08, PA03, PM02a, Qid07, Qua06, Ram03, Rao06, RM08a, RA06, RM09, RCR04, RS03, RRD^{+}09, Rut01, SLC^{+}09, SNA09, SSC00a, Shu05, SHVP05, SL08b, Sob00, SL03, Ste01, SK05, TCK09, TSC^{+}02, Way05, WB05]. engineering [Xu08, YM02, GKP06, MMC01]. engineers [Haz05, NSH^{+}09]. English [NSW08]. enhance [ABE^{+}09a, BKW05, CKM07, CKMW02, Gha05a, IL04, KCI03, LW07, Rob00c]. enhanced [CNCL09, LJ08, TIKT04]. enhancement [CT09]. enhances [GMHS01]. Enhancing [GM08, Kla04, KB04a, RF09, RHE04, RMMH01, Rob00a, RTM^{+}04, RJM^{+}08, SDLdQ05, Sur02, WM00, BLH01]. ENIAC [Har04]. enough [BC04c, BDM01b, Got09a, Noy02]. enquiry [AC08]. enquiry-based [AC08]. enrich [MC08]. Enriching [PPE03, BGN09]. enrollment [BLOM06, PTC^{+}01, Raf08]. enrollments [Man07]. Ensemble [BGN09]. ensure [BED^{+}03, Jac00b]. Enter [FC00, WB02]. entering [TT05]. enterprise [Moo02]. Entertainment [KvLA09]. Enthusing [CMCB09, CBKM07]. enumeration [Gin05b]. EnViDiA [HKL^{+}02]. environment [ACS02, APR07, AAC^{+}05, AS02, BC07b, BMK03, BGNM04, BDF^{+}08, BROB00, Bry08, CWHH05, CPJ07, Cha02, CWHL06, CNCL09, Cle05a, CR02, CGI05, DLPQ04, Dix04, EDES05, ET02, Fei07, Fos05, FHBD06, Gib00, GL05, GD08, HW06, HA04, HKL^{+}02, HKF00, Hua03, HS03, KP05, KH01, KPT00, KPT03, LL06b, ML04, MD06b, MR09, NEN00, Ols05, Pan09, PHZ^{+}03, PW01, PP04, PNF00, RM01, Rio02, Ros08, SFPFV08, SBR^{+}07, SKM02, TSO05, TL01, TBMB09, Tru05, Udo06, WK06, Wal05d, WSP02, WGS05, vTNC08]. environment-passing [Fos05]. environments [Arn07, BCB09, Cha01, GDB03, Gol01, HKL^{+}02, LM03, Oli05, PM02b, SPAC^{+}05, WB05]. epiphany [Got02]. epsilon [Pop09a, Pop08a]. equal [HF04]. Equilibrating [RWP03]. equitable [TS02]. equity [Huf02]. Equivalence [SP03, Rou04, RS08]. equivalences [GET06]. ER-diagrams [TWS05, TWS06]. era [BB03b, Del00, Mos03]. Ergo [Aha00, Ast09a]. EROSI [Geo00]. error [Chu01, ESS^{+}07, Edw04, LC01, NPM08, Qua07, Sma07, SB07b]. Errors [HMRM03, ABH^{+}07, AL07b, Chu03, Geo00, KYC07, Lom07, MKKC08, PWH00, Sar08, TL01, YK07]. essays [MMK02]. essential [GA06, Gup06]. established [SS00]. establishment [She04a]. estimates [PB03b]. Estimating [KN09]. Estrin [EN05]. etc [CM05b]. Ethical [Del00, Poh05, ADM^{+}09, Fle06, GY00a, GY00b, MK06, Wal05d]. Ethics [BHM06a, Hep06, SSA08, App06, Bot05, Bow00, Bri09, CG05, CL08a, Eps06, GC05, Got03, Got09a, GGKT04, LB00, Mar01a, Mar07a, Mar07b, Mar08, Mar09a, MMK02, PBM07, Qui06, RMPD09, San05]. ethnography [WS02]. eTutorials [RBG02]. ETV [Ter05]. Euler [Par03b]. Europe [Kla05a].

European [Cun07, MTBS09, OCM01, Por09]. evaluate [CJ01b]. Evaluating [BCB^{+}08, Cha02, CB00, CMD08, CB04, DLHAK08, ER09, HEM09, Lax02, MLC04, NCK^{+}03, WL00, Car05b, CBP03, Cra07, FKO07, Kum09b, MMD^{+}09, Ran07b, RCE^{+}00]. Evaluation [BHY09, DLPQ04, Xiz01, AKT07, BCM05, Bel02, BHN^{+}09, CH09, CFH^{+}01, DSSP06, EB09, Gri03, Kum05d, LAKE07, Mat06, MSTS02, NvdH09, NW00, Pap03c, PTT06, SHW^{+}09, Soj03c, Tho03b, VT01]. Evaluations [TY06, KNT02]. evaluative [ES05a]. EVEGA [KH01]. even [BGL^{+}07]. Evening [DHWH03]. Event [BDM01b, CC02, Soo01a]. Event-driven [BDM01b, CC02]. Events [HF04, Law01, AMW06, Poh05]. every [AB08]. everyone [LC08, ZF04]. Everything [GGH03, PD06]. evidence [DJS08, LFT09]. evil [Cle06a, HLL^{+}08]. Evolution [Lar05, TYT^{+}09, DK06, ESW09, Luq05, RR03, YD04, YA05]. Evolutionary [dMBM08, JA03]. Evolving [RD08, ADM^{+}09]. exact [MG04]. exam [BG08b, BDMY05, CG02, HHRA06, HKN^{+}02, DHRH05]. examination [BDW^{+}05, Cle09a, Eng02, LL04, LCS^{+}09, MC04, MR01]. Examinations [Tho03a, Tho03b]. example [AAD^{+}07, BCB^{+}08, BHN^{+}09, DWT00, GH01, God06, Han04]. example-based [DWT00]. Examples [Bur03, Ros02, ACD07, ALH07, BLN06, Fek05, Gin00a, dOG07, LCL08, RJAH04, Ste09b]. exams [BED^{+}03, Gha05b, Jac00b]. Excel [Hil04, Fon01, LS06, Lor00]. excellence [Pop07a]. exceptional [Bec03a]. exchange [DUM^{+}05]. excursion [BSV04, Min08]. execute [Cha03b]. execution [AYWM08, Ext00, LD01, Ort03, Rob07a, War01, Yeh03]. exercise [BVPE06, Bla09c, KMMS02, Rol06a, WW04, Wal05a]. Exercises [Kop07, BH07, CHMG01, EMM00, Erl06, GM09, GRLDPG^{+}01, HLNT01, RH08b, SMK01, YH03]. EXercita [GRLDPG^{+}01]. exist [BEM^{+}07, GG06a]. existing [MWF^{+}02]. exon [FMR^{+}06]. exotic [GS05a]. Expanding [Rob03b, ABB03a, Vat00]. expansion [BCLW08]. expectancy [Jen01a]. expectations [HY09]. expected [SM00b]. Experience [ES00, Eng02, LS04b, Old08, PRK05, ABE^{+}09a, BvdH09, BGD03b, Bar05, BUCP02, BTGD00, Cas09, CL02, CLP06, CN07, Cun05, DFS03, DLM09, FM02, GY01a, GCF^{+}01, GAL^{+}06, GK02, HM00, HAB^{+}00, HP07, JWC08, Lap02, LHS04, LS01, NWF^{+}03, NCA^{+}02, PHW07, Pol01, PMC^{+}04, Qid07, Ram03, RSK06, Reb05, Sch02a, SDLdQ05, Sta05, SM02c, Sur01, UD01, VR04, WWWG08, XSD07, YM06]. experienced [Pil09]. Experiences [ADP01, AAS^{+}04, BCF^{+}00, Dav09, DTZZ05, Doo07, GY00a, GY00b, KV08, MKS02, MPQB^{+}00, MDD06, MHK^{+}03b, NV05, RF00b, SM02a, Sha06, SHP^{+}06, Ste01, SMF09, SF04, Aba08, APR06, Coh07a, DBR^{+}08, DPBP09, Dei07, Dei08, DAHA09, DJWD05, DKN08, FSAF03, GCW^{+}03, Ham06, HHB05, HEM09, wH07, KCF^{+}00, KDP06, KI08, Lax02, LOP^{+}07, MF07, OJJ00, PQEA^{+}04, Rah08, SAHS04, SS00, TQCPQE08, WG07, WI09, WWGT07, ZIBG05]. Experiencing [SSR07]. experiential [BP09a]. Experiment [Bem02, Che00, CLK00, ES05a, HS04, Lax01, Mat06, MAFLR06, OCM01, PY05, SJA04]. Experiment-based [Bem02]. Experimental [EB09, BCM05, BC08, BGNM04, Bor04, BCF^{+}00, Bry08, Tre02, WTTT^{+}08]. Experimentation [Hun05, Rob00b, Wal04a, DEJ^{+}06, DJLA07, FNW02, Rob01a, Rol03b, SHW^{+}09, VIPC09]. Experimenting [MHW03]. Experiments [BNSW05, GT00, JKS00, TWS05, Bry09, HCMM00, SAM09]. expert [CL04, DDL03, LRPQ04]. expertise [CHK^{+}01, CWP09, NRDP07]. experts [KV00, LRPQ04]. explain [Fon01]. explaining [LFT09]. explanation [PP07, SK09]. explanatory [BL06b]. explanograms [PPE03]. explicit [HAG05]. explicitly [CR03, Kop01]. exploit [TE08]. Exploiting [DeP06, ZJ03, PT03]. Exploration [Rob02, CNP^{+}09, DK01, EG08, Gin00a, Hut04]. Explorations [Lis01a]. explorative [Ker04, OGMA07]. Exploratorium [BSL^{+}08]. exploratory [Mar07c, MH08]. ExploreCSEd [SBM05]. explorer [CPJ07]. Exploring [AL09, BMK09, CFT08, CPT09a, CHUB08, DUM^{+}05, God06, HNC08, LDK^{+}07, NRA^{+}03, PGK05, PT03, Ras01, RSHL07, AEPQ09, BMK03, CKMP09, Haw09a, KH08, Rol09a, Rol09b, SBM05, Tur03, Wol02]. export [AGHM02]. expose [PB09a]. Exposure [ENO09, BS00]. expressed [SBE^{+}08]. expression [BH07, Kum05d]. expressions [AOMC07]. extended [PRK05, Ste04]. Extending [DDS08, GGR03, HSB09, KvdHR05a, RK09b, TLS07, WC04, BSA^{+}05]. extends [PA06]. extensibility [SD09]. extensible [HCB04, KPT00, RF01, SD08]. extension [PH04b, PHN06]. extent [CMD08]. External [Jam05]. extra [Joy01]. extraction [POMA08]. Extreme [LR05, BCD^{+}04, Car02a, HR03, MPK08b]. eye [NTTHT09]. eyes [Har05].

Fabrication [RB09]. face [Got06b, Har01b]. facets [BHYA05]. facilitate [AHJ^{+}07, CS08, PP03a, PKB06]. Facilitated [Cra07]. facilitates [LNR05]. Facilitating [FAL^{+}05]. facility [Rob00c]. fact [SKM02]. factor [Rad09]. Factors [EH05, GR00a, VS07a, BMK09, BR05a, HALW01, LML07, Ni09, RRRH04, SD03b, WS01b]. Faculty [CSC03, Sal03a, Sal03b, BGSL06, Car05b, CLP^{+}01, DHGR04, FKO07, KCF^{+}00, LDA^{+}00, LB00, Lee01b, Lew07a, MR00, Per09, Sam05, SPW^{+}08]. FAdo [MR05]. fail [Edm07b]. failed [ABK^{+}05]. Failure [BC07c, Cle07a, Got06a, RRR02, RRRH04]. Failures [KT07]. Fair [Bec08]. false [Reg08]. familiar [Fon01]. familiarization [FS09]. family [DKN08, DMKN02, GMJDA09]. family-based [DKN08]. fashion [LNCC09]. Fast [GMGM06, ABL07, GS06b, KvLA09, RKS07]. fault [MAFO08, SD03c]. favour [Mas05]. feature [ELC07]. feature-based [ELC07]. features [Qua07, Ric00, SSMN04]. February [ACM01, Ano03, DNTB05, GWG02, GKJD03, Imp02b]. feedback [ALR08, AKA09, Bib06, CBH09, Fon02, FPQ^{+}08, Gri01, HD04, KLOT05, LHS09, OHZ01, PBDWL09, TIKT04, Wal06b]. Fellowship [Joh02]. fellowships [SLPG07]. Female [Mos02, BAAM06, Fra02, Haz05, Kum08a, RIP^{+}09, SM02a]. few [Ort03]. FGA [BL05b]. Fibonacci [Rol03b, RSHL07]. field [Gib00, Goo08, PRC06, PRW04, SCE07, TMS07]. field-dependence [PRW04]. field-dependent [Gib00]. field-dependent/independent [Gib00]. fields [Cle05a]. Fifteen [FM06, SCHS09]. fifth [JKDN04, MSS^{+}04]. fifth-year [MSS^{+}04]. fighting [Poh05]. file [BP00, Sha04]. Filling [CNCL09]. Final [Mar06, BT06, McC06a, McC06b, OBLH03, RBC^{+}06]. finally [McC02]. finding [RJAH04]. Findings [Ran07a]. Fine [KA09, BGDR05]. Finite [SW06a, Boz06, Gri02c, Gri03, WW06a]. First [DAHA09, JT04, AWS^{+}09, BS08, Bel02, BC04a, BKM^{+}01, BW05, BLOM06, BCS^{+}07, Bla09c, BDM01a, BZ04b, CS02, CDWH00, Chr05a, CDP03, Cun00, DGMW04, DR07, DAKLH07, DLHAK08, DLM09, Edm07a, Edm09, EGMS08, Fle07, Fos05, Gol04a, GL08a, GL09, GR00a, Gri08, HAB^{+}00, HDH06a, Hen07b, HV07, HRS06, Hov05, HTW04, How05, HBCW03, HD07, KR00, Las09, LE07, Lis01a, LS08c, LKPC04, MGS07, MS05a, MAD^{+}01a, MAD^{+}01b, MB05, MWPB05, Mur01, Mur07b, Noy02, PPP06, PSW03, Pil04, Pow08, Rad06, Ras00, RSK06, RRKS06, RM01, Rio02, Ros06, Rou02, Sae09, SM00a, SL08a, SB09, Sei09, Sha03, SDM^{+}02, SHVP05, SLSM02, STS09, Ste01, Sur02, TRT04, VW00a, VI01, WB06, Sna06a]. first-time [Lis01a]. first-year [Edm07a, Edm09, GR00a, HV07, MGS07, MAD^{+}01a, MAD^{+}01b, Pow08, Rio02, Sur02]. fit [FHBD06, HB09b]. Five [BH03, Lis03a]. fixed [MT08]. flash [PB05b]. flat [AKWZ07, FIG07]. Flavius [Hen06a]. flavor [BS06a]. flex [Kag09]. flexibility [MMC01, ZRY05]. Flexible [KH06, Lea01, AFR08, KvD02, KLOT05, Lar05, LM06, RKS07]. Flo [Ber03a]. Flo-and-Mac [Ber03a]. flocking [Gre00b]. Florida [GCW^{+}03]. FLOW [PHMS08]. flowcharts [Bou03, CM05b]. Fluency [AKM00, DDC^{+}03b, WAS^{+}02]. fly [BA04b, BA05a, Esp08, Gin09d, RL09]. focus [CG05, DKR09, KvdHR05b]. foes [Gin08c]. foil [YM06]. font’ [Dic07a]. Food [DR07]. Food-first [DR07]. foolish [Rol08a]. force [RCE^{+}00, RBC^{+}05, RBC^{+}06]. forensics [GLD^{+}05]. forest [LST^{+}06, Wal00a]. forever [RH07]. forgetting [Haw07]. Forging [Sud08]. fork [Rob02]. form [Cla05, JJOA07, Rut01]. Formal [AOMC07, DG09, Mai09, ADG^{+}01, BLA^{+}09e, CCY03, CGM04, Eng02, HW06, LTHN09, Pil09, Rod06, RLR07, Sig07, SF02, Sob00, TN02, WD05, Zin08]. format [Coh07b, EBC^{+}08, San05]. Formation [CF02, Bow07b, Dei05, SR07]. Formative [HB06, MSTS02, Sol07a]. formed [SSS07]. former [GG06a]. Forming [Bec07]. forms [AOMC07]. formula [Rol03b]. Forthcoming [Cam05b]. forum [Reg02b, Wal02a]. forums [Dia05, SMR03]. forward [SNH08]. Foss [MdL09]. foster [WB09]. Fostering [How03, LJG05, Mic00, YOB08, HL07]. Foundation [Joh02, BT01, BBT07, CMT04a, CMT04b, CGT05, Cun05, Hum04, PMT03, Sta05, BP02, MR00]. Foundational [Boe06]. foundations [TBS^{+}00]. Four [Kur08a, Dei08, HGEB08, Lis06b]. Fourth [Ano03, Fro07]. FPGAs [JHLT06]. frame [CTBS06]. Frames [Hab02]. Framework [Hun05, MvW01, Wic03, ABL07, CTLW03, DW07, Etl06, Gri00, Gui05, Haz03b, HCB04, Hu03, Kag09, Kol08a, LRD09, MMP05, POM07, Phe07, RC04a, SBZ09, Ten03, Way05]. Frameworks [CC02, Chr04, NP02b, PV08, VA02]. Free [McC00b, MTBS09, LC01]. Freedom [Rad09]. Frenzy [CCM07]. frequently [VT01]. fresh [LCG07]. freshman [GD07]. freshmen [Dew05, Osb06, TT00, VI05]. friendly [Lar05, MWG08]. frontier [Haw09a, LA08]. frustration [HE07]. Fulbright [FSAF03]. full [LL06a, Rou06]. Fully [CN07, SMK01]. Fun [SK05, Cur07, GCD^{+}09, HLCS08, MRGS08]. function [RSVI09]. Functional [TV08, AMW06, Cox05, Man09, MG04, PL04a, Soo00, VW00a]. functions [Man05, Pra05b, Rou04]. Fundamental [VZGE07]. fundamentals [BCE06, CKM07, JDFV08, Lud05, NL04, NS02, SM00a, SC07, Shu06b]. fundamentals-based [SM00a]. funded [Cun05, PBWS06]. Funding [BBT07, BT01, BP02, CIL^{+}03, CFT08, CPT09a, CMT04a, CGT05, MR00, PMT03, TP03]. funnel [PN08]. Further [GM06b, LFT09, BT06, McC06c, McC06d, Rol09b, Yai08]. futile [Gin07d]. Future [LB00, Sut00, AWS^{+}09, AHPS09, BCS02, Bor02, CAK09, Cas08, DG09, FS09, Haw07, Kum09c, Laz08, Lee01b, LR07, Man04, MUPVPF^{+}02, MGFK06, OAB07, TS08, TW09b]. fuzzy [WZ00].

Game [CKLW05, KH05b, RM08b, Ros02, SPW^{+}08, ABAE^{+}09, BL06a, BSA^{+}05, Bri09, CD08, Cha06a, CC05, DW07, EB08, EB09, FNW02, GGH03, Gin09a, GD07, Gou08, Gou09, Han04, Hua03, KF03, KAPG07, KR06, Kur09a, Kur09b, LV08, LR04, LM06, Lim07, LS08a, McG09, NL04, NvdH04, NL08, PRK05, PKR06, Pra05a, RI08, RMY08, Sei09, Soh04, Wic05, WP07, BGNM04, Sco03]. game-based [NL04]. game-building [KAPG07]. game-like [Gou09]. Game-themed [SPW^{+}08]. Game2Learn [BRP^{+}07]. Gameboy [KK09]. Games [BS06a, CM08a, HLL^{+}08, LH02, RM08b, BRP^{+}07, Bay09, BSA^{+}05, BGL^{+}07, Car08, Fro08, Ges07, Gig03, Gin02b, Gin07c, GMJDA09, Got08, HRBC03, Jon00, KSE02, LE07, LLP^{+}09, LRD09, NW02, Raj05, Ree00, Sch07, SHS^{+}08, SSR07, SdSK05, SK05, Wal03a, WKC^{+}09, WW09, WBBC09]. games-based [Raj05]. gaming [MP09b, WBPW06]. gap [BGD03b, Bar05, BHLW02, CNCL09, GG06a, Kre09, Wic07, dBGdMB09]. gaps [HL04b]. Garden [MSK09]. gates [War04]. gauntlet [DMKS01]. geared [BHRB08]. GeekOS [HHB04]. Gender [BRP^{+}03, CJ01a, CJ02, CSD02, GG06a, HV07, Huf02, MGS07, Spe02, ABB03a, BDW^{+}05, CV08, GESZ09, IK06, Ira04, LZL08, MNBK03, MRM^{+}06, RPH^{+}03, SM02b, TS02, VZ06, VHF07]. gender-equitable [TS02]. Gendered [Coh07a, CPT02]. gene [FMR^{+}06]. General [MFL01, GM08, Gor03a, Les04]. Generalised [TWS09]. generalization [RSVI09]. Generals [TBAK04]. generate [AAF^{+}06]. generated [CMS06b, DLRHP09, GM09, KS00]. generating [BGKT00b, Kum00, TG05].

Generation [Ker04, CSA^{+}07, Cli04, KLOT05, MP09a, RKS07, SPAC^{+}05, SC08, pYB08]. generator [For04a]. generic [AFR08, BGNM04, GGÁC03, RCF00]. genetic [BDMY05, Lan01, PB03b, Pil04, WP02]. genomic [LD04, MHW07]. genre [FU02]. gentle [RSUFPF08]. geo [FCA05]. geo-referenced [FCA05]. Geographic [ZB08]. geographically [HWH^{+}05]. geographically-distributed [HWH^{+}05]. geometry [And00, LS01, SL00a]. Georgia [BBE^{+}09]. gestalts [MRH09]. Get [BA00, BZ04a, BGL^{+}07, Kac09a]. Getting [Bib06, Cla00, Sur01, HBCW03, MEM^{+}07]. giants [LSG05]. gimmick [Gol02]. GINI [MMN^{+}09]. girl [Oli05, PMC^{+}04]. Girls [Car08, Kla02b, WBBC09, CFKS02, CH09, DLM07, GL03a, SD03c, SB06]. Give [HBH07, Har01a]. GK [SLPG07]. GK-12 [SLPG07]. glass [DVE06, PEH07]. Global [Cle08a, CK08, DBK^{+}09, Pea02, PTT06, PTT08, RVP08]. Globalization [Gab03]. globally [CB04]. glut [Lis08b]. go [CWL08, Dic07b, Hua03, Lee02c, dR04]. goal [MH05]. goes [AGGR04, GLR^{+}01]. Going [SCL^{+}08]. GoJava [Wis06]. gone [vdK07]. Good [BA00, BMT^{+}07, BMT^{+}08, BMD^{+}09, Got09a, HLL^{+}08, MD05, MLM^{+}08, Rol01]. good/evil [HLL^{+}08]. Google [LW07, May08]. Google^{T M} [Cle06a]. goVista [CKMW02]. GPU [BC07b]. Gr8 [CH09]. Grace [TBMC08]. grade [AR08, Fro07, Haz03c, HKN^{+}02, Lor00]. grader [Hil04, HSD05]. graders [Jac01, KB08]. Grading [Bec03b, MMK02, Wal09b, ALR08, AK09, BG08b, BGKT00a, ES05a, EPQ08, Ell06a, FPQ^{+}08, Geh01, Hel07b, KLOT05, LCB^{+}03, LW06, Noo06, PL04b, ST07, Wal00b]. gradual [VI00]. Graduate [AHM^{+}04, Gor06a, Imp09, Joh02, BZ04a, BP09a, Coh07a, CA02, DBR^{+}08, GV09, Gor00b, GGS^{+}06b, HE06, JKS00, Jen08, JY01a, KT01, LdLM^{+}08, Sha06, SD03b, SW04]. graduated [BS00]. graduates [BS08, RP06, Tha07, YPdMA^{+}08]. graduating [BBY08, EMM^{+}06b]. Grand [Lis05a]. Graph [Gin06a, BP09b, KH01, Lim07, LNR03, STTBC07, She03b, Sna06b, Wu05]. Graphical [Cun02, LM06, Sco03, BBE06, BN09, CNCL09, DSCU01, GL05, GH06, Köl08b, VS07b]. Graphics [Par04b, Par04a, Sch00, ACSS06, BDM01a, Cun00, Cun02, CS03b, Cun07, GL08a, GL09, GMGM06, Gou06, LM05, LS06, MD06a, MD08, NS02, Nod03, NH05, Pra05b, PC08, SL00b, STS09, SSR07, TF07a]. graphics-based [BDM01a, MD08]. GraphicsMentor [NS02]. graphing [LM05, LS06]. GraphPad [BP09b]. graphs [CMS06c]. graying [Got06b]. Great [Den04, DDRG04, FG08, GP00]. Greece [DSF^{+}03, Fin03]. greedy [Gin03b, VIPC09]. green [TW09a]. Greenfoot [Hen06b, Köl08b]. grew [PN08]. Grid [MAFW08, HWH^{+}05, WF08]. grief [Lis09a]. ground [LJM05]. Group [FMKQ08, JP03, Joy02, Bar05, Bec07, CCM07, CD01b, Cle02b, Dei05, DWJ02, JY01b, LS02, MH00, OCLL00, RF00a, SS06, TM04, WJDL04]. grouping [OCLL00]. Groups [CF02, Hob00, HRB^{+}09, Kum09b, MS05b, Red01, SB06]. groupware [RBG02, Tow09]. Groupwork [Bow07a]. Growing [BK03]. Growlroarare [HHRA06]. growth [How03, KV00, Lap07, Rou04]. GSPIM [BBE06]. GUI [BH04, BK03, Eng01, Eng04, Hei03, KW01, PB05a, PB05b, SLC^{+}09, TETPQ08]. GUI-like [KW01]. guidance [Ekm06, HSB09]. guide [Imp06a, MDD06]. Guided [PBDWL09, PCW09]. Guidelines [Kle02, KR01, MH05, BS02, GGS^{+}06b, HKCW08, Haw08b, HKCW09, PG07, Rou02, San09a]. guides [McC03b]. Guiding [Ros02, CGB08, Had05]. GUIs [HF04, Les03, PRR02, Röß06].

hack [Ast09a, LC05]. hackday [RHC07]. Hacking [Cla05, LW07]. Hadoop [Bro09]. half [BSV04]. half-adder [BSV04]. halls [KSE02]. Hamming [Chu01]. hand [NTTHT09]. hand-eye [NTTHT09]. handheld [All03, JKDS01, RW02, Ric02]. Hands [PF03, BC07b, BAMSK05, Bry09, Ens09, QLT09, RBFR06, SHW^{+}09, SM05a, SM05b, VMD01]. Hands-on [PF03, BC07b, BAMSK05, Bry09, Ens09, QLT09, RBFR06, SHW^{+}09, SM05a, SM05b]. Handwritten [Gol04b]. handy [Imb03]. Hanoi [Min07, Min08, Min09]. happenings [Pop08a, Pop09a, Tur08b]. Hard [CLR09, LBA07]. hardware [CWM05, CFKS02, Dic00, GDM05, JMS02, SC04, SC07, SNA09]. Hardy’ [AF03a]. harmful [BS00, GEVZ09a, Got00a, HGH^{+}01, Mas03]. Hashing [Ham02a]. Hasty [Gin07d]. Having [Cut07]. HCI [BP09a, Cha02, Cox05, FC00, KvD06, KvdMvDvdV00, Lif04a, Lif04b, MCS04, Pas05, PA06, Vat01]. Head [Ast07b, Rob04a]. health [SMF09]. Hear [McC02]. Hearing [CLR09]. heart [Bri09]. held [MFRW07, MFR^{+}08, MFR^{+}09]. Helipad [Gin04d]. help [Cha04a, Ekm06, HM00, HSM06, Jin08, NPM08, OHZ01, Rut01, SSA08, Wal06b, Wic09]. helpful [MMK02, VVV04]. Helping [CGB08, DL06, MKLH09]. here [DPBP09, MJ03]. heuristic [GA06, Gin07c]. Hidden [Gin06b, BS09]. Hide [WSS05]. hierarchies [PG06]. High [BBC06, CCH^{+}03c, HL06, Oli05, fCST02, ABAE^{+}09, AGE04, BM05, BAB^{+}08, BC07d, BCLW08, BEE07, Cas09, CNCL09, Chr05b, CTBS06, EH05, EH07, GESZ09, GEHR09, Gin00a, HGEB08, LV08, MVD01, Mar08, Mar09a, PMC^{+}04, PHZ^{+}03, PC08, RH03, RRKS06, RHY07, RIP^{+}09, SBZ09, TH06, Wos06, YM06]. high-school [Gin00a, RH03]. High-tech [CCH^{+}03c, PHZ^{+}03]. higher [ABS08, GLD^{+}05, MvW01, Sal03a, Sal03b, SRGF08]. highlighter [BC09]. Highlighting [Phi04]. Highly [HH07, Köl08b]. highway [Alv05]. Hilbert [Ras01]. Hillsborough [CRPZ06]. him [CGB08]. hiring [CV08, Wal01a]. historic [Lee00a]. Historical [PRL09, DMI^{+}09, Kum08b, MWM01, PPRV06, Whi09]. historically [WTTT^{+}08]. History [Lee00b, Lee01a, ABO^{+}05, AAO^{+}08, AAEI08, BAK^{+}09, Cam07, CM06b, DKR09, GM06a, GM08, Gür02b, HB01, HILS05, Hem05, Imp02a, IL04, Imp05c, Lee02d, Lee02c, OAB07, Wil00, Wil03, Kum09d, ASA08a, ASA08b]. hockey [BM07a]. Home [See08, Bro09, HHB05, MDD06]. homebrew [KK09]. honesty [Wal04b]. Honeynet [Tal06]. Hong [WD03]. Honor [Pop05b, Pop02]. honors [Sor05]. Hopper [TBMC08]. hops [Gin08a]. horse [SMH05]. hours [Mal09]. Houston [ACM06, BTHR06, MD04]. HTML [MDS04, ZJ03]. HTTPExplorer [Tur03]. Huffman [BP00, Wic03]. human [BBMR03, Haz01b, LLB02, McC03a, MR09, Rya01, WS02, WP06b, DTL^{+}02, Shu06a]. human-computer [BBMR03, McC03a, Rya01, WS02, DTL^{+}02]. human/computer [LLB02]. humanistic [LR01]. humanitarian [EMdL^{+}07]. humanities [Lin05]. humanity [HEM09]. Humans [EA07, SPAC^{+}05]. hunt [TPK^{+}06]. hunting [GY02]. Hybrid [ESW09, Ros01, CW03, Min09, Qua06, SJS^{+}07]. hybrids [PB03b]. hypermedia [AD01, MMC01]. hypertext [Tur03]. Hypertextbook [Ros08]. hypertextbooks [BGGR01, RNH^{+}06, Ros08].

I/O [Gri00, Rob06a]. Iambic [Sie00]. IAPPGA [Wu05]. IBBA [WD03]. IBM [Sie00]. Ice [BM07a, BC04b]. ICE/TTM [BC04b]. icebreaker [CD01b]. ICETM [BC04b]. Iconic [CM05b, CCG05]. ICPL [Har02]. ICT [BPS09, CB04, FS09, ICL02, KV00, LML07, LMS00, SNS07]. IDE [CH06, FMKQ08, HCB04, Kac09b, Köl08b, PH03, PH04b, PHBM05, PHN06, RC04b, WKB02, Xu07]. IDE-based [Xu07]. idea [BMT^{+}07, BMT^{+}08, BMD^{+}09, BBS06, ABL07]. ideal [Han04, Hua03]. Ideas [Gha07, BLOM06, Haz03a, Haz08, Our04, SH07, WKB02]. IDEFIX [GGÁC03]. Identifying [GGH^{+}08, HMRM03, Hui01, PHM^{+}01, TRT06, KB09, MGH^{+}06]. IDEs [MHK^{+}03b, OPS^{+}02]. idol [FCKU09]. IEEE [IHL^{+}04, SL03]. IEEE-CS [IHL^{+}04, MMLT09, RCE^{+}00, SL03]. IEEE-CS/ACM [IHL^{+}04, SL03]. If [FLS04b, MBD04, KMMS02]. IFIP [MvW01, NP02a, Tur06, Tur07a, Tur07b, Tur08a, Tur08b, Tur09]. IFIP/UNESCO [MvW01]. ignorance [Lis03a]. II [WZ00, BL06b, CHMG01, Dal00, McC05a, MR01, SJ06, Wal02b, Whi03]. III [WP07]. ILE [FCKU09, HKF00]. ILE-idol [FCKU09]. Illinois [CCM08]. illuminating [Gin02a]. illustrate [AYWM08, MDGM05]. Illustrating [RW02, Ric02, SC04]. illustrations [GGAB01]. Image [Bur03, Can06, AD08, EG08, Hun03, KH00, Kla05a, MV05, Wal07b, WN05]. Imagery [RTH^{+}08, Lew07b]. images [Fri05]. imaginary [Ada07]. IMC [LR01]. immersive [DBR^{+}08]. immigrant [DKN08]. Impact [Sok00, ADM^{+}09, BR05b, BPW^{+}09, CP00, FG08, Fra02, HFHK05, LB00, MN01, NCK^{+}03, SM08, Wic07, Wol07]. impacts [LAKE07]. impaired [AM07, CGB08, CMS06c]. impairments [LR08, NL08]. imperative [Lau07, Ras00, Sav05]. implement [SM05b]. Implementation [Don08, DHRH05, TMD^{+}04, Wan01, AC07, AL01, BMW04, BL05a, BC04b, CCK01, Hu05, Jon00, Kaw07, Mar05b, Mar06, NLCvL09, OG05, Ort08, PH07, Pow03, Sch09b, She04b, SBZ09, WJJD06]. implementations [Mus06]. implemented [HHRA06, Nef04]. Implementing [CP00, PSW03, Soh06a, SP03, WKB02, RJM^{+}05, RCD^{+}02, SSP^{+}05c, WW04]. Implications [Cap02, Chr05a, DJC09, ABA09, CSD02, Hol05, MT08, PCW09]. Implicit [Pra05b]. Import [AGHM02]. importance [BC07a, CCMS05, CR03, Tha07]. important [ALH07, GGH^{+}08, GS03, Hal08]. improve [Bar01, BZ04a, BYS09, CLM^{+}07, CS06, HCMM00, MFR^{+}08]. improved [EB09]. improvement [BS06b, Cla05, Hui01, LDHW02, OCM01, WLCL08]. Improving [ALR08, DBR^{+}08, DD08, EGB07, Fon02, GP07, GY01a, GHLL08, Kur08b, MFR^{+}09, NWF^{+}03, Pow08, BBE^{+}09, EGB05, JKM07, Luq05, PW01, WK07]. in-class [Dei05, DKL^{+}07]. In-person [ES05a]. in-process [KB09]. in-service [KP04]. Inclusion [KW07, Lew02]. Inclusive [DW07, Han07, Dei07, IMRJ09, Kan07, LR08, MMRI07, MICM08]. inclusivity [SNS07]. Incorporating [D’A03, DTL^{+}02, GD05, KKLK04, LJ08, MS03, Soh06b, DO07, Sob00, Sta05]. incorporation [CG05]. incorrect [Cle04b]. increase [CL02, HRB^{+}09, WC04]. Increasing [Dal02, Goo08, HR00, Kla05b, KF03, PSH^{+}07, PTC^{+}01, PMC^{+}04, RLR07, RWL^{+}09, WAS^{+}02, Dei05, Par06b, WGS05]. incredible [Cam02a, RPR03]. Incremental [RMT09, Jan01]. Independent [Bow07b, CCH03a, FG08, Gib00]. indexed [DAH08]. Indiana [Hen03b]. indicator [BC06a]. individual [BEW08, HD04, LCPQR05, Rob00e, YOB08]. individuals [Bur08, WL00]. Inducing [TBMB09]. induction [PPA08]. Inductive [Hen03a, Hen03c]. Industrial [PRC06, RC02]. Industry [Hit05, HMS08, PRK05, PM02b, Reg02b, RRD^{+}09, San08, UCJ^{+}04, YH05]. industry-driven [PRK05]. inexpensive [NTTHT09, Xu08]. inexperienced [NLMN06, PJ01]. infant [SA07]. Influence [FS09, BMK09, BR05a, LML07, LBA08, MGH07, PHW07, SNS07]. Influences [SD03b]. influencing [EH05, Ni09]. informal [Gri01]. informatic [Bla09b]. Informatics [KvdHR05b, BF06, CCD^{+}08a, DH06, Fle06, God06, KvdHR05a, SB09, WB09, vVML04, MvW01]. Information [Cam06b, CDGS01, DDC^{+}03b, FLL^{+}01, Gor02, Gor03b, Gor04b, Gor05, Gra03, Gup01, JCG06, LLP04, Spo00, Wei05b, AE01, AKM00, ASPS04, Alv05, AT01, BCS02, BA08, BAMSK05, Bry09, CIK^{+}03, CMS06a, CNP^{+}09, CSD02, Dic07a, DS08, Ekm06, Eps06, FCA05, Gha07, GLF01, Gor04a, Gor06b, GGS^{+}06b, Ham07, HKCW08, Lea01, LFL06, LKGY00, LGA^{+}01, LC05, MA01, MM00, ND08, PG07, Pop02, Rob00d, Spa04, Top08b, Top09a, Tut00, VKV01, VKV02, Wal05a, WB02, YSLL00, YLYX06, ZB08, LFGY02]. informed [Hol08]. informing [CBKM07]. InfoTraffic [Arn07, ALH07]. infrastructure [DWJ02, LM06, TLW^{+}04, TLW^{+}09]. Infusing [cLwW07, Rao05]. inhabited [SPAC^{+}05]. inheritance [Par09, WSP02]. initial [Ber01, CCC07, CP00, HY09, Har02, Utt06]. Initiative [BPT^{+}04, BPW^{+}09, CIL^{+}03, CMD^{+}07, MTBS09, PBWS06, PB06, SSS^{+}09, Wal02a]. initiatives [Lee00b, Lee01a]. Injectable [Got03]. Injecting [LHS09]. injuries [BS09]. ink [PB09a]. inking [Wol07]. Innovating [AD08, Tha07]. Innovation [ALRY08, BCK04, CJGU02, CFPL05, DSF^{+}03, DGS06, FKCM01, Fin01, Fin02, Fin03, HPT07, May08, YM02, Kac08a, Ram02, Röß08, Shu06b, TFG05, To04]. innovations [Ni09]. Innovative [HY01, BHBG07, BMP06, DSBK01, LMS00, Mik05, Ram03, SSS06]. input [MDGM05]. input-output [MDGM05]. inquiry [BPK^{+}01, FM09, Ham04a]. inquiry-based [BPK^{+}01]. inroads [Owe08]. ins [NBL09]. insight [LW05]. insights [BW06]. INSPIRED [DLM09]. Inspiring [BAB^{+}08, CMCB09]. instant [BAB^{+}08, FPQ^{+}08]. instantiation [SFD08]. instead [HK01]. institute [Mar01b]. institution [GE06]. institutional [BGD03a, HHB05, KDH^{+}00, MAD^{+}01a, MAD^{+}01b, MMW^{+}05]. institutions [ZIBG05]. instructed [Bow08a]. instructed-teacher [Bow08a]. instruction [ADP01, AAD^{+}07, AWS^{+}09, CTP09, CO00, CLM^{+}07, CNCL09, Hol08, Kla04, LAP08, MHA04, MGH07, NTTHT09, PDS07, Sco02, SDLdQ05, Sta06, WW00]. Instructional [OCLL00, BCB09, CSA^{+}07, DSSP06, DTW07, HLS02, Hui01, HNC08, LR01, MLC04, PRB09, RWP03, SSC01, SSP^{+}05b, SC08]. instructions [Dic07b]. instructive [Gin04c]. instructivist [ML04]. instructor [BPW^{+}09, Kur08a, MSP^{+}06, VA02]. instructors [CTR^{+}03, CGY^{+}01, Kay00, MKLH09, SHS^{+}08, WP06a]. instrument [LCS^{+}09]. Instrumentation [TAHA00]. instruments [OCLL00]. integral [RK09a]. integrate [Gol02, JHLT06, WW06b]. Integrated [Hel07a, BHR07, BDM^{+}07, BSL^{+}08, CGI05, NM02, RD08, Rio02, TAHA00, Udo06, WP07]. Integrating [Gra07, HWRC08, HAFT09, Imp02a, Jon01, KD07, KPT03, LGA^{+}01, Pas05, Pol01, RBC^{+}02, RKD02, RS03, SF02, SS00, SM07, TN02, WM03a, Wal05d, AHJ^{+}07, BUCP02, CCG05, CN07, DJC09, FC00, HH05b, LJM05, MNN^{+}09, MK06, MWF^{+}02, Pho01, SS07, TA08, WW09]. integration [BP05, JS06, LSMB00, Put01]. integrity [Joy06b, RMPD09]. Intellectual [Kay00, WBGM02]. intelligence [BC07e, Bur06, HBH07, HA04, Imb03, Kla02a, MF07, Mer01, SS06]. intelligences [Bec03c]. intelligent [MR09, MWPB05, Pil03, SPAC^{+}05, Soh06b, TG05]. intensive [Pol01]. intent [To04]. Intention [LV05]. Intention-based [LV05]. Inter [KFE^{+}07, Dew05, KDH^{+}00]. inter-institutional [KDH^{+}00]. inter-object [Dew05]. Inter-university [KFE^{+}07]. Interacting [RRRH04]. Interaction [HALW01, RBM04, And00, BHR07, BBMR03, CS00b, Dei05, GDB03, HGK09, HR00, Kum01, Lau06, McC03a, NRB^{+}05, Rob02, RLR07, TBMB09, WS02, DTL^{+}02, Shu06a]. interactions [DFGM08]. Interactive [BLN06, Han02, KH00, LS04b, MR05, SB07a, Soj03b, VIPCUF09, ABL07, Arn07, Bac03, BGKT00a, BH07, BP09a, CCH03a, CCM07, Cas09, CSA^{+}07, CFR04, CC01, Est01, HH07, HMS04, JFH00, KT01, KK03, LHS04, LLB02, LRD09, MSK09, MS05b, PBK04, Pho01, Rob06b, RNH^{+}06, STTBC07, Sei09, SS00, SC08, VIPC09, Ver05, VCJ09, McD02]. interactively [Rod06]. interactivity [KW01]. interdisciplinarity [Kac07b]. Interdisciplinary [MH00, ZLL^{+}07, BBN07, DH06, FJ06, HGK09, Kac08a, KvLA09, RF09, Spo00, SSW08, Wal05d, ZE01]. interest [AAF^{+}06, Dal02, GL03a, LJG05, Oli05, SM02b]. Interesting [GELS04a, Hem05, Hen02, SWR05]. interests [Djo08, RPG04]. Interface [Hel07b, Sco03, BL01a, CMS06c, DSCU01, GS05b, Guz01b, HTSW07, Hu06, KA09, LBC04, MMRI07, PP07, Rya01, SB07b, Tra07]. Interface-based [Hel07b]. interfaces [Ber09b, BCF^{+}00, GH06]. Interleaved [Gin09c]. intermediate [KYC07, PB05a, vTNC08]. internals [Sci07]. International [AHLY03, VKV01, VKV02, AL01, CCK01, FSAF03, FAL^{+}05, LDA^{+}00, LDHW02, Pop02, Pre07, TCK09]. Internationalization [Lee02b, Röß06]. Internet [ACd08, CD08, GT01, GL03b, HTBS01, HKW^{+}02, KPN02, LS08b, Mar01a, Mar08, Mat03, MTHS00, Par07a, Pul00, Sha00, She06, SMTH00, Tho03a]. Internet-based [Pul00, Sha00]. Internet-centric [HKW^{+}02]. Internets [MMN^{+}09]. internetworking [YH04]. interoperability [DG09]. interoperable [BdMG08]. interpolation [FLS04a]. interpreter [Fos05, MPD^{+}05, Ort08]. Interpreting [Han05b]. interprocess [Ree00]. intervention [OCLL00]. interventions [Kum09b]. interview [EN05]. interviews [Dic05, WPS01]. Intra [FK05]. Intra-curriculum [FK05]. intrinsic [JS06]. intro [BYS09]. introduce [Gig03, LS08c, Lud05, MAMC05, WS06]. introduced [Sic09]. Introducing [Arn07, BC09, BO01, Cas09, GELS04b, GL07a, Ham05, HRS06, KK09, Les04, Lif04b, Pap03d, PKR03, Rod02, Ros05, SD03c, Soo09, TMSZ07, VW00a, WTTT^{+}08, Wir08, BL01a, BSV04, CG02, CC02, DMKN02, Gri00, Han04, HLM01, Kru03, MS07, NR05, RW05, SD03a, Wal02a, ZJ03]. Introduction [BCL00, BKM^{+}01, Bou03, BZ04b, Cor07c, Gol04a, MR06, PG06, Rad06, RIP^{+}09, RSUFPF08, RS08, SMK09, Wos06]. Introductory [Cha06a, ES05b, HMRM03, LLP^{+}09, LL03, MDS04, MFL01, Rob04c, Sav05, AGGR04, ABAA09, ADP01, BMK09, Bar02a, BUCP02, BED^{+}03, Bay09, BC07c, Ber03b, Ber09b, BVPE06, Bla09a, BLA^{+}09e, BHN^{+}09, BDM^{+}07, Car03, CBP03, CB00, CM02, CO00, Cig03, CTR^{+}03, CFGL05, Coh07b, CDP03, EB09, ES05c, ET02, FP09, FKKW00, FG07, Gel00, Gib00, GGH^{+}08, HPC^{+}03, HPCZ04, His09, HRB^{+}09, HB09b, Jac08, KEK^{+}07, KH05b, LDB^{+}03b, LB08a, LE07, LRPQ04, LFT09, LNR05, MWBF02, MF07, MBK09, MdL09, MH05, NLCvL09, Osb02, PSM^{+}07, PQEA^{+}04, PSW03, PGC^{+}06, Pro00, Pro09, Rad07, Raj05, RMZG06, Reb05, Rob00e, Roi05, Ros09, RF00b, Sch04, ST08, SSP^{+}05c, Sol07b, Sta05, SSR07, TFG05, TRWJ02, TATL07, WS01b, WMP08]. introspective [GH06]. introverted [Tow09]. intRusion [TW05]. intuition [Gin01d]. intuitive [ABE^{+}09a]. inventories [AHH^{+}06, Rut01]. inverted [PM06]. Investigating [JKKL04, MFRW07, MAFLR05, Pre07]. investigation [HTSW07, MMW^{+}05, PJ05, RMB00]. investigative [Pap03d]. invigorate [JG08]. Invited [Tur01]. invoke [Kac08a]. involvement [Ell06b, Wal02a]. involving [CTP^{+}04]. IO [PR04]. IP [FMJM02]. iPodLinux [LB08a]. IPP [Est01]. IREEL [DEJ^{+}06, DJLA07]. IRONCODE [Bai05]. irresponsible [Got04]. IS/IT [Gor00a]. IS2001 [Gor01a]. ISA [PT03]. Ishmael [Lis06a]. Island [ESS^{+}07]. Isn’t [LFG08, Noy02]. isolated [HCHP01]. Israel [GE06]. Israeli [EH05, EH07, GEHR09, LA07, LA08, fCST02]. ISSAC [Hen03b]. issue [Got06b, Imp05b, Imp06b, Kno00a]. Issues [SFD08, CMGR07, Cle08b, Cle09a, Deb04, Fle06, GESZ09, GY00a, GY00b, Gol01, Gor01b, GHG^{+}08, HL06, IK06, Joy02, JY04, LAP08, LGA^{+}01, LC05, MKS02, MK06, Mos03, NP02a, PHS^{+}07, PQEA^{+}04, Phi04, Wal08a, Wal09a, Wal09b, Cle09b]. IT4PS [ASPS04]. Italy [DGS06, Car02b]. iteration [Man02]. Iterative [BL05a, Boi04, Boi06]. ITEST [PBWS06, PB06]. ITiCSE [ALRY08, BCK04, CJGU02, CFPL05, DSF^{+}03, DGS06, FKCM01, Fin01, Fin02, Fin03, HPT07, Cas08, CMT04b, JKM07, TP03]. ITiCSE-01 [Fin01]. ITiCSE-02 [Fin02]. ITiCSE-03 [Fin03]. IV [McC06a, McC06b]. ivory [Reg02b].

J# [GS05a]. J2ME [UCJ^{+}04, Utt06]. J2SE [Utt06]. Jackson [Our04]. Janus [Ada06]. Japan [NSH^{+}09]. Java [RBC^{+}05, RBC^{+}06, OPS^{+}02, AF03b, ACS02, BC07a, Bar02a, Bec01, BZ05, BZ07, BVPE06, BB00, BD01, BB03a, Bur03, CM05b, Chr05b, CF02, CR02, CL08b, CLP06, CHUB08, Cul00, DMU02, DSCU01, DUK02, Djo08, DHWH03, DHRH05, DMKN02, ET05, Edm09, Ell00a, ES05b, ET02, EGST08, Fek08, Fle00, Fle01, Fre04, Fro08, Gag02, Gri00, Hag00, Han05b, Hei03, Hel07b, HCMM00, HKF00, HMRM03, HSSC05, HSB09, Hun05, JT04, JMS02, JBMP03, JKKL04, JCOP07, KW01, KR00, KR01, KB04a, Kum04, Kum05b, LDB^{+}03a, Les03, Liu08, LHS04, LS04b, LH02, MDS04, Mer00, Mil09, MWM01, MB05, MHK^{+}03b, NW03, NB01, OJJ00, Par04b, Par04a, PH03, PH04a, PR04, PJ09, RRP00, RM07, RRP01, Reg00, Reg02a, RM08b, Rob00c, Rob07a, Rob04c, Rob06b, Ros02, Röß06]. Java [SL07b, Sch00, Sch02b, Sci07, Sco03, Soo01a, SP03, TZ01, Utt06, WVMN05, WW06a, Wic03, Wis06, Wu05, XSD07]. Java-based [CR02, CL08b, HKF00, JT04, JKKL04, Röß06, Sci07]. java.net [Gag02]. JavaScript [Lab09, MDS04, Ree01, Soj03b, ZJ03]. Javiva [TZ01]. JAWAA [AFJ^{+}03]. JDBC [DUK02]. Jeli [Rob00c]. Jeliot [MFR^{+}09, MMBAS04]. Jenuity [vTNC08]. jeopardy [Ber05]. JERIC [Kno00a]. Jeroo [SD03a]. JERPA [ET02]. JEWL [Eng01, Eng04]. jFAST [WW06a]. JFLAP [CFR04, LJ08, Rod06, RWL^{+}09]. jGRASP [CH06]. JHAVÉ [NEN00]. jigsaw [Soh06a]. JKarelRobot [BS01]. JLS [PK08]. Job [GY02, BS08, Wal01a]. Job-hunting [GY02]. joint [CF04b]. Josephus [Hen06a]. journal [Kno00b]. joy [GCD^{+}09, MRGS08]. judging [GMFA09]. July [ALRY08, DSF^{+}03, Fin03]. June [BCK04, CJGU02, CFPL05, DSF^{+}03, DGS06, FKCM01, Fin01, Fin02, Fin03, HPT07, Wal01b]. junior [CNCL09, TH06]. JUnit [WSW05]. Just [BF05, OG01, Dav09, HKN^{+}02, Mar01b, PSW03, RJAH04]. Just-in-time [BF05, Dav09, Mar01b, PSW03].

K-12 [BT06, Goo08, HEB08, OM08, CS06, HH05b, LZBO09, TDJ^{+}03, TMD^{+}04, VCC^{+}05, VCC^{+}07]. K-8 [BT06]. K.I.C [Raf08]. Kaleidoscope [Wic01]. Kangaroo [Gin08a]. Karel [Bec01, Ber06]. Kaya [GDM05]. KDF [Wil04a]. KDF-9 [Wil04a]. Keeping [Kül09]. keeps [Tea02]. Kentucky [GWG02, Imp02b, RHDR07]. kernel [Hol00, PRB09]. kernel-based [Hol00]. key [DS08, HAuR04, HGEB08, NM02]. keyboard [Dic07b, NL07]. Keynote [Rob03b]. keys [Ber09a]. keyword [Gin03c]. Khepera [HLM01]. Kids [Raf08]. Killer [ACD07]. kinaesthetic [CMCB09]. kinase [CSM^{+}07a]. kind [GS09]. kindergarten [PD06]. Kinesthetic [BGW04, SP07]. kit [MNN^{+}09]. knapsack [Rol07]. know [BB06, CLM^{+}06, FLS04b, Fit09, GGH03, HV05, MT05, PD06, WG07]. Knowledge [BF06, BP05, Mik05, SJH06, CG02, CTP^{+}04, Del00, FLS04a, GP07, Gha05b, HL04b, KN09, Kop07, Kre09, MT05, MMW^{+}05, Pho01, RC04a, RF00c, SV03, SNH08, SDLdQ05, Yue07]. Knowledge-based [BP05]. knowledge-driven [SNH08]. Kong [WD03]. Korea [PRC06, YSLL00]. Korean [CBJL08]. kRobot [HLM01]. KwaZulu [GM07].

Lab [Erl06, AL05a, BED^{+}03, BDM^{+}07, CJ01b, Che00, CTR^{+}03, GP00, HLNT01, HMS04, Lar05, MSP^{+}06, MR02, Pra05b, Rad06, Ram03, RKDM01, Rou02, Wal04a]. lab-based [CTR^{+}03, Rad06]. laboratories [ATNT04, BR05b, DTW07, GGS06a, JKDS01, RKDM01, SSP^{+}05b]. Laboratory [Dor07, Way06, BP03, BCF^{+}00, Bry08, CG02, CD08, CFW01, DKP^{+}06, FMJM02, HCHP01, LCL08, Mat03, McC05c, MD06b, O’L06, PMF08, Rio02, WW04, Wea04, Wic05, WWWG08, WWGT07, YH04]. laboratory-based [Mat03]. Labs [Les03, See08, HM05, Lis01a, MD05, PF03, QLT09, SSP^{+}05a, Soh06a]. lacking [vVML04]. lacks [Mar09b]. LAMP [Qid07]. LAMP-LEGO [Qid07]. Language [CCH03a, Ort08, Ran00, RRKS06, VVV04, Xu08, AMUFVI09, App01, BdMG08, Bea03, Bel02, Ber01, BD01, CMS06b, CH07, DLPQ04, DWJ04, Fle07, GGÁC03, Got01, Gou09, Har02, Hen05c, HH05a, McC01b, Phi04, RM01, Rob04b, RF01, Ruc07, SL08a, SFPFV08, Sig07, Sta06, Sto06, TF07a, VAFM03, VS06, Yeh03, Zil05]. Language-independent [CCH03a]. languages [ABB^{+}06, AC07, AOMC07, BBF^{+}09, CCY03, CGM04, FK04, GET06, GS05a, HW06, KBC01, Kum05c, Mau02, Mea06, MSK09, Pil09, Rod06, RLR07, SK02, TN02, VI05, WD05, WS01a]. laptop [ACd08]. Laptops [CP03]. Large [Cle01a, CHS05, AR00b, Bro09, CM02, Cle05a, LCB^{+}03, Par06b, RF00a, RBM04, SHL^{+}00, TM04, Wol07, Wol02]. large-scale [Bro09, RF00a, RBM04]. last [WW09]. late [Got09b]. launching [PC08]. ‘Laurel [AF03a]. law [GPP00, Kay00, Lee03b, Par03a]. lazy [NW00]. LC [BN09]. LC-3 [BN09]. LC3uArch [BN09]. lead [CHMG01, SMH05]. Leadership [Mar07b]. leading [LA07, Lap07, RH03]. leading-teachers [RH03]. leak [MD04]. learn [BG08a, BGDR05, BW06, CT09, Ere00, GM06a, GKL^{+}02, HDH06b, HP06, Jin08, Tho01]. learn-by-doing [BG08a]. Learned [DHRH05, Bro08, DLM09, Fit09, HV05, Jam05, MIS^{+}07, MR02, MKBW07, PPJH09, PD06, Wil07]. Learner [Lau06, App06, KB09]. learner-centered [App06]. Learning [Aba08, Bur08, CS05, CT07, Fri04, GSyC07, GMN02, Gin08b, Gra05, dOG07, Jac08, KPCM06, KB04a, KAPG07, Kum01, Kum04, LDHW02, LNCC09, Les03, LR04, Lom07, Mar05a, MPJC09, Nod03, PFF04, Pil09, PM02a, PWA06, RW05, Rod06, SHVP05, Son09, TRWJ02, Tho00, TBR05, VA02, ZTS^{+}09, AEPQ06, AKCC06, AL07a, ABE^{+}09a, APR07, AAC^{+}05, AO06a, AAD^{+}07, AL09, AHM^{+}04, Arn07, BC07a, BHRB08, BRP^{+}07, BCM05, BC08, BGW04, BC06a, BYS09, BC06b, BL00, BPK^{+}01, BGGR01, Bot05, Bow08a, BPW^{+}09, BROB00, BGKT00a, Bro08, Bud06, BCB09, BAdC07, BdCGB07, BB01, CHMG01, CS02, Cas02, CSA^{+}07, Cha00a, Cha01, Cha02, CBP03, Cha00b, CMD08, CO00, CGM04, CS08, CC01, CRPZ06, CM01, CHM^{+}01, CK08, CK04, Con00, CT09, Dav01a, Dav06, DW05a, DTP03, DDS08, Dic05]. learning [Dix04, DFS03, Dra09, EB09, ESP03, EMM00, EA00, Ell06a, ET02, FMJM02, FLLM06, For04a, FAL^{+}05, FJA^{+}07, GE07, Gar04, GDB03, GGÁC03, GDHW06, GCDJ08, GM09, GCW^{+}03, Gib00, Gin01c, Gin03b, Gin09c, Goe08, Gon06, GKL^{+}02, GS06b, GA03, Ham07, HAM00, HR02, Haz02, HW06, HRBC03, Hob00, HP07, HRB^{+}09, HYN03, HP06, Hum04, HLCS08, wH09a, JKS00, Jam02, JS06, JS08b, JG08, Kac03, KK02, Kaw07, Ker04, KH00, Kol08a, Köl08b, Kum09a, KMN^{+}03, cL05, Lap02, Las09, Lau07, LCW06, Lee01b, LR07, LW05, LOP^{+}07, Lin07, LDK^{+}07, LHL03, Liu06, LR05, Lud05, LNR05, LLB02, LM03, MMN^{+}09, MPK^{+}08a, Mar09b, Mat06, MA01, McC03a, McC05b, McC05a, McC06c, McC06a, McC06d, McC06b, MF07, ML04, MC04, MR09, MBD04, MHK03a, MSK09]. learning [MICM08, MRM^{+}06, MPK08b, NS02, NSS^{+}05, Nyg02, OG06, Ola00, OGMA07, Ols05, OCLL00, PBK04, PBDWL09, Pan09, PR02, PPE03, PMF08, Pet09, PCW09, PA03, PA06, PJ01, PMG^{+}04, PW01, Pre07, Put01, Qua06, Rad07, Rad09, RH03, RLW04, RHE04, Rio02, RCF00, Ros01, Ros07, RJM^{+}08, RMN06, SOSL04, STTBC07, SV00, SB07a, SNH08, Sha00, She04a, SYLL00, SFPFV08, SAHS04, SJ04, SP07, Ska02, SJA04, SS00, SMS08, SKM02, SC08, Sur02, Tal06, TLH^{+}09, TSO05, TL01, Tra03, TM03, Tru05, THH09, Van04, VIPC09, VVV04, VCJ09, Vin05, WAS^{+}02, WC04, WHCT09, WI09, WB09, Wir03, WM00, WBPW06, YOB08, dBGdMB09]. learning-based [RH03]. leave [Pop08b]. leavers [BBY08]. Lecture [Cha07, ZH02, AAS^{+}04, MSP^{+}06, O’N05, Par06b, Pul00, RLM08, Wir03, Wol02]. Lecturelets [Cul00]. lecturers [BAdC07]. lectures [BHR07, Cul00, DAH08, LRKM07, RF00b, RTM^{+}04, RBM04, Sca00]. lecturing [Wal06a, WGS05]. led [BYS09, DW05a, HRB^{+}09, HLCS08]. Lee [Lee03b]. Leeds [BCK04]. left [BT05, CMRR04]. LEGO [Bar02a, GMN02, Imb03, JJ05, Kla02a, Kum05a, Qid07, JCOP07, MGFK06, SS07, Wol01]. LEGO(R) [MO09]. LEGOs [BMK03, Law01, LDB^{+}03a]. LEGOs(R) [HH05a]. LEGO^{T M} [LDB^{+}03b]. lemma [GET05]. lemonade [Wol02]. lengthy [Gin09b]. less [OHZ01]. lesson [SH08]. Lessons [DHRH05, Mar04b, Wal03b, Wil07, Bro08, DLM09, Ell06b, HMS04, Jam05, cLwW07, cLwW09, MR02, MKBW07, PPJH09]. Let [PPP06]. letters [PT02]. level [AL05a, ACSS06, BT06, Bou09, CCD^{+}08a, Cha04b, DCP^{+}09, Gor06b, GLC07, Ham05, HKCW08, HKCW09, Haz02, HPC^{+}03, Jen08, Joy02, JY04, Joy06a, Kla02a, LM03, MAFW08, MMN^{+}09, MVD01, RH03, SD08, Yeh03, vTNC08]. Leveling [TMS07]. Levels [DK01, PK06, BS00, BS01, PGK05]. Leveraging [HS02, LCL08, AJD09]. lexical [SNH08]. liberal [ABB^{+}03b, BBN^{+}04, Cli06, Con00, DH06, Doo03, Gup01, HS03, MKBW07]. libraries [CTLW03, Fek02, ICL02]. Library [CIL^{+}03, AHJ^{+}07, AH08b, BDM01a, CS00a, CFJ^{+}02, CIK^{+}03, Fro08, Mur07a, PKB06, WK06]. licensing [Got00a]. life [Cle02a, Got03, LV08, Sud08, Wic05]. light [Sam02]. Lightweight [DKL^{+}07, ACS02, CH06, HCB04, vTNC08]. like [BMT^{+}07, BMT^{+}08, BMD^{+}09, Gou09, KW01, TT05, WBBC09]. likelihood [PB03b]. likely [Deb02]. limit [JDC00]. limited [BAMSK05, BPD03, RS04, WYOH02]. line [ABK^{+}01, CW03, DeP06, FH01, GMFA09, Hui01, PHC08, PW01, RCF00, Sol07a, WPS01, WM00]. linear [Bla09d, Min07, Min09, Yan08]. linearization [JWC07]. lines [Chr05b]. linguistics [Hol05]. link [FM09, RB09]. linked [DMU02, Rol06a, She04b]. Linking [BR09, HEB08, Nod03]. links [McC00a, McC04b]. Linux [Dav04, NV05]. Linuxgym [Sol07a]. Lisp [Kla04, Man09]. List [Rol05, BMG03, Bru04, Bru05a, Par02b, Rol06a, She04b, She08]. lists [DMU02, Dia05, Hu05]. lite [AKWZ07]. literacy [Gup06, HV05, HV07, MGS07]. literature [GC02, Joe06b, PSE^{+}05, PSM^{+}07]. little [Imp05c, Pat06, SH08, YB01]. live [COR07a, Cle03b]. lives [RH07]. Living [Joe06b, Bou03, GS03]. LL [GOGPGP08, Pij00, Tay02a]. LMS [Qua07]. load [Sok02]. Local [BHRB08, PFVI00, AM07, PTT08, Rol03a, RRD^{+}09]. local-remote [RRD^{+}09]. location [Gor05]. logging [Rob00c, Rob03a]. logic [ACL^{+}01, Arn07, Eth04, Kur08b, LLB02, PK08, RC00, SD08, SD09, War04]. logo [NL07]. logs [Cha00b]. Long [HY09, Cle03b, LM06, ESS^{+}07]. Long-term [HY09]. look [DMI^{+}09, FNB^{+}09, La 06, LLLC06, MV05, Whi09]. looking [DVE06, PEH07]. looks [GAGG04]. loop [PHMS08, WP06a]. loops [EB08]. Louis [DNTB05]. Loving [GKL^{+}02]. low [SC07, ZF04]. low-cost [SC07, ZF04]. lower [BPS09, Ham05]. lower-level [Ham05]. LR [GOGPGP08, Pij00, Tay02a]. Ludwig [Sha05]. LYDIAN [KPT00, KPT03].

M [Sor05]. MaBL [AC07]. Mac [Ber03a]. Macchine [GM06a]. Machine [Con00, Boy08, GCW^{+}03, NBL09, RMN06, SWR05, Sto06, Yeh03, BD01]. machines [BBS06, GOMSJVGP08, JKDS01, Lee00a, Sch09b]. machinima [vLK09]. Machshava [fCST02]. Made [YLS05, DW07, PR04, RH08b]. Madrid [ALRY08]. Magic [Kru03, CM08b]. Mailing [San09b, Bru04, Bru05a, Dia05]. maintaining [Bro08, HHS02]. maintenance [HBIG03, TMSZ07]. major [BMK09, BMP06, Car06, CIL^{+}03, FK02, PN08, PT02, SSW08, Tut00]. majoring [Kla05b]. majors [AE01, AKM00, Coh02b, Cor07c, Etl06, Goe08, GD07, Gou06, Guz03, GF05, HHK^{+}09, Kac03, KH05a, SL09b, ST08, Sur02, TATL07]. Make [Abo09, La 06, ABH^{+}07, Bru05b, DKL^{+}07, HGEB08, KMMS02, LWS07, Lee02d, Tow02]. makes [Ni09, VZ06]. Making [DCP00, Deb02, Dem05, Hen02, Kur09b, PJ01, Ros07, SV03, Wol02, Yeh03, Zel00, Bri09, DWJ02, GCD^{+}09, Kno02, LC08, MRGS08, Ruc07, SSA08]. male [Laz02b, Laz02a]. Man [YB01, ADM^{+}09, Pau08]. manage [FH01, RW05]. manageable [Lee03a]. Management [CM02, BAET09, BCC^{+}05, CKM07, FH01, Fek05, HH06, LMS00, Mer00, Nør00, PDS07, Par02a, POM^{+}08, Pho01, RW07, RNH^{+}06, RJM^{+}08, She04a, VT01, Wol01, WP07, Xiz01]. managerial [BC07e]. Managing [AGGR04, For04b, Hov05, MDL^{+}02]. mania [Hen04a]. manifesto [Lis03b]. manipulating [DSCU01, Hab02]. manipulation [KH05b, MR05, SWR05]. manual [Che00, Kac09a]. many [BHYA05]. map [KRTT01, LDB^{+}03b, OTUW07b]. Maple [Phi04]. Mapping [YLS05, AC07]. mappings [UD03]. maps [Hoo08, Ols05, SBE^{+}08]. March [ACM06, BTHR06, CDWH00, DRFG08, GWG02, Imp02b, Imp04, JKDN04, RHDR07]. marine [BCL00, NRW04]. marker [SD00]. markers [PB03b]. market [Wal01a, WW06b]. marking [BGNM04, GH06, HST02, Tho03b, TWS05, TWS06, TWS09, TLS07]. Marmoset [SHP^{+}06]. MARS [VS06, Lis05b]. mart [SL06]. Mass [HD04, MKS02]. masses [Cha07]. master [dMBG^{+}08, VKV01]. MasterMind^{T M} [LC06]. Masters [CCD^{+}08a, CCD^{+}08b, CDS09, DCP^{+}09, Gor06b, TLW^{+}04, TLW^{+}09]. mastery [Jac08]. matching [FNW02, RS09]. material [BL01b, Joy01, KK03, POMA08]. Materials [HHF^{+}03, ABK^{+}01, CSA^{+}07, DMI^{+}09, HHLS06, Soj03b, SC08]. Math [HBEM03, Hen08b, Hen09a, Hen09b, ALH07, Bea02, Hen02, HDH06a, Hen07a, Hen07b, HD07, JJK03, KTH^{+}00, LL06a, MAR07d, Noy02, Pap03a, TKB01]. math-phobic [KTH^{+}00, TKB01]. mathematical [Gin07c, GMHS01, Haz03a, HBD^{+}01, HHF^{+}03, MRH09]. mathematician [DV04]. Mathematics [CIL^{+}03, ELE^{+}02, Hen05a, Hut04, Wal05b, AGHM02, AHH^{+}06, BMW04, CF04b, FM09, Gol04b, HEB08, HL07, HDA^{+}02, Kos00, MBPB03, Mar05b, Mar06, Pap03b, Ral05, Sur02, TBS^{+}00]. mathematics/computer [CF04b]. MATLAB [JDFV08]. Matrix [KMS01]. matter [Abo09, JM02, LS05]. mature [Ras03]. maturity [Tho00]. Mauritius [ABB03a]. maximum [PB03b]. maximum-likelihood [PB03b]. may [Kac06, Kac08a]. maybe [HKN^{+}02]. Maze [RRP02, Gig02, Sna06b]. MBA [Lea01]. MBTI [Cap02]. McKinley [BA06]. MCQ [DLRHP09, Fon04]. MCQs [TG05]. MDT [PA03]. me [Har01a, Lis06a]. mean [BABBH04, DGSW06]. meaning [Cle01b]. meaningful [LWS07]. means [Bec08, OHZ01, VIPC09, WPS01]. measurable [MGH^{+}06]. measure [Ric09]. measurement [CTBS06]. measures [Imp09]. Measuring [FM03a, PB03a, Cle08b, Kac03, LV05]. Mechanics [MKBW07]. mechanism [WP06a]. mechanisms [Ree00, WC04]. media [ABB^{+}03b, DTP03, Guz03, PR02, RWP03, Shu06b, TRT05, WBS04]. MediaMime [SKM02]. medical [AL07b, TLW^{+}04, TLW^{+}09]. Meet [Imp05b, Imp06b, Bec03a, Cun07, RW07]. meeting [Lis04b]. meetings [Lis06b]. meets [Lis06a]. Melbourne [Ell00b]. Mellon [FM02, FB02]. member [Wal02a, WSP04]. members [Bru04, Bru05a]. memory [HL04c, NR05, PWH00]. Men [AL06]. Mental [CLB07, GSG03, Mir09, SGG06, MFRW07, MFR^{+}08, MFR^{+}09, RLW04, SG07, YBAD05]. mentor [MK02]. Mentoring [Kno01]. mentors [Tow02]. Merging [RNH^{+}06, LH05]. merits [Dav06]. message [Har00, PKB06]. messages [NPM08]. messaging [BAB^{+}08]. Meta [Fle06, Val04]. meta-analysis [Val04]. Meta-informatics [Fle06]. Metacognitive [Gin01c]. metadata [OB05]. metal [HHB04, NC07]. metaphor [Mil09]. Method [SP03, BR02, CMRR04, DL06, GBO08, HK01, Hab02, Jac01, Ker04, Mat06, MCS04, Mér09b, MPJC09, RD00, WP02]. Methodological [Ran07a]. Methodology [MKC09, Ari00, BZ05, God06, Gri02b, Luq05, PSS09]. Methods [HL06, LH03, TW09a, AD01, ADG^{+}01, AHJ^{+}07, AH08b, BCD^{+}04, BH03, CSM^{+}07a, Cle03a, Cle04b, DG09, Dei05, Got04, HL04a, HAH^{+}06, HAJ07, Lis05b, LTHN09, Mai09, McC01a, Our04, Pho01, RBC^{+}02, Sha06, SF02, Sob00, SSS07, TG05, WS02, Zin08]. metrics [Con04, TW09a]. metrics-based [Con04]. micro [MMN^{+}09]. microarchitecture [BN09]. microbreweries [Pop03a]. microcode [VS07b]. microprocessor [Ska02]. Microsoft [FH01]. MicroTiger [VS07b]. microworld [SM05a, Xin03]. Microwulf [AB08]. Mid [KNT02]. Mid-term [KNT02]. middle [Ada07, CFKS02, CH09, DLM07, DKN08, LNCC09, RHL^{+}09, SD03c, SL08b, TBAA09, vLK09]. middle-school [CH09, TBAA09]. MIG21 [MS05b]. milestones [MGH^{+}06]. millennium [ANS00, YM06]. Mind [Kac09b, Kre09, dMFM09]. MindMapX [Dav05]. mindset [MT08, VB02]. mindshare [Bri06]. MINDSTORMS [Bar02a, EBG^{+}05, GMN02, JJ05, JCOP07, KC07, LDB^{+}03b, MGFK06, SS07]. Mindstorms(R) [MO09]. Mindstorms^{T M} [Kla02a]. Minesweeper [BA05b]. miniconference [SW04]. MiniJava [Rob01b]. minimal [Sco02]. MiniMIPS [BP03]. mining [LZL08, MR06, Mus06, Rah08, Roi05, WK06]. Minix [How02]. minor [BJ07]. minority [AJF07, Qid07]. minute [Hen05b]. MIPS [BBE06, PT03, VS06]. MISC [Sco02]. misconceptions [ST07, Tra03]. Misleading [Gin01d]. mismanagement [AEPQ09]. missing [Got02]. Missouri [DNTB05]. mistakes [Gin03b]. Misunderstandings [TQCPQE08]. Mixed [Lis05b, HPC^{+}03, Phi04]. MLE [HR03]. Mobile [BHBG07, RBTC03, Dra09, Esp08, HLN07, Kur09a, Kur09b, LDK^{+}07, MNN^{+}09, MLH05, WW09]. Moby [Lis06a]. Mock [CL08a, BFMC05]. modal [CS00b]. Mode [Dav04]. Model [AO06a, DDLR01, NSH^{+}09, PA03, Way06, AF03a, BHRB08, BC04a, BBN^{+}04, BF06, CS02, CCK01, DSE02, EH07, EGB05, FJ06, Fon01, Fon03, Gor00b, GGS^{+}06b, HGEB08, Hen08a, KDP06, KvD02, LV08, Ort03, PHC09, Pho01, Pre07, Qua04, RH08a, RF09, RRP01, RMPD09, RCF00, Sav05, SB08, Sch02a, Sch04, SW04, Soh06a, Soo01a, TYT^{+}09, Top07a, Top07b, WSP02]. model-driven [PHC09]. model-first [BC04a]. modeler [Ing09a, CK04]. Modeling [Hen04a, Boz06, HR02, TSO05, Wag05, War05]. Models [SGC^{+}01, AGE04, AGE06, Bar02a, CLB07, Cle00a, FKO07, GE09, GSG03, HNC08, IDD06, MFRW07, MFR^{+}08, MFR^{+}09, MAFO08, Mir09, RLW04, SCHS09, SGG06, SG07, SW06a, Tow02, TM03, TN02, YBAD05]. Modern [TS00, RC00, WJJD06, Xu07, ZSW04]. modes [FM09, JCK03]. modified [BYS09, LW00]. modify [PPP06]. modular [Gin09d]. modulators [CSM^{+}07a]. module [AEPQ06, CGR^{+}08, SL08b]. modules [GKL^{+}02, HPC^{+}03, HPCZ04]. Monash [ASA08a, ASA08b]. monitor [Rob01a]. monitoring [DDL03, YYLH06]. monster [GS09]. MOODLE [CGR^{+}08, DDS08, Doo09, Mor08, Rad08, RK09b]. Moon [ADM^{+}09]. moral [JM02]. Morning [DHWH03]. Most [Dal06, Deb02, Sur01]. motivate [MO09, SJ05, SJ06, WW06b]. motivated [Djo08]. Motivating [BVPE06, She06, Qid07, SM05b, War04]. Motivation [Jen01a, RMY08, ATNT04, Ges07, Jen01b, KF03, TW09a]. motivational [Rad09]. motor [NL07, NL08]. move [Edw04]. movement [MdL09]. movies [And00, PB05b]. Moving [Dia05, Doo04, TA08]. MPI [Rol08b]. MPS [GDM05]. MS [ND08]. MSIS [GGS^{+}06b]. much [Bec06, Par09, Reb05, SB06]. muddy [KNT02]. MulE [BSA^{+}05]. Multi [BHH^{+}08, NvdH09, AL05a, AHJ^{+}07, AH08b, AMW06, Bor07, CL00, CS00b, Con08, HHB05, HPC^{+}03, LAF^{+}04, LLLC06, MAD^{+}01a, MAD^{+}01b, MMW^{+}05, NRDP07, Par02a, PBM03, PNF00, RH03, RF00a, SD08, Sha00, SR07, SSC00b, TAHA00]. multi-agent [Par02a, SSC00b]. multi-continental [Sha00]. multi-course [AL05a]. multi-disciplinary [LLLC06, PBM03]. multi-discipline [SR07]. multi-expertise [NRDP07]. multi-functional [AMW06]. multi-group [RF00a]. multi-institutional [HHB05, MAD^{+}01a, MAD^{+}01b, MMW^{+}05]. multi-level [HPC^{+}03, RH03, SD08]. multi-modal [CS00b]. multi-national [LAF^{+}04, MAD^{+}01b]. multi-perspective [AHJ^{+}07, AH08b]. multi-platform [Con08]. Multi-player [BHH^{+}08]. multi-science [TAHA00]. Multi-site [NvdH09]. multi-threaded [CL00]. multi-user [Bor07, PNF00]. multiagent [Soh04, VB02]. multicast [Bri06]. multicore [ES08]. Multidisciplinary [EGMS08, SHS^{+}08, WWK^{+}05, BJ07, BPD03, HHK^{+}09, dMBG^{+}08]. multilanguage [Sha02]. Multimedia [Ros00, BPK^{+}01, CPJ07, CC01, Ciz03, CK04, DW05a, EA00, Ell00a, GSVG05, Guz01a, HTBS01, MTHS00, MMRI07, MS05b, MWPB05, PBK04, Ros05, Rya01, Sca00, SMTH00, SSW08]. multimedia-based [Sca00]. Multimodal [DGSW06]. multinational [MAD^{+}01a]. multiparadigm [GGÁC03]. multiplatform [Sha02]. multiplatform/multilanguage [Sha02]. multiplayer [Ree00]. multiple [AC07, Bec03c, DTP03, Fek02, Fon02, FS06, Jac01, MH05, Soj03c]. multiple-goal [MH05]. multipurpose [PA03]. multithreaded [BCHS00, CS00a, She02]. multiuser [Sci07]. murder [Got08]. MUSE [Ram03]. Museum [ASA08a, ASA08b, GM06a, Wil03]. museums [Imp01]. Music [HGK09, MBK09]. musical [Ham02b, Ham04b]. Musicomputation [CKMP09]. must [RWGM00]. mutatio [Con08]. mutual [RSUFPF08]. MUVE [MT09]. mystery [Reg08]. myth [Got00a, Yan09a]. MythematiCS [Pap03a, Pap03b]. Myths [Lew00].

narrative [ABB^{+}03b, War04]. narratives [ABO^{+}05]. narrowing [Dic07c]. Natal [GM07]. National [CIL^{+}03, Gor01b, BT01, BBT07, CMT04a, CMT04b, CGT05, Cun05, LAF^{+}04, MAD^{+}01b, PMT03, VCC^{+}05, VCC^{+}07, BP02, Joh02, MR00]. natural [Cli04]. Naturally [LCS^{+}09, Lis09a, Rol05]. nature [Cle01b]. navigation [Dic07b, JB07]. necessary [PW08]. necessity [Cle06b]. Need [Kum09b, Bea02, Bea03, Cle01a, Fit09, HAG05, Hil00, Kac01b, Kac01a, Mar01a, OHZ01, Ral05, RC00, SSA08]. needed [PD06]. needing [GP00]. needs [Bec03a, HMS08, RW07]. neighborhood [TBMC08]. Neither [Boy08]. Net [Bel02, KTKP02]. NetPrIDE [CGI05]. nets [Bar01, Bar02b]. NetWire [CD01a]. Network [Chu01, Chu03, ES05b, JBMP03, Lin07, Pul07, Str00, AL07a, BBS06, CM05a, Cas02, CD08, Cig03, CGI05, DEJ^{+}06, DJLA07, Dav01b, DKP^{+}06, ES05c, FMJM02, GL07a, HCHP01, JMS02, MA01, MV01, PFJC05, Ric00, Ste03, VMD01, Wal05a, WWGT07]. Networking [JBMP03, SLSM02, AE01, Bog04, Bor07, BCF^{+}00, CW03, CWM05, Cha04a, Erl06, FMJM02, Gag02, GD08, HL08, Jan01, MMN^{+}09, NN00, RW02, Ric02, RS04, RMMH01, Str00, WWWG08]. Networks [JKKL04, AAF^{+}06, CD01a, Cha04a, Cle07b, ELC07, ENO09, Fon01, HCHP01, Imb03, LOP^{+}07, TT06, TY06, TBM09, Yan09a]. neural [ELC07, Fon01, Imb03, Yan09a]. Nevada [Ano03, GKJD03]. never [Deb02]. newcomers [BP05]. Next [Fre04, Cli04, Haw09a, McC00a, SRR00]. nexus [TM04]. Niftiness [Par04c]. Nifty [HAB^{+}08, PCR^{+}01, Par01, PER^{+}02, PPR^{+}03, PML^{+}04, Par04a, PLA^{+}05, PWM^{+}06, PCS^{+}07, Par07b, Par08, PMS^{+}09, Par09]. nine [LW00]. nine-tails [LW00]. No [LSGB07, Bib06, Cha06a, Cle06a, VR04]. nomenclature [MHW07]. Non [BKM^{+}01, KKR06, AKM00, Cor07c, DV04, Gin09d, Goe08, Gou06, Guz03, GF05, LFL06, OG06, Pap03d, SL09b, Str00, Wis05]. non-CIS [AKM00]. non-computer [Pap03d]. non-distributed [Wis05]. non-majors [Cor07c, Goe08, Gou06, Guz03, GF05, SL09b]. non-mathematician [DV04]. non-modular [Gin09d]. non-prescriptive [OG06]. non-profit [LFL06]. Non-programming [BKM^{+}01]. non-specialist [Str00]. Non-traditional [KKR06]. noncompact [WZ00]. nondeterminism [ABA09, BA07a]. nondeterministically [ALBA08]. nor [Boy08]. Norfolk [Imp04, JKDN04]. normalizing [Jac01]. North [ACM01, WMGR01, GE08b, Kla05a, Wil07].

Northern [GWG02, Imp02b]. notation [CCG05, Rou04, War05]. Note [LWS07, SM00b, SDG^{+}08]. Noteblogging [SDG^{+}08]. Notes [Wal00b, Wal01a, O’N05, Wir03]. notions [POMA08]. novel [HEB08, KAPG07, Moh09, OM08]. Novice [ET05, Yue07, AEH05, CMS06b, CL04, Cle09a, DSE02, DLPQ04, ES06, FNB^{+}09, Gin03c, Gou09, HB09a, LAMJ05, LV03, LAF^{+}04, LST^{+}06, LCS^{+}09, MFRW07, MFR^{+}08, MFR^{+}09, MLM^{+}08, Pil03, PJ05, RBJ^{+}09, SDE01, SCL^{+}08, Soo09, TRT06, VS07a, YK07]. novices [BC07a, DCP01, EDES05, HK01, NSW08, NPM08, SFM^{+}07, ZTS^{+}09]. NP [BA05b]. NP-complete [BA05b]. NSDL [CIL^{+}03, KCI03]. NSF [Pre01, CIL^{+}03, CTP^{+}04, CFT08, CPT09a, PBWS06, SLPG07, TP03]. NSF-DGE [SLPG07]. NSF-funded [PBWS06]. number [Bec03c, Kla05b, Kru03, Man05, Nau06, WZ00]. numbers [GH01, Kru03, PPRV06, Rol03b, RSHL07]. Numeri [GM06a]. numerical [Rol03b]. NURBS [FLS04b]. nuts [CLP^{+}01].

O [Gri00, LR04, Rob06a, Rou04]. O-O [LR04]. Object [AF03b, AV02b, BB00, Ing09a, LBB^{+}01, PH03, PH04a, Sax03, SP03, Wic03, AL02, Ari00, BC06a, BKHL02, BHN^{+}09, BDM01a, BHLW02, Bur08, CC01, Chr05a, CK04, CF04b, CH06, CHJB07, Dew05, DSCU01, ET05, Ext00, Fei07, For04a, GP07, Gel00, GL07b, GL08b, Ham05, HR02, Hol05, HP06, Jan01, KR01, Köl08b, Lew00, Lin05, LS08c, Mar07c, MSK09, MWM01, Noo00, NSS^{+}05, Nyg02, OBL03, OBL04, PHC09, RBA05, RD08, Ran00, RRP01, Ras03, RC00, RMT09, SD03a, ST07, SBE^{+}08, SSS07, TH06, TRT04, TRT06, TQCPQE08, UD03, VZGE07, Wan01, WSP04, Xin03, Ran00]. object-first [BDM01a]. Object-orientation [BB00, Lew00, Ran00]. Object-Oriented [PH03, Wic03, Ing09a, LBB^{+}01, Sax03, AL02, Ari00, BC06a, BHLW02, CF04b, DSCU01, Fei07, GP07, Gel00, GL07b, GL08b, Ham05, HR02, Hol05, HP06, Jan01, Köl08b, MWM01, Noo00, Nyg02, RBA05, RD08, RRP01, Ras03, RC00, RMT09, SD03a, ST07, SBE^{+}08, SSS07, TH06, TQCPQE08, UD03, Xin03]. object-relational [UD03, Wan01]. object-understanding [Lin05]. objections [JS08a]. objective [ALR08, Lis01b]. Objectives [Lis01b, Cas03, CR03, SMS08]. objectKarel [SM05a, Xin03]. Objects [BBK05, DDC^{+}03a, KR00, LS08c, NW03, PRR02, Aba08, AHL04, ABK^{+}05, BKHL02, BCS^{+}07, CBP03, Chr05a, CDP03, ES06, Ere00, Fos05, GA03, HAB^{+}08, JJOA07, LBC^{+}06, MR05, MWPB05, PPP06, PQEA^{+}04, Rou02, VCJ09, Vin05]. Objects-early [BBK05, LBC^{+}06, PQEA^{+}04]. objects-first [BCS^{+}07, CDP03, MWPB05, PPP06, Rou02]. Observational [TL01]. Observations [JCK03, BM05, CCC07]. observed [MKLH09]. obsolescence [Kla00]. occam [JJ05]. occam-pi [JJ05]. occupational [Huf02]. occurring [Lis09a, LCS^{+}09]. Octave [MAMC05]. ODIN [WPS01]. off [ABK^{+}01, LCB^{+}03]. off-the [ABK^{+}01]. off-the-shelf [LCB^{+}03]. Offering [PW01, Old08]. office [Mal09]. officers [Bry09]. offline [CBJL08]. Offshore [FKMR04, AMVZ06, HFHK05]. oft [Rol08a]. okay [DSM01]. older [GS05b]. olympiad [KH08]. on-campus [FH01, HE06]. on-computer [Jac00b]. On-line [WPS01, ABK^{+}01, CW03, DeP06, FH01, GMFA09, Hui01, PW01, RCF00, WM00]. on-the-fly [Esp08, Gin09d, RL09]. Once [VZ06, Bai05, DSE02, Got03]. One [BT06, Hen05b, Lis06b, Rol02, BCS^{+}07, Ber03b, GG06a, HB09b, Mer01, Yeh03]. One-minute [Hen05b]. One-time [Rol02]. one-to-one [Ber03b]. Online [GE08b, Kum05b, AR08, AL07a, AKA09, BFD^{+}09, BSA^{+}05, Bow07a, Bow07b, Bow08a, BL01b, Dal06, DDL03, DFS03, Esh07, HH07, Hel07a, HPC^{+}03, HPCZ04, Jac00a, Kum05d, LL06b, LS08b, MDM04, OTUW07b, PPE03, Pre07, QLT09, RH08b, SOSL04, Sha05, SJ02, TE08, WC04, WM03b, Wos06, Wu05, ZJ03]. only [Liu06]. onto [LSG05]. ontology [CSS05, CDF^{+}07, CSD^{+}07, CLMW09, DCT06]. OO [Ast01, BCB^{+}08, GLR^{+}01, Gri08, PG06, Pro09]. OOA&OOD [PY05]. OOD [AF03b]. OOP [Ada06, BVPE06, Bor04, BB03b, MH05, Nod04, SH08, SM05b]. OOPtutor [Gel00]. Open [Mer01, PBOP07, WBGM02, BSL^{+}08, CFW01, CSP09, DKP^{+}06, DDA^{+}07, DD08, EMdL^{+}07, GSW08, HW08, HPM^{+}08, HETD09, KMN^{+}03, Liu08, McA08, MTBS09, MWG08, MdL09, NN00, PL05, PT02, Ros01, SBG^{+}08, GE06]. open-ended [HW08, KMN^{+}03]. open-source [CSP09, EMdL^{+}07, Liu08, McA08, MWG08]. OpenGL [LM05, TF07a]. Opening [Har05, TF07b, GE08a, HK01, Yai08]. Operating [Dic00, How02, AE01, AS02, Bha00, Bla09a, Bry08, CFW01, Dav04, Don01, Don08, GGS06a, HRBC03, HLS02, Hol00, wH07, wH09a, wH09b, Lab09, MMP05, NB01, NV05, PMF08, PRB09, PP06, Rob02, She07, Wis05]. Operational [WWWG08]. operations [Cha03b, Rol06a, WK06, War05]. operative [SM02c]. opinions [FS09]. opnet [HL08]. opportunities [BT01, BP02, BBT07, BDDL04, CFT08, CPT09a, CMT04a, CGT05, FLLM06, GPL^{+}05, HILS05, Jon05, MR00, NRRR06, PMT03]. opportunity [Jon02, Sie06]. ’opt [Got09b]. ’opt-out [Got09b]. optimization [Rol03a, VT01]. Oracle [Udo06]. oral [AAO^{+}08, BAK^{+}09, Gha05b, Mar06, OAB07]. Order [PP07]. orders [Lis03a]. ordinary [Mar09a]. organisation [Ram02]. organization [BSV04, BT06, BP05, CHK^{+}01, CK02, DLM07, Fri04, FB02, HB01, JHLT06, She03a, SSC01, To04, WYOH02, ZSW04]. organization/architecture [WYOH02]. organizations [LFL06, MC08, Tho06]. organize [CIK^{+}03]. organized [CA02, Hoo08]. orientation [AV02b, BB00, KR01, Lew00, OBL03, OBL04, Ran00, TPK^{+}06]. Oriented [PH03, Wic03, AL02, Ari00, BC06a, BHN^{+}09, BHLW02, Bur08, CIK^{+}03, CC01, CF04b, CSP09, DSCU01, Ext00, Fei07, GP07, Gel00, GL07b, GL08b, Ham05, HR02, Hol05, HP06, Ing09a, Jan01, KP04, Köl08b, LBB^{+}01, LL04, Mar07c, MC04, MSK09, MWM01, MGH07, Noo00, Nyg02, OBLH03, PHC09, RBA05, RD08, RRP01, Ras03, RC00, RMT09, SD03a, ST07, SBE^{+}08, Sax03, SSS07, TH06, TRT06, TQCPQE08, UD03, VZGE07, VS06, Xin03]. originality [Lis08b]. orthodox [JJOA07]. other [CWP09, Dic07b, Haz08]. Our [RD00, TKB01, AD08, BMP06, Dei07, GY03, GS03, Imp05b, Imp06b, KTH^{+}00, McC04c, PN08, Pet09, Wal07b, YPdMA^{+}08]. Out-of-the-box [Ast09a]. outcome [Imp07]. outcome-based [Imp07]. Outcomes [CCMC05, YLRZ07, ZRY06, Cas03, FMK^{+}00, MN01]. Outcomes-based [CCMC05, YLRZ07, ZRY06]. Outgoing [McC04b]. outlook [Yan08]. output [MDGM05, OB07]. outreach [BC07d, BCLW08, DK06, LV08, LR08, PPJH09, ZLOB07]. outsourcing [AMVZ06, FKMR04, HP07, HFHK05]. outstanding [Pop05a]. OVAL [Ran00]. overcome [CM01]. overcoming [JS08a, WPS01]. overflow [PT03]. overkill [Ast01]. overloading [BS09]. overlooked [Yan07]. overriding [PP07]. oversight [WP06b]. Overview [PB06, SMS^{+}06, Gor04b, Imp05a, Rob01b, SCC^{+}04]. owners [GPP00]. oxygen [MD04].

p [Cle09b]. pace [Kül09]. pack [Rad08]. package [Fon01, GL08a, GL09, KW01, Röß06, Wu05]. packages [KH00]. PACT [CCC07]. Pages [DUK02]. pain [COR07a]. Pair [JS08a, BW05, BEW08, GDHW06, HMDK04, Han05a, Han06, KWW^{+}04, MWBF02, MHW03, MAFLR05, MAFLR06, NWF^{+}03, TRR03, Van04, WU01, Wil07]. pair-programming [BW05, BEW08, MWBF02, MAFLR05, MAFLR06, WU01]. Paired [PSS09]. Pale [Laz02b, Laz02a]. palette [Ste06]. palindrome [Sie00]. Pan [Dav01a]. Pander [Ast09b]. Panel [DDRG04, DHGR04, HBIG03, MWF^{+}02, UHFT09, ACK^{+}00, FMK^{+}00, HAB^{+}00, Joy01, KCF^{+}00, KDH^{+}00, KDM^{+}00, LKGY00, LSMB00, MDH^{+}00, MPQB^{+}00, MFC00, PCR^{+}01, RWGM00, RCE^{+}00, SHN^{+}00, SHL^{+}00, Sut00, TBS^{+}00]. panelists [BRB^{+}01]. pantomime [PY05]. pantomime-based [PY05]. paper [BG08b, Fon09, SPM00, Ter00]. paper-based [SPM00]. paper-slide-show [Ter00]. papers [WMM02]. parables [Dic07a]. paradigm [Bow08c, CF04b, GEVZ09a, MH06, Rob07b, Sta06, SH07, VZGE07]. paradigms [Cle01b, KEK^{+}07, VAFM03]. Parallel [ANS00, Gra07, JSJ08, Kes04, MK00, NLMN06, Ort03, PKB06, PJ01, Rol08b]. parallelism [CLMW09, EWH^{+}09, Rol09a, Rol09b]. parameter [SK02]. parameterized [HSB09]. parameters [JY01b]. parking [Wir08]. parsers [KS00]. parsing [GOGPGP08, NRW05, Pij00, Sax03, Tay02a]. part [ASA08a, ASA08b, CT09, Dal00, GY00b, Mar02, Mar03, McC05b, McC05a, McC06c, McC06a, McC06d, McC06b, MBD04, CHMG01, SJ05, SJ06]. participants [WLCL08]. participation [AH08b, BGD03a, BKW05, Cla00, CL02, DFS03, DLM09, FHBD06, HRB^{+}09, PHS^{+}07, RI08, SSS^{+}09]. Participatory [WS02]. partners [Pat06]. partnership [LL06a]. Partonomy [TH06]. Pascal [NC07]. PASS [JY01b]. passing [Fos05, PKB06, SK02]. passion [GCD^{+}09, MRGS08, Pau08]. passive [TLH^{+}09]. past [Dal00, Haw07, Laz08]. paste [NW06]. patching [Gin07d]. path [SDE01]. pathway [KR08]. pathways [ABO^{+}05]. patient [SA07]. Pattern [MGH07, SP03, AGE06, Ber01, Bow08a, Gin09c, Hab02, Ham05, MLH05, MHA04, Nef04, NKI^{+}09, Ort08]. pattern-based [Hab02, Ham05, MLH05, MHA04]. Pattern-oriented [MGH07]. Patterns [DH05, JSJ08, NW03, AEH05, ACD07, CCC07, Cha00b, Chr04, DFGM08, Dew05, FU02, Ges07, Gin04a, GMJDA09, Ham02b, Ham04b, Han04, LBB^{+}01, MKB05, Mer09a, NW00, NW01, NW02, NRW04, NRW05, ÓT04, PH04b, PHN06, PHC09, PPP06, Pro00, Ree04, SF04, TWS06, Wag05, Wei05a, Wic01, WP02, Wic05]. pay [Kac01b, Kac01a]. PB&J [DR07]. PBL [CB04, Mar09b]. PC [AAV^{+}03, AAS^{+}04, AS02, CTP09, GEVZ09b, Gol01, KHPC08, MDGM05, Par06a, SAHS04, SJS^{+}07]. PC-based [CTP09, Gol01]. PCs [TEP06]. PDA [Gui05]. PDF [Soj03a, Soj03b]. PDP [HB01, She03a]. PDP-11 [HB01]. PDP-11/10 [HB01]. PDP-8 [She03a]. PE [OTUW07a]. pedagogic [ACS02, Jon05, Lis09a, MMP05, NM04]. Pedagogical [FU02, RRP00, RA06, ALR08, Bow08a, CCC07, Gri00, HAFT09, Kop07, ML04, MHA04, Ras00, Ras03, RTET03, RN02, SL00b, SMH05]. Pedagogically [Gre00a, JSJ08, PK08]. pedagogy [AAC^{+}05, BN00, Bow06, DD08, HCJ^{+}08, HAM00, Haw09a, KP04, LH05, Lew00, MKB05, PW08]. Peer [Chi05, Doo09, KMN^{+}03, Pap03c, PA06, BFD^{+}09, BYS09, Cha03a, CO00, DKL^{+}07, Geh01, GCPQA05, HRB^{+}09, HLCS08, SJ04, Son09, Try05, WLCL08]. peer-led [BYS09, HRB^{+}09, HLCS08]. Peer-to-peer [KMN^{+}03, Cha03a]. peers [Son09]. PeerWise [DLRH08]. pen [ABL07, BJJ01, BBW04, CS08]. pen-based [BJJ01, BBW04, CS08]. Penny [Hen04b]. People [Tow02, NL07, NL08]. Pep8CPU [WO07]. Perceived [LBA08]. perception [BGSL06, Cox05, HL04a, PHW07]. perceptions [BBY08, Edm08b, Kac03, Lew07b, MGS07, SL08a, SDM^{+}02, SLPG07, TH06, Van04]. Percolations [Kac08a, Kac09a, Kac09b]. perfect [BGL^{+}07]. perfected [Got00a]. perfectly [Sch07]. performamatics [HGK09]. performance [ABAA09, ACL^{+}03, Cha06b, CMD08, CSM07b, CMS07, Cle02b, Cle08b, Cle09b, DSSP06, Deb04, GR00a, Gov07, Han05a, HALW01, Hau05, HL04b, Imp09, JP03, Kur08b, LCPQR05, MDM04, MWBF02, PR02, PCW09, PJ05, RRKS06, Sok00, TRWJ02, WR07, YA05]. performed [MPJC09]. performing [EA07, RF00c]. permutations [Gin01a]. permuter [Ste06]. persist [BMK09, Kac07a, SBLH07]. Persistence [PH04a, CWC09, Pow08]. person [ES05a, Mar09a]. personal [Cha04a, Mar05d, MC08, Sur01]. personalised [YH03]. Personality [LCW06, GSyC07, Rut01]. personalized [AKA09, FCA05, HSB09, RBC07]. Personalizing [SKO^{+}09]. perspective [AHLY03, AHJ^{+}07, AH08b, Arm08, BO01, Chr04, Chr05a, CK08, CM06b, DBK^{+}09, FPA^{+}06, GEHR09, Gin02a, His09, KTH^{+}00, Kla05a, KA09, LRPQ04, LS05, OHRL01, PK06, PV08, San08, Sha00, SMR03, SB06, SPW^{+}08, Vat00]. Perspectives [AAEI08, FLK^{+}09, PM02b, Gin02c, LBC^{+}06]. Peru [ABS08, GS09]. Pervasive [Bur04, PWC^{+}05]. Perverse [Rol08a]. Petri [Bar01, Bar02b]. Ph.D [Bal00]. phase [AL01, DGMW07, EA00]. PhD [Sae09]. phenomena [LL00]. phenomenographic [DTP03]. phenomenography [Lis03b]. philosophers [Rob01a]. philosophy [Had05, TS00]. phish [RR06]. phobes [TRR03]. phobic [KTH^{+}00, TKB01]. phone [Dra09]. phones [LDK^{+}07]. Photon [YLS05]. phylogenetic [MHW07]. Pi [Pop08a, JJ05, Pop09a]. PIC [SC04, SC07]. PIC-based [SC04]. Picture [LD01, Wal08a]. pictures [Lew07b]. pieces [PD03]. piles [Hen04b]. Pilot [Bou03, DLM07, BBC06, CKMP09, LNCC09, SJS^{+}07, BGKT00a]. pintos [PRB09]. Pioneering [Gür02a]. pipe [Rob02]. pipe-fork [Rob02]. pipeline [ABB03a, BBE^{+}09, Cam02a, EH05, JP02, RPR03]. pitfalls [CL00]. Piuh [BT04, BL06a]. PKAL [Mar01b]. PL [DWJ04, DJWD05]. PL-detective [DWJ04, DJWD05]. place [FU02, Hit05, LS06]. Placement [Gin00b, ACK^{+}00, AWS^{+}09, MR01, OHRL01, SSP^{+}05c]. Plagiarism [ELC07, Car00, CCH^{+}03c, DH05, Gib09, Got06b, LC01, NW06, SDM^{+}02, SGKL^{+}01]. plagiarized [JWC07, JWC08]. plain [NSW08]. Plan [SYLL00, FB02, San09a]. Planning [Ran07b, KPCM06, LV03, OAB07]. plans [CFL^{+}04, ES06]. platform [AO06b, CBKA09, Con08, Gri02c, HLM01, MW09, OG05, Par02a, SPAC^{+}05, THH09, Xu08]. platforms [NV05, Sok00]. Platonic [Par03b]. play [ACK08, Eps01, JDGAGMGC05, Sch07]. player [BHH^{+}08, Shu06b, Ter05]. playing [AL02, BM07a, BSA^{+}05, CL08a, Pop09b, TMS07, WBBC09]. PlayToLearn [BL06a]. plug [Kag09, NBL09]. plug-in [Kag09]. plug-ins [NBL09]. PNG [Ste06]. pocket [MG00]. Podcasting [Mal07]. podcasts [SRGF08]. point [Lew02]. pointer [ABH^{+}07, AEPQ09, War05]. pointers [Kum01, Kum09a, PWH00, Rol03a]. points [Pij00, Wal09b]. poised [McC01a]. Poker [SS06]. policies [GL03b, KCF^{+}00]. policy [RMPD09, SBG^{+}08]. political [Cle04b]. polygon [Lew02]. Polymorphic [BRB^{+}01]. polymorphism [BWC^{+}05, Han04]. ponder [Ast09b]. Poogle [HW08]. Pool [KSE02]. pools [Sha04]. poor [CHMG01, VS07a]. pop [Cic09]. popular [GT00, Xu08]. populations [AL09]. populous [GG06b]. portable [CS00a]. portal [RW07]. portals [RC04a]. portfolio [Est01, Sam05]. Portland [DRFG08]. PortOS [AS02]. Portugal [CFPL05]. positioned [LD04]. positions [DHGR04]. positive [PMC^{+}04, SK09]. positivists [Lis05b]. possible [KT07, PB05a]. Post [AS02, JY01a, LD04, Ste04]. post-classical [Ste04]. post-genomic [LD04]. Post-PC [AS02]. poster [ACL^{+}01, And00, BP00, Bar01, BL00, BPK^{+}01, CJ01a, Cas01, Cha00b, CC01, CS00b, CY01, CCK01, DWT00, Del00, Eng01, Ere00, Fow01, GR00b, Gri01, Hag00, HY01, Imp01, Jen01a, JY01a, JY01b, Kac01a, KV00, Kno01, KPT00, Kop01, Kos00, KB01, LR01, Law01, Lax01, LD01, Man00, MA01, MTHS00, OCM01, Ode00, Put01, Ras01, RF00c, SD00, Sca00, SPM00, SL00a, SL00b, SSC01, SYLL00, Soo00, Soo01a, Str00, Ter00, Tho01, Tho00, VT01, VI01, VKV01, WPS01, Xiz01]. postgraduate [DSM01, Joy02, JY04, Joy06a]. postroom [Osb02]. postscript [HT00, Ruc07]. Potential [Sar08, CBKM07, GS03, LDK^{+}07]. potpourri [ABB^{+}06]. power [KR06, RRP00, RRP01, War05]. powerful [Lap02]. Powers [Cun00]. practica [CJ01b]. Practical [GNPL01, SR07, SRGF08, Gol01, Kes04, PTT06, Qui06, TYT^{+}09, TV08]. practically [ML04]. Practice [Rou06, BED^{+}03, BR09, Doo03, DDC^{+}03b, FK08, Gib09, HTW04, MA01, Mik05, Nod03, PM06, SD03b, TP03, UCJ^{+}04, Vin05, YLYX06]. practices [BUCP02, BAET09, BKR^{+}05, Cle05b, Con08, CCM08, Dei07, DGHC03, HWRC08, HY01, LW06, PTT08, SDM^{+}02]. practicing [Bou09]. practicum [GP02, HL04b]. Practitioner [Cle00b]. practitioners [AKWZ07]. pragmatic [Cox05]. praise [Pap03a]. pratfalls [CHB06]. Pre [Gig03, Hen07a, Jin08, BDMY05, CKMP09, cL05, NP02a]. pre-assessment [BDMY05]. Pre-college [Hen07a, CKMP09, NP02a]. Pre-games [Gig03]. Pre-programming [Jin08]. pre-school [cL05]. preach [Rou06]. predict [RRRH04]. predicting [BDMY05]. Prediction [CSM^{+}07a, JFH00]. predictions [Cha06b]. Predictive [TLH^{+}09]. predictor [LC06]. Predictors [JJK03, RRR02, RVWB04, RBJ^{+}09]. prefer [CM08a]. preference [Sca00]. Prefuse [EVS07]. Preliminary [SAHS04, BBF^{+}09, BBN^{+}04, CPT02, DKL^{+}07, GCW^{+}03, GLD^{+}05, Gri03, Sca00]. Prelude [Soo01a]. premier [Kno00a]. Preparation [GEHR09, ELE^{+}02, EPM09]. prepare [PBWS06, PB06]. preparedness [MN01]. Preparing [EWH^{+}09, HTBS01, HDH06a, HMS08, Las03, LFGY02, Mar05c, Can06, ES08, MW09]. prescriptive [OG06]. presence [CMD08, TRT05]. Present [YSLL00, AHPS09, CIK^{+}03, Laz08, Wal07b]. presentation [AAV^{+}03, AAS^{+}04, Haz01a, Haz03a, MMRI07]. presentations [BU09, DO07, DAH08, Kno01, TEP06, TRT05]. presenter [GS06b, WGS05]. presenters [RKDM01]. press [GT00]. pressure [Bib06]. prevention [SA07]. preview [Kno00a]. primal [PP03b]. primary [Mus06, PPRV06]. primer [CLP^{+}01, SV00]. Priming [JP02]. primitives [AF03a]. principle [Mar09a, Ada06]. principled [Gri08]. Principles [RJM^{+}05, BH04, Cor07c, Den04, DDRG04, HTW04, MR02]. printed [Dic07a]. prior [RSK06, SV03]. priori [CSM^{+}07a]. priority [BGSL06]. private [LL02]. probability [And07, GAGR05]. Problem [CSM07b, PW08, ASPS04, Ars09, BLH01, BA05b, Ber03a, BC06b, BN00, CWHH05, CWHL06, CHP^{+}05, Cun02, DWJ02, ESP03, FP09, GGAB01, Gin01c, Gin01d, Gin02c, Got08, GMHS01, HB09a, Hen06a, HSB09, KB09, Kum08a, cL05, cLwW07, LV05, LW00, MD04, Min07, Min09, MBD04, Mir09, MHA04, MH05, MGH07, ÓT04, OG06, PFF04, PP03b, Raf08, RP04, Rol09a, Rol09b, RM09, RS08, SDE01, SW06a, Sok00, VI01, WZ00, Zil05, dRTW04]. problem-based [BC06b, ESP03, ÓT04, OG06, Zil05]. problem-solving [BN00, Kum08a, MHA04, MH05, SDE01]. problem-tracing [HSB09]. Problems [Gri02a, Soo01b, Utt06, Bha00, BCC^{+}07, Cas01, EGB07, FM09, GELS04a, Haz01a, HILS05, KB08, KB09, Kum00, RBlR01]. procedural [VZGE07]. Proceeding [ACM01]. proceedings [CDWH00, GWG02, LB08b, RHDR07, Val04, ACM06, ALRY08, Ano03, BTHR06, BCK04, CJGU02, CFPL05, DSF^{+}03, DNTB05, DGS06, DRFG08, Ell00b, FKCM01, Fin01, Fin02, Fin03, GKJD03, HPT07, Imp02b, Imp04, JKDN04, WMGR01]. Process [WLCL08, AF03a, BC05, BT05, Boi09, BPD03, CHDH02, Car05b, Con04, Fil01, GGH^{+}08, Goo03, GGR03, GF05, Hil00, JKM07, KB09, KJGW08, LC01, Lea08, MK00, NvdH04, OCM01, OTUW07a, RTET03, Rob02, Tho00, VVV04, Wag05, FPA^{+}06, Por09]. processes [Aha00, KD07, MS03, PB09a, PM02a, Van04]. Processing [Bur03, FMR^{+}06, Hun03, KMSB08, MV05, Rol05, Sha04, SFPFV08, VKV01, WO07, WN05]. processors [AMUFVI09, Ste03]. produced [Bur08]. producing [KK03]. Production [ACR03, DKG06, HD04, vLK09]. productively [Kac07a]. products [BC04d]. Professional [Haz05, BHM06a, BAAM06, Cle01a, FK08, FLK^{+}09, Got09a, HL04a, Jon02, JY04, KV00, MSS^{+}04, MC08, PSS09, Pre07, RC04b, TMSZ07]. professionalism [BHM06a, Got02, Got05, Got06a]. professionally [Got08, Got09a, Got09b]. professionals [CV08, CPT02, Got00a, Joy06a, Mar07b, Wal05d]. professor [GEVZ09b, LRKM07]. profile [LZ07]. profiles [BL00, YH03]. profit [LFL06]. Proglets [Edm09]. prognosis [BKR^{+}05]. Program [HMDK04, Han05b, MMBAS04, Mor08, Rol03a, SM03, Ter00, ALR04, Abu04, ASA00, BLA^{+}09e, BC07d, BCLW08, BDMY05, Boi04, Bro08, BGSL06, Cam06b, CHB06, CRPZ06, Cle09a, CH09, CT09, DK06, Eth04, Fri04, GHBG^{+}03a, GHBG^{+}03b, Geo02, Gin04c, GA06, Gor05, Gri02c, HM00, HCMM00, HE06, HHS02, JKM07, KvLA09, LJMV07, LV03, LMS00, LD01, LHL03, LM05, LLB02, MR00, MK02, Per09, PPJH09, PG06, Pul06, RLW04, RRP02, RF00a, Red01, Ric09, SDE01, SM02c, Sut00, TW09a, Ter05, Tho01, Tru05, TBR05, TW09b, UFVI05, VKV01, YLRZ07, Yeh03, dMBG^{+}08, FSAF03, Joh02]. Programmable [JBMP03, JKKL04, WO07]. Programmatic [WB02]. programmer [BO01, DSE02, ET05, Gin03c, KYC07, LCS^{+}09, RBJ^{+}09, VS07a, YK07]. programmers [BB00, Cle09a, FNB^{+}09, Gou09, HB09a, Kan07, KWW^{+}04, LAMJ05, Lis01a, LAF^{+}04, LST^{+}06, MFRW07, MFR^{+}08, MFR^{+}09, NLMN06, Pil03, PJ01, SDE01, SCL^{+}08, Soo09]. ’Programming [Sta06, AKWZ07, AT01, BC04a, BR05a, Cle01a, Fle00, Hei03, HGH^{+}01, HMRM03, JT04, Kum04, LDB^{+}03a, MDS04, MPK^{+}08a, MFL01, Pro00, Ros02, Sco03, Shu06b, WK07, ABB^{+}06, ABAA09, ABAE^{+}09, ACR03, AL07a, ABB^{+}03b, ADP01, AHL04, Ari00, Ars09, Ast04, BC07a, Bai05, BC07b, BBF^{+}09, BCE06, BUCP02, Bar01, Bar02b, BED^{+}03, Bec03b, BCHS00, BZ05, BC05, BC06a, BC07c, BKM^{+}01, BCD^{+}04, BW05, BLOM06, BVPE06, BH04, BLA^{+}09e, Boi09, BHN^{+}09, BBE06, BN00, Bou03, Bow08c, BDF^{+}08, BEW08, BDM01b, BHLW02, Bur08, BB03b, BS06b, BL01c, CG05, CHMG01, Car02a, CWHH05, CS00a, CFJ^{+}02, CJ01a, CJ02, CLB07, CBP03, CB00, CJ01b, Cha06a, CM05b, CMS06b, CNCL09, CC02, CCG05, CTR^{+}03, CKLW05, CFKS02, Cox05, CBH09, DW04, DCP00, DLPQ04, DeP06, DWJ02, DWJ04, DL06]. programming [DMKN02, ES06, Edm07a, Edm07b, Edm09, EBC^{+}08, EPQ08, EDES05, Ell00a, ET02, Eng01, Eng02, Est01, Fei07, FK04, FS09, Fle07, FL05, For04a, For04b, Gar04, GGÁC03, GDHW06, Gel00, Gib00, Gig03, GL07a, GL07b, GSW08, Gou06, Gov07, Gra07, Gre00b, GRLDPG^{+}01, Gri02b, GA03, HM00, HMDK04, Han05a, Han06, HE07, Han09, Har02, Hel07b, HPC^{+}03, HPCZ04, Hic04, His09, HR03, HH05a, Hov05, How05, HP06, Hui01, IS09, Jac08, Jac00b, JS08a, JSJ08, JS08b, Jen01b, JWC08, Jin08, KK09, Kag09, KN09, KBC01, KH05b, Köl08b, KS04, KK03, Kum05b, Kum05c, LBB^{+}01, Las09, LNCC09, Law01, LDB^{+}03b, Les04, LE07, LLP^{+}09, LS08a, LL03, LFT09, Liu08, LR05, Lom07, Lor03, LC06, LSGB07, LKPC04, LCL08, Man09, Mar05a, MAMC05, MS05a, MUPVPF^{+}02]. programming [Mau02, MAD^{+}01a, MAD^{+}01b, MWBF02, MHW03, MAR07d, MGH^{+}06, Mea06, MDGM05, MAFLR05, MAFLR06, MS07, MSK09, Mos06, MMW^{+}05, MRM^{+}06, MPK08b, MKLH09, NWF^{+}03, NLCvL09, Nod03, OG06, Ort03, PB05b, PRK05, PKR06, PHC09, PL04a, PSM^{+}07, PBOP07, Pet09, Phi04, Pil03, Pil04, PJ05, PJ01, PMG^{+}04, PGC^{+}06, PP04, Pro09, Raj05, RSK06, RC00, RHKS01, RM01, RBlR01, RMT09, Rol07, Ros09, RH08b, Sae09, SMK01, SD03a, SL08a, SBE^{+}08, SD00, SL09b, Sav05, SM02a, Sei09, Sha05, SK02, She02, Sie06, SJ04, SDR09, SHP^{+}06, SBM05, SM05a, SMK09, SW06b, SGKL^{+}01, SH07, SSR07, SPW^{+}08, TIKT04, TRWJ02, TRR03, TRT04, Top08a, TN02, Utt06, Van04, VAFM03, VW00a, VI05, VCJ09, Wal04a, WW00, WS01a, Wic07, Wic09, WU01, Wil07, Xin03, YA05]. programming [ZJ03, Zil05, dR04, vTNC08]. Programming-lite [AKWZ07]. ProgrammingLand [HSD05, HSM06]. programs [BGD03a, BCB^{+}08, BHN^{+}09, BM07b, Bur08, CL00, CL02, Coh07a, CWC09, CF04b, DCP^{+}09, Deb04, Edm08a, Eng04, ER09, FH05, GOB^{+}07, Gol08, Gor03a, GGS^{+}06b, GLD^{+}05, Hel07b, HS02, Imp07, Imp09, JWC07, KvdHR05a, Kle02, KDM^{+}00, LdLM^{+}08, Lau07, MIS^{+}07, MPQB^{+}00, MK01, ND08, OHZ01, Ode00, PSH^{+}07, RBA05, RPR03, Ran07b, ST07, SSA08, TP03, TETPQ08, Top08a, Top09a, Tur01, TZ01, TSC^{+}02, War01]. progress [EGB07, KPN02, YYLH06]. Progressive [Djo09, KV00]. Project [MC04, WD03, Aba08, ALR04, Abu04, AC08, AAO^{+}08, AO06a, AL01, BGD03b, BAK^{+}09, BA08, BAET09, Bem02, BP03, BDA03, BFMC05, Bla09a, CHDH02, CSD^{+}07, CCH^{+}03b, CTP^{+}04, COR07a, Cle09b, Cut07, CFH^{+}01, DGHC03, DAJDRG08, Don01, Don08, Ell06b, EG08, FH01, FPQ^{+}08, GGÁC03, GDM05, GD05, Gou09, HBH07, HY01, HH06, IMRJ09, Imp05a, JY01b, KPCM06, Lab09, LDA^{+}00, Lax02, Lee02b, LBC04, LM06, Lif04a, Lin07, Liu06, MK06, MG00, Mat06, MWM01, MH00, NM02, OAB07, PDS07, PPRV06, PB06, Pre01, Qua06, Ram03, RHY07, RF00a, Red01, RW07, Rob00a, Sae09, SS06, SCC^{+}04, Sha02, Sie00, Ste03, SM08, SJS^{+}07, Top07b, Whi03, Wol01, WP07, YH05, ZSW04]. project-based [DAJDRG08, Qua06]. Project-centered [WD03]. Projects [Kum05c, PH04a, AL05a, AJF07, Bar05, Bec08, BBMW03, Ber09b, BA07b, BKS^{+}04, BPD03, CHK03a, CHK05, COR07a, CFW01, Cle02b, Cle07a, CHP^{+}05, Con04, CHS05, CSP09, Cun05, DeP06, Dem05, Dic00, DWJ02, Djo08, EGMS08, FH01, FLLM06, Fin04, Fin05, FAL^{+}05, GDHW06, Gha07, Goo03, GKP06, HY09, HE07, Han09, HPM^{+}08, HL08, How02, Hua03, Jam05, Jan01, JBB^{+}03, KBC01, KvD06, KvdMvDvdV00, KKR06, Kur08a, LFL06, LR04, LSGB07, MWG08, OBLH03, PBOP07, PM02a, Pra05a, RN03, RMY08, Ruc07, RMN06, RRD^{+}09, Rut01, SC04, Sha04, SR07, SHL^{+}00, TM04, WN05, WL00]. Prolog [Fle07, Fow01, Moh09, WD05, WS06]. promotes [LR05, Qua06]. Promoting [MHR06, Qua04, CGM04, Coh07b, HLL03, MA01, RH08a, SSS^{+}09]. proof [Bla09d, Ens09, GA06, She08, SJA04]. proof-of-concept [SJA04]. ProofChecker [SBR^{+}07]. proofs [Arm06, Haz03a, Hen09a, SBR^{+}07]. properties [WSP04]. property [Kay00]. proposal [BdMG08, BT06, MMP05, MKB05, MICM08, ND08, RS03]. proposals [Bar02b]. Proposed [LKGY00, DRKG02, Gor00a, MVD01, PH04b, RBlR01]. propositional [Arn07, LLB02]. prospective [LH03, RH08a]. prospects [MUPVPF^{+}02]. protection [SA07]. protocol [Tur03]. protocols [AA04, Chu03, CGI05, DEJ^{+}06, DJLA07, Ric00, Str00]. provide [CRPZ06, Ekm06, GHBG^{+}03b]. Providers [GL03b]. Providing [GL05, Goo03, PH04c, Röß08, BH07, HCB04, SM08]. ProxmapSearch [SJ05, SJ06]. ProxmapSort [SJ05, SJ06]. proxy [DDL03]. pruning [Mas05]. pseudocode [AC07]. Psychology [Vat00]. public [MKC09, SDG^{+}08, Wal09a]. Publish [Lor00]. Publishing [TFIJ08, CF04a, CF06, GRLDPG^{+}01]. puffy [Kac08b]. pumping [GET05]. pupils [MPH09]. Puppy [Mar00a]. purposes [DCT06, VT01, pYB08]. pursue [BAB^{+}08]. push [GOMSJVGP08]. push-down [GOMSJVGP08]. Pushdown [McD02]. Putting [EMM^{+}06a, Sol07b, TP03, Chr04]. Puzzles [HRBC03, Ros02, LP02, Lev05, Rao06, Sna06a]. Puzzles-First [Sna06a]. PyLighter [BC09]. pyruvate [CSM^{+}07a]. Python [BMK03, EPM09, GL07b, GL08a, GL08b, HRS06, Rad06, Rad08, RMZG06, SL08a, Sha03].

qualifications [DS08, McG09]. Qualitative [HDE^{+}06, Med01, MLM^{+}08, PK06]. Quality [DDLFS01, CTBS06, GSW08, HMDK04, HT00, How03, LdLM^{+}08, MS03, PG07, PC08, Sur02]. quantifying [MNBK03]. quantitative [Arm08, FH05, NBF^{+}08]. Queens [Gin05a, Rol08b, Rol06b]. Queensland [CL02]. query [AYWM08, Hol08]. querying [BA08]. Question [RLM08, PBDWL09, WP06a]. questionnaire [MH08]. questions [BS05, Cle09a, Gov07, HSB09, VT01]. queue [She04b]. queueing [Arn07]. QuickDraw [STS09]. quickest [Dic07b]. QuickTime [And00]. quirky [MLM^{+}08]. Quiver [EFK04]. quizzes [Cic09, WM00]. quo [McC01a, WBGM02].

race [Got00b, MH00]. Radford [COR07a, CTP^{+}04, FLL^{+}01]. radical [Reg00]. radio [BA07b, Sko07]. railroad [McC05c]. raise [GHG^{+}08]. Raising [ATNT04, Joy06b]. Randolph [Lis07a]. random [Bra05, JDFV08, Nau06, SM00b]. randomization [KR06]. Randomness [GAGR05]. ranking [AKT07]. Rapid [Soj03c, LHS09]. RAPTOR [CWHH05]. rarely [Lee02d]. raster [ACSS06]. raster-level [ACSS06]. rate [BM05, CTBS06]. rated [FG07]. rates [BC07c]. ratings [WMM02].

Rational [Shu05]. rationale [CP00, Pow03]. rationalism [Gre00a]. RCX [Wol01]. RCXLisp [Kla04]. Re [DMRS06, ISC^{+}09]. Re-centering [DMRS06]. reaching [OM08]. reactive [HM05]. read [Dic07a, Zel00]. readers [FdMFPT07]. readiness [GM07]. Reading [CMRR04, Cas01, Fon09, Wal07a, GBO08, LAF^{+}04, Sud08]. readings [Eis03]. Readn [Boo07]. ready [BC04d, DW07, HHRA06]. ready-made [DW07]. Real [BN00, Edm08a, HLNT01, Imp01, RF00a, ATNT04, ALH07, BA08, CMGR07, Cli01, DEJ^{+}06, DJLA07, DDL03, EA07, Got08, GKP06, JG08, Kaw07, KW07, La 06, McC05c, QLT09, RBG02, SW06b, Wal05a, WSS05, ZB08]. Real-time [HLNT01, ATNT04, Cli01, EA07, KW07, McC05c, QLT09, RBG02, Wal05a]. Real-world [BN00, RF00a, ALH07, BA08, GKP06, SW06b]. Realism [DSSP06]. realistic [GD05, KvD06]. Reality [Gri01, Dic07c, Har04, MGS07, Xu07, ZE01]. really [BLW01, Cha03b, DSE02, ESP03, Fit09, HAG05, NBF^{+}08, SMH05, VI00]. realm [BT04, BL06a]. reasoning [Gin07c, Hen03a, Hen03c, Lau07, LHS09, SHW^{+}09, Yan07]. rebuilding [Kum09d]. recognition [Ber09b]. recognize [HA02]. Recognizing [Pop05a]. Recommendations [ACK^{+}00, MBPB03, MMLT09, RBC07, RCD^{+}02]. reconfigurable [VS07b]. reconnecting [Lon08]. reconsidered [Lev00]. Reconstructed [CTBS06]. reconstruction [Gin03a]. record [Gib02]. recorded [LRKM07]. recording [DAHA09, TL01, ZH02]. recruit [BBN07]. Recruiting [Coh02b, GOB^{+}07, MKC09, OTUW07a, Tho06, Wal08b]. Recruitment [CA02, DBR^{+}08]. recurrences [BCC^{+}07]. recurring [BCC^{+}07]. Recursion [VI00, BDM05, DCP01, Geo00, Gin04b, GSG03, GMS09, HA02, Man09, Man02, Min08, Par02b, Ras01, Rol01, Ros05, RSHL07, RSUFPF08, RSVI09, SGG06, SG07, Soo00, Soo01b, Ste09b, VIPCUF08, VIPCUF09, Wir08, Yan08]. recursive [LL00, LL02, Man05, Mir09, Sax03, SN02]. Recursively [LL00]. redesign [HPCZ04]. Redesigning [HPC^{+}03, MDS04]. Rediscovering [GCD^{+}09, MRGS08]. Reducing [AL07b, Haz02, Par06b, PW01]. Reduction [AGE06, IS09]. reductionist [WS01a]. Reductive [Arm04, AGE05, AGEH06, Arm08]. redundancy [Chu01]. Redundant [Qua07]. redux [Dor07]. reengineer [LBK08]. reentry [HS02]. Reevaluating [Lau08]. Refactoring [Wic03, Cle05b, SSS06, SSS07]. referee [WMM02]. referenced [FCA05]. referencing [LL03]. refinement [BGG^{+}00, Mér09b]. refinement-based [Mér09b]. refinements [Imp09]. Refining [PHMS08, Gin01e, Lau08]. reflection [BHY09, Edw04, FKKW00]. reflection-in-action [Edw04]. Reflections [BAAM06, Cam07, Dal00, GE07, Haz08, Hen07b, HD07, Kum08b, Kum09c, Kum09d, Wil05, DVE06, Mar05d, Par07b, Tur01]. reflective [SM07]. refreshing [Rag04]. RegeXeX [BH07]. registration [AR08]. regular [AOMC07, BH07, GET06, Min09, MR05]. regularities [Gin00a]. reinforce [FdMFPT07, Soo00, SWR05]. reinforcing [Cox05]. Reinvigorating [Our04]. related [EH07, HAuR04, Poh05, RPH^{+}03, Sur02]. Relating [WI09, Cas01]. relational [Ber03b, MG00, UD03, Wan01]. relations [Wal09a, WPS01]. Relationship [CMS06b, Ber03b, GCF04, GCFP04, Haz03c, LFT09]. relationships [Cle07b, CHUB08, Goe08, SSP^{+}05a]. relative [CCMS05]. RELAX [MG00]. relevance [ARVV06, Gao00, Lis03b, OM08, PBM07]. relevant [BKS^{+}04, BNS08, Deb02, Dem05]. reliable [Lew02]. religion [Wil00]. religious [Lis04b]. remedies [KT07]. remember [Wil04a]. Remote [AM07, Rob00c, YH04, Bor07, DEJ^{+}06, DJLA07, Rob03a, RRD^{+}09, TL01]. Removing [Mar05d, Cle06b]. Renaissance [SSS^{+}09]. Reno [Ano03, GKJD03]. repetition [Mar05a]. Replace [Reg02a]. Replacing [His09, Utt06]. replay [TL01]. replicated [MAFLR06]. replication [ENO09]. Report [BBF^{+}09, CIL^{+}03, JKM07, LS04b, MMLT09, RBC^{+}05, SMS^{+}06, BvdH09, BMW04, BPT^{+}04, BT06, BBN^{+}04, Cam03a, CK02, CCD^{+}08b, CTP^{+}04, DDLR01, FKO07, Gor04b, HAJ07, KPN02, Lee02b, LHS04, Mar05b, Mar06, Mar05c, RCE^{+}00, RBC^{+}06, SM03, Sca00, SCE07, SJS^{+}07, TIKT04, Wal01b]. reporting [SB07b]. reports [BBMR03, GCF^{+}01, HAB^{+}00, NCA^{+}02]. Repositories [JA03]. repository [DLRHP09, MWG08, SFM^{+}07, TYC^{+}07]. Representation [Ric09, CMS06a, Kru03, SWR05]. representations [LZL08, SN02]. representatives [RCE^{+}00]. represented [HRB^{+}09, PB06]. republic [GG06a]. request [Dic07c, Jan01]. required [Cha06a, Cle05a, Sch02a, VR04]. Requirements [Fil01, CM06a, Cun07, Hut04, MS03, RVP08, RN02]. rescue [ACK08]. Research [Cle01b, GAL^{+}06, HHB05, HAH^{+}06, Joh02, LBC^{+}06, Lis09a, Ran07a, SW04, Way06, AGHM02, AGHC03, AHGP05, AHJ^{+}07, AH08b, AHM^{+}04, BM07a, BBN07, BABBH04, BTGD00, Ber03a, Can06, CK08, CA02, DBK^{+}09, DBR^{+}08, Dal00, Dal02, DW05a, Ell06b, ENO09, FG08, Fon09, GCF04, GCFP04, GK02, HLL03, HDE^{+}06, HAJ07, Hua03, JY01b, Kno01, KDP06, LdLM^{+}08, Lan01, Lar05, Lis03b, Lis04a, Lis07a, Lis08a, Lis09b, LCS^{+}09, MDL^{+}02, Med01, MT08, MKBW07, OBLH03, Pas05, Pea02, PD03, PSE^{+}05, PSH^{+}07, PFJC05, PWA06, Qua04, Rah08, RF09, Sch02a, Sha06, SSP^{+}05b, SGC^{+}01, TFG05, TM04, Val04, Wen04, WI09, WB06]. research-led [DW05a]. research-oriented [OBLH03]. researchers [Bri06, CY01, Ols06]. Resist [Kac07a]. resistance [Zin08]. resisting [Zin08]. resizable [Bla09d]. Resolved [ABK^{+}05]. resource [BCS02, BPD03, Cam04a, CK04, MM00, McC02, PW01]. Resources [CGY^{+}01, McC03a, Rob04c, BKM^{+}01, BAMSK05, BGGR01, Bow00, BZ04b, Cas03, CF04a, CF06, CI03, HCK07, KDH^{+}00, Kno00b, Kno02, MMD^{+}09, MICM08, PPE03, RS04, SPM00, VCC^{+}05, VCC^{+}07, WYOH02, Yai08]. respect [GM07]. Responding [GGKT04, MBPB03]. Response [Imp08, AMVZ06]. responses [Gre00a, LCS^{+}09, PBDWL09, VHF07]. responsibilities [GL03b]. responsibility [Ing09b, Med04]. responsibility-driven [Ing09b]. rest [CY01]. Restricted [MS00]. restructure [FIG07]. restructuring [Gar04]. Results [Kum05d, CPT02, CA02, Dal06, DJWD05, GG06a, LZL08, Pap03d, Sae09, SB09, Sob00, TRT04]. Resurrecting [Rob07b]. retain [BBN07]. Retaining [TATL07, Coh02b, GOB^{+}07, Tho06]. Retention [HE06, AJF07, BRP^{+}07, BBY08, BYS09, Coh07b, CCM08, CA02, DBR^{+}08, PSH^{+}07, TPK^{+}06]. Rethinking [Mun09, Ree01, BT05]. Retina [MKLH09]. retrieval [FMR^{+}06]. retrieving [BGKT00b]. retrospective [SCHS09]. returned [BBS06]. REU [RF09]. reusability [CLK00, Sma07]. Reusable [UFVI05, Vin05, Con04, CMCB09, DSCU01]. reuse [Fle01, Gib09]. reveal [Mos02]. Revealing [BC05, FHBD06]. Reverse [Gha05a, Lor03, Mar02, Mar03]. Reversing [GA06]. Review [BA00, DHRH05, Ran07a, DKL^{+}07, Geh01, GCPQA05, GC02, HHRA06, Lis04a, MUPVPF^{+}02, Try05, WLCL08, Zel00]. Reviewing [KJGW08, Son09]. reviews [BFD^{+}09, HAFT09, TQCPQE08]. Revised [HKCW09]. Revising [Haw08b, San09a, Top07a]. revision [RJM^{+}05, TWS09, Top07b]. revisited [Ham02a, How02, SGG06, SMS08]. Revisiting [YK07]. revitalized [UHFT09]. revitalizing [Ada09]. revolution [Luq05]. reward [Haz03c]. Rewrite [NCA^{+}02]. RFID [AL07b, SA07]. Rich [NBL09]. ride [Lim07]. right [CMRR04, CS08, Cle06a, DR07, GAGG04, Pat06, RJAH04, Tha07]. rigor [Gin07c, Gin07d]. rigorous [DW05b]. risks [AM05]. ’rithmeticand [Boo07]. RMI [HHJ08]. road [LDB^{+}03b, Mar08, Mar09a, Wir02]. roadmap [TS08]. robocode [Liu08, OG06]. RoboCup [HS03, VB02]. robot [Bec01, HLM01, Lar05, LNH09, XM06, Xu08]. robotics [BM07a, BMK03, CPP07, Fag00, GP02, HLM01, HM05, Kla02a, Kla04, KC07, NTTHT09, WWK^{+}05, WW09, WTTT^{+}08]. robots [ACK08, Cha05, FM03a, GCF^{+}01, Imb03, JCOP07, Kum05a, Law01, LDB^{+}03b, Les04, MGFK06, SM02a, SKO^{+}09, WB09, Wol01]. RobotStudio [Xu07]. robust [Gou06]. rock [GL03a]. rocket [BA01, Boy08]. Role [AL02, Top08b, ACK08, Arm06, BSA^{+}05, BS04, CHK03b, CL08a, CD08, Cic09, Cle05b, Cle07a, CTLW03, DKG06, GLA^{+}08, GKHPQ07, JDGAGMGC05, LOP^{+}07, Lom07, NRA^{+}03, Ola00, Top08a, Top09b, Tow02, VAFM03, Wal03a, WHCT09]. role-game [CD08]. role-play [JDGAGMGC05]. role-playing [BSA^{+}05, CL08a]. Roles [BAS04, BS06b, ABAA09, CTR^{+}03, GS06a, KS04, Kur08a, RF00a, Son09]. roller [GP00]. room [Lis06b, MV01]. root [SB07b]. ROSE [MWG08, Shu05]. Rough [BA08]. route [Ros01]. routing [AL07a, Cas02, LL06b]. RS [Jan01]. RS-232 [Jan01]. RSA [Ber09a]. rubric [AK09]. rubric-based [AK09]. Rubrics [McC03b, Bec03b, MMK02]. Ruby [Ort08, She07]. Rubyric [AKA09]. rules [Fle00, Lis09b, WSP04]. run [Mar09a]. Runestone [LDA^{+}00]. Running [HHB04, OBLH03, Ros09]. runtime [MKKC08]. Runtimes [Han05b]. safe [Fek08, Ros01]. said [Car00]. SAIL [BGKT00b]. salesman [Sok00]. same [MSP^{+}06]. sample [BNSW05]. sandwich [KV00]. satellite [YM06]. Saudi [ASA00]. Saving [LKPC04]. say [CV08, LRPQ04, PT02]. SBTa [GEVZ03]. scaffolded [Gin09c, RIP^{+}09]. Scaffolding [Cha01, TRT04]. Scalable [Lon08, GS06b, RI08, SD08]. scale [Bro09, KMSB08, LCB^{+}03, RF00a, RBM04]. scaled [Mea06]. Scaling [Pul06]. Scavenger [TPK^{+}06]. scenario [ATNT04, Vat01]. scenario-based [Vat01]. scheduling [CHP^{+}05, Rob04a]. scheme [BDRH09, MSTS02, WS06, Hic04, RSK06]. Scheme-based [Hic04, RSK06]. schemes [HAuR04, vVML04]. Scholarship [Pop03a, FKO07]. Scholarships [Pop03b]. School [CRPZ06, HL06, fCST02, ABAE^{+}09, AGE04, BZ04a, BBC06, BAB^{+}08, BC07d, BCLW08, BEE07, Cas09, CBJL08, CFKS02, CH09, DK06, DLM07, DKN08, EH05, GESZ09, GEHR09, Gin00a, GM07, HGEB08, LV08, cL05, LNCC09, MJ03, Oli05, PMC^{+}04, RH03, RRKS06, RHY07, RHL^{+}09, RIP^{+}09, SD03c, SL08b, TBAA09, TH06, Wos06]. schoolers [Ada07]. schools [Alb09, DDLR01, EH05, EH07, FG07, Sko07, SBZ09]. Science [ACM01, ACM06, ALRY08, Ano03, BTHR06, BPT^{+}04, BP02, BCK04, BZ04b, CJGU02, CDWH00, CIL^{+}03, CFPL05, DSF^{+}03, DNTB05, DGS06, DRFG08, Ell00b, FKCM01, Fin01, Fin02, Fin03, GWG02, GKJD03, HMRM03, HPT07, Imp02b, Joh02, JKDN04, MR00, Ran07a, Rob04c, RHDR07, Spo00, WMGR01, AHLY03, AEH05, ASA00, ACL^{+}03, AKWZ07, AGHM02, AGHC03, AHGP05, AHH^{+}06, AAF^{+}06, APR06, AAS^{+}04, ARVV06, ALH07, ACK^{+}00, AWS^{+}09, BGDJ02, BGD03a, BGDR05, BMK09, Bea02, Bea03, BM05, BM07a, Bec07, BBN07, BABBH04, BKM^{+}01, BKR^{+}05, BT01, BBW04, BRP^{+}03, BBY08, BAB^{+}08, Bot05, BEM^{+}07, Bou09, Bow08a, BDM^{+}07, Boy08, BBN^{+}04, BR02, BMR04, Bra05, BPS09, Bro08, BS01, BKS^{+}04, BCB09, BR09, BBT07, BDDL04, BS04, CGB08, Car05a, Car08, Car02b, Car06, Cas09, CB00, CM02, Cha06b]. science [CO00, Cha04b, CBJL08, Chr03, CR03, CFGL05, CH07, CKMP09, CKLW05, CF04b, CDP03, CCMC05, Cor07c, CS03a, CT07, CT09, CPST02, CS03b, CMT04a, CMT04b, CGT05, Cun05, CA02, Cur03, Cur07, CM08b, CS06, CBKM07, DKG06, DR07, DG05, DV04, DMRS06, DH06, DDA^{+}07, DD08, DW07, DTL^{+}02, EMM^{+}06a, Ega07, EH07, Eis03, ELE^{+}02, Eps06, Etl06, Fag00, FM03a, FKKW00, FM09, Fon09, FG07, FH05, FS02, FB02, Fri05, Fro07, FJA^{+}07, FIG07, GT00, GESZ09, GEHR09, GSyC07, GKHPQ07, GS00, Geh01, GG06a, Gol04a, Goo08, GR00b, Got00a, GD07, GL03a, Gra07, GK02, GMHS01, GHLL08, GBO08, Gür02a, GS09, Guz01a, GS03, Hal08, HHK^{+}09, Han07, HEB08, HALW01, HL07, HKCW08, Haw08b, HKCW09, Haz01a, Haz03a, Haz03c, HL04a, HL04b, HDE^{+}06, HL06]. science [HGEB08, Hel07a, HDA^{+}02, HBEM03, Hep06, HE06, HETD09, HDH06b, HAJ07, HRB^{+}09, HB09b, Hug09, IK06, Imp04, Jam02, Joe06a, Jon00, Kac03, KK02, Kay00, KvdHR05b, KH05a, KH08, Ker04, KW07, Khu08, Kla05a, Kla05b, Kle02, KH06, Kli05, KR08, Kol08a, KAPG07, KSRD05, LC08, LH03, LH05, LA08, LML^{+}01, Lax01, Laz08, LD04, Lee00b, Lee01a, Lee02a, Lee02b, LJG05, LS05, Lew07a, LZBO09, LHL03, Lis04a, LZL08, Luq05, Mal07, MHR06, MIS^{+}07, MAMC05, Mar02, Mar03, MA01, MDH^{+}00, McC00a, MM00, McC02, MCS04, MDM04, McG00, Med04, MAFLR05, MAFLR06, Mic00, MR01, MPH09, MK01, MJ03, MPJC09, MKC09, Mos02, MBE^{+}09, MT05, MMW^{+}05, MT08, Mus06, NRA^{+}03, Noy02, Ola00, Oli05, OHRL01, OM08, Pan09, Pap03a, Pap03b, Pap03d, PKR06, Par06b, PS06]. science [PBM03, Par05b, PDC01, Pas05, PH04c, PSH^{+}07, Per09, PTC^{+}01, PSS09, PB03b, PV08, PD06, Pol01, PPA08, Pow08, PM02b, PT02, PMT03, Pro00, Pul06, Rad06, Raf08, RH03, RH08a, Ral05, Ran07b, RMZG06, Ras00, RRKS06, RHY07, Reb05, RBC^{+}02, RB09, RMPD09, Rio02, Rob00e, RKI02, Rob03b, Rob04b, Rod02, Ros01, Ros09, RDCE06, RF00b, RJM^{+}08, RS03, Sam02, SM00a, SM03, SV00, SB08, SBG^{+}08, SSC00a, SLPG07, SSC00b, SHN^{+}00, Shi02, SAHS04, SF02, ST08, Sna06a, Sol07b, SSA08, SMS08, SBZ09, SHL^{+}00, SGKL^{+}01, SH07, SJS^{+}07, SB06, Sur02, TW09a, TPK^{+}06, TF07b, TAHA00, Tha07, Tho06, TBS^{+}00, TM03, THH09, TDJ^{+}03, TMD^{+}04, TYC^{+}07, TSC^{+}02, VW00a, Vau00, VCC^{+}05, VCC^{+}07, VZ06, WJDL04, Wal09a, WHCT09, Wil00, WS01b, Wir02, Wir03, WBS04]. science [WR07, Wos06, pYB08, YH05, ZLL^{+}07, ZB08, ZV04, ZIBG05, dMBG^{+}08, HKW^{+}02, LdLM^{+}08]. science-specific [FJA^{+}07]. sciences [LS02, PBWS06, PB06, Shi02, SMF09]. Scientific [FNW02, BB00, BR02, Hoo08, MC08, Ola00, RHY07, WSC03]. scientist [Mar04a, PPR^{+}02]. scientists [BBCD09, DAKLH07, DLHAK08, DRKG02, HS02, Kap04, ML07, Ros07, TT05, WS02, WAC^{+}08, vdK07]. scope [FK04, Kum00, Kum01]. scores [Jac01]. scoring [LV05]. Scotland [HPT07]. Scrambling [YPdMA^{+}08]. ’Scratch [WMP08, ML07, Bla09a, FS09, MPK^{+}08a, WLMM09]. Scratching [SL08b]. screen [FdMFPT07, Jon05]. Scribbles [Mar07c]. scripting [RF01, Sol07a]. SCS [FB02]. SE [Cli01, JS06]. Search [AKT07, ACK08, BC06b, CKW04, LRKM07, NM02, Rob04b, Rol02]. Searching [Ham04a, Lee03a, Man09, She08]. SEATS [DWT00]. Seattle [CBKA09]. Second [App01, CPP07, ACM01, AKM00, Bla09b, Can06, SHS^{+}08, UD01]. second-year [SHS^{+}08]. secondary [BPS09, DK06, ESP03, EGB05, EGB07, GM07, MV05, MJ03, Sae09, SB08, SB09]. section [Boz06]. sector [EH05]. sector-based [EH05]. secure [AL07b]. Security [JCG06, Kre09, TS08, AA04, ALO^{+}03, BBMW03, BAMSK05, Bor07, BBS06, CHK03b, Cam04a, Car04, CD08, Cle02a, DFGH04, Eps06, GV06, Gha07, HCHP01, HMS04, LS04a, Lin07, LC05, Mat03, MV01, MWF^{+}02, ND08, O’L06, PT03, SGMJ09, Spa04, Tal06, TA08, TW05, Vau00, WW04, Yan09b, YLYX06, DTW07]. see [CGB08, GAGG04, Har05, Kac08b]. SEED [DTW07]. seeing [LST^{+}06]. seeking [AAO^{+}08]. seemed [BMT^{+}07, BMT^{+}08, BMD^{+}09]. seen [BAS04, Gre00b]. selecting [LW05]. Self [Ell06a, Hoo08, Lap02, RLW04, AD08, BS05, Cla05, HY09, Kum08a, LWS07, MT08, NLCvL09, PHW07, PMC^{+}04]. self-appreciation [HY09].

Self-assessment [Lap02, BS05, NLCvL09]. self-confidence [Kum08a]. self-directed [Ell06a]. Self-efficacy [RLW04]. Self-grading [Ell06a]. self-improvement [Cla05]. Self-organized [Hoo08]. self-perception [PHW07]. self-theories [MT08]. semantic [LRKM07, RLM08, DAJDRG08]. semantics [Kum09a, PHC08]. semaphore [Ree02, SW06a]. semaphores [KBAP00, Ree04]. semester [LM06, Whi03]. semesters [OJJ00]. semi [Jac00a]. semi-automated [Jac00a]. seminar [BCL00, Kay00, KR00, LB00, MR00, Rag04, SW04, Röß08]. seminars [AMW06, LA07, LA08, Röß08]. semiotics [Fon03]. senior [Gin04b, MMW^{+}05, Sch02a, Shu06a]. seniors [BBY08]. sense [ABO^{+}05, DKL^{+}07, Mar07a, Wal01c, Wal02b]. sensor [TY06, TBM09]. Seoul [Tur09]. Separating [Gin08c]. separation [Rou06]. sequel [FNW08]. sequence [CKM07, Don01, FMR^{+}06, How05, PQEA^{+}04, Sch04, TAHA00, TRWJ02, TATL07, Wic05]. sequences [CTBS06]. sequencing [dMBM08]. Serious [Cur07, HLCS08, RM09]. Server [DUK02, Sha02]. Service [FLLM06, GL03b, Jam02, RC04a, TM03, Bro08, CRPZ06, GS06a, Kaw07, KP04, LHL03, Pop05b, Ros07, SV00, SL07a, SW04]. Service-learning [FLLM06, Bro08, LHL03]. services [AFR08, AO06b, HHJ08, Hum04, KFE^{+}07, LJM05, Str00]. serving [DWY03]. Servlets [DUK02]. SeSF [ES05b]. SeSFJava [ES05c]. session [ACL^{+}01, And00, ADM^{+}09, ACK^{+}00, BP00, Bar01, BL00, BPK^{+}01, BCL00, CJ01a, Cas01, Cha00b, CC01, CS00b, CY01, CCK01, DWT00, Del00, DDRG04, Eng01, Ere00, FMK^{+}00, Fow01, GR00b, Gri01, Hag00, HAB^{+}00, HY01, Imp01, Jen01a, JY01a, JY01b, Joy01, Kac01a, KV00, Kay00, KCF^{+}00, KDH^{+}00, KDM^{+}00, KPT00, KR00, Kop01, Kos00, KB01, LR01, Law01, Lax01, LB00, LKGY00, LSMB00, LD01, Man00, MR00, MA01, MDH^{+}00, MPQB^{+}00, MTHS00, MFC00, OCM01, Ode00, Put01, Ras01, RWGM00, RCE^{+}00, RF00c, SD00, Sca00, SPM00, SL00a, SL00b, SSC01, SHN^{+}00, SYLL00, Soo00, Soo01a, SHL^{+}00, Str00, Sut00, Ter00, Tho01, Tho00, TBS^{+}00, VT01, VI01, VKV01, WPS01, Xiz01]. sessions [CT07]. set [ABE^{+}02, DLP08, RHL^{+}09, Sco02, Han04]. setting [Cam05b, CKW04, PTT08]. Seven [LML07, WSP04, Wil07]. seventh [ACM06]. sexier [YPdMA^{+}08]. Sexism [CWC09]. sexual [God06]. shading [TF07a]. Shall [CAMF^{+}03, GY01b]. shallow [Car03]. sharable [HW08]. share [CFL^{+}04]. Shared [LL02]. Sharing [CF04a, CF06, HHLS06, EBC^{+}08, Fin09, KDH^{+}00, Kop07, Mik05, Pho01]. sharpening [Gin01e]. sheets [Lau07]. shelf [ABK^{+}01, LCB^{+}03]. shift [GEVZ09a, PW08]. shinning [Got06b]. shock [Got01, Got05]. shoestring [GV06]. shopping [SL06]. Short [LS04b, LHS04]. Shortchanging [BCS02]. shortcut [Dic07b]. should [BDM05, BGSL06, Chr03, HAH^{+}06, PP07, Sic09, SNS07, Wal04c, Wal05c]. shoulders [LSG05]. show [Ter00, WSS05]. showcase [Fin04, Fin05, Pra05a, Pre01]. showing [Gib02]. shows [CM08b]. shrinking [Cam02a, EH05, RPR03]. Si [BT04, BL06a]. side [Mar00b, MDD06, Wol02].

SIGCSE [ACM01, ACM06, ALRY08, Ano03, BTHR06, BCK04, CJGU02, CDWH00, CFPL05, DSF^{+}03, DNTB05, DGS06, DRFG08, Ell00b, FKCM01, Fin01, Fin02, GWG02, GKJD03, HPT07, Imp02b, Imp04, JKDN04, Pre01, RHDR07, WMGR01, Wal01b, ADM^{+}09, BMW04, Bru04, Bru05a, CAK09, Fin04, Fin05, FKO07, GCF04, GCFP04, Hen08b, Hen09b, HAJ07, Imp08, JKM07, Kno00b, KJGW08, LB08b, Mar05b, Mar06, RRP02, Reg02b, RHDR07, Val04, Wal01b, Wal02a, Whi09]. SIGCSE-01 [ACM01]. SIGCSE-02 [Imp02b]. SIGCSE-04 [Imp04]. SIGCSE-member [Wal02a]. SIGCSE-members [Bru04, Bru05a]. SIGCSE/ITiCSE [JKM07]. signal [ZSW04]. significant [CGM04, MF07, RHY07]. SIGPLAN [BBF^{+}09]. SIMD [Ort03]. Simple [ZF04, Ast01, BDM01b, BB01, CL00, Dic00, KW01, KR06, LRD09, Mér09b, PR04, Sci07, She04b, TLS07, WKB02]. SimpleDB [Sci07]. simpler [MP09a]. simplex [PP03b]. simplicity [DSSP06, Lee03a, May08, Rob04b]. SimSE [NvdH04, NvdH09]. simulate [ACK08]. simulated [Kur08b]. simulates [BB01]. Simulation [Sok02, AYWM08, And07, BP03, BBE06, BROB00, CW03, DSSP06, IDD06, KPT03, KM00, KMS01, MGE^{+}05, NvdH04, NRW04, OG05, PCW09, PK08, RC02, RKD02, SD09, ZSW04]. simulation-visualisation [KPT03]. simulations [HH06, Ibb02, KC07]. simulator [BN09, Dic00, Har04, MDGM05, Rob02, Rob04a, Rob05, Rob06a, Rob07a, SD08, She03a, Sto06, VS07b, VS06, WO07, WW06a, YB01]. simulators [CCY03, WYOH02]. sing [Pat06]. single [Dix04]. singular [FS06]. site [NvdH09]. sites [HBIG03]. situated [SYLL00]. sixth [DNTB05]. size [BDM^{+}07, HB09b]. SKA [HTK02]. skates [GP00]. sketch [Mar07c]. sketching [ABL07]. skill [BEW08, MGH^{+}06]. skills [BLH01, BU09, BBCD09, BMR04, Bra05, BS06b, CG02, Cle05a, CLP06, DWJ02, Esh07, FP09, Gra05, HH07, Ham04a, HL07, Hen07a, JB07, KP05, cLwW07, Lie05, LAF^{+}04, LFT09, MAD^{+}01a, MAD^{+}01b, MD05, MD06b, MC08, Mic00, Ort03, Pol01, PL04b, RH08a, Rao05, RKD02, Sol07a, SBM05, Tha07, VS07a, VMD01, VCJ09, WK07]. skin [Shu06b]. slacker [LSGB07]. slaying [Lab09]. SLic [KBC01]. slide [Ter00]. slides [Gol04b]. sliding [Gin04a, RTM^{+}04]. Small [AV02a, BPT^{+}04, Alb09, Bro09, COR07a, DCP00, Doo04, Doo07, Jon01, LS04a]. Small-college [AV02a]. smaller [BDM^{+}07]. Smart [SSC01, HHB05]. SMILES [CSM^{+}07a]. SML [GS05a]. Sneaking [Joy01]. Sniffer [JBMP03, JKKL04]. snippets [Mot06]. SNITCH [NW06]. soccer [ATNT04, BHH^{+}08]. sociability [CMD08]. Social [AAF^{+}06, HL06, RM09, ADM^{+}09, CMD08, Cle07b, God06, Haz03b, LB00, LOP^{+}07, Med04, OM08, PBM07]. Socially [BNS08, BKS^{+}04]. socially-relevant [BKS^{+}04]. societal [MK06]. societies [Pop05b]. Society [Pop07b, Joe06b, Pop02, Wei05b]. socio [Kol08a]. socio-cultural [Kol08a]. socket [BM07b]. sockets [HHJ08]. SOCS [BZ04b]. SoDIS(R) [Hit05]. soft [Haz08, SH07]. SOFTICE [GGS06a]. Software [BA07b, Ciz03, Cle04a, Doo03, Ere00, Gib09, Hen03b, HT00, JDGAGMGC05, Liu06, NvdH04, OCM01, Vat00, AD01, AEPQ06, All03, BvdH09, BG08a, BL01a, BHY09, Bec08, BAET09, BS08, BCD^{+}04, BFMC05, Boe06, Cam05a, CHDH02, Cap02, Cha00a, CB02, CCMS05, CV08, CC01, Chr09, CC05, Cle05b, Cli01, CLK00, Con04, Con08, CHS05, CSP09, Cur03, CFH^{+}01, DPM00, DDA^{+}07, Doo08, EMM^{+}06b, Edw04, EMdL^{+}07, FH01, FC00, FK02, FK05, Fil01, FLS04a, GMN02, Gha05a, Gol02, GSW08, Got00a, GD05, Guz01b, HY09, Had01, Had05, HALW01, Hau05, Haz01b, Haz03b, Haz05, Hen08a, Hen08b, HPM^{+}08, HHLS06, HEM09, HETD09, HL08, HP07, HH06, How03, HS03, Huf02, Hum04, HMS08, Imp02a, IHL^{+}04, JKS00, Jen08, JDC00, JBB^{+}03, KvdHR05a, KH05a, Kum08a, Kur08a]. software [KFE^{+}07, Lar06, LCB^{+}03, LCW06, Lea08, Lee03a, LBA07, LBA08, Lie05, LL06b, LHL03, LTHN09, Lud05, LNR05, Mat06, McA08, MDH^{+}00, McC00b, Med01, MTBS09, MAFLR05, MAFLR06, MW09, NN00, NW06, NSH^{+}09, NH07, NBF^{+}08, Our04, PRC06, PDS07, Par06a, PDC01, PHM^{+}01, PTT06, PA03, PM02a, Qua06, Ram03, Rao06, RM08a, RRP01, RA06, RW07, Rob00c, RRD^{+}09, Rut01, SD09, SLC^{+}09, SL09a, SN08, SNA09, SYLL00, Shu05, SL08b, Sob00, SL03, Sol07a, SMH05, Ste01, SK05, TCK09, TMSZ07, TS00, Van08, Way05, Wea04, WB05, WBGM02, Xiz01, YM02]. Software-defined [BA07b]. soil [Got07]. solids [Par03b]. SOLO [LST^{+}06, Rus09, SCL^{+}08]. solution [BGG^{+}00, MBD04, MGH07, Raf08, RBlR01, Rol07]. solutions [BL05b, DSB^{+}03, Soo01b]. solve [KB08, Raf08, WVMN05]. solvers [dRTW04]. Solving [BCC^{+}07, RKDM01, ASPS04, Ars09, BLH01, BN00, CWHH05, CWHL06, CSM07b, Cun02, DWJ02, FP09, GGAB01, Gin01c, Gin01d, GMHS01, HB09a, KB09, Kum08a, cLwW07, LV05, MHA04, MH05, PFF04, PW08, SDE01]. Some [Bal00, Ham06, Tur08b, BM05, Edm08b, Gol01, HM00, HAuR04, Joy02, OG01, PBM03, Pij00, TRT04]. something [CM08a]. sometimes [BGD03b]. Song [LH05]. sons [Pop08b]. soon [Got09b, Reb05]. sort [Ast03, Rol05, WW00]. sorted [Hu05]. Sorting [FNW08, Gin03d, Man09, CLM^{+}07, Lax01, McC04a, NW01]. SOsim [MMP05]. SOTA [GCGBUFVI05]. sound [BABBH04, BR09, Gre00a, WSP04]. SoundSpaces [GSVG05]. source [CFW01, CHP^{+}05, CSP09, DDA^{+}07, DD08, EMdL^{+}07, GSW08, HPM^{+}08, HETD09, HS02, JWC07, JWC08, JB07, Liu08, McA08, McC00b, MWG08, MdL09, NN00, PBOP07, PPRV06, PL05, WBGM02]. sources [DeP06, Mus06, PRL09, TE08]. South [GM07, GSyC07]. Soviet [GG06a]. space [CM01, Kum09a, RC02, Shi06, WZ00, Haw09a]. spaces [Bow07a]. Spain [ALRY08]. Spamulator [ACd08]. Spanish [IMRJ09]. sparking [GL03a]. Spatial [JB07]. speak [HMS^{+}09, SRR00]. speaking [HL07, LL00, Osb06]. special [CIL^{+}03, Fin04, Fin05]. specialist [Str00]. Specialized [MK01, BGKT00b]. specializing [Vas02]. Specific [Bar02b, FJA^{+}07, Hen05c]. specification [BLA^{+}09e, BGG^{+}00, SHW^{+}09]. specifications [LHS09, LBC04]. specifying [SMS08]. specimen [Rol08b]. speech [KW07]. SPIMbot [Zil05]. spin [BA07a]. splines [FLS04b, LS03]. sponsored [Jam05, JBB^{+}03]. spot [TBM09]. Spreadsheet [Rol03b, Fon01, PFF04]. Spreadsheet-aided [Rol03b]. SQL [Cal08, Fek05, MG04, PL04b, SOSL04]. SQL-based [MG04]. SQLator [SOSL04]. square [DG05]. squares [Gin06b]. squeak [Guz01b]. Squint [Mur07a]. SRec [VIPCUF08, VIPCUF09]. SSM [DDL03]. St [DNTB05]. stability [Rob04b]. staff [SMR03]. stages [Lis09a]. standards [BPS09, Gor01b, MTBS09, Wil05]. Stanislavski [Kli05]. start [ANS00, LCG07, Mar01a]. start-ups [Mar01a]. started [APGR07]. Starting [Gin01e, WLMM09, DR07, MPQB^{+}00]. Starving [Rob01a]. State [Wil07, Boz06, CK02, CM01, CNP^{+}09, DDC^{+}03a, Gri03, KH05b, Reg03, SBG^{+}08, SCE07]. statement [Mas04]. states [dMFM09]. static [Kum00]. statics [RBA05]. stating [CR03]. statistical [LZL08, MDM04]. statistics [And07]. Status [BMW04, FKO07, GLF01, HAJ07, Mar05b, RBC^{+}05, DDL03, McC01a, WBGM02, YSLL00]. stay [CWL08]. staying [Wal08a]. stealth [Wei05a]. Steganography [Hun05, SWR05]. STEM [Mar05c]. step [BT06, DGMW04, SNH08, Yai08]. stepping [BDA03]. steps [RR06, VI00]. stereographics [ZF04]. stereotypes [GL03a]. Still [MJ03, DPBP09]. Stimulating [AJF07]. stock [She06]. stones [BDA03]. Stopping [Yan09a]. Storage [Hei03]. stories [CM08a, Dic07b]. story [BG08a, Chr09]. story-centered [BG08a]. story-telling [Chr09]. Storytelling [Rao06, Pap03a, Pap03b]. Stout [Rol02]. stranger [MJ03]. Strangers [Lis07b]. strategic [dRTW04]. Strategies [Cle09a, Rob00e, AHM^{+}04, CGM04, CL02, CT07, Goo08, KCF^{+}00, KB09, LW07, LB00, MEM^{+}07, MLM^{+}08, PWA06, She04a, Wai06, Wal01a]. strategy [OCLL00, WP02]. Strawman [RCE^{+}00]. strayed [Rol08a]. strength [Con04]. striking [Cle03a]. strings [Ast07a]. Striving [HBD^{+}01]. stroke [Ber09b]. stroke-based [Ber09b]. Structural [DWT00, BDM05]. structure [BS00, CB02, EVS07, HTK02, HCMM00, HCB04, MT05, RTH^{+}08, ZRY05]. structured [HWH^{+}05, Lis01a]. structures [Ada09, Aha00, BMG03, Bud06, CCL^{+}07, CSM07b, CHJB07, DV04, ESW09, Fek02, GEVZ09a, GMFA09, Gig02, LBM^{+}04, MS00, Mar05a, MD08, NM04, NM02, Oka05, PHBM05, Sha04, She06, SWR05, Ten03, WW06b]. Structuring [BTGD00, PD03]. Struggles [BS08]. Student [AMW06, BBY08, DDL03, DABJ00, DLRH08, Dic05, FMK^{+}00, GM09, Goe03, Han05a, Han06, Hen03b, KLOT05, LL04, OBL03, SG07, SL08a, SBE^{+}08, Sca00, SH07, WJDL04, WB05, YYLH06, ABAA09, AEPQ09, ABE^{+}09a, ABE^{+}09b, And00, AL09, AJF07, Bar05, BFD^{+}09, BU09, BTGD00, BDW^{+}05, BKW05, BYS09, BPW^{+}09, BM07b, BCC^{+}07, BL01c, CGB08, Cas01, Cha06b, CLM^{+}07, CSM07b, CMS07, CM01, CHP^{+}05, Cra07, CFH^{+}01, DTP03, DLRHP09, DSB^{+}03, Dic07a, Djo08, DFS03, Edm08b, ER09, FH01, Fle00, FB02, FAL^{+}05, FK08, Gha05b, Gib02, GPL^{+}05, GSW08, GD05, Gov07, HA02, HLL03, HY09, Ham06, HCJ^{+}08, HALW01, Haz03c, HEM09, Jac00b, Jac01, JY01b, JG08, Kac01b, Kac01a, Kac03, Kan07, KWW^{+}04, Kno01, KYC07, Kur08b, LML07, LDA^{+}00, Lax02, LL00, Lew07b, LOP^{+}07, LCL08, LM03]. student [MAFO08, MEM^{+}07, MDM04, MC04, MC08, MN01, MBE^{+}09, Ode00, Pap03d, Par06b, PB09a, PSH^{+}07, PGK05, PK06, PCW09, PJ05, Pop03a, RHE04, RHY07, RA06, Red01, RW05, Sal03a, Sal03b, SD00, SV03, SPM00, SLPG07, SSS^{+}09, Son09, SB06, TBAA09, TH06, TL01, Tho06, TETPQ08, TBMB09, TZ01, VT01, Van04, VHF07, Wal06b, Wic07, WGS05, WU01, WPS01, WM00, WR07, WKB02, YH03, Zin08]. Student-built [KLOT05]. student-constructed [Fle00]. Student-generated [GM09]. Student-oriented [LL04]. student-written [TETPQ08, TZ01]. student/faculty [LDA^{+}00]. Students [BW06, Car05b, Cle04b, DO07, Ega05, Gre00b, HMRM03, Ros02, ABH^{+}07, Aha00, AEH05, ABAE^{+}09, AA04, ABB^{+}03b, AR00b, AHM^{+}04, AT01, AGE04, ARVV06, ALBA08, Ars09, BGD03b, BMK09, BLH01, Bea02, Bea03, BM05, BBN07, Bec03a, BZ04a, BSV04, BRP^{+}03, BAB^{+}08, BGGR01, BC04d, BS05, BAdC07, BdCGB07, CCL^{+}07, CPP07, Can06, CS02, Car00, Car06, CMD^{+}07, CLM^{+}06, CNCL09, CV08, CTR^{+}03, CM08a, CMS06c, CBC09, CT09, CA02, CLP^{+}01, CBKM07, Deb02, Dei07, Dei08, DSM01, DK06, DKN08, EMM^{+}06b, Edw04, Ekm06, EMdL^{+}07, Eps06, ES08, EWH^{+}09, Eth04, Fek08, Fle01, Fra02, FS02, GSyC07, GY03, GM08, Gib00, Gin04b, GAGG04, GL05, GLC07, HMS^{+}09, HBH07, HTBS01, HL07, HSM06, HE06, HETD09, HV05, HV07, HR03, Hug09, HP06, HMS08, HKN^{+}02, Jen01b, Jin08, JCOP07]. students [KW07, KMMS02, Kum08a, Kur09a, LV08, LDHW02, Las03, LNCC09, LBK08, Lew07a, Lim07, Lin05, LC05, Lon08, LNR03, Lud05, LR08, LKPC04, MGS07, MAMC05, MAD^{+}01a, MAD^{+}01b, MO09, MW09, MS07, MJ03, MPJC09, MdL09, MT05, MKLH09, NEN00, OHZ01, OBLH03, OB05, OM08, Per09, PBWS06, PB06, Pil09, Qid07, Qin09, RBG02, RC00, Reb05, RMY08, Rio02, RKDM01, RHL^{+}09, RIP^{+}09, SDM^{+}02, SD03b, SR07, Sig07, SHW^{+}09, SL08b, SJ05, SJ06, SM05b, SMH05, SSS07, SHS^{+}08, Ter05, Tho01, TT06, TV08, Tow09, TMSZ07, TT05, VR04, Wal07b, WG07, WSP04, Wic09, WB06, WD03, Wos06, YPdMA^{+}08, YA05, Zel00]. studied [LS02]. Studies [NW03, ACL^{+}03, Bor04, Dei08, EH07, IK06, KH06, KT07, PD03, RCR04, TL01]. Studio [KTKP02, CS02, DSBK01, HNC08, Ram03, WK07]. studio-based [CS02, DSBK01, HNC08]. study [ASA00, BC08, BBC06, BW06, BBY08, Bou03, BPW^{+}09, BCL00, BS05, BS01, Cal08, Cam06b, CL00, CCM08, DTP03, ES06, ER09, FAL^{+}05, GE08b, GEVZ09a, GG06a, GLD^{+}05, HH07, Har05, Haz01b, HAFT09, HLN07, Kla02a, KPT03, KYC07, LAMJ05, LAF^{+}04, LCS^{+}09, MHW07, Mar07c, MAD^{+}01a, MAD^{+}01b, MBD04, MN01, MT05, OBL04, Pap03d, PRW04, PB03b, Qid07, Rad08, RJM^{+}05, RKD02, SB09, SM02b, Sok02, Sol07b, SS00, Sud08, TRR03, TT00, To04, VZ06, VZGE07, WS01b]. Studying [Dei07, GHBG^{+}03a, GHBG^{+}03b, AGE04, RVWB04]. style [Ras03]. styles [BL00, Bur08, CS05, DFS03, GSyC07, HRBC03, LCW06, PR02, RWP03, TRWJ02, ZTS^{+}09]. Subject [SPM00, Lea01]. subjective [Car05b]. subjects [RHL^{+}09]. submission [RMT09, SHP^{+}06]. subsetting [DLPQ04]. subversion [SLC^{+}09]. Subverting [CKM07]. success [BC06a, BR05a, BYS09, BDMY05, BL01c, DTZZ05, HRB^{+}09, JJK03, JY01b, LV05, RRKS06, Ric09, RRR02, RRRH04, RVWB04, Wen04, WS01b]. successes [GAL^{+}06, NRRR06]. Successful [HB09a, MEM^{+}07, RMZG06, DPBP09, MKC09]. such [SB06]. Sudoku [NL08, SH08]. suggested [Soo01b]. suggestions [DWY03, LBC04, TS02]. suitability [Kla02a, SP07]. suite [BA04a, BZ04b, DTW07, Nør00]. sum [Aha00]. summaries [Cas01]. summary [Imp08]. Summer [Sko07, CPP07, DLM07, GHLL08, KDP06, LA07, LA08, LNCC09, Mar01b, PMC^{+}04, CLR09]. summit [Haw09b]. sun [TBM09]. Supercomputing [NSW08, AV02a]. Support [ADG^{+}01, CF02, GM02, Rob04c, BHR07, Bec07, BDM01a, BP05, CFJ^{+}02, Cas03, COR07a, CB04, CRPZ06, Ekm06, FMKQ08, GDB03, Got04, HR02, HHF^{+}03, Ing09b, KCF^{+}00, LV03, BGN09, Mar07c, Mar09b, MA01, Mer09a, Mur07a, NRB^{+}05, Rad07, RBG02, RLR07, RBM04, RJM^{+}08, RK09b, SAHS04, SB06, TLS07, TLW^{+}09, VA02, VCC^{+}05, VCC^{+}07, Wal02a, WC04, WU01, Wir03]. supported [JDGAGMGC05, RA06]. Supporting [AAD^{+}07, BCC^{+}05, FKKW00, Lap07, PBK04, TETPQ08, BS01, Ell06a, GD08, HTK02, KK02, PH07, SFPFV08, Tho01, TLW^{+}04]. supports [VT01]. surface [FLS04a, SL08b]. Surprise [Mot06]. Survey [Wil04b, Bla09b, Car02b, CPT02, Dal06, DJS08, FG07, Goe03, Mos02, PSM^{+}07, Ska02, pYB08]. surveys [Lew07b, MNBK03]. survival [Esh07]. Survivor [HBCW03]. sustainability [MSM08, MMD^{+}09, YM06]. sustainable [RK09a]. sustaining [Lis09b]. SW [PTT08]. swarm [ACK08]. switch [Mas04]. syllabi [RS09]. syllabus [TYC^{+}07, Wal05c]. symbol [GCGBUFVI05, UFGCGBVI06]. Symposium [ACM01, ACM06, Ano03, BTHR06, CDWH00, DNTB05, DRFG08, GWG02, GKJD03, Imp02b, JKDN04, RHDR07, WMGR01, Imp04, Val04]. synchronization [AF03a, Bha00, Hol00, Rob00b, Rob01a]. synchronous [Bow07a, Bow07b, BDF^{+}08, KV08, MW09, PC08]. syndrome [Ega05, Gin03c]. synergies [CF04b]. synergy [HPM^{+}08]. synoptic [GLC07]. syntax [LL02]. Synthesis [TG05, CHM^{+}05, CSS05, SNA09, Shu06a, Vat01]. SysRel [Bac03]. syst [Sci07]. system [AL05a, AH08a, AR08, AYWM08, Alv05, AAS^{+}04, AS02, BHR07, BLH01, BCHS00, Bla09a, BG08b, Bor07, BCC^{+}05, BGKT00b, BH07, CM05a, Cha03a, CWHL06, DWT00, DLRH08, DAHA09, DWJ04, Don01, EH05, Eng06, Ere00, FPQ^{+}08, GGÁC03, Gel00, GSVG05, GSW08, Gou09, HST02, HLS02, Hum04, KHPC08, KvdHR05a, KT01, KMS01, MR06, NB01, NTTHT09, Noo06, PWC^{+}05, PL05, PRB09, QLT09, Rob06b, RK09a, SDE01, Sha05, SK02, She02, SSC00b, SAHS04, Soh04, Soh06b, SHP^{+}06, SKM02, TIKT04, Tow09, VT01, VIPCUF08, ZSW04, CTP^{+}04, EFK04]. system-level [AL05a]. Systematic [Chr03, Gon06]. Systems [AR00a, Bur06, Aba08, AE01, AD01, Alb09, BA08, Bec03c, BAMSK05, Bha00, Bla09a, BHH^{+}08, BO01, Bry08, CFW01, Cle02a, Cli04, Cli01, Dav09, Dav04, Dic05, Dic00, Don08, EMM^{+}06b, GGS06a, GSVG05, GLF01, Gor04a, Gor04b, Gor05, Gor06b, GGS^{+}06b, Got09b, HRBC03, Hol00, How02, wH07, wH09a, wH09b, Imp02a, KK02, KC07, KPT03, KT07, Lab09, Lea01, LFL06, LKGY00, MMP05, McC05c, Mer00, Moo02, NV05, OCLL00, PMF08, PG07, PP06, Rob02, RJM^{+}08, SA07, She07, TT06, Top08b, Top09a, Tut00, VB02, WM03a, WSC03, WKC^{+}09, Wis05, ZB08, vW00b, LFGY02].

table [UFGCGBVI06]. tables [BL01a, BS02, GCGBUFVI05]. Tablet [SJS^{+}07, AAV^{+}03, AAS^{+}04, BG08b, CTP09, KHPC08, Par06a, SAHS04, TEP06]. tablet-based [BG08b]. tabletPC [Gol04b]. tactics [HAM00]. tactile [Dic07c]. Tag [CBH09]. Tail [RSVI09, For04b]. tailing [To04]. tails [LW00]. Take [Pop08b, Fri05, Kac09a, LZ07]. taken [Wir02]. Taking [Alv05, BT01, BP02, CMT04a, CGT05, Lim07, Mar08, Mar09a, PMT03, SDG^{+}08]. tale [JA03]. tales [Edm08b]. Taming [Fre04, HSSC05, RC04b]. tangible [RB09]. tank [Gin04e, MD04]. target [Ruc07, TY06]. targeted [DTZZ05, Par06a, PK08]. Targeting [LJMV07]. Task [RBC^{+}05, RBC^{+}06, Bow08b, HA04, PCW09, RCE^{+}00, SB09, LRKM07]. tasks [Geo02, SDR09]. taught [BDM05, FG08, MVC00, RWGM00, Shu06b]. taxonomies [Lis07b]. taxonomy [Bow08b, FJA^{+}07, Lau08, LST^{+}06, MLK07, Rus09, SMS08, TH06, Yeh02]. TDD [MS07]. Teach [JBMP03, LM05, LS06, QLT09, SL09b, AEPQ06, Arm04, Ast04, Bru04, Bru05a, BLW01, Cha03a, DCP01, DLP08, DAJDRG08, DL06, Etl06, Fag00, Hag00, HB01, HDH06a, HCHP01, HAH^{+}06, Hun03, Lim07, Pet09, Ree00, RD00, RH07, Sae09, Soh04, Sto06, TW05, Tre02, VB02, WSC03, War04, WN05, WSP02, Wil00, dMFM09]. Teacher [VHF07, Bow08a, Esh07, FS09, God06, HLL03, HLNT01, KP04, Ni09, Ros01, SB06, WPS01]. teachers [ABAE^{+}09, BC07d, BCLW08, CBKM07, DBBK02, GEHR09, GM07, cL05, LR01, LH03, LA07, Lap07, LBA07, Ni09, OHZ01, Pap03d, RH03, RH08a, RHL^{+}09, Wal04c, fCST02]. Teaching [AE01, AF03b, Aho08, ABK^{+}01, ABB^{+}03b, ACSS06, AM05, ALBA08, Ars09, AO06b, APGR07, Bar02a, BBMR03, BL05b, Bec01, BMG03, BA07a, Ber02, BCD^{+}04, BWC^{+}05, BJJ01, BAMSK05, Bog04, Bor04, Bot05, Bou09, Bra05, BB03a, Bur03, BB03b, Cal08, CCL^{+}07, CM06a, CHDH02, Car04, CWM05, Cha04a, Cha04b, CC05, Cli04, CMS06c, CR02, CDP03, CSP09, Cur03, Dav04, Dew05, DPM00, DV07, Edm08b, Ega07, ESP03, ES05b, Esh07, Fek02, Fek05, Fek08, FL05, Fre04, GEVZ03, GOMSJVGP08, GL07b, GL08b, GMJDA09, GT01, GC05, GSW08, Haz01b, HL06, Hen05c, Hol08, HH05a, HH06, HS03, wH07, Imb03, Jen08, JCG06, JCOP07, Kap04, KCF^{+}00, KDH^{+}00, Kop01, KR06, KS04, cL05, LS04a, Lee01b, Lie05, LS08a, LTHN09, LC05, LS03, MP09a, MFL01, Mas04, MD06a, McA08]. Teaching [MK00, Mer00, MHK^{+}03b, Mur01, NN00, NW03, Ort03, PH03, PMC05, PTT06, PP06, PNF00, Qin09, Ran00, RBlR01, RS04, Roi05, Rou04, Ruc07, RMN06, San02, SN08, Sch00, Sch09a, Sch02b, Sco03, She07, She03a, Sna06b, Ste04, SSS07, TF07a, TT05, Udo06, Van08, Vat00, Vat01, Wag05, Wal01c, Wal02b, Wei05a, Wic05, WYOH02, Wol01, WWGT07, Wu05, XSD07, Yan09b, YLYX06, ZV04, ADG^{+}03, ABB^{+}06, ABAE^{+}09, ADG^{+}01, AKCC06, App06, AHJ^{+}07, Ari00, ABA09, ALH07, BF05, BC07b, Bal00, Bal03, BGDR05, Bec03c, BA04a, BZ05, Bha00, Boe06, Bow00, BR02, BL01b, Bry08, Bud06, CS02, CWHH05, CD01a, CS00a, Cas09, CHK^{+}01, Cha02, CC01, Chr05a, Cle06b, Con08, Cra07, CPST02, DO07, Dav09, DVE06, DFGM08, Deu08a, DWJ02, DWJ04, Dix04, Dou00, DHGR04, EB08, ES05c]. teaching [FM03a, FSAF03, Fon03, Gag02, GET05, GOGPGP08, GV09, GP07, GCPQA05, God06, Gri08, Gri02c, GNPL01, GGKT04, GA03, Guz01b, HK01, HS04, Ham02b, Ham04b, HR02, HL04a, Haz08, Hen07b, HW06, HRBC03, HA04, HY01, Hol05, HAJ07, HD07, wH09a, wH09b, Ing09b, Joe06b, JDFV08, JCK03, KK02, KCI03, KR01, LH03, LDB^{+}03b, LB00, LHS09, LE07, LP02, LRPQ04, LL06b, Lis03a, LBM^{+}04, Luq05, MMN^{+}09, MA01, McC03a, McC04c, MCS04, MMC01, Mer09a, MR02, MBK09, MV01, MT09, MICM08, MB05, MPK08b, NL04, NvdH04, NLMN06, NSW08, NV05, NBL09, ÓT04, OG06, Ola00, Osb02, PB05b, Pap03a, Pap03b, PP03a, PDS07, PBM03, PHC08, PHC09, PKB06, PSM^{+}07, PRL09, PA06, PD06, PPA08, PGC^{+}06, PEH07, Pul00, Qui06, RRP00, RH08a, Raj05]. teaching [RMZG06, RM08a, RRP01, RA06, Reg03, RM08b, Ric00, Rob03a, RCR04, RTH^{+}08, Sam05, SC07, Sav05, Sci07, Sha06, She02, SLPG07, SYLL00, Shu04, SW04, SP07, SNS07, SSS06, Soj03b, Ste01, Str00, SF04, Ten03, TM04, UD01, Utt06, VAFM03, Wal06b, WKC^{+}09, WS02, WVMN05, WSS05, WI09, WSP04, WF08, WD03, Xin03, XM06, Zil05]. Team [Dei05, KvdMvDvdV00, AC08, AO06a, BYS09, CHDH02, CHS05, FH01, GDHW06, GY01a, GD05, Hau05, HRB^{+}09, HLCS08, KvD06, Las09, Lax02, LSGB07, MD05, SR07, SDLdQ05, Try05, WL00]. team-based [Las09]. Teams [Hil00, MFC00, ABE^{+}09b, CB04, GPL^{+}05, HALW01, RRD^{+}09, Rut01]. teamwork [Las03, Mar07b, WWK^{+}05]. tech [CCH^{+}03c, PHZ^{+}03, DGHC03]. Technical [ACM01, ACM06, Ano03, BTHR06, CDWH00, DNTB05, DRFG08, GWG02, GKJD03, Imp02b, JKDN04, RHDR07, WMGR01, DKR09, Etl06, Gra05, Imp04, Kac03, MAMC05, Val04, WW09]. Technically [HL07]. technique [Cle02b, Gha05a, JWC07, Xiz01]. techniques [BCM05, CPST02, GGAB01, Gra07, HR03, Lar06, LCB^{+}03, MDGM05, PD06, RA06, SL09a, dBGdMB09]. Technological [BHM^{+}06b, MGS07]. technologies [AL09, CI03, Dra09, Ell02, HHB05, KDP06, Mik05, RK09a, SJ02]. Technology [ALRY08, BCK04, CJGU02, CIL^{+}03, CFPL05, DSF^{+}03, DGS06, FKCM01, Fin01, Fin02, Fin03, HPT07, LS04b, TLW^{+}04, TLW^{+}09, AKM00, ASPS04, All03, AAD^{+}07, AL07b, AM05, BCS02, Cam06b, CS08, CDGS01, CH07, CSD02, Dic07a, DS08, DDC^{+}03b, Eps06, GE09, GPP00, Gor02, Gor03b, Gup01, HTBS01, HKCW08, LML^{+}01, LLP04, Lif04a, Lif04b, LGA^{+}01, Lon08, LHS04, MVC00, MVD01, MHR06, MTBS09, MTHS00, MT09, OGMA07, Pho01, Pul00, Put01, Ram02, Rob00d, SSC00a, SMTH00, VMD01, VKV02, WB02, Wol07, FLL^{+}01, Spo00]. technology-based [GE09]. Tele [HMS04, LRKM07]. Tele-lab [HMS04]. tele-tASK [LRKM07]. telecommunications [LL06b, VKV01]. Telemachus [SD00]. telling [Chr09]. tempered [Ree02]. Template [SP03, PKB06, SFD08, WP02]. Ten [CL02, TS02, BFMC05]. ten-week [BFMC05]. tendencies [CHMG01]. tennis [BM07a]. Term [LBC04, HY09, KNT02, Las09]. terminology [LL02]. terms [Lis07b]. ternary [Goe08]. terrain [OG05]. terrorism [Poh05]. terrorists [Poh05]. Test [Ada09, BW05, JS06, JS08b, KI08, Pro09, DJS08, DJC09, Fon04, Gor00a, GHG^{+}08, JMS02, Lew02, Sch09a, SSP^{+}05c, WP06a]. Test-driven [Ada09, JS06, JS08b, KI08, Pro09, DJS08, DJC09, Sch09a]. Test-first [BW05]. testbed [RN02, Ste09a]. Testing [CG02, MS05a, AEPQ06, BUCP02, Chr03, Edw04, Gol02, GSS00, Hep06, JS06, Jon01, KH05a, Kre09, Kur08b, NP02b, PKR03, PR04, PJ09, SHP^{+}06, TV08, TLS07, WSW05]. tests [Fon02, Gha05b, RF00c, Soj03c, TETPQ08]. tethered [ACK08]. Tetris [Par01]. TeX [Soj03c]. Texas [BTHR06, CDWH00, Dav01a]. text [CSM^{+}07a, CNCL09, Dic07a, WK06]. text-enhanced [CNCL09]. textbook [BAWRMM02, DMKS01, MDD06]. textbooks [BHN^{+}09, Car03]. textiles [BEE07]. TextMOLE [WK06]. th [Laz02a]. their [BS08, BZ04a, CT09, CBKM07, FU02, GKHPQ07, JY01a, Kac09b, YA05]. Thelma [EN05]. them [HAH^{+}06, Pop08b, SM05b, Tea02, Tra03, WVMN05]. theme [GWG02, Ras03, Ree02]. themed [SPW^{+}08]. themselves [Boy00]. theorem [JDC00]. theorems [Haz03a]. theoretical [CCY03, HL04b, KAPG07, OG01, SB08, Tra03]. theoretically [BABBH04]. theories [MT08, To04]. Theory [RM08b, ARVV06, Arn07, BJJ01, CFR04, CCY03, CGM04, Cle09a, Doo03, ET05, GET05, GGH03, GKL^{+}02, Ham07, Han02, Haz02, HW06, HR00, Kes04, Lim07, MRH09, MMC01, Mer09a, MH06, Nod03, Pil09, RBFR06, RLR07, RWL^{+}09, Shi06, Sig07, SBR^{+}07, Ver05, VVV04, Vin05]. there [Cle08a, MD04, Pop08b, Raf08, RP06, SB06, VZ06]. these [GET06, Gri02b]. thesis [Lis07a]. Thessaloniki [DSF^{+}03, Fin03]. they’d [APGR07]. thing [Eps01, SB06]. Things [PS06, BVPE06]. think [AKCC06, ALBA08, Ars09, Bai05, CS02, MPH09, TT05]. think-alouds [Ars09]. think-twice [Bai05]. Thinking [Cle08b, Cle09a, Cle09b, Got08, Got09a, Got09b, LF09, McC04c, Osb06, Arm04, AGE05, AGE06, AGEH06, Arm08, AHPS09, Bur06, CPT^{+}09b, FG08, GMHS01, HHK^{+}09, Han07, HBD^{+}01, HHF^{+}03, HCW07, Mas05, PB09b, Qin09, Rao05, TH06]. third [Doo08, GWG02, Imp02b, LR07, PH04c]. Thirty [ACM01, Ano03, Imp02b, RHDR07, ACM06, CDWH00, DNTB05, GWG02, JKDN04]. Thirty-Eighth [RHDR07]. thirty-fifth [JKDN04]. thirty-first [CDWH00]. Thirty-Fourth [Ano03]. Thirty-second [ACM01]. thirty-seventh [ACM06]. thirty-sixth [DNTB05]. Thirty-third [Imp02b, GWG02]. those [Dic07c, Har05]. thought [PB09a]. Thoughts [Wal06a, Wal06b, Bal00, BM05, GM06b]. thread [Fek08]. thread-safe [Fek08]. threaded [CL00]. ThreadMentor [She02]. threads [BZ07, Don08, OB07]. ThreadsTM [FIG07]. Three [OJJ00, Cle09b, CLP06, LS01, PQEA^{+}04, Sch04, SRR00, Wai06]. three-course [PQEA^{+}04]. three-year [CLP06, LS01]. Threshold [BEM^{+}07, EMM^{+}06a]. throughout [Gol02, Gou08, MP09b, RMB00]. Tiger [Fre04, HHRA06]. tight [Boe06]. Time [JT04, ABE^{+}09b, ATNT04, BF05, BDW^{+}05, Boe06, BMT^{+}07, BMT^{+}08, BMD^{+}09, BBS06, Cha03b, Cli01, Dav09, Doo08, EA07, HDH06a, Hen07b, HLNT01, HBCW03, HD07, KW07, LFG08, Lis01a, Mar01b, McC05c, PSW03, QLT09, RBG02, Rol02, Ste01, Wal05a, Wan06]. times [Hem05]. tips [APGR07, Bay09, GNPL01, McC05b, McC05a, McC06c, McC06a, McC06d, McC06b, SR07, SRGF08]. TLA [HY01]. to/from [AGHM02, AGHC03]. today [Gup06, Wai06]. together [CS06, HBEM03, MAR07d]. told [APGR07]. tolerant [SD03c]. tomorrow [PBWS06, PB06]. too [Bec06, Got09b, Par09, Reb05, SB06]. Tool [HR02, Ing09b, AK09, AC07, AKT07, AMUFVI09, AKA09, Ber03a, Bha00, Boi09, BBE06, BGKT00a, BP09b, BB01, CD01a, CTP09, CKMW02, CKW04, CTBS06, CI03, DDL03, DFGM08, Deu08a, Dic05, DL06, Doo09, Ext00, FH01, FLS04a, FH05, FCA05, GCGBUFVI05, GOGPGP08, Gar04, Gri01, HLL03, KS00, KB08, Kop07, LR01, Lar05, Liu06, Lor03, MLH05, MKKC08, NW06, NS02, PP03a, PA03, PRL09, PK08, Qui06, RN03, RSF00, RF00c, SD03a, SL00b, Soo00, TL01, TLS07, TEP06, TZ01, Wis06, YYLH06]. tool-based [KB08]. Toolbook [Ell00a, PP03a]. Toolkit [Sch02b, MMN^{+}09, VVV04, WD05]. Toolkits [Ras00]. Tools [LCB^{+}03, PGC^{+}06, AL06, ABE^{+}09a, BP07, BA04a, BBK05, BC04b, CPST02, DLP08, DPM00, GKHPQ07, Ham07, Han02, Lar06, LBA07, LBA08, MHW07, McC01b, NRB^{+}05, Nør00, PFF04, RRP00, RRP01, STTBC07, SM03, SF02, TLH^{+}09, TWS09, VA02]. top [FG07, Gin01e, LdLM^{+}08, WF08]. top-down [Gin01e, WF08]. top-rated [FG07]. topic [AHH^{+}06, Chr03, Ols05, Ros00]. topics [BGSL06, Dal06, DLRHP09, Hui01, MCS04, SL08b, Wal05b, WN05, WW06b]. TOPKAPI [RF00c]. topologies [MK00]. topology [CCJ^{+}02, EGST08]. torture [Got08]. TOS [NB01]. touch [Jon05, RH07]. touch-screen [Jon05]. ’Touching [Dic07c]. tournament [KF03]. tower [Reg02b]. Towers [Min07]. toxic [CWC09]. toy [GMN02]. toys [AL06]. trace [Ing09a, Lor03, Ter05]. Trace4J [Ing09a]. Tracing [War01, Gol08, HSB09, KN09, LAF^{+}04, LFT09, VS07a]. track [ALO^{+}03, KvLA09, Khu08, ZLOB07]. Tracking [TFG05, AWS^{+}09, Hui01, TY06]. tracks [TA08, ZLL^{+}07]. tradeoffs [CB02]. Traditional [GS05a, AHL04, CM08a, GE09, Gra07, KD07, KKR06, Pho01, RC00, ZH02]. traffic [ZSW04]. TRAILS [DGHC03]. Training [SBLH07, TT06, dRTW04, Bac03, KV00, KP05, KP04, PBK04, PY05, RTH^{+}08, SSA08]. transaction [BS02, Fek05]. transactional [Cle07b]. transcription [KW07]. Transfer [AGHC03, Cam05a, Cam06a, Haw08b, Tur03]. transformation [Cha00b]. transformations [LJ08, MBE^{+}09]. Transforming [RHY07]. transition [PH07, Pop07b]. Transitioning [PQEA^{+}04]. translating [Dic07b]. translation [Ber03b, McC01b, Rob05]. Translator [Röß06]. transmission [Chu01, Chu03, Shi06]. transportability [LZ07]. transversal [Sic09]. trap [Gin03b]. traveling [Sok00]. treasure [Gig02]. Treasurer [Wal01b]. Treating [GY03, Lax01, Sol07b]. treatment [Car03]. Tree [Par02b, BK03, Rol02]. Tree-list [Par02b]. Treemap [KB04b]. trees [BNSW05, Hab02, LST^{+}06, SM00b, TBAK04, Wal00a]. treeview [YYLH06]. Treisman [CMS07]. trenches [BBMR03, Bay09, SM03]. trend [TSO05]. Trends [BAK^{+}09, RR03, AAO^{+}08, Gor01b, HV07, WR07]. trial [DGMW07, Edw04]. trial-and-error [Edw04]. trials [CL08a]. tricks [McC05b, McC05a, McC06c, McC06a, McC06d, McC06b]. TrucStudio [POM^{+}08]. truth [Cle01b]. TS [Chr05b]. TS-05 [Chr05b]. TTM [BC04b]. Tuple [Gin05b]. Turing [BL01a, CM05b, GOMSJVGP08, Sor05]. turned [Got05]. Turning [RBFR06, Got07]. turnover [CSC03]. tutor [ACL^{+}01, FK04, Kum05d]. tutored [ADP01, LAP08]. Tutorial [Car02a, BW05, Ber09b, LLB02, Mea06, Rob06b]. tutorials [vW00b]. Tutoring [RH08a, BPW^{+}09, DWT00, KP05, MWPB05, Sha05, SK02, Soh06b]. Tutors [Kum04, BL00, Dic07c, Jin08, Kum05b, Kum08a, Pil03]. twelve [WS01b]. twice [Bai05]. twin [Min07]. Two [Bur03, Cam01, HCK07, MRH09, Moh09, PB05a, vLK09, BvdH09, BS05, Cam02b, Cam03a, CHK03a, CHK03b, Cam04a, Cam04b, CHK05, Cam05b, CHK06, Chu03, CN07, HM05, Jon02, JCK03, Kle01, Oka05, PFVI00, TFG05, XSD07, dMFM09]. two-body [Jon02]. Two-Dimensional [Bur03]. Two-Year [HCK07, Cam01, Cam02b, Cam03a, CHK03a, CHK03b, Cam04a, Cam04b, CHK05, Cam05b, CHK06, CN07, Kle01, TFG05, XSD07]. TX [ACM06]. type [ABE^{+}02, TE08]. types [Bow08b, CMRR04, CM08a, GSyC07, LCW06]. typical [Tra03].

U.S. [LZL08]. UAE [BHM06a, BAAM06]. Ubiquitous [GS06b, WGS05, BHBG07, BHRB08, Cha03a, EWH^{+}09, THH09]. ubiquity [Kac06]. Ucigame [Fro08]. UCITArian [Got00b]. ugly [MD05, Rol01]. UK [BCK04, FKCM01, HPT07, Fin01]. UML [GKHPQ07, Got04, PY05, UD03, dMFM09]. uncomfortable [Cle04a]. uncomputability [IS09]. undecidability [GELS04b]. under-represented [HRB^{+}09, PB06]. underdeveloped [PV08]. underestimated [RM09]. Undergraduate [BBMW03, DBK^{+}09, Ell06b, GK02, HL08, IK06, Lan01, Rah08, SHN^{+}00, TSC^{+}02, Way06, AA04, AGE05, AGEH06, ALO^{+}03, Bal00, BHBG07, BBN07, Bem02, BZ04a, Ber03a, BA07b, BP09a, Bur04, BDDL04, BS04, Cha05, Coh02b, CS03b, Deb04, DVE06, DH06, DUM^{+}05, DDA^{+}07, DRKG02, DTL^{+}02, ELE^{+}02, Fra02, Fri05, GGS06a, GKHPQ07, Gha07, GCF^{+}01, Ham05, HBH07, HLM01, HWH^{+}05, Hua03, HS03, JCG06, Jon00, JCK03, Kap04, KvdHR05a, KDP06, KKR06, Lar06, LS01, LNR03, MAFW08, MDH^{+}00, MMLT09, MDL^{+}02, MKC09, MKBW07, NTTHT09, Ols06, Osb02, PDC01, PHM^{+}01, PWA06, PD06, Pow08, PP04, Qua04, RF09, RHKS01, Sha04, SL00a, SF02, Top07a, Top08a, UHFT09, Wen04, WF08, YD04, ZLL^{+}07]. Undergraduates [GR00b, Alb09, Cha04b, DBR^{+}08, DFGH04, Fon01, GAL^{+}06, Les04, LZL08, MR02, O’L06, Pap03d, PSH^{+}07, Reg03, SC07, Sta05, TBM09, UD01, WK07]. underrepresented [CPP07, PBWS06]. understand [Wic09]. Understanding [BZ07, Gov07, Ira04, LW05, PCW09, Can06, CD01a, CKMW02, HLL03, IS09, JDGAGMGC05, KWW^{+}04, Lin05, LJ08, OBL03, PGK05, PK06, RBA05, SBE^{+}08, SH07, Sur02]. UNESCO [MvW01]. unexpected [TRT04]. Unicast [Bri06]. unified [HB09b, MCS04]. unique [Bur08, CM06b]. unistroke [Ber09b]. Unit [NP02b, PJ09, BUCP02, FLS04a, PKR03, SC07, WO07]. units [VMD01]. universal [GL05, MH06]. universe [Ber06, MS00]. universities [ESS^{+}07, LCPQR05, MKC09, WTTT^{+}08]. University [COR07a, GE06, ASA00, Bou09, CT07, GSyC07, KH06, KFE^{+}07, LMS00, OHRL01, PH04c, Reg03, TFG05, ZH02, CTP^{+}04, CL02, CCM08, Dav01a, FLL^{+}01, GCW^{+}03, RHC07, Wil07]. UNIX [Rob06a, SHVP05, Gol01, Sol07a]. Unix-centric [Gol01]. unknown [HW08]. unknown-answer [HW08]. UNLOCK [BLH01]. Unlocking [FM02, FM03b]. unplugged [Hen09b, NKI^{+}09, TBAA09]. unstuck [HH07, MEM^{+}07]. unsuccessful [HB09a]. unsuspecting [CM06a]. untapped [BCS02]. unteachable [Tre02]. unusual [Ruc07]. Upcoming [Haw09b]. Update [Cam04b, Kle01, Top07b, Cam01, CCD^{+}08b]. UPE [Pop04, Pop08a]. ups [Mar01a]. upset [McC01a]. upside [Got05, MVC00]. Upsilon [Pop09a, Pop08a]. urban [ACK08, MPK^{+}08a]. Urbana [CCM08]. Urbana-Champaign [CCM08]. USA [ACM06, Ano03, BTHR06, DRFG08, GKJD03, JKDN04, WMGR01, Imp02b, Imp04, RHDR07, Sal03a, Sal03b]. usability [AAC^{+}05, Cle02a, CI03, FC00, Hep06, HPM^{+}08, LBC04, RCR04, RS03, Shu05]. USAR [ACK08]. USC [Dav01b]. Use [ABE^{+}09a, Bar01, DW05b, Dra09, Guz01a, Imp05c, JDFV08, PB05b, Rob04c, SL09a, TM04, AK09, ACK08, Bar02b, Bow00, Cha02, Cha05, Cle02a, DTP03, DLRH08, GT00, Gar04, GCF^{+}01, GLC07, HBIG03, Hu06, JY01a, Kül09, LFL06, LBA07, Lew07b, MAFO08, MT09, O’N05, OPS^{+}02, PL05, Phe07, RA06, Ste09a, SMH05, ZH02]. Useful [Fin09, BEM^{+}07, BC04d]. User [GS05b, ABE^{+}09a, BL01a, Bor07, DSCU01, GH06, Guz01b, Kac09a, KA09, Lar05, LBC04, MMN^{+}09, NRB^{+}05, PNF00, Qua07, RR06, Tra07, Dav04]. user-centered [Tra07]. user-level [MMN^{+}09]. users [GS05b]. Using [AT04, ABAE^{+}09, BED^{+}03, Bay09, BBN07, BU09, BBW04, BKW05, BYS09, BDMY05, Bru05b, BB03a, Cas03, CIK^{+}03, CFL^{+}04, CJ01b, Cha06b, COR07a, CCY03, CC01, CM01, CF04b, DCP01, DCT06, Deu08b, Edw04, ES05b, ES05c, Fag00, FH01, FM09, Fon01, Fon03, FdMFPT07, GCPQA05, Geo02, Guz01b, Hag00, HB01, HILS05, HCHP01, HETD09, HRB^{+}09, HM02, Hun03, IL04, Imp07, Jac00b, JKDS01, JMS02, JBMP03, JKKL04, JG08, KCI03, KT01, KTKP02, LW07, LML^{+}01, LB08a, Les03, LP02, Liu08, LHS04, LS04b, MFR^{+}08, MAMC05, MDGM05, Med01, MNBK03, MGE^{+}05, MICM08, NW03, OGMA07, PP03a, PB09a, PH03, PHBM05, PHC09, RBC07, Ree00, Reg03, Rob03a, Rut01, SD08, Sam05, Soh04, Soo00, SJ05, SJ06, SM05b, SHL^{+}00, SP03, Sto06, TWS06, TW05, Tow09, UD03]. Using [VB02, WSC03, War04, WN05, WSP02, WSW05, WW06b, Wic09, ABK^{+}01, And00, AL07b, Ars09, AO06b, BL05b, Bec03b, BMG03, BJJ01, BH04, Bra05, BGKT00b, Bry09, BKS^{+}04, CMRR04, CHDH02, CSM^{+}07a, Cha00a, CLM^{+}07, CKM07, CMS06c, Con08, Cor07c, CSP09, DVE06, DSCU01, DUK02, DW07, Doo09, ELC07, Eng04, ER09, Fow01, Gag02, GMFA09, GMN02, God06, GGH^{+}08, GMJDA09, Gui05, Ham04b, HAuR04, Hen05c, HPC^{+}03, HPCZ04, HL08, HL04c, HR03, HH05a, HH06, HS03, Imb03, JFH00, JCOP07, JCK03, KW07, KF03, KPT03, KR06, Kum08a, Kur08b, KFE^{+}07, LCB^{+}03, Law01, LDB^{+}03b, Les04, LR04, Lim07, LS08a, LS06, LS08c, MFR^{+}09, MMP05, Man09, MDD06, Nef04, NN00, NV05, OJJ00, Ort08, PB03b, PT03, Phi04, RC02, RMT09, Rob00c, Ros05, RF00b, Rou04, RSVI09]. using [Rus09, RRD^{+}09, SC04, SM02a, SV03, SJH06, Sha03, She03a, Soj03b, Soj03c, SHP^{+}06, Sud08, TL01, TQCPQE08, Utt06, VW00a, VCJ09, WSS05, Wic01, Wic03, WS06, Wir03, WYOH02, Wu05, XM06, YYLH06, dBGdMB09]. utensil [PA03]. utility [BP00]. utilized [Gin01c].

validating [TZ01]. validation [NSS^{+}05]. validity [Hit05]. valuable [McC02]. Value [Ros02, Gib02, Hit05, Jen01a, Rob00a, Wic09, ZJ03]. value-added [ZJ03]. values [FK08, FLK^{+}09, HWRC08]. Variability [WMM02]. variables [ABAA09, BAS04, BS06b, GS06a, JDFV08, KS04]. variation [ET05, LBM^{+}04]. variations [Kum09b, Ree02]. varying [Gin02c]. VAST [AMUFVI09]. VB [GS05a]. VDE [GD08]. vector [Shi06]. Vegas [Rol06b]. vehicle [SS07, WBPW06]. vehicles [MH00]. veil [Mar05d]. Venus [Lis05b]. verge [Gin04e]. verification [TIKT04]. VERKKOKE [AL07a, LL06b]. Versatile [KRTT01, BP00]. Version [Fre04, CKM07, RW05]. versus [Cle00a, Mai09, PFVI00, Ras03]. Very [Cas02]. veterans [DWY03]. via [FNW02, Joe06b, KK09, Kaw07, Nod04, NR05, PBDWL09, Rol03a, WW09, ZJ03]. viability [MFRW07]. vibes [Tur08a, Tur09]. Victoria [Ell00b]. Video [Bow00, wH09b, ADP01, Cha07, Got08, LAP08, SFM^{+}07, SJS^{+}07, WP07]. videos [RLM08]. view [BFD^{+}09, CPT02, Noo00, PHM^{+}01, Pea02, Pij00]. viewed [Fle01]. viewers [CH06, CHJB07]. viewing [KRTT01, Lau06]. viewpoints [KI08]. views [BH04, DSM01, TBAA09, Wal05d]. Virginia [Imp04, JKDN04, CTP^{+}04]. ’Virtual [DMKN02, AC08, Bow06, BD01, BBS06, CF02, DG05, GPL^{+}05, Mon02, Sch09b, TBAK04, Way06, BR05b, BCB09, CM05a, Cha01, Cha02, CD01b, FMJM02, GEVZ09b, GM06a, GC05, Har04, HKL^{+}02, Imp01, KV08, LRKM07, MKS02, NV05, NL07, PNF00, RBC07, Rod02, SPAC^{+}05, SAM09, TBMB09, Wal05a, WWWG08, WWGT07, ZE01]. virtual-reality [ZE01]. virtualization [AJD09, Bor07, GLA^{+}08]. Virtualized [WKC^{+}09]. Virtualizing [Mal09]. Viruses [AB05]. visibility [CHUB08]. visible [Mur07a]. Visio [Shu05]. vision [AM07, Ols06]. vision-impaired [AM07]. Visions [Haw07]. visitor [Nef04]. visits [Gor00a, TBMC08]. Visual [Bel02, Bha00, Sie06, Ste09b, SN02, Bla09d, CWHH05, CSA^{+}07, CFR04, Cun02, EG08, Fei07, FS02, LR08, Mos03, Pan09, RM07, She08, SC08, Ver05, HM02, KTKP02]. VisualGraph [LNR03]. visualisation [Boi09, Ibb02, KPT03, MFR^{+}08]. visualisations [MFR^{+}09]. visualising [Geo00]. Visualization [Ran00, All03, AMUFVI09, And00, BCHS00, BBE06, BL06b, BGSL06, CCH03a, CCJ^{+}02, CN07, DCP01, DDC^{+}03a, EVS07, FCA05, GCGBUFVI05, Geo02, HKL^{+}02, HR00, JDGAGMGC05, KS00, KH01, MLK07, NCK^{+}03, NRA^{+}03, NRB^{+}05, NRRR06, NR05, PP03b, SDE01, SB07a, SCE07, She03b, Sut00, VIPCUF09]. visualization-based [AMUFVI09]. visualizations [Bur08, HCMM00, HCB04, KB04b, KLOT05, LJMV07, Lau06, NEN00, RN02, RNH^{+}06, SSMN04, SJH06, Yeh02]. visualized [BGSL06]. Visualizing [BM07b, RL09, UFGCGBVI06, WR07, ABAA09, CPJ07, CGI05, TZ01]. visually [CGB08, CMS06c]. voice [Cun05, RBG02]. void’ [Dic07c]. volume [SL03]. volumes [IHL^{+}04]. volunteers [CRPZ06]. votes [Mar08]. voting [AM05, Esp08, Got06a, Jon05, PL05, VT01]. VR [AH08a, Eps01, Sta05, ZF04]. vs [Bri06, Car03, DFS03, Gha05b, Hen07a, Lau06, Mir09, PM02b, SDLdQ05, SGKL^{+}01, SM02c, TLH^{+}09, VZGE07, WJDL04]. Vulnerabilities [SA07, PT03].

W [GOB^{+}07, GC02]. WADEIn [BL06b]. walk [Bot05, CMRR04]. walking [Bot05, DMKS01]. walks [Bra05]. wander [SA07]. Wanted [VR04, GGH03]. WAPPEN [Kag09]. war [KSE02]. warfare [Wal05a]. Wari [Sco03]. Warren [Rol02]. warrior [WW00]. warriors [TRR03]. washing [Boy08]. watch [Got01]. water [BSV04, SMH05]. waterfall [Cle03b]. way [CC02, Dic07b, God06, Got09b, KPCM06, Luq05, MC08, RW05, Shu04, WMP08]. ways [BW06, Jon02]. WCCE [Tur08a]. weak [LKPC04]. wealth [CF04a, CF06]. wearable [LNCC09]. Weathering [EBG^{+}05]. Weaving [Mur07b, CPT^{+}09b]. Web [AFJ^{+}03, AFR08, AOMC07, AO06b, Ber02, Boi06, BGGR01, BB01, Cha00b, Che00, Dia05, DAJDRG08, DV07, EPQ08, Ekm06, EMM00, Ell02, Est01, Fow01, FdMFPT07, FPQ^{+}08, GS06a, GSW08, Gou06, GRLDPG^{+}01, Had01, Hal08, Ham07, Ham04a, HBIG03, Hic04, HHJ08, Hum04, JFH00, Kag09, KRTT01, KK02, KMSB08, Kla00, KvD02, KMMS02, Kum04, KMN^{+}03, KFE^{+}07, LCB^{+}03, Lee03a, LJM05, Lor00, MSTS02, NEN00, NM02, OCLL00, PDS07, Pap03d, PMC05, RN03, RC02, RJ03, Ros01, Ros06, Rya01, SL06, Sch09a, SNH08, SPM00, SMR03, SMK09, SS07, TIKT04, TSO05, TT00, Tre02, TBR05, UFVI05, pYB08, YLYX06, YD04, YB01, Cul00]. Web-Based [Kum04, AFJ^{+}03, BB01, Che00, Dia05, Ekm06, Est01, Fow01, GS06a, GSW08, Had01, Ham07, JFH00, Kag09, KK02, KvD02, LCB^{+}03, MSTS02, NEN00, Ros01, SNH08, SPM00, SMR03, TIKT04, TT00, Tre02, UFVI05, YB01]. Web-CAT [EPQ08]. Web-centric [DV07]. Web-environment [TSO05]. Web-scale [KMSB08]. Webreg [AR08]. Website [ESS^{+}07, RIP^{+}09]. WebTasks [RH08b]. Webware [BCF^{+}00]. WeBWorK [BCE06]. Webworlds [Cha00a]. week [BFMC05]. Weighted [McG09]. WEKA [MR06]. Well [Lee02d, DABJ00, FLS04b, LD04, Ree02, SSS07]. well-being [DABJ00]. well-tempered [Ree02]. we’re [AKCC06, GGH03, PB09b]. We’ve [McC05c]. Where [Gri02b, Lee02c, vdK07, Lis08a]. Which [ABH^{+}07, CMD08, Etl06]. while [PW01]. whistle [Bri09]. whistle-blower [Bri09]. white [Kac08b]. whiteboard [BJJ01, Cas09]. Whither [BDRH09]. who [CCL^{+}07, Deb02, Har05, Tow02]. whole [GS06b]. whom [HV05]. whose [BBS06]. wide [GP02]. widen [BGD03b, Bar05]. Wiki [Ber02, O’N05]. Wikis [GCDJ08]. Will [CWL08, Deb02, LR07, REC^{+}01]. Windows [wH09a, wH09b, PP06, Shu06b, TW05, wH07]. winds [LM03]. winning [Got00b, Kno01]. WIPE [EDES05]. Wired [JKKL04]. Wireless [JKKL04, BHH^{+}08, Bri06, RW02, Ric02, RS04, Ste09a, SAM09, TY06]. wireshark [HL08]. wish [APGR07]. within [CKLW05, CT07, Kum09b, Lee02b]. without [ACSS06, Bec03a, COR07a, GP00, Haw07, KC07, MFL01, PF03]. Wits [HRS06, RSK06, SG07, SL08a]. Women [Coh02a, Gal02, Gür02b, HDA^{+}02, MRM^{+}06, PHW07, RPR03, Tay02b, Tea02, AHLY03, ABO^{+}05, AL05b, AL06, BGD03a, Bec07, CL02, Coh02b, CWC09, CA02, DPBP09, Edm08a, Ega07, FM03b, Fri05, FHBD06, GM02, GOB^{+}07, GC02, Gür02a, KSE02, Kla05b, Lee02d, LS02, MHR06, Mar05d, PTC^{+}01, PPJH09, Pow08, PM02b, PT02, RPG04, Ric09, RKI02, WB02, ZIBG05, FB02, FHBD06]. women-CS [FHBD06]. words [Lew07b]. Work [SM02c, BS05, BC07e, Dei05, Dic07c, FMKQ08, Joy02, LR07, LBA07, LCL08, MC04, MFC00, NBF^{+}08, Pop08b, REC^{+}01, WJDL04, Wal07a]. workbench [MSK09, SOSL04]. worked [LCL08]. workflow [BCC^{+}05]. workforce [CHK03b, GPF^{+}01, Mar05c, Rob00d, WB02]. Working [ABE^{+}09b, CCC07, CS06, Cle05a, HBEM03, McC05c]. workloads [Dou00]. Works [MHK^{+}03b, FMK^{+}00, Kla05b, PGC^{+}06]. workshop [BBF^{+}09, BZ05, CA02, Doo09, Haz05, SL08b, ZV04]. workshops [CMS07]. world [AKWZ07, AL02, ALH07, BA08, BN00, DCP00, DDL03, ES08, EA07, FIG07, Gal02, GKP06, GNPL01, Gup06, Har05, Joe06b, Kaw07, Laz02b, Laz02a, LD04, PH04c, RF00a, SW06b, SAM09, ZB08]. worlds [Ada07, AJD09, Rod02, ZE01]. worldwide [TT00]. worth [Lew07b]. woven [MH05]. writable [KMN^{+}03]. Write [LH02, Edm08a, GY01b]. WriteOn [TEP06]. WriteOn1.0 [CTP09]. Writing [BAWRMM02, DMKS01, Deb04, HDH06b, Hol08, Kac03, KKLK04, LFT09, Osb06, SL09a, SM07, Van04]. writ’n [Boo07]. written [Gha05b, TETPQ08, TZ01]. wrong [Gin08b]. Wu [EB08]. WWW [CI03, EA00, GP02, Ibb02, Nør00, RF00c]. WWW-based [EA00, Nør00]. x [Jen01a]. Xen [AJD09]. XML [AT04, BK03, DSCU01, DUM^{+}05, Hei03, KB01, NRB^{+}05, PMC05, WM03a]. XML-based [NRB^{+}05]. XNA [LS08a]. XP [LW06]. XQuery [MPD^{+}05].

Yahoo [RHC07]. ye [McC02]. yeah [Cle06a]. Year [DHRH05, HCK07, AWS^{+}09, Cam01, Cam02b, Cam03a, CHK03a, CHK03b, Cam04a, Cam04b, CHK05, Cam05b, CHK06, CS02, CLP06, CN07, DLM09, Edm07a, Edm09, EGMS08, GR00a, HAB^{+}00, HV07, HRS06, Hov05, How05, Kle01, LL06a, LS01, MGS07, MAD^{+}01a, MAD^{+}01b, MAFLR05, MAFLR06, MSS^{+}04, OBLH03, Pop04, Pow08, Ras00, RRKS06, Rio02, SCHS09, SM00a, SDM^{+}02, SHVP05, STS09, SHS^{+}08, Sur02, TFG05, TRT04, WB06, XSD07]. years [CL02, Gor00b, Wil07]. yield [Cle07b]. York [ESS^{+}07]. You’d [ABE^{+}02]. Young [PPJH09]. yours [YPdMA^{+}08]. youth [MPK^{+}08a]. YouTube [Mar07a].

zipper [HK01]. zones [ABE^{+}09b]. zoom [Gin01e]. zoom-in [Gin01e].

References

Aly:2004:CSP

[AA04] Alaaeldin A. Aly and Shakil Akhtar. Cryptography and security protocols course for undergraduate IT students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(2):44–47, June 2004. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

               tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
               ftp.ira.uka.de/bibliography/Misc/DBLP/2004.bib.

Amershi:2005:DCP

[AAC^{+}05] S. Amershi, N. Arksey, G. Carenini, C. Conati, A. Mackworth, H. Maclaren, and D. Poole. Designing CIspace: pedagogy and usability in a learning environment for AI. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):178–182, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Anderson:2007:SAL

[AAD^{+}07] Richard Anderson, Ruth Anderson, K. M. Davis, Natalie Linnell, Craig Prince, and Valentin Razmov. Supporting active learning and example based instruction with classroom technology. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):69–73, March 2007.

CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Almstrum:2008:PHC

[AAEI08] Vicki L. Almstrum, Bill Aspray, Mary Anne Egan, and John Impagliazzo. Perspectives on history in computing and education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):2–3, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Alt:2006:SNG

[AAF^{+}06] Casey Alt, Owen Astrachan, Jeffrey Forbes, Richard Lucic, and Susan Rodger. Social networks generate interest in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):438–442, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Almstrum:2008:CEO

[AAO^{+}08] Vicki L. Almstrum, E. Anne G. Applin, Barbara Boucher Owens, Elizabeth Adams, Lecia J. Barker, John Impagliazzo, Patricia A. Joseph, Amardeep Kahlon, Mary Z. Last, Andrea Lawrence, and Alison Young. Computing educators oral history project: seeking the trends. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (4):122–141, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Anderson:2004:ETP

[AAS^{+}04] Richard Anderson, Ruth Anderson, Beth Simon, Steven A. Wolfman, Tammy VanDeGrift, and Ken Yasuhara. Experiences with a tablet PC based lecture presentation system in computer science courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):56– 60, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Anderson:2003:CPT

[AAV^{+}03] Richard J. Anderson, Ruth Anderson, Tammy VanDeGriff, Steven A. Wolfman, and Ken Yasuhara. Classroom presentation from the tablet PC. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):238, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Aycock:2005:V

[AB05] J. Aycock and K. Barker. Viruses 101. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):152–156, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Adams:2008:MBC

[AB08] Joel C. Adams and Tim H. Brom. Microwulf: a Beowulf cluster for every desk. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):121–125, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Abad:2008:LTC

[Aba08] Cristina L. Abad. Learning through creating learning objects: experiences with a class project in a distributed systems course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):255–259, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Armoni:2009:CND

[ABA09] Michal Armoni and Mordechai Ben-Ari. The concept of nondeterminism: its development and implications for teaching. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):141–160, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Al-Barakati:2009:EVR

[ABAA09] Nouf M. Al-Barakati and Arwa Y. Al-Aama. The effect of visualizing roles of variables on student performance in an introductory programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (3):228–232, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Al-Bow:2009:UGC

[ABAE^{+}09] Mohammed Al-Bow, Debra Austin, Jeffrey Edgington, Rafael Fajardo, Joshua Fishburn, Carlos Lara, Scott Leutenegger, and Susan Meyer. Using game creation for teaching computer programming to high school students and teachers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):104–108, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Adams:2003:EPG

[ABB03a] Joel C. Adams, Vimala Bauer, and Shakuntala Baichoo. An expanding pipeline: gender in Mauritius. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):59–63, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Andersen:2003:TPL

[ABB^{+}03b] Peter Bøgh Andersen, Jens Bennedsen, Steffen Brandorff, Michael E. Caspersen, and Jesper Mosegaard. Teaching programming to liberal arts students: a narrative media approach. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):109–113, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Adams:2006:ATP

[ABB^{+}06] Elizabeth Adams, Doug Baldwin, Judith Bishop, John English, Pamela Lawhead, and Daniel Stevenson. Approaches to teaching the programming languages course: a potpourri. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):299–300, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Alm:2002:YBS

[ABE^{+}02] Jacob Alm, Robert Baber, Shauna Eggers, Christopher O’Toole, and Abin Shahab. You’d better set down for this!: creating a set type for CS1 & CS2 in C#. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):14–18, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Altin:2009:UIT

[ABE^{+}09a] Rukiye Altin, Mustafà Bektik, Nuray Ekşioğlu, Can Koray, Ömer Canbek Öner, Merve Sadeta, Hilal Şener, Duygu Simsek, Ching-Chen Ma, Christine Price, and Christopher R.

Routh. Use of intuitive tools to enhance student learning and user experience. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):365, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Altin:2009:WAT

[ABE^{+}09b] Rukiye Altin, Mustafa Bektik, Nuray Eksioglu, Can Koray, Omer Canbek Oner, Merve Sadetas, Hilal Sener, Duygu Simsek, Ching-Chen Ma, Christine Price, and Christopher Routh. Working across time zones in cross-cultural student teams. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):360, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Adcock:2007:WPE

[ABH^{+}07] Bruce Adcock, Paolo Bucci, Wayne D. Heym, Joseph E. Hollingsworth, Timothy Long, and Bruce W. Weide. Which pointer errors do students make? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (1):9–13, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Alphonce:2001:TUS

[ABK^{+}01] Carl Alphonce, Debra Burhans, Helene Kershner, Barbara Sherman, Deborah Walters, Erica Eddy, Gloria Melara, Pete DePasquale, J. Philip East, Fred Springsteel, and Kurt F. Lauckner. Teaching using off-the shelf on-line materials. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):429–430, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Astrachan:2005:ROE

[ABK^{+}05] Owen Astrachan, Kim Bruce, Elliot Koffman, Michael Kölling, and Stuart Reges. Resolved: objects early has failed. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):451–452, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Adelmann:2007:IFF

[ABL07] Robert Adelmann, Tobias Bischoff, and Tobias Lauer. IDEA: a framework for the fast creation of interactive animations by pen sketching. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):291–295, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Almstrum:2005:BSH

[ABO^{+}05] Vicki L. Almstrum, Lecia Jane Barker, Barbara Boucher Owens, Elizabeth Adams, William Aspray, Nell B. Dale, Wanda Dann, Andrea Lawrence, and Leslie Schwartzman. Building a sense of history: narratives and pathways of women computing educators. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4): 173–189, December 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2005.bib.

Abowd:2009:MIM

[Abo09] Gregory D. Abowd. Make IT matter: how computing can make a difference. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):553, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Alvarez:2008:CHE

[ABS08] Marco A. Alvarez, José Baiocchi, and José Antonio Pow Sang. Computing and higher education in Peru. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):35–39, June 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2008.bib.

Abunawass:2004:CCPb

[Abu04] Adel Abunawass. COMPASS: a CS program assessment project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):269, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Albin-Clark:2007:MTM

[AC07] Adrian Albin-Clark. MaBL: a tool for mapping pseudocode to multiple implementation languages. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (3):315, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Albin-Clark:2008:VCE

[AC08] Adrian Albin-Clark. Virtual chat in an enquiry-based team project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):153–157, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Alphonce:2007:KKE

[ACD07] Carl Alphonce, Michael Caspersen, and Adrienne Decker. Killer “killer examples” for design patterns. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):228–232, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Aycock:2008:SIL

[ACd08] John Aycock, Heather Crawford, and Rennie deGraaf. Spamulator: the Internet on a laptop. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (3):142–147, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Astrachan:2000:RCA

[ACK^{+}00] Owen Astrachan, Robert Cartwight, Rich Kick, Cay Horstmann, Fran Trees, Gail Chapman, David Gries, Henry Walkers, and Ursula Wolz. Recommendations for changes in advanced placement computer science (panel session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):416, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Albin-Clark:2008:URP

[ACK08] Adrian Albin-Clark and T. R. Vishnu Arun Kumar. The use of role play to simulate a tethered swarm of robots for urban search and rescue (USAR). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):335, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Abraham:2001:LTP

[ACL^{+}01] David Abraham, Liz Crawford, Leanna Lesta, Agathe Merceron, and Kalina Yacef. The logic tutor (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):184, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Alexander:2003:CSA

[ACL^{+}03] Sylvia Alexander, Martyn Clark, Ken Loose, June Amillo, Mats Daniels, Roger D. Boyle, Cary Laxer, and Dermot Shinners-Kennedy. Case studies in admissions to and early performance in computer science degrees. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):137–147, December 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

ACM:2001:PTS

[ACM01] ACM, editor. Proceeding of the Thirty-second SIGCSE Technical Symposium on Computer Science Education (SIGCSE-01), 21–25 February 2001, Charlotte, North Carolina, volume 33(1) of SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

ACM:2006:PST

[ACM06] ACM, editor. Proceedings of the thirty-seventh SIGCSE Technical Symposium on Computer Science Education: SIGCSE 2006: Houston, TX, USA, March 1–5, 2006, SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 2006. CO-DEN SIGSD3. ISBN 1-59593-259-3. ISSN 0097-8418 (print), 2331-3927 (electronic). LCCN QA76.27. ACM order number 457060.

Allen:2003:PPC

[ACR03] Eric Allen, Robert Cartwright, and Charles Reis. Production programming in the classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35 (1):89–93, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Allen:2002:DLP

[ACS02] Eric Allen, Robert Cartwright, and Brian Stoler. DrJava: a lightweight pedagogic environment for Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):137–141, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Angel:2006:TCG

[ACSS06] Edward Angel, Steve Cunningham, Peter Shirley, and Kelvin Sung. Teaching computer graphics without raster-level algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):266–267, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Aedo:2001:ASE

[AD01] Ignacio Aedo and Paloma Díaz. Applying software engineering methods for hypermedia systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):5–8, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Astrachan:2008:IOS

[AD08] Owen Astrachan and Peter Denning. Innovating our self image. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):178–179, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Adams:2006:OJP

[Ada06] Joel C. Adams. OOP and the Janus Principle. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):359–363, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Adams:2007:AMS

[Ada07] Joel C. Adams. Alice, middle schoolers & the imaginary worlds camps. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):307–311, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Adams:2009:TDD

[Ada09] Joel Adams. Test-driven data structures: revitalizing CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):143–147, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Almstrum:2001:STF

[ADG^{+}01] Vicki L. Almstrum, C. Neville Dean, Don Goelman, Thomas B. Hilburn, and Jan Smith. Support for teaching formal methods. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(2):71– 88, June 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2001.bib.

Adams:2003:CTC

[ADG^{+}03] Liz Adams, Mats Daniels, Annegret Goold, Orit Hazzan, Kathy Lynch, and Ian Newman. Challenges in teaching capstone courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):219–220, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Appel:2009:MMB

[ADM^{+}09] Florence A. Appel, Katherine Deibel, C. Dianne Martin, Joseph D. Oldham, Tarsem S. Purewal, Jr., and Carol L. Spradling. From the man on the Moon to 2001 and beyond: the evolving social and ethical impact of computers a session to commemorate SIGCSE’s 40th anniversary. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):333–334, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Anderson:2001:ETV

[ADP01] Richard Anderson, Martin Dickey, and Hal Perkins. Experiences with tutored video instruction for introductory programming courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):347–351, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Adams:2001:TNO

[AE01] D. Robert Adams and Carl Erickson. Teaching networking and operating systems to information systems majors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):85–89, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ahmadzadeh:2005:APD

[AEH05] Marzieh Ahmadzadeh, Dave Elliman, and Colin Higgins. An analysis of patterns of debugging among novice computer science students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):84–88, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Agarwal:2006:DAL

[AEPQ06] Rahul Agarwal, Stephen H. Edwards, and Manuel A. Pérez-Quiñones. Designing an adaptive learning module to teach software testing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):259–263, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Allevato:2009:DEP

[AEPQ09] Anthony Allevato, Stephen H. Edwards, and Manuel A. Pérez-Quiñones. Dereferee: exploring pointer mismanagement in student code. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):173–177, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Acharya:2003:LHM

[AF03a] Mithun Acharya and Robert Funderlic. ‘Laurel and Hardy’ model for analyzing process synchronization algorithms and primitives. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):107–110, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

Adams:2003:OCD

[AF03b] Joel Adams and Jeremy Frens. Object centered design for Java: Teaching OOD in CS-1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):273–277, January 2003. CODEN SIGSD3. ISBN 1-58113-648-X. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Compiler/java.bib.

Akingbade:2003:JEW

[AFJ^{+}03] Ayonike Akingbade, Thomas Finley, Diana Jackson, Pretesh Patel, and Susan H. Rodger. JAWAA: easy Web-based animation from CS 0 to advanced CS courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):162–166, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Amelung:2008:TGF

[AFR08] Mario Amelung, Peter Forbrig, and Dietmar Rösner. Towards generic and flexible Web services for e-assessment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):219–224, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Armoni:2004:AHS

[AGE04] Michal Armoni and Judith Gal-Ezer. On the achievements of high school students studying computational models. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):17–21, September 2004. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Armoni:2005:RTU

[AGE05] Michal Armoni and Judith Gal-Ezer. Reductive thinking in undergraduate CS courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):362, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Armoni:2006:RAT

[AGE06] Michal Armoni and Judith Gal-Ezer. Reduction — an abstract thinking pattern: the case of the computational models course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):389–393, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Armoni:2006:RTU

[AGEH06] Michal Armoni, Judith Gal-Ezer, and Orit Hazzan. Reductive thinking in undergraduate CS courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):133–137, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Adams:2004:MID

[AGGR04] Elizabeth S. Adams, Mary Granger, Don Goelman, and Catherine Ricardo. Managing the introductory database course: what goes in and what comes out? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):497–498, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Almstrum:2003:TCS

[AGHC03] Vicki L. Almstrum, David Ginat, Orit Hazzan, and John M. Clement. Transfer to/from computing science education: the case of science education research. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35 (1):303–304, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Almstrum:2002:IEC

[AGHM02] Vicki L. Almstrum, David Ginat, Orit Hazzan, and Tom Morley. Import and export to/from computing science education: the case of mathematics education research. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):193–194, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Adams:2008:BEV

[AH08a] Joel C. Adams and Joshua Hotrop. Building an economical VR system for CS education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):148– 152, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Applin:2008:CRM

[AH08b] Anne G. Applin and Hilary J. Holz. Computing research methods multi-perspective digital library: a call for participation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):337, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Aharoni:2000:CES

[Aha00] Dan Aharoni. Cogito, Ergo sum! cognitive processes of students dealing with data structures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (1):26–30, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Almstrum:2005:CCS

[AHGP05] Vicki L. Almstrum, Orit Hazzan, Mark Guzdial, and Marian Petre. Challenges to computer science education research. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):191–192, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Almstrum:2006:CIC

[AHH^{+}06] Vicki L. Almstrum, Peter B. Henderson, Valerie J. Harvey, Cinda Heeren, William A. Marion, Charles Riedesel, Leen-Kiat Soh, and Allison Elliott Tew. Concept inventories in computer science for the topic discrete mathematics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):132–145, December 2006. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2006.bib.

Applin:2007:MPD

[AHJ^{+}07] Anne Gates Applin, Hilary J. Holz, William Joel, Ifeyinwa Okoye, Katherine Deibel, Becky Grasser, Briony J. Oates, and Gwendolyne Wood. A multi-perspective digital library to facilitate integrating teaching research methods across the computing curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):184– 203, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2007.bib.

Andrianoff:2004:AOT

[AHL04] Steven K. Andrianoff, Dalton R. Hunkins, and David B. Levine. Adding objects to the traditional ACM programming contest. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):105–109, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Adams:2003:IPW

[AHLY03] Elizabeth S. Adams, Orit Hazzan, Hrafn Loftsson, and Alison Young. International perspective of women and computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):45–46, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Argollo:2004:GSL

[AHM^{+}04] Eduardo Argollo, Mauricio Hanzich, Diego Mostaccio, Germán Bianchini, Paula Fritzsche, Ferran Bonàs, Emilio Luque, Juan Carlos Moure, and Dolores Rexachs. Graduate students learning strategies through research collaboration. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):262, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Aho:2008:TCC

[Aho08] Alfred V. Aho. Teaching the compilers course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):6–8, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Astrachan:2009:PFC

[AHPS09] Owen Astrachan, Susanne Hambrusch, Joan Peckham, and Amber Settle. The present and future of computational thinking. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):549–550, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Anderson:2009:XWL

[AJD09] Benjamin R. Anderson, Amy K. Joines, and Thomas E. Daniels. Xen worlds: leveraging virtualization in distance education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):293–297, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Austin:2007:SMS

[AJF07] Shermane A. Austin, Leon P. Johnson, and John M. Flowers. Stimulating minority student retention with BalloonSAT projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):320, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Ahoniemi:2009:AUR

[AK09] Tuukka Ahoniemi and Ville Karavirta. Analyzing the use of a rubric-based grading tool. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):333– 337, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Auvinen:2009:ROA

[AKA09] Tapio Auvinen, Ville Karavirta, and Tuukka Ahoniemi. Rubyric: an online assessment tool for effortless authoring of personalized feedback. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):377, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Almstrum:2006:TLW

[AKCC06] Vicki L. Almstrum, David Klappholz, Steven Condly, and John Clement. Are they learning what (we think) we’re teaching? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):195, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Aiken:2000:FIT

[AKM00] Robert Aiken, Ned Kock, and Munir Mandviwalla. Fluency in information technology: a second course for non-CIS majors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):280–284, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Alhalabi:2007:SER

[AKT07] Wadee S. Alhalabi, Miroslav Kubat, and Moiez A. Tapia. Search engine ranking efficiency evaluation tool. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):97–101, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2007.bib.

Allen:2007:PLD

[AKWZ07] Robert B. Allen, David Klappholz, Michael R. Wick, and Carol Zander. Programming-lite: a dialog on educating computer science practitioners in a “flat world”. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):243–244, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Azadegan:2001:ICP

[AL01] Shiva Azadegan and Chao Lu. An international common project: implementation phase. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3): 125–128, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Andrianoff:2002:RPO

[AL02] Steven K. Andrianoff and David B. Levine. Role playing in an object-oriented world. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):121– 125, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Adams:2005:CMC

[AL05a] Joel C. Adams and W. David Laverell. Configuring a multicourse lab for system-level projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37 (1):525–529, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Almstrum:2005:WAW

[AL05b] Vicki L. Almstrum and Mary Z. Last. What attracts women to CS? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):378, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Almstrum:2006:MTW

[AL06] Vicki L. Almstrum and Mary Z. Last. Men are from toys: women are from tools. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):313, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Alstes:2007:VLR

[AL07a] Anton Alstes and Janne Lindqvist. VERKKOKE: learning routing and network programming online. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):91–95, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Anshel:2007:RME

[AL07b] Michael Anshel and Sarah Levitan. Reducing medical errors using secure RFID technology. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (2):157–159, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2007.bib.

Anderson:2009:ETB

[AL09] Nicole Anderson and Chi-Cheng Lin. Exploring technologies for building collaborative learning communities among diverse student populations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3): 243–247, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Albrecht:2009:BBS

[Alb09] Jeannie R. Albrecht. Bringing big systems to small schools: distributed systems for undergraduates. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):101–105, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Armoni:2008:TST

[ALBA08] Michal Armoni, Noa Lewenstein, and Mordechai Ben-Ari. Teaching students to think nondeterministically. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):4–8, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Arnold:2007:ITI

[ALH07] Ruedi Arnold, Marc Langheinrich, and Werner Hartmann. InfoTraffic: teaching important concepts of computer science and math through real-world examples. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):105–109, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Allert:2003:CTA

[All03] James Allert. A companion technology approach to CS1: handheld computers with concept visualization software. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):134–138, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Azadegan:2003:UTC

[ALO^{+}03] S. Azadegan, M. Lavine, M. O’Leary, A. Wijesinha, and M. Zimand. An undergraduate track in computer security.

SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):207–210, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Abunawass:2004:CCPa

[ALR04] Adel Abunawass, Will Lloyd, and Edwin Rudolph. COM-PASS: a CS program assessment project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):127–131, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ahoniemi:2008:IPF

[ALR08] Tuukka Ahoniemi, Essi Lahtinen, and Tommi Reinikainen. Improving pedagogical feedback and objective grading. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):72–76, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Amillo:2008:PAS

[ALRY08] June Amillo, Cary Laxer, Ernestina Menasalvas Ruiz, and Alison Young, editors. Proceedings of the 13th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education, ITiCSE 2008, Madrid, Spain, June 30–July 2, 2008, SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 2008. CODEN SIGSD3. ISBN 1-60558-078-3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Alvim:2005:TBI

[Alv05] Duarte Alvim. Taking the bus in the information highway system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):403, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Armen:2005:TAR

[AM05] Chris Armen and Ralph Morelli. Teaching about the risks of electronic voting technology. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):227– 231, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Armstrong:2007:RLD

[AM07] Helen Armstrong and Iain Murray. Remote and local delivery of Cisco education for the vision-impaired. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):78–81, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Almeida-Martinez:2009:VVB

[AMUFVI09] Francisco J. Almeida-Martínez, Jaime Urquiza-Fuentes, and J. Ángel Velázquez-Iturbide. VAST: a visualization-based educational tool for language processors courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):342, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Aspray:2006:ERO

[AMVZ06] William Aspray, A. Frank Mayadas, Moshe Y. Vardi, and Stuart H. Zweben. educational response to offshore outsourcing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):330–331, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Axelsson:2006:SAS

[AMW06] Karin Axelsson, Ulf Melin, and Tommy Wedlund. Student activity in seminars: designing multi-functional assessment events. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):93–97, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Anderson:2000:AVU

[And00] Jay Martin Anderson. Algorithm visualization using Quick-Time movies for student interaction (poster session).: algorithms from computational geometry. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):185, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Anderson:2007:CSP

[And07] Scott D. Anderson. A course on simulation, probability and statistics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):110–114, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Anonymous:2003:PTF

[Ano03] Anonymous, editor. Proceedings of the Thirty-Fourth SIGCSE Technical Symposium on Computer Science Education: SIGCSE 2003: Reno, Nevada, USA, February 19–22, 2003. ACM Press, New York, NY 10036, USA, 2003. CO-DEN SIGSD3. ISBN 1-58113-648-X. ISSN 0097-8418 (print), 2331-3927 (electronic). LCCN QA76.27 .A34 2003.

Adams:2000:PCS

[ANS00] Joel Adams, Chris Nevison, and Nan C. Schaller. Parallel computing to start the millennium. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (1):65–69, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Amzad:2006:MBP

[AO06a] Imtiaz Amzad and Arturo Jose Ortiz. Model based project centered team learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):328, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Assuncao:2006:TWS

[AO06b] Luís Assunção and António Luís Osório. Teaching Web services using .NET platform. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):339, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Arbe:2007:FLT

[AOMC07] José Miguel Blanco Arbe, Ana Sánchez Ortega, and Jesús Ibáñez Mártinez-Conde. Formal languages through Web forms and regular expressions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (4):100–104, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2007.bib.

Astrachan:2007:TTW

[APGR07] Owen Astrachan, Nick Parlante, Daniel D. Garcia, and Stuart Reges. Teaching tips we wish they’d told us before we started. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):2–3, March 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Applin:2001:SLA

[App01] Anne Gates Applin. Second language acquisition and CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):174–178, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Applin:2006:LCA

[App06] Anne G. Applin. A learner-centered approach to teaching ethics in computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):530–534, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Amelung:2006:EEA

[APR06] Mario Amelung, Michael Piotrowski, and Dietmar Rösner. EduComponents: experiences in e-assessment in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):88–92, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Amelung:2007:ECB

[APR07] Mario Amelung, Michael Piotrowski, and Dietmar Rösner. eduComponents: a component-based e-learning environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):352, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Abi-Raad:2000:SAA

[AR00a] Maurice Abi-Raad. Systems analysis with attitude! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):57–60, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Andersson:2000:ESL

[AR00b] Roy Andersson and Torgny Roxå. Encouraging students in large classes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):176–179, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Aitenbichler:2008:WOS

[AR08] Erwin Aitenbichler and Guido Rößling. Webreg: an online system for registration and grade distribution. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):336, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Arif:2000:MTO

[Ari00] Essam M. Arif. A methodology for teaching object-oriented programming concepts in an advanced programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(2):30–34, June 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Armoni:2004:RTH

[Arm04] Michal Armoni. Reductive thinking: how to teach it? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):280, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Armoni:2006:RPC

[Arm06] Michal Armoni. On the role of proofs in a course on design and analysis of algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4): 39–42, December 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2006.bib.

Armoni:2008:RTQ

[Arm08] Michal Armoni. Reductive thinking in a quantitative perspective: the case of the algorithm course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):53–57, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Arnold:2007:IPL

[Arn07] Ruedi Arnold. Introducing propositional logic and queueing theory with the InfoTraffic interactive learning environments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):356, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Arshad:2009:TPP

[Ars09] Naveed Arshad. Teaching programming and problem solving to CS2 students using think-alouds. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (1):372–376, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Armoni:2006:ATR

[ARVV06] M. Armoni, S. Rodger, M. Vardi, and R. Verma. automata theory: its relevance to computer science students and course contents. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):197–198, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Atkin:2002:PEO

[AS02] Benjamin Atkin and Emin Gün Sirer. PortOS: an educational operating system for the post-PC environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):116–120, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Al-Salman:2000:CSE

[ASA00] Abdulmalik S. Al-Salman and Jacob Adeniyi. Computer science education in a Saudi Arabian university: a comparative study of its B.Sc. program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(4): 34–39, December 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ainsworth:2008:MMCa

[ASA08a] A. Barbara Ainsworth, Judithe Sheard, and Chris Avram. The Monash Museum of Computing History: part 1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):31–34, June 2008. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2008.bib.

Ainsworth:2008:MMCb

[ASA08b] A. Barbara Ainsworth, Judithe Sheard, and Chris Avram. The Monash Museum of Computing History: part 2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):31–34, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Alfonsi:2004:IIT

[ASPS04] C. Alfonsi, N. Scarabottolo, D. Pedreschi, and M. Simi. IT4PS: information technology for problem solving. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):241, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Astrachan:2001:OOW

[Ast01] Owen Astrachan. OO overkill: when simple is better than not. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):302–306, March 2001. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Astrachan:2003:BSA

[Ast03] Owen Astrachan. Bubble sort: an archaeological algorithmic analysis. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):1–5, January 2003.

CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Astrachan:2004:WCA

[Ast04] Owen Astrachan. Why i care about programming and how to teach it. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):235, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Astrachan:2007:CAS

[Ast07a] Owen L. Astrachan. Cutting apron strings. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):19–20, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2007.bib.

Astrachan:2007:HC

[Ast07b] Owen L. Astrachan. Head in the clouds. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):23, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2007.bib.

Astrachan:2009:BCE

[Ast09a] Owen Astrachan. Out-of-the-box: cogito ergo hack. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):80–81, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Astrachan:2009:PP

[Ast09b] Owen Astrachan. Pander to ponder. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (1):192–196, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Ariga:2001:PSI

[AT01] Taeko Ariga and Hideki Tsuiki. Programming for students of information design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(4):59–63, December 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2001.bib.

Adams:2004:UXC

[AT04] D. Robert Adams and Christian Trefftz. Using XML in a compiler course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):4–6, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Amirijoo:2004:RMR

[ATNT04] Mehdi Amirijoo, Aleksandra Tešanović, and Simin Nadjm-Tehrani. Raising motivation in real-time laboratories: the soccer scenario. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):265–269, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Aycock:2005:CC

[AU05] John Aycock and Jim Uhl. Choice in the classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4):84–88, December 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2005.bib.

Adams:2002:SCS

[AV02a] Joel Adams and David Vos. Small-college supercomputing: building a Beowulf cluster at a comprehensive college. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):411–415, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Alphonce:2002:OOC

[AV02b] Carl Alphonce and Phil Ventura. Object orientation in CS1-CS2 by design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):70–74, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Astrachan:2009:APC

[AWS^{+}09] Owen Astrachan, Henry Walker, Chris Stephenson, Lien Diaz, and Jan Cuny. Advanced placement computer science: the future of tracking the first year of instruction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):397–398, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Allenstein:2008:QSS

[AYWM08] Brett Allenstein, Andrew Yost, Paul Wagner, and Joline Morrison. A query simulation system to illustrate database query execution. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):493–497, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Ben-Ari:2000:HGG

[BA00] Mordechai Ben-Ari. How to get a good review. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(2):4–6, June 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ben-Ari:2001:BDR

[BA01] Mordechai Ben-Ari. The bug that destroyed a rocket. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(2):58–59, June 2001. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2001.bib.

Ben-Ari:2004:STT

[BA04a] Mordechai Ben-Ari. A suite of tools for teaching concurrency. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):251, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ben-Ari:2004:CDF

[BA04b] Mordechai (Moti) Ben-Ari. The Concorde doesn’t fly anymore. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):1, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ben-Ari:2005:CDF

[BA05a] Mordechai (Moti) Ben-Ari. The Concorde doesn’t fly anymore. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):196, March 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ben-Ari:2005:MNC

[BA05b] Mordechai (Moti) Ben-Ari. Minesweeper as an NP-complete problem. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4):39–40, December 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2005.bib.

Ben-Ari:2006:MA

[BA06] Mordechai (Moti) Ben-Ari. McKinley’s Amazon. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):75–77, December 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2006.bib.

Ben-Ari:2007:TCN

[BA07a] Mordechai Ben-Ari. Teaching concurrency and nondeterminism with spin. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):363–364, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Birmingham:2007:SDR

[BA07b] William Birmingham and Leah Acker. Software-defined radio for undergraduate projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):293– 297, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Beaubouef:2008:RQR

[BA08] Theresa Beaubouef and Ghassan Alkadi. Rough querying: a real-world information systems project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):87–91, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bilal:2006:RBN

[BAAM06] Sara Saeed Bilal, Habib Al Mutawa, Fatma Hassan Al Amir, and C. Dianne Martin. Reflections on being a new female IT professional in the UAE. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):6, December 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2006.bib.

Bigham:2008:IBH

[BAB^{+}08] Jeffrey P. Bigham, Maxwell B. Aller, Jeremy T. Brudvik, Jessica O. Leung, Lindsay A. Yazzolino, and Richard E. Ladner. Inspiring blind high school students to pursue computer science with instant messaging chatbots. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):449–453, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Ben-Ari:2004:WDW

[BABBH04] Mordechai Ben-Ari, Anders Berglund, Shirley Booth, and Christian Holmboe. What do we mean by theoretically sound research in computer science education? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):230–231, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Baciu:2003:SIA

[Bac03] Rodica Baciu. SysRel: an interactive application for training. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):268, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bueno:2007:ALA

[BAdC07] F. Javier Bueno, M. Goretti Alonso, and J. Raul Fernandez del Castillo. Assisting lecturers to adapt e-learning content for deaf students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):335, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Beck:2009:BPS

[BAET09] Jon Beck, Vicki L. Almstrum, Heidi J. C. Ellis, and Massood Towhidnejad. Best practices in software engineering project class management. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):201– 202, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Bailey:2005:ITT

[Bai05] Mark W. Bailey. IRONCODE: think-twice, code-once programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):181–185, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barker:2009:TDC

[BAK^{+}09] Lecia J. Barker, Elizabeth Adams, Amardeep Kahlon, Andrea Lawrence, and Allison Young. Trends and discoveries of the computing educators oral history project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):399–400, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Baldwin:2000:STU

[Bal00] Doug Baldwin. Some thoughts on undergraduate teaching and the Ph.D. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(4):60–62, December 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Baldwin:2003:CTA

[Bal03] Doug Baldwin. A compiler for teaching about compilers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):220–223, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bhagyavati:2005:THC

[BAMSK05] Bhagyavati, Stephen O. Agyei-Mensah, Rose Shumba, and Iretta B. C. Kearse. Teaching hands-on computer and information systems security despite limited resources. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):325–326, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barros:2001:UPN

[Bar01] João Paulo Barros. Use Petri nets to improve your concurrent programming course (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):192, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barnes:2002:TIJ

[Bar02a] David J. Barnes. Teaching introductory Java through LEGO MINDSTORMS models. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):147– 151, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Barros:2002:SPU

[Bar02b] João Paulo Barros. Specific proposals for the use of Petri nets in a concurrent programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (3):165–167, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barker:2005:WDG

[Bar05] Lecia J. Barker. When do group projects widen the student experience gap? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):276–280, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ben-Ari:2004:RVS

[BAS04] Mordechai Ben-Ari and Jorma Sajaniemi. Roles of variables as seen by CS educators. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):52–56, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ben-Ari:2002:WT

[BAWRMM02] Mordechai (Moti) Ben-Ari, Henry M. Walker, Gaynor Redvers-Mutton, and Keith Mansfield. Writing a textbook. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):94–95, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bayliss:2009:UGI

[Bay09] Jessica D. Bayliss. Using games in introductory courses: tips from the trenches. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):337–341, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Bishop:2000:OOJ

[BB00] Judith Bishop and Nigel Bishop. Object-orientation in Java for scientific programmers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):357– 361, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Butler:2001:WBL

[BB01] Jucain E. Butler and Jay B. Brockman. A Web-based learning tool that simulates a simple computer architecture. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(2):47–50, June 2001. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2001.bib.

Bruhn:2003:ATJ

[BB03a] Russel E. Bruhn and Philip J. Burton. An approach to teaching Java using computers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35 (4):94–99, December 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Compiler/java.bib;   ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/2003.
bib.

Burton:2003:TPO

[BB03b] Philip J. Burton and Russel E. Bruhn. Teaching programming in the OOP era. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):111– 114, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2003.bib.

Boyle:2006:WDY

[BB06] Roger Boyle and Joanna Briggs. What don’t you know? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):343, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bennett:2006:HSC

[BBC06] Andrew Bennett, Joanna Briggs, and Martyn Clark. High school computing clubs: a pilot study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):38–42, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Blume:2009:CSC

[BBCD09] Lil Blume, Ron Baecker, Christopher Collins, and Aran Donohue. A “communication skills for computer scientists” course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):65–69, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Borunda:2006:GGV

[BBE06] Patrick Borunda, Chris Brewer, and Cesim Erten. GSPIM: graphical visualization tool for MIPS assembly programming and simulation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):244–248, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bruckman:2009:GCI

[BBE^{+}09] Amy Bruckman, Maureen Biggers, Barbara Ericson, Tom McKlin, Jill Dimond, Betsy DiSalvo, Mike Hewner, Lijun Ni, and Sarita Yardi. “Georgia computes!”: improving the computing education pipeline. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):86– 90, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Bailey:2009:RSP

[BBF^{+}09] Mark Bailey, Kim Bruce, Kathleen Fisher, Robert Harper, and Stuart Reges. Report of the 2008 SIGPLAN programming languages curriculum workshop: preliminary report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):132–133, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Bergin:2005:OET

[BBK05] Joe Bergin, Kim Bruce, and Michael Kölling. Objects-early tools: a demonstration. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):390– 391, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barnes:2003:THC

[BBMR03] Julie Barnes, Rob Bryant, Daniel D. McCracken, and Susan Reiser. Teaching human-computer interaction: reports from the trenches. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):125–126, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Benjamin:2003:UCS

[BBMW03] D. Paul Benjamin, Charles Border, Robert Montante, and Paul J. Wagner. Undergraduate cyber security course projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):351–352, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brady:2004:MCL

[BBN^{+}04] Alyce Brady, Kim Bruce, Robert Noonan, Allen Tucker, and Henry Walker. The 2003 model curriculum for a liberal arts degree in computer science: preliminary report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):282–283, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Beck:2007:UIB

[BBN07] Jon Beck, Brent Buckner, and Olga Nikolova. Using interdisciplinary bioinformatics undergraduate research to recruit and retain computer science students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):358–361, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bullers:2006:VMI

[BBS06] William I. Bullers, Jr., Stephen Burd, and Alessandro F. Seazzu. Virtual machines — an idea whose time has returned: application to network, security, and database courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):102–106, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Burge:2007:FOC

[BBT07] Mark Burge, Diana Burley, and Harriet Taylor. Funding opportunities in computer science education at the national science foundation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):49, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Berque:2004:UPB

[BBW04] Dave Berque, Terri Bonebright, and Michael Whitesell. Using pen-based computers across the computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):61–65, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Biggers:2008:SPC

[BBY08] Maureen Biggers, Anne Brauer, and Tuba Yilmaz. Student perceptions of computer science: a retention study comparing graduating seniors with CS leavers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (1):402–406, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Bennedsen:2004:PCM

[BC04a] Jens Bennedsen and Michael E. Caspersen. Programming in context: a model-first approach to CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):477–481, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Boyer:2004:IIT

[BC04b] Truman Parks Boyer and Mohsen Chitsaz. ICETM and ICE/TTM: tools to assist in compiler design and implementation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(4):55–57, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2004.bib.

Boyle:2004:CCE

[BC04c] Roger Boyle and Martyn Clark. CS++: content is not enough. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):422–426, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bruhn:2004:CCC

[BC04d] Russel E. Bruhn and Judy Camp. Capstone course creates useful business products and corporate-ready students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(2):87–92, June 2004. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2004.bib.

Bennedsen:2005:RPP

[BC05] Jens Bennedsen and Michael E. Caspersen. Revealing the programming process. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):186–190, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bennedsen:2006:AAI

[BC06a] Jens Bennedsen and Michael E. Caspersen. Abstraction ability as an indicator of success for learning object-oriented programming? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):39–43, June 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2006.bib.

Bird:2006:BSE

[BC06b] Steven Bird and James R. Curran. Building a search engine to drive problem-based learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (3):153–157, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bagley:2007:CIN

[BC07a] Carole A. Bagley and C. Candace Chou. Collaboration and the importance for novices in learning Java computer programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):211–215, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Bailey:2007:HET

[BC07b] Mike Bailey and Steve Cunningham. A hands-on environment for teaching GPU programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (1):254–258, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bennedsen:2007:FRI

[BC07c] Jens Bennedsen and Michael E. Caspersen. Failure rates in introductory programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):32– 36, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2007.bib.

Blum:2007:COP

[BC07d] Lenore Blum and Thomas J. Cortina. CS4HS: an outreach program for high school CS teachers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (1):19–23, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Burea:2007:CAI

[BC07e] Vladimír Burea and Pavel Cech. Complexity of ambient intelligence in managerial work. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3): 325, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Beck:2008:ESC

[BC08] Leland L. Beck and Alexander W. Chizhik. An experimental study of cooperative learning in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (1):205–209, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Boland:2009:IPD

[BC09] Michael G. Boland and Curtis Clifton. Introducing Py-Lighter: dynamic code highlighter. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (1):489–493, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Borstler:2008:EOE

[BCB^{+}08] Jürgen Börstler, Henrik B. Christensen, Jens Bennedsen, Marie Nordström, Lena Kallin Westin, Jan Erik Moström, and Michael E. Caspersen. Evaluating OO example programs for CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):47–52, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Buendia:2009:IAD

[BCB09] Félix Buendía, Juan Carlos Cano, and José Vicente Benlloch. An instructional approach to drive computer science courses through virtual learning environments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):6–10, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Botev:2005:SWC

[BCC^{+}05] Chavdar Botev, Hubert Chao, Theodore Chao, Yim Cheng, Raymond Doyle, Sergey Grankin, Jon Guarino, Saikat Guha, Pei-Chen Lee, Dan Perry, Christopher Re, Ilya Rifkin, Tingyan Yuan, Dora Abdullah, Kathy Carpenter, David Gries, Dexter Kozen, Andrew Myers, David Schwartz, and Jayavel Shanmugasundaram. Supporting workflow in a course management system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):262–266, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Byrd:2007:SRS

[BCC^{+}07] Jonathan Byrd, Monty Carter, David Cook, Philip Kirkham, and Martha J. Kosa. Solving recurring student problems with recurrences. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):343, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Bergin:2004:TSD

[BCD^{+}04] Joe Bergin, James Caristi, Yael Dubinsky, Orit Hazzan, and Laurie Williams. Teaching software development methods: the case of extreme programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (1):448–449, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bailey:2008:DAD

[BCD08] Mark W. Bailey, Clark L. Coleman, and Jack W. Davidson. Defense against the dark arts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1): 315–319, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Baldwin:2006:WPF

[BCE06] Jacqueline Baldwin, Eileen Crupi, and Tabitha Estrellado. WeBWorK for programming fundamentals. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):361, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brown:2000:EWI

[BCF^{+}00] David C. Brown, Isabel F. Cruz, David Finkel, Robert E. Kinicki, and Craig E. Wills. Experiences with the Webware, interfaces and networking experimental laboratory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):387–391, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bedy:2000:VSM

[BCHS00] Michael Bedy, Steve Carr, Xianlong Huang, and Ching-Kuang Shene. A visualization system for multithreaded programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):1–5, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Boyle:2004:PAS

[BCK04] Roger D. Boyle, Martyn Clark, and Amruth Kumar, editors. Proceedings of the 9th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education, ITiCSE 2004, Leeds, UK, June 28–30, 2004, SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 2004. CO-DEN SIGSD3. ISBN 1-58113-836-9. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brady:2000:IMB

[BCL00] Alyce Brady, Michael J. Clancy, and Kathleen Larson. Introduction to the marine biology case study (seminar session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):427, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Blum:2008:ECO

[BCLW08] Lenore Blum, Thomas J. Cortina, Ed Lazowska, and Joe Wise. The expansion of CS4HS: an outreach program for high school teachers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):377–378, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Beck:2005:CLT

[BCM05] Leland L. Beck, Alexander W. Chizhik, and Amy C. McElroy. Cooperative learning techniques in CS1: design and experimental evaluation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):470–474, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Balcita:2002:SFI

[BCS02] Angela M. Balcita, Doris L. Carver, and Mary Lou Soffa. Shortchanging the future of information technology: the untapped resource. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):32–35, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/

2002.bib. Women and Computing.

Bergin:2007:DOO

[BCS^{+}07] Joe Bergin, Mike Clancy, Don Slater, Michael Goldweber, and David B. Levine. Day one of the objects-first first course: what to do. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):264–265, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bredlau:2001:ALT

[BD01] Carl Bredlau and Dorothy Deremer. Assembly language through the Java Virtual Machine. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33 (1):194–198, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bentley:2003:SSP

[BDA03] Hilary Bentley, Jenny Davies, and Jo Allan. The stepping stones project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):229, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bueno:2007:LCA

[BdCGB07] F. Javier Bueno, J. Raul Fernández del Castillo, Soledad Garcia, and Reca Borrego. E-learning content adaptation for deaf students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):271–275, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Burhans:2004:BUC

[BDDL04] Debra T. Burhans, Matt DeJongh, Travis E. Doom, and Mark LeBlanc. Bioinformatics in the undergraduate curriculum: opportunities for computer science educators. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):229–230, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Boyer:2008:DED

[BDF^{+}08] Kristy Elizabeth Boyer, August A. Dwight, R. Taylor Fondren, Mladen A. Vouk, and James C. Lester. A development environment for distributed synchronous collaborative programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):158–162, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Bruce:2001:LSG

[BDM01a] Kim B. Bruce, Andrea Danyluk, and Thomas Murtagh. A library to support a graphics-based object-first approach to CS 1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):6–10, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bruce:2001:EDP

[BDM01b] Kim B. Bruce, Andrea P. Danyluk, and Thomas P. Murtagh. Event-driven programming is simple enough for CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):1–4, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bruce:2005:WSR

[BDM05] Kim B. Bruce, Andrea Danyluk, and Thomas Murtagh. Why structural recursion should be taught before arrays in CS 1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):246–250, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Boyer:2007:CSC

[BDM^{+}07] Kristy Elizabeth Boyer, Rachael S. Dwight, Carolyn S. Miller, C. Dianne Raubenheimer, Matthias F. Stallmann, and Mladen A. Vouk. A case for smaller class size with integrated lab for introductory computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):341–345, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barchino:2008:IAL

[BdMG08] Roberto Barchino, Luis de Marcos, and Jose M. Gutierrez. An interoperable assessment language proposal. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):357, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Boetticher:2005:UPA

[BDMY05] Gary D. Boetticher, Wei Ding, Charles Moen, and Kwok-Bun Yue. Using a pre-assessment exam to construct an effective concept-based genetic program for predicting course success. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):500–504, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brown:2009:WSS

[BDRH09] Richard Brown, Janet Davis, Samuel A. Rebelsky, and Brian Harvey. Whither scheme?: 21st century approaches to scheme in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):551–552, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Beyer:2005:CCS

[BDW^{+}05] Sylvia Beyer, Michelle DeKeuster, Kathleen Walter, Michelle Colar, and Christina Holcomb. Changes in CS students’ attitudes towards CS over time: an examination of gender differences. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):392–396, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Beaubouef:2002:WCS

[Bea02] Theresa Beaubouef. Why computer science students need math. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):57–59, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2002.bib.

Beaubouef:2003:WCS

[Bea03] Theresa Beaubouef. Why computer science students need language. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):51–54, December 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

Becker:2001:TCK

[Bec01] Byron Weber Becker. Teaching CS1 with Karel the robot in Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):50–54, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Becker:2003:AMN

[Bec03a] Katrin Becker. Assignments that meet the needs of exceptional students without disadvantaging the average. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):271, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Becker:2003:GPA

[Bec03b] Katrin Becker. Grading programming assignments using rubrics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):253, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Becker:2003:MIA

[Bec03c] Katrin Becker. A multiple intelligences approach to teaching number systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):250, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Becker:2006:HMC

[Bec06] K. Becker. How much choice is too much? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):78–82, December 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2006.bib.

Beck:2007:FWC

[Bec07] Jon Beck. Forming a women’s computer science support group. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):400–404, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Beck:2008:FDM

[Bec08] Jon Beck. Fair division as a means of apportioning software engineering class projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):68– 71, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Barros:2003:ULE

[BED^{+}03] João Paulo Barros, Luís Estevens, Rui Dias, Rui Pais, and Elisabete Soeiro. Using lab exams to ensure programming practice in an introductory programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):16–20, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Buechley:2007:TCE

[BEE07] Leah Buechley, Mike Eisenberg, and Nwanua Elumeze. Towards a curriculum for electronic textiles in the high school classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):28–32, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Bell:2002:VBN

[Bel02] Douglas Bell. Visual basic. net as a first language: an evaluation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):107–108, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2002.bib.

Bem:2002:EBP

[Bem02] Ewa Z. Bem. Experiment-based project in undergraduate computer architecture. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):171– 175, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Boustedt:2007:TCC

[BEM^{+}07] Jonas Boustedt, Anna Eckerdal, Robert McCartney, Jan Erik Moström, Mark Ratcliffe, Kate Sanders, and Carol Zander. Threshold concepts in computer science: do they exist and are they useful? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):504–508, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bergin:2001:PLI

[Ber01] Joseph Bergin. A pattern language for initial course design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):282–286, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bergin:2002:TWW

[Ber02] Joseph Bergin. Teaching on the Wiki Web. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):195, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bernick:2003:FMP

[Ber03a] Jonathan P. Bernick. The Flo-and-Mac problem: a tool for encouraging undergraduate research. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):103–106, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2003.bib.

Bernick:2003:TOO

[Ber03b] Jonathan P. Bernick. A translation of the one-to-one relationship for introductory relational database courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):66–67, December 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

Bergin:2005:AJ

[Ber05] Joseph Bergin. Academic jeopardy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37 (3):389, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bergin:2006:KUD

[Ber06] Joe Bergin. Karel universe drag & drop editor. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):307, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bergmann:2009:DKR

[Ber09a] Seth D. Bergmann. Degenerate keys for RSA encryption. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):95–98, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Berque:2009:TSB

[Ber09b] Dave A. Berque. A tutorial on stroke-based interfaces: unistroke recognition algorithms appropriate for compelling projects in introductory courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (1):485–486, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Braught:2008:EPP

[BEW08] Grant Braught, L. Martin Eby, and Tim Wahls. The effects of pair-programming on individual programming skill. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):200–204, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Bailey:2005:JTT

[BF05] Tammy Bailey and Jeffrey Forbes. Just-in-time teaching for CS0. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):366–370, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Buerck:2006:KDD

[BF06] John Buerck and David Feig. Knowledge discovery and dissemination: a curriculum model for informatics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):48–51, December 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2006.bib.

Bauer:2009:SVO

[BFD^{+}09] Christine Bauer, Kathrin Figl, Michael Derntl, Peter Paul Beran, and Sonja Kabicher. The student view on online peer reviews. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):26–30, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Bernstein:2005:APT

[BFMC05] Marc Bernstein, Kelly M. FitzGerald, James P. Macdonell, and Arturo I. Concepcion. AlgorithmA project: the ten-week mock software company. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):142– 146, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bareiss:2008:SCL

[BG08a] Ray Bareiss and Martin Griss. A story-centered, learn-by-doing approach to software engineering education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):221–225, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Bloomfield:2008:TBP

[BG08b] Aaron Bloomfield and James F. Groves. A tablet-based paper exam grading system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):83–87, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Barker:2003:EIC

[BGD03a] Lecia J. Barker and Kathy Garvin-Doxas. The effect of institutional characteristics on participation of women in computer science Bachelors degree programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):242, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barker:2003:WPC

[BGD03b] Lecia J. Barker and Kathy Garvin-Doxas. Why project courses sometimes widen the experience gap among students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):258, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barker:2002:DCC

[BGDJ02] Lecia Jane Barker, Kathy Garvin-Doxas, and Michele Jackson. Defensive climate in the computer science classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):43–47, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Barker:2005:WCC

[BGDR05] Lecia J. Barker, Kathy Garvin-Doxas, and Eric Roberts. What can computer science learn from a fine arts approach to teaching? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):421–425, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Burgos:2000:ASD

[BGG^{+}00] J. M. Burgos, J. Galve, J. García, J. J. Moreno, S. Muñoz, and D. Villén. Abstract solution design by specification refinement. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):186, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Boroni:2001:ESA

[BGGR01] Christopher M. Boroni, Frances W. Goosey, Michael T. Grinder, and Rockford J. Ross. Engaging students with active learning resources: hypertextbooks for the Web. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):65–69, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bridgeman:2000:PIT

[BGKT00a] Stina Bridgeman, Michael T. Goodrich, Stephen G. Kobourov, and Roberto Tamassia. PILOT: an interactive tool for learning and grading. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):139–143, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bridgeman:2000:SSG

[BGKT00b] Stina Bridgeman, Michael T. Goodrich, Stephen G. Kobourov, and Roberto Tamassia. SAIL: a system for generating, archiving, and retrieving specialized assignments using L^{A}TEX. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):300–304, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Burd:2007:CGC

[BGL^{+}07] Bary Burd, John Goulden, Brian Ladd, Michael Rogers, and Kris Stewart. Computer games in the classroom, or, how to get perfect attendance, even at 8 AM. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):496, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lillian:2009:EEC

[BGN09] Cassel Lillian (Boots), Hislop Gregory, and Bharath Nadella. Ensemble: enriching communities and collections to support education in computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):355, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Blumenstein:2004:EAG

[BGNM04] Michael Blumenstein, Steven Green, Ann Nguyen, and Vallipuram Muthukkumarasamy. An experimental analysis of GAME: a generic automated marking environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):67–71, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brusilovsky:2006:WSV

[BGSL06] Peter Brusilovsky, Jonathan Grady, Michael Spring, and Chul-Hwan Lee. What should be visualized?: faculty perception of priority topics for program visualization. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):44–48, June 2006. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2006.bib.

Begel:2004:KLC

[BGW04] Andrew Begel, Daniel D. Garcia, and Steven A. Wolfman. Kinesthetic learning in the classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):183–184, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Blandford:2003:FEE

[BH03] Dick K. Blandford and Deborah J. Hwang. Five easy but effective assessment methods. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):41– 44, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bishop:2004:DPG

[BH04] Judith Bishop and Nigel Horspool. Developing principles of GUI programming using views. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1): 373–377, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brown:2007:RIS

[BH07] Christopher W. Brown and Eric A. Hardisty. RegeXeX: an interactive system providing regular expression exercises. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):445–449, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bhatti:2000:VTT

[Bha00] M. Afzal Bhatti. Visual tool for teaching synchronization problems in operating systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(2): 44–45, June 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barbosa:2007:MUC

[BHBG07] Jorge Barbosa, Rodrigo Hahn, Débora N. F. Barbosa, and Cláudio F. R. Geyer. Mobile and ubiquitous computing in an innovative undergraduate course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (1):379–383, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Borriello:2008:MPS

[BHH^{+}08] Gaetano Borriello, Carl Hartung, Bruce Hemingway, Karl Koscher, and Brian Mayton. Multi-player soccer and wireless embedded systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):82–86, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Bucci:2002:AOO

[BHLW02] Paolo Bucci, Wayne Heym, Timothy J. Long, and Bruce W. Weide. Algorithms and object-oriented programming: bridging the gap. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):302–306, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Bilal:2006:EPI

[BHM06a] Sara Saeed Bilal, Fatma Hassan, and C. Dianne Martin. Ethics and professionalism for an IT professional in the UAE. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):6–7, June 2006. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2006.bib.

Byrnes:2006:TCE

[BHM^{+}06b] Jessica Byrnes, Thomas Huffner, Todd McNeal, Lauren Pisciotta, and Zackary Zweber. Technological confidence at an early age. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):365, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Borstler:2009:EOO

[BHN^{+}09] Jürgen Börstler, Mark S. Hall, Marie Nordström, James H. Paterson, Kate Sanders, Carsten Schulte, and Lynda Thomas. An evaluation of object oriented example programs in introductory programming textbooks. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4):126–143, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bar:2007:ISI

[BHR07] Henning Bär, Gina Häussge, and Guido Rößling. An integrated system for interaction support in lectures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):281–285, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Barbosa:2008:LMG

[BHRB08] Jorge Barbosa, Rodrigo Hahn, Solon Rabello, and Débora Barbosa. Local: a model geared towards ubiquitous learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):432–436, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Barzilay:2009:ESE

[BHY09] Ohad Barzilay, Orit Hazzan, and Amiram Yehudai. Evaluation of a software engineering course by reflection. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):273–277, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Beidler:2005:MFD

[BHYA05] Jack Beidler, Hilary Holz, Ken Yasuhara, and Evans J. Adams. The many facets of diversity. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37 (1):558–559, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bibby:2006:GFN

[Bib06] Pete Bibby. Getting feedback: no pressure! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):346, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bruhn:2007:MBM

[BJ07] Russel Bruhn and Steven F. Jennings. A multidisciplinary bioinformatics minor. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):348–352, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Berque:2001:TTC

[BJJ01] Dave Berque, David K. Johnson, and Larry Jovanovic. Teaching theory of computation using pen-based computers and an electronic whiteboard. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):169–172, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Boshart:2003:GGX

[BK03] Mark A. Boshart and Martha J. Kosa. Growing a GUI from an XML tree. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):223, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bergin:2002:OO

[BKHL02] Joseph Bergin, Richard Kick, Judith Hromcik, and Kathleen Larson. The object is objects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1): 251, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Bergin:2001:NPR

[BKM^{+}01] Joseph Bergin, Charles Kelemen, Myles F. McNally, Thomas L. Naps, Michael Goldweber, Chris Power, and Stephen J. Hartley. Non-programming resources for an introduction to CS: a collection of resources for the first courses in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(2):89– 100, June 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2001.bib.

Bergin:2005:ADC

[BKR^{+}05] Joseph Bergin, Clifton Kussmaul, Thomas Reichlmayr, James Caristi, and Gary Pollice. Agile development in computer science education: practices and prognosis. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):130–131, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Buckley:2004:BUS

[BKS^{+}04] Michael Buckley, Helene Kershner, Kris Schindler, Carl Alphonce, and Jennifer Braswell. Benefits of using sociallyrelevant projects in computer science and engineering education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):482–486, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bhagyavati:2005:UAD

[BKW05] Bhagyavati, Stan Kurkovsky, and Christopher C. Whitehead. Using asynchronous discussions to enhance student participation in CS courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):111–115, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bland:2000:APL

[BL00] Constance G. Bland and Pamela B. Lawhead. Agents, profiles, learning styles and tutors (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):185, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barnes:2001:TTI

[BL01a] Julie Barnes and Laura Leventhal. Turing the tables: introducing software engineering concepts in a user interface design course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):214–218, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brown:2001:DBO

[BL01b] Judy Brown and Jiayun Lu. Designing better online teaching material. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):352–356, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Byrne:2001:ESA

[BL01c] Pat Byrne and Gerry Lyons. The effect of student attributes on success in programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3): 49–52, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bass:2005:IID

[BL05a] Ilana Bass and Dvir Lanzberg. Iterative implementation of DFS. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):393, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bass:2005:TAS

[BL05b] Ilana Bass and Dvir Lanzberg. Teaching to analyze solutions using FGA. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):360, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bianco:2006:PGA

[BL06a] Giovanni M. Bianco and Ignazio Locatelli. PlayToLearn: a game adventure in the realm of Si Piuh. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):331, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brusilovsky:2006:WIC

[BL06b] Peter Brusilovsky and Tomasz D. Loboda. WADEIn II: a case for adaptive explanatory visualization. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):48–52, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Black:2009:BOS

[Bla09a] Michael D. Black. Build an operating system from scratch: a project for an introductory operating systems course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):448–452, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Blaheta:2009:CBS

[Bla09b] Don Blaheta. CS262: a breadth-second survey of informatic CS. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):49–53, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Blaheta:2009:DCE

[Bla09c] Don Blaheta. Democracy in the classroom: an exercise for the first days of CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):36–39, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Blaheta:2009:VPA

[Bla09d] Don Blaheta. A visual proof of amortised-linear resizable arrays. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):338, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Blanco:2009:ICP

[BLA^{+}09e] Javier Blanco, Leticia Losano, Nazareno Aguirre, María Marta Novaira, Sonia Permigiani, and Gastón Scilingo. An introductory course on programming based on formal specification and program calculation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):31–37, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Beaubouef:2001:USE

[BLH01] Theresa Beaubouef, Richard Lucas, and James Howatt. The UNLOCK system: enhancing problem solving skills in CS-1 students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(2):43–46, June 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2001.bib.

Bruce-Lockhart:2006:IEE

[BLN06] Michael Bruce-Lockhart and Theodore S. Norvell. Interactive embedded examples: a demonstration. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):357, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bergin:2006:FPC

[BLOM06] Joe Bergin, Raymond Lister, Barbara Boucher Owens, and Myles McNally. The first programming course: ideas to end the enrollment decline. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):301–302, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bucci:2001:DWR

[BLW01] Paolo Bucci, Timothy J. Long, and Bruce W. Weide. Do we really teach abstraction? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):26– 30, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Beaubouef:2005:WHA

[BM05] Theresa Beaubouef and John Mason. Why the high attrition rate for computer science students: some thoughts and observations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):103–106, June 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2005.bib.

Beaubouef:2007:IHT

[BM07a] Theresa Beaubouef and Patrick McDowell. Ice hockey and tennis balls: playing at computer science research with robotics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):121–124, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2007.bib.

Brown:2007:VBS

[BM07b] Christian Brown and Chris McDonald. Visualizing Berkeley socket calls in students’ programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (3):101–105, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Boustedt:2009:ISL

[BMD^{+}09] Jonas Boustedt, Robert McCartney, Katherine Deibel, Jim Huggins, Beth Simon, Suzanne Westbrook, and The Mystery Presenter. It seemed like a good idea at the time. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):265–266, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Bell:2003:TDS

[BMG03] Douglas Bell and Mehdi Mir-Ghasemi. Teaching data structures using list boxes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):79–81, December 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2003.bib.

Blank:2003:PRE

[BMK03] Douglas Blank, Lisa Meeden, and Deepak Kumar. Python robotics: an environment for exploring robotics beyond LE-GOs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):317–321, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barker:2009:EFI

[BMK09] Lecia J. Barker, Charlie McDowell, and Kimberly Kalahar. Exploring factors that influence computer science introductory course students to persist in the major. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):153–157, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Blahnik:2006:ACC

[BMP06] James Blahnik, Bonnie McVey, and David Pankratz. Adding concentrations to the CS major: our dean calls us ‘innovative’. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):191–194, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Braught:2004:CEC

[BMR04] Grant Braught, Craig S. Miller, and David Reed. Core empirical concepts and skills for computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):245–249, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Boustedt:2007:ISL

[BMT^{+}07] Jonas Boustedt, Robert McCartney, Josh Tenenberg, Titus Winters, Stephen Edwards, Briana B. Morrison, David R. Musicant, Ian Utting, and Carol Zander. It seemed like a good idea at the time. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):346–347, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Boustedt:2008:ISL

[BMT^{+}08] Jonas Boustedt, Robert McCartney, Josh Tenenberg, Scott D. Anderson, Caroline M. Eastman, Daniel D. Garcia, Paul V. Gestwicki, and Margaret S. Menzin. It seemed like a good idea at the time. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1): 528–529, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Baldwin:2004:SRS

[BMW04] Doug Baldwin, Bill Marion, and Henry Walker. Status report on the SIGCSE committee on the implementation of a discrete mathematics course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):98– 99, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bosse:2000:RWP

[BN00] Michael J. Bossé and N. R. Nandakumar. Real-world problem-solving, pedagogy, and efficient programming algorithms in computer education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(4): 66–69, December 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brownfield:2009:LGS

[BN09] Andrew Brownfield and Cindy Norris. LC3uArch: a graphical simulator of the LC-3 microarchitecture. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):413–417, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Buckley:2008:SRC

[BNS08] Michael Buckley, John Nordlinger, and Devika Subramanian. Socially relevant computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):347– 351, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Barnes:2005:EBS

[BNSW05] G. Michael Barnes, John Noga, Peter D. Smith, and Jeff Wiegley. Experiments with balanced-sample binary trees. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):166–170, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bryant:2001:ICS

[BO01] Randal E. Bryant and David R. O’Hallaron. Introducing computer systems from a programmer’s perspective. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):90–94, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Boetje:2006:FAT

[Boe06] Jerry Boetje. Foundational actions: teaching software engineering when time is tight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):285– 288, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bogoyavlenskaya:2004:TNC

[Bog04] O. Yu. Bogoyavlenskaya. Teaching networking congestion control. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(4):35–41, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2004.bib.

Boisvert:2004:CCI

[Boi04] Charles Boisvert. eL-CID: communicating iterative program design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):253, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Boisvert:2006:WAC

[Boi06] Charles Boisvert. Web animation to communicate iterative development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):173–177, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Boisvert:2009:VTP

[Boi09] Charles R. Boisvert. A visualisation tool for the programming process. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):328–332, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Booch:2007:RWR

[Boo07] Grady Booch. Readn’, writ’n, ’rithmetic… and code’n. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):197, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Borg:2002:CDE

[Bor02] Anita Borg. Computing 2002: democracy, education, and the future. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):13–14, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/

2002.bib. Women and Computing.

Borge:2004:TOE

[Bor04] Richard Borge. Teaching OOP: experimental studies. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):252, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Border:2007:DDM

[Bor07] Charles Border. The development and deployment of a multiuser, remote access virtualization system for networking, security, and system administration classes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):576–580, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Botting:2005:TLE

[Bot05] Richard J. Botting. Teaching and learning ethics in computer science: walking the walk. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):342– 346, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bouvier:2003:PSL

[Bou03] Dennis J. Bouvier. Pilot study: living flowcharts in an introduction to programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35 (1):293–295, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Boute:2009:TPC

[Bou09] Raymond Boute. Teaching and practicing computer science at the university level. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):24–30, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bowyer:2000:VRU

[Bow00] Kevin Bowyer. Video resources for use in teaching ethics and computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):217–221, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bower:2006:VCP

[Bow06] Matt Bower. Virtual classroom pedagogy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):148–152, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bower:2007:GAS

[Bow07a] Matt Bower. Groupwork activities in synchronous online classroom spaces. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):91–95, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bower:2007:ISA

[Bow07b] Matt Bower. Independent, synchronous and asynchronous an analysis of approaches to online concept formation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):141–145, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Bower:2008:ITC

[Bow08a] Matt Bower. The “instructed-teacher”: a computer science online learning pedagogical pattern. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (3):189–193, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Bower:2008:TTT

[Bow08b] Matt Bower. A taxonomy of task types in computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):281–285, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Bowring:2008:NPP

[Bow08c] James F. Bowring. A new paradigm for programming competitions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):87–91, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Boyle:2000:DIT

[Boy00] Robert Boyle. Do it themselves. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (3):180, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Boyle:2008:NRS

[Boy08] Roger D. Boyle. Neither rocket science nor washing machine science, but computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3): 5–6, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Bozic:2006:AFS

[Boz06] Vladan Bozic. Application of finite state automata to the bakery algorithm in critical section modeling. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):362, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barros:2000:VAC

[BP00] João Paulo Barros and Rui Pais. A versatile assignment in CS 2 (poster session): a file compression utility based on the Huffman code. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):185, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bernat:2002:TAN

[BP02] Andrew Bernat and Jane Prey. Taking advantage of National Science Foundation funding opportunities. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):52, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Bem:2003:MSP

[BP03] Ewa Z. Bem and Luke Petelczyc. MiniMIPS: a simulation project for the computer architecture laboratory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):64–68, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bures:2005:KBS

[BP05] Vladimir Bures and Daniela Ponce. Knowledge-based support of newcomers integration into an organization. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):372, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barker:2007:AT

[BP07] Guy Barker and Jane Prey. Accessibility tools. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):354, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Brown:2009:CDG

[BP09a] Christopher Brown and Robert Pastel. Combining distinct graduate and undergraduate HCI courses: an experiential and interactive approach. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):392– 396, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Bryfczynski:2009:GGC

[BP09b] Samuel P. Bryfczynski and Roy P. Pargas. GraphPad: a graph creation tool for CS2/CS7. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (3):389, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Burnell:2003:ARL

[BPD03] Lisa J. Burnell, John W. Priest, and John R. Durrett. Assessment of a resource limited process for multidisciplinary projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):68–71, December 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

Blank:2001:CPS

[BPK^{+}01] Glenn David Blank, William M. Pottenger, G. Drew Kessler, Martin Herr, Harriet Jaffe, and Soma Roy. CIMEL (poster session): constructive, collaborative inquiry-based multimedia E-learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):179, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brinda:2009:BIC

[BPS09] Torsten Brinda, Hermann Puhlmann, and Carsten Schulte. Bridging ICT and CS: educational standards for computer science in lower secondary education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):288–292, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Bareiss:2004:CSS

[BPT^{+}04] Cathy Bareiss, Kris Powers, Scott Thede, Marsha Meredith, Christine Shannon, and Judy Williams. The Computer Science Small Department Initiative (CS SDI) report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):332–333, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Boyer:2009:III

[BPW^{+}09] Kristy Elizabeth Boyer, Robert Phillips, Michael D. Wallis, Mladen A. Vouk, and James C. Lester. The impact of instructor initiative on student learning: a tutoring study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):14–18, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Braught:2002:DTS

[BR02] Grant Braught and David Reed. Disequilibration for teaching the scientific method in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):106–110, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Bergin:2005:PFI

[BR05a] Susan Bergin and Ronan Reilly. Programming: factors that influence success. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):411–415, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bower:2005:IVC

[BR05b] Matt Bower and Debbie Richards. The impact of virtual classroom laboratories in CSE. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1): 292–296, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Burg:2009:LCS

[BR09] Jennifer Burg and Jason Romney. Linking computer science, art, and practice through digital sound. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):473–477, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Braught:2005:TES

[Bra05] Grant Braught. Teaching empirical skills and concepts in computer science using random walks. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):41–45, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Becker:2001:PP

[BRB^{+}01] Byron Weber Becker, Richard Rasala, Joseph Bergin, Christine Shannon, and Eugene Wallingford. Polymorphic panelists. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):410–411, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bristow:2006:UVM

[Bri06] Patrick Bristow. Unicast vs. multicast over wireless: a cross-disciplinary mindshare for educational application researchers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):242–244, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brinkman:2009:HWB

[Bri09] Bo Brinkman. The heart of a whistle-blower: a corporate decision-making game for computer ethics classes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):316–320, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Brooks:2008:CCL

[Bro08] Christopher H. Brooks. Community connections: lessons learned developing and maintaining a computer science service-learning program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):352– 356, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Brown:2009:HHL

[Bro09] Richard A. Brown. Hadoop at home: large-scale computing at a small college. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):106–110, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Bravo:2000:DSC

[BROB00] C. Bravo, M. A. Redondo, M. Ortega, and J. Bravo. DO-MOSIMCOL: a simulation collaborative environment for the learning of domotic design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (2):65–67, June 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
Contrib.bib.

Beyer:2003:GDC

[BRP^{+}03] Sylvia Beyer, Kristina Rynes, Julie Perrault, Kelly Hay, and Susan Haller. Gender differences in computer science students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):49–53, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Barnes:2007:GBC

[BRP^{+}07] Tiffany Barnes, Heather Richter, Eve Powell, Amanda Chaffin, and Alex Godwin. Game2Learn: building CS1 learning games for retention. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):121– 125, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Bruce:2004:CHT

[Bru04] Kim B. Bruce. Controversy on how to teach CS 1: a discussion on the SIGCSE-members mailing list. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(4):29–34, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2004.bib.

Bruce:2005:CHT

[Bru05a] Kim B. Bruce. Controversy on how to teach CS 1: a discussion on the SIGCSE-members mailing list. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):111–117, June 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2005.bib.

Bruce:2005:UAM

[Bru05b] Kim B. Bruce. Using abstractions to make concepts concrete. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):1, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brylow:2008:ELE

[Bry08] Dennis Brylow. An experimental laboratory environment for teaching embedded operating systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):192–196, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Brynielsson:2009:IAC

[Bry09] Joel Brynielsson. An information assurance curriculum for commanding officers using hands-on experiments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):236–240, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Buck:2000:DEC

[BS00] Duane Buck and David J. Stucki. Design early considered harmful: graduated exposure to complexity and structure based on levels of cognitive development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):75–79, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Buck:2001:JCS

[BS01] Duane Buck and David J. Stucki. JKarelRobot: a case study in supporting levels of cognitive development in the computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):16–20, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bock:2002:DGB

[BS02] Douglas B. Bock and John F. Schrage. Denormalization guidelines for base and transaction tables. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):129–133, December 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2002.bib.

Burhans:2004:RCS

[BS04] Debra T. Burhans and Gary R. Skuse. The role of computer science in undergraduate bioinformatics education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):417–421, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brusilovsky:2005:ESW

[BS05] Peter Brusilovsky and Sergey Sosnovsky. Engaging students to work with self-assessment questions: a study of two approaches. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):251–255, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bayliss:2006:GFC

[BS06a] Jessica D. Bayliss and Sean Strout. Games as a “flavor” of CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):500–504, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Byckling:2006:RVP

[BS06b] Pauli Byckling and Jorma Sajaniemi. Roles of variables and programming skills improvement. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (1):413–417, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Begel:2008:SNC

[BS08] Andrew Begel and Beth Simon. Struggles of new college graduates in their first software development job. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):226–230, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Buck:2009:HIO

[BS09] Duane Buck and David J. Stucki. The hidden injuries of overloading ’ADT’. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):256–259, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Boavida:2005:MGE

[BSA^{+}05] Miguel Boavida, João Paulo Santos, Pedro Assunção, Luis Soares, João Luz, and Ricardo Viegas. The MulE game engine extending online role-playing games. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37 (3):404, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Brusilovsky:2008:OIE

[BSL^{+}08] Peter Brusilovsky, Sergey Sosnovsky, Danielle H. Lee, Michael Yudelson, Vladimir Zadorozhny, and Xin Zhou. An open integrated Exploratorium for database courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):22–26, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Berque:2004:BWE

[BSV04] Dave A. Berque, Ian Serlin, and Atanas Vlahov. A brief water excursion: introducing computer organization students to a water driven 1-bit half-adder. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(2): 52–56, June 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2004.bib.

Bernat:2001:TAN

[BT01] Andrew Bernat and Harriet Taylor. Taking advantage of national science foundation funding opportunities. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):420, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bianco:2004:RSP

[BT04] Giovanni M. Bianco and Simonetta Tinazzi. In the realm of Si Piuh. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):247, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bianco:2005:WCL

[BT05] Giovanni M. Bianco and Simonetta Tinazzi. Why are children left out of the CS rethinking process? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):352, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bianco:2006:OSF

[BT06] Giovanni M. Bianco and Simonetta Tinazzi. One step further the ACM K-12 final report: a proposal for level 1: computer organization for K-8. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):207–211, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bernat:2000:SSR

[BTGD00] Andrew Bernat, Patricia J. Teller, Ann Gates, and Nellie Delgado. Structuring the student research experience. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):17–20, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Baldwin:2006:PST

[BTHR06] Doug Baldwin, Paul T. Tymann, Susan M. Haller, and Ingrid Russell, editors. Proceedings of the 39th SIGCSE Technical Symposium on Computer Science Education, SIGCSE 2006, Houston, Texas, USA, March 3–5, 2006, SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 2006. CO-DEN SIGSD3. ISBN 1-59593-259-3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bennett:2009:UDS

[BU09] Chris Bennett and Timothy Urness. Using daily student presentations to address attitudes and communication skills in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):76–80, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Barriocanal:2002:EIA

[BUCP02] Elena García Barriocanal, Miguel-Ángel Sicilia Urbán, Ignacio Aedo Cuevas, and Paloma Díaz Pérez. An experience in integrating automated unit testing practices in an introductory programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4): 125–128, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2002.bib.

Budd:2006:ALA

[Bud06] Timothy A. Budd. An active learning approach to teaching the data structures course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):143– 147, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Burger:2003:TTD

[Bur03] Kevin R. Burger. Teaching two-dimensional array concepts in Java with image processing examples. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):205–209, January 2003. CO-DEN SIGSD3. ISBN 1-58113-648-X. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Compiler/java.bib;   ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/2003.
bib.

Burge:2004:PCU

[Bur04] Mark Burge. Pervasive computing in the undergraduate curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):181–182, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bures:2006:STB

[Bur06] Vladimír Bures. Systems thinking as a basis for ambient intelligence. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):318, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Burrell:2008:LOO

[Bur08] Christopher J. Burrell. Learning object oriented programming: unique visualizations of individuals learning styles, activities and the programs produced. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (3):339, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Baker:2009:ERD

[BvdH09] Alex Baker and André van der Hoek. An experience report on the design and delivery of two new software design courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):519–523, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Bierre:2006:MOB

[BVPE06] Kevin Bierre, Phil Ventura, Andrew Phelps, and Christopher Egert. Motivating OOP by blowing things up: an exercise in cooperation and competition in an introductory Java programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):354–358, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bergin:2005:TFP

[BW05] Joseph Bergin and Eugene Wallingford. Test-first pairprogramming tutorial. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):416, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Berglund:2006:SLC

[BW06] Anders Berglund and Mattias Wiggberg. Students learn CS in different ways: insights from an empirical study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):265–269, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bergin:2005:TPE

[BWC^{+}05] Joseph Bergin, Eugene Wallingford, Michael Caspersen, Michael Goldweber, and Michael Kolling. Teaching polymorphism early. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):342–343, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Biggers:2009:UCM

[BYS09] Maureen Biggers, Tuba Yilmaz, and Monica Sweat. Using collaborative, modified peer-led team learning to improve student success and retention in intro CS. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):9–13, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Benaya:2004:CSI

[BZ04a] Tamar Benaya and Ela Zur. Can students improve their undergraduate achievements and get accepted to graduate school? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):246, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Burch:2004:SCS

[BZ04b] Carl Burch and Lynn Ziegler. Science of computing suite (SOCS): resources for a breadth-first introduction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):437–441, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Benaya:2005:APJ

[BZ05] Tamar Benaya and Ela Zur. Advanced programming in Java workshop: teaching methodology. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37 (3):348, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Benaya:2007:UTA

[BZ07] Tamar Benaya and Ela Zur. Understanding threads in an advanced Java course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):323, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Cuny:2002:RRW

[CA02] Janice Cuny and William Aspray. Recruitment and retention of women graduate students in computer science and engineering: results of a workshop organized by the computing research association. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):168– 174, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/

2002.bib. Women and Computing.

Cassel:2009:SBB

[CAK09] Lillian N. Cassel, Richard H. Austing, and Elliott Koffman. SIGCSE: from the beginnings to a bright future. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):64, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Caldeira:2008:TSC

[Cal08] Carlos Pampulim Caldeira. Teaching SQL: a case study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):340, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Campbell:2001:TYC

[Cam01] Robert D. Campbell. Two-year college education committee update. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(4):24–25, December 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2001.bib.

Camp:2002:ISP

[Cam02a] Tracy Camp. The incredible shrinking pipeline. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):129–134, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2002.bib. Women and

Computing.

Campbell:2002:NCC

[Cam02b] Robert Campbell. New computing curriculum for two-year colleges. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):21–22, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2002.bib.

Campbell:2003:ATY

[Cam03a] Robert D. Campbell. ACM two-year college education committee report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):20–21, December 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2003.bib.

Campbell:2003:C

[Cam03b] Robert D. Campbell. Cybersecurity. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35 (2):24–26, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2003.bib.

Campbell:2004:ISD

[Cam04a] Robert D. Campbell. IT security and data assurance: a new resource for two-year colleges. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (4):20, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2004.bib.

Campbell:2004:UTY

[Cam04b] Robert D. Campbell. Update on two-year college activity. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(2):21–23, June 2004. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2004.bib.

Campbell:2005:ADT

[Cam05a] Robert D. Campbell. Associate-degree transfer curriculum in software engineering. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):17– 18, June 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2005.bib.

Campbell:2005:FAT

[Cam05b] Robert D. Campbell. Forthcoming activities in the twoyear college setting. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4):21–22, December 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2005.bib.

Campbell:2006:ADT

[Cam06a] Robert D. Campbell. Associate-degree transfer curriculum in computer engineering. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):15, December 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2006.bib.

Campbell:2006:ITP

[Cam06b] Robert D. Campbell. Information technology: … a program of study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):15–16, June 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2006.bib.

Campbell:2007:RH

[Cam07] Robert D. Campbell. Reflections on a history. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):15, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2007.bib.

Carter:2003:HSW

[CAMF^{+}03] Janet Carter, Kirsti Ala-Mutka, Ursula Fuller, Martin Dick, John English, William Fone, and Judy Sheard. How shall we assess this? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):107– 123, December 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2003.bib;   ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Techreports/UKent.
bib;   http://www.cs.kent.ac.uk/pubs/2003/1756.

Canosa:2006:IUS

[Can06] Roxanne L. Canosa. Image understanding as a second course in AI: preparing students for research. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):543–547, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Capretz:2002:IMS

[Cap02] Luiz Fernando Capretz. Implications of MBTI in software engineering education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):134–137, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2002.bib.

Carter:2000:WSS

[Car00] Janet Carter. What the students said about plagiarism. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):186, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Caristi:2002:TEP

[Car02a] James Caristi. Tutorial on extreme programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):100, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Cartelli:2002:CSE

[Car02b] Antonio Cartelli. Computer science education in Italy: a survey. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):36–39, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2002.bib.

Carl:2003:TDV

[Car03] Stephen P. Carl. The treatment of deep vs. shallow copy in introductory C++ textbooks. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):99– 102, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2003.bib.

Carlson:2004:TCS

[Car04] David Carlson. Teaching computer security. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(2):64–67, June 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2004.bib.

Cardoso:2005:NCC

[Car05a] João M. P. Cardoso. New challenges in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):203–207, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carrasquel:2005:SEF

[Car05b] Jacobo Carrasquel. Students evaluating faculty: a subjective process. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):346, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carter:2006:WSA

[Car06] Lori Carter. Why students with an apparent aptitude for computer science don’t choose to major in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):27–31, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carmichael:2008:GCS

[Car08] Gail Carmichael. Girls, computer science, and games. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):107–110, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cassel:2001:RSP

[Cas01] Lillian N. Cassel. Reading summaries (poster session): relating class to student’s problems with the current reading assignment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):188, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cassel:2002:VAL

[Cas02] Lillian N. Cassel. Very active learning of network routing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):195, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cassel:2003:UCR

[Cas03] Lillian N. Cassel. Using CITIDEL resources to support documenting objectives and outcomes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35 (3):255, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cassell:2008:FI

[Cas08] Lillian (Boots) Cassell. The future of ITiCSE. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):331–332, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Cassapu:2009:IIW

[Cas09] Anita Cassapu. Introducing the interactive whiteboard in computer science teaching: a high school experience. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):364, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Chamillard:2000:EPA

[CB00] A. T. Chamillard and Kim A. Braun. Evaluating programming ability in an introductory computer science course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):212–216, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chamillard:2002:SEC

[CB02] A. T. Chamillard and Kim A. Braun. The software engineering capstone: structure and tradeoffs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):227–231, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Cheng:2004:EEI

[CB04] Chew Swee Cheng and Chris Beaumont. Evaluating the effectiveness of ICT to support globally distributed PBL teams.

SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):47–51, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Courte:2009:DSD

[CBC09] Jill Courte and Cathy Bishop-Clark. Do students differentiate between computing disciplines? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (1):29–33, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Cummins:2009:TBF

[CBH09] Stephen Cummins, Liz Burd, and Andrew Hatch. Tag based feedback for programming courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (4):62–65, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Choi:2008:DOC

[CBJL08] Sook Kyoung Choi, Tim Bell, Soo Jin Jun, and Won Gyu Lee. Designing offline computer science activities for the Korean elementary school curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (3):338, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Cappos:2009:SPE

[CBKA09] Justin Cappos, Ivan Beschastnikh, Arvind Krishnamurthy, and Tom Anderson. Seattle: a platform for educational cloud computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):111–115, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Cutts:2007:EIP

[CBKM07] Quintin I. Cutts, Margaret I. Brown, Lynsey Kemp, and Calum Matheson. Enthusing and informing potential computer science students and their teachers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):196–200, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Chalk:2003:DEL

[CBP03] Peter Chalk, Claire Bradley, and Poppy Pickard. Designing and evaluating learning objects for introductory programming education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):240, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Choi:2001:UIM

[CC01] Sun-Hea Choi and Sandra Cairncross. Using interactive multimedia for teaching and learning object oriented software design (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):176, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Christensen:2002:FCD

[CC02] Henrik Bærbak Christensen and Michael E. Caspersen. Frameworks in CS1: a different way of introducing eventdriven programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):75–79, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Claypool:2005:TSE

[CC05] Kajal Claypool and Mark Claypool. Teaching software engineering through game design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3): 123–127, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carle:2007:WPP

[CCC07] Andy Carle, Michael Clancy, and John Canny. Working with pedagogical patterns in PACT: initial applications and observations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):238–242, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Caspersen:2008:WML

[CCD^{+}08a] Michael Caspersen, Lillian Cassell, Gordon Davies, Arnold Pears, Stephen Seidman, and Heikki Topi. What is Masters level education in informatics? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (3):341, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Cassel:2008:CUA

[CCD^{+}08b] Lillian N. Cassel, Michael Caspersen, Gordon Davies, Renee McCauley, Andrew McGettrick, Art Pyster, and Robert Sloan. Curriculum update from the ACM Education Board: CS2008 and a report on Masters degrees. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):530–531, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Cilliers:2005:EII

[CCG05] Charmain Cilliers, André Calitz, and Jéan Greyling. The effect of integrating an Iconic programming notation into CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):108–112, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Campbell:2003:LII

[CCH03a] Alistair E. R. Campbell, Geoffrey L. Catto, and Eric E. Hansen. Language-independent interactive data visualization. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):215–219, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cera:2003:DP

[CCH^{+}03b] Christopher D. Cera, Bruce Char, Nira Herrmann, Robert N. Lass, Aparna Nanjappa, Jeffrey L. Popyack, and Paul Zoski. The DUPLEX project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):259, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cera:2003:HTD

[CCH^{+}03c] Christopher D. Cera, Bruce Char, Nira Herrmann, Robert N. Lass, Aparna Nanjappa, Jeffrey L. Popyack, and Paul Zoski. High-tech dishonesty: cheating, plagiarism and detection. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):244, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carr:2002:CVT

[CCJ^{+}02] Steve Carr, Ping Chen, Timothy R. Jozwowski, Jean Mayo, and Ching-Kuang Shene. Channels, visualization, and topology editor. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):106–110, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cover:2001:IMC

[CCK01] Comfort Fay Cover, Robert D. Campbell, and Karl J. Klee. An international model for curriculum adaptation and implementation (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):179, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Calder:2007:TDS

[CCL^{+}07] Matt Calder, Robert F. Cohen, Jessica Lanzoni, Neal Landry, and Joelle Skaff. Teaching data structures to students who are blind. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):87–90, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Caristi:2007:AFI

[CCM07] James Caristi, Stephen Cooper, and Judy Mullins. Alice Frenzy: an interactive group activity. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):394, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Crenshaw:2008:CSR

[CCM08] Tanya L. Crenshaw, Erin Wolf Chambers, and Heather Metcalf. A case study of retention practices at the University of Illinois at Urbana-Champaign. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1): 412–416, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Cooper:2005:OBC

[CCMC05] Stephen Cooper, Lillian Cassel, Barbara Moskal, and Steve Cunningham. Outcomes-based computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):260–261, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chen:2005:RIS

[CCMS05] Tzu-Yi Chen, Stephen Cooper, Robert McCartney, and Leslie Schwartzman. The (relative) importance of software design criteria. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):34–38, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chesnevar:2003:UTC

[CCY03] Carlos Iván Chesñevar, María L. Cobo, and William Yurcik. Using theoretical computer simulators for formal languages and automata theory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):33– 37, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2003.bib.

Carniani:2001:NET

[CD01a] Enrico Carniani and Renzo Davoli. The NetWire emulator: a tool for teaching and understanding networks. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):153–156, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clear:2001:CIE

[CD01b] Tony Clear and Mats Daniels. A cyber-icebreaker for an effective virtual group? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):121–124, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Catuogno:2008:IRG

[CD08] Luigi Catuogno and Alfredo De Santis. An Internet role-game for the laboratory of network security course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):240–244, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Cassel:2007:COA

[CDF^{+}07] Lillian N. Cassel, Gordon Davies, William Fone, Anneke Hacquebard, John Impagliazzo, Richard J. LeBlanc, Joyce Currie Little, Andrew D. McGettrick, and Michela Pedrona. The computing ontology: application in education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):171–183, December 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Chu:2001:ITC

[CDGS01] Bei-Tseng (Bill) Chu, Venu Dasigi, John Gorgone, and David Spooner. Information technology curriculum development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):400–401, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cooper:2003:TOF

[CDP03] Stephen Cooper, Wanda Dann, and Randy Pausch. Teaching objects-first in introductory computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):191–195, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cassel:2009:MDC

[CDS09] Lillian N. Cassel, Gordon Davies, and Stephen Seidman. Masters degrees in computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):366, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Cassel:2000:PST

[CDWH00] Boots Cassel, Nell B. Dale, Henry MacKay Walker, and Susan M. Haller, editors. The proceedings of the thirty-first SIGCSE Technical Symposium on Computer Science Education: SIGCSE 2000, Austin, Texas, March 8–12, 2000, volume 32(1) of SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 2000. CODEN SIGSD3. ISBN 1-58113-213-1. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clear:2002:ACJ

[CF02] Tony Clear and Graeme Foot. Avatars in cyberspace — a Java 3D application to support formation of virtual groups. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):222, September 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Compiler/java.bib.

Cassel:2004:SWP

[CF04a] Lillian N. Cassel and Edward A. Fox. Sharing the wealth: publishing electronic resources. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3): 234, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Conrad:2004:USB

[CF04b] Marc Conrad and Tim French. Using the synergies between the object-oriented paradigm and mathematics in joint mathematics/computer science programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (3):254, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cassel:2006:SWP

[CF06] Lillian N. Cassel and Edward A. Fox. Sharing the wealth: publishing electronic resources. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3): 327, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cohen:2005:AIC

[CFGL05] Robert F. Cohen, Alexander V. Fairley, David Gerry, and Gustavo R. Lima. Accessibility in introductory computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):17–21, March 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Czajkowski:2001:SPS

[CFH^{+}01] Michael F. Czajkowski, Cheryl V. Foster, Thomas T. Hewett, Joseph A. Casacio, William C. Regli, and Heike A. Sperber. A student project in software evaluation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):13–16, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carr:2002:CLS

[CFJ^{+}02] Steve Carr, Changpeng Fang, Tim Jozwowski, Jean Mayo, and Ching-Kuang Shene. A communication library to support concurrent programming courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (1):360–364, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Countryman:2002:DHP

[CFKS02] Jeri Countryman, Alegra Feldman, Linda Kekelis, and Ellen Spertus. Developing a hardware and programming curriculum for middle school girls. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):44– 47, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/

2002.bib. Women and Computing.

Cassel:2004:UCD

[CFL^{+}04] Lillian N. Cassel, Edward A. Fox, JAN Lee, Manuel A. Pérez-Quiñones, Deborah Knox, John Impagliazzo, and Lee Giles. Using CITIDEL to develop and share class plans. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):270, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cunha:2005:PAS

[CFPL05] José C. Cunha, William M. Fleischman, Viera K. Proulx, and João Lourenço, editors. Proceedings of the 10th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education, ITiCSE 2005, Caparica, Portugal, June 27–29, 2005, SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 2005. CODEN SIGSD3. ISBN 1-59593-024-8. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cavalcante:2004:VIA

[CFR04] Ryan Cavalcante, Thomas Finley, and Susan H. Rodger. A visual and interactive automata theory course with JFLAP 4.0. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):140–144, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cooper:2008:ENF

[CFT08] Stephen Cooper, Timothy Fossum, and Harriet Taylor. Exploring NSF funding opportunities. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (1):197–198, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Claypool:2001:OSL

[CFW01] Mark Claypool, David Finkel, and Craig Wills. An open source laboratory for operating systems projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):145–148, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Califf:2002:TSK

[CG02] Mary Elaine Califf and Mary Goodwin. Testing skills and knowledge: introducing a laboratory exam in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):217–221, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Califf:2005:EIE

[CG05] Mary Elaine Califf and Mary Goodwin. Effective incorporation of ethics into courses that focus on programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):347–351, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Califf:2008:HHS

[CGB08] Mary Elaine Califf, Mary M. Goodwin, and Jake Brownell. Helping him see: guiding a visually impaired student through the computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (1):444–448, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Crescenzi:2005:NIE

[CGI05] Pilu Crescenzi, Giorgio Gambosi, and Gaia Innocenti. Net-PrIDE an integrated environment for developing and visualizing computer network protocols. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37 (3):306–310, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chesnevar:2004:DSP

[CGM04] Carlos I. Chesñevar, Maria P. González, and Ana G. Maguitman. Didactic strategies for promoting significant learning in formal languages and automata theory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):7–11, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Corbera:2008:DNM

[CGR^{+}08] Francisco Corbera, Eladio Gutiérrez, Julián Ramos, Sergio Romero, and María A. Trenas. Development of a new MOO-DLE module for a basic course on computer architecture. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):349, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Cunningham:2005:TAN

[CGT05] Steve Cunningham, Diana Gant, and Harriet Taylor. Taking advantage of national science foundation funding opportunities. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):193, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clear:2001:RIC

[CGY^{+}01] Tony Clear, Michael Goldweber, Frank H. Young, Paul M. Leidig, and Kirk Scott. Resources for instructors of capstone courses in computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(4):93–113, December 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2001.bib.

Cross:2006:JLI

[CH06] James H. Cross II and T. Dean Hendrix. jGRASP: a lightweight IDE with dynamic object viewers for CS1 and CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):356, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cohen:2007:CSL

[CH07] Avi Cohen and Bruria Haberman. Computer science: a language of technology. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):65–69, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2007.bib.

Craig:2009:GDG

[CH09] Michelle Craig and Diane Horton. Gr8 designs for Gr8 girls: a middle-school program and its evaluation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):221–225, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Chalk:2000:ALS

[Cha00a] Peter Chalk. Apprenticeship learning of software engineering using Webworlds. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):112–115, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chang:2000:DLP

[Cha00b] Chih-Kai Chang. Discovering learning patterns from Web logs by concept transformation analysis (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):186–187, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chalk:2001:SLV

[Cha01] Peter Chalk. Scaffolding learning in virtual environments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):85–88, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chalk:2002:EUV

[Cha02] Peter D. Chalk. Evaluating the use of a virtual learning environment for teaching aspects of HCI. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (3):125–129, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chandra:2003:BPP

[Cha03a] Surendar Chandra. Beacond: a peer-to-peer system to teach ubiquitous computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):257–261, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chaudhuri:2003:DAO

[Cha03b] Ranjan Chaudhuri. Do the arithmetic operations really execute in constant time? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):43– 44, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2003.bib.

Chang:2004:TCN

[Cha04a] Rocky K. C. Chang. Teaching computer networking with the help of personal computer networks. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (3):208–212, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chaudhuri:2004:TBL

[Cha04b] Ranjan Chaudhuri. Teaching bit-level algorithm analysis to the undergraduates in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(2):62–63, June 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2004.bib.

Challinger:2005:EUR

[Cha05] Judith Challinger. Efficient use of robots in the undergraduate curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):436–440, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chamillard:2006:IGC

[Cha06a] A. T. Chamillard. Introductory game creation: no programming required. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):515–519, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chamillard:2006:USP

[Cha06b] A. T. Chamillard. Using student performance predictions in a computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (3):260–264, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chandra:2007:LVC

[Cha07] Surendar Chandra. Lecture video capture for the masses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):276–280, September 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Chewar:2006:APP

[CHB06] Christa M. Chewar, Kevin L. Huggins, and Jean R. S. Blair. Avoiding the pratfalls of program assessment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):29–33, December 2006. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2006.bib.

Cannon:2002:TSP

[CHDH02] Robert Cannon, Thomas B. Hilburn, and Jorge Diaz-Herrera. Teaching a software project course using the team software process. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):369–370, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Cheatham:2000:WBL

[Che00] Thomas J. Cheatham. A Web-based lab manual for CS 1: an experiment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):105–108, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chinn:2005:PAA

[Chi05] Donald Chinn. Peer assessment in the algorithms course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):69–73, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cross:2007:DOV

[CHJB07] James H. Cross II, T. Dean Hendrix, Jhilmil Jain, and Larry A. Barowski. Dynamic object viewers for data structures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):4–8, March 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cassel:2001:DET

[CHK^{+}01] Lillian (Boots) Cassel, Mark A. Holliday, Deepak Kumar, John Impagliazzo, Kevin Bolding, Murray Pearson, Jim Davies, Gregory S. Wolffe, and William Yurcik. Distributed expertise for teaching computer organization & architecture. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(2):111–126, June 2001. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2001.bib.

Campbell:2003:CPA

[CHK03a] Robert D. Campbell, Elizabeth K. Hawthorne, and Karl J. Klee. Curricular projects of the ACM two-year college education committee. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):222, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Campbell:2003:RTY

[CHK03b] Robert D. Campbell, Elizabeth K. Hawthorne, and Karl J. Klee. The role of two-year colleges in educating the cybersecurity workforce. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):235, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Campbell:2005:CPA

[CHK05] Robert D. Campbell, Elizabeth K. Hawthorne, and Karl J. Klee. Curricular projects of the ACM two-year college education committee. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):361, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Campbell:2006:AAT

[CHK06] Robert D. Campbell, Elizabeth K. Hawthorne, and Karl J. Klee. Activities of the ACM two-year college education committee. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):321, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clear:2001:DDL

[CHM^{+}01] Tony Clear, Arto Haataja, Jeanine Meyer, Jarkko Suhonen, and Stuart A. Varden. Dimensions of distance learning for computer education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(2):101– 110, June 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2001.bib.

Cassel:2005:SCC

[CHM^{+}05] Lillian N. Cassel, Anneke Hacquebard, Andrew D. McGettrick, Gordon Davies, Richard J. LeBlanc, Charles Riedesel, Yaakov L. Varol, Gail T. Finley, Samuel Mann, and Robert H. Sloan. A synthesis of computing concepts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4):162–172, December 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2005.bib.

Carbone:2001:CPE

[CHMG01] Angela Carbone, John Hurst, Ian Mitchell, and Dick Gunstone. Characteristics of programming exercises that lead to poor learning tendencies: Part II. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33 (3):93–96, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Combs:2005:CSP

[CHP^{+}05] William Combs, Robert Hawkins, Thomas Pore, Arik Schechet, Tim Wahls, and Louis Ziantz. The course scheduling problem as a source of student projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):81–85, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Christensen:2003:STS

[Chr03] Henrik Bærbak Christensen. Systematic testing should not be a topic in the computer science curriculum! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):7–10, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Christensen:2004:FPD

[Chr04] Henrik Bærbak Christensen. Frameworks: putting design patterns into perspective. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):142– 145, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Christensen:2005:IPT

[Chr05a] Henrik Baerbak Christensen. Implications of perspective in teaching objects first and object design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):94–98, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Christensen:2005:TLJ

[Chr05b] Henrik Bærbak Christensen. TS-05: 150 lines of Java with high architectural complexity. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3): 396, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Christensen:2009:STA

[Chr09] Henrik Bærbak Christensen. A story-telling approach for a software engineering course design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (3):60–64, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Coppit:2005:LTP

[CHS05] David Coppit and Jennifer M. Haddox-Schatz. Large team projects in software engineering courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):137–141, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chung:2001:NAH

[Chu01] Sei-Jong Chung. Network architecture: Hamming codes and cyclic redundancy for transmission error correction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(4):48–50, December 2001. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2001.bib.

Chung:2003:NPC

[Chu03] Sei-Jong Chung. Network protocols: correcting transmission errors of up to two bits. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):60–62, December 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2003.bib.

Cross:2008:EAV

[CHUB08] James H. Cross II, T. Dean Hendrix, David A. Umphress, and Larry A. Barowski. Exploring accessibility and visibility relationships in Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):103–108, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Crescenzi:2003:TDE

[CI03] Pilu Crescenzi and Gaia Innocenti. A tool to develop electronic course books based on WWW technologies, resources and usability criteria. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):163–167, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cicirello:2009:REP

[Cic09] Vincent A. Cicirello. On the role and effectiveness of pop quizzes in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):286–290, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Cigas:2003:ICN

[Cig03] John Cigas. An introductory course in network administration. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):113–116, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cassel:2003:UEO

[CIK^{+}03] Lillian N. Cassel, John Impagliazzo, Deborah Knox, C. Lee Giles, Edward A. Fox, JAN Lee, and Manuel A. Pérez-Quiñones. Using an education oriented digital library to organize and present classes in computing and information. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):260, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cassel:2003:RNM

[CIL^{+}03] Lillian N. Cassel, John Impagliazzo, Ann Lally, Edward A. Fox, Manuel A. Pérez-Quiñones, and Jane Prey. Report on the NSF major educational funding initiative for a national science, technology, engineering, and mathematics education digital library (NSDL) with special emphasis on the computing education component. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1): 301–302, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cizmar:2003:SEC

[Ciz03] Dawn Cizmar. Software engineering concepts for multimedia development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):230, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carter:2001:GDP

[CJ01a] Janet Carter and Tony Jenkins. Gender differences in programming? (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):173, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chamillard:2001:ULP

[CJ01b] A. T. Chamillard and Jay K. Joiner. Using lab practica to evaluate programming ability. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1): 159–163, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carter:2002:GDP

[CJ02] Janet Carter and Tony Jenkins. Gender differences in programming? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):188–192, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Caspersen:2002:PAS

[CJGU02] Michael E. Caspersen, Dan Joyce, Don Goelman, and Ian Utting, editors. Proceedings of the 7th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education, ITiCSE 2002, Aarhus, Denmark, June 24–28, 2002, SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 2002. CODEN SIGSD3. ISBN 1-58113-499-1. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cassel:2002:SCR

[CK02] Lillian (Boots) Cassel and Deepak Kumar. A state of the course report: computer organization & architecture. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):175–177, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Coffey:2004:MMO

[CK04] John W. Coffey and Robert Koonce. MODeLeR: multimedia object design learning resource. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (1):472–476, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clear:2008:CCC

[CK08] Tony Clear and Diana Kassabova. A course in collaborative computing: collaborative learning and research with a global perspective. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):63–67, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Coleman:2005:GDP

[CKLW05] Ron Coleman, Mary Krembs, Alan Labouseur, and Jim Weir. Game design & programming concentration within the computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):545– 550, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clifton:2007:SFS

[CKM07] Curtis Clifton, Lisa C. Kaczmarczyk, and Michael Mrozek. Subverting the fundamentals sequence: using version control to enhance course management. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (1):86–90, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cole:2009:MPC

[CKMP09] Marilyn C. Cole, Evan Korth, Adam Meyers, and Sam Pluta. Musicomputation: a pilot course exploring a pre-college computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):381, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Collberg:2002:GTE

[CKMW02] Christian Collberg, Stephen G. Kobourov, Jessica Miller, and Suzanne Westbrook. Aλ goVista: a tool to enhance algorithm design and understanding. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):228, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Collberg:2004:AAS

[CKW04] Christian Collberg, Stephen G. Kobourov, and Suzanne Westbrook. AlgoVista: an algorithmic search tool in an educational setting. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):462–466, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Choi:2000:SCP

[CL00] Sung-Eun Choi and E. Christopher Lewis. A study of common pitfalls in simple multi-threaded programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):325–329, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clayton:2002:TYS

[CL02] Debbie Clayton and Teresa Lynch. Ten years of strategies to increase participation of women in computing programs: the Central Queensland University experience: 1999–2001. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):89–93, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.

uka.de/bibliography/Misc/DBLP/2002.bib. Women and

Computing.

Chmiel:2004:DNE

[CL04] Ryan Chmiel and Michael C. Loui. Debugging: from novice to expert. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):17–21, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lai:2005:TCA

[cL05] Yiu chi Lai. Teaching computer applications to preschool teachers through problem based learning approach. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4):89–92, December 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2005.bib.

Canosa:2008:MTR

[CL08a] Roxanne L. Canosa and Joan M. Lucas. Mock trials and roleplaying in computer ethics courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (1):148–152, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Corliss:2008:BCJ

[CL08b] Marc L. Corliss and E. Christopher Lewis. Bantam: a customizable, Java-based, classroom compiler. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):38–42, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Clark:2000:GPT

[Cla00] Martyn Clark. Getting participation through discussion. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):129–133, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clark:2005:HFS

[Cla05] Eric Clark. Hacking as a form of “self-improvement”. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):397, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Caspersen:2007:MMP

[CLB07] Michael E. Caspersen, Kasper Dalgaard Larsen, and Jens Bennedsen. Mental models and programming aptitude. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):206–210, September 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Clear:2000:CVC

[Cle00a] Tony Clear. Competition versus cooperation: models for computer education? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(4):11–12, December 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clear:2000:PED

[Cle00b] Tony Clear. Practitioner education — degrees of difference? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(2):11–12, June 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clear:2001:PLN

[Cle01a] Tony Clear. “programming in the large” and the need for professional discrimination. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(4):9–10, December 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2001.bib.

Clear:2001:RPN

[Cle01b] Tony Clear. Research paradigms and the nature of meaning and truth. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(2):9–10, June 2001. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2001.bib.

Clear:2002:DUS

[Cle02a] Tony Clear. Design and usability in security systems: daily life as a context of use? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):13–14, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2002.bib.

Clear:2002:DTA

[Cle02b] Tony Clear. A diagnostic technique for addressing group performance in capstone projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3): 196, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clear:2003:DAM

[Cle03a] Tony Clear. Documentation and agile methods: striking a balance. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):12–13, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

Clear:2003:WDL

[Cle03b] Tony Clear. The waterfall is dead..: long live the waterfall!! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):13–14, December 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

Clear:2004:SEA

[Cle04a] Tony Clear. Software engineering and the academy: uncomfortable bedfellows? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(2):14– 15, June 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2004.bib.

Clear:2004:SBP

[Cle04b] Tony Clear. Students becoming political and “incorrect” through agile methods. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(4):13–15, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2004.bib.

Clear:2005:CLC

[Cle05a] Tony Clear. Comprehending large code bases — the skills required for working in a “brown fields” environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):12–14, June 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2005.bib.

Clear:2005:DDP

[Cle05b] Tony Clear. Disciplined design practices: a role for refactoring in software engineering? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4): 15–16, December 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2005.bib.

Clear:2006:GDN

[Cle06a] Tony Clear. Google^{T M} — “do no evil”: yeah right! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):8–10, December 2006. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2006.bib.

Clear:2006:NRC

[Cle06b] Tony Clear. On the necessity of removing “cruelty” from the teaching of computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2): 8–10, June 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2006.bib.

Clear:2007:CCP

[Cle07a] Tony Clear. Computing capstone projects and the role of failure in education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):13–15, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2007.bib.

Clear:2007:CRT

[Cle07b] Tony Clear. Computing relationships: transactional algorithms yield to social networks. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2): 9–10, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2007.bib.

Clear:2008:GCC

[Cle08a] Tony Clear. Global collaboration in course delivery: are we there yet? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):11–12, June 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2008.bib.

Clear:2008:TIA

[Cle08b] Tony Clear. Thinking issues: assessment in computing education: measuring performance or conformance? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):13–15, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clear:2009:TIS

[Cle09a] Tony Clear. Thinking issues: Strategies for answering examination questions: how do novice programmers build a theory of the program? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4):9–12, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clear:2009:TIT

[Cle09b] Tony Clear. Thinking ISsues: the three p’s of capstone project performance. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):69–70, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Clifton:2001:CSA

[Cli01] Joseph M. Clifton. A CS/SE approach to a real-time embedded systems software development course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):278–281, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cliff:2004:TAN

[Cli04] David Cliff. Teaching about natural systems, the next generation of computers, and the generation after that? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):1, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cliburn:2006:CCL

[Cli06] Daniel C. Cliburn. A CS0 course for the liberal arts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):77–81, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Concepcion:2000:AER

[CLK00] Arturo I. Concepcion, Nathan Leach, and Allan Knight. Algorithma 99: an experiment in reusability & component based software engineering. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):162– 166, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chen:2006:WDB

[CLM^{+}06] Tzu-Yi Chen, Gary Lewandowski, Robert McCartney, Kate Sanders, and Beth Simon. What do beginning students know, and what can they do? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):329, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chen:2007:CCU

[CLM^{+}07] Tzu-Yi Chen, Gary Lewandowski, Robert McCartney, Kate Sanders, and Beth Simon. Commonsense computing: using student sorting abilities to improve instruction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):276–280, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cassel:2009:CPC

[CLMW09] Lillian Cassel, Richard LeBlanc, Andrew McGettrick, and Michael Wrinn. Concurrency and parallelism in the computing ontology. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):402, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Curtin:2001:NBA

[CLP^{+}01] Dan Curtin, Gary Lewandowski, Carla Purdy, Dennis Gibson, and Lisa Meeden. The nuts and bolts of academic careers: a primer for students and beginning faculty. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):414–415, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Crescenzi:2006:ACJ

[CLP06] Pilu Crescenzi, Michele Loreti, and Rosario Pugliese. Assessing CS1 Java skills: a three-year experience. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):348, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cavender:2009:SAA

[CLR09] Anna C. Cavender, Richard E. Ladner, and Robert I. Roth. The Summer Academy for Advancing Deaf and Hard of Hearing in Computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):514–518, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Lai:2007:IPS

[cLwW07] Yiu chi Lai and Tak wah Wong. Infusing problem solving skills into computer lessons. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4): 84–86, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2007.bib.

Lai:2009:DCC

[cLwW09] Yiu chi Lai and Tak wah Wong. Developing creativity in computer lessons. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):132–135, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ciesielski:2001:UAS

[CM01] Vic Ciesielski and Peter McDonald. Using animation of state space algorithms to overcome student learning difficulties. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):97–100, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chamillard:2002:MCL

[CM02] A. T. Chamillard and Laurence D. Merkle. Management challenges in a large introductory computer science course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):252–256, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Casado:2005:VNS

[CM05a] Martin Casado and Nick McKeown. The virtual network system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):76–80, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chen:2005:IPF

[CM05b] Stephen Chen and Stephen Morris. Iconic programming for flowcharts, Java, Turing, etc. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):104– 107, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Callele:2006:TRE

[CM06a] David Callele and Dwight Makaroff. Teaching requirements engineering to an unsuspecting audience. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):433–437, March 2006. CODEN SIGSD3. ISBN 1-59593-259-3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2006.bib;
ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.uka.de/
bibliography/SE/alspaugh.bib.

Cortina:2006:DHC

[CM06b] Thomas J. Cortina and Richard McKenna. The design of a history of computing course with a unique perspective. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):67–71, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cliburn:2008:GSS

[CM08a] Daniel C. Cliburn and Susan Miller. Games, stories, or something more traditional: the types of assignments college students prefer. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):138–142, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Curzon:2008:ECS

[CM08b] Paul Curzon and Peter W. McOwan. Engaging with computer science through magic shows. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (3):179–183, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Curzon:2009:EIR

[CMCB09] Paul Curzon, Peter W. McOwan, Quintin I. Cutts, and Tim Bell. Enthusing & inspiring with reusable kinaesthetic activities. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):94–98, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Cassel:2007:IAS

[CMD^{+}07] Lillian N. Cassel, Andrew McGettrick, Gordon Davies, Heikki Topi, and Bob Sloan. An initiative to attract students to computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):133–134, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Charlton:2008:EEW

[CMD08] Terence Charlton, Lindsay Marshall, and Marie Devlin. Evaluating the extent to which sociability and social presence affects learning performance. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):342, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Cassel:2007:CCC

[CMGR07] Lillian (Boots) Cassel, Andrew McGettrick, Mark Guzdial, and Eric Roberts. The current crisis in computing: what are the real issues? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):329–330, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Canning:2004:RTC

[CMRR04] James Canning, William Moloney, Ali Rafieymehr, and Demetrio Rey. Reading types in C using the right left walk method. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(4):52–54, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2004.bib.

Cassel:2006:CRC

[CMS06a] Lillian (Boots) Cassel, Andrew McGettrick, and Robert H. Sloan. A comprehensive representation of the computing and information disciplines. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1): 199–200, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chen:2006:REP

[CMS06b] Tzu-Yi Chen, Alvaro Monge, and Beth Simon. Relationship of early programming language to novice generated design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):495–499, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cohen:2006:TGV

[CMS06c] Robert F. Cohen, Arthur Meacham, and Joelle Skaff. Teaching graphs to visually impaired students using an active auditory interface. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):279–282, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chinn:2007:TWS

[CMS07] Donald Chinn, Kristofer Martin, and Catherine Spencer. Treisman workshops and student performance in CS. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):203–207, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cunningham:2004:TAN

[CMT04a] Steve Cunningham, Ernest McDuffie, and Harriet Taylor. Taking advantage of national science foundation funding opportunities. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):48, March 2004. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cunningham:2004:UNS

[CMT04b] Steve Cunningham, Ernest McDuffie, and Harriet Taylor. The US national science foundation and the ITiCSE effort. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):258, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Crescenzi:2007:FIA

[CN07] Pilu Crescenzi and Carlo Nocentini. Fully integrating algorithm visualization into a CS2 course.: a two-year experience. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):296–300, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Cheung:2009:FGP

[CNCL09] Joey C. Y. Cheung, Grace Ngai, Stephen C. F. Chan, and Winnie W. Y. Lau. Filling the gap in programming instruction: a text-enhanced graphical programming environment for junior high students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):276– 280, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Cooper:2009:ECS

[CNP^{+}09] Stephen Cooper, Christine Nickell, Victor Piotrowski, Brenda Oldfield, Ali Abdallah, Matt Bishop, Bill Caelli, Melissa Dark, E. K. Hawthorne, Lance Hoffman, Lance C. Pérez, Charles Pfleeger, Richard Raines, Corey Schou, and Joel Brynielsson. An exploration of the current state of information assurance education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4): 109–125, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chase:2000:CCL

[CO00] J. D. Chase and Edward G. Okie. Combining cooperative learning and peer instruction in introductory computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):372–376, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cohoon:2002:WCB

[Coh02a] J. McGrath Cohoon. Women in CS and biology. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):82–86, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Cohoon:2002:RRW

[Coh02b] Joanne McGrath Cohoon. Recruiting and retaining women in undergraduate computing majors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (2):48–52, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/

DBLP/2002.bib. Women and Computing.

Cohoon:2007:GEC

[Coh07a] J. McGrath Cohoon. Gendered experiences of computing graduate programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):546–550, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cohoon:2007:ICF

[Coh07b] James P. Cohoon. An introductory course format for promoting diversity and retention. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):395– 399, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Congdon:2000:MLL

[Con00] Clare Bates Congdon. Machine learning in the liberal arts curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):100–104, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Conn:2004:RAS

[Con04] Richard Conn. A reusable, academic-strength, metricsbased software engineering process for capstone courses and projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):492–496, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Connolly:2008:CMT

[Con08] Randy W. Connolly. Complecto mutatio: teaching software design best practices using multi-platform development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):345, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Chase:2007:ULP

[COR07a] J. D. Chase, Ed Oakes, and Sean Ramsey. Using live projects without pain: the development of the small project support center at radford university. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):469– 473, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cornell:2007:CC

[Cor07b] Lee D. Cornell. CS2 for an IS curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):340, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Cortina:2007:ICS

[Cor07c] Thomas J. Cortina. An introduction to computer science for non-majors using principles of computation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):218–222, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cox:2005:PHA

[Cox05] David Cox. A pragmatic HCI approach: engagement by reinforcing perception with functional design and programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):39–43, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Coy:2004:BD

[Coy04] Wolfgang Coy. Between the disciplines. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(2):7–10, June 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2004.bib.

Countermine:2000:IIC

[CP00] Terry Countermine and Phil Pfeiffer. Implementing an IT concentration in a CS department: content, rationale, and initial impact. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):275–279, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Campbell:2003:LC

[CP03] Andrea Beth Campbell and Roy P. Pargas. Laptops in the classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):98–102, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carson:2007:AEV

[CPJ07] Erik Carson, Ian Parberry, and Bradley Jensen. Algorithm explorer: visualizing algorithms in a 3D multimedia environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):155–159, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cannon:2007:SAR

[CPP07] Kelly R. Cannon, Katherine A. Panciera, and Nikolaos P. Papanikolopoulos. Second annual robotics summer camp for underrepresented students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):14– 18, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Cunningham:2002:CSE

[CPST02] Steve Cunningham, Sylvia Clark Pulliam, Charles D. Swanson, and Peter R. Turner. Computational science and engineering: tools and techniques for teaching. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):135–136, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Craig:2002:GVC

[CPT02] Annemieke Craig, Rose Paradis, and Eva Turner. A gendered view of computer professionals: preliminary results of a survey. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):101–104, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2002.bib. Women and

Computing.

Cooper:2009:ENF

[CPT09a] Stephen Cooper, Joan Peckham, and Harriet Taylor. Exploring NSF funding opportunities in EHR and CISE. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):67–68, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Curzon:2009:CTC

[CPT^{+}09b] Paul Curzon, Joan Peckham, Harriet Taylor, Amber Settle, and Eric Roberts. Computational thinking (CT): on weaving it in. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):201–202, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Comer:2002:TJB

[CR02] James Comer and Robert Roggio. Teaching a Java-based CS1 course in an academically-diverse environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):142–146, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Clarke:2003:IES

[CR03] Faith Clarke and Han Reichgelt. The importance of explicitly stating educational objectives in computer science curricula. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):47–50, December 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

Craig:2007:FSD

[Cra07] Michelle Craig. Facilitated student discussions for evaluating teaching. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):190–194, March 2007.

CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Christensen:2006:CVS

[CRPZ06] K. Christensen, D. Rundus, G. Perera, and S. Zulli. CSE volunteers: a service learning program to provide IT support to the Hillsborough County School District. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):229–233, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carr:2000:PCL

[CS00a] Steve Carr and Ching-Kuang Shene. A portable class library for teaching multithreaded programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):124–127, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chong:2000:MMC

[CS00b] Ng S. T. Chong and Masao Sakauchi. A multi-modal chat for coordinated interaction (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):187, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Carbone:2002:SBT

[CS02] Angela Carbone and Judy Sheard. A studio-based teaching and learning model in IT: what do first year students think? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):213–217, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Crouch:2003:CSA

[CS03a] Donald B. Crouch and Leslie Schwartzman. Computer science accreditation: the advantages of being different. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):36–40, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cunningham:2003:CGU

[CS03b] Steve Cunningham and Angela B. Shiflet. Computer graphics in undergraduate computational science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):372–375, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chamillard:2005:LSA

[CS05] A. T. Chamillard and Ricky E. Sward. Learning styles across the curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):241–245, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cutler:2006:WTI

[CS06] Robb Cutler and Chris Stephenson. Working together to improve K-12 computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):201, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chidanandan:2008:APB

[CS08] Archana Chidanandan and Shannon Sexton. Adopting penbased technology to facilitate active learning in the classroom: is it right for you? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):343, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Cassen:2007:VLE

[CSA^{+}07] T. Cassen, K. R. Subramanian, Jeffrey Alexander, Drew Linderman, and A. Nasipuri. A visual learning engine for interactive generation of instructional materials. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):319, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Cohoon:2003:FTC

[CSC03] J. McGrath Cohoon, Rebecca Shwalb, and Lih-Yuan Chen. Faculty turnover in CS departments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35 (1):108–112, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cukier:2002:GIT

[CSD02] Wendy L. Cukier, Denise Shortt, and Irene Devine. Gender and information technology: implications of definitions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):142–148, December 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2002.bib.

Cassel:2007:COP

[CSD^{+}07] Lillian N. Cassel, Robert H. Sloan, Gordon Davies, Heikki Topi, and Andrew McGettrick. The computing ontology project: the computing education application. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):519–520, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Caronna:2007:PMP

[CSM^{+}07a] Jason S. Caronna, Rojita Sharma, Jonathan D. Marra, Virginia L. Iuorno, Katherine G. Herbert, and Jeffrey H. Toney. Prediction of modulators of pyruvate kinase in SMILES text using a priori methods. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3): 348, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Chinn:2007:PSS

[CSM07b] Donald Chinn, Catherine Spencer, and Kristofer Martin. Problem solving and student performance in data structures and algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):241– 245, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Costa-Soria:2009:TSA

[CSP09] Cristóbal Costa-Soria and Jennifer Pérez. Teaching software architectures and aspect-oriented software development using open-source projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):385, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Cassel:2005:SOA

[CSS05] Lillian (Boots) Cassel, Russell Shackelford, and Robert H. Sloan. A synthesis and ontology of all of computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):65–66, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cukierman:2007:LSS

[CT07] Diana Cukierman and Donna McGee Thompson. Learning strategies sessions within the classroom in computing science university courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):341, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Cukierman:2009:AEP

[CT09] Diana Cukierman and Donna McGee Thompson. The academic enhancement program: encouraging students to learn about learning as part of their computing science courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):171–175, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Costa:2006:RHF

[CTBS06] Monica I. Costa, Carmem Tavares, João Barroso, and Salviano Soares. Reconstructed high frame rate sequences quality measurement tool. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):333, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Collins:2003:RFL

[CTLW03] William Collins, Josh Tenenberg, Raymond Lister, and Suzanne Westbrook. The role for framework libraries in CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):403–404, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chase:2004:ATB

[CTP^{+}04] J. D. Chase, Dennie Templeton, Michael Peterson, Carlotta Eaton, Diane Wolff, and Edward G. Okie. Articulation through a body of knowledge: a report on an NSF ATE project involving Radford University and the Virginia Community College System. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):149– 150, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chandrasekar:2009:WTP

[CTP09] Samantha Chandrasekar, Joseph G. Tront, and Jane C. Prey. WriteOn1.0: a tablet PC-based tool for effective classroom instruction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):323–327, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Clancy:2003:NRS

[CTR^{+}03] Michael Clancy, Nate Titterton, Clint Ryan, Jim Slotta, and Marcia Linn. New roles for students, instructors, and computers in a lab-based introductory programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):132–136, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Culwin:2000:LWB

[Cul00] Fintan Culwin. Lecturelets: web based Java enabled lectures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):5–8, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cunningham:2000:PCC

[Cun00] Steve Cunningham. Powers of 10: the case for changing the first course in computer graphics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (1):46–49, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cunningham:2002:GPS

[Cun02] Steve Cunningham. Graphical problem solving and visual communication in the beginning computer graphics course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):181–185, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Cunningham:2005:VEN

[Cun05] Steve Cunningham. The voice of experience: national science foundation funded projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):128– 129, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cunningham:2007:CGC

[Cun07] Steve Cunningham. A computer graphics curriculum to meet the European Bologna requirements. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (3):310, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Curran:2003:TSE

[Cur03] William S. Curran. Teaching software engineering in the computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):72–75, December 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2003.bib.

Curzon:2007:SFC

[Cur07] Paul Curzon. Serious fun in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):1, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Cutter:2007:HBB

[Cut07] Pamela Cutter. Having a BLAST: a bioinformatics project in CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):353–357, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chinn:2008:WSS

[CV08] Donald Chinn and Tammy VanDeGrift. What students say about gender in hiring software professionals. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):344, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Cameron:2003:ESH

[CW03] Brian H. Cameron and Kay Wijekumar. The effectiveness of simulation in a hybrid and on-line networking course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):117–119, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cohoon:2009:STW

[CWC09] J. McGrath Cohoon, Zhen Wu, and Jie Chao. Sexism: toxic to women’s persistence in CSE doctoral programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):158–162, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Carlisle:2005:RVP

[CWHH05] Martin C. Carlisle, Terry A. Wilson, Jeffrey W. Humphries, and Steven M. Hadfield. RAPTOR: a visual programming environment for teaching algorithmic problem solving. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):176–180, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Chen:2006:DAA

[CWHL06] Ming-Yu Chen, Jyh-Da Wei, Jeng-Hung Huang, and D. T. Lee. Design and applications of an algorithm benchmark system in a computational problem solving environment.

SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):123–127, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cohoon:2008:WTS

[CWL08] J. McGrath Cohoon, Zhen Wu, and Luo Luo. Will they stay or will they go? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):397–401, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Casado:2005:TNH

[CWM05] Martin Casado, Gregory Watson, and Nick McKeown. Teaching networking hardware. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):208– 212, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Cassel:2009:CDE

[CWP09] Lillian N. Cassel, Thomas Way, and Sridhara Potluri. CPATH: distributed expertise — collaborating with other disciplines. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):382, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Clear:2001:CER

[CY01] Tony Clear and Alison Young. Are computing educators and researchers different from the rest? (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):173, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

DAntonio:2003:IBA

[D’A03] Lawrence D’Antonio. Incorporating bioinformatics in an algorithms course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):211–214, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Davy:2000:SWB

[DABJ00] J. R. Davy, K. Audin, M. Barkham, and C. Joyner. Student well-being in a computing department. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):136–139, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dickson:2008:ACI

[DAH08] Paul E. Dickson, W. Richards Adrion, and Allen R. Hanson. Automatic creation of indexed presentations from classroom lectures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):12–16, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Dickson:2009:FEC

[DAHA09] Paul E. Dickson, W. Richards Adrion, Allen R. Hanson, and David T. Arbour. First experiences with a classroom recording system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):298–302, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Diaz-Agudo:2008:HTS

[DAJDRG08] Belén Diaz-Agudo, Guillermo Jiménez-Diaz, and Juan A. Recio-Garcia. How to teach Semantic Web?: a project-based approach. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):98–102, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Dodds:2007:BFC

[DAKLH07] Zachary Dodds, Christine Alvarado, Geoff Kuenning, and Ran Libeskind-Hadas. Breadth-first CS 1 for scientists. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):23–27, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Dale:2000:RPR

[Dal00] Nell Dale. Reflections on past research: part II. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(2):14–16, June 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dale:2001:A

[Dal01] Nell Dale. ACE 2000. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(2):12– 14, June 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2001.bib.

Dale:2002:IIC

[Dal02] Nell Dale. Increasing interest in CS ed research. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):16–17, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2002.bib.

Dale:2005:CEC

[Dal05] Nell Dale. Content and emphasis in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4):69–73, December 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2005.bib.

Dale:2006:MDT

[Dal06] Nell B. Dale. Most difficult topics in CS1: results of an online survey of educators. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):49– 53, June 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2006.bib.

Davies:2001:DLU

[Dav01a] Gordon Davies. Distance learning at the University of Texas–Pan American. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(2):22– 23, June 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2001.bib.

Davies:2001:UDE

[Dav01b] Gordon Davies. USC’s distance education network (DEN). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(4):22–23, December 2001. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2001.bib.

Davoli:2004:TOS

[Dav04] Renzo Davoli. Teaching operating systems administration with User Mode Linux. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):112– 116, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Davis:2005:M

[Dav05] Leon Davis. MindMapX. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):405, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Davies:2006:ML

[Dav06] Gordon Davies. On the merits of e-learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):16–17, June 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2006.bib.

Davis:2009:EJT

[Dav09] Janet Davis. Experiences with just-in-time teaching in systems and design courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):71– 75, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Dale:2002:DBA

[DBBK02] Nell Dale, Judith Bishop, David J. Barnes, and Christoph Kessler. A dialog between authors and teachers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):44–45, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

deBlas:2009:ALC

[dBGdMB09] Juan Manuel de Blas, José María Gutiérrez, Luis de Marcos, and Roberto Barchino. Automatic E-learning contents composition by using gap analysis techniques. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):369, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

DAntonio:2009:URC

[DBK^{+}09] Lawrence D’Antonio, Roger Boyle, Amruth Kumar, Logan Muller, Claudia Roda, and Matti Tedre. Undergraduate research in CS: a global perspective. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (3):139–140, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Dahlberg:2008:IRG

[DBR^{+}08] Teresa Dahlberg, Tiffany Barnes, Audrey Rorrer, Eve Powell, and Lauren Cairco. Improving retention and graduate recruitment through immersive research experiences for undergraduates. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):466–470, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Dann:2000:MCP

[DCP00] Wanda Dann, Stephen Cooper, and Randy Pausch. Making the connection: programming with animated small world. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):41–44, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dann:2001:UVT

[DCP01] Wanda Dann, Stephen Cooper, and Randy Pausch. Using visualization to teach novices recursion. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):109–112, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Davies:2009:AEB

[DCP^{+}09] Gordon Davies, Lillian N. Cassel, Arthur Pyster, Michael Caspersen, and Heikki Topi. ACM Education Board and Masters level programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):487–488, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Davies:2006:UCO

[DCT06] Gordon Davies, Lillian N. Cassel, and Heikki Topi. Using a computing ontology for educational purposes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):334, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dionisio:2008:ICS

[DD08] John David N. Dionisio and Kam D. Dahlquist. Improving the computer science in bioinformatics through open source pedagogy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):115–119, June 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2008.bib.

Dionisio:2007:OSS

[DDA^{+}07] John David N. Dionisio, Caskey L. Dickson, Stephanie E. August, Philip M. Dorin, and Ray Toal. An open source software culture in the undergraduate computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):70–74, June 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Dann:2003:OVB

[DDC^{+}03a] Wanda Dann, Toby Dragon, Stephen Cooper, Kevin Dietzler, Kathleen Ryan, and Randy Pausch. Objects: visualization of behavior and state. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):84–88, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dougherty:2003:ITF

[DDC^{+}03b] John P. Dougherty, Tom Dececchi, Tony Clear, Brad Richards, Stephen Cooper, and Tadeusz Wilusz. Information technology fluency in practice. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35 (2):153–171, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2003.bib.

Darbhamulla:2003:SSM

[DDL03] Ravikumar Darbhamulla, Manirupa Das, and Pamela Lawhead. Student status monitoring tool (SSM): proxy for the real world expert in online course delivery. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):257, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Davies:2001:QDE

[DDLFS01] Gordon Davies, Wendy Doube, Wendy Lawrence-Fowler, and Dale Shaffer. Quality in distance education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):392–393, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Denning:2001:MCI

[DDLR01] Peter J. Denning, Wayne Dyksen, Richard LeBlanc, and Edward Robertson. Model curricula for IT schools: report of a curriculum committee. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):431–432, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Denning:2004:PSG

[DDRG04] Peter J. Denning, Rudolph Darken, Eric Roberts, and Mark Guzdial. Panel session: great principles in computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):447, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

DiCerbo:2008:EMC

[DDS08] Francesco Di Cerbo, Gabriella Dodero, and Giancarlo Succi. Extending Moodle for collaborative learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):324, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Debray:2002:MCD

[Deb02] Saumya Debray. Making compiler design relevant for students who will (most likely) never design a compiler. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):341–345, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Debray:2004:WEP

[Deb04] Saumya Debray. Writing efficient programs: performance issues in an undergraduate CS curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):275–279, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Deibel:2005:TFM

[Dei05] Katherine Deibel. Team formation methods for increasing interaction during in-class group work. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37 (3):291–295, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Deibel:2007:SOI

[Dei07] Katherine Deibel. Studying our inclusive practices: course experiences of students with disabilities. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):266–270, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Deibel:2008:CEC

[Dei08] Katherine Deibel. Course experiences of computing students with disabilities: four case studies. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (1):454–458, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Dairaine:2006:IRE

[DEJ^{+}06] L. Dairaine, E. Exposito, G. Jourjon, P. Casenove, F. Tan, and E. Lochin. IREEL: remote experimentation with real protocols and applications over emulated network. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):314, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Delgado:2000:EKE

[Del00] Gladys García Delgado. Ethical knowledge for an electronic era (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):187–188, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Demaille:2005:MCC

[Dem05] Akim Demaille. Making compiler construction projects relevant to core curriculums. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):266– 270, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Denning:2004:GPC

[Den04] Peter J. Denning. Great principles in computing curricula. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):336–341, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

DePasquale:2006:ELD

[DeP06] Peter DePasquale. Exploiting on-line data sources as the basis of programming projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):283– 287, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Deugo:2008:ETT

[Deu08a] Dwight Deugo. Eclipse as a teaching tool. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):318, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Deugo:2008:UEC

[Deu08b] Dwight Deugo. Using Eclipse in the classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):322, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Dewan:2005:TIO

[Dew05] Prasun Dewan. Teaching inter-object design patterns to freshmen. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):482–486, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

DePalma:2004:CCS

[DFGH04] Paul De Palma, Charles Frank, Suzanne Gladfelter, and Joshua Holden. Cryptography and computer security for undergraduates. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):94–95, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Denegri:2008:TTI

[DFGM08] Ezequiel Denegri, Guillermo Frontera, Antonio Gavilanes, and Pedro J. Martín. A tool for teaching interactions between design patterns. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):371, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Dodero:2003:ESP

[DFS03] Juan Manuel Dodero, Camino Fernández, and Daniel Sanz. An experience on students’ participation in blended vs. online styles of learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):39–42, December 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2003.bib.

Davoli:2005:VSV

[DG05] Renzo Davoli and Michael Goldweber. Virtual square (V^{2}) in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):301– 305, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Davies:2009:FMF

[DG09] Jim Davies and Jeremy Gibbons. Formal methods for future interoperability. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):60–64, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

DiGiano:2003:CET

[DGHC03] Chris DiGiano, Mike Griffin, Jeff Huang, and Mark Chung. Consolidating Ed-tech co-design best practices through the TRAILS project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):245, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Davis:2004:FS

[DGMW04] Timothy Davis, Robert Geist, Sarah Matzko, and James Westall. τϵ́χνη: a first step. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):125– 129, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Davis:2007:TPN

[DGMW07] Timothy Davis, Robert Geist, Sarah Matzko, and James Westall. τϵχνη: trial phase for the new curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):415–419, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Davoli:2006:PAS

[DGS06] Renzo Davoli, Michael Goldweber, and Paola Salomoni, editors. Proceedings of the 11th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education, ITiCSE 2006, Bologna, Italy, June 26–28, 2006, SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 2006. CODEN SIGSD3. ISBN 1-59593-055-8. ISSN 0097-8418 (print), 2331-3927 (electronic).

Denning:2006:MCC

[DGSW06] Tamara Denning, William G. Griswold, Beth Simon, and Michelle Wilkerson. Multimodal communication in the classroom: what does it mean for us? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (1):219–223, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Daly:2005:PP

[DH05] Charlie Daly and Jane Horgan. Patterns of plagiarism. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):383–387, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Deremer:2006:IUS

[DH06] Dorothy Deremer and Katherine G. Herbert. An interdisciplinary undergraduate science informatics degree in a liberal arts context. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):179–183, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dougherty:2004:PTF

[DHGR04] John P. Dougherty, Thomas B. Horton, Daniel D. Garcia, and Susan H. Rodger. Panel on teaching faculty positions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):231–232, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Drysdale:2005:YRC

[DHRH05] Scot Drysdale, Judith Hromcik, David Reed, and Reg Hahne. The year in review: Changes and lessons learned in the design and implementation of the AP CS Exam in Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):323–324, March 2005. CODEN SIGSD3. ISBN 1-58113-997-7. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Compiler/java.bib;   ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/2005.
bib.

Drysdale:2003:JMJ

[DHWH03] Scot Drysdale, Judith Hromcik, Mark Allen Weiss, and Reg Hahne. Java in the morning… Java in the evening… Java in 2004. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):271–272, January 2003. CODEN SIGSD3. ISBN 1-58113-648-X. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Compiler/java.bib;   ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/2003.
bib.

DiCosmo:2006:EC

[Di 06] Roberto Di Cosmo. Educating the e-citizen. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):1, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dias:2005:MML

[Dia05] Artur Miguel Dias. Moving from mailing lists to Web-based discussion forums. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):373, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dickinson:2000:OSP

[Dic00] John Dickinson. Operating systems projects built on a simple hardware simulator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):320–324, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dick:2005:SIT

[Dic05] Martin Dick. Student interviews as a tool for assessment and learning in a systems analysis and design course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):24–28, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dickinson:2007:DRT

[Dic07a] Anne Dickinson. ’but I don’t read text printed in Braille font’…: parables from a business and information technology student. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):336, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Dickinson:2007:SQW

[Dic07b] Anne Dickinson. Is the shortcut the quickest way to go?: translating instructions for keyboard navigation and other stories. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):358, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Dickinson:2007:TVN

[Dic07c] Anne Dickinson. ’touching the void’…: narrowing the crevasse between the tactile diagrams that tutors request and those that work in reality. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3): 339, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Dixon:2004:SCE

[Dix04] Mark Dixon. A single CASE environment for teaching and learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):271, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Desai:2009:IIT

[DJC09] Chetan Desai, David S. Janzen, and John Clements. Implications of integrating test-driven development into CS1/CS2 curricula. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):148–152, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Dairaine:2007:IRE

[DJLA07] Laurent Dairaine, Guillaume Jourjon, Emmanuel Lochin, and Sebastien Ardon. IREEL: remote experimentation with real protocols and applications over an emulated network. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):92–96, June 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Djordjevic:2008:JPM

[Djo08] Mirela Djordjevic. Java projects motivated by student interests. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):321, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Djordjevic:2009:PAC

[Djo09] Mirela Djordjević. Progressive assignment in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):346, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Desai:2008:SET

[DJS08] Chetan Desai, David Janzen, and Kyle Savage. A survey of evidence for test-driven development in academia. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):97–101, June 2008. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2008.bib.

Diwan:2005:PDE

[DJWD05] Amer Diwan, Michele H. Jackson, William M. Waite, and Jacob Dickerson. PL-detective: experiences and results. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):221–225, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Diehl:2001:LE

[DK01] Stephan Diehl and Andreas Kerren. Levels of exploration. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):60–64, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dodds:2006:ECO

[DK06] Zachary Dodds and Leslie Karp. The evolution of a computational outreach program to secondary school students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):448–452, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Davis:2006:RCS

[DKG06] Timothy A. Davis and John Kundert-Gibbs. The role of computer science in digital production arts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):73–77, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Denning:2007:LPP

[DKL^{+}07] Tamara Denning, Michael Kelly, David Lindquist, Roshni Malani, William G. Griswold, and Beth Simon. Lightweight preliminary peer review: does in-class peer review make sense? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):266–270, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Doyle:2008:ECF

[DKN08] Maureen Doyle, Kevin G. Kirby, and Gary Newell. Engaging constructions: family-based computing experiences for immigrant middle school students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):58– 62, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

DeHart:2006:ONL

[DKP^{+}06] John DeHart, Fred Kuhns, Jyoti Parwatikar, Jonathan Turner, Charlie Wiseman, and Ken Wong. The open network laboratory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):107–111, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Draper:2009:HCC

[DKR09] Geoffrey M. Draper, Robert R. Kessler, and Richard F. Riesenfeld. A history of computing course with a technical focus. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):458–462, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Dony:2006:THT

[DL06] Isabelle Dony and Baudouin Le Charlier. A tool for helping teach a programming method. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3): 212–216, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dodds:2008:EBF

[DLHAK08] Zachary Dodds, Ran Libeskind-Hadas, Christine Alvarado, and Geoff Kuenning. Evaluating a breadth-first CS 1 for scientists. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):266–270, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Doerschuk:2007:PSC

[DLM07] Peggy Doerschuk, Jiangjiang Liu, and Judith Mann. Pilot summer camps in computing for middle school girls: from organization through assessment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (3):4–8, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Doerschuk:2009:IBP

[DLM09] Peggy Doerschuk, Jiangjiang Liu, and Judith Mann. IN-SPIRED broadening participation: first year experience and lessons learned. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):238–242, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Demaille:2008:STT

[DLP08] Akim Demaille, Roland Levillain, and Benoît Perrot. A set of tools to teach compiler construction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (3):68–72, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

DePasquale:2004:ESP

[DLPQ04] Peter DePasquale, John A. N. Lee, and Manuel A. Pérez-Quiñones. Evaluation of subsetting programming language elements in a novice’s programming environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):260–264, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Denny:2008:SUP

[DLRH08] Paul Denny, Andrew Luxton-Reilly, and John Hamer. Student use of the PeerWise system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (3):73–77, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Denny:2009:CCT

[DLRHP09] Paul Denny, Andrew Luxton-Reilly, John Hamer, and Helen Purchase. Coverage of course topics in a student generated MCQ repository. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):11–15, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

deMarcos:2008:MCS

[dMBG^{+}08] Luis de Marcos, Roberto Barchino, José-Antonio Gutiérrez, Juan-Manuel de Blas, José-Ramón Hilera, and Salvador Oton. A multidisciplinary computer science master program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):358, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

deMarcos:2008:EAC

[dMBM08] Luis de Marcos, Roberto Barchino, and José-Javier Martinez. Evolutionary approaches for curriculum sequencing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):373, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

de-Marcos:2009:TSM

[dMFM09] Luis de Marcos, Fernando Flores, and José-Javier Martínez. The two states of the mind to teach UML. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):361, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Dale:2009:HLC

[DMI^{+}09] Nell Dale, Andrew McGettrick, John Impagliazzo, Robert Aiken, Elliot Koffman, and Jim Leisy. A historical look at curricula and materials. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):197–198, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Duplantis:2002:VFA

[DMKN02] Willa Duplantis, Eve MacGregor, Maria M. Klawe, and Michele Ng. ’virtual family’: an approach to introducing Java programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):40–43, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/

2002.bib. Women and Computing.

Dale:2001:WTW

[DMKS01] Nell Dale, Rick Mercer, Elliot Koffman, and Walter Savitch. Writing a textbook: walking the gauntlet. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):408–409, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Denning:2006:RCC

[DMRS06] Peter Denning, Andrew McGettrick, Paul Rosenbloom, and Larry Snyder. Re-centering computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):65–66, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dershem:2002:AJL

[DMU02] Herbert L. Dershem, Ryan L. McFall, and Ngozi Uti. Animation of Java linked lists. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):53– 57, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Dann:2005:PST

[DNTB05] Wanda Dann, Thomas L. Naps, Paul T. Tymann, and Doug Baldwin, editors. Proceedings of the thirty-sixth SIGCSE Technical Symposium on Computer Science Education: SIGCSE 2005, St. Louis, Missouri, February 23– 27, 2005, SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 2005. CODEN SIGSD3. ISBN 1-58113-997-7 (paperback). ISSN 0097-8418 (print), 2331-3927 (electronic). LCCN QA76.27 .A34 2005.

Dankel:2007:STS

[DO07] Douglas D. Dankel II and Jonathan Ohlrich. Students teaching students: incorporating presentations into a course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):96–99, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Guimaraes:2007:LCC

[dOG07] José de Oliveira Guimarães. Learning compiler construction by examples. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):70–74, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2007.bib.

Donaldson:2001:ADO

[Don01] John L. Donaldson. Architecture-dependent operating system project sequence. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):322–326, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Donaldson:2008:ITO

[Don08] John L. Donaldson. Implementation of threads as an operating systems project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):187–191, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Dooley:2003:SEL

[Doo03] John F. Dooley. Software engineering in the liberal arts: combining theory and practice. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):48– 51, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2003.bib.

Dooley:2004:MCS

[Doo04] John F. Dooley. Moving to CC2001 at a small college. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):196–198, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dooley:2007:ECS

[Doo07] John F. Dooley. Experiences with CC2001 at a small college. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):314, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Dooley:2008:SDC

[Doo08] John F. Dooley. A software development course for CC2001: the third time is charming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):346, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Dooley:2009:PAU

[Doo09] John F. Dooley. Peer assessments using the Moodle workshop tool. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):344, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Dorin:2007:LR

[Dor07] Philip M. Dorin. Laboratory redux. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (2):84–87, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2007.bib.

Doube:2000:DTW

[Dou00] Wendy Doube. Distance teaching workloads. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):347–351, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dee:2009:WWS

[DPBP09] Hannah M. Dee, Karen E. Petrie, Roger D. Boyle, and Reena Pau. Why are we still here?: experiences of successful women in computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):233–237, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Dick:2000:TTS

[DPM00] Martin Dick, Margot Postema, and Jan Miller. Teaching tools for software engineering education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (3):49–52, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

delRio:2004:HGA

[dR04] Agustín Cernuda del Río. How not to go about a programming assignment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(2):97–100, June 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2004.bib.

Davis:2007:FFC

[DR07] Janet Davis and Samuel A. Rebelsky. Food-first computer science: starting the first course right with PB&J. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):372–376, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Draganova:2009:UMP

[Dra09] Chrisina Draganova. Use of mobile phone technologies in learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):399, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Dougherty:2008:PST

[DRFG08] J. D. Dougherty, Susan H. Rodger, Sue Fitzgerald, and Mark Guzdial, editors. Proceedings of the 39th SIGCSE Technical Symposium on Computer Science Education, SIGCSE 2008, Portland, OR, USA, March 12–15, 2008, SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 2008. CO-DEN SIGSD3. ISBN 1-59593-799-4. ISSN 0097-8418 (print), 2331-3927 (electronic).

Doom:2002:PUB

[DRKG02] Travis Doom, Michael Raymer, Dan Krane, and Oscar Garcia. A proposed undergraduate bioinformatics curriculum for computer scientists. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):78–81, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

deRaadt:2004:TSP

[dRTW04] Michael de Raadt, Mark A. Toleman, and Richard Watson. Training strategic problem solvers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (2):48–51, June 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2004.bib.

Dorge:2008:WIT

[DS08] Christina Dörge and Carsten Schulte. What are information technology’s key qualifications? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3): 296–300, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Dick:2003:ASC

[DSB^{+}03] Martin Dick, Judy Sheard, Catherine C. Bareiss, Janet Carter, Donald Joyce, Trevor Harding, and Cary Laxer. Addressing student cheating: definitions and solutions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):172–184, June 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib;
ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.uka.de/
bibliography/Techreports/UKent.bib.

Docherty:2001:IDS

[DSBK01] Michael Docherty, Peter Sutton, Margot Brereton, and Simon Kaplan. An innovative design and studio-based CS degree. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):233–237, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dietrich:2001:RGU

[DSCU01] Suzanne W. Dietrich, Dan Suceava, Chakrapani Cherukuri, and Susan D. Urban. A reusable graphical user interface for manipulating object-oriented databases using Java and XML. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):362–366, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dagdilelis:2002:WTR

[DSE02] Vassilios Dagdilelis, Maya Satratzemi, and Georgios Evangelidis. What they really do?: attempting (once again) to model novice programmers’ behavior. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3): 244, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dagdilelis:2003:PAS

[DSF^{+}03] Vassilios Dagdilelis, Maya Satratzemi, David Finkel, Roger D. Boyle, and Georgios Evangelidis, editors. Proceedings of the 8th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education, ITiCSE 2003, Thessaloniki, Greece, June 30 — July 2, 2003, SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 2003. CO-DEN SIGSD3. ISBN 1-58113-672-2. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dick:2001:IOC

[DSM01] Martin Dick, Judy Sheard, and Selby Markham. Is it okay to cheat? — the views of postgraduate students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):61–64, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

DeRosa:2006:RSD

[DSSP06] Peter DeRosa, Kai Shen, Christopher Stewart, and Jonathan Pearson. Realism and simplicity: disk simulation for instructional OS performance evaluation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (1):308–312, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Douglas:2002:IHC

[DTL^{+}02] Sarah Douglas, Marilyn Tremaine, Laura Leventhal, Craig E. Wills, and Bill Manaris. Incorporating Human-Computer Interaction into the undergraduate computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):211–212, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Demetriadis:2003:PSS

[DTP03] Stavros Demetriadis, Evangelos Triantfillou, and Andreas Pombortsis. A phenomenographic study of students’ attitudes toward the use of multiple media for learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):183–187, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Du:2007:SSI

[DTW07] Wenliang Du, Zhouxuan Teng, and Ronghua Wang. SEED: a suite of instructional laboratories for computer SEcurity EDucation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):486–490, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dierbach:2005:ECC

[DTZZ05] Charles Dierbach, Blair Taylor, Harry Zhou, and Iliana Zimand. Experiences with a CS0 course targeted for CS1 success. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):317–320, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dietrich:2002:JDC

[DUK02] Suzanne W. Dietrich, Susan D. Urban, and Ion Kyriakides. JDBC demonstration courseware using Servlets and Java Server Pages. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):266–270, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Dietrich:2005:EXD

[DUM^{+}05] Suzanne W. Dietrich, Susan D. Urban, Hua Ma, Yang Xiao, and Shama Patel. Exploring XML for data exchange in the context of an undergraduate database curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):53–57, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Decker:2004:WCC

[DV04] Adrienne Decker and Phil Ventura. We claim this class for computer science: a non-mathematician’s discrete structures course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):442–446, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Draganova:2007:TAW

[DV07] Chrisina Draganova and Vassil Vassilev. Teaching AJAX in Web-centric courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):311, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Decker:2006:TLG

[DVE06] Adrienne Decker, Phil Ventura, and Christopher Egert. Through the looking glass: reflections on using undergraduate teaching assistants in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (1):46–50, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Daly:2004:AAP

[DW04] Charlie Daly and John Waldron. Assessing the assessment of programming ability. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):210–213, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Davis:2005:RLC

[DW05a] Hugh C. Davis and Su White. A research-led curriculum in multimedia: learning about convergence. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):29–33, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dougherty:2005:UAR

[DW05b] John P. Dougherty and David G. Wonnacott. Use and assessment of a rigorous approach to CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37 (1):251–255, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Distasio:2007:ICS

[DW07] Joseph Distasio and Thomas Way. Inclusive computer science education using a ready-made computer game framework. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):116–120, September 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Diwan:2002:ITS

[DWJ02] Amer Diwan, William M. Waite, and Michele H. Jackson. An infrastructure for teaching skills for group decision making and problem solving in programming projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):276–280, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Diwan:2004:PDS

[DWJ04] Amer Diwan, William M. Waite, and Michele H. Jackson. PL-detective: a system for teaching programming language concepts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):80–84, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Davidovic:2000:SEB

[DWT00] Alex Davidovic, James Warren, and Elena Tricina. Structural example-based adaptive tutoring system (poster session) (SEATS). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):187, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

DeLoatch:2003:SDC

[DWY03] Sandra DeLoatch, Henry Walker, and Frank H. Young. On serving as department chair: suggestions from veterans. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):187–188, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Elkateeb:2000:WBM

[EA00] Ali Elkateeb and Ala Awad. A WWW-based multimedia center for learning data communications — phase 1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(2):68–73, June 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Eshet-Alkalai:2007:HUB

[EA07] Yoram Eshet-Alkalai. Humans under bombardment: performing in a real-time world. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4): 57–60, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2007.bib.

Eagle:2008:WCT

[EB08] Michael Eagle and Tiffany Barnes. Wu’s castle: teaching arrays and loops in a game. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):245– 249, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Eagle:2009:EEE

[EB09] Michael Eagle and Tiffany Barnes. Experimental evaluation of an educational game for improved learning in introductory computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):321–325, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Edwards:2008:DCF

[EBC^{+}08] Stephen H. Edwards, Jürgen Börstler, Lillian N. Cassel, Mark S. Hall, and Joseph Hollingsworth. Developing a common format for sharing programming assignments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):167–182, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ernest:2005:WMD

[EBG^{+}05] J. C. Ernest, A. S. Bowser, S. Ghule, S. Sudireddy, J. P. Porter, D. A. Talbert, and M. J. Kosa. Weathering Mind-Storms with drizzle and DIODE in CS0. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):353, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Efopoulos:2005:WPE

[EDES05] Vassilios Efopoulos, Vassilios Dagdilelis, Georgios Evangelidis, and Maya Satratzemi. WIPE: a programming environment for novices. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):113–117, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Edmondson:2007:DDA

[Edm07a] Carol Edmondson. Dear Dorothy: agonies from first-year programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):125–126, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2007.bib.

Edmondson:2007:HFP

[Edm07b] Carol Edmondson. How to fail a programming assignment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):54–56, June 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Edmondson:2008:RWD

[Edm08a] Carol Edmondson. Real women don’t write programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):112–114, June 2008. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2008.bib.

Edmondson:2008:TTS

[Edm08b] Carol Edmondson. Teaching tales: some student perceptions of computing education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4): 103–106, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Edmondson:2009:PFY

[Edm09] Carol Edmondson. Proglets for first-year programming in Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):108–112, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Edwards:2004:UST

[Edw04] Stephen H. Edwards. Using software testing to move students from trial-and-error to reflection-in-action. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):26–30, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ellsworth:2004:QS

[EFK04] Christopher C. Ellsworth, James B. Fenwick, Jr., and Barry L. Kurtz. The Quiver System. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (1):205–209, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Erkan:2008:IBD

[EG08] Ali Erkan and Diyan Gochev. An image background detection project for a visual exploration of DFS and BFS. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):483–487, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Egan:2005:SAS

[Ega05] Mary Anne L. Egan. Students with Asperger’s syndrome in the CS classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):27–30, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Egan:2007:TWC

[Ega07] Mary Anne L. Egan. Teaching a “women in computer science” course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):216–220, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Ericson:2005:MIS

[EGB05] Barbara Ericson, Mark Guzdial, and Maureen Biggers. A model for improving secondary CS education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):332–336, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ericson:2007:ISC

[EGB07] Barbara Ericson, Mark Guzdial, and Maureen Biggers. Improving secondary CS education: progress and problems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):298–301, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Escribano:2008:MPF

[EGMS08] Carmen Escribano, Antonio Giraldo, Águeda Mata, and María Asunción Sastre. Multidisciplinary projects for first year engineering courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):366, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Escribano:2008:DTJ

[EGST08] Carmen Escribano, Antonio Giraldo, María Asunción Sastre, and Mario Trotta. Digital topology Java applet. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):329, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Eidelman:2005:FIS

[EH05] Larisa Eidelman and Orit Hazzan. Factors influencing the shrinking pipeline in high schools: a sector-based analysis of the Israeli high school system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37 (1):406–410, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Eidelman:2007:EMA

[EH07] Larisa Eidelman and Orit Hazzan. Eccles’ model of achieventrelated choices: the case of computer science studies in Israeli high schools. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):29–33, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Eisenberg:2003:CCS

[Eis03] Michael Eisenberg. Creating a computer science canon: a course of “classic” readings in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):336–340, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ekman:2006:WBI

[Ekm06] Josephine Ekman. A Web-based information centre to provide help, guidance and support for students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):363–364, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Engels:2007:PDU

[ELC07] Steve Engels, Vivek Lakshmanan, and Michelle Craig. Plagiarism detection using feature-based neural networks. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):34–38, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Elenbogen:2002:MPU

[ELE^{+}02] Bruce S. Elenbogen, John Laird, Richard Enbody, Chris Mc-Donald, Peter B. Henderson, Richard Nau, and Steve Tanimoto. Mathematics preparation for undergraduate degrees in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):98–99, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Ellis:2000:TMD

[Ell00a] Ainslie Ellis. Toolbook multimedia demonstrations for Java programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):181, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ellis:2000:PAS

[Ell00b] Ainslie E. Ellis, editor. Proceedings of the ACM SIGCSE 4th Australasian Conference on Computer Science Education, ACSE 2000, Melbourne, Victoria, Australia, 2000, volume 8 of SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 2000. CODEN SIGSD3. ISBN 1-58113-271-9. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ellis:2002:AWT

[Ell02] Heidi J. C. Ellis. Andragogy in a Web technologies course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):206–210, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Ellis:2006:SGA

[Ell06a] Heidi J. C. Ellis. Self-grading: an approach to supporting selfdirected learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):349, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ellis:2006:UIB

[Ell06b] Heidi J. C. Ellis. Undergraduate involvement in bioinformatics research: lessons from the CONNJUR project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):322, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ellis:2007:CHO

[EMdL^{+}07] Heidi J. C. Ellis, Ralph A. Morelli, Trishan R. de Lanerolle, Jonathan Damon, and Jonathan Raye. Can humanitarian open-source software development draw new students to CS?

SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):551–555, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Elenbogen:2000:YMW

[EMM00] Bruce S. Elenbogen, Bruce R. Maxim, and Chris McDonald. Yet, more Web exercises for learning C++. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):290–294, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Eckerdal:2006:PTC

[EMM^{+}06a] Anna Eckerdal, Robert McCartney, Jan Erik Moström, Mark Ratcliffe, Kate Sanders, and Carol Zander. Putting threshold concepts into context in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):103–107, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Eckerdal:2006:CGS

[EMM^{+}06b] Anna Eckerdal, Robert McCartney, Jan Erik Moström, Mark Ratcliffe, and Carol Zander. Can graduating students design software systems? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):403–407, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Estrin:2005:ITE

[EN05] Thelma Estrin and Frederik Nebeker. An interview with Thelma Estrin. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4):6–13, December 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2005.bib.

English:2001:JPS

[Eng01] John English. JEWL (poster session): GUI programming for complete beginners. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):176, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

English:2002:ECA

[Eng02] John English. Experience with a computer-assisted formal programming examination. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):51– 54, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

English:2004:AAG

[Eng04] John English. Automated assessment of GUI programs using JEWL. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):137–141, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

English:2006:CAA

[Eng06] John English. The checkpoint automated assessment system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):337, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Erkan:2009:ERT

[ENO09] Ali Erkan, Sam Newmark, and Nicolas Ommen. Exposure to research through replication of research: a case in complex networks. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):114–118, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Ensley:2009:HAP

[Ens09] Douglas Ensley. A hands-on approach to proof and abstraction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4):45–47, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Enbody:2009:PCP

[EPM09] Richard J. Enbody, William F. Punch, and Mark McCullen. Python CS1 as preparation for C++ CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):116–120, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Edwards:2008:WCA

[EPQ08] Stephen H. Edwards and Manuel A. Perez-Quinones. Web-CAT: automatically grading programming assignments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):328, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Epstein:2001:PTE

[Eps01] Richard G. Epstein. The play’s the thing: ElderCare VR. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):428, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Epstein:2006:ESC

[Eps06] Richard G. Epstein. An ethics and security course for students in computer science and information technology. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):535–537, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

English:2009:ESP

[ER09] John English and Tammy Rosenthal. Evaluating students’ programs using automated assessment: a case study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):371, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Eremin:2000:SSL

[Ere00] Evgeny Eremin. Software system to learn objects (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):188, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Erlinger:2006:LEC

[Erl06] Michael Erlinger. Lab exercises for computer networking courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):305, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

English:2000:EAA

[ES00] John English and Phil Siviter. Experience with an automatically assessed course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):168–171, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

East:2005:PGE

[ES05a] J. Philip East and J. Ben Schafer. In-person grading: an evaluative experiment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):378–382, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Elsharnouby:2005:USJ

[ES05b] T. Elsharnouby and A. U. Shankar. Using SeSF Java in teaching introductory network courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):67–71, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Compiler/java.bib.

Elsharnouby:2005:UST

[ES05c] Tamer Elsharnouby and A. Udaya Shankar. Using SeSFJava in teaching introductory network courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):67–71, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ebrahimi:2006:ESN

[ES06] Alireza Ebrahimi and Christina Schweikert. Empirical study of novice programming with plans and objects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):52–54, December 2006. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2006.bib.

Ernst:2008:CCP

[ES08] Daniel J. Ernst and Daniel E. Stevenson. Concurrent CS: preparing students for a multicore world. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):230–234, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Eshet:2007:TOS

[Esh07] Yoram Eshet. Teaching online: survival skills for the effective teacher. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):16–18, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Ekaterini:2003:TIS

[ESP03] Georgouli Ekaterini, Bakoyannis Spyros, and Giannakoulias Panagiotis. Teaching IT in secondary education through problem-based learning could be really beneficial. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):243, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Esponda:2008:EVF

[Esp08] Margarita Esponda. Electronic voting on-the-fly with mobile devices. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):93–97, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Ebrahimi:2007:WEA

[ESS^{+}07] Alireza Ebrahimi, Christina Schweikert, S. Sayeed, S. Parham, H. Akibu, A. Saeed, and W. Parris. Website error analysis of colleges and universities on Long Island in New York. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):171–176, June 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Estell:2001:IWB

[Est01] John K. Estell. IPP: a Web-based interactive programming portfolio. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):149–153, March 2001.

CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ernst:2009:HCD

[ESW09] Daniel J. Ernst, Daniel E. Stevenson, and Paul J. Wagner. Hybrid and custom data structures: evolution of the data structures course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):213–217, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Emory:2002:JDL

[ET02] David Emory and Roberto Tamassia. JERPA: a distancelearning environment for introductory Java programming courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):307–311, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Eckerdal:2005:NJP

[ET05] Anna Eckerdal and Michael Thuné. Novice Java programmers’ conceptions of “object” and “class”, and variation theory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):89–93, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Etheredge:2004:CPL

[Eth04] Jim Etheredge. CMeRun: program logic debugging courseware for CS1/CS2 students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):22– 25, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Etlinger:2006:FWT

[Etl06] Henry A. Etlinger. A framework in which to teach (technical) communication to computer science majors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):122–126, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Erkan:2007:DSV

[EVS07] Ali S. Erkan, T. J. VanSlyke, and Timothy M. Scaffidi. Data structure visualization with L^{A}TEX and Prefuse. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):301–305, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Ernst:2009:PSU

[EWH^{+}09] Daniel Ernst, Barry Wittman, Brian Harvey, Tom Murphy, and Michael Wrinn. Preparing students for ubiquitous parallelism. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):136–137, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Exton:2000:ETA

[Ext00] Chris Exton. Elucidate: a tool to aid comprehension of concurrent object oriented execution. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (3):33–36, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fagin:2000:UAB

[Fag00] Barry Fagin. Using Ada-based robotics to teach computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):148–151, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fuller:2005:FSL

[FAL^{+}05] Ursula Fuller, June Amillo, Cary Laxer, W. Michael Mc-Cracken, and Joseph Mertz. Facilitating student learning through study abroad and international projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4):139–151, December 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2005.bib.

Frieze:2002:BEC

[FB02] Carol Frieze and Lenore Blum. Building an effective computer science student organization: the Carnegie Mellon women@SCS action plan. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):74– 78, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/

2002.bib. Women and Computing.

Faulkner:2000:EUE

[FC00] Xristine Faulkner and Fintan Culwin. Enter the usability engineer: integrating HCI and software engineering. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):61–64, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Freitas:2005:PVT

[FCA05] Sérgio Freitas, Maria Beatriz Carmo, and Ana Paula Afonso. A personalized visualization tool for geo-referenced information. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):398, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fincher:2009:II

[FCKU09] Sally Fincher, Stephen Cooper, Michael Kölling, and Ian Utting. ILE-idol. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):4–5, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Teachers:2002:MIN

[fCST02] Israel National Center for Computer Science Teachers. “Machshava”: the Israeli National Center for High School Computer Science Teachers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):234, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Freire:2007:USR

[FdMFPT07] André Pimenta Freire, Renata Pontin de Mattos Fortes, Debora Maria Barroso Paiva, and Marcelo Augusto Santos Turine. Using screen readers to reinforce Web accessibility education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):82–86, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Feinberg:2007:VOO

[Fei07] Dave Feinberg. A visual object-oriented programming environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):140–144, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fekete:2002:TDS

[Fek02] Alan Fekete. Teaching data structures with multiple collection class libraries. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):396–400, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Fekete:2005:TTM

[Fek05] Alan Fekete. Teaching transaction management with SQL examples. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):163–167, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fekete:2008:TSD

[Fek08] Alan D. Fekete. Teaching students to develop thread-safe Java classes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):119–123, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Forbes:2007:WDT

[FG07] Jeffrey Forbes and Daniel D. Garcia. “… but what do the toprated schools do?”: a survey of introductory computer science curricula. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):245–246, March 2007.

CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Feamster:2008:CGR

[FG08] Nick Feamster and Alexander Gray. Can great research be taught?: independent research with cross-disciplinary thinking and broader impact. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):471– 475, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Farrior:2001:UPM

[FH01] Debra Burton Farrior and Daniel E. Hallock. Using project management concepts and Microsoft project software as a tool to develop and manage both on-line and on-campus courses and student team projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33 (1):427, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fossum:2005:NQA

[FH05] T. V. Fossum and S. M. Haller. A new quantitative assessment tool for computer science programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):153–157, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Frieze:2006:CED

[FHBD06] Carol Frieze, Orit Hazzan, Lenore Blum, and M. Bernardine Dias. Culture and environment as determinants of women’s participation in computing: revealing the “women-CS fit”. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):22–26, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Furst:2007:THR

[FIG07] Merrick Furst, Charles Isbell, and Mark Guzdial. ThreadsTM: how to restructure a computer science curriculum for a flat world. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):420–424, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Filho:2001:RES

[Fil01] Wilson P. Paula Filho. Requirements for an educational software development process. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):65– 68, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Finckel:2001:PAS

[Fin01] David Finckel, editor. Proceedings of the 6th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE-01), 25–27 June 2001, Canterbury, UK, volume 33(3) of SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Finkel:2002:PAS

[Fin02] David Finkel, editor. Proceedings of the 7th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE-02), 24–28 June 2002, Aarhus, Denmark, volume 34(3) of SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Finkel:2003:PAC

[Fin03] David Finkel, editor. Proceedings of the 8th Annual Conference on Innovation and Technology in Computer Science Education (ITiCSE-03), 30 June–2 July 2003, Thessaloniki, Greece, volume 35(3) of SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fincher:2004:SSP

[Fin04] Sally Fincher. SIGCSE special projects showcase. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):96–97, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fincher:2005:SSP

[Fin05] Sally Fincher. SIGCSE special projects showcase. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):475–476, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fincher:2009:US

[Fin09] Sally Fincher. Useful sharing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):2, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Fitzgerald:2009:ARN

[Fit09] Sue Fitzgerald. All I really need to know I learned in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):1, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Fetrow:2006:BCC

[FJ06] Jacquelyn S. Fetrow and David J. John. Bioinformatics and computing curriculum: a new model for interdisciplinary courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):185–189, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fuller:2007:DCS

[FJA^{+}07] Ursula Fuller, Colin G. Johnson, Tuukka Ahoniemi, Diana Cukierman, Isidoro Hernán-Losada, Jana Jackova, Essi Lahtinen, Tracy L. Lewis, Donna McGee Thompson, Charles Riedesel, and Errol Thompson. Developing a computer science-specific learning taxonomy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):152–170, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2007.bib.

Fekete:2002:DMS

[FK02] Alan Fekete and Bob Kummerfeld. Design of a major in software development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):73–77, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Fernandes:2004:TSP

[FK04] Eric Fernandes and Amruth N. Kumar. A tutor on scope for the programming languages course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (1):90–93, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fenwick:2005:ICS

[FK05] James B. Fenwick, Jr. and Barry L. Kurtz. Intra-curriculum software engineering education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1): 540–544, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fuller:2008:ASP

[FK08] Ursula Fuller and Bob Keim. Assessing students’ practice of professional values. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):88–92, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Fincher:2001:PAS

[FKCM01] Sally Fincher, Bruce J. Klein, Fintan Culwin, and Michael McCracken, editors. Proceedings of the 6th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education, ITiCSE 2001, Canterbury, UK, June 25– 27, 2001, SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 2001. CODEN SIGSD3. ISBN 1-58113-330-8. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fekete:2000:SRI

[FKKW00] Alan Fekete, Judy Kay, Jeff Kingston, and Kapila Wimalaratne. Supporting reflection in introductory computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):144–148, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ferguson:2004:OOC

[FKMR04] Ernest Ferguson, Clifton Kussmaul, Daniel D. McCracken, and Mary Ann Robbert. Offshore outsourcing: current conditions & diagnosis. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):330–331, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fink:2007:SRS

[FKO07] Shmuel Fink, Gerald Kruse, and Keith Olson. Status report on the SIGCSE committee on models for evaluating faculty scholarship. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):84–85, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Flood:2005:TPC

[FL05] Raymond Flood and Bob Lockhart. Teaching programming collaboratively. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):321–324, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fleury:2000:PJS

[Fle00] Ann E. Fleury. Programming in Java: student-constructed rules. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):197–201, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fleury:2001:ERV

[Fle01] Ann E. Fleury. Encapsulation and reuse as viewed by Java students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):189–193, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fleischman:2006:MIE

[Fle06] William M. Fleischman. Meta-informatics and ethical issues in computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):232–236, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fleck:2007:PFP

[Fle07] Arthur Fleck. Prolog as the first programming language. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):61–64, December 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Fuller:2009:PDA

[FLK^{+}09] Ursula Fuller, Joyce Currie Little, Bob Keim, Charles Riedesel, Diana Fitch, and Su White. Perspectives on developing and assessing professional values in computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4):174–194, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ferren:2001:DDC

[FLL^{+}01] Ann S. Ferren, J. A. N. Lee, Ivan Liss, J. D. Chase, and Robert Phillips. The design and development of the College of Information Technology at Radford University. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):416–417, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ferguson:2006:SLP

[FLLM06] Roger Ferguson, Chang Liu, Mary Last, and Joe Mertz. Service-learning projects: opportunities and challenges. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):127–128, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fisher:2004:CSI

[FLS04a] John Fisher, John Lowther, and Ching-Kuang Shene. Curve and surface interpolation and approximation: knowledge unit and software tool. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):146–150, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fisher:2004:IYK

[FLS04b] John Fisher, John Lowther, and Ching-Kuang Shene. If you know B-splines well, you also know NURBS! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):343–347, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fisher:2002:UCC

[FM02] Allan Fisher and Jane Margolis. Unlocking the clubhouse: the Carnegie Mellon experience. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):79– 83, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/

2002.bib. Women and Computing.

Fagin:2003:MER

[FM03a] Barry Fagin and Laurence Merkle. Measuring the effectiveness of robots in teaching computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):307–311, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fisher:2003:UCW

[FM03b] Allan Fisher and Jane Margolis. Unlocking the clubhouse: women in computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):.23, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Frens:2006:FCF

[FM06] Jeremy D. Frens and Andrew Meneely. Fifteen compilers in fifteen days. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):92–96, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Flatland:2009:UMI

[FM09] Robin Y. Flatland and James R. Matthews. Using modes of inquiry and engaging problems to link computer science and mathematics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):387–391, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Fabrega:2002:VNL

[FMJM02] Lluís Fàbrega, Jordi Massaguer, Teodor Jové, and David Mérida. A virtual network laboratory for learning IP networking. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):161–164, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fitzgerald:2000:SOA

[FMK^{+}00] Sue Fitzgerald, Merry McDonald, Norbert J. Kubilus, Mark Fienup, and Dian Lopez. Student outcomes assessment (panel session): what works and what doesn’t. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):405–406, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fisker:2008:GWS

[FMKQ08] Kasper Fisker, Davin McCall, Michael Kölling, and Bruce Quig. Group work support for the BlueJ IDE. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):163–168, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Fatima:2006:AGP

[FMR^{+}06] Tazeen Fatima, Jonathan Marra, Ronald Realubit, Georgiy Schegolev, and Katherine G. Herbert. Automated gene processing and exon sequence retrieval. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (3):366, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fenwick:2009:ALB

[FNB^{+}09] James B. Fenwick, Jr., Cindy Norris, Frank E. Barry, Josh Rountree, Cole J. Spicer, and Scott D. Cheek. Another look at the behaviors of novice programmers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):296–300, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Fenwick:2002:SEM

[FNW02] James B. Fenwick, Cindy Norris, and James Wilkes. Scientific experimentation via the matching game. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):326–330, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Furcy:2008:SSS

[FNW08] David Furcy, Thomas Naps, and Jason Wentworth. Sorting out sorting: the sequel. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):174– 178, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Fone:2001:UFP

[Fon01] William Fone. Using a familiar package to demonstrate a difficult concept: using an excel spreadsheet model to explain the concepts of neural networks to undergraduates. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):165–168, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fone:2002:IFM

[Fon02] William Fone. Improving feedback from multiple choice tests. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):196, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fone:2003:USA

[Fon03] William Fone. Using semiotics to analyze the efficiency of a model in teaching. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):232, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fone:2004:DMT

[Fon04] William Fone. Design of MCQ test. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (3):250, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fong:2009:RCS

[Fon09] Philip W. L. Fong. Reading a computer science research paper. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):138–140, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ford:2004:LOG

[For04a] Lindsey Ford. A learning object generator for programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):268, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ford:2004:MPT

[For04b] Lindsey Ford. Managing the programming tail. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):249, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fossum:2005:CFC

[Fos05] T. V. Fossum. Classes as first class objects in an environmentpassing interpreter. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):261–265, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fowler:2001:WBC

[Fow01] Aliy Fowler. Web-based CALL using Prolog (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):181, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Falkner:2009:DAP

[FP09] Katrina Falkner and Edward Palmer. Developing authentic problem solving skills in introductory computing classes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):4–8, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Fuller:2006:CPB

[FPA^{+}06] Ursula Fuller, Arnold Pears, June Amillo, Chris Avram, and Linda Mannila. A computing perspective on the Bologna Process. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):115–131, December 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2006.bib.

Fu:2008:AAP

[FPQ^{+}08] Xiang Fu, Boris Peltsverger, Kai Qian, Lixin Tao, and Jigang Liu. APOGEE: automated project grading and instant feedback system for Web based computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):77–81, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Francioni:2002:CIU

[Fra02] Joan M. Francioni. A conference’s impact on undergraduate female students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):66– 69, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/

2002.bib. Women and Computing.

Frens:2004:TTT

[Fre04] Jeremy D. Frens. Taming the tiger: Teaching the next version of Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):151–155, March 2004. CODEN SIGSD3. ISBN 1-58113-798-2. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Compiler/java.bib;   ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/2004.
bib.

Friedman:2004:LPO

[Fri04] William H. Friedman. Learning program organization through COBOL. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(2):77–81, June 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2004.bib.

Frieze:2005:DIC

[Fri05] Carol Frieze. Diversifying the images of computer science: undergraduate women take on the challenge! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):397–400, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Frost:2007:FGC

[Fro07] Daniel Frost. Fourth grade computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):302–306, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Frost:2008:UJL

[Fro08] Daniel Frost. Ucigame, a Java library for games. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):310–314, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Francioni:2002:CSA

[FS02] Joan M. Francioni and Ann C. Smith. Computer science accessibility for students with visual disabilities. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):91–95, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Frandsen:2006:SCM

[FS06] Gudmund Skovbjerg Frandsen and Michael I. Schwartzbach. A singular choice for multiple choice. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (4):34–38, December 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2006.bib.

Fesakis:2009:IFS

[FS09] Georgios Fesakis and Kiriaki Serafeim. Influence of the familiarization with “scratch” on future teachers’ opinions and attitudes about programming and ICT in education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):258–262, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Fitzgerald:2003:FPI

[FSAF03] Sue Fitzgerald, G. Michael Schneider, Robert M. Aiken, and Barry Fagin. The Fulbright Program: international teaching and curriculum development experiences. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):267–268, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fincher:2002:PPT

[FU02] Sally Fincher and Ian Utting. Pedagogical patterns: their place in the genre. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):199–202, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gunawardena:2003:CLO

[GA03] Ananda Gunawardena and Victor Adamchik. A customized learning objects approach to teaching programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):264, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ginat:2006:REH

[GA06] David Ginat and Michal Armoni. Reversing: an essential heuristic in program and proof design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):469–473, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gabbert:2003:GCC

[Gab03] Paula Gabbert. Globalization and the computing curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):61–65, June 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

Gagne:2002:JNB

[Gag02] Greg Gagne. To java.net and beyond: teaching networking concepts using the Java networking API. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):406–410, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Ginat:2004:ILR

[GAGG04] David Ginat, Owen Astrachan, Daniel D. Garcia, and Mark Guzdial. “but it looks right!”: the bugs students don’t see. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):284–285, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ginat:2005:RPE

[GAGR05] David Ginat, Richard Anderson, Daniel D. Garcia, and Richard Rasala. Randomness and probability in the early CS courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):556–557, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Galpin:2002:WCA

[Gal02] Vashti Galpin. Women in computing around the world. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):94–100, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2002.bib. Women and

Computing.

Granger:2006:REU

[GAL^{+}06] Mary J. Granger, Guy-Alain Amoussou, Miguel A. Labrador, Sue Perry, and Kelly M. Van Busum. Research experience for undergraduates: successes and challenges. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):558–559, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gaona:2000:RDC

[Gao00] Amparo López Gaona. The relevance of design in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(2):53–55, June 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Garner:2004:UCR

[Gar04] Stuart Garner. The use of a code restructuring tool in the learning of programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):277, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gunawardena:2008:MAR

[GBO08] Ananda Gunawardena, John Barr, and Andrew Owens. A method for analyzing reading comprehension in computer science courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):348, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Gurer:2002:AWL

[GC02] Denise W. Gürer and Tracy Camp. An ACM-W literature review on women in computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (2):121–127, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/

DBLP/2002.bib. Women and Computing.

Goold:2005:TEV

[GC05] Annegret Goold and Jo Coldwell. Teaching ethics in a virtual classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):232–236, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Garcia:2009:RPB

[GCD^{+}09] Daniel D. Garcia, Robb Cutler, Zachary Dodds, Eric Roberts, and Alison Young. Rediscovering the passion, beauty, joy, and awe: making computing fun again, continued. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):65–66, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Gehringer:2008:WCL

[GCDJ08] Edward F. Gehringer, Lillian Cassel, Katherine Deibel, and William Joel. Wikis: collaborative learning for CS education.

SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):379–380, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Goldweber:2001:URU

[GCF^{+}01] Michael Goldweber, Clare Congdon, Barry Fagin, Deborah Hwang, and Frank Klassner. The use of robots in the undergraduate curriculum: experience reports. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):404–405, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goldweber:2004:RBCa

[GCF04] Michael Goldweber, Martyn Clark, and Sally Fincher. The relationship between CS education research and the SIGCSE community. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):147–148, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goldweber:2004:RBCb

[GCFP04] Michael Goldweber, Martyn Clark, Sally Fincher, and Arnold Pears. The relationship between CS education research and the SIGCSE community. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):228– 229, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gallego-Carrillo:2005:SVT

[GCGBUFVI05] Micael Gallego-Carrillo, Francisco Gortázar-Bellas, Jaime Urquiza-Fuentes, and J. Ángel Velázquez-Iturbide. SOTA: a visualization tool for symbol tables. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37 (3):385, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gehringer:2005:UPR

[GCPQA05] Edward F. Gehringer, Donald D. Chinn, Manuel A. Pérez-Quiñones, and Mark A. Ardis. Using peer review in teaching computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):321–322, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Georgiopoulos:2003:CML

[GCW^{+}03] M. Georgiopoulos, J. Castro, A. Wu, R. DeMara, E. Gelenbe, A. Gonzalez, M. Kysilka, and M. Mollaghasemi. CRCD in machine learning at the University of Central Florida: preliminary experiences. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):249, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goulding:2005:IRC

[GD05] Tom Goulding and Rita DiTrolio. Incorporating realistic constraints into a student team software project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4):54–58, December 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2005.bib.

Goulding:2007:CGD

[GD07] Tom Goulding and Rita DiTrolio. Complex game development by freshman computer science majors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):92–99, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2007.bib.

Goldweber:2008:VEE

[GD08] Michael Goldweber and Renzo Davoli. VDE: an emulation environment for supporting computer networking courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):138–142, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Garvin-Doxas:2003:CLE

[GDB03] Kathy Garvin-Doxas and Lecia J. Barker. Creating learning environments that support interaction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):276, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gehringer:2006:CLB

[GDHW06] Edward F. Gehringer, Katherine Deibel, John Hamer, and Keith J. Whittington. cooperative learning: beyond pair programming and team projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (1):458–459, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goldweber:2005:KPM

[GDM05] Michael Goldweber, Renzo Davoli, and Mauro Morsiani. The Kaya OS project and the μMPS hardware emulator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):49–53, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gal-Ezer:2006:OUI

[GE06] Judith Gal-Ezer. The Open University of Israel: a distance education institution. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):16–17, December 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2006.bib.

Gal-Ezer:2007:RDL

[GE07] Judith Gal-Ezer. Reflections on distance learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):21–22, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2007.bib.

Gal-Ezer:2008:DEO

[GE08a] Judith Gal-Ezer. Distance education: opening books. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):21, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gal-Ezer:2008:OCN

[GE08b] Judith Gal-Ezer. Online courses: North Carolina business and IT courses: a case study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2): 18–19, June 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2008.bib.

Gal-Ezer:2009:DED

[GE09] Judith Gal-Ezer. Distance education: different models of course development: from traditional distance education to technology-based education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):77– 78, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gehringer:2001:EPR

[Geh01] Edward F. Gehringer. Electronic peer review and peer grading in computer-science courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1): 139–143, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gal-Ezer:2009:PHS

[GEHR09] Judith Gal-Ezer, Orit Hazzan, and Noa Ragonis. Preparation of high school computer science teachers: the Israeli perspective. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):269–270, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Gelderblom:2000:OCS

[Gel00] J. Helene Gelderblom. OOPtutor: a CBL system for introductory object-oriented programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(2):35–38, June 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gal-Ezer:2004:IBP

[GELS04a] Judith Gal-Ezer, Dvir Lanzberg, and Daphna Shahak. Interesting basic problems for CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (3):275, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gal-Ezer:2004:IU

[GELS04b] Judith Gal-Ezer, Dvir Lanzberg, and Daphna Shahak. Introducing undecidability. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):276, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

George:2000:EVR

[Geo00] Carlisle E. George. EROSI — visualising recursion and discovering new errors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):305–309, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

George:2002:UVA

[Geo02] Carlisle E. George. Using visualization to aid program construction tasks. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):191–195, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Gestwicki:2007:CGM

[Ges07] Paul V. Gestwicki. Computer games as motivation for design patterns. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):233–237, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gal-Ezer:2009:CSI

[GESZ09] Judith Gal-Ezer, Daphna Shahak, and Ela Zur. Computer science issues in high school: gender and more. … SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):278–282, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Gal-Ezer:2005:CTP

[GET05] Judith Gal-Ezer and Mark Trakhtenbrot. Challenges in teaching the pumping lemma in automata theory course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):369, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gal-Ezer:2006:ACR

[GET06] Judith Gal-Ezer and Mark Trakhtenbrot. Algebraic characterization of regular languages: how to cope with all these equivalences? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):325, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gal-Ezer:2003:TAE

[GEVZ03] Judith Gal-Ezer, Tamar Vilner, and Ela Zur. Teaching algorithm efficiency in a CS1 course SBTa different approach. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):256, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gal-Ezer:2009:PSC

[GEVZ09a] Judith Gal-Ezer, Tamar Vilner, and Ela Zur. Has the paradigm shift in CS1 a harmful effect on data structures courses: a case study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):126–130, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Gal-Ezer:2009:PYP

[GEVZ09b] Judith Gal-Ezer, Tamar Vilner, and Ela Zur. The professor on your PC: a virtual CS1 course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (3):191–195, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Guzdial:2005:DPN

[GF05] Mark Guzdial and Andrea Forte. Design process for a nonmajors computing course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):361– 365, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gharibyan:2006:GGC

[GG06a] Hasmik Gharibyan and Stephan Gunsaulus. Gender gap in computer science does not exist in one former Soviet republic: results of a study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):222–226, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Guerreiro:2006:CAP

[GG06b] Pedro Guerreiro and Katerina Georgouli. Combating anonymousness in populous CS1 and CS2 courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):8–12, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ginat:2001:CIA

[GGAB01] David Ginat, Dan Garcia, Owen Astrachan, and Joseph Bergin. Colorful illustrations of algorithmic design techniques and problem solving. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):425–426, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gayo:2003:GLM

[GGÁC03] J. E. Labra Gayo, J. M. Morales Gil, A. M. Fernández Álvarez, and H. Sagastegui Chigne. A generic e-learning multiparadigm programming language system: IDEFIX project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):391–395, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Garcia:2003:EYA

[GGH03] Daniel D. Garcia, David Ginat, and Peter Henderson. Everything you always wanted to know about game theory: but were afraid to ask. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):96–97, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goldman:2008:IID

[GGH^{+}08] Ken Goldman, Paul Gross, Cinda Heeren, Geoffrey Herman, Lisa Kaczmarczyk, Michael C. Loui, and Craig Zilles. Identifying important and difficult concepts in introductory computing courses using a Delphi process. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):256–260, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Grodzinsky:2004:RCT

[GGKT04] Frances Grodzinsky, Ed Gehringer, Laurie S. King, and Herman Tavani. Responding to the challenges of teaching computer ethics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):280–281, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gray:2003:ECC

[GGR03] Kathleen Arnold Gray, Mark Guzdial, and Spencer Rugaber. Extending CRC cards into a complete design process. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):226, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gaspar:2006:SUO

[GGS06a] Alessio Gaspar, Clark Godwin, and Joe Stanaback. SOFT-ICE undergraduate operating systems laboratories. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):354, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gorgone:2006:MMC

[GGS^{+}06b] John T. Gorgone, Paul Gray, Edward A. Stohr, Joseph S. Valacich, and Rolf T. Wigand. MSIS 2006: model curriculum and guidelines for graduate degree programs in information systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):121–196, June 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2006.bib.

Gegg-Harrison:2001:AEN

[GH01] Timothy S. Gegg-Harrison. Ancient Egyptian numbers: a CS-complete example. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):268–272, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gray:2006:IAM

[GH06] Geoffrey R. Gray and Colin A. Higgins. An introspective approach to marking graphical user interfaces. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):43–47, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ghafarian:2005:RET

[Gha05a] Ahmad K. Ghafarian. Reverse engineering technique to enhance software engineering education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):354, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gharibyan:2005:ASK

[Gha05b] Hasmik Gharibyan. Assessing students’ knowledge: oral exams vs. written tests. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):143–147, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ghafarian:2007:IPU

[Gha07] Ahmad Ghafarian. Ideas for projects in undergraduate information assurance and security courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):322, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Gegg-Harrison:2003:SPCa

[GHBG^{+}03a] Timothy S. Gegg-Harrison, Gary R. Bunce, Rebecca D. Ganetzky, Christina M. Olson, and Joshua D. Wilson. Studying program correctness by constructing contracts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):129–133, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gegg-Harrison:2003:SPCb

[GHBG^{+}03b] Timothy S. Gegg-Harrison, Gary R. Bunce, Rebecca D. Ganetzky, Christina M. Olson, and Joshua D. Wilson. Studying program correctness in ProVIDE. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):262, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gray:2008:CBT

[GHG^{+}08] John Gray, Gill Harrison, Andrea Gorra, Jakki Sheridan-Ross, and Janet Finlay. A computer-based test to raise awareness of disability issues. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3): 347, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Groth:2008:ICS

[GHLL08] Dennis P. Groth, Helen H. Hu, Betty Lauer, and Hwajung Lee. Improving computer science diversity through summer camps. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):180–181, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Gibbs:2000:ECL

[Gib00] David C. Gibbs. The effect of a constructivist learning environment for field-dependent/independent students on achievement in introductory computer programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):207–211, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gibson:2002:SRB

[Gib02] Robyn Gibson. The student record book: showing the value of documentation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):197, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gibson:2009:SRP

[Gib09] J. Paul Gibson. Software reuse and plagiarism: a code of practice. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):55–59, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Giguette:2002:CAT

[Gig02] Ray Giguette. The Crawfish and the Aztec treasure maze: adventures in data structures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4): 89–93, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2002.bib.

Giguette:2003:PGG

[Gig03] Ray Giguette. Pre-games: games designed to introduce CS1 and CS2 programming assignments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35 (1):288–292, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ginat:2000:CEE

[Gin00a] David Ginat. Colorful examples for elaborating exploration of regularities in high-school CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (3):81–84, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ginat:2000:PC

[Gin00b] David Ginat. Placement calculations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(4):20–21, December 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ginat:2001:CP

[Gin01a] David Ginat. Chain of permutations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33 (4):20–21, December 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2001.bib.

Ginat:2001:CC

[Gin01b] David Ginat. Color conversion. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(2): 20–21, June 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2001.bib.

Ginat:2001:MAU

[Gin01c] David Ginat. Metacognitive awareness utilized for learning control elements in algorithmic problem solving. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):81–84, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ginat:2001:MIA

[Gin01d] David Ginat. Misleading intuition in algorithmic problem solving. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):21–25, March 2001. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ginat:2001:STR

[Gin01e] David Ginat. Starting top-down, refining bottom-up, sharpening by zoom-in. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(4):28–31, December 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2001.bib.

Ginat:2002:AIP

[Gin02a] David Ginat. Aha! an illuminating perspective. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):1–2, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Ginat:2002:DG

[Gin02b] David Ginat. Divisor games. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4): 28–29, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2002.bib.

Ginat:2002:VPP

[Gin02c] David Ginat. On varying perspectives of problem decomposition. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):331–335, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Ginat:2003:BR

[Gin03a] David Ginat. Board reconstruction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35 (4):25–26, December 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2003.bib.

Ginat:2003:GTL

[Gin03b] David Ginat. The greedy trap and learning from mistakes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):11–15, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ginat:2003:NPS

[Gin03c] David Ginat. The novice programmers’ syndrome of designby-keyword. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):154–157, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ginat:2003:SD

[Gin03d] David Ginat. Sorting and disorders. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35 (2):29–30, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2003.bib.

Ginat:2004:APC

[Gin04a] David Ginat. Algorithmic patterns and the case of the sliding delta. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(2):29–33, June 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2004.bib.

Ginat:2004:DSC

[Gin04b] David Ginat. Do senior CS students capitalize on recursion? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):82–86, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ginat:2004:EIA

[Gin04c] David Ginat. Embedding instructive assertions in program design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):62–66, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ginat:2004:H

[Gin04d] David Ginat. Helipad. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(4):25–26, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2004.bib.

Ginat:2004:VET

[Gin04e] David Ginat. On the verge of an empty tank. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(2):24–26, June 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.

math.utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2004.bib.

Ginat:2005:QC

[Gin05a] David Ginat. Queens on a chessboard. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4):26–27, December 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2005.bib.

Ginat:2005:TE

[Gin05b] David Ginat. Tuple enumeration. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37 (2):22–23, June 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2005.bib.

Ginat:2006:GC

[Gin06a] David Ginat. Graph cycles. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):20– 22, June 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2006.bib.

Ginat:2006:HS

[Gin06b] David Ginat. Hidden squares. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4): 22–23, December 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2006.bib.

Ginat:2007:AA

[Gin07a] David Ginat. Alternating accumulation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):28–29, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2007.bib.

Ginat:2007:DA

[Gin07b] David Ginat. Domino arrangements. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (2):24–25, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2007.bib.

Ginat:2007:EHR

[Gin07c] David Ginat. Elaborating heuristic reasoning and rigor with mathematical games. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):32–36, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2007.bib.

Ginat:2007:HDF

[Gin07d] David Ginat. Hasty design, futile patching and the elaboration of rigor. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):161–165, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Ginat:2008:CCK

[Gin08a] David Ginat. Colorful challenges: Kangaroo hops. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):28–29, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ginat:2008:LWC

[Gin08b] David Ginat. Learning from wrong and creative algorithm design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):26–30, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Ginat:2008:SBF

[Gin08c] David Ginat. Separating between foes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):24–25, June 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2008.bib.

Ginat:2009:CCC

[Gin09a] David Ginat. Colorful challenges: chips game. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):82–83, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ginat:2009:CCL

[Gin09b] David Ginat. Colorful challenges: The lengthy 0–1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4):22–23, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ginat:2009:IPC

[Gin09c] David Ginat. Interleaved pattern composition and scaffolded learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):109–113, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Ginat:2009:NMD

[Gin09d] David Ginat. On the non-modular design of on-the-fly computations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4):35–39, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Greening:2002:URE

[GK02] Tony Greening and Judy Kay. Undergraduate research experience in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):151–155, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gayler:2007:UTW

[GKHPQ07] Dick Gayler, David Klappholz, Valerie J. Harvey, and Manuel A. Pérez-Quiñones. UML tools: what is their role in undergraduate computer science courses? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):129–130, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Grissom:2003:PST

[GKJD03] Scott Grissom, Deborah Knox, Dan Joyce, and Wanda Dann, editors. Proceedings of the 34th SIGCSE Technical Symposium on Computer Science Education, 2003, Reno, Nevada, USA, February 19–23, 2003, SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 2003. CODEN SIGSD3. ISBN 1-58113-648-X. ISSN 0097-8418 (print), 2331-3927 (electronic).

Grinder:2002:LLT

[GKL^{+}02] Michael T. Grinder, Seong B. Kim, Teresa L. Lutey, Rockford J. Ross, and Kathleen F. Walsh. Loving to learn theory: active learning modules for the theory of computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):371–375, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Grisham:2006:DEE

[GKP06] Paul S. Grisham, Herb Krasner, and Dewayne E. Perry. Data Engineering education with real-world projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):64–68, June 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.

math.utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2006.bib.

Graham:2003:CGR

[GL03a] Sandy Graham and Celine Latulipe. CS girls rock: sparking interest in computer science and debunking the stereotypes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):322–326, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Granger:2003:CDP

[GL03b] Mary J. Granger and Joyce Currie Little. Classroom discussions: policies and responsibilities of Internet Service Providers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):99–103, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goldwasser:2005:PSU

[GL05] Michael H. Goldwasser and David Letscher. Providing students universal access to a centralized, graphical computing environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):79–83, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goldwasser:2007:INP

[GL07a] Michael H. Goldwasser and David Letscher. Introducing network programming into a CS1 course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):19–22, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Goldwasser:2007:TOO

[GL07b] Michael H. Goldwasser and David Letscher. Teaching objectoriented programming in Python. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (3):365–366, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Goldwasser:2008:PGP

[GL08a] Michael H. Goldwasser and David Letscher. A Python graphics package for the first day and beyond. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):326, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Goldwasser:2008:TOO

[GL08b] Michael H. Goldwasser and David Letscher. Teaching an object-oriented CS1 -: with Python. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (3):42–46, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Goldwasser:2009:GPF

[GL09] Michael H. Goldwasser and David Letscher. A graphics package for the first day and beyond. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (1):206–210, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Gaspar:2008:RVC

[GLA^{+}08] Alessio Gaspar, Sarah Langevin, William Armitage, R. Sekar, and T. Daniels. The role of virtualization in computing education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):131–132, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Gorra:2007:AUS

[GLC07] Andrea Gorra, Sanela Lazarevski, and Jackie Campbell. An account of the use of synoptic assessment for students in the area of databases at level 2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3): 334, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Gottschalk:2005:CFP

[GLD^{+}05] Larry Gottschalk, Jigang Liu, Brahma Dathan, Sue Fitzgerald, and Michael Stein. Computer forensics programs in higher education: a preliminary study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):147–151, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gorgone:2001:SIS

[GLF01] John T. Gorgone, Doris K. Lidtke, and David Feinstein. Status of information systems accreditation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):421–422, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gries:2001:ACG

[GLR^{+}01] David Gries, Kathleen Larson, Susan H. Rodger, Mark A. Weiss, and Ursula Wolz. AP CS goes OO. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):423–24, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gabbert:2002:SCW

[GM02] Paula Gabbert and Paige H. Meeker. Support communities for women in computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):62– 65, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/

2002.bib. Women and Computing.

Giangrandi:2006:NMV

[GM06a] Paolo Giangrandi and Claudio Mirolo. “Numeri e Macchine”: a virtual museum to learn the history of computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):78–82, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gonzalez-Morris:2006:FTB

[GM06b] Germán González-Morris. Further thoughts on backtracking and bounding. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):46–47, December 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2006.bib.

Govender:2007:CRR

[GM07] Desmond Wesley Govender and Manoj Maharaj. Challenges with respect to the e-readiness of secondary school teachers in KwaZulu–Natal, South Africa. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (3):191–195, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Giangrandi:2008:EGB

[GM08] Paolo Giangrandi and Claudio Mirolo. Enhancing the general background of CS students through a computing history course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):301–305, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Gehringer:2009:SGA

[GM09] Edward F. Gehringer and Carolyn S. Miller. Studentgenerated active-learning exercises. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (1):81–85, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Garcia-Mateos:2009:CAD

[GMFA09] Ginés Gárcia-Mateos and José Luis Fernández-Alemán. A course on algorithms and data structures using on-line judging. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):45–49, September 2009. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Gomez-Martin:2006:FAD

[GMGM06] Pedro Pablo Gómez-Martín and Marco Antonio Gómez-Martín. Fast application development to demonstrate computer graphics concepts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):250– 254, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gries:2001:HMT

[GMHS01] David Gries, Bill Marion, Peter Henderson, and Diane Schwartz. How mathematical thinking enhances computer science problem solving. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):390– 391, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gomez-Martin:2009:TDP

[GMJDA09] Marco Antonio Gómez-Martín, Guillermo Jiménez-Díaz, and Javier Arroyo. Teaching design patterns using a family of games. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):268–272, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Garcia:2002:LHD

[GMN02] Mario A. Garcia and Holly Patterson Mc-Neill. Learning how to develop software using the toy LEGO Mindstorms.

SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):239, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gunion:2009:CRA

[GMS09] Katherine Gunion, Todd Milford, and Ulrike Stege. Curing recursion aversion. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):124–128, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Grissom:2001:PTT

[GNPL01] Scott Grissom, Tom Naps, Nick Parlante, and Pamela Lawhead. Practical teaching tips from around the world. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):419, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gabbert:2007:AWN

[GOB^{+}07] Paula Gabbert, Barbara Boucher Owens, Bettina Bair, Chris Stephenson, Joanne McGrath Cohoon, and Gloria Childress Townsend. ACM-W’s new programs for recruiting and retaining women in computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1): 247–248, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Godejord:2006:ETM

[God06] Per Arne Godejord. Exploring teacher methodology: using the example of sexual abuse of children as a way of teaching social informatics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):340, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goelman:2003:SES

[Goe03] Don Goelman. Student empowerment in a survey course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):270, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goelman:2008:DNM

[Goe08] Don Goelman. Databases, non-majors and collaborative learning: a ternary relationships. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (3):27–31, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Garcia-Osorio:2008:TTL

[GOGPGP08] César García-Osorio, Carlos Gómez-Palacios, and Nicolás García-Pedrajas. A tool for teaching LL and LR parsing algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):317, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Golub:2001:PBD

[Gol01] Evan Golub. PC-based development environments and a Unix-centric curriculum: some practical issues. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(2):51–54, June 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.

math.utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2001.bib.

Goldwasser:2002:GIS

[Gol02] Michael H. Goldwasser. A gimmick to integrate software testing throughout the curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1): 271–275, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Goldman:2004:CFI

[Gol04a] Kenneth J. Goldman. A concepts-first introduction to computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):432–436, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Golub:2004:HST

[Gol04b] Evan Golub. Handwritten slides on a tabletPC in a discrete mathematics course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):51–55, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goluboff:2008:DTD

[Gol08] Barry Goluboff. A denotational tracing domain for C++ programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):56–61, June 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2008.bib.

Garcia-Osorio:2008:TPA

[GOMSJVGP08] César García-Osorio, Iñigo Mediavilla-Sáiz, Javier Jimeno-Visitación, and Nicolás García-Pedrajas. Teaching pushdown automata and Turing machines. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):316, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Gonzalez:2006:SAA

[Gon06] Graciela Gonzalez. A systematic approach to active and cooperative learning in CS1 and its effects on CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):133–137, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goold:2003:PPP

[Goo03] Annegret Goold. Providing process for projects in capstone courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):26–29, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goode:2008:IDK

[Goo08] Joanna Goode. Increasing diversity in K-12 computer science: strategies from the field. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):362– 366, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Gorgone:2000:CAV

[Gor00a] John T. Gorgone. CSAB authorizes visits to test IS/IT proposed accreditation criteria. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(4): 13–14, December 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gorgone:2000:NGC

[Gor00b] John T. Gorgone. A new IS graduate curriculum model — after eighteen years. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(2):13–14, June 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gorgone:2001:ICC

[Gor01a] John T. Gorgone. The IS2001 curriculum in CC2001 computing compendium. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(4):11–12, December 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2001.bib.

Gorgone:2001:NIC

[Gor01b] John T. Gorgone. National IT curricula: issues, definition, trends, and standards. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(2):11– 12, June 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2001.bib.

Gorgone:2002:ITA

[Gor02] John T. Gorgone. Information technology: an adolescent in the arena. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):15–16, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2002.bib.

Gorgone:2003:AGA

[Gor03a] John T. Gorgone. ABET’s general accreditation criteria to apply to all computing programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35 (4):14–16, December 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2003.bib.

Gorgone:2003:ITA

[Gor03b] John T. Gorgone. Information technology accreditation criteria. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):13–15, June 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

Gorgone:2004:DIS

[Gor04a] John T. Gorgone. Draft information systems accreditation criteria for 2006. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(2):15–17, June 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2004.bib.

Gorgone:2004:ISO

[Gor04b] John T. Gorgone. Information systems and the overview report for computing curricula 2004. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (4):15–16, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2004.bib.

Gorgone:2005:ISP

[Gor05] John T. Gorgone. Information systems: program location, business component, and accreditation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4):16–18, December 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2005.bib.

Gorgone:2006:GCN

[Gor06a] John T. Gorgone. Graduate curriculum now available. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):10–11, June 2006. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2006.bib.

Gorgone:2006:MLA

[Gor06b] John T. Gorgone. Masters level accreditation for information systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):10–11, December 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2006.bib.

Gotterbarn:2000:ELS

[Got00a] Don Gotterbarn. The education and licensing of software professionals: the myth of “a perfected science” considered harmful. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(2):8–9, June 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gotterbarn:2000:BUW

[Got00b] Don Gotterbarn. On being a UCITArian: winning the race to the bottom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(4):6–7, December 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gotterbarn:2001:ASW

[Got01] Don Gotterbarn. Antipodal shock: watch your language! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(4):7–8, December 2001. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2001.bib.

Gotterbarn:2002:BBE

[Got02] Don Gotterbarn. Black and blue epiphany: the missing elements of professionalism. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):8–9, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2002.bib.

Gotterbarn:2003:ICO

[Got03] Don Gotterbarn. Injectable computers: once more into the breach! the life cycle of computer ethics awareness. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):10–11, December 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

Gotterbarn:2004:UAM

[Got04] Don Gotterbarn. UML and agile methods: in support of irresponsible development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(2):11– 13, June 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2004.bib.

Gotterbarn:2005:ASP

[Got05] Don Gotterbarn. Antipodal shock: professionalism turned upside down. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4):14–15, December 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2005.bib.

Gotterbarn:2006:VFP

[Got06a] Don Gotterbarn. E-voting: a failure of professionalism? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):7–8, December 2006. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2006.bib.

Gotterbarn:2006:SFP

[Got06b] Donald Gotterbarn. The shinning face of plagiarism: the graying of an issue. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):7–8, June 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2006.bib.

Gotterbarn:2007:DSC

[Got07] Donald Gotterbarn. Don’t soil the commons: turning the blogosphere into the clogosphere. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4): 12–13, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2007.bib.

Gotterbarn:2008:TPR

[Got08] Don Gotterbarn. Thinking professionally: a real problem with video games; not murder, not torture. … SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):9–10, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gotterbarn:2009:TPP

[Got09a] Don Gotterbarn. Thinking professionally: professional computer ethics: “i didn’t do it” is not good enough. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):65–66, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gotterbarn:2009:TPW

[Got09b] Don Gotterbarn. Thinking professionally: When soon after is way too late: the deception of ’opt-out’ systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4):6–8, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gousie:2006:RWP

[Gou06] Michael B. Gousie. A robust Web programming and graphics course for non-majors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):72– 76, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goulding:2008:CGD

[Gou08] Tom Goulding. Complex game development throughout the college curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):68–71, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goulding:2009:ESA

[Gou09] Tom Goulding. An encryption system in assembly language: a game-like project for novice programmers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4):40–44, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Govender:2007:USP

[Gov07] Irene Govender. Understanding students’ performance in programming assessment questions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (3):326, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Gerhardt-Powals:2000:GLN

[GP00] Jill Gerhardt-Powals. Have a great lab without needing roller skates. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):182, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gallagher:2002:WAR

[GP02] John C. Gallagher and Steven Perretta. WWW autonomous robotics: enabling wide area access to a computer engineering practicum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):13–17, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Garzas:2007:ITO

[GP07] Javier Garzás and Mario Piattini. Improving the teaching of object-oriented design knowledge. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (4):108–112, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2007.bib.

Golshani:2001:CCI

[GPF^{+}01] F. Golshani, S. Panchanathan, O. Friesen, Y. C. Park, and J. J. Song. A comprehensive curriculum for IT education and workforce development: an engineering approach. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):238–242, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goold:2005:VST

[GPL^{+}05] Annegret Goold, Arnold Pears, Mary Z. Last, Chris Beaumont, and Swee Cheng Chew. Virtual student teams: opportunities and challenges. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):340– 341, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gerhardt-Powals:2000:DEL

[GPP00] Jill Gerhardt-Powals and Matthew H. Powals. Distance education: law attempts to catch up with technology (battle between copyright owners and academics). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):335–342, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goold:2000:FAP

[GR00a] Annagret Goold and Russell Rimmer. Factors affecting performance in first-year computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (2):39–43, June 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goold:2000:UBC

[GR00b] Annegret Goold and Russell Rimmer. Undergraduates in business computing and computer science (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):188, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Grable:2003:ICC

[Gra03] Ross Grable. Information characteristics for the curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):74–77, June 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

Granger:2005:LTC

[Gra05] Mary J. Granger. Learning technical concepts with collaboration and communication skills. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37 (3):391, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Graham:2007:IPP

[Gra07] John R. Graham. Integrating parallel programming techniques into traditional computer science curricula. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):75–78, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2007.bib.

Greening:2000:PSR

[Gre00a] Tony Greening. Pedagogically sound responses to economic rationalism. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):149–156, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Greening:2000:SSF

[Gre00b] Tony Greening. Students seen flocking in programming assignments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):93–96, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Grissom:2000:PFI

[Gri00] Scott Grissom. A pedagogical framework for introducing Java I/O in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(4):57–59, December 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Grissom:2001:RCP

[Gri01] Scott Grissom. Reality check (poster session): an informal feedback tool. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):195, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gries:2002:PCE

[Gri02a] David Gries. Problems with CS education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):135, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gries:2002:WPM

[Gri02b] David Gries. Where is programming methodology these days? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):5–7, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2002.bib.

Grinder:2002:AAC

[Gri02c] Michael T. Grinder. Animating automata: a cross-platform program for teaching finite automata. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):63–67, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Grinder:2003:PEE

[Gri03] Michael T. Grinder. A preliminary empirical evaluation of the effectiveness of a finite state automaton animator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):157–161, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gries:2008:PAT

[Gri08] David Gries. A principled approach to teaching OO first. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):31–35, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Gregorio-Rodriguez:2001:EAW

[GRLDPG^{+}01] C. Gregorio-Rodríguez, L. Llana-Díaz, P. Palao-Gostanza, C. Pareja-Flores, R. Martínez-Unanue, and J. Á. Velázquez-Iturbide. EXercita: automatic Web publishing of programming exercises. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):161–164, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ge:2000:CCS

[GS00] Yuzhen Ge and Jiangeng Sun. E-commerce and computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):250–255, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Guzdial:2003:CSM

[GS03] Mark Guzdial and Elliot Soloway. Computer science is more important than calculus: the challenge of living up to our potential. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):5–8, June 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

Geller:2005:TME

[GS05a] Viktor Geller and Christelle Scharff. Traditional and more “exotic” .NET languages: VB .NET, J#, C# and SML .NET. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):406, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gould:2005:UIC

[GS05b] John Gould and Matthew Schaefer. User interface considerations for older users. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):408, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gerdt:2006:WBS

[GS06a] Petri Gerdt and Jorma Sajaniemi. A Web-based service for the automatic detection of roles of variables. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):178–182, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Griswold:2006:UPF

[GS06b] William G. Griswold and Beth Simon. Ubiquitous presenter: fast, scalable active learning for the whole classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):358, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gutierrez:2009:CSE

[GS09] Juan M. Gutiérrez and Ian D. Sanders. Computer science education in Peru: a new kind of monster? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):86–89, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gotschi:2003:MMR

[GSG03] Tina Götschi, Ian Sanders, and Vashti Galpin. Mental models of recursion. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):346–350, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Greening:2000:CDD

[GSS00] Tony Greening, Glenn Stevens, and David Stratton. A case for data-driven testing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):172– 175, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gomes:2005:SBB

[GSVG05] Raquel Gomes, Ricardo Seabra, João Ventura, and Nuno Guedes. SoundSpaces: behavior based control system for multimedia systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):407, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gotel:2008:TSQ

[GSW08] Olly Gotel, Christelle Scharff, and Andrew Wildenberg. Teaching software quality assurance by encouraging student contributions to an open source Web-based system for the assessment of programming assignments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):214–218, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Galpin:2007:LSP

[GSyC07] Vashti C. Galpin, Ian D. Sanders, and Pei yu Chen. Learning styles and personality types of computer science students at a South African university. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3): 201–205, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Gabbert:2000:EUP

[GT00] Paula Gabbert and Kevin Treu. Experiments with the use of popular press in the computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):124–128, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Goodrich:2001:TIA

[GT01] Michael T. Goodrich and Roberto Tamassia. Teaching Internet algorithmics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):129–133, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Guimaraes:2005:CDA

[Gui05] Mario Guimaraes. Constructing database applications for PDA using the .NET framework. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37 (3):417, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gupta:2001:ITL

[Gup01] Gireesh K. Gupta. Information technology and liberal arts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(2):55–57, June 2001. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2001.bib.

Gupta:2006:CLE

[Gup06] Gireesh K. Gupta. Computer literacy: essential in today’s computer-centric world. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):115– 119, June 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2006.bib.

Gurer:2002:PWC

[Gür02a] Denise W. Gürer. Pioneering women in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):175–180, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2002.bib. Women and

Computing.

Gurer:2002:WCH

[Gür02b] Denise W. Gürer. Women in computing history. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):116–120, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2002.bib. Women and

Computing.

Guzdial:2001:UCM

[Guz01a] Mark Guzdial. Use of collaborative multimedia in computer science classes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):17–20, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Guzdial:2001:UST

[Guz01b] Mark Guzdial. Using squeak for teaching user interface software. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):219–223, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Guzdial:2003:MCC

[Guz03] Mark Guzdial. A media computation course for non-majors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):104–108, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

George:2006:DSC

[GV06] Binto George and Anna Valeva. A database security course on a shoestring. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):7–11, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Garera:2009:CTG

[GV09] Sujata Garera and Jorge Vasconcelos. Challenges in teaching a graduate course in applied cryptography. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):103–107, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gersting:2002:PST

[GWG02] Judith L. Gersting, Henry MacKay Walker, and Scott Grissom, editors. The proceedings of the thirty-third SIGCSE Technical Symposium on Computer Science Education: SIGCSE 2002, Northern Kentucky, February 27–March 3, 2002: theme, At the center of change, SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 2002. CO-DEN SIGSD3. ISBN 1-58113-473-8. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gersting:2000:EEIa

[GY00a] Judith L. Gersting and Frank H. Young. Experiences with ethical issues. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(2):20–21, June 2000.

CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gersting:2000:EEIb

[GY00b] Judith L. Gersting and Frank H. Young. Experiences with ethical issues: part 2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(4):18–19, December 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gersting:2001:ITE

[GY01a] Judith L. Gersting and Frank H. Young. Improving the team experience. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(4):18–19, December 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2001.bib.

Gersting:2001:SWW

[GY01b] Judith L. Gersting and Frank H. Young. Shall we write? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(2):18–19, June 2001. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2001.bib.

Gersting:2002:JH

[GY02] Judith L. Gersting and Frank H. Young. Job-hunting 101. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):20–21, December 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2002.bib.

Gersting:2003:TOS

[GY03] Judith L. Gersting and Frank H. Young. Treating our students as adults. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):21–22, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2003.bib.

Haberman:2002:CBC

[HA02] Bruria Haberman and Haim Averbuch. The case of base cases: why are they so difficult to recognize? student difficulties with recursion. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):84–88, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hill:2004:DTE

[HA04] John M. D. Hill and Kenneth L. Alford. A distributed task environment for teaching artificial intelligence with agents. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):224–228, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hailperin:2000:CFY

[HAB^{+}00] Max Hailperin, David Arnow, Judith Bishop, Chester Lund, and Lynn Andrea Stein. Concurrency the first year (panel session): experience reports. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):407– 408, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Haberman:2002:FBP

[Hab02] Bruria Haberman. Frames and boxes: a pattern-based method for manipulating binary trees. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):60–64, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2002.bib.

Hummel:2008:NOC

[HAB^{+}08] Joe Hummel, Carl Alphonce, Joe Bergin, Michael Caspersen, Stuart Hansen, James Heliotis, and Michael Kölling. Nifty objects for CS0 and CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):437– 438, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Hadjerrouit:2001:WBA

[Had01] Said Hadjerrouit. Web-based application development: a software engineering approach. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(2): 31–34, June 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2001.bib.

Hadjerrouit:2005:CGP

[Had05] Said Hadjerrouit. Constructivism as guiding philosophy for software engineering education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4): 45–49, December 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2005.bib.

Hundhausen:2009:IPC

[HAFT09] Christopher Hundhausen, Anukrati Agrawal, Dana Fairbrother, and Michael Trevisan. Integrating pedagogical code reviews into a CS 1 course: an empirical study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):291–295, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Hagan:2000:UBT

[Hag00] Dianne Hagan. Using BlueJ to teach Java (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):188–189, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Haberman:2005:IRA

[HAG05] Bruria Haberman, Haim Averbuch, and David Ginat. Is it really an algorithm: the need for explicit discourse. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):74–78, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Holz:2006:RMC

[HAH^{+}06] Hilary J. Holz, Anne Gates Applin, Bruria Haberman, Donald Joyce, Helen C. Purchase, and Catherine Reed. Research methods in computing: what are they, and how should we teach them? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):96–114, December 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2006.bib.

Holz:2007:SRS

[HAJ07] Hilary Holz, Anne Applin, and William Joel. Status report of the SIGCSE committee on teaching computer science research methods. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):327–328, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hall:2008:WWS

[Hal08] Wendy Hall. What is Web science and why is it important to CSE. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):1–2, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Hause:2001:IFS

[HALW01] Martha L. Hause, Vicki L. Almstrum, Mary Z. Last, and M. R. Woodroffe. Interaction factors in software development performance in distributed student teams in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):69–72, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hamza:2000:CPC

[HAM00] M. Khalid Hamza, Bassem Alhalabi, and David M. Marcovitz. Creative pedagogy for computer learning: eight effective tactics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(4):70–73, December 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hamer:2002:HR

[Ham02a] John Hamer. Hashing revisited. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (3):80–83, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hamer:2002:MAT

[Ham02b] John Hamer. A musical approach to teaching design patterns. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):197, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hamelin:2004:SWD

[Ham04a] Denis Hamelin. Searching the Web to develop inquiry and collaborative skills. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(4):76–79, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2004.bib.

Hamer:2004:ATD

[Ham04b] John Hamer. An approach to teaching design patterns using musical composition. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):156–160, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hamza:2005:IOO

[Ham05] Haitham S. Hamza. Introducing object-oriented analysis and design in lower-level undergraduate courses: a pattern-based approach. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):350, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hamer:2006:SEC

[Ham06] John Hamer. Some experiences with the “contributing student approach”. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):68–72, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hamada:2007:WBT

[Ham07] Mohamed Hamada. Web-based tools for active learning in information theory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):60–64, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hannay:2002:ITC

[Han02] David G. Hannay. Interactive tools for computation theory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):68–70, December 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2002.bib.

Hansen:2004:GSI

[Han04] Stuart Hansen. The game of set(R): an ideal example for introducing polymorphism and design patterns. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):110–114, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hanks:2005:SPC

[Han05a] Brian Hanks. Student performance in CS1 with distributed pair programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):316–320, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hansen:2005:IJP

[Han05b] Stuart A. Hansen. Interpreting Java program runtimes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):36–40, March 2005. CO-DEN SIGSD3. ISBN 1-58113-997-7. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Compiler/java.bib;   ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/2005.
bib.

Hanks:2006:SAT

[Han06] Brian Hanks. Student attitudes toward pair programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):113–117, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hanson:2007:ITC

[Han07] Vicki L. Hanson. Inclusive thinking in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):3, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Hansen:2009:APP

[Han09] Stuart A. Hansen. Analyzing programming projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):377–381, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Hartley:2000:AYM

[Har00] Stephen J. Hartley. “Alfonse, you have a message!”. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):60–64, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hartley:2001:AGM

[Har01a] Stephen J. Hartley. “Alfonse, give me a call!”. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):229–232, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hartman:2001:CFC

[Har01b] Janet Hartman. The changing face of computing accreditation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(4):13–14, December 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2001.bib.

Harrison:2002:IIC

[Har02] Chris J. Harrison. ICPL: an initial concurrent programming language. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):101–105, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2002.bib.

Hartman:2003:CC

[Har03] Janet Hartman. Corporate clips. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35 (2):17–19, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2003.bib.

Harms:2004:VRS

[Har04] Douglas E. Harms. A virtual reality simulator of the ENIAC. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):239, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Harrison:2005:OEW

[Har05] Susan M. Harrison. Opening the eyes of those who can see to the world of those who can’t: a case study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):22–26, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hause:2005:DTP

[Hau05] Martha L. Hause. Distributed team performance in software development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):357, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hayat:2004:CSE

[HAuR04] Khawaja Amer Hayat, Umar Waqar Anis, and S. Tauseef ur Rehman. Cryptanalysis of some encryption/cipher schemes using related key attack. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (4):85–87, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). NOTE FROM ACM: It has been determined that the authors of this article plagiarized the contents from a previously published paper. Therefore ACM has shut off access to this paper.

Hawthorne:2007:VFF

[Haw07] Elizabeth K. Hawthorne. Visions of a future … without forgetting the past. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):19–20, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2007.bib.

Hawthorne:2008:CCC

[Haw08a] Elizabeth K. Hawthorne. Community college corner: crafting a compendium for associated-degree computing curricula. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):20–21, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hawthorne:2008:RGA

[Haw08b] Elizabeth K. Hawthorne. Revising the guidelines for associate-degree transfer curriculum in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):18, June 2008. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2008.bib.

Hawthorne:2009:CCCa

[Haw09a] Elizabeth K. Hawthorne. Community college corner: exploring CAP-space: the next frontier in curricula, assessment and pedagogy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):76–77, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hawthorne:2009:CCCb

[Haw09b] Elizabeth K. Hawthorne. Community college corner: Upcoming computing education summit for community colleges. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4):16–17, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hazzan:2001:PCS

[Haz01a] Orit Hazzan. On the presentation of computer science problems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(4):55–58, December 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2001.bib.

Hazzan:2001:THA

[Haz01b] Orit Hazzan. Teaching the human aspect of software engineering — a case study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):124– 128, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hazzan:2002:RAL

[Haz02] Orit Hazzan. Reducing abstraction level when learning computability theory concepts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):156– 160, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hazzan:2003:ACS

[Haz03a] Orit Hazzan. Application of computer science ideas to the presentation of mathematical theorems and proofs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):38–42, June 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

Hazzan:2003:CSA

[Haz03b] Orit Hazzan. Cognitive and social aspects of software engineering: a course framework. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):3–6, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hazzan:2003:CSS

[Haz03c] Orit Hazzan. Computer science students’ conception of the relationship between reward (grade) and cooperation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):178–182, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hazzan:2005:PDW

[Haz05] Orit Hazzan. Professional development workshop for female software engineers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):75– 79, June 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2005.bib.

Hazzan:2008:RTA

[Haz08] Orit Hazzan. Reflections on teaching abstraction and other soft ideas. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):40–43, June 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2008.bib.

Harms:2001:UPT

[HB01] Douglas Harms and Dave Berque. Using a PDP-11/10 to teach content and history in computer organization courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):209–213, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Higgins:2006:FCB

[HB06] Colin A. Higgins and Brett Bligh. Formative computer based assessment in diagram based domains. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):98–102, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hanks:2009:SUP

[HB09a] Brian Hanks and Matt Brandt. Successful and unsuccessful problem solving approaches of novice programmers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):24–28, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Huang:2009:UAI

[HB09b] Timothy Huang and Amy Briggs. A unified approach to introductory computer science: can one size fit all? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):253–257, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Huggins:2003:SGT

[HBCW03] James Huggins, Joseph Bergin, James Caristi, and Ellen Walker. Survivor: getting through that class the first time. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):234–235, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Henderson:2001:SMT

[HBD^{+}01] Peter B. Henderson, Doug Baldwin, Venu Dasigi, Marcel Dupras, Jane Fritz, David Ginat, Don Goelman, John Hamer, Lew Hitchner, Will Lloyd, Bill Marion, Jr., Charles Riedesel, and Henry Walker. Striving for mathematical thinking. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(4):114–124, December 2001. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2001.bib.

Henderson:2003:MEC

[HBEM03] Peter B. Henderson, William Barker, Susanna Epp, and William Marion. Math educators, computer science educators: working together. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1): 236–237, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hansen:2007:GSC

[HBH07] David M. Hansen, Joseph Bruce, and David Harrison. Give students a clue: a course-project for undergraduate artificial intelligence. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):44–48, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Heines:2003:PDM

[HBIG03] Jesse Heines, Katy Börner, Melody Y. Ivory, and Edward F. Gehringer. Panel on the development, maintenance, and use of course Web sites. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):94–95, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hendrix:2004:EFP

[HCB04] T. Dean Hendrix, James H. Cross II, and Larry A. Barowski. An extensible framework for providing dynamic data structure visualizations in a lightweight IDE. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):387–391, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hill:2001:UIN

[HCHP01] John M. D. Hill, Curtis A. Carver, Jr., Jeffrey W. Humphries, and Udo W. Pooch. Using an isolated network laboratory to teach advanced networks and security. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):36–40, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hamer:2008:CSP

[HCJ^{+}08] John Hamer, Quintin Cutts, Jana Jackova, Andrew Luxton-Reilly, Robert McCartney, Helen Purchase, Charles Riedesel, Mara Saeli, Kate Sanders, and Judithe Sheard. Contributing student pedagogy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):194–212, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hawthorne:2007:CRA

[HCK07] Elizabeth K. Hawthorne, Robert D. Campbell, and Karl J. Klee. Curricular resources from the ACM Two-Year College Education Committee. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3): 309, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Hendrix:2000:DVI

[HCMM00] T. Dean Hendrix, James H. Cross II, Saeed Maghsoodloo, and Matthew L. McKinney. Do visualizations improve program comprehensibility? experiments with control structure diagrams for Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):382–386, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Henderson:2007:CT

[HCW07] Peter B. Henderson, Thomas J. Cortina, and Jeannette M. Wing. Computational thinking. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1): 195–196, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Heaney:2004:MPI

[HD04] David Heaney and Charlie Daly. Mass production of individual feedback. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):117–121, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Huggins:2007:RTD

[HD07] Kevin L. Huggins and Rachelle DeCoste. Reflections on teaching discrete math for the first time. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):28–31, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2007.bib.

Henderson:2002:WMC

[HDA^{+}02] Peter B. Henderson, Paul De Palma, Vicki L. Almstrum, Orit Hazzan, and Kim Potter Kihlstrom. Women, mathematics and computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):131–132, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Hazzan:2006:QRC

[HDE^{+}06] Orit Hazzan, Yael Dubinsky, Larisa Eidelman, Victoria Sakhnini, and Mariana Teif. Qualitative research in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):408– 412, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Henderson:2006:PTD

[HDH06a] Peter B. Henderson, Rachelle DeCoste, and Kevin L. Huggins. Preparing to teach discrete math for the first time. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):20–21, December 2006. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2006.bib.

Hoffman:2006:BWL

[HDH06b] Mark E. Hoffman, Timothy Dansdill, and David S. Herscovici. Bridging writing to learn and writing in the discipline in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1): 117–121, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hislop:2006:RDC

[HE06] Gregory W. Hislop and Heidi J. C. Ellis. Retention of distance and on-campus students in a graduate computer science degree program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):342, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hansen:2007:EFP

[HE07] Stuart Hansen and Erica Eddy. Engagement and frustration in programming projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):271– 275, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hart:2008:NAK

[HEB08] Mindy Hart, James P. Early, and Dennis Brylow. A novel approach to K-12 CS education: linking mathematics and computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):286–290, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Heines:2003:EXS

[Hei03] Jesse M. Heines. Enabling XML storage from Java applets in a GUI programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):88–93, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Compiler/java.bib;   ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/2003.
bib.

Helmick:2007:IOC

[Hel07a] Michael T. Helmick. Integrated online courseware for computer science courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):146– 150, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Helmick:2007:IBP

[Hel07b] Michael T. Helmick. Interface-based programming assignments and automatic grading of Java programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):63–67, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Hemmendinger:2005:CHI

[Hem05] David Hemmendinger. Computing history: interesting times. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):366, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hislop:2009:ESE

[HEM09] Gregory W. Hislop, Heidi J. C. Ellis, and Ralph A. Morelli. Evaluating student experiences in developing software for humanity. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):263–267, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Henderson:2002:MMI

[Hen02] Peter B. Henderson. Making math interesting. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):26–28, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2002.bib.

Henderson:2003:IR

[Hen03a] Peter B. Henderson. Inductive reasoning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):27–29, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2003.bib.

Henderson:2003:IIS

[Hen03b] Peter B. Henderson. ISSAC: Indiana Student Software Awards Competition. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):332–335, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Henderson:2003:MIR

[Hen03c] Peter B. Henderson. More on inductive reasoning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):23–25, December 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2003.bib.

Henderson:2004:MM

[Hen04a] Peter B. Henderson. Modeling mania. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(2):23–24, June 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2004.bib.

Henderson:2004:PP

[Hen04b] Peter B. Henderson. Penny piles. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (4):22–24, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2004.bib.

Henderson:2005:MC

[Hen05a]  Peter B. Henderson.  Mathematics in the curricula.

SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):20–22, June 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2005.bib.

Henderson:2005:OMC

[Hen05b] Peter B. Henderson. One-minute column. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4):24–26, December 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2005.bib.

Henry:2005:TCC

[Hen05c] Tyson R. Henry. Teaching compiler construction using a domain specific language. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):7–11, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Henderson:2006:JFP

[Hen06a] Peter B. Henderson. The Josephus Flavius’ problem. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):17–20, June 2006. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2006.bib.

Henriksen:2006:GD

[Hen06b] Poul Henriksen. Greenfoot demonstration. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):355, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Henderson:2007:PCM

[Hen07a] Peter B. Henderson. Pre-college math concepts vs. skills. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):26–28, December 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Henderson:2007:RTD

[Hen07b] Peter B. Henderson. Reflections on teaching discrete math for the first time. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):24, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Henderson:2008:AMC

[Hen08a] Peter B. Henderson. Abstraction, model checking and software correctness. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):23–24, June 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2008.bib.

Henderson:2008:MCS

[Hen08b] Peter B. Henderson. Math counts: software correctness and a SIGCSE 2008 BoF. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):27–28, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Henderson:2009:MCA

[Hen09a] Peter B. Henderson. Math counts: Arguments, proofs, and …. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4):21–22, December 2009. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Henderson:2009:MCS

[Hen09b] Peter B. Henderson. Math CountS: SIGCSE 2009 and CS unplugged. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):81–82, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hepting:2006:EUT

[Hep06] Daryl H. Hepting. Ethics and usability testing in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):76– 80, June 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2006.bib.

Hislop:2009:UOS

[HETD09] Gregory W. Hislop, Heidi J. C. Ellis, Allen B. Tucker, and Scott Dexter. Using open source software to engage students in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1): 134–135, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Hansen:2004:EEG

[HF04] Stuart Hansen and Timothy Fossum. Events not equal to GUIs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):378–381, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Huen:2005:IOO

[HFHK05] Wing Huen, Ernest Ferguson, Peter B. Henderson, and Clifton Kussmaul. IT offshore outsourcing: impact on CS/IS curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):258–259, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hazzan:2008:MHS

[HGEB08] Orit Hazzan, Judith Gal-Ezer, and Lenore Blum. A model for high school computer science education: the four key elements that make it! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):281–285, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Hitchner:2001:PEC

[HGH^{+}01] Lewis E. Hitchner, Judith Gersting, Peter B. Henderson, Philip Machanick, and Yale N. Patt. Programming early considered harmful. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):402–403, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Heines:2009:MPI

[HGK09] Jesse M. Heines, Gena R. Greher, and Sarah Kuhn. Music performamatics: interdisciplinary interaction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):478–482, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Hood:2005:TPL

[HH05a] Cynthia S. Hood and Dennis J. Hood. Teaching programming and language concepts using LEGOs(R). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):19–23, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hood:2005:TIC

[HH05b] Cynthia S. Hood and Dennis J. Hood. Toward integrating computing concepts into the K-12 curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):375, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hood:2006:TSP

[HH06] Dennis J. Hood and Cynthia S. Hood. Teaching software project management using simulations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):289–293, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Halstensen:2007:HIO

[HH07] Maths Halstensen and Frode Haug. Highly interactive online study skills course: “unstuck”. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (3):332, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Hovemeyer:2004:RBM

[HHB04] David Hovemeyer, Jeffrey K. Hollingsworth, and Bobby Bhattacharjee. Running on the bare metal with GeekOS. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):315–319, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hannon:2005:RCE

[HHB05] Charles Hannon, Manfred Huber, and Lisa Burnell. Research to classroom: experiences from a multi-institutional course in smart home technologies. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):121– 125, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Henderson:2003:MDS

[HHF^{+}03] Peter B. Henderson, Lewis E. Hitchner, Sister Jane Fritz, Bill Marion, Christelle Scharff, John Hamer, and Charles Riedesel. Materials development in support of mathematical thinking. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):185–190, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

Holliday:2008:SRW

[HHJ08] Mark A. Holliday, J. Traynham Houston, and E. Matthew Jones. From sockets and RMI to Web services. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):236–240, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Hambrusch:2009:MAT

[HHK^{+}09] Susanne Hambrusch, Christoph Hoffmann, John T. Korb, Mark Haugan, and Antony L. Hosking. A multidisciplinary approach towards computational thinking for science majors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):183–187, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Hislop:2006:SSE

[HHLS06] Gregory W. Hislop, Thomas B. Hilburn, Michael J. Lutz, and Mark J. Sebern. Sharing software engineering curriculum materials. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):332, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hahne:2006:GRW

[HHRA06] Reg Hahne, Cay Horstmann, David Reed, and Don Allen. Growl… roar… are we ready for tiger?: review of the current climate and changes to be implemented for the 2007 AP CS exam. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):378, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hogan:2002:DME

[HHS02] Thomas R. Hogan, Patrick R. Harrison, and Kay G. Schulze. Developing and maintaining an effective assessment program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):52–56, December 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2002.bib.

Hickey:2004:SBW

[Hic04] Timothy J. Hickey. Scheme-based Web programming as a basis for a CS0 curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):353– 357, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hilburn:2000:TNP

[Hil00] Thomas B. Hilburn. Teams need a process! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):53–56, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hill:2004:EGA

[Hil04] Thomas G. Hill. Excel grader and access grader. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(2):101–105, June 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2004.bib.

Hazzan:2005:UHC

[HILS05] Orit Hazzan, John Impagliazzo, Raymond Lister, and Shimon Schocken. Using history of computing to address problems and opportunities. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):126–127, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hislop:2009:RIP

[His09] Gregory W. Hislop. Replacing introductory programming courses with a broader perspective on computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):390, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Hitchcock:2005:ICV

[Hit05] Leo Hitchcock. Industry certification: value, validity, and a place for SoDIS(R). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4):59–63, December 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2005.bib.

Haberman:2001:ABB

[HK01] Bruria Haberman and Yifat Ben-David Kolikant. Activating “black boxes” instead of opening “zipper” — a method of teaching novices basic CS concepts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33 (3):41–44, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hawthorne:2008:CSI

[HKCW08] Elizabeth K. Hawthorne, Karl J. Klee, Robert D. Campbell, and Anita M. Wright. Computer science and information technology associate-level curricular guidelines. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):350, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Hawthorne:2009:RAL

[HKCW09] Elizabeth K. Hawthorne, Karl J. Klee, Robert D. Campbell, and Anita M. Wright. Revised associate-level curricular guidelines in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (3):374, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Hosny:2000:IJB

[HKF00] H. Hosny, O. Khaled, and M. E. Fathalla. ILE: a Java-based environment for CS courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):152– 155, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Holdfeldt:2002:EEE

[HKL^{+}02] Peter Holdfeldt, Boris Koldehofe, Carina Lindskog, Torbjörn Olsson, Wanja Petersson, Jonas Svensson, and Linus Valtersson. EnViDiA: an educational environment for visualization of distributed algorithms in virtual environments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):226, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hunt:2002:HDG

[HKN^{+}02] Fran Hunt, Joe Kmoch, Chris Nevison, Susan Rodger, and Julie Zelenski. How to develop and grade an exam for 20,000 students (or maybe just 200 or 20). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (1):285–286, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Hickey:2002:ICC

[HKW^{+}02] Timothy Hickey, Amruth Kumar, Linda Wilkens, Andrew Beiderman, Aparna Mahadev, and Heidi Ellis. Internetcentric computing in the Computer Science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):50–51, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Hazzan:2004:CPP

[HL04a] Orit Hazzan and Tami Lapidot. Construction of a professional perception in the “methods of teaching computer science” course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(2):57–61, June 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2004.bib.

Hazzan:2004:PCS

[HL04b] Orit Hazzan and Tami Lapidot. The practicum in computer science education: bridging gaps between theoretical knowledge and actual performance. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(4): 47–51, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2004.bib.

Holliday:2004:CAU

[HL04c] Mark A. Holliday and David Luginbuhl. CS1 assessment using memory diagrams. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):200–204, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hazzan:2006:SIC

[HL06] Orit Hazzan and Tami Lapidot. Social issues of computer science in the “Methods of Teaching Computer Science in the High School” course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):72– 75, June 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2006.bib.

Havill:2007:TSF

[HL07] Jessen T. Havill and Lewis D. Ludwig. Technically speaking: fostering the communication skills of computer science and mathematics students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):185– 189, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hnatyshin:2008:UDC

[HL08] Vasil Y. Hnatyshin and Andrea F. Lobo. Undergraduate data communications and networking projects using opnet and wireshark software. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):241–245, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Huss-Lederman:2008:SFP

[HLCS08] Steven Huss-Lederman, Donald Chinn, and James Skrentny. Serious fun: peer-led team learning in CS. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):330–331, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Haberman:2003:ART

[HLL03] Bruria Haberman, Ela Lev, and Dorothy Langley. Action research as a tool for promoting teacher awareness of students’ conceptual understanding. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):144– 148, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Haller:2008:GGE

[HLL^{+}08] Susan Haller, Brian Ladd, Scott Leutenegger, John Nordlinger, Jody Paul, Henry Walker, and Carol Zander. Games: good/ evil. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):219–220, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Harlan:2001:KRK

[HLM01] Robert M. Harlan, David B. Levine, and Shelley McClarigan. The Khepera robot and the kRobot class: a platform for introducing robotics in the undergraduate curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):105–109, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hurst:2007:SAA

[HLN07] Wolfgang Hürst, Tobias Lauer, and Eveline Nold. A study of algorithm animations on mobile devices. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):160–164, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Herzog:2001:RTL

[HLNT01] Erik Herzog, Peter Loborg, and Simin Nadjm-Tehrani. Realtime lab exercises: a teacher’s dilemma. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):273–277, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Holland:2002:NIO

[HLS02] David A. Holland, Ada T. Lim, and Margo I. Seltzer. A new instructional operating system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1): 111–115, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Hagan:2000:DIH

[HM00] Dianne Hagan and Selby Markham. Does it help to have some programming experience before beginning a computing degree program? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):25–28, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hummel:2002:UVB

[HM02] Joe Hummel and Jean Mehta. Using Visual Basic in the CS curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):283–284, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Harlan:2005:CEB

[HM05] Robert M. Harlan and Shelley McClarigan. Creating emergent behaviors: two robotics labs that combine reactive behaviors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):441–445, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hanks:2004:PQP

[HMDK04] Brian Hanks, Charlie McDowell, David Draper, and Milovan Krnjajic. Program quality with pair programming in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):176–180, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hristova:2003:ICJ

[HMRM03] Maria Hristova, Ananya Misra, Megan Rutter, and Rebecca Mercuri. Identifying and correcting Java programming errors for introductory computer science students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):153–156, January 2003. CODEN SIGSD3. ISBN 1-58113-648-X. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Compiler/java.bib;   ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/2003.
bib.

Hu:2004:TLI

[HMS04] Ji Hu, Christoph Meinel, and Michael Schmitt. Tele-lab IT security: an architecture for interactive lessons for security education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):412–416, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Humphrey:2008:PSI

[HMS08] Watts Humphrey, Robert Musson, and Rafael Salazar. Preparing students for industry’s software engineering needs.

SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):481–482, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Hanks:2009:CSS

[HMS^{+}09] Brian Hanks, Laurie Murphy, Beth Simon, Renée McCauley, and Carol Zander. CS1 students speak: advice for students by students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):19–23, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Hundhausen:2008:ESB

[HNC08] Christopher D. Hundhausen, N. Hari Narayanan, and Martha E. Crosby. Exploring studio-based instructional models for computing education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):392– 396, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Hobbs:2000:EGL

[Hob00] Mike Hobbs. Email groups for learning and assessment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):183, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Holliday:2000:KBS

[Hol00] Mark A. Holliday. A kernel-based synchronization assignment for the operating systems course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (3):184, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Holmboe:2005:LOO

[Hol05] Christian Holmboe. The linguistics of object-oriented design: implications for teaching. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):188– 192, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hollingsworth:2008:TQW

[Hol08] Joseph E. Hollingsworth. Teaching query writing: an informed instruction approach. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):351, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Hoo:2008:SOM

[Hoo08] Jason T. Soo Hoo. Self-organized maps in scientific data analysis. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):372, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Hovis:2005:MCF

[Hov05] Robert A. Hovis. Managing the complexity in first year programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):394, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Howatt:2002:OSP

[How02] James Howatt. Operating systems projects: Minix revisited. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):109–111, December 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2002.bib.

Howles:2003:FGS

[How03] Trudy Howles. Fostering the growth of a software quality culture. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):45–47, June 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

Howles:2005:CAF

[How05] Trudy Howles. Community and accountability in a first year programming sequence. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):99– 102, June 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2005.bib.

Hughes:2006:ACS

[HP06] Janet Hughes and D. Ramanee Peiris. ASSISTing CS1 students to learn: learning approaches and object-oriented programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):275–279, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Honig:2007:COE

[HP07] William L. Honig and Tejasvini Prasad. A classroom outsourcing experience for software engineering learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):181–185, September 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Herrmann:2003:RIC

[HPC^{+}03] Nira Herrmann, Jeffrey L. Popyack, Bruce Char, Paul Zoski, Christopher D. Cera, Robert N. Lass, and Aparna Nanjappa. Redesigning introductory computer programming using multi-level online modules for a mixed audience. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):196–200, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Herrmann:2004:ACR

[HPCZ04] Nira Herrmann, Jeffrey L. Popyack, Bruce Char, and Paul Zoski. Assessment of a course redesign: introductory computer programming using online modules. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):66–70, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hepting:2008:CSB

[HPM^{+}08] Daryl H. Hepting, Lijuan Peng, Timothy J. Maciag, David Gerhard, and Brien Maguire. Creating synergy between usability courses and open source software projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):120–123, June 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2008.bib.

Hughes:2007:PAS

[HPT07] Janet Hughes, D. Ramanee Peiris, and Paul T. Tymann, editors. Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education, ITiCSE 2007, Dundee, Scotland, UK, June 25–27, 2007, SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 2007. CODEN SIGSD3. ISBN 1-59593-610-6. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hung:2000:IVI

[HR00] Ted Hung and Susan H. Rodger. Increasing visualization and interaction in the automata theory course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):6–10, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hansen:2002:TSC

[HR02] Klaus Marius Hansen and Anne Vinter Ratzer. Tool support for collaborative teaching and learning of object-oriented modeling. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):146–150, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Holloway:2003:DMS

[HR03] Jonathan Holloway and Mark Ratcliffe. Development of a MLE by students for students using extreme programming techniques. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):225, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Horwitz:2009:UPL

[HRB^{+}09] Susan Horwitz, Susan H. Rodger, Maureen Biggers, David Binkley, C. Kolin Frantz, Dawn Gundermann, Susanne Hambrusch, Steven Huss-Lederman, Ethan Munson, Barbara Ryder, and Monica Sweat. Using peer-led team learning to increase participation and success of under-represented groups in introductory computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (1):163–167, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Hill:2003:PGA

[HRBC03] John M. D. Hill, Clark K. Ray, Jean R. S. Blair, and Curtis A. Carver, Jr. Puzzles and games: addressing different learning styles in teaching operating systems concepts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):182–186, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Holt:2006:IPF

[HRS06] Alexander Holt, Sarah Rauchas, and Ian Sanders. Introducing Python into the first year curriculum at Wits. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):335, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Humphreys:2002:LAS

[HS02] Sheila Humphreys and Ellen Spertus. Leveraging an alternative source of computer scientists: reentry programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):53–56, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2002.bib. Women and

Computing.

Huang:2003:TUS

[HS03] Timothy Huang and Frank Swenton. Teaching undergraduate software design in a liberal arts environment using RoboCup. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):114–118, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hamada:2004:CET

[HS04] Mohamed Hamada and Kazuhiko Shiina. A classroom experiment for teaching automata. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):261, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hsiao:2009:EPP

[HSB09] I-Han Hsiao, Sergey Sosnovsky, and Peter Brusilovsky. Extending parameterized problem-tracing questions for Java with personalized guidance. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):392, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Hill:2005:GP

[HSD05] Curt Hill, Brian M. Slator, and Lisa M. Daniels. The grader in ProgrammingLand. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):211–215, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hill:2006:AHS

[HSM06] Curt Hill, Vijayakumar Shanmugasundaram, and Martina Miteva. Agents help students in ProgrammingLand. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):183–187, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hsia:2005:TJC

[HSSC05] James I. Hsia, Elspeth Simpson, Daniel Smith, and Robert Cartwright. Taming Java for the classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):327–331, March 2005. CO-DEN SIGSD3. ISBN 1-58113-997-7. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Compiler/java.bib;   ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/2005.
bib.

Higgins:2002:MSC

[HST02] Colin Higgins, Pavlos Symeonidis, and Athanasios Tsintsifas. The marking system for CourseMaster. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):46–50, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hilburn:2000:SQC

[HT00] Thomas B. Hilburn and Massood Townhidnejad. Software quality: a curriculum postscript? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (1):167–171, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Harmeyer:2001:PSI

[HTBS01] Kathleen Harmeyer, Donna Tupper, William Beck, and Sylvia Sorkin. Preparing students for Internet and multimedia technology careers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):376– 380, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hamilton-Taylor:2002:SSA

[HTK02] Ashley George Hamilton-Taylor and Eileen Kraemer. SKA: supporting algorithm and data structure discussion. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):58–62, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Hayes:2007:IAA

[HTSW07] Alan Hayes, Pete Thomas, Neil Smith, and Kevin Waugh. An investigation into the automated assessment of the designcode interface. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):324, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Howe:2004:CFA

[HTW04] Emily Howe, Matthew Thornton, and Bruce W. Weide. Components-first approaches to CS1/CS2: principles and practice. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):291–295, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hu:2003:FAA

[Hu03] Chenglie Hu. A framework for applet animations with controls. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):90–93, December 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

Hu:2005:AIS

[Hu05] Chenglie Hu. 2D array implementation of sorted lists. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):36–40, June 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2005.bib.

Hu:2006:WUI

[Hu06] Chenglie Hu. When to use an interface? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):86–90, June 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2006.bib.

Huang:2003:GGI

[Hua03] Timothy Huang. The game of go: an ideal environment for capstone and undergraduate research projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):84–88, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Huff:2002:GSD

[Huf02] Chuck Huff. Gender, software design, and occupational equity. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):112–115, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2002.bib. Women and

Computing.

Huggins:2009:ECS

[Hug09] James K. Huggins. Engaging computer science students through cooperative education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4): 90–94, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Huizinga:2001:ITI

[Hui01] Dorota M. Huizinga. Identifying topics for instructional improvement through on-line tracking of programming assignments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):129–132, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Humphrey:2004:WSF

[Hum04] Marty Humphrey. Web services as the foundation for learning complex software system development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):457–461, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hunt:2003:UIP

[Hun03] Kenny Hunt. Using image processing to teach CS1 and CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):86–89, December 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

Hunt:2005:JFE

[Hun05] Kenny Hunt. A Java framework for experimentation with steganography. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):282– 286, March 2005. CODEN SIGSD3. ISBN 1-58113-997-7. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Compiler/java.bib;   ftp://ftp.
math.utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2005.bib.

Huth:2004:MER

[Hut04] Michael Huth. Mathematics for the exploration of requirements. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(2):34–39, June 2004. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2004.bib.

Hoffman:2005:CLW

[HV05] Mark E. Hoffman and David R. Vance. Computer literacy: what students know and from whom they learned it. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):356–360, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hoffman:2007:GDT

[HV07] Mark E. Hoffman and David R. Vance. Gender difference trends in computer literacy of first-year students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):405–409, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hielscher:2006:ALE

[HW06] Michael Hielscher and Christian Wagenknecht. AtoCC: learning environment for teaching theory of automata and formal languages. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):306, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Head:2008:PUA

[HW08] Christopher C. D. Head and Steven A. Wolfman. Poogle and the unknown-answer assignment: open-ended, sharable CS1 assignments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):133–137, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Holliday:2005:GDA

[HWH^{+}05] Mark A. Holliday, Barry Wilkinson, Jeffrey House, Samir Daoud, and Clayton Ferner. A geographically-distributed, assignment-structured undergraduate grid computing course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):206–210, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hanks:2008:IAC

[HWRC08] Brian Hanks, Carol Wellington, Tom Reichlmayr, and Cyril Coupal. Integrating agility in the CS curriculum: practices through values. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):19–20, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Holden:2001:ITP

[HY01] Jan Holden and Alison Young. Innovative teaching practices in computing education (poster session): the TLA project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):174, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Haberman:2009:LTS

[HY09] Bruria Haberman and Cecile Yehezkel. Long-term software projects development: the affect of students’ self-appreciation and initial expectations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):363, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Hubscher-Younger:2003:CCL

[HYN03] Teresa Hübscher-Younger and N. Hari Narayanan. Constructive and collaborative learning of algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):6–10, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ibbett:2002:WVC

[Ibb02] Roland N. Ibbett. WWW visualisation of computer architecture simulations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):247, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Impagliazzo:2002:IDL

[ICL02] John Impagliazzo, Lillian (Boots) Cassel, and John A. N. Lee. ICT and digital libraries. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):237, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ibbett:2006:CAS

[IDD06] Roland N. Ibbett, J. C. Diaz y Carballo, and D. A. W. Dolman. Computer architecture simulation models. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):353, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Impagliazzo:2004:ICA

[IHL^{+}04] John Impagliazzo, Esther A. Hughes, Richard LeBlanc, Tim Lethbridge, Andrew McGettrick, Ann E. K. Sobel, Pradip K. Srimani, and Mitchell D. Theys. IEEE-CS/ACM computing curricula: computer engineering & software engineering volumes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):450–452, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ilias:2006:USC

[IK06] Aristidis Ilias and Maria Kordaki. Undergraduate studies in computer science and engineering: gender issues. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):81–85, June 2006. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2006.bib.

Impagliazzo:2004:UHE

[IL04] John Impagliazzo and John A. N. Lee. Using history to enhance computing courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):238, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Imberman:2003:TNN

[Imb03] Susan P. Imberman. Teaching neural networks using LEGO handy board robots in an artificial intelligence course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):312–316, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Impagliazzo:2001:RVC

[Imp01] John Impagliazzo. Real and virtual computing museums (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):185, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Impagliazzo:2002:IHS

[Imp02a] John Impagliazzo. Integrating the history of systems software in the computing curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):221, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Impagliazzo:2002:PTT

[Imp02b] John Impagliazzo, editor. Proceedings of the Thirty-third SIGCSE Technical Symposium on Computer Science Education (SIGCSE-02), 27 February–3 March 2002, Northern Kentucky, USA, volume 34(1) of SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Impagliazzo:2003:CECb

[Imp03] John Impagliazzo. Computer engineering curriculum development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):221, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Impagliazzo:2004:PST

[Imp04] John Impagliazzo, editor. Proceedings of the 35th SIGCSE technical symposium on Computer science education (SIGCSE-04), 3–7 March 2004, Norfolk, Virginia, USA, volume 36(1) of SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Impagliazzo:2005:CCO

[Imp05a] John Impagliazzo. Computing curricula overview project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):347, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Impagliazzo:2005:MOC

[Imp05b] John Impagliazzo. Meet our columnists for this issue. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):26, June 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2005.bib.

Impagliazzo:2005:ULH

[Imp05c] John Impagliazzo. Use a little history. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):379, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Impagliazzo:2006:CCG

[Imp06a] John Impagliazzo. Computing curricula 2005: and guide. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):311, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Impagliazzo:2006:MOC

[Imp06b] John Impagliazzo. Meet our columnists for this issue. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):23, June 2006. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2006.bib.

Impagliazzo:2007:UOB

[Imp07] John Impagliazzo. Using an outcome-based approach to assess computing programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3): 344, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Impagliazzo:2008:RSS

[Imp08] John Impagliazzo. Response summary from the SIGCSE community. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):113, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Impagliazzo:2009:GAP

[Imp09] John Impagliazzo. Graduate attributes and performance measures: refinements in assessing programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):370, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Iglesias:2009:ASE

[IMRJ09] Ana Iglesias, Lourdes Moreno, Pablo Revuelta, and Javier Jiménez. APEINTA: a Spanish educational project aiming for inclusive education in and out of the classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):393, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Inghelbrecht:2009:OOD

[Ing09a] Yanic Inghelbrecht. Object-oriented design with trace modeler and Trace4J. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):375, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Inghelbrecht:2009:TST

[Ing09b] Yanic Inghelbrecht. Tool support for teaching responsibilitydriven design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):351, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Irani:2004:UGC

[Ira04] Lilly Irani. Understanding gender and confidence in CS course culture. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):195–199, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ibanez:2009:CRU

[IS09] Jesús Ibáñez and Ana Sánchez. Constructive reduction: understanding uncomputability through programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):90–94, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Isbell:2009:RDC

[ISC^{+}09] Charles L. Isbell, Lynn Andrea Stein, Robb Cutler, Jeffrey Forbes, Linda Fraser, John Impagliazzo, Viera Proulx, Steve Russ, Richard Thomas, and Yan Xu. (re)defining computing curricula by (re)defining computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (4):195–207, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Impagliazzo:2003:CECa

[ISMS03] John Impagliazzo, Robert Sloan, Andrew McGettrick, and Pradip K. Srimani. Computer engineering computing curricula. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):355–356, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jones:2003:RCC

[JA03] Edward L. Jones and Clement S. Allen. Repositories for CS courses: an evolutionary tale. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3): 119–123, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jackson:2000:SAA

[Jac00a] David Jackson. A semi-automated approach to online assessment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):164–167, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jacobson:2000:UCE

[Jac00b] Norman Jacobson. Using on-computer exams to ensure beginning students’ programming competency. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(4):53–56, December 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jacobson:2001:MNS

[Jac01] Norman Jacobson. A method for normalizing students’ scores when employing multiple graders. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33 (4):35–38, December 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2001.bib.

Jackova:2008:LMI

[Jac08] Jana Jacková. Learning for mastery in an introductory programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):352, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Jamieson:2002:SLC

[Jam02] Leah H. Jamieson. Service learning in computer science and engineering. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):133–134, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

James:2005:ESP

[Jam05] Richard H. James. External sponsored projects: lessons learned. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):94–98, June 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2005.bib.

Janzen:2001:ROR

[Jan01] David Janzen. From RS-232 to object request brokers: incremental object-oriented networking projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):41–44, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jones:2007:SSN

[JB07] Sue Jane Jones and Gary E. Burnett. Spatial skills and navigation of source code. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):231– 235, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Judith:2003:CSP

[JBB^{+}03] Williams C. Judith, Bettina Bair, Jürgen Börstler, Lethbridge C. Timothy, and Ken Surendran. Client sponsored projects in software engineering courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):401–402, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jipping:2003:UJT

[JBMP03] Michael J. Jipping, Agata Bugaj, Liliyana Mihalkova, and Donald E. Porter. Using Java to teach networking concepts with a programmable network sniffer. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):120–124, January 2003. CO-DEN SIGSD3. ISBN 1-58113-648-X. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Compiler/java.bib;   ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/2003.
bib.

Jensen:2006:TUC

[JCG06] Bradley K. Jensen, Melinda Cline, and Carl Stephen Guynes. Teaching the undergraduate CS information security course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):61–63, June 2006. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2006.bib.

Juliano:2003:OUT

[JCK03] Benjoe A. Juliano, Chi-Hui Chen, and Elena Kroumova. Observations from using two modes of teaching undergraduate computer architecture. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):248, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jipping:2007:TSJ

[JCOP07] Michael J. Jipping, Cameron Calka, Brian O’Neill, and Christopher R. Padilla. Teaching students Java bytecode using Lego Mindstorms robots. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):170– 174, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jovanovic-Dolecek:2000:CES

[JDC00] Gordana Jovanovic-Dolecek and Victor H. Champac. CGT-DEMO — educational software for the central limit theorem. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(2):46–48, June 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jovanovic-Dolecek:2008:UMT

[JDFV08] Gordana Jovanovic-Dolecek and Alfonso Fernandez-Vazquez. Use of MATLAB in teaching the fundamentals of random variables. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):46–51, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jimenez-Diaz:2005:SBU

[JDGAGMGC05] Guillermo Jiménez-Díaz, Mercedes Gómez-Albarrán, Marco A. Gómez-Martin, and Pedro A. González-Calero. Software behaviour understanding supported by dynamic visualization and role-play. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):54–58, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jenkins:2001:MVX

[Jen01a] Tony Jenkins. Motivation = value x expectancy (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):174, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jenkins:2001:MSP

[Jen01b] Tony Jenkins. The motivation of students of programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):53–56, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jenkins:2008:TCA

[Jen08] Marcelo Jenkins. Teaching computer aided software engineering at the graduate level. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):63–67, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Jarc:2000:ABI

[JFH00] Duane J. Jarc, Michael B. Feldman, and Rachelle S. Heller. Assessing the benefits of interactive prediction using Webbased algorithm animation courseware. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):377–381, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jukic:2008:URD

[JG08] Nenad Jukic and Paul Gray. Using real data to invigorate student learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):6–10, June 2008. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2008.bib.

Jipping:2006:HIF

[JHLT06] Michael J. Jipping, Sara Henry, Kathleen Ludewig, and Leslie Tableman. How to integrate FPGAs into a computer organization course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):234–238, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jin:2008:PPA

[Jin08] Wei Jin. Pre-programming analysis tutors help students learn basic programming concepts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):276– 280, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Jacobsen:2005:TCC

[JJ05] Christian L. Jacobsen and Matthew C. Jadud. Towards concrete concurrency: occam-pi on the LEGO Mindstorms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):431–435, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jamba-Joyner:2003:PSD

[JJK03] Lisa Jamba-Joyner and William F. Klostermeyer. Predictors for success in a discrete math course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35 (2):66–69, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2003.bib.

Jimeno:2007:CYO

[JJOA07] Roberto Jimeno, M. S. Jorge, and L. Ortega-Arjona. Curb your objects!: an orthodox form for C# classes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):138–141, June 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Joyce:2004:PST

[JKDN04] Dan Joyce, Deborah Knox, Wanda Dann, and Thomas L. Naps, editors. Proceedings of the thirty-fifth SIGCSE Technical Symposium on Computer Science Education: SIGCSE 2004: Norfolk, Virginia, USA, March 3–7, 2004, SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 2004. CODEN SIGSD3. ISBN 1-58113-798-2. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jipping:2001:UHC

[JKDS01] Michael J. Jipping, Joshua Krikke, Sarah Dieter, and Samantha Sandro. Using handheld computers in the classroom: laboratories and collaboration on handheld machines. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):169–173, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jipping:2004:IWW

[JKKL04] Michael J. Jipping, Andrew Kalafut, Nathan Kooistra, and Kathleen Ludewig. Investigating wired and wireless networks using a Java-based programmable sniffer. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):12–16, September 2004. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Compiler/java.bib.

Joyce:2007:RIS

[JKM07] Dan Joyce, Bruce Klein, and Renée McCauley. Report on improving SIGCSE/ITiCSE program creation process. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):128–129, December 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Jacobsen:2000:EDL

[JKS00] Michele Jacobsen, Rob Kremer, and Mildred L. G. Shaw. Experiments with distance learning in software engineering graduate courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(2):56–59, June 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Johnson:2002:DCM

[JM02] Deborah G. Johnson and Keith W. Miller. Is diversity in computing a moral matter? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2): 9–10, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/

2002.bib. Women and Computing.

Jipping:2002:UJD

[JMS02] Michael J. Jipping, Steve Marlowe, and Alexander Sherstov. Using Java to design and test hardware circuits over a classroom network. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):162–166, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Joel:2006:ECS

[Joe06a] William J. Joel. Engaging computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):316, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Joel:2006:LDW

[Joe06b] William J. Joel. Living in a digital world: teaching computers and society via literature. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):303, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Johnson:2002:NSF

[Joh02] Jeffrey Johnson. National Science Foundation Graduate Research Fellowship Program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):215– 216, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Jones:2000:DIC

[Jon00] Randolph M. Jones. Design and implementation of computer games: a capstone course for undergraduate computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):260–264, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jones:2001:ITC

[Jon01] Edward L. Jones. Integrating testing into the curriculum — arsenic in small doses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):337– 341, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jones:2002:CWP

[Jon02] Anita Jones. The curious ways of professional cultures and the “two-body opportunity”. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):24– 25, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/

2002.bib. Women and Computing.

Jones:2005:POT

[Jon05] Mark Jones. The pedagogic opportunities of touch-screen voting. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):223–226, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Joyce:2001:SEM

[Joy01] Daniel Joyce. Sneaking in extra material (panel session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):190, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Joyce:2002:GWP

[Joy02] Donald Joyce. Group work at postgraduate level: some issues. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):220, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Joyce:2006:ECP

[Joy06a] Donald Joyce. Educating computing professionals at postgraduate level. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):315, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Joyce:2006:RAA

[Joy06b] Donald Joyce. Raising awareness about academic integrity. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):350, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jepson:2002:PP

[JP02] Andrea Jepson and Teri Perl. Priming the pipeline. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):36–39, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2002.bib. Women and

Computing.

Joseph:2003:GDC

[JP03] Anthony Joseph and Mabel Payne. Group dynamics and collaborative group performance. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1): 368–371, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Janzen:2006:TDL

[JS06] David S. Janzen and Hossein Saiedian. Test-driven learning: intrinsic integration of testing into the CS/SE curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):254–258, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Jacobson:2008:PPC

[JS08a] Norman Jacobson and Suzanne K. Schaefer. Pair programming in CS1: overcoming objections to its adoption. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):93–96, June 2008. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2008.bib.

Janzen:2008:TDL

[JS08b] David Janzen and Hossein Saiedian. Test-driven learning in early programming courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):532– 536, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Jadud:2008:PPP

[JSJ08] Matthew C. Jadud, Jon Simpson, and Christian L. Jacobsen. Patterns for programming in parallel, pedagogically. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):231–235, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Jacobson:2004:ITE

[JT04] Norman Jacobson and Alex Thornton. It is time to emphasize ArrayLists over Arrays in Java-based first programming courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(4):88– 92, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Compiler/java.bib;   ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/2004.
bib.

Ji:2007:SCL

[JWC07] Jeong-Hoon Ji, Gyun Woo, and Hwan-Gue Cho. A source code linearization technique for detecting plagiarized programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):73–77, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Ji:2008:EDP

[JWC08] Jeong-Hoon Ji, Gyun Woo, and Hwan-Gue Cho. An experience of detecting plagiarized source codes in competitive programming contests. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):369, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Joyce:2001:EDB

[JY01a] Donald Joyce and Alison Young. Electronic discussion boards (poster session): their use in post graduate computing courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):182, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Joyce:2001:PPP

[JY01b] Donald Joyce and Alison Young. The PASS project (poster session): group research into parameters affecting student success. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):177, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Joyce:2004:API

[JY04] Donald Joyce and Alison Young. Addressing professional issues at postgraduate level. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):257, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kultur:2009:FAP

[KA09] Can Kültür and Ilgim Veryeri Alaca. Fine arts perspective in user interface design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):386, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Kaczmarczyk:2001:ASAb

[Kac01a] Lisa C. Kaczmarczyk. Accreditation and student assessment in distance education (poster session): why we all need to pay attention. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):113–116, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kaczmarczyk:2001:ASAa

[Kac01b] Lisa C. Kaczmarczyk. Accreditation and student assessment in distance education: why we all need to pay attention. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):113–116, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kaczmarczyk:2003:TWC

[Kac03] Lisa C. Kaczmarczyk. A technical writing class for computer science majors: measuring student perceptions of learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):341–345, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kaczmarczyk:2006:UCM

[Kac06] Lisa C. Kaczmarczyk. The ubiquity of computing may be education’s biggest challenge. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4): 18–19, December 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2006.bib.

Kaczmarczyk:2007:RPP

[Kac07a] Lisa C. Kaczmarczyk. Resist and persist (productively). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):23–24, December 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Kaczmarczyk:2007:WCD

[Kac07b] Lisa C. Kaczmarczyk. We can do interdisciplinarity. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):20–21, June 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Kaczmarczyk:2008:PII

[Kac08a] Lisa Kaczmarczyk. Percolations: interdisciplinary innovation may invoke carnivorous colleagues. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (4):23–24, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kaczmarczyk:2008:DCS

[Kac08b] Lisa C. Kaczmarczyk. Drink the coffee and see the puffy white clouds. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):20–22, June 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2008.bib.

Kaczmarczyk:2009:PGC

[Kac09a] Lisa Kaczmarczyk. Percolations: get a clue: ditch the manual and take the users’ cues. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):79– 80, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kaczmarczyk:2009:PMA

[Kac09b] Lisa Kaczmarczyk. Percolations: Mind alterations: the IDE and their conceptual development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (4):19–20, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kagawa:2009:WWB

[Kag09] Koji Kagawa. WAPPEN: a Web-based application framework for programming and its bison/flex plug-in. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):383, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Kane:2007:ESP

[Kan07] Shaun K. Kane. Engaging student programmers as inclusive designers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):342, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Kaplan:2004:TCU

[Kap04] Daniel T. Kaplan. Teaching computation to undergraduate scientists. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):358–362, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Korte:2007:LGB

[KAPG07] Laura Korte, Stuart Anderson, Helen Pain, and Judith Good. Learning by game-building: a novel approach to theoretical computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):

53–57, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Kawell:2007:CRW

[Kaw07] Greg Kawell. Concepts to real world implementation via service learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):113–116, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2007.bib.

Kay:2000:IPL

[Kay00] David G. Kay. Intellectual property law basics for computer science instructors (seminar session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (1):425, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kosa:2001:XBD

[KB01] Martha J. Kosa and Mark A. Boshart. XML and browser development activities in CS2 (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):182, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kolling:2004:EAB

[KB04a] Michael Kölling and David J. Barnes. Enhancing apprenticebased learning of Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (1):286–290, March 2004. CODEN SIGSD3. ISBN 1-58113-798-2. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Compiler/java.bib;   ftp://ftp.
math.utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2004.bib.

Kosa:2004:TVC

[KB04b] Martha J. Kosa and Mark A. Boshart. Treemap visualizations for CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):259, September 2004.

CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kiesmueller:2008:HDG

[KB08] Ulrich Kiesmueller and Torsten Brinda. How do 7th graders solve algorithmic problems?: a tool-based analysis. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):353, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Kiesmueller:2009:AIL

[KB09] Ulrich Kiesmueller and Torsten Brinda. Automatically identifying learners’ problem solving strategies in-process solving algorithmic problems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):354, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Kolikant:2000:AS

[KBAP00] Yifat Ben-David Kolikant, Mordechai Ben-Ari, and Sarah Pollack. The anthropology semaphores. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):21–24, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

King:2001:WCM

[KBC01] L. A. Smith King, John Barr, and Ben Coleman. What could be more SLic?: projects for the programming languages course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):248–252, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Klassner:2007:MRA

[KC07] Frank Klassner and Christopher Continanza. Mindstorms without robotics: an alternative to simulations in systems courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):175–179, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kay:2000:TAS

[KCF^{+}00] David G. Kay, Clare Bates Congdon, Sue Fitzgerald, Merle S. King, and Pat Semmes. Teaching advice and support for new and adjunct faculty (panel session): experiences, policies, and strategies. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):414–415, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Knox:2003:UNC

[KCI03] Deborah Knox, Lillian N. Cassel, and John Impagliazzo. Using the NSDL and CITIDEL to enhance teaching. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):246, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kessler:2007:ITA

[KD07] Robert Kessler and Nathan Dykman. Integrating traditional and agile processes in the classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (1):312–316, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Klein:2000:TII

[KDH^{+}00] Bruce J. Klein, Mats Daniels, Dianne Hagan, Anders Berglund, Annegret Goold, Mary Last, Tony Clear, and Erkki Sutinen. Teaching inter-institutional courses (panel session): sharing challenges and resources. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (3):178–179, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Koffman:2000:IPC

[KDM^{+}00] Elliot Koffman, Dorothy Deremer, Chris McDonald, Loren Rhodes, Rebecca Thomas, A. Joe Turner, and Curt White. IT programs and CS departments (panel session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):417–418, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Knox:2006:MSU

[KDP06] Deborah L. Knox, Peter J. DePasquale, and Sarah M. Pulimood. A model for summer undergraduate research experiences in emerging technologies. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (1):214–218, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Koffman:2007:NPI

[KEK^{+}07] Elliot Koffman, Heidi Ellis, Charles Kelemen, Curt White, and Steven Wolfman. New paradigms for introductory computing courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):67–68, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kerren:2004:GME

[Ker04] Andreas Kerren. Generation as method for explorative learning in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (3):77–81, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kessler:2004:PAT

[Kes04] Christoph Kessler. A practical access to the theory of parallel algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):397–401, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kolaas:2003:IAM

[KF03] Øyvind Kolås and Ivar Farup. Increasing assignment motivation using a game AI tournament. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35 (3):269, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kurtz:2007:IUS

[KFE^{+}07] Barry L. Kurtz, James B. Fenwick, Jr., Christopher C. Ellsworth, Xiaohong Yuan, Adam Steele, and Xiaoping Jia. Inter-university software engineering using Web services. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):464–468, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Khuri:2000:IPL

[KH00] Sami Khuri and Hsiu-Chin Hsu. Interactive packages for learning image compression algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):73–76, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Khuri:2001:EEV

[KH01] Sami Khuri and Klaus Holzapfel. EVEGA: an educational visualization environment for graph algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):101–104, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kazemian:2005:STC

[KH05a] Fereydoun Kazemian and Trudy Howles. A software testing course for computer science majors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37 (4):50–53, December 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2005.bib.

Kolling:2005:GPI

[KH05b] Michael Kölling and Poul Henriksen. Game programming in introductory courses with direct state manipulation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):59–63, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Klimes:2006:FCS

[KH06] Cyril Klimes and Hashim Habiballa. Flexible computer science university studies. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):338, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kearse:2008:CSO

[KH08] Iretta B. Kearse and Charles R. Hardnett. Computer science olympiad: exploring computer science through competition. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):92–96, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Kamin:2008:SDT

[KHPC08] Sam Kamin, Michael Hines, Chad Peiper, and Boris Capitanu. A system for developing tablet PC applications for education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):422–426, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Khuri:2008:BTC

[Khu08] Sami Khuri. A bioinformatics track in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):508–512, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Kollanus:2008:TDD

[KI08] Sami Kollanus and Ville Isomöttönen. Test-driven development in education: experiences with critical viewpoints. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):124–127, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Kumar:2008:RSR

[KJGW08] Amruth Kumar, Patricia A. Joseph, Michael Goldweber, and Paul J. Wagner. Reviewing the SIGCSE reviewing process. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):84–89, June 2008. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2008.bib.

Kasyanov:2002:WBS

[KK02] Victor N. Kasyanov and Elena V. Kasianova. Web-based systems for supporting computer-science teaching and learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):238, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kujansuu:2003:CPI

[KK03] Esa Kujansuu and Marko Kulmala. Codewitz: producing interactive elearning material for beginners in programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):266, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kacmarcik:2009:ICP

[KK09] Gary Kacmarcik and Sylvie Giral Kacmarcik. Introducing computer programming via Gameboy advance homebrew. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):281–285, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Kaczmarczyk:2004:IWC

[KKLK04] Lisa Kaczmarczyk, Gerald Kruse, Dian Rae Lopez, and Deepak Kumar. Incorporating writing into the CS curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):179–180, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kumar:2006:NTP

[KKR06] Amruth Kumar, Deepak Kumar, and Ingrid Russell. Nontraditional projects in the undergraduate AI course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):479–480, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Klassner:2000:CWD

[Kla00] Frank Klassner. Can Web development courses avoid obsolescence? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):77–80, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Klassner:2002:CSL

[Kla02a] Frank Klassner. A case study of LEGO Mindstorms^{T M} suitability for artificial intelligence and robotics courses at the college level. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):8–12, March 2002.

CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Klawe:2002:GBC

[Kla02b] Maria M. Klawe. Girls, boys, and computers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):16–17, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.

math.utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/

Misc/DBLP/2002.bib. Women and Computing.

Klassner:2004:ELI

[Kla04] Frank Klassner. Enhancing Lisp instruction with RCXLisp and robotics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):214–218, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Klawe:2005:CIC

[Kla05a] Maria Klawe. Changing the image of computer science: a North American perspective in conversation with Europe. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):3, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Klawe:2005:INW

[Kla05b] Maria Klawe. Increasing the number of women majoring in computer science: what works? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1): 562, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Klee:2001:UTY

[Kle01] Karl J. Klee. Update on two-year college activities. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(2):23–24, June 2001. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2001.bib.

Klee:2002:GAD

[Kle02] Karl J. Klee. Guidelines for associate-degree programs in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):246, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Klinger:2005:SCS

[Kli05] William J. Klinger. Stanislavski and computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4):111–114, December 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2005.bib.

Krebs:2005:SBA

[KLOT05] Markus Krebs, Tobias Lauer, Thomas Ottmann, and Stephan Trahasch. Student-built algorithm visualizations for assessment: flexible generation, feedback and grading. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):281–285, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Korhonen:2000:ASA

[KM00] Ari Korhonen and Lauri Malmi. Algorithm simulation with automatic assessment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):160–163, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Korhonen:2002:DIM

[KMMS02] Ari Korhonen, Lauri Malmi, Pertti Myllyselkä, and Patrik Scheinin. Does it make a difference if students exercise on the Web or in the classroom? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):121– 124, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kurhila:2003:PPL

[KMN^{+}03] Jakko Kurhila, Miikka Miettinen, Petri Nokelainen, Patrik Floréen, and Henry Tirri. Peer-to-peer learning with openended writable Web. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):173–177, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Korhonen:2001:MCA

[KMS01] Ari Korhonen, Lauri Malmi, and Riku Saikkonen. Matrix — concept animation and algorithm simulation system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):180, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kimball:2008:CCW

[KMSB08] Aaron Kimball, Sierra Michels-Slettvet, and Christophe Bisciglia. Cluster computing for Web-scale data processing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):116–120, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Kasurinen:2009:EPK

[KN09] Jussi Kasurinen and Uolevi Nikula. Estimating programming knowledge with Bayesian knowledge tracing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):313–317, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Knox:2000:PPI

[Kno00a] Deborah L. Knox. A preview of the premier issue of JERIC. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(4):15–16, December 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Knox:2000:SEN

[Kno00b] Deborah L. Knox. SIGCSE endorses a new journal on educational resources in computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (2):17–18, June 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Knox:2001:MSR

[Kno01] Deborah L. Knox. Mentoring student research: award winning poster presentations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(2):15– 17, June 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2001.bib.

Knox:2002:CMR

[Kno02] Deborah Knox. CITIDEL: making resources available. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):225, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kessler:2002:MTC

[KNT02] Christoph Kessler and Simin Nadjm-Tehrani. Mid-term course evaluations with muddy cards. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):233, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Knuth:2003:BE

[Knu03] Donald E. Knuth. Bottom-up education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):2, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kolikant:2008:CSE

[Kol08a] YifatBen-David Kolikant. Computer-science education as a cultural encounter: a socio-cultural framework for articulating learning difficulties. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):291– 295, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Kolling:2008:GHG

[Köl08b] Michael Kölling. Greenfoot: a highly graphical IDE for learning object-oriented programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (3):327, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Koppelman:2001:TAE

[Kop01] Herman Koppelman. Teaching abstraction explicitly (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):191, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Koppelman:2007:ETS

[Kop07] Herman Koppelman. Exercises as a tool for sharing pedagogical knowledge. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):361, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Kosa:2000:DAD

[Kos00] Martha J. Kosa. Distributed algorithms in the discrete mathematics course (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3): 189–190, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kolikant:2004:COP

[KP04] Yifat Ben-David Kolikant and Sarah Pollack. Communityoriented pedagogy for in-service CS teacher training. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):191–195, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kenny:2005:ATD

[KP05] Claire Kenny and Claus Pahl. Automated tutoring for a database skills training environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):58–62, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Keenan:2006:LPP

[KPCM06] Frank Keenan, Sarah Powell, Gerry Coleman, and Kevin Mc-Daid. Learning project planning the agile way. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):324, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kurtz:2002:EIE

[KPN02] Barry L. Kurtz, Dee Parks, and Eric Nicholson. Effective Internet education: a progress report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (1):312–316, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Koldehofe:2000:LPS

[KPT00] Boris Koldehofe, Marina Papatriantafilou, and Philippas Tsigas. LYDIAN (poster session): an extensible educational animation environment for distributed algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):189, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Koldehofe:2003:ISV

[KPT03] Boris Koldehofe, Marina Papatriantafilou, and Philippas Tsigas. Integrating a simulation-visualisation environment in a basic distributed systems course: a case study using LY-DIAN. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):35–39, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kolling:2000:OFJ

[KR00] Michael Kölling and John Rosenberg. Objects first with Java and BlueJ (seminar session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):429, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kolling:2001:GTO

[KR01] Michael Kölling and John Rosenberg. Guidelines for teaching object orientation with Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3): 33–36, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kortsarts:2006:TPR

[KR06] Yana Kortsarts and Jeffrey Rufinus. Teaching the power of randomization using a simple game. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (1):460–463, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Knobelsdorf:2008:CPC

[KR08] Maria Knobelsdorf and Ralf Romeike. Creativity as a pathway to computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):286– 290, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Kreeger:2009:STM

[Kre09] Matthew Nicolas Kreeger. Security testing: mind the knowledge gap. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):99–102, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Karvonen:2001:VCM

[KRTT01] Antti Karvonen, Erkki Rautama, Jorma Tarhio, and Jari Turkia. Versatile concept map viewing on the Web. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):105–108, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kruse:2003:MNA

[Kru03] Gerald Kruse. “magic numbers” approach to introducing binary number representation in CS0. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35 (3):272, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kaplan:2000:CVT

[KS00] Alan Kaplan and Denise Shoup. CUPV — a visualization tool for generated parsers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):11– 15, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kuittinen:2004:TRV

[KS04] Marja Kuittinen and Jorma Sajaniemi. Teaching roles of variables in elementary programming courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):57–61, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kiesler:2002:PHC

[KSE02] Sara B. Kiesler, Lee S. Sproull, and Jacquelynne S. Eccles. Pool halls, chips, and war games: women in the culture of computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):159– 164, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/

2002.bib. Women and Computing.

Kumar:2005:EAC

[KSRD05] Amruth N. Kumar, Rose K. Shumba, Bina Ramamurthy, and Lawrence D’Antonio. Emerging areas in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):453–454, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Koldehofe:2001:UAI

[KT01] Boris Koldehofe and Philippas Tsigas. Using actors in an interactive animation in a graduate course on distributed system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):149–152, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kopec:2007:FCS

[KT07] Danny Kopec and Suzanne Tamang. Failures in complex systems: case studies, causes, and possible remedies. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):180–184, June 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Kelemen:2000:OCB

[KTH^{+}00] Charles Kelemen, Allen Tucker, Peter Henderson, Owen Astrachan, and Kim Bruce. Has our curriculum become mathphobic? (an American perspective). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (3):132–135, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Krane:2002:UVS

[KTKP02] Roseann Krane, Alfred Thompson, Harriette Kisilinsky, and Pat Phillips. Using Visual Studio. Net in the classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):247–248, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Kultur:2009:KPC

[Kül09] Can Kültür. Keeping the pace in CS-1 through the use of CMS. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):398, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Kumar:2000:DGP

[Kum00] Amruth Kumar. Dynamically generating problems on static scope. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):9–12, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kumar:2001:LIB

[Kum01] Amruth N. Kumar. Learning the interaction between pointers and scope in C++. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):45–48, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kumar:2004:WBT

[Kum04] Amruth Kumar. Web-based tutors for learning programming in C++/Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):266, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-

3927 (electronic). URL ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Compiler/java.
bib.

Kumar:2005:LRA

[Kum05a] Amruth N. Kumar. LEGO robots and AI. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):418, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kumar:2005:OTC

[Kum05b] Amruth N. Kumar. Online tutors for C++/Java programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):387, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kumar:2005:PPL

[Kum05c] Amruth N. Kumar. Projects in the programming languages course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):395, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kumar:2005:REE

[Kum05d] Amruth N. Kumar. Results from the evaluation of the effectiveness of an online tutor on expression evaluation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):216–220, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kumar:2008:EUP

[Kum08a] Amruth N. Kumar. The effect of using problem-solving software tutors on the self-confidence of female students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):523–527, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Kumar:2008:RHC

[Kum08b] Deepak Kumar. Reflections: historical cheesecakes …. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):11–13, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kumar:2009:DSA

[Kum09a] Amruth N. Kumar. Data space animation for learning the semantics of C++ pointers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):499– 503, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Kumar:2009:NCV

[Kum09b] Amruth N. Kumar. Need to consider variations within demographic groups when evaluating educational interventions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):176–180, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Kumar:2009:RBF

[Kum09c] Deepak Kumar. Reflections: Back to the future 100? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4):8–9, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kumar:2009:RRH

[Kum09d] Deepak Kumar. Reflections: rebuilding history… again! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):67–69, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kurkovsky:2008:FRI

[Kur08a] Stan Kurkovsky. Four roles of instructor in software engineering projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):354, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Kurmas:2008:ISP

[Kur08b] Zachary Kurmas. Improving student performance using automated testing of simulated digital logic circuits. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):265–270, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Kurkovsky:2009:EST

[Kur09a] Stan Kurkovsky. Engaging students through mobile game development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):44–48, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Kurkovsky:2009:MCM

[Kur09b] Stan Kurkovsky. Making the case for mobile game development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):401, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Kajaval:2000:PGI

[KV00] Jorma Kajaval and Rauno Varonen. The professional growth of ICT experts through progressive sandwich training (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):189, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Koppelman:2008:ESV

[KV08] Herman Koppelman and Harald Vranken. Experiences with a synchronous virtual classroom in distance education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):194–198, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Koppelman:2002:MFW

[KvD02] Herman Koppelman and Betsy van Dijk. A model for a flexible, Web-based course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):231, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Koppelman:2006:CRC

[KvD06] Herman Koppelman and Betsy van Dijk. Creating a realistic context for team projects in HCI. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (3):58–62, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kay:2005:EUC

[KvdHR05a] David G. Kay, André van der Hoek, and Debra J. Richardson. Extending undergraduate CS programs with informatics: emphasizing software and system design in context. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):367, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kay:2005:IFC

[KvdHR05b] David G. Kay, André van der Hoek, and Debra J. Richardson. Informatics: a focus on computer science in context. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):551–555, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Koppleman:2000:TPD

[KvdMvDvdV00] Herman Koppleman, Charles P. A. G. van der Mast, Elisabeth M. A. G. van Dijk, and Gerrit C. van der Veer. Team projects in distance education: a case in HCI design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):97–100, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kessler:2009:EAE

[KvLA09] Robert Kessler, Mark van Langeveld, and Roger Altizer. Entertainment arts and engineering (or how to fast track a new interdisciplinary program). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):539– 543, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Koffman:2001:SJP

[KW01] Elliot Koffman and Ursula Wolz. A simple Java package for GUI-like interactivity. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):11–15, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kheir:2007:IDS

[KW07] Richard Kheir and Thomas Way. Inclusion of deaf students in computer science classes using real-time speech transcription. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):261–265, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Katira:2004:UCS

[KWW^{+}04] Neha Katira, Laurie Williams, Eric Wiebe, Carol Miller, Suzanne Balik, and Ed Gehringer. On understanding compatibility of student pair programmers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (1):7–11, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kopec:2007:DSI

[KYC07] Danny Kopec, Gavriel Yarmish, and Patrick Cheung. A description and study of intermediate student programmer errors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):146–156, June 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

LaBarge:2006:MIL

[La 06] Bill La Barge. Make it look real, make it look cool. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):564, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lapidot:2007:ISS

[LA07] Tami Lapidot and Dan Aharoni. The Israeli summer seminars for CS leading teachers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3): 318, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Lapidot:2008:FCS

[LA08] Tami Lapidot and Dan Aharoni. On the frontier of computer science: Israeli summer seminars. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (4):72–74, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Labouseur:2009:BBO

[Lab09] Alan G. Labouseur. A browser-based operating systems project: JavaScript adventures in dinosaur slaying. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4):71–75, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lister:2004:MNS

[LAF^{+}04] Raymond Lister, Elizabeth S. Adams, Sue Fitzgerald, William Fone, John Hamer, Morten Lindholm, Robert Mc-Cartney, Jan Erik Moström, Kate Sanders, Otto Seppälä, Beth Simon, and Lynda Thomas. A multi-national study of reading and tracing skills in novice programmers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(4):119–150, December 2004. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2004.bib.

Larsson:2007:WIC

[LAKE07] Erik Larsson, Medhi Amirijoo, Daniel Karlsson, and Petru Eles. What impacts course evaluation? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):333, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Lahtinen:2005:SDN

[LAMJ05] Essi Lahtinen, Kirsti Ala-Mutka, and Hannu-Matti Järvinen. A study of the difficulties of novice programmers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):14–18, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lankewicsz:2001:URG

[Lan01] Linda Bright Lankewicsz. Undergraduate research in genetic algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):114–118, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lapidot:2002:SAP

[Lap02] Tami Lapidot. Self-assessment as a powerful learning experience. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):198, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lapidot:2007:SGC

[Lap07] Tami Lapidot. Supporting the growth of CS leading teachers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):327, September 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Linnell:2008:CCI

[LAP08] Natalie Linnell, Richard Anderson, and Jane Prey. Crosscultural issues in a tutored video instruction course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):53–57, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Larkworthy:2005:ELF

[Lar05] Thomas Larkworthy. Evolution lab: the flexible user friendly robot research tool. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):409, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Larson:2006:UCS

[Lar06] Eric Larson. An undergraduate course on software bug detection tools and techniques. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):249– 253, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Last:2003:PSD

[Las03] Mary Z. Last. Preparing students for distributed teamwork. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):239, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lasserre:2009:ATB

[Las09] Patricia Lasserre. Adaptation of team-based learning on a first term programming class. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3): 186–190, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Lauer:2006:LIA

[Lau06] Tobias Lauer. Learner interaction with algorithm visualizations: viewing vs. changing vs. constructing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):202–206, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lau:2007:ALS

[Lau07] Kung-Kiu Lau. Active learning sheets for a beginner’s course on reasoning about imperative programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):198–202, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lauer:2008:RRE

[Lau08] Tobias Lauer. Reevaluating and refining the engagement taxonomy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):355, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Lawhead:2001:ERP

[Law01] Pamela Lawhead. Events robots and programming using Legos in CS1 (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):183, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lawrence:2005:XXX

[Law05] Richard Lawrence. [4]cite. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):410, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Laxer:2001:TCS

[Lax01] Cary Laxer. Treating computer science as science as: an experiment with sorting (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):189, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Laxer:2002:EST

[Lax02] Cary Laxer. Evaluating student team project experiences. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):198, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lazowska:2002:PMT

[Laz02a] Ed Lazowska. Pale and male: 19 th century design in a 21st century world. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):11–12, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Women and Computing.

Lazowska:2002:PMC

[Laz02b] Edward D. Lazowska. Pale and male: 19^{th} century design in a 21^{st} century world. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):11– 12, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/

2002.bib. Women and Computing.

Lazowska:2008:CSP

[Laz08] Ed Lazowska. Computer science: past, present, and future. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):549, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Lee:2000:FFD

[LB00] John A. N. Lee and Kevin Bowyer. Future faculty development seminar in ethics, social impact and alternative teaching strategies (seminar session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):426, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lawson:2008:UII

[LB08a] Barry Lawson and Lewis Barnett. Using iPodLinux in an introductory OS course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):182–186, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Lister:2008:CAS

[LB08b] Raymond Lister and Ilona Box. A citation analysis of the SIGCSE 2007 proceedings. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):476– 480, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Levy:2007:WWH

[LBA07] Ronit Ben-Bassat Levy and Mordechai Ben-Ari. We work so hard and they don’t use it: acceptance of software tools by teachers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):246–250, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Levy:2008:PBC

[LBA08] Ronit Ben-Bassat Levy and Mordechai Ben-Ari. Perceived behavior control and its influence on the adoption of software tools. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):169–173, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Lang:2001:OOP

[LBB^{+}01] Joseph E. Lang, Brian R. Bogovich, Sean C. Barry, Brian G. Durkin, Michael R. Katchmar, Jonathan H. Kelly, J. Michael McCollum, and Michael Potts. Object-oriented programming and design patterns. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(4):68–70, December 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2001.bib.

Leventhal:2004:TPU

[LBC04] L. M. Leventhal, J. Barnes, and J. Chao. Term project user interface specifications in a usability engineering course: challenges and suggestions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):41–45, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lister:2006:RPO

[LBC^{+}06] Raymond Lister, Anders Berglund, Tony Clear, Joe Bergin, Kathy Garvin-Doxas, Brian Hanks, Lewis E. Hitchner, Andrew Luxton-Reilly, Kate Sanders, Carsten Schulte, and Jacqueline L. Whalley. Research perspectives on the objectsearly debate. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):146–165, December 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2006.bib.

Leach:2008:CSR

[LBK08] Ronald J. Leach, Legand L. Burge, and Harry N. Keeling. Can students reengineer? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):102– 106, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Lister:2004:DVT

[LBM^{+}04] Raymond Lister, Ilona Box, Briana Morrison, Josh Tenenberg, and D. Suzanne Westbrook. The dimensions of variation in the teaching of data structures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):92–96, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lancaster:2001:TEF

[LC01] Thomas Lancaster and Fintan Culwin. Towards an error free plagiarism detection process. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3): 57–60, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Logan:2005:TSH

[LC05] Patricia Y. Logan and Allen Clarkson. Teaching students to hack: curriculum issues in information security. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):157–161, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lorenzen:2006:MPC

[LC06] Torben Lorenzen and Hang-Ling Chang. MasterMind^{T M}: a predictor of computer programming aptitude. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):69–71, June 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.

math.utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2006.bib.

Ladner:2008:CSE

[LC08] Richard E. Ladner and Dan Comden. Computer science for everyone: making your computing classes and departments accessible. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):547–548, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Lass:2003:TTL

[LCB^{+}03] Robert N. Lass, Christopher D. Cera, Nathaniel T. Bomberger, Bruce Char, Jeffrey L. Popyack, Nira Herrmann, and Paul Zoski. Tools and techniques for large scale grading using Web-based commercial off-the-shelf software. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):168–172, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Liu:2007:BFS

[LCG07] Haifeng Liu, Xianglan Chen, and Yuchang Gong. BabyOS: a fresh start. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):566–570, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lui:2008:LSP

[LCL08] Andrew K. Lui, Yannie H. Y. Cheung, and Siu Cheung Li. Leveraging students’ programming laboratory work as worked examples. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):69–73, June 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2008.bib.

Lewis:2005:EID

[LCPQR05] Tracy L. Lewis, J. D. Chase, Manuel A. Pérez-Quiñones, and Mary Beth Rosson. The effects of individual differences on CS2 course performance across universities. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):426–430, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lister:2009:NOD

[LCS^{+}09] Raymond Lister, Tony Clear, Simon, Dennis J. Bouvier, Paul Carter, Anna Eckerdal, Jana Jacková, Mike Lopez, Robert McCartney, Phil Robbins, Otto Seppälä, and Errol Thompson. Naturally occurring data as research instrument: analyzing examination responses to study the novice programmer. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4):156–173, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Layman:2006:PTL

[LCW06] Lucas Layman, Travis Cornwell, and Laurie Williams. Personality types, learning styles, and an agile approach to software engineering education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):428– 432, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Linington:2001:PPE

[LD01] Janet Linington and Mark Dixon. Picture program execution (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):175, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

LeBlanc:2004:BCC

[LD04] Mark D. LeBlanc and Betsey D. Dyer. Bioinformatics and computing curricula 2001: why computer science is well positioned in a post-genomic world. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(4):64–68, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2004.bib.

Last:2000:ISF

[LDA^{+}00] Mary Z. Last, Mats Daniels, Vicki L. Almstrum, Carl Erickson, and Bruce Klein. An international student/faculty collaboration: the runestone project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (3):128–131, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lawhead:2003:LJP

[LDB^{+}03a] Pam Lawhead, Michaele E. Duncan, Constance G. Bland, Michael Goldweber, Madeleine Schep, and David J. Barnes. Legos, Java, programming assignments and CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):47–48, January 2003. CODEN SIGSD3. ISBN 1-58113-648-X. ISSN 0097-8418 (print), 2331-

3927 (electronic). URL ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Compiler/java.
bib;   ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2003.bib.

Lawhead:2003:RMT

[LDB^{+}03b] Pamela B. Lawhead, Michaele E. Duncan, Constance G. Bland, Michael Goldweber, Madeleine Schep, David J. Barnes, and Ralph G. Hollingsworth. A road map for teaching introductory programming using LEGO^{T M} Mindstorms robots. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):191– 201, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2003.bib;   ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Techreports/UKent.
bib;   http://www.cs.kent.ac.uk/pubs/2003/1666.

Last:2002:LSC

[LDHW02] Mary Z. Last, Mats Daniels, Martha L. Hause, and Mark R. Woodroffe. Learning from students: continuous improvement in international collaboration. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3): 136–140, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lindquist:2007:EPM

[LDK^{+}07] David Lindquist, Tamara Denning, Michael Kelly, Roshni Malani, William G. Griswold, and Beth Simon. Exploring the potential of mobile phones for active learning in the classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):384–388, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Laender:2008:ARE

[LdLM^{+}08] Alberto H. F. Laender, Carlos José Pereira de Lucena, José Carlos Maldonado, Edmundo de Souza e Silva, and Nivio Ziviani. Assessing the research and education quality of the top Brazilian Computer Science graduate programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):135–145, June 2008. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2008.bib.

Leutenegger:2007:GFA

[LE07] Scott Leutenegger and Jeffrey Edgington. A games first approach to teaching introductory programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):115–118, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Learmonth:2001:FDI

[Lea01] Rod Learmonth. Flexible delivery of information systems as a core MBA subject. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):29–32, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Leach:2008:AAA

[Lea08] Ronald J. Leach. Analysis of ABET accreditation as a software process. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):356, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Lee:2000:EHM

[Lee00a] John A. N. Lee. Emulators of “historic machines”. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(2):10–11, June 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lee:2000:HCS

[Lee00b] John A. N. Lee. History in computer science education: across the curriculum initiatives. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (4):9–10, December 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lee:2001:HCS

[Lee01a] John A. N. Lee. History in computer science education: across the curriculum initiatives. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33 (2):8, June 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2001.bib.

Lee:2001:TLC

[Lee01b] John A. N. Lee. Teaching and learning in the 21st century: the development of “future CS faculty”. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(2):25–30, June 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2001.bib.

Lee:2002:EAC

[Lee02a] J. A. N. Lee. The emporium approach to computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):203–207, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lee:2002:ICR

[Lee02b] J. A. N. Lee. Internationalization of the curriculum report of a project within computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (1):68–72, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Lee:2002:WDH

[Lee02c] J. A. N. Lee. Where did history go? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (4):11–12, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2002.bib.

Lee:2002:WBW

[Lee02d] John A. N. (JAN) Lee. Well behaved women rarely make history! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):14–15, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/

2002.bib. Women and Computing.

Lee:2003:MWS

[Lee03a] Arthur H. Lee. A manageable Web software architecture: searching for simplicity. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):229– 233, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lee:2003:LL

[Lee03b] John A. N. Lee. “Lee’s law”. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):10– 12, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2003.bib.

Leska:2003:LDG

[Les03] Chuck Leska. Learning to develop GUIs in Java using closed labs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):228, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Compiler/java.bib.

Leska:2004:IUP

[Les04] Chuck Leska. Introducing undergraduates to programming using robots in the general education curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):263, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Levitin:2000:DAA

[Lev00] Anany Levitin. Design and analysis of algorithms reconsidered. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):16–20, March 2000. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Levitin:2005:APA

[Lev05] Anany Levitin. Analyze that: puzzles and analysis of algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):171–175, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lewis:2000:MAO

[Lew00] John Lewis. Myths about object-orientation and its pedagogy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):245–249, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lewis:2002:RTI

[Lew02] Jerome L. Lewis. A reliable test for inclusion of a point in a polygon. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):81–84, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2002.bib.

Lewis:2007:ABA

[Lew07a] Clayton Lewis. Attitudes and beliefs about computer science among students and faculty. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):37– 41, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2007.bib.

Lewis:2007:PWW

[Lew07b] Tracy L. Lewis. Are pictures worth 1000 words? the use of affective imagery surveys to capture student perceptions of the computing discipline. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3): 312, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Lu:2009:TAC

[LF09] James J. Lu and George H. L. Fletcher. Thinking about computational thinking. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):260–264, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Laverell:2008:IIT

[LFG08] W. David Laverell, Zongming Fei, and James N. Griffioen. Isn’t it time you had an Emulab? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (1):246–250, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Lidtke:2002:PIS

[LFGY02] Doris K. Lidtke, David Feinstein, John Gorgone, and Gayle Yaverbaum. Preparing for Information Systems accreditation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):169–170, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Leidig:2006:UCB

[LFL06] Paul M. Leidig, Roger Ferguson, and Jonathan Leidig. The use of community-based non-profit organizations in information systems capstone projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3): 148–152, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lister:2009:FER

[LFT09] Raymond Lister, Colin Fidge, and Donna Teague. Further evidence of a relationship between explaining, tracing and writing skills in introductory programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):161–165, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Little:2001:ICI

[LGA^{+}01] Joyce Currie Little, Mary J. Granger, Elizabeth S. Adams, Jaana Holvikivi, Susan K. Lippert, Henry MacKay Walker, and Alison Young. Integrating cultural issues into the computer and information technology curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(2):136–154, June 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2001.bib.

Lorenzen:2002:CCW

[LH02] Torben Lorenzen and Ward Heilman. CS1 and CS2: Write computer games in Java! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (4):99–100, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Compiler/java.bib;   ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/2002.
bib.

Lapidot:2003:MTC

[LH03] Tami Lapidot and Orit Hazzan. Methods of teaching a computer science course for prospective teachers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):29–34, December 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2003.bib.

Lapidot:2005:SDM

[LH05] Tami Lapidot and Orit Hazzan. Song debugging: merging content and pedagogy in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4):79–83, December 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2005.bib.

Linos:2003:SLP

[LHL03] Panagiotis K. Linos, Stephanie Herman, and Julie Lally. A service-learning program for computer science and software engineering. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):30–34, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lopez-Herrejon:2004:UIT

[LHS04] Roberto E. Lopez-Herrejon and Morrie Schulman. Using interactive technology in a short Java course: an experience report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):203–207, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Leonard:2009:IRF

[LHS09] Dana P. Leonard, Jason O. Hallstrom, and Murali Sitaraman. Injecting rapid feedback and collaborative reasoning in teaching specifications. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):524– 528, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Liew:2005:TSD

[Lie05] C. W. Liew. Teaching software development skills early in the curriculum through software engineering. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):133–137, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Liffick:2004:ATP

[Lif04a] Blaise W. Liffick. An assistive technology project for an HCI course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):273, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Liffick:2004:IAT

[Lif04b] Blaise W. Liffick. Introducing assistive technology in an HCI course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):232, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lim:2007:TSR

[Lim07] Darren Lim. Taking students out for a ride: using a board game to teach graph theory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):367– 371, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lindholm:2005:DOU

[Lin05] Morten Lindholm. Development of object-understanding among students in the humanities. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37 (3):382, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lincke:2007:NSA

[Lin07] Susan J. Lincke. Network security auditing as a communitybased learning project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):476–480, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lischner:2001:ESL

[Lis01a] Ray Lischner. Explorations: structured labs for firsttime programmers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):154–158, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lister:2001:OOA

[Lis01b] Raymond Lister. Objectives and objective assessment in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):292–296, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lister:2003:FOT

[Lis03a] Raymond Lister. The five orders of teaching ignorance. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):16–17, December 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

Lister:2003:RMR

[Lis03b] Raymond Lister. A research manifesto, and the relevance of phenomenography. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):15– 16, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2003.bib.

Lister:2004:BRC

[Lis04a] Raymond Lister. Book review: computer science education research. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(2):17–18, June 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2004.bib.

Lister:2004:CRM

[Lis04b] Raymond Lister. A clandestine religious meeting. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(4):16–17, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2004.bib.

Lister:2005:GC

[Lis05a] Raymond Lister. Grand challenges. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37 (2):14–15, June 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2005.bib.

Lister:2005:MMP

[Lis05b] Raymond Lister. Mixed methods: positivists are from Mars, constructivists are from Venus. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4): 18–19, December 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2005.bib.

Lister:2006:CMI

[Lis06a] Raymond Lister. Call me Ishmael: Charles Dickens meets Moby Book. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):11–13, June 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2006.bib.

Lister:2006:ORF

[Lis06b] Raymond Lister. One room, four meetings. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):11–13, December 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2006.bib.

Lister:2007:RTC

[Lis07a] Raymond Lister. The Randolph thesis: CSEd research at the crossroads. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):16–18, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2007.bib.

Lister:2007:SET

[Lis07b] Raymond Lister. Strangers, enemies, terms and taxonomies. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):12–13, June 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Lister:2008:CRW

[Lis08a] Raymond Lister. CS research: We are what we cite — so where are we? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):16–18, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lister:2008:OG

[Lis08b] Raymond Lister. The originality glut. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):14–15, June 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2008.bib.

Lister:2009:CRB

[Lis09a] Raymond Lister. CS research: Book burning, naturally occurring data, and the stages of pedagogic grief. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4):13–14, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lister:2009:CRR

[Lis09b] Raymond Lister. CS research: rules for sustaining the discourse — engage! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):72–74, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Liu:2006:SPD

[Liu06] Chang Liu. Software project demonstrations as not only an assessment tool but also a learning tool. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):423–427, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Liu:2008:UOS

[Liu08] Peter L. Liu. Using open-source robocode as a Java programming assignment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):63–67, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lucas:2008:ITJ

[LJ08] Joan M. Lucas and Jonathan Jarvis. Incorporating transformations into JFLAP for enhanced understanding of automata. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):14–18, March 2008. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Lewandowski:2005:FCI

[LJG05] Gary Lewandowski, Elizabeth Johnson, and Michael Goldweber. Fostering a creative interest in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):535–539, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lim:2005:IWS

[LJM05] Billy B. L. Lim, Chu Jong, and Pruthikrai Mahatanankoon. On integrating Web services from the ground up into CS1/ CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):241–245, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lahtinen:2007:TPV

[LJMV07] Essi Lahtinen, Hannu-Matti Järvinen, and Suvi Melakoski-Vistbacka. Targeting program visualizations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):256–260, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Lidtke:2000:PIS

[LKGY00] Doris K. Lidtke, Willis King, John Gorgone, and Gayle Yaverbaum. Proposed information systems accreditation criteria (panel session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):409–410, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lui:2004:SWP

[LKPC04] Andrew K. Lui, Reggie Kwan, Maria Poon, and Yannie H. Y. Cheung. Saving weak programming students: applying constructivism in a first programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(2):72–76, June 2004. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2004.bib.

Levy:2000:RSA

[LL00] Dalit Levy and Tami Lapidot. Recursively speaking: analyzing students’ discourse of recursive phenomena. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):315–319, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Levy:2002:STP

[LL02] Dalit Levy and Tami Lapidot. Shared terminology, private syntax: the case of recursive descriptions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):89–93, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lister:2003:IPC

[LL03] Raymond Lister and John Leaney. Introductory programming, criterion-referencing, and bloom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):143–147, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Larsson:2004:SOE

[LL04] Erik Larsson and Anders Larsson. Student-oriented examination in a computer architecture course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):245, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

LeBlanc:2006:DPC

[LL06a] Mark D. LeBlanc and Rochelle Leibowitz. Discrete partnership: a case for a full year of discrete math. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):313–317, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lindqvist:2006:VOT

[LL06b] Janne Lindqvist and Sanna Liimatainen. VERKKOKE: online teaching environment for telecommunications software and routing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):319, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lukins:2002:TPP

[LLB02] Stacy Lukins, Alan Levicki, and Jennifer Burg. A tutorial program for propositional logic with human/computer interactive learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):381–385, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Lopez:2006:MDL

[LLLC06] Antonio M. Lopez, Jr., Robert Lent, Frederick G. Lopez, and Madonna G. Constantine. A multi-disciplinary look at the computing disciplines. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):264–265, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lawson:2004:ITA

[LLP04] Eydie Lawson, Doris K. Lidtke, and Barbara Price. Information technology accreditation activities. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):49–50, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lewis:2009:IPC

[LLP^{+}09] Mark Lewis, Scott Leutenegger Leutenegger, Michael Panitz, Kelvin Sung, and Scott A. Wallace. Introductory programming courses and computer games. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (1):204–205, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Lynch:2003:WCS

[LM03] Kathy Lynch and Selby Markham. The winds of change: students’ comfort level in different learning environments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):70–73, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lorenzen:2005:TGC

[LM05] Torben Lorenzen and Joseph Michaud. Teach 2D graphics concepts with an OpenGL graphing program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):54–55, June 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.

math.utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2005.bib.

Lewis:2006:GGD

[LM06] Mark C. Lewis and Berna Massingill. Graphical game development in CS2: a flexible infrastructure for a semester long project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):505–509, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lavonen:2001:UCS

[LML^{+}01] Jari Lavonen, Veijo Meisalo, Matti Lattu, Liisa Leinonen, and Tadeusz Wilusz. Using computers in science and technology education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(2):127–135, June 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2001.bib.

Lang:2007:SFI

[LML07] Catherine Lang, Judy McKay, and Sue Lewis. Seven factors that influence ICT student achievement. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):221–225, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Lemmen:2000:IUP

[LMS00] Karel Lemmen, Fred Mulder, and Wim Smit. An innovative university program on management and ICT. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):55–59, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lau:2009:LPT

[LNCC09] Winnie W. Y. Lau, Grace Ngai, Stephen C. F. Chan, and Joey C. Y. Cheung. Learning programming through fashion and design: a pilot summer course in wearable computing for middle school students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):504– 508, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Lauwers:2009:CDD

[LNH09] Tom Lauwers, Illah Nourbakhsh, and Emily Hamner. CSbots: design and deployment of a robot designed for the CS1 classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):428–432, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Lucas:2003:VGC

[LNR03] Jeff Lucas, Thomas L. Naps, and Guido Rößling. Visual-Graph: a graph class designed for both undergraduate students and educators. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):167–171, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ludi:2005:ISE

[LNR05] Stephanie Ludi, Swaminathan Natarajan, and Thomas Reichlmayr. An introductory software engineering course that facilitates active learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):302– 306, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lomako:2007:LCP

[Lom07] Gennady Lomako. Learning computer programming and role of errors in design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):142– 145, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2007.bib.

Long:2008:SAR

[Lon08] Philip D. Long. Scalable apprenticeships: reconnecting students through technology. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):3–4, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Liccardi:2007:RSN

[LOP^{+}07] Ilaria Liccardi, Asma Ounnas, Reena Pau, Elizabeth Massey, Päivi Kinnunen, Sarah Lewthwaite, Marie-Anne Midy, and Chandan Sarkar. The role of social networks in students’ learning experiences. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):224–237, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2007.bib.

Lorenzen:2000:PYE

[Lor00] Torben Lorenzen. Publish your Excel grade book on the Web. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(2):77–78, June 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lorenzen:2003:RTP

[Lor03] Torben Lorenzen. The reverse trace: a programming tool. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):115–116, June 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

Levitin:2002:UPT

[LP02] Anany Levitin and Mary-Angela Papalaskari. Using puzzles in teaching algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):292– 296, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Lanari:2001:IPS

[LR01] David Lanari and Stefano Roccetti. e-IMC (poster session): an authoring tool for humanistic teachers aimed to develop and distribute customized instructional courseware. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):183, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Leska:2004:LCC

[LR04] Chuck Leska and John Rabung. Learning O-O concepts in CS I using game projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):237, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Loftus:2005:EPP

[LR05] Chris Loftus and Mark Ratcliffe. Extreme programming promotes extreme learning? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):311– 315, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Leidl:2007:HWF

[LR07] Martin Leidl and Guido Rößling. How will future learning work in the third dimension? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3): 329, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Ludi:2008:DIO

[LR08] Stephanie Ann Ludi and Thomas Reichlmayr. Developing inclusive outreach activities for students with visual impairments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):439–443, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Luxton-Reilly:2009:SFI

[LRD09] Andrew Luxton-Reilly and Paul Denny. A simple framework for interactive games in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):216– 220, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Linckels:2007:VTT

[LRKM07] Serge Linckels, Stephan Repp, Naouel Karam, and Christoph Meinel. The virtual tele-tASK professor: semantic search in recorded lectures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):50–54, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lewis:2004:WDE

[LRPQ04] Tracy L. Lewis, Mary Beth Rosson, and Manuel A. Pérez-Quiñones. What do the experts say?: teaching introductory design from an expert’s perspective. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (1):296–300, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lowther:2001:CGU

[LS01] John L. Lowther and Ching-Kuang Shene. Computing with geometry as an undergraduate course: a three-year experience. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):119–123, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lopez:2002:AAW

[LS02] Antonio M. Lopez, Jr. and Lisa J. Schulte. African American women in the computing sciences: a group to be studied. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):87–90, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Lowther:2003:TBS

[LS03] John Lowther and Ching-Kuang Shene. Teaching B-splines is not difficult! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):381–385, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

LeBlanc:2004:TCS

[LS04a] Cathie LeBlanc and Evelyn Stiller. Teaching computer security at a small college. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):407– 411, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

LopezHerrejon:2004:UIT

[LS04b] R. E. Lopez Herrejon and M. Schulman. Using interactive technology in a short Java course: An experience report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):203–207, September 2004. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Compiler/java.bib.

Lewis:2005:CSD

[LS05] Tracy L. Lewis and Wanda J. Smith. The computer science debate: it’s a matter of perspective. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37 (2):80–84, June 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2005.bib.

Lorenzen:2006:TGU

[LS06] Torben Lorenzen and Abdul Sattar. Teach graphics using excel in place of a graphing calculator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):61–63, December 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2006.bib.

Linhoff:2008:TGP

[LS08a] Joe Linhoff and Amber Settle. Teaching game programming using XNA. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):250–254, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Lorenzen:2008:HCO

[LS08b] Torben Lorenzen and Abdul Sattar. How to create an online Internet course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):100–102, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lorenzen:2008:OFU

[LS08c] Torben Lorenzen and Abdul Sattar. Objects first using Alice to introduce object constructs in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):62–64, June 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2008.bib.

Lopez:2005:CSG

[LSG05] Antonio M. Lopez, Jr., Lisa J. Schulte, and Marguerite S. Giguette. Climbing onto the shoulders of giants. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):401–405, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lorenzen:2007:NST

[LSGB07] Torben Lorenzen, John Santore, David Glassman, and Juozas Baltikauskas. No slacker on team programming projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):117–118, December 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Lidtke:2000:WNA

[LSMB00] Doris K. Lidtke, Lee Saperstein, Kenneth Martin, and Della Bonnette. What’s new with ABET/CSAB integration (panel session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):413, March 2000. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lister:2006:SFT

[LST^{+}06] Raymond Lister, Beth Simon, Errol Thompson, Jacqueline L. Whalley, and Christine Prasad. Not seeing the forest for the trees: novice programmers and the SOLO taxonomy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):118–122, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Liu:2009:TFM

[LTHN09] Shaoying Liu, Kazuhiro Takahashi, Toshinori Hayashi, and Toshihiro Nakayama. Teaching formal methods in the context of software engineering. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):17– 23, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ludi:2005:ALA

[Lud05] Stephanie Ludi. Active-learning activities that introduce students to software engineering fundamentals. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):128–132, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Luque:2005:ERW

[Luq05] Emilio Luque. Is evolution or revolution the way for improving the teaching methodology in computer science? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):2, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lane:2003:CPP

[LV03] H. Chad Lane and Kurt VanLehn. Coached program planning: dialogue-based support for novice program design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):148–152, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lane:2005:IBS

[LV05] H. Chad Lane and Kurt VanLehn. Intention-based scoring: an approach to measuring success at solving the composition problem. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):373–377, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ladner:2008:GLO

[LV08] Richard E. Ladner and Tammy VanDeGrift. The game of life: an outreach model for high school students with disabilities. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):296–297, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Li:2000:MNT

[LW00] Ming-Sun Li and Marcus Wright. On a modified nine-tails problem. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(2):82, June 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Liccardi:2005:UDD

[LW05] Ilaria Liccardi and Su White. Understanding disciplinary differences: an insight into selecting effective e-learning approaches. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):411, September 2005.

CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Levine:2006:XPA

[LW06] David B. Levine and Henry M. Walker. XP practices applied to grading. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):173–177, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Lancor:2007:UGH

[LW07] Lisa Lancor and Robert Workman. Using Google hacking to enhance defense strategies. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):491– 495, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Layman:2007:NSM

[LWS07] Lucas Layman, Laurie Williams, and Kelli Slaten. Note to self: make assignments meaningful. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (1):459–463, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Liffick:2007:YCT

[LZ07] Blaise W. Liffick and Gary Zoppetti. You can take it with you: profile transportability. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3): 106–110, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Lin:2009:ECS

[LZBO09] Chi-Cheng Lin, Mingrui Zhang, Barbara Beck, and Gayle Olsen. Embedding computer science concepts in K-12 science curricula. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):539–543, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Lopez:2008:CRG

[LZL08] Antonio M. Lopez, Jr., Kun Zhang, and Frederick G. Lopez. Cultural representations of gender among U.S. computer science undergraduates: statistical and data mining results. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):407–411, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

McCaughey:2001:LTS

[MA01] Aine McCaughey and Sylvia Alexander. The learning and teaching support network promoting best practice in the information and computer science academic community (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):175, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCracken:2001:MMI

[MAD^{+}01a] M. McCracken, V. Almstrum, D. Diaz, M. Guzdial, D. Hagan, Y. B.-D. Kolikant, C. Laxer, L. Thomas, I. Utting, and T. Wilusz. A multinational, multi-institutional study of assessment of programming skills of first-year CS students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(4):125–140, December 2001. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Techreports/UKent.bib;
 http://www.cs.kent.ac.uk/pubs/2001/1365.

McCracken:2001:MNM

[MAD^{+}01b] Michael McCracken, Vicki L. Almstrum, Danny Diaz, Mark Guzdial, Dianne Hagan, Yifat Ben-David Kolikant, Cary Laxer, Lynda Thomas, Ian Utting, and Tadeusz Wilusz. A multi-national, multi-institutional study of assessment of programming skills of first-year CS students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(4):125–180, December 2001. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2001.bib.

Mendes:2005:IPP

[MAFLR05] Emilia Mendes, Lubna Basil Al-Fakhri, and Andrew Luxton-Reilly. Investigating pair-programming in a 2nd-year software development and design computer science course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):296–300, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mendes:2006:REP

[MAFLR06] Emilia Mendes, Lubna Al-Fakhri, and Andrew Luxton-Reilly. A replicated experiment of pair-programming in a 2nd-year software development and design computer science course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):108–112, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Masuck:2008:AFM

[MAFO08] Carol Masuck, Jim Alves-Foss, and Paul W. Oman. Analysis of fault models for student use. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2): 79–83, June 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2008.bib.

Mache:2008:GCU

[MAFW08] Jens Mache, Amy Apon, Thomas Feilhauer, and Barry Wilkinson. Grid computing at the undergraduate level: can we do it? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):381–382, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Maibaum:2009:FMV

[Mai09] Tom Maibaum. Formal methods versus engineering. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):6–12, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Malan:2007:PCS

[Mal07] David J. Malan. Podcasting computer science E-1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):389–393, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Malan:2009:VOH

[Mal09] David J. Malan. Virtualizing office hours in CS 50. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):303–307, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Marques:2005:UOI

[MAMC05] Nuno C. Marques, Francisco Azevedo, Carmen Morgado, and Jorge F. Custódio. Using Octave to introduce programming to technical science students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):198– 202, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Manne:2000:CCP

[Man00] Fredrik Manne. Competing in computing (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):190, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Manolopoulos:2002:BCC

[Man02] Yannis Manolopoulos. Binomial coefficient computation: recursion or iteration? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):65–67, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2002.bib.

Mander:2004:BF

[Man04] Keith Mander. Back to the future? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (3):2, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Manolopoulos:2005:NRC

[Man05] Yannis Manolopoulos. On the number of recursive calls of recursive functions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):61–64, June 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2005.bib.

Manaris:2007:DCE

[Man07] Bill Z. Manaris. Dropping CS enrollments: or the emperor’s new clothes? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):6–10, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2007.bib.

Maniccam:2009:SSU

[Man09] S. Maniccam. Sorting and searching using Lisp, functional programming, and recursion. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4): 53–56, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Martin:2000:DPB

[Mar00a] C. Dianne Martin. Debunking the Puppy Baron culture. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(2):7, June 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Martin:2000:MDS

[Mar00b] C. Dianne Martin. More on the “dark side” of computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(4):8–9, December 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Martin:2001:ECW

[Mar01a] C. Dianne Martin. .Ethics @ .coms: why Internet startups need ethics codes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(2):7– 8, June 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2001.bib.

Martin:2001:PSI

[Mar01b] C. Dianne Martin. PKAL summer institute 2001: justin-time computer education for the 21st century. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(4):5–6, December 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2001.bib.

Martin:2002:CCR

[Mar02] C. Dianne Martin. Computing curricula 2001: reverse engineering a computer science curriculum (part 1). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):10–11, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2002.bib.

Martin:2003:CCR

[Mar03] C. Dianne Martin. Computing curricula 2001: reverse engineering a computer science curriculum (part 2). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):9–10, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2003.bib.

Martin:2004:DCS

[Mar04a] C. Dianne Martin. Draw a computer scientist. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(4):11–12, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2004.bib.

Martin:2004:LCI

[Mar04b] C. Dianne Martin. Lessons from Chernobyl for IT. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(2):11, June 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2004.bib.

Marcelino:2005:LRS

[Mar05a] M. J. Marcelino. Learning repetition structures in programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):351, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Marion:2005:SRS

[Mar05b] Bill Marion. Status report on the SIGCSE committee on the implementation of a discrete mathematics course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):194–195, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Martin:2005:PSW

[Mar05c] C. Dianne Martin. Preparing the STEM workforce of the 21^{s}t century: report on a report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):10– 11, June 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2005.bib.

Martin:2005:RVP

[Mar05d] C. Dianne Martin. Removing the veil: personal reflections on educating women in Dubai. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4): 30–33, December 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2005.bib.

Marion:2006:FOR

[Mar06] Bill Marion. Final oral report of the SIGCSE committee on the implementation of a discrete mathematics course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):268–269, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Martin:2007:BEY

[Mar07a] C. Dianne Martin. Blogger ethics and YouTube common sense. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):11–12, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Martin:2007:LTE

[Mar07b] C. Dianne Martin. Leadership, teamwork, and ethics in the development of IT professionals. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (2):8–9, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2007.bib.

Martin:2007:SES

[Mar07c] Christopher James Martin. Scribbles: an exploratory study of sketch based support for early collaborative object oriented design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):286–290, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

McMaster:2007:DMP

[MAR07d] Kirby McMaster, Nicole Anderson, and Brian Rague. Discrete math with programming: better together. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):100–104, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Martin:2008:THR

[Mar08] C. Dianne Martin. Taking the high road: Blogging for votes: the ethics of Internet campaigning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (4):10–11, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Martin:2009:THR

[Mar09a] C. Dianne Martin. Taking the high road: ethics on the run: the principle of the ordinary person. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (2):66–67, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Martin:2009:CLS

[Mar09b] Javier Garcia Martin. Cooperative learning to support the lacks of PBL. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):343, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Mason:2003:CCH

[Mas03] John Mason. Comments considered harmful. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):120–122, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2003.bib.

Mason:2004:TAS

[Mas04] John Mason. Teaching by analogy: the switch statement. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(4):105–107, December 2004. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2004.bib.

Massarelli:2005:DAP

[Mas05] Riccardo Massarelli. The difficult art of pruning in favour of creative thinking. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):368, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mateti:2003:LBC

[Mat03] Prabhaker Mateti. A laboratory-based course on Internet security. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):252–256, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Matsuura:2006:EMP

[Mat06] Saeko Matsuura. An evaluation method of project based learning on software development experiment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):163–167, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Maurer:2002:CPL

[Mau02] W. Douglas Maurer. The comparative programming languages course: a new chain of development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):336–340, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Mayer:2008:IDS

[May08] Marissa Mayer. Innovation, design, and simplicity at Google. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):199, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Moritz:2005:DFC

[MB05] Sally H. Moritz and Glenn D. Blank. A design-first curriculum for teaching Java in a CS1 course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):89–93, June 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2005.bib.

Minton:2004:IDP

[MBD04] Liz Minton, Roger Boyle, and Vania Dimitrova. If diversity is a problem could e-learning be part of the solution?: a case study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):42–46, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mostrom:2009:CSS

[MBE^{+}09] Jan Erik Moström, Jonas Boustedt, Anna Eckerdal, Robert McCartney, Kate Sanders, Lynda Thomas, and Carol Zander. Computer science student transformations: changes and causes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):181–185, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Misra:2009:MCT

[MBK09] Ananya Misra, Douglas Blank, and Deepak Kumar. A music context for teaching introductory computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):248–252, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Marion:2003:HDR

[MBPB03] William Marion, Adrienne Bloss, Kris Powers, and Doug Baldwin. How departments are responding to the mathematics recommendations in CC2001. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35 (1):189–190, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Melin:2004:POS

[MC04] Ulf Melin and Stefan Cronholm. Project oriented student work: learning & examination. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3): 87–91, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Metrolho:2008:BPO

[MC08] José Carlos Metrolho and Mónica Isabel Teixeira Costa. Branches of professional organizations a way to enrich student’s scientific and personal skills. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (3):360, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

McAndrew:2008:TCO

[McA08] Alasdair McAndrew. Teaching cryptography with opensource software. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):325–329, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

McCauley:2000:CSE

[McC00a] Renée McCauley. Computer science education links — what next? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(4):14–15, December 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCauley:2000:FSS

[McC00b] Renée McCauley. “free source” software — what a blessing! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(2):16–17, June 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCauley:2001:ADM

[McC01a] Renée McCauley. Agile development methods poised to upset status quo. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(4):14–15, December 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2001.bib.

McCauley:2001:BAL

[McC01b] Renée McCauley. A bounty of accessible language translation tools. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(2):14–15, June 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2001.bib.

McCauley:2002:HYH

[McC02] Renée McCauley. Hear ye, hear ye!: a valuable new resource for computer science educators finally available. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):17–18, December 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2002.bib.

McCauley:2003:RTL

[McC03a] Renée McCauley. Resources for teaching and learning about human-computer interaction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):

16–17, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2003.bib.

McCauley:2003:RAG

[McC03b] Renée McCauley. Rubrics as assessment guides. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):17–18, December 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2003.bib.

McCann:2004:CSC

[McC04a] Lester I. McCann. Contemplate sorting with columnsort. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):166–169, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCauley:2004:OL

[McC04b] Renée McCauley. Outgoing links. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (4):17–18, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2004.bib.

McCauley:2004:TAO

[McC04c] Renée McCauley. Thinking about our teaching. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(2):18–19, June 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.

math.utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2004.bib.

McConnell:2005:ACLb

[McC05a] Jeffrey J. McConnell. Active and cooperative learning: more tips and tricks (part II). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4):34–38, December 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2005.bib.

McConnell:2005:ACLa

[McC05b] Jeffrey J. McConnell. Active and cooperative learning: tips and tricks (part I). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):27– 30, June 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2005.bib.

McCormick:2005:WBW

[McC05c] John W. McCormick. We’ve been working on the railroad: a laboratory for real-time embedded systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):530–534, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McConnell:2006:ACLb

[McC06a] Jeffrey J. McConnell. Active and cooperative learning: final tips and tricks (part IV). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):25–28, December 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McConnell:2006:ACLd

[McC06b] Jeffrey J. McConnell. Active and cooperative learning: final tips and tricks (part IV). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):25–28, December 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2006.bib.

McConnell:2006:ACLa

[McC06c] Jeffrey J. McConnell. Active and cooperative learning: further tips and tricks (part 3). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):24– 28, June 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McConnell:2006:ACLc

[McC06d] Jeffrey J. McConnell. Active and cooperative learning: further tips and tricks (part 3). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):24– 28, June 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2006.bib.

McDonald:2002:IPA

[McD02] Jennifer McDonald. Interactive Pushdown Automata Animation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):376–380, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

McGuffee:2000:DCS

[McG00] James W. McGuffee. Defining computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(2):74–76, June 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McGill:2009:WGD

[McG09] Monica McGill. Weighted game developer qualifications for consideration in curriculum development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):347–351, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

McCrickard:2004:DSE

[MCS04] D. Scott McCrickard, C. M. Chewar, and Jacob Somervell. Design, science, and engineering topics?: teaching HCI with a unified method. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):31–35, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McKinney:2004:HWP

[MD04] Dawn McKinney and Leo F. Denton. Houston, we have a problem: there’s a leak in the CS1 affective oxygen tank. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):236–239, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McKinney:2005:AAT

[MD05] Dawn McKinney and Leo F. Denton. Affective assessment of team skills in agile CS1 labs: the good, the bad, and the ugly. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):465–469, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Matzko:2006:TCG

[MD06a] Sarah Matzko and Timothy A. Davis. Teaching CS1 with graphics and C. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):168–172, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McKinney:2006:DCS

[MD06b] Dawn McKinney and Leo F. Denton. Developing collaborative skills early in the CS curriculum in a laboratory environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):138–142, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Matzko:2008:GBA

[MD08] Sarah Matzko and Timothy A. Davis. A graphics-based approach to data structures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):109– 113, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

McFall:2006:EUC

[MDD06] Ryan McFall, Herbert Dershem, and Darcy Davis. Experiences using a collaborative electronic textbook: bringing the “guide on the side” home with you. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (1):339–343, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Medeiros:2005:UPS

[MDGM05] Pedro D. Medeiros, Vitor A. Duarte, M. Cecilia Gomes, and Rui F. Marques. Using a PC simulator to illustrate inputoutput programming techniques. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37 (3):388, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCauley:2000:ASE

[MDH^{+}00] Reneée McCauley, Nell Dale, Thomas Hilburn, Susan Mengel, and Branson W. Murrill. The assimilation of software engineering into the undergraduate computer science curriculum (panel session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):423–424, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McGuffee:2002:MUC

[MDL^{+}02] James W. McGuffee, Herbert L. Dershem, Linda B. Lankewicz, Gary Lewandowski, Dian Lopez, and Oberta A. Slotterbeck. Managing undergraduate CS research. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):416–417, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Morelli:2009:FEI

[MdL09] Ralph Morelli and Trishan de Lanerolle. Foss 101: engaging introductory students in the open source movement. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):311–315, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

McDonald:2004:SAS

[MDM04] Merry McDonald, Brian Dorn, and Gary McDonald. A statistical analysis of student performance in online computer science courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):71–74, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mahmoud:2004:RIC

[MDS04] Qusay H. Mahmoud, Włodek Dobosiewicz, and David A. Swayne. Redesigning introductory computer programming with HTML, JavaScript, and Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):120–124, March 2004. CODEN SIGSD3. ISBN 1-58113-798-2. ISSN 0097-8418 (print), 2331-

3927 (electronic). URL ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Compiler/java.
bib;   ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2004.bib.

Mead:2006:CTS

[Mea06] Jerry Mead. A compiler tutorial scaled for the programming languages course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):32–36, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Medley:2001:UQR

[Med01] M. Dee Medley. Using qualitative research software for CS education research. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):141–144, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Medina:2004:BBB

[Med04] Eden Miller Medina. Beyond the ballot box: computer science education and social responsibility. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(4):7–10, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2004.bib.

McCartney:2007:SSS

[MEM^{+}07] Robert McCartney, Anna Eckerdal, Jan Erik Mostrom, Kate Sanders, and Carol Zander. Successful students’ strategies for getting unstuck. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):156– 160, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Merzbacher:2000:TDM

[Mer00] Matthew Merzbacher. Teaching database management systems with Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):31–35, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Merzbacher:2001:OAI

[Mer01] Matthew Merzbacher. Open artificial intelligence — one course for all. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):110–113, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Merceron:2009:DPS

[Mer09a] Agathe Merceron. Design patterns to support teaching of automata theory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):341, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Mery:2009:SRB

[Mér09b] Dominique Méry. A simple refinement-based method for constructing algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):51–59, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McGovern:2007:CSL

[MF07] Amy McGovern and Jason Fager. Creating significant learning experiences in introductory artificial intelligence. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):39–43, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Moses:2000:TWP

[MFC00] Louise Moses, Sally Fincher, and James Caristi. Teams work (panel session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):421–422, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Marks:2001:TACb

[MFL01] Joe W. Marks, William T. Freeman, and Henry H. Leitner. Teaching applied computing without programming: a case-based introductory course for general education.

SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):80–84, March 2001. CO-DEN SIGSD3. ISBN 1-58113-329-4. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2001.bib;   ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Techreports/MERL.

bib. ”Teaching Applied Computing without Programming: A Case-Based Introductory Course for General Education”, ACM Technical Symposium on Computer Science Education (SIGCSE), ISBN: 1-58113-329-4, Vol 33, Issue 1, pages 80–84, February 2001, Proc ACM Press (http://portal.acm.org/citation.cfm?id=364547&coll=ACM&dl=ACM&CFID=3033011&CFTOKEN=83234373), Contact: Joe Marks (www.merl.com/people/marks/), William Freeman (www.merl.com/people/freeman/).

Ma:2008:UCC

[MFR^{+}08] Linxiao Ma, John D. Ferguson, Marc Roper, Isla Ross, and Murray Wood. Using cognitive conflict and visualisation to improve mental models held by novice programmers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):342–346, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Ma:2009:IMM

[MFR^{+}09] Linxiao Ma, John Ferguson, Marc Roper, Isla Ross, and Murray Wood. Improving the mental models held by novice programmers using cognitive conflict and Jeliot visualisations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):166–170, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Ma:2007:IVM

[MFRW07] Linxiao Ma, John Ferguson, Marc Roper, and Murray Wood. Investigating the viability of mental models held by novice programmers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):499–503, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Matos:2000:RRA

[MG00] Victor Matos and Rebecca Grasser. RELAX — the relational algebra pocket calculator project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (4):40–44, December 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Matos:2004:SBD

[MG04] Victor Matos and Becky Grasser. SQL-based discovery of exact and approximate functional dependencies. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(4):58–63, December 2004. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2004.bib.

Mendes:2005:USC

[MGE^{+}05] António José Mendes, Anabela Gomes, Micaela Esteves, Maria José Marcelino, Crescencio Bravo, and Miguel Angel Redondo. Using simulation and collaboration in CS1 and CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):193–197, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McNally:2006:DLM

[MGFK06] Myles McNally, Michael Goldweber, Barry Fagin, and Frank Klassner. Do Lego Mindstorms robots have a future in CS education? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):61–62, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mead:2006:CAI

[MGH^{+}06] Jerry Mead, Simon Gray, John Hamer, Richard James, Juha Sorva, Caroline St. Clair, and Lynda Thomas. A cognitive approach to identifying measurable milestones for programming skill acquisition. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):182–194, December 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2006.bib.

Muller:2007:POI

[MGH07] Orna Muller, David Ginat, and Bruria Haberman. Patternoriented instruction and its influence on problem decomposition and solution construction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3): 151–155, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Madigan:2007:GPR

[MGS07] Elinor M. Madigan, Marianne Goodfellow, and Jeffrey A. Stone. Gender, perceptions, and reality: technological literacy among first-year students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1): 410–414, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mosiman:2000:ICG

[MH00] Steve Mosiman and Christoph Hiemcke. Interdisciplinary capstone group project: designing autonomous race vehicles. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):270–274, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Muller:2005:GMG

[MH05] Orna Muller and Bruria Haberman. Guidelines for a multiplegoal CS introductory course: algorithmic problem-solving woven into OOP. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):356, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Miliszewska:2006:ETU

[MH06] Iwona Miliszewska and John Horwood. Engagement theory: a universal paradigm? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):158–162, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Martin:2008:CDQ

[MH08] Chris J. Martin and Janet M. Hughes. Cognitive dimensions questionnaire applied to exploratory algorithm design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):184–188, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Muller:2004:APP

[MHA04] Orna Muller, Bruria Haberman, and Haim Averbuch. (an almost) pedagogical pattern for pattern-based problem-solving instruction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):102–106, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mirmotahari:2003:DLC

[MHK03a] Omid Mirmotahari, Christian Holmboe, and Jens Kaasbøll. Difficulties learning computer architecture. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):247, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Murray:2003:EIJ

[MHK^{+}03b] Keitha A. Murray, Jesse M. Heines, Michael Kölling, Tom Moore, Paul J. Wagner, Nan C. Schaller, and John A. Trono. Experiences with IDEs and Java teaching: What works and what doesn’t. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):215–216, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-

3927 (electronic). URL ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Compiler/java.
bib.

Mano:2006:PWS

[MHR06] Meirav Mano, Bruria Haberman, and Tammy Rosenthal. Promoting women in science and technology. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):304, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McDowell:2003:EPP

[MHW03] Charlie McDowell, Brian Hanks, and Linda Werner. Experimenting with pair programming in the classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):60–64, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Marra:2007:SPT

[MHW07] Jonathan D. Marra, Katherine G. Herbert, and Jason T. L. Wang. A study of phylogenetic tools for genomic nomenclature data cleaning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):347, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Michael:2000:FAC

[Mic00] Mark Michael. Fostering and assessing communication skills in the computer science context. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (1):119–123, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Moreno:2008:UAD

[MICM08] Lourdes Moreno, Ana M. Iglesias, Elena Castro, and Paloma Martinez. Using accessible digital resources for teaching database design: towards an inclusive distance learning proposal. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):32–36, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Mikulecky:2005:KSI

[Mik05] Peter Mikulecky. Knowledge sharing as an innovative educational practice: educating knowledge technologies. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):355, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Milner:2009:BMJ

[Mil09] Walter W. Milner. A broken metaphor in Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4):76–77, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Minsker:2007:LTT

[Min07] Steven Minsker. The linear twin Towers of Hanoi problem. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):37–40, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Minsker:2008:ABR

[Min08] Steven Minsker. Another brief recursion excursion to Hanoi. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):35–37, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Minsker:2009:CLH

[Min09] Steven Minsker. The classical/linear Hanoi hybrid problem: regular configurations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4):57–61, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mirolo:2009:MMR

[Mir09] Claudio Mirolo. Mental models of recursive computations vs. recursive analysis in the problem domain. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):397, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Marion:2007:ACS

[MIS^{+}07] Bill Marion, John Impagliazzo, Caroline St. Clair, Barry Soroka, and Deborah Whitfield. Assessing computer science programs: what have we learned. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (1):131–132, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Moorman:2003:SSH

[MJ03] Phoenix Moorman and Elizabeth Johnson. Still a stranger here: attitudes among secondary school students towards computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):193–197, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McDonald:2000:TPA

[MK00] Chris McDonald and Kamran Kazemi. Teaching parallel algorithm with process topologies. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (1):70–74, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Montante:2001:SCP

[MK01] Robert Montante and Zahira Khan. Specialized certification programs in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1): 371–375, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Miller:2002:MPC

[MK02] Amanda Miller and Judy Kay. A mentor program in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):9–13, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Martin:2006:CCI

[MK06] Fred G. Martin and Sarah Kuhn. Computing in context: integrating an embedded computing project into a course on ethical and societal issues. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):525– 529, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Marlowe:2005:DPD

[MKB05] Thomas J. Marlowe, Cyril S. Ku, and James W. Benham. Design patterns for database pedagogy: a proposal. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):48–52, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Musicant:2007:MUR

[MKBW07] David Musicant, Amruth Kumar, Doug Baldwin, and Ellen Walker. Mechanics of undergraduate research at liberal arts colleges: lessons learned. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):65– 66, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Morreale:2009:MSU

[MKC09] Patricia Morreale, Stan Kurkovsky, and George Chang. Methodology for successful undergraduate recruiting in computer science at comprehensive public universities. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):91–95, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Murphy:2008:BTD

[MKKC08] Christian Murphy, Eunhee Kim, Gail Kaiser, and Adam Cannon. Backstop: a tool for debugging runtime errors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):173–177, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Murphy:2009:RHS

[MKLH09] Christian Murphy, Gail Kaiser, Kristin Loveland, and Sahar Hasan. Retina: helping students and instructors based on observed programming activities. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (1):178–182, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Malmi:2002:EAA

[MKS02] Lauri Malmi, Ari Korhonen, and Riku Saikkonen. Experiences in automatic assessment on mass courses and issues for designing virtual courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):55–59, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McKenna:2004:CIP

[ML04] Peter McKenna and Ben Laycock. Constructivist or instructivist: pedagogical concepts practically applied to a computer learning environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):166–170, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Malan:2007:SBC

[ML07] David J. Malan and Henry H. Leitner. Scratch for budding computer scientists. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):223–227, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Moskal:2004:EEN

[MLC04] Barbara Moskal, Deborah Lurie, and Stephen Cooper. Evaluating the effectiveness of a new instructional approach. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):75–79, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Modak:2005:PBD

[MLH05] Vishal D. Modak, David D. Langan, and Thomas F. Hain. A pattern-based development tool for mobile agents. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):72–75, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Myller:2007:AET

[MLK07] Niko Myller, Mikko Laakso, and Ari Korhonen. Analyzing engagement taxonomy in collaborative algorithm visualization. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):251–255, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Murphy:2008:DGB

[MLM^{+}08] Laurie Murphy, Gary Lewandowski, Renée McCauley, Beth Simon, Lynda Thomas, and Carol Zander. Debugging: the good, the bad, and the quirky — a qualitative analysis of novices’ strategies. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):163–167, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

McCauley:2000:IRC

[MM00] Renée McCauley and Bill Manaris. An information resource for computer science educators. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(2): 25–29, June 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Moreno:2004:PAJ

[MMBAS04] Andrés Moreno, Niko Myller, Mordechai Ben-Ari, and Erkki Sutinen. Program animation in Jeliot 3. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):265, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mendes:2001:CFT

[MMC01] Emilia Mendes, Nile Mosley, and Steve Counsell. The cognitive flexibility theory: an approach for teaching Hypermedia Engineering. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):21–24, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mann:2009:CSE

[MMD^{+}09] Samuel Mann, Logan Muller, Janet Davis, Claudia Roda, and Alison Young. Computing and sustainability: evaluating resources for educators. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4):144–155, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Moskal:2002:GEC

[MMK02] Barbara Moskal, Keith Miller, and L. A. Smith King. Grading essays in computer ethics: rubrics considered helpful. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):101–105, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

McGettrick:2009:RAI

[MMLT09] Andrew McGettrick, Renée McCauley, Richard LeBlanc, and Heikki Topi. Report on the ACM/IEEE-CS undergraduate curricula recommendations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):267– 268, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Maheswaran:2009:GUL

[MMN^{+}09] Muthucumaru Maheswaran, Alexis Malozemoff, Daniel Ng, Sheng Liao, Song Gu, Balasubramaneyam Maniymaran, Julie Raymond, Reehan Shaikh, and Yuanyuan Gao. GINI: a userlevel toolkit for creating micro Internets for teaching & learning computer networking. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):39– 43, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Maia:2005:CFO

[MMP05] Luiz Paulo Maia, Francis Berenger Machado, and Ageu C. Pacheco, Jr. A constructivist framework for operating systems education: a pedagogic proposal using the SOsim. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):218–222, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Moreno:2007:AIM

[MMRI07] Lourdes Moreno, Paloma Martínez, Belén Ruíz, and Ana Iglesias. Accessible interface for multimedia presentation in inclusive education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):321, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Murphy:2005:MII

[MMW^{+}05] Laurie Murphy, Renée McCauley, Suzanne Westbrook, Timothy Fossum, Susan Haller, Briana B. Morrison, Brad Richards, Kate Sanders, Carol Zander, and Ruth E. Anderson. A multi-institutional investigation of computer science seniors’ knowledge of programming concepts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):510–514, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Morrison:2001:SIS

[MN01] Malcolm Morrison and Timothy S. Newman. A study of the impact of student background and preparedness on outcomes in CS I. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):179–183, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Meeden:2003:UDS

[MNBK03] Lisa Meeden, Tia Newhall, Doug Blank, and Deepak Kumar. Using departmental surveys to assess computing culture: quantifying gender differences in the classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):188–192, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mahmoud:2009:AKI

[MNN^{+}09] Qusay H. Mahmoud, Thanh Ngo, Razieh Niazi, Pawel Popowicz, Robert Sydoryshyn, Matthew Wilks, and Dave Dietz. An academic kit for integrating mobile devices into the CS curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):40–44, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

McWhorter:2009:DLM

[MO09] William Isaac McWhorter and Brian C. O’Connor. Do LEGO(R) Mindstorms(R) motivate students in CS1? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):438–442, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Mohr:2009:TNP

[Moh09] Jonathan Mohr. Two novel Prolog assignments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):350, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Montano:2002:VDC

[Mon02] Gloria Montano. Virtual development center. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):70–73, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.

math.utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/

Misc/DBLP/2002.bib. Women and Computing.

Moore:2002:BED

[Moo02] Thomas K. Moore. Bringing the enterprise into a database systems course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):262–265, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Moreno:2008:PAA

[Mor08] Andrés Moreno. Program animation activities in Moodle. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):361, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Moskal:2002:FCS

[Mos02] Barbara Moskal. Female computer science doctorates: what does the survey of earned doctorates reveal? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):105–111, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2002.bib. Women and

Computing.

Moses:2003:DIV

[Mos03]  Louise E. Moses.  Design issues in the visual era.

SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):52–56, June 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

Moses:2006:APA

[Mos06] Louise Moses. Animation programming: an alternative approach to CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):317, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Motil:2006:SS

[Mot06] John Motil. Surprise snippets. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3): 351, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mahoney:2009:TCC

[MP09a] William Mahoney and Jay Pedersen. Teaching compiler code generation: simpler is better. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4): 30–34, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Morrison:2009:EGT

[MP09b] Briana B. Morrison and Jon A. Preston. Engagement: gaming throughout the curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1): 342–346, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

More:2005:BXI

[MPD^{+}05] Sara Miner More, Tim Pevzner, Alin Deutsch, Scott Baden, and Paul Kube. Building an XQuery interpreter in a compiler construction course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):2–6, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mitchell:2009:CSW

[MPH09] Alison Mitchell, Helen C. Purchase, and John Hamer. Computing science: what do pupils think? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):353, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Mora:2009:LMB

[MPJC09] Higinio Mora Mora, María Teresa Signes Pont, Rafael Camps Jordá, and Juan Manuel García Chamizo. Learning method based on collaborative assessment performed by the students: an application to computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (3):372, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Maloney:2008:PCU

[MPK^{+}08a] John H. Maloney, Kylie Peppler, Yasmin Kafai, Mitchel Resnick, and Natalie Rusk. Programming by choice: urban youth learning programming with scratch. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):367–371, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Murphy:2008:DLA

[MPK08b] Christian Murphy, Dan Phung, and Gail Kaiser. A distance learning approach to teaching eXtreme programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):199–203, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

McCracken:2000:ESC

[MPQB^{+}00] Daniel D. McCracken, Manuel A. Pérez-Quiñones, Robert Bryant, Fred Springsteel, and Anne-Louise Radimsky. Experiences in starting computer engineering programs (panel session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):401–402, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Martin:2000:NSF

[MR00] C. Dianne Martin and Margaret Reek. The National Science Foundation (seminar session): funding opportunities for CS faculty through the CCLI program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):428, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Miller:2001:PEC

[MR01] Nancy E. Miller and Donna S. Reese. A placement examination for computer science II. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(2):39– 42, June 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2001.bib.

Micco:2002:BCL

[MR02] Mary Micco and Hart Rossman. Building a cyberwar lab: lessons learned: teaching cybersecurity principles to undergraduates. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):23–27, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Moreira:2005:IMR

[MR05] Nelma Moreira and Rogério Reis. Interactive manipulation of regular objects with FAdo. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):335– 339, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Markov:2006:IWD

[MR06] Zdravko Markov and Ingrid Russell. An introduction to the WEKA data mining system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):367– 368, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mhiri:2009:ADI

[MR09] Faten Mhiri and Sylvie Ratté. AARTIC: development of an intelligent environment for human learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):359, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

McGettrick:2008:RPB

[MRGS08] Andrew McGettrick, Eric Roberts, Daniel D. Garcia, and Chris Stevenson. Rediscovering the passion, beauty, joy and awe: making computing fun again. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (1):217–218, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

McMaster:2009:TMG

[MRH09] Kirby McMaster, Brian Rague, and Steven Hadfield. Two mathematical gestalts for computer theory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):251–255, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Murphy:2006:WCG

[MRM^{+}06] Laurie Murphy, Brad Richards, Renée McCauley, Briana B. Morrison, Suzanne Westbrook, and Timothy Fossum. Women catch up: gender differences in learning programming concepts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):17–21, March 2006. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Makinen:2000:RUD

[MS00] Erkki Mäkinen and Markku Siermala. Restricted universe data structures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(2):22–24, June 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Moody:2003:IQA

[MS03] Daniel L. Moody and Guttorm Sindre. Incorporating quality assurance processes into requirements analysis education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):74–78, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Marrero:2005:TFE

[MS05a] Will Marrero and Amber Settle. Testing first: emphasizing testing in early programming courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):4–8, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Morgado:2005:MAM

[MS05b] Carmen Morgado and Luís Soares. MIG21 API: multimedia interactive groups API. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):399, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Miller:2007:TAI

[MS07] James Miller and Michael Smith. A TDD approach to introducing students to embedded programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):33–37, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Miura:2009:AGI

[MSK09] Motoki Miura, Taro Sugihara, and Susumu Kunifuji. Anchor Garden: an interactive workbench for basic data concept learning in object oriented programming languages. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):141–145, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Mann:2008:CES

[MSM08] Samuel Mann, Lesley Smith, and Logan Muller. Computing education for sustainability. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4): 183–193, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McCauley:2006:CBS

[MSP^{+}06] Renée McCauley, Christopher W. Starr, Walter Pharr, Rox-Ann H. Stalvey, and George Pothering. Is CS1 better with the same lecture and lab instructor? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (2):54–60, June 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2006.bib.

Merritt:2004:DPD

[MSS^{+}04] Susan M. Merritt, Allen Stix, Judith E. Sullivan, Fred Grossman, Charles C. Tappert, and David A. Sachs. Developing a professional doctorate in computing: a fifth-year assessment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(4):42–46, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2004.bib.

Meisalo:2002:FES

[MSTS02] Veijo Meisalo, Jarkko Suhonen, Sirpa Torvinen, and Erkki Sutinen. Formative evaluation scheme for a Web-based course design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):130–134, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Murphy:2005:DCS

[MT05] Laurie Murphy and Josh Tenenberg. Do computer science students know what they know?: a calibration study of data structure knowledge. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):148–152, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Murphy:2008:DFM

[MT08] Laurie Murphy and Lynda Thomas. Dangers of a fixed mindset: implications of self-theories research for computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):271–275, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Moffat:2009:UMT

[MT09] David C. Moffat and Kathryn Trinder. The use of MUVE technology in teaching AI algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):396, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Megias:2009:FTA

[MTBS09] David Megias, Wouter Tebbens, Lex Bijlsma, and Francesc Santanach. Free technology academy: a European initiative for distance education about free software and open standards. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):70–74, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Mento:2000:IMT

[MTHS00] Barbara Mento, Donna Tupper, Kathleen Harmeyer, and Sylvia Sorkin. Internet and multimedia technology curriculum development (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (3):190–191, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mundie:2009:RC

[Mun09] Craig Mundie. Rethinking computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):203, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Martinez-Unanue:2002:EBP

[MUPVPF^{+}02] R. Martínez-Unanue, M. Paredes-Velasco, C. Pareja-Flores, J. Urquiza-Fuentes, and J. Á. Velázquez-Iturbide. Electronic books for programming education: a review and future prospects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):34–38, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Murtagh:2001:TBF

[Mur01] Thomas P. Murtagh. Teaching breadth-first depth-first. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):37–40, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Murtagh:2007:SBV

[Mur07a] Thomas P. Murtagh. Squint: barely visible library support for CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):526–530, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Murtagh:2007:WCC

[Mur07b] Thomas P. Murtagh. Weaving CS into CS1: a doubly depth-first approach. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):336–340, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Musicant:2006:DMC

[Mus06] David R. Musicant. A data mining course for computer science: primary sources and implementations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):538–542, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mitchener:2001:CRA

[MV01] W. Garrett Mitchener and Amin Vahdat. A chat room assignment for teaching network security. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33 (1):31–35, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McAndrew:2005:SLD

[MV05] Alasdair McAndrew and Anne Venables. A “secondary” look at digital image processing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):337– 341, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Maj:2000:CTT

[MVC00] S. P. Maj, D. Veal, and P. Charlesworth. Is computer technology taught upside down? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):140– 143, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Maj:2001:PNH

[MVD01] S. P. Maj, D. Veal, and R. Duley. A proposed new high level abstraction for computer technology. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33 (1):199–203, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Mulder:2001:IUI

[MvW01] Fred Mulder and Tom J. van Weert. IFIP/UNESCO’s Informatics Curriculum Framework 2000 for higher education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(4):75–83, December 2001. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2001.bib.

Meneely:2009:PSD

[MW09] Andrew Meneely and Laurie Williams. On preparing students for distributed software development with a synchronous, collaborative development platform. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (1):529–533, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

McDowell:2002:EPP

[MWBF02] Charlie McDowell, Linda Werner, Heather Bullock, and Julian Fernald. The effects of pair-programming on performance in an introductory programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):38–42, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Mullins:2002:PIS

[MWF^{+}02] Paul Mullins, Jim Wolfe, Michael Fry, Erik Wynters, William Calhoun, Robert Montante, and William Oblitey. Panel on integrating security concepts into existing computer courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):365–366, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Meneely:2008:RRE

[MWG08] Andrew Meneely, Laurie Williams, and Edward F. Gehringer. ROSE: a repository of education-friendly open-source projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):7–11, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Morelli:2001:JAH

[MWM01] Ralph Morelli, Ralph Walde, and Gregg Marcuccio. A Java API for historical ciphers: an object-oriented design project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):307–311, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Moritz:2005:OFD

[MWPB05] Sally H. Moritz, Fang Wei, Shahida M. Parvez, and Glenn D. Blank. From objects-first to design-first with multimedia and intelligent tutoring. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):99–103, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nau:2006:RN

[Nau06] Richard W. Nau. A random number. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (3):345, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nicholas:2001:TED

[NB01] Tyrone Nicholas and Jerzy A. Barchanski. TOS: an educational distributed operating system in Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):312–316, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Norris:2008:CCQ

[NBF^{+}08] Cindy Norris, Frank Barry, James B. Fenwick Jr., Kathryn Reid, and Josh Rountree. ClockIt: collecting quantitative data on how beginning software developers really work. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):37–41, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Norvell:2009:RCP

[NBL09] Theodore S. Norvell and Michael P. Bruce-Lockhart. Rich content plug-ins for the teaching machine. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):348, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Nair:2007:MCP

[NC07] Achuth Sankar S. Nair and Vinod Chandra. A MetaL for C and Pascal. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):87–91, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2007.bib.

Null:2002:RCC

[NCA^{+}02] Linda Null, Mike Ciaraldi, Liz Adams, Ursula Wolz, and Max Hailperin. Rewrite cycles in CS courses: experience reports. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):249–250, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Naps:2003:EEI

[NCK^{+}03] Thomas L. Naps, Stephen Cooper, Boris Koldehofe, Charles Leska, Guido Rößling, Wanda Dann, Ari Korhonen, Lauri Malmi, Jarmo Rantakokko, Rockford J. Ross, Jay Anderson, Rudolf Fleischer, Marja Kuittinen, and Myles F. Mc-Nally. Evaluating the educational impact of visualization. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):124–136, December 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

Narasimhan:2008:DIS

[ND08] V. Lakshmi Narasimhan and Manik Lal Das. Data and information security (DIS) for BS and MS programs: a proposal. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):95–99, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Neff:2004:ABC

[Nef04] Norman Neff. Attribute based compiler implemented using visitor pattern. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):130–134, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Naps:2000:JEA

[NEN00] Thomas L. Naps, James R. Eagan, and Laura L. Norton. JHAVÉ — an environment to actively engage students in Web-based algorithm visualizations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (1):109–113, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nodelman:2005:CAC

[NH05] Vladimir Nodelman and Bruria Haberman. Complex analysis in computer graphics course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):381, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Noonan:2007:CSD

[NH07] Robert E. Noonan and John R. Hott. A course in software development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):135–139, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ni:2009:WMC

[Ni09] Lijun Ni. What makes CS teachers change?: factors influencing CS teachers’ adoption of curriculum innovations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):544–548, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Nishida:2009:CUD

[NKI^{+}09] Tomohiro Nishida, Susumu Kanemune, Yukio Idosaka, Mitaro Namiki, Tim Bell, and Yasushi Kuno. A CS unplugged design pattern. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):231– 235, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Natvig:2004:ACG

[NL04] Lasse Natvig and Steinar Line. Age of computers: gamebased teaching of computer fundamentals. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):107–111, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Norte:2007:VLK

[NL07] Stéphane Norte and Fernando Graça Lobo. A virtual logo keyboard for people with motor disabilities. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):111–115, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Norte:2008:SGP

[NL08] Stéphane Norte and Fernando Lobo. A Sudoku game for people with motor impairments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):319, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Ngai:2009:ISA

[NLCvL09] Grace Ngai, Winnie W. Y. Lau, Stephen C. F. Chan, and Hong va Leong. On the implementation of self-assessment in an introductory programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4):85–89, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Neeman:2006:ATP

[NLMN06] Henry Neeman, Lloyd Lee, Julia Mullen, and Gerard Newman. Analogies for teaching parallel computing to inexperienced programmers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):64–67, December 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2006.bib.

Newhall:2002:CPC

[NM02] Tia Newhall and Lisa Meeden. A comprehensive project for CS2: combining key data structures and algorithms into an integrated Web browser and search engine. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):386–390, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Nair:2004:CDS

[NM04] Achuth Sankar S. Nair and T. Mahalakshmi. Conceptualizing data structures: a pedagogic approach. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(4):97–100, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2004.bib.

Nelson:2000:TCN

[NN00] Daniel Nelson and Yau Man Ng. Teaching computer networking using open source software. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (3):13–16, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nodelman:2003:LCG

[Nod03] Vladimir Nodelman. Learning computer graphics by programming: linking theory and practice. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):261, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nodelman:2004:OCC

[Nod04] Vladimir Nodelman. OOP via C++, C#, … ? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):255, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Noonan:2000:OOV

[Noo00] Robert E. Noonan. An object-oriented view of backtracking. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):362–366, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Noonan:2006:BEG

[Noo06] Robert E. Noonan. The back end of a grading system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):56–60, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Normark:2000:SWB

[Nør00] Kurt Nørmark. A suite of WWW-based tools for advanced course management. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):65–68, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Noyes:2002:FCC

[Noy02] James L. Noyes. A first course in computational science: (why a math book isn’t enough). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (1):18–22, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Neuwirth:2002:IIP

[NP02a] Erich Neuwirth and Viera K. Proulx. IFIP and issues in pre-college education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):23–24, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2002.bib.

Noonan:2002:UTF

[NP02b] Robert E. Noonan and Richard H. Prosl. Unit testing frameworks. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):232–236, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Nienaltowski:2008:CEM

[NPM08] Marie-Hélène Nienaltowski, Michela Pedroni, and Bertrand Meyer. Compiler error messages: what can help novices? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):168–172, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Null:2005:CIM

[NR05] Linda Null and Karishma Rao. CAMERA: introducing memory concepts via visualization. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1): 96–100, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Naps:2003:ERV

[NRA^{+}03] Thomas L. Naps, Guido Rößling, Vicki L. Almstrum, Wanda Dann, Rudolf Fleischer, Christopher D. Hundhausen, Ari Korhonen, Lauri Malmi, Myles F. McNally, Susan H. Rodger, and J. Ángel Velázquez-Iturbide. Exploring the role of visualization and engagement in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):131–152, June 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

Naps:2005:DXB

[NRB^{+}05] Thomas L. Naps, Guido Rößling, Peter Brusilovsky, John English, Duane J. Jarc, Ville Karavirta, Charles Leska, Myles F. McNally, Andrés Moreno, Rockford J. Ross, and Jaime Urquiza-Fuentes. Development of XML-based tools to support user interaction with algorithm visualization. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4):123–138, December 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2005.bib.

Nita-Rotaru:2007:MEA

[NRDP07] Cristina Nita-Rotaru, Melissa Dark, and Voicu Popescu. A multi-expertise application-driven class. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):119–123, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Naps:2006:AVC

[NRRR06] Thomas Naps, Susan Rodger, Guido Rößling, and Rockford Ross. Animation and visualization in the curriculum: opportunities, challenges, and successes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (1):328–329, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nguyen:2004:DPM

[NRW04] Dung (‘Zung’) Nguyen, Mathias Ricken, and Stephen Wong. Design patterns for marine biology simulation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):467–471, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nguyen:2005:DPP

[NRW05] Dung (‘Zung’) Nguyen, Mathias Ricken, and Stephen Wong. Design patterns for parsing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):477– 481, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nikolic:2002:GTL

[NS02] Dejan Nikolic and Ching-Kuang Shene. GraphicsMentor: a tool for learning graphics fundamentals. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):242–246, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Nishihara:2009:MCE

[NSH^{+}09] Hideaki Nishihara, Koichi Shinozaki, Koji Hayamizu, Toshiaki Aoki, Kenji Taguchi, and Fumihiro Kumeno. Model checking education for software engineers in Japan. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):45–50, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nugent:2005:DDV

[NSS^{+}05] Gwen Nugent, Leen-Kiat Soh, Ashok Samal, Suzette Person, and Jeff Lang. Design, development, and validation of a learning object for CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):370, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Neeman:2008:SPE

[NSW08] Henry Neeman, Horst Severini, and Dee Wu. Supercomputing in plain English: teaching cyberinfrastructure to computing novices. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):27–30, June 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2008.bib.

Nickens:2009:IHE

[NTTHT09] Glenn V. Nickens, Ethan J. Tira-Thompson, Thorna Humphries, and David S. Touretzky. An inexpensive handeye system for undergraduate robotics instruction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):423–427, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Nieh:2005:ETO

[NV05] Jason Nieh and Chris Vaill. Experiences teaching operating systems using virtual platforms and Linux. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):520–524, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Navarro:2004:SES

[NvdH04] Emily Oh Navarro and André van der Hoek. SimSE: an educational simulation game for teaching the software engineering process. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):233, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Navarro:2009:MSE

[NvdH09] Emily Navarro and André van der Hoek. Multi-site evaluation of SimSE. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):326–330, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Nguyen:2000:DPL

[NW00] Dung Nguyen and Stephen B. Wong. Design patterns for lazy evaluation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):21–25, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nguyen:2001:DPS

[NW01] Dung (‘Zung’) Nguyen and Stephen B. Wong. Design patterns for sorting. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):263–267, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nguyen:2002:DPG

[NW02] Dung (‘Zung’) Nguyen and Stephen B. Wong. Design patterns for games. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):126–130, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Nevison:2003:TOE

[NW03] Chris Nevison and Barbara Wells. Teaching objects early and design patterns in Java using case studies. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):94–98, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Compiler/java.bib.

Niezgoda:2006:SST

[NW06] Sebastian Niezgoda and Thomas P. Way. SNITCH: a software tool for detecting cut and paste plagiarism. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):51–55, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nagappan:2003:ICE

[NWF^{+}03] Nachiappan Nagappan, Laurie Williams, Miriam Ferzli, Eric Wiebe, Kai Yang, Carol Miller, and Suzanne Balik. Improving the CS1 experience with pair programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):359–362, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Nygaard:2002:CCO

[Nyg02] Kristen Nygaard. COOL (comprehensive object-oriented learning). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):218, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Owens:2007:CEO

[OAB07] Barbara Boucher Owens, Vicki L. Almstrum, and Lecia J. Barker. The computing educators oral history project: planning for the future. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):328, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Or-Bach:2005:EBM

[OB05] Rachel Or-Bach. Educational benefits of metadata creation by students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4):93–97, December 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2005.bib.

Oechsle:2007:CAO

[OB07] Rainer Oechsle and Kay Barzen. Checking automatically the output of concurrent threads. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3): 43–47, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Or-Bach:2003:SUO

[OBL03] Rachel Or-Bach and Ilana Lavy. Students’ understanding of object orientation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):251, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Or-Bach:2004:CAA

[OBL04] Rachel Or-Bach and Ilana Lavy. Cognitive activities of abstraction in object orientation: an empirical study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(2):82–86, June 2004. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2004.bib.

Olsson:2003:RRO

[OBLH03] B. Olsson, M. Berndtsson, B. Lundell, and J. Hansson. Running research-oriented final year projects for CS and IS students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):79–83, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ou:2000:IIW

[OCLL00] Kuo-Liang Ou, Gwo-Dong Chen, Chen-Chung Liu, and Baw-Jhiune Liu. Instructional instruments for Web group learning systems: the grouping, intervention, and strategy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):69–72, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

OConnor:2001:SPI

[OCM01] Rory O’Connor, Gerry Coleman, and Maurizio Morisio. Software process improvement education (poster session): a European experiment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):180, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Odekirk:2000:ASP

[Ode00] Elizabeth Odekirk. Analyzing student programs (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):191, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Olagunju:2001:JCS

[OG01] Amos O. Olagunju and Katrenia Geiger. Just clicking some theoretical aspects of computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33 (2):37–38, June 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2001.bib.

Oechsle:2005:DDA

[OG05] Rainer Oechsle and Tim Gottwald. DisASTer (distributed algorithms simulation terrain): a platform for the implementation of distributed algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3): 44–48, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

OKelly:2006:RPB

[OG06] Jackie O’Kelly and J. Paul Gibson. RoboCode & problembased learning: a non-prescriptive approach to teaching programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):217–221, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Oliver:2007:UDT

[OGMA07] Iain Oliver, Kristoffer Getchell, Alan Miller, and Colin Allison. Using disruptive technology for explorative learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):96–100, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Owens:2001:UPA

[OHRL01] Barbara Boucher Owens, Fran Hunt, Stuart Reges, and Cary Laxer. The university perspective: awarding credit for advanced placement (AP) in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):396–397, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Odekirk-Hash:2001:AFP

[OHZ01] Elizabeth Odekirk-Hash and Joseph L. Zachary. Automated feedback on programs means students need less help from teachers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):55–59, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Odekirk:2000:TSC

[OJJ00] Elizabeth Odekirk, Dominic Jones, and Peter Jensen. Three semesters of CSO using Java: assignments and experiences. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):144–147, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Okasaki:2005:ATC

[Oka05] Chris Okasaki. Alternatives to two classic data structures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):162–165, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

OLeary:2006:LBC

[O’L06] Mike O’Leary. A laboratory based capstone course in computer security for undergraduates. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (1):2–6, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Olagunju:2000:RSD

[Ola00] Amos O. Olagunju. The role of scientific discovery in teaching and learning of computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (4):28–31, December 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Oldham:2008:EOC

[Old08] Joseph D. Oldham. Experience offering a course centered on cyberporn. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):158–162, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Olivieri:2005:HSE

[Oli05] Lisa M. Olivieri. High school environments and girls’ interest in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):85– 88, June 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2005.bib.

Olsevicova:2005:ATM

[Ols05] Kamila Olsevicova. Application of topic maps in e-learning environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):363, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Olson:2006:EDU

[Ols06] Clark F. Olson. Encouraging the development of undergraduate researchers in computer vision. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (3):255–259, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Owens:2008:CNA

[OM08] Jim Owens and Jeanna Matthews. Cybercivics: a novel approach to reaching K-12 students with the social relevance of computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):372–376, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

ONeill:2005:AUW

[O’N05] Melissa E. O’Neill. Automated use of a Wiki for collaborative lecture notes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):267–271, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Owens:2002:JIW

[OPS^{+}02] Barbara Boucher Owens, Rich Pattis, Chris Stephenson, Jack Rehder, and Dean Sanders. JAVA IDEs: why and how we use what we do. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):213–214, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Ortiz:2003:TSE

[Ort03] Ariel Ortiz. Teaching the SIMD execution model:: assembling a few parallel programming skills. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35 (1):74–78, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ortiz:2008:LDI

[Ort08] Ariel Ortiz. Language design and implementation using Ruby and the interpreter pattern. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):48– 52, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Osborne:2002:PCT

[Osb02] Hugh Osborne. The postroom computer: teaching introductory undergraduate computer architecture. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):157–161, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Osborne:2006:TSW

[Osb06] Lawrence J. Osborne. Thinking, speaking, and writing for freshmen. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):112–116, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

OCinneide:2004:PBA

[ÓT04] Mel Ó Cinnéide and Richard Tynan. A problem-based approach to teaching design patterns. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (4):80–82, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2004.bib.

Omwega:2007:APR

[OTUW07a] Eric Omwega, Jonathan Tsu, Monica Ugwi, and Christine Wang. Automation of PE recruiting process. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):350, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Omwega:2007:DOC

[OTUW07b] Eric Omwega, Jonathan Tsu, Monica Ugwi, and Christine Wang. Development of an online campus map. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):351, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Ourusoff:2004:RSE

[Our04] Nicholas Ourusoff. Reinvigorating the software engineering curriculum with Jackson’s methods and ideas. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(2):93–96, June 2004. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2004.bib;
ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.uka.de/
bibliography/SE/alspaugh.bib.

Owens:2008:NDI

[Owe08] Barbara Boucher Owens. New directions for inroads. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):112, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pickard:2003:MDT

[PA03] Michael M. Pickard and Jason R. Adams. Model determination tool (MDT): a multipurpose software engineering learning utensil. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):76–78, December 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2003.bib.

Plimmer:2006:PTE

[PA06] Beryl Plimmer and Robert Amor. Peer teaching extends HCI learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):53–57, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pantaleev:2009:DVC

[Pan09] Alex Pantaleev. Dzver: a visual computer science learning environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):387, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Papadimitriou:2003:MPS

[Pap03a] Christos H. Papadimitriou. MythematiCS: in praise of storytelling in the teaching of computer science and math. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):7–9, December 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

Papadimitriou:2003:MST

[Pap03b] Christos H. Papadimitriou. Mythematics: storytelling in the teaching of computer science and mathematics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):1, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 8th Annual Conference on Innovation and Technology in Computer Science Education.

Papalaskari:2003:PEA

[Pap03c] Mary-Angela Papalaskari. Peer evaluation in an algorithms course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):274, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Papastergiou:2003:INC

[Pap03d] Marina Papastergiou. Introducing non-computer science undergraduates to Web development: results of an investigative study on student teachers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):231, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Parlante:2001:NAT

[Par01] Nick Parlante. Nifty assignments: Tetris on the brain. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(4):25–27, December 2001. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2001.bib.

Pardo:2002:MAP

[Par02a] Abelardo Pardo. A multi-agent platform for automatic assignment management. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):60–64, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Parlante:2002:TLR

[Par02b] Nick Parlante. Tree-list recursion. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (4):30–31, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2002.bib.

Parlante:2003:AL

[Par03a] Nick Parlante. Astrachan’s law. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4): 26–27, December 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2003.bib.

Parlante:2003:PES

[Par03b] Nick Parlante. Platonic Euler solids. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35 (2):30–31, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2003.bib.

Parlante:2004:NAG

[Par04a] N. Parlante. Nifty assignments: Graphics and Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(2):26–27, June 2004. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Compiler/java.bib.

Parlante:2004:GJ

[Par04b] Nick Parlante. Graphics and Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (2):26–27, June 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2004.bib.

Parlante:2004:N

[Par04c] Nick Parlante. Niftiness. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(4):26–27, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2004.bib.

Parlante:2005:DCE

[Par05a] Nick Parlante. A day for CS education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4):27–28, December 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2005.bib.

Parlante:2005:WCS

[Par05b] Nick Parlante. What is computer science? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):24–25, June 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2005.bib.

Pargas:2006:TTP

[Par06a] Roy Pargas. A targeted tablet PC software development course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):336, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pargas:2006:RLI

[Par06b] Roy P. Pargas. Reducing lecture and increasing student activity in large computer science courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):3–7, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Parlante:2007:CI

[Par07a] Nick Parlante. Cheating and the Internet. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):29–30, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2007.bib.

Parlante:2007:NR

[Par07b] Nick Parlante. Nifty reflections. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2): 25–26, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2007.bib.

Parlante:2008:NA

[Par08] Nick Parlante. Nifty assignments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (1):112–113, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Parlante:2009:NAT

[Par09] Nick Parlante. Nifty assignments: too much inheritance. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):83–84, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pastel:2005:ISR

[Pas05] Robert Pastel. Integrating science and research in a HCI design course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):31–35, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pattis:2006:CSC

[Pat06] Richard Pattis. Can’t sing, can’t act, can dance a little: (on choosing the right dancing partners). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (1):1, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pausch:2008:ADM

[Pau08] Randy Pausch. Alice: a dying man’s passion. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):1, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Parker:2003:MEC

[PB03a] J. R. Parker and Katrin Becker. Measuring effectiveness of constructivist and behaviourist assignments in CS102. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):40–44, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Phillips:2003:CSC

[PB03b] Andrew T. Phillips and C. Alex Buerkle. A computational science case study: classification of hybrids using genetic markers and maximum-likelihood estimates. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):94–98, June 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

Pais:2005:TPA

[PB05a] Rui Pais and João Paulo Barros. Two possible approaches for an intermediate GUI course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):365, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pais:2005:UFM

[PB05b] Rui Pais and João Paulo Barros. Use of flash movies for teaching GUI programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):390, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Peterson:2006:OCI

[PB06] Laurence I. Peterson and Dale Benham. Overview of the cybertech-ITEST project: an initiative to attract and prepare under-represented students for tomorrow’s careers in the computing sciences. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):453–455, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pargas:2009:UIE

[PB09a] Roy P. Pargas and Samuel Bryfczynski. Using ink to expose students’ thought processes in CS2/CS7. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):168–172, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Pargas:2009:WWT

[PB09b] Roy P. Pargas and Samuel Bryfczynski. What were they thinking? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):134–138, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Palmer-Brown:2009:GLD

[PBDWL09] Dominic Palmer-Brown, Chrisina Draganova, and Sin Wee-Lee. Guided learning via diagnostic feedback to question responses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):362, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Pahl:2004:SAD

[PBK04] Claus Pahl, Ronan Barrett, and Claire Kenny. Supporting active database learning and training through interactive multimedia. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):27–31, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Parker:2003:CSA

[PBM03] J. R. Parker, Katrin Becker, and Douglas McCullough. Computer science and the arts: some multi-disciplinary teaching collaborations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):252, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Purewal:2007:ESR

[PBM07] Tarsem S. Purewal, Jr., Chris Bennett, and Frederick Maier. Embracing the social relevance: computing, ethics and the community. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):556–560, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pedroni:2007:OSP

[PBOP07] Michela Pedroni, Till Bay, Manuel Oriol, and Andreas Pedroni. Open source projects in programming courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):454–458, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Peterson:2006:CIN

[PBWS06] Laurence I. Peterson, Dale Benham, Amy B. Woszczynski, and Rich Schlesinger. CyberTech-ITEST: an NSF-funded initiative to attract and prepare underrepresented students for tomorrow’s careers in the computing sciences. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):332–333, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pullen:2008:DAL

[PC08] J. Mark Pullen and Jim X. Chen. Distributed application launching for high quality graphics in synchronous distance education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):204–208, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Parlante:2001:NAP

[PCR^{+}01] Nick Parlante, Mike Clancy, Stuart Reges, Julie Zelenski, and Owen Astrachan. Nifty assignments panel. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):412–413, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Parlante:2007:NA

[PCS^{+}07] Nick Parlante, John Cigas, Angela B. Shiflet, Raja Sooriamurthi, Mike Clancy, Bob Noonan, and David Reed. Nifty assignments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):497–498, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Philpott:2009:USP

[PCW09] Anne Philpott, Tony Clear, and Jacqueline Whalley. Understanding student performance on an algorithm simulation task: implications for guided learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):408–412, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Pears:2003:SCR

[PD03] Arnold N. Pears and Mats Daniels. Structuring CSed research studies: connecting the pieces. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3): 149–153, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pollard:2006:ENK

[PD06] Shannon Pollard and Robert C. Duvall. Everything I needed to know about teaching I learned in kindergarten: bringing elementary education techniques to undergraduate computer science classes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):224–228, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Parrish:2001:BSC

[PDC01] Allen Parrish, Brandon Dixon, and David Cordes. Binary software components in the undergraduate computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):332–336, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Papadopoulos:2007:CBI

[PDS07] Pantelis M. Papadopoulos, Stavros N. Demetriadis, and Ioannis G. Stamelos. Case-based instruction on the Web for teaching software project management. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (3):136–140, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Pears:2002:DGV

[Pea02] Arnold Pears. Defining a global view of CS education research. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):227, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Powers:2007:TLG

[PEH07] Kris Powers, Stacey Ecott, and Leanne M. Hirshfield. Through the looking glass: teaching CS0 with Alice. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):213–217, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Parlante:2002:NA

[PER^{+}02] Nick Parlante, John K. Estell, David Reed, David Levine, Dan Garcia, and Julie Zelenski. Nifty assignments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):319–320, March 2002. CODEN SIGSD3.

ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Perrenet:2009:DBA

[Per09] Jacob Perrenet. Differences in beliefs and attitudes about computer science among students and faculty of the Bachelor program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):129–133, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Petre:2009:DEW

[Pet09] Marian Petre. Distance education: What our children can teach us about distance learning and learning programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4):17–18, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pollard:2003:HLC

[PF03] Shannon Pollard and Jeffrey Forbes. Hands-on labs without computers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):296–300, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pacini:2004:LPS

[PFF04] Giuliano Pacini, Giuseppe Fiorentino, and Annalina Fabrizio. Learning problem solving with spreadsheet and database tools. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):267, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pisa:2005:BRN

[PFJC05] Ricardo Pisa, Vasco Ferreira, Elisabete Jesus, and Bruno Carlos. BRaiN research network. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):400, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pareja-Flores:2000:LVC

[PFVI00] Cristóbal Pareja-Flores and J. Ángel Velázquez-Iturbide. Local versus comprehensive assignments: two complementary approaches. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(4):48–51, December 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Proulx:2006:DCH

[PG06] Viera K. Proulx and Kathryn E. Gray. Design of class hierarchies: an introduction to OO program design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):288–292, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Piattini:2007:ACQ

[PG07] Mario Piattini and Félix García. Adapting the course “quality of information systems” to E.H.E.A. guidelines. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):50–53, June 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Powers:2006:TTI

[PGC^{+}06] Kris Powers, Paul Gross, Steve Cooper, Myles McNally, Kenneth J. Goldman, Viera Proulx, and Martin Carlisle. Tools for teaching introductory programming: what works? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):560–561, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Perrenet:2005:ESU

[PGK05] Jacob Perrenet, Jan Friso Groote, and Eric Kaasenbrood. Exploring students’ understanding of the concept of algorithm: levels of abstraction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):64–68, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Paterson:2003:TJU

[PH03] James H. Paterson and John Haddow. Teaching Java: Using an object-oriented database and the BlueJ IDE. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):273, September 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Compiler/java.bib.

Paterson:2004:AOP

[PH04a] James H. Paterson and John Haddow. Approaches to object persistence in Java projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):256, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Compiler/
java.bib.

Paterson:2004:PDP

[PH04b] James H. Paterson and John Haddow. A proposed design patterns extension for the blueJ IDE. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (3):279, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Paxton:2004:PCA

[PH04c] John Paxton and Karim Heredia. Providing curricular assistance to a third world university computer science department. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):199–202, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Paterson:2007:CCS

[PH07] James H. Paterson and John Haddow. From classes to code: supporting the transition from design to implementation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):362, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Paterson:2005:UBI

[PHBM05] James H. Paterson, John Haddow, Miriam Birch, and Alex Monaghan. Using the BlueJ IDE in a data structures course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):349, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Paterson:2008:DLT

[PHC08] James H. Paterson, John Haddow, and Ka Fai Cheng. Drawing the line: teaching the semantics of binary class associations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):362, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Paterson:2009:UCP

[PHC09] James H. Paterson, John Haddow, and Ka Fai Cheng. Using coding patterns in a model-driven approach to teaching object oriented programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):358, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Pheatt:2007:EUD

[Phe07] Chuck Pheatt. An easy to use distributed computing framework. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):571–575, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Phillips:2004:HPL

[Phi04] Andrew T. Phillips. Highlighting programming language issues using mixed language programming in Maple and C. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):85–89, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Parrish:2001:IAV

[PHM^{+}01] Allen Parrish, Joe Hollingsworth, Peter Maurer, Benjamin Shults, and Bruce Weide. Identifying an appropriate view of software components for undergraduate education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):394–395, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Poe:2008:REC

[PHMS08] David Poe, Christine Hansen, Kellie McGowan, and Gautam Singh. Refining educational content through a closedloop FLOW approach. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):102– 106, June 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2008.bib.

Paterson:2006:DPE

[PHN06] James H. Paterson, John Haddow, and Michael Nairn. A design patterns extension for the BlueJ IDE. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):280–284, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Phoha:2001:IDM

[Pho01] Vir V. Phoha. An interactive dynamic model for integrating knowledge management methods and knowledge sharing technology in a traditional classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):144–148, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Peckham:2007:BPC

[PHS^{+}07] Joan Peckham, Lisa L. Harlow, David A. Stuart, Barbara Silver, Helen Mederer, and Peter D. Stephenson. Broadening participation in computing: issues and challenges. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):9–13, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Pau:2007:WCH

[PHW07] Reena Pau, Wendy Hall, and Su White. Women in computing: how does experience influence self-perception of computing careers? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):349, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Popyack:2003:ADH

[PHZ^{+}03] Jeffrey L. Popyack, Nira Herrmann, Paul Zoski, Bruce Char, Chris Cera, and Robert N. Lass. Academic dishonesty in a high-tech environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):357– 358, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pijls:2000:LLP

[Pij00] Wim Pijls. LR and LL parsing: some new points of view. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(4):24–27, December 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pillay:2003:DIP

[Pil03] Nelishia Pillay. Developing intelligent programming tutors for novice programmers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):78– 82, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2003.bib.

Pillay:2004:FCG

[Pil04] Nelishia Pillay. A first course in genetic programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(4):93–96, December 2004. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2004.bib.

Pillay:2009:LDE

[Pil09] Nelishia Pillay. Learning difficulties experienced by students in a course on formal languages and automata theory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4):48–52, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pollock:2001:MPP

[PJ01] Lori Pollock and Mike Jochen. Making parallel programming accessible to inexperienced programmers through cooperative learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):224–228, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pillay:2005:ISC

[PJ05] Nelishia Pillay and Vikash R. Jugoo. An investigation into student characteristics affecting novice programming performance. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4):107–110, December 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2005.bib.

Proulx:2009:UTJ

[PJ09] Viera K. Proulx and Weston Jossey. Unit testing in Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):349, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Perrenet:2006:LAS

[PK06] Jacob Perrenet and Eric Kaasenbrood. Levels of abstraction in students’ understanding of the concept of algorithm: the qualitative perspective. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):270– 274, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Poplawski:2008:JPT

[PK08] David A. Poplawski and Zachary Kurmas. JLS: a pedagogically targeted logic design and simulation tool. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):314, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Paul:2006:TLF

[PKB06] Jerome L. Paul, Michal Kouril, and Kenneth A. Berman. A template library to facilitate teaching message passing parallel computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):464–468, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Patterson:2003:IUT

[PKR03] Andrew Patterson, Michael Kölling, and John Rosenberg. Introducing unit testing with BlueJ. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35 (3):11–15, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Parberry:2006:ASG

[PKR06] Ian Parberry, Max B. Kazemzadeh, and Timothy Roden. The art and science of game programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):510–514, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Paz:2004:EAA

[PL04a] Tamar Paz and Tami Lapidot. Emergence of automated assignment conceptions in a functional programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):181–185, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Prior:2004:BES

[PL04b] Julia Coleman Prior and Raymond Lister. The backwash effect on SQL skills grading. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3): 32–36, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Penha-Lopes:2005:WUO

[PL05] Joana Matos Penha-Lopes. Why use an open source e-voting system? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):412, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Parlante:2005:NA

[PLA^{+}05] Nick Parlante, David Levine, Steven Andrianoff, Aaron J. Gordon, Alyce Brady, Pamela Cutter, Paul Kube, Jefferson Ng, and Richard E. Pattis. Nifty assignment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):371–372, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Plekhanova:2002:LPS

[PM02a] Valentina Plekhanova and Walter Middleton. Learning processes in software engineering projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):230, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pratt:2002:PAV

[PM02b] Lorien Y. Pratt and Manavendra Misra. Perspectives on academic vs. industry environments for women in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):20–22, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/

2002.bib. Women and Computing.

Pedroni:2006:ICP

[PM06] Michela Pedroni and Bertrand Meyer. The inverted curriculum in practice. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):481–485, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pollock:2004:IHS

[PMC^{+}04] Lori Pollock, Kathleen McCoy, Sandra Carberry, Namratha Hundigopal, and Xiaoxin You. Increasing high school girls’ self confidence and awareness of CS through a positive summer experience. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):185–189, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Paterson:2005:TXW

[PMC05] James H. Paterson, Frances McCormick, and Gerry Creechan. Teaching XML in a Web development context. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):359, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Perez:2008:CLO

[PMF08] Jorge E. Pérez, Javier García Martín, and Isabel Muñoz Fernández. Cooperative learning in operating systems laboratory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):323, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Powell:2004:DLC

[PMG^{+}04] Norman Powell, David Moore, John Gray, Janet Finlay, and John Reaney. Dyslexia and learning computer programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):242, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Parlante:2004:NA

[PML^{+}04] Nick Parlante, David Matuszek, Jeff Lehman, David Reed, John K. Estell, and Donald Chinn. Nifty assignments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):46–47, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Parlante:2009:NA

[PMS^{+}09] Nick Parlante, Thomas P. Murtagh, Mehran Sahami, Owen Astrachan, David Reed, Christopher A. Stone, Brent Heeringa, and Karen Reid. Nifty assignments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):483–484, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Prey:2003:TAN

[PMT03] Jane Prey, Ernest McDuffie, and Harriet Taylor. Taking advantage of national science foundation funding opportunities. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):142, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pearce:2008:FGO

[PN08] Janice Pearce and Mario Nakazawa. The funnel that grew our CIS major in the CS desert. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (1):503–507, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Pullen:2000:TCM

[PNF00] J. Mark Pullen, Eugene Norris, and Mark Fix. Teaching C++ in a multi-user virtual environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (2):60–64, June 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pohlhaus:2005:EED

[Poh05] William Pohlhaus. Ethical engagement with data collection efforts related to fighting terrorists and terrorism in the context of recent events. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):401, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pollock:2001:IIE

[Pol01] Lori Pollock. Integrating an intensive experience with communication skills development into a computer science course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):287–291, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pedroni:2007:FDC

[POM07] Michela Pedroni, Manuel Oriol, and Bertrand Meyer. A framework for describing and comparing courses and curricula. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):131–135, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Pedroni:2008:CMT

[POM^{+}08] Michela Pedroni, Manuel Oriol, Bertrand Meyer, Enrico Albonico, and Lukas Angerer. Course management with TrucStudio. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):260–264, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Pedroni:2008:AEN

[POMA08] Michela Pedroni, Manuel Oriol, Bertrand Meyer, and Lukas Angerer. Automatic extraction of notions from course material. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):251–255, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Popyack:2002:IHS

[Pop02] Jeffrey L. Popyack. The international honor society for the computing and information disciplines. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):25–26, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2002.bib.

Popyack:2003:SSA

[Pop03a] Jeffrey Popyack. Scholarship, student awards, microbreweries, and baseball. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):26–27, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2003.bib.

Popyack:2003:SAA

[Pop03b] Jeffrey Popyack. Scholarships, awards, advice, and the abacus. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):21–23, December 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

Popyack:2004:BYU

[Pop04] Jeffrey L. Popyack. Banner year for UPE. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(4):21–22, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2004.bib.

Popyack:2005:ROA

[Pop05a] Jeffrey Popyack. Recognizing outstanding achievement. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4):22–24, December 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2005.bib.

Popyack:2005:HSC

[Pop05b] Jeffrey L. Popyack. Honor societies, college and service. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):18–20, June 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2005.bib.

Popyack:2007:NAC

[Pop07a] Jeffrey L. Popyack. New alliances and continuing excellence. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):21–23, June 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Popyack:2007:ST

[Pop07b] Jeffrey L. Popyack. Society in transition. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):25–26, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2007.bib.

Popyack:2008:UPE

[Pop08a] Jeffrey Popyack. Upsilon Pi Epsilon: UPE happenings. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):25–26, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Popyack:2008:TYD

[Pop08b] Jeffrey L. Popyack. Take your daughters (and sons) to work: and leave them there. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):22– 23, June 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2008.bib.

Popyack:2009:UPE

[Pop09a] Jeffrey Popyack. Upsilon pi epsilon: New happenings. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4):20–21, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Popyack:2009:BPA

[Pop09b] Jeffrey L. Popyack. Blackjack-playing agents in an advanced AI course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):208–212, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Porcheron:2009:BPE

[Por09] Patrick Porcheron. The Bologna Process in European education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):3, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Powers:2003:BRI

[Pow03] Kris D. Powers. Breadth-also: a rationale and implementation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):243–247, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Powell:2008:IPF

[Pow08] Rita Manco Powell. Improving the persistence of first-year undergraduate women in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):518–522, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Papadopoulos:2003:UTA

[PP03a] George Papadopoulos and Hariton M. Polatoglou. Using ToolBook authoring tool to facilitate the teaching of algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):277, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Papamanthou:2003:VPS

[PP03b] Charalampos Papamanthou and Konstantinos Paparrizos. A visualization of the primal simplex algorithm for the assignment problem. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):267, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Price:2004:CEU

[PP04] Colin Price and John Price. Common environment for undergraduate computer programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (3):248, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Polze:2006:TOS

[PP06] Andreas Polze and Dave Probert. Teaching operating systems: the windows case. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):298– 302, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pecinovsky:2007:OES

[PP07] Rudolf Pecinovsky and Jarmila Pavlickova. Order of explanation should be interface: abstract classes — overriding. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):338, September 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Polycarpou:2008:CAT

[PPA08] Irene Polycarpou, Ana Pasztor, and Malek Adjouadi. A conceptual approach to teaching induction for computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):9–13, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Pears:2003:EOL

[PPE03] Arnold N. Pears, Lars Pettersson, and Carl Erickson. Enriching online learning resources with “explanograms”. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):237, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pivkina:2009:YWC

[PPJH09] Inna Pivkina, Enrico Pontelli, Rachel Jensen, and Jessica Haebe. Young women in computing: lessons learned from an educational & outreach program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (1):509–513, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Pecinovsky:2006:LMO

[PPP06] Rudolf Pecinovský, Jarmila Pavlícková, and Lubos Pavlícek. Let’s modify the objects-first approach into design-patterns-first. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):188–192, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pearl:2002:BCS

[PPR^{+}02] Amy Pearl, Martha E. Pollack, Eve A. Riskin, Becky Thomas, Elizabeth Wolf, and Alice Wu. Becoming a computer scientist. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):135–143, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2002.bib. Women and

Computing.

Parlante:2003:NA

[PPR^{+}03] Nick Parlante, Jeffrey Popyack, Stuart Reges, Stephen Weiss, Scott Dexter, Chaya Gurwitz, Joseph Zachary, and Grant Braught. Nifty assignments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1): 353–354, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pengelley:2006:PAB

[PPRV06] David Pengelley, Inna Pivkina, Desh Ranjan, and Karen Villaverde. A project in algorithms based on a primary historical source about Catalan numbers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (1):318–322, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Perez-Quinones:2004:TOE

[PQEA^{+}04] M. A. Pérez-Quiñones, Steven Edwards, Claude Anderson, Doug Baldwin, James Caristi, and Paul J. Wagner. Transitioning to an objects-early three-course introductory sequence: issues and experiences. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1): 499–500, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Parkinson:2002:DCS

[PR02] Adrian Parkinson and James A. Redmond. Do cognitive styles affect learning performance in different computer media? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):39–43, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Proulx:2004:JIT

[PR04] Viera K. Proulx and Richard Rasala. Java IO and testing made simple. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):161–165, March 2004. CODEN SIGSD3. ISBN 1-58113-798-2. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Compiler/java.bib;   ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/2004.
bib.

Prakash:2005:CGA

[Pra05a] Edmond C. Prakash. Computer game arena: a showcase for computer game projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):358, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Prakash:2005:IFL

[Pra05b] Edmond C. Prakash. Implicit functions lab: a collaboratory for computer graphics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):376, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pfaff:2009:PIO

[PRB09] Ben Pfaff, Anthony Romano, and Godmar Back. The pintos instructional operating system kernel. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):453–457, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Pak:2006:IDD

[PRC06] Suehee Pak, Eunha Rho, and Juno Chang. Industrial demand-driven curriculums for computer-software field in Korea. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):63–67, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Prey:2001:ASN

[Pre01] Jane Prey. ACM SIGCSE NSF CCLI project showcase. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):418, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Preston:2007:BLI

[Pre07] Christina Preston. Braided learning: investigating an emerging model of online professional debate between international educators. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):62–64, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Parberry:2005:EID

[PRK05] Ian Parberry, Timothy Roden, and Max B. Kazemzadeh. Experience with an industry-driven capstone course on game programming: extended abstract. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37 (1):91–95, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pivkina:2009:HST

[PRL09] Inna Pivkina, Desh Ranjan, and Jerry Lodder. Historical sources as a teaching tool. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):401– 402, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Proulx:2000:PPD

[Pro00] Viera K. Proulx. Programming patterns and design patterns in the introductory computer science course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):80–84, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Proulx:2009:TDD

[Pro09] Viera K. Proulx. Test-driven design for introductory OO programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):138–142, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Proulx:2002:OBG

[PRR02] Viera K. Proulx, Jeff Raab, and Richard Rasala. Objects from the beginning — with GUIs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (3):65–69, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Parkinson:2004:AFD

[PRW04] Adrian Parkinson, James A. Redmond, and Cathal Walsh. Accommodating field-dependence: a cross-over study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):72–76, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pargas:2006:TCC

[PS06] Roy P. Pargas and Dhaval M. Shah. Things are clicking in computer science courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):474– 478, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pears:2005:CCL

[PSE^{+}05] Arnold Pears, Stephen Seidman, Crystal Eney, Päivi Kinnunen, and Lauri Malmi. Constructing a core literature for computing education research. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4): 152–161, December 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2005.bib.

Peckham:2007:ISR

[PSH^{+}07] Joan Peckham, Peter Stephenson, Jean-Yves Hervé, Ron Hutt, and Miguel Encarnação. Increasing student retention in computer science through research programs for undergraduates. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):124–128, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pears:2007:SLT

[PSM^{+}07] Arnold Pears, Stephen Seidman, Lauri Malmi, Linda Mannila, Elizabeth Adams, Jens Bennedsen, Marie Devlin, and James Paterson. A survey of literature on the teaching of introductory programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4): 204–223, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2007.bib.

Pharr:2009:PPD

[PSS09] Walter M. Pharr, Christopher W. Starr, and Caroline S. Starr. Paired professional development: a methodology for continued professional development in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):218–222, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Phillips:2003:ICB

[PSW03] Andrew T. Phillips, Daniel E. Stevenson, and Michael R. Wick. Implementing CC2001: a breadth-first introductory course for a just-in-time curriculum design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):238–242, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Prey:2002:WDY

[PT02] Jane Prey and Kevin Treu. What do you say?: open letters to women considering a computer science major. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):18–20, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2002.bib. Women and

Computing.

Phillips:2003:ESV

[PT03] Andrew T. Phillips and Jack S. E. Tan. Exploring security vulnerabilities by exploiting buffer overflow using the MIPS ISA. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):172–176, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pfleeger:2001:IEW

[PTC^{+}01] Shari Lawrence Pfleeger, Pat Teller, Sheila E. Castaneda, Manda Wilson, and Rowan Lindley. Increasing the enrollment of women in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33 (1):386–387, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Petkovic:2006:TPS

[PTT06] Dragutin Petkovic, Gary Thompson, and Rainer Todtenhoefer. Teaching practical software engineering and global software engineering: evaluation and comparison. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):294–298, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Petkovic:2008:ACL

[PTT08] Dragutin Petkovic, Gary D. Thompson, and Rainer Todtenhoefer. Assessment and comparison of local and global SW engineering practices in a classroom setting. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):78–82, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Pullen:2000:IBL

[Pul00] J. Mark Pullen. The Internet-based lecture: converging teaching and technology. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):101– 104, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pullen:2006:SDE

[Pul06] J. Mark Pullen. Scaling up a distance education program in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):33–37, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pullen:2007:NED

[Pul07] J. Mark Pullen. Network educationware demonstration. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):355, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Putnik:2001:ILT

[Put01] Zoran Putnik. On integration of learning and technology (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):185, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Plane:2008:CCB

[PV08] Jandelyn D. Plane and Isabella Venter. Comparing capacity building frameworks for computer science education in underdeveloped countries: an Asian and African perspective.

SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):306–310, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Preston:2001:OCL

[PW01] Jon A. Preston and Laura Wilson. Offering CS1 on-line reducing campus resource demand while improving the learning environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):342–346, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pulimood:2008:PSC

[PW08] Sarah Monisha Pulimood and Ursula Wolz. Problem solving in community: a necessary shift in CS pedagogy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):210–214, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Polack-Wahl:2006:LSU

[PWA06] Jennifer A. Polack-Wahl and Karen Anewalt. Learning strategies and undergraduate research. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):209–213, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Peiper:2005:EDP

[PWC^{+}05] Chad Peiper, David Warden, Ellick Chan, Boris Capitanu, and Sam Kamin. eFuzion: development of a pervasive educational system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):237–240, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pike:2000:CCC

[PWH00] Scott M. Pike, Bruce W. Weide, and Joseph E. Hollingsworth. Checkmate: cornering C++ dynamic memory errors with checked pointers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):352–356, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Parlante:2006:NA

[PWM^{+}06] Nick Parlante, Steven A. Wolfman, Lester I. McCann, Eric Roberts, Chris Nevison, John Motil, Jerry Cain, and Stuart Reges. Nifty assignments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):562– 563, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Pavlov:2005:BEA

[PY05] Vladimir L. Pavlov and Anton Yatsenko. “The Babel experiment”: an advanced pantomime-based training in OOA&OOD with UML. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):231– 235, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Yang:2008:AGW

[pYB08] Kuo pao Yang and Theresa Beaubouef. Automatic generation of Web survey for assessment purposes in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):78–82, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Qidwai:2007:LLE

[Qid07] Uvais A. Qidwai. A LAMP-LEGO experience of motivating minority students to study engineering. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):41–44, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2007.bib.

Qin:2009:TCT

[Qin09] Hong Qin. Teaching computational thinking through bioinformatics to biology students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1): 188–191, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Qian:2009:TRT

[QLT09] Kai Qian, Jigang Liu, and Lixin Tao. Teach real-time embedded system online with real hands-on labs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):367, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Quade:2004:PUR

[Qua04] Ann M. Quade. Promoting undergraduate research: a classroom model. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):235, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Quade:2006:DHS

[Qua06] Ann M. Quade. Developing a hybrid software engineering curse that promotes project-based active learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):308, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Quarless:2007:RFD

[Qua07] D. A. Quarless. Redundant features of design in blackboard (LMS) and user error. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):177– 179, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2007.bib.

Quinn:2006:CBA

[Qui06] Michael J. Quinn. Case-based analysis: a practical tool for teaching computer ethics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):520– 524, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Razmov:2006:PTS

[RA06] Valentin Razmov and Richard Anderson. Pedagogical techniques supported by the use of student devices in teaching software engineering. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):344–348, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Radenski:2006:PFL

[Rad06] Atanas Radenski. “Python first”: a lab-based digital introduction to computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):197– 201, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Radenski:2007:DSA

[Rad07] Atanas Radenski. Digital support for abductive learning in introductory computing courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (1):14–18, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Radenski:2008:DCS

[Rad08] Atanas Radenski. Digital CS1 study pack based on Moodle and Python. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):325, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Radenski:2009:FCM

[Rad09] Atanas Radenski. Freedom of choice as motivational factor for active learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):21–25, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Rafieymehr:2008:KCK

[Raf08] Ali Rafieymehr. Kids in computing (K.I.C.): is there a solution to solve the computer science enrollment problem? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):107–111, June 2008. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2008.bib.

Ragonis:2004:RAA

[Rag04] Noa Ragonis. A refreshing approach to an academic seminar course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):236, September 2004.

CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rahal:2008:URE

[Rah08] Imad Rahal. Undergraduate research experiences in data mining. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):461–465, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Rajaravivarma:2005:GBA

[Raj05] Rathika Rajaravivarma. A games-based approach for teaching the introductory programming course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4):98–102, December 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2005.bib.

Ralston:2005:DWN

[Ral05] Anthony Ralston. Do we need ANY mathematics in computer science curricula? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):6– 9, June 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2005.bib.

Ramakrishnan:2002:DDC

[Ram02] Sita Ramakrishnan. DoIT: dynamic curriculum organisation by innovation through technology. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (3):242, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ramakrishan:2003:MSL

[Ram03] Sita Ramakrishan. MUSE studio lab and innovative software engineering capstone project experience. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):21–25, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Raner:2000:TOO

[Ran00] Mirko Raner. Teaching object-orientation with the Object Visualization and Annotation Language (OVAL). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):45–48, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Randolph:2007:FMR

[Ran07a] Justus J. Randolph. Findings from “A Methodological Review of the Computer Science Education Research: 2000– 2005”. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):130, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Randolph:2007:PEP

[Ran07b] Justus J. Randolph. Planning and evaluating programs in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):131, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2007.bib.

Rao:2005:ICT

[Rao05] M. R. K. Krishna Rao. Infusing critical thinking skills into content of AI course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):173–177, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rao:2006:SPS

[Rao06] M. R. K. Krishna Rao. Storytelling and puzzles in a software engineering course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):418–422, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rasala:2000:TFY

[Ras00] Richard Rasala. Toolkits in first year computer science: a pedagogical imperative. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):185– 191, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rasala:2001:ERH

[Ras01] Richard Rasala. Exploring recursion in Hilbert curves (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):194, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rasala:2003:EOV

[Ras03] Richard Rasala. Embryonic object versus mature object: object-oriented style and pedagogical theme. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):89–93, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reiser:2009:FTL

[RB09] Susan L. Reiser and Rebecca F. Bruce. Fabrication: a tangible link between computer science and creativity. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):382–386, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Ragonis:2005:USD

[RBA05] Noa Ragonis and Mordechai Ben-Ari. On understanding the statics and dynamics of object-oriented programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):226–230, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reed:2002:IEM

[RBC^{+}02] David Reed, Doug Baldwin, Michael Clancy, Allen Downey, and Stuart Hansen. Integrating empirical methods into computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):48–49, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Roberts:2005:AJT

[RBC^{+}05] Eric S. Roberts, Kim B. Bruce, Robb Cutler, James H. Cross II, Scott B. Grissom, Karl Klee, Susan H. Rodger, Fran Trees, Ian Utting, and Frank Yellin. The ACM Java Task Force: Status report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):46–47, March 2005. CODEN SIGSD3. ISBN 1-58113-997-7. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Compiler/java.bib;   ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/2005.
bib.

Roberts:2006:AJT

[RBC^{+}06] Eric Roberts, Kim Bruce, James H. Cross II, Robb Cutler, Scott Grissom, Karl Klee, Susan Rodger, Fran Trees, Ian Utting, and Frank Yellin. The ACM Java Task Force: final report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):131–132, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reategui:2007:UVC

[RBC07] Eliseo Reategui, Elisa Boff, and John A. Campbell. Using virtual characters in personalized recommendations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):180–184, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rodger:2006:TAT

[RBFR06] Susan H. Rodger, Bart Bressler, Thomas Finley, and Stephen Reading. Turning automata theory into a hands-on course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):379–383, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rapanotti:2002:EVG

[RBG02] Lucia Rapanotti, Canan Tosunoglu Blake, and Robert Griffiths. eTutorials with voice groupware: real-time conferencing to support computing students at a distance. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):116–120, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rodrigo:2009:ABP

[RBJ^{+}09] Ma. Mercedes T. Rodrigo, Ryan S. Baker, Matthew C. Jadud, Anna Christine M. Amarra, Thomas Dy, Maria Beatriz V. Espejo-Lahoz, Sheryl Ann L. Lim, Sheila A. M. S. Pascua, Jessica O. Sugay, and Emily S. Tabanao. Affective and behavioral predictors of novice programmer achievement. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):156–160, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Renaud:2001:TPD

[RBlR01] Karen Renaud, John Barrow, and Petra le Roux. Teaching programming from a distance: problems and a proposed solution. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(4):39–42, December 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2001.bib.

Rossling:2004:ISL

[RBM04] Guido Rößling, Henning Bär, and Max Mühlhäuser. Interaction support in large-scale lectures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (3):244, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rossling:2003:MCE

[RBTC03] Guido Rößling, Henning Bär, Christoph Trompler, and Chin-Man Choi. Mobile computing in education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):227, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rayside:2000:AOO

[RC00] Derek Rayside and Gerard T. Campbell. Aristotle and object-oriented programming: why modern students need traditional logic. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):237–244, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ratte:2002:UWC

[RC02] Sylvie Ratté and Jocelyne Caron. On using the Web as a collaboration space in the context of an industrial simulation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):141–145, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ramakrishnan:2004:SBF

[RC04a] S. Ramakrishnan and A. Cambrell. Service based framework for knowledge portals. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):260, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reis:2004:TPI

[RC04b] Charles Reis and Robert Cartwright. Taming a professional IDE for the classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):156–160, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roberts:2002:CCI

[RCD^{+}02] Eric Roberts, C. Fay Cover, Gordon Davies, Michael Schneider, and Robert Sloan. Computing Curricula 2001 implementing the recommendations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1): 167–168, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Roberts:2000:CPS

[RCE^{+}00] Eric Roberts, C. Fay Cover, Gerald Engel, Carl Chang, James H. Cross II, and Russ Shackelford. Curriculum 2001 (panel session): evaluating the Strawman report representatives of the ACM/IEEE-CS task force. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):411–412, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rosbottom:2000:GML

[RCF00] John Rosbottom, Jonathan Crellin, and Dave Fysh. A generic model for on-line learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):108– 111, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rosson:2004:CST

[RCR04] Mary Beth Rosson, John M. Carroll, and Con M. Rodi. Case studies for teaching usability engineering. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):36–40, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rosso:2000:OMT

[RD00] Ana Rosso and Marcela Daniele. Our method to teach algorithmic development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(2):49–52, June 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ramnath:2008:EIC

[RD08] Sarnath Ramnath and Brahma Dathan. Evolving an integrated curriculum for object-oriented analysis and design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):337–341, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Rosmaita:2006:ACS

[RDCE06] Brian J. Rosmaita, Katherine Deibel, Robert F. Cohen, and Mary Anne L. Egan. Accessibility and computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):63–64, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rebelsky:2005:NSS

[Reb05] Samuel A. Rebelsky. The new science students in too much, too soon an abbreviated, accelerated, constructivist, collaborative, introductory experience in CS. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):312–316, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roberts:2001:CCH

[REC^{+}01] Eric Roberts, Gerald Engel, C. Fay Cover, Andrew McGettrick, Carl Chang, and Ursula Wolz. Computing curricula 2001 how will it work for you? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1): 433–434, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Redmond:2001:CPA

[Red01] Michael A. Redmond. A computer program to aid assignment of student project groups. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):134– 138, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reese:2000:UMG

[Ree00] Donna S. Reese. Using multiplayer games to teach interprocess communication mechanisms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (4):45–47, December 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reed:2001:RCJ

[Ree01] David Reed. Rethinking CS0 with JavaScript. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):100–104, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reek:2002:WTS

[Ree02] Kenneth A. Reek. The well-tempered semaphore: theme with variations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):356–359, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Reek:2004:DPS

[Ree04] Kenneth A. Reek. Design patterns for semaphores. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):320–324, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reges:2000:CRJ

[Reg00] Stuart Reges. Conservatively radical Java in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):85–89, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reges:2002:CCR

[Reg02a] Stuart Reges. Can C# replace Java in CS1 and CS2? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):4–8, September 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Compiler/java.bib.

Reges:2002:SFI

[Reg02b] Stuart Reges. SIGCSE 2002 forum: industry in the ivory tower. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):367, March 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Reges:2003:UUT

[Reg03] Stuart Reges. Using undergraduates as teaching assistants at a state university. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):103–107, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reges:2006:BBC

[Reg06] Stuart Reges. Back to basics in CS1 and CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):293–297, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reges:2008:MBB

[Reg08] Stuart Reges. The mystery of “b := (b = false)”. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):21–25, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Rebelsky:2000:RWP

[RF00a] Samuel A. Rebelsky and Clif Flynt. Real-world program design in CS2: the roles of a large-scale, multi-group class project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):192–196, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rossling:2000:EUA

[RF00b] Guido Rößling and Bernd Freisleben. Experiences in using animations in introductory computer science lectures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):134–138, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rossling:2000:TPS

[RF00c] Guido Rößling and Bernd Freisleben. TOPKAPI (poster session): a tool for performing knowledge tests over the WWW. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):191, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rossling:2001:AES

[RF01] Guido Rößling and Bernd Freisleben. AnimalScript: an extensible scripting language for algorithm animation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):70–74, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Raicu:2009:EUE

[RF09] Daniela Stan Raicu and Jacob David Furst. Enhancing undergraduate education: a REU model for interdisciplinary research. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):468–472, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Ragonis:2003:MLD

[RH03] Noa Ragonis and Bruria Haberman. A multi-level distance learning-based course for high-school computer science leading-teachers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):224, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Russell:2007:TTL

[RH07] Ingrid Russell and Susan Haller. To teach is to touch lives forever. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):1, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ragonis:2008:TMP

[RH08a] Noa Ragonis and Orit Hazzan. Tutoring model for promoting teaching skills of computer science prospective teachers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):276–280, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Rossling:2008:WOP

[RH08b] Guido Rößling and Sebastian Hartte. WebTasks: online programming exercises made easy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3): 363, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Rowan:2007:YUH

[RHC07] Murray Rowan, Tim Hawkins, and Andy Cobley. Yahoo University hackday. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):306, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Russell:2007:PST

[RHDR07] Ingrid Russell, Susan M. Haller, J. D. Dougherty, and Susan H. Rodger, editors. SIGCSE 2007: proceedings of the Thirty-Eighth SIGCSE Technical Symposium on Computer Science Education: March 7–10, 2007, Covington, Kentucky, USA, SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 2007. CODEN SIGSD3. ISBN 1-59593-361-1 (paperback). ISSN 0097-8418 (print), 2331-3927 (electronic). LCCN QA76.27 .A34 2007.

Ratcliffe:2004:ESL

[RHE04] Mark Ratcliffe, Jonathan Holloway, and Wayne Ellis. Enhancing student learning through collaboration. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):272, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rebelsky:2001:WDD

[RHKS01] Samuel A. Rebelsky, Peter B. Henderson, Amruth N. Kumar, and F. N. (Fred) Springsteel. Why I do declare!: declarative programming in the undergraduate curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):398–399, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rodger:2009:EMS

[RHL^{+}09] Susan H. Rodger, Jenna Hayes, Gaetjens Lezin, Henry Qin, Deborah Nelson, Ruth Tucker, Mercedes Lopez, Stephen Cooper, Wanda Dann, and Don Slater. Engaging middle school teachers and students with Alice in a diverse set of subjects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):271–275, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Raveh:2007:THS

[RHY07] Barak Raveh, Bruria Haberman, and Cecile Yehezkel. Transforming a high school student project in computer science into a significant scientific achievement. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (3):331, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Repenning:2008:BPT

[RI08] Alexander Repenning and Andri Ioannidou. Broadening participation through scalable game design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):305–309, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Richards:2000:BFT

[Ric00] Brad Richards. Bugs as features: teaching network protocols through debugging. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):256–259, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Richards:2002:INCb

[Ric02] Brad Richards. Illustrating networking concepts with wireless handheld devices. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):240, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Richards:2009:RWC

[Ric09] Brad Richards. Representation of women in CS: how do we measure a program’s success? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1): 96–100, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Riordan:2002:TIL

[Rio02] Denis Riordan. Towards an integrated learning laboratory environment for first-year computer science students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):112–116, December 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2002.bib.

Rosson:2009:SID

[RIP^{+}09] Mary Beth Rosson, Anastassia Ioujanina, Timothy Paone, Gretel Sheasley, Hansa Sinha, Craig Ganoe, John M. Carroll, and Jan Mahar. A scaffolded introduction to dynamic Website development for female high school students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):226–230, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Reed:2003:WA

[RJ03] Dale Reed and Sam John. Web annotator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):386–390, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reed:2004:CFJ

[RJAH04] Dale Reed, Sam John, Ryan Aviles, and Feihong Hsu. CFX: finding just the right examples for CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):363–367, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rao:2005:PCD

[RJM^{+}05] M. R. K. Krishna Rao, S. Junaidu, T. Maghrabi, M. Shafique, M. Ahmed, and K. Faisal. Principles of curriculum design and revision: a case study in implementing computing curricula CC2001. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):256–260, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rossling:2008:ELM

[RJM^{+}08] Guido Rößling, Mike Joy, Andrés Moreno, Atanas Radenski, Lauri Malmi, Andreas Kerren, Thomas Naps, Rockford J. Ross, Michael Clancy, Ari Korhonen, Rainer Oechsle, and J. Ángel Velázquez Iturbide. Enhancing learning management systems to better support computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):142–166, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rossling:2009:SIE

[RK09a] Guido Rößling and Yavor Kolarov. A system for integral efficiency analysis of sustainable technologies. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):395, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Rossling:2009:EMB

[RK09b] Guido Rößling and Andreas Kothe. Extending Moodle to better support computing education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (3):146–150, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Robbins:2002:ISC

[RKD02] Kay A. Robbins, Catherine Sauls Key, and Keith Dickinson. Integrating a simulation case study into CS2: developing design, empirical and analysis skills. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (1):391–395, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Robbins:2001:SCC

[RKDM01] Kay A. Robbins, Catherine Sauls Key, Keith Dickinson, and John Montgomery. Solving the CS1/CS2 lab dilemma: students as presenters in CS1/CS2 laboratories. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):164–168, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roberts:2002:EWC

[RKI02] Eric S. Roberts, Marina Kassianidou, and Lilly Irani. Encouraging women in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (2):84–88, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/

DBLP/2002.bib. Women and Computing.

Rossling:2007:EFF

[RKS07] Guido Rößling, Simon Kulessa, and Silke Schneider. Easy, fast, and flexible algorithm animation generation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):357, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Rossling:2009:VCA

[RL09] Guido Rößling and Florian Lindner. Visualizing compression algorithms on-the-fly. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):376, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Repp:2008:QAL

[RLM08] Stephan Repp, Serge Linckels, and Christoph Meinel. Question answering from lecture videos based on an automatic semantic annotation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):17–21, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Rodger:2007:IIS

[RLR07] Susan H. Rodger, Jinghui Lim, and Stephen Reading. Increasing interaction and support in the formal languages and automata theory course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):58– 62, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Ramalingam:2004:SEM

[RLW04] Vennila Ramalingam, Deborah LaBelle, and Susan Wiedenbeck. Self-efficacy and mental models in learning to program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):171–175, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reis:2001:AEF

[RM01] Rogério Reis and Nelma Moreira. Apoo: an environment for a first course in assembly language programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(4):43–47, December 2001. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2001.bib.

Rafieymehr:2007:JVD

[RM07] Ali Rafieymehr and Richard McKeever. Java visual debugger. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):75–79, June 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Rao:2008:ESE

[RM08a] T. M. Rao and Sandeep Mitra. An early software engineering approach to teaching CS1, CS2 and AI. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):143–147, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Reyes:2008:GDJ

[RM08b] Miguel Reyes and Águeda Mata. Games developed in Java for teaching “combinatorial game theory”. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):359, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Rossling:2009:SES

[RM09] Guido Rößling and Marius Müller. Social engineering: a serious underestimated problem. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):384, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Reed:2000:EIT

[RMB00] David Reed, Craig Miller, and Grant Braught. Empirical investigation throughout the CS curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):202–206, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rickman:2001:ECN

[RMMH01] Jon Rickman, Merry McDonald, Gary McDonald, and Phillip Heeler. Enhancing the computer networking curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):157–160, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Russell:2006:TAT

[RMN06] Ingrid Russell, Zdravko Markov, and Todd Neller. Teaching AI through machine learning projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):323, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Riedesel:2009:MAE

[RMPD09] Charles P. Riedesel, Eric D. Manley, Susan Poser, and Jitender S. Deogun. A model academic ethics and integrity policy for computer science departments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (1):357–361, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Ritter:2009:ISP

[RMT09] Nicola Ritter, Tanya McGill, and Nik Thompson. Incremental submission of programming code using object-oriented classes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4):66–70, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Repp:2008:MSG

[RMY08] Stephan Repp, Christoph Meinel, and Sevil Yakhyayeva. Motivation of the students in game development projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):368, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Ranum:2006:SAT

[RMZG06] David Ranum, Bradley Miller, John Zelle, and Mark Guzdial. Successful approaches to teaching introductory computer science courses with Python. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):396– 397, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rossling:2002:TPR

[RN02] Guido Rößling and Thomas L. Naps. A testbed for pedagogical requirements in algorithm visualizations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):96–100, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ramakrishnan:2003:DCW

[RN03] Sub Ramakrishnan and Emeka Nwosu. DBMS course: Web based database administration tool and class projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):16–20, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rossling:2006:MIV

[RNH^{+}06] Guido Rößling, Thomas L. Naps, Mark S. Hall, Ville Karavirta, Andreas Kerren, Charles Leska, Andrés Moreno, Rainer Oechsle, Susan H. Rodger, Jaime Urquiza-Fuentes, and J. Ángel Velázquez-Iturbide. Merging interactive visualizations with hypertextbooks and course management. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):166–181, December 2006. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2006.bib.

Robbert:2000:EVP

[Rob00a] Mary Ann Robbert. Enhancing the value of a project in the database course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):36–40, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Robbins:2000:EBB

[Rob00b] Steven Robbins. Experimentation with bounded buffer synchronization. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):330–334, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Robbins:2000:RLJ

[Rob00c] Steven Robbins. Remote logging in Java using Jeli: a facility to enhance development of accessible educational software. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):114–118, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roberts:2000:CEI

[Rob00d] Eric Roberts. Computing education and the information technology workforce. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(2):83–90, June 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roberts:2000:SEI

[Rob00e] Eric Roberts. Strategies for encouraging individual achievement in introductory computer science courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):295–299, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Robbins:2001:SPE

[Rob01a] Steven Robbins. Starving philosophers: experimentation with monitor synchronization. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1): 317–321, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roberts:2001:OM

[Rob01b] Eric Roberts. An overview of MiniJava. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):1–5, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Robbins:2002:EPI

[Rob02] Steven Robbins. Exploration of process interaction in operating systems: a pipe-fork simulator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (1):351–355, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Robbins:2003:URL

[Rob03a] Steven Robbins. Using remote logging for teaching concurrency. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):177–181, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roberts:2003:KAE

[Rob03b] Eric Roberts. Keynote address: Expanding the audience for computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):.23, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Robbins:2004:DHS

[Rob04a] Steven Robbins. A disk head scheduling simulator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):325–329, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roberts:2004:DCL

[Rob04b] Eric Roberts. The dream of a common language: the search for simplicity and stability in computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):115–119, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roberts:2004:RSU

[Rob04c] Eric Roberts. Resources to support the use of Java in introductory computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):233–234, March 2004. CODEN SIGSD3. ISBN 1-58113-798-2. ISSN 0097-8418 (print), 2331-

3927 (electronic). URL ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Compiler/java.
bib;   ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2004.bib.

Robbins:2005:ATS

[Rob05] Steven Robbins. An address translation simulator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):515–519, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Robbins:2006:UCS

[Rob06a] Steven Robbins. A UNIX concurrent I/O simulator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):303–307, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roberts:2006:ITS

[Rob06b] Eric Roberts. An interactive tutorial system for Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):334–338, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Robbins:2007:JES

[Rob07a] Steven Robbins. A Java execution simulator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):536–540, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roberts:2007:RAP

[Rob07b] Eric Roberts. Resurrecting the applet paradigm. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):521–525, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rodger:2002:ICS

[Rod02] Susan H. Rodger. Introducing computer science through animation and virtual worlds. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):186– 190, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Rodger:2006:LAF

[Rod06] Susan Rodger. Learning automata and formal languages interactively with JFLAP. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):360, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roiger:2005:TIC

[Roi05] Richard J. Roiger. Teaching an introductory course in data mining. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):415, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rolfe:2001:BCR

[Rol01] Timothy J. Rolfe. Binomial coefficient recursion: the good, and the bad and ugly. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(2):35– 36, June 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2001.bib.

Rolfe:2002:OTB

[Rol02] Timothy J. Rolfe. One-time binary search tree balancing: the Day/Stout/Warren (DSW) algorithm. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):85–88, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2002.bib.

Rolfe:2003:POE

[Rol03a] Timothy J. Rolfe. Program optimization: enforcement of local access and array access via pointers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):63–65, December 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2003.bib.

Rolfe:2003:SAN

[Rol03b] Timothy J. Rolfe. Spreadsheet-aided numerical experimentation: analytic formula for Fibonacci numbers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):117–119, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2003.bib.

Rolfe:2005:LPS

[Rol05] Timothy J. Rolfe. List processing: sort again, naturally. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):46–48, June 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2005.bib.

Rolfe:2006:CED

[Rol06a] Timothy J. Rolfe. Classroom exercise demonstrating linked list operations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):83–84, December 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2006.bib.

Rolfe:2006:VDQ

[Rol06b] Timothy J. Rolfe. Las Vegas does n-queens. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):37–38, June 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2006.bib.

Rolfe:2007:ADP

[Rol07] Timothy J. Rolfe. An alternative dynamic programming solution for the 0/1 knapsack. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):54–56, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2007.bib.

Rolfe:2008:PFO

[Rol08a] Timothy J. Rolfe. Perverse and foolish oft I strayed. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):52–55, June 2008. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2008.bib.

Rolfe:2008:SMA

[Rol08b] Timothy J. Rolfe. A specimen MPI application: N-queens in parallel. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):42–45, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rolfe:2009:APE

[Rol09a] Timothy J. Rolfe. The assignment problem: exploring parallelism. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):127–131, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rolfe:2009:APF

[Rol09b] Timothy J. Rolfe. The assignment problem: further exploring parallelism. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4):78–81, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ross:2000:MTC

[Ros00] John Minor Ross. Multimedia: from topic to course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):343–346, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rosbottom:2001:HLS

[Ros01] John Rosbottom. Hybrid learning — a safe route into Webbased open and distance learning for the computer science teacher. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):89–92, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ross:2002:GST

[Ros02] John Minor Ross. Guiding students through programming puzzles: Value and examples of Java game assignments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):94–98, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Compiler/java.bib;   ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/2002.
bib.

Rosenthal:2005:IRU

[Ros05] Tammy Rosenthal. Introducing recursion by using multimedia. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):374, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rosmaita:2006:AFN

[Ros06] Brian J. Rosmaita. Accessibility first!: a new approach to Web design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):270–274, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rosmaita:2007:MSL

[Ros07] Brian J. Rosmaita. Making service learning accessible to computer scientists. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):541–545, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ross:2008:HHA

[Ros08] Rockford J. Ross. Hypertextbooks and a Hypertextbook authoring environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):133–137, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Rosenbloom:2009:RPC

[Ros09] Arnold Rosenbloom. Running a programming contest in an introductory computer science course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):347, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Rossling:2006:TPI

[Röß06] Guido Rößling. Translator: a package for internationalization for Java-based applications and GUIs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):312, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rossling:2008:PSI

[Röß08] Guido Rößling. Providing a Seminar++: innovation seminars. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):312, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Roumani:2002:DGL

[Rou02] Hamzeh Roumani. Design guidelines for the lab component of objects-first CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):222–226, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Roussos:2004:TGF

[Rou04] Constantine Roussos. Teaching growth of functions using equivalence classes: an alternative to big O notation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):170–174, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roumani:2006:PWY

[Rou06] Hamzeh Roumani. Practice what you preach: full separation of concerns in CS1/CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):491– 494, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rolfe:2004:APB

[RP04] Timothy J. Rolfe and Paul W. Purdom. An alternative problem for backtracking and bounding. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(4):83–84, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2004.bib.

Randall:2006:DAC

[RP06] Cindy H. Randall and Barbara A. Price. Desirable attributes for computing graduates: is there agreement? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):91–96, June 2006. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2006.bib.

Rich:2004:CCD

[RPG04] Lauren Rich, Heather Perry, and Mark Guzdial. A CS1 course designed to address interests of women. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):190–194, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rowell:2003:CRG

[RPH^{+}03] Ginger Holmes Rowell, Diane G. Perhac, Judith A. Hankins, Brenda C. Parker, Chrisila C. Pettey, and Judith M. Iriarte-Gross. Computer-related gender differences. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):54–58, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Randall:2003:WCP

[RPR03] Cindy H. Randall, Barbara Price, and Han Reichgelt. Women in computing programs: does the incredible shrinking pipeline apply to all computing programs? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):55–59, December 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2003.bib.

Robbert:2003:TED

[RR03] Mary Ann Robbert and Catherine M. Ricardo. Trends in the evolution of the database curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):139–143, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Robila:2006:DPS

[RR06] Stefan A. Robila and James W. Ragucci. Don’t be a phish: steps in user education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):237– 241, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rusu:2009:AAI

[RRD^{+}09] Adrian Rusu, Amalia Rusu, Rebecca Docimo, Confesor Santiago, and Mike Paglione. Academia-academia-industry collaborations on software engineering projects using local-remote teams. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):301–305, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Rauchas:2006:LPH

[RRKS06] Sarah Rauchas, Benjamin Rosman, George Konidaris, and Ian Sanders. Language performance at high school and success in first year computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (1):398–402, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Raab:2000:PPT

[RRP00] Jeff Raab, Richard Rasala, and Viera K. Proulx. Pedagogical power tools for teaching Java. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3): 156–159, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rasala:2001:JPT

[RRP01] Richard Rasala, Jeff Raab, and Viera K. Proulx. Java power tools: model software for teaching object-oriented design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):297–301, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rasala:2002:SMD

[RRP02] Richard Rasala, Jeff Raab, and Viera K. Proulx. The SIGCSE 2001 Maze Demonstration program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (1):287–291, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Rountree:2002:PSF

[RRR02] Nathan Rountree, Janet Rountree, and Anthony Robins. Predictors of success and failure in a CS1 course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):121–124, December 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2002.bib.

Rountree:2004:IFP

[RRRH04] Nathan Rountree, Janet Rountree, Anthony Robins, and Robert Hannah. Interacting factors that predict success and failure in a CS1 course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(4):101–104, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2004.bib.

Rozanski:2003:IUE

[RS03] Evelyn P. Rozanski and Nan C. Schaller. Integrating usability engineering into the computer science curriculum: a proposal. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):202–206, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Richards:2004:TWN

[RS04] Brad Richards and Benjamin Stull. Teaching wireless networking with limited resources. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1): 306–310, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rubio-Sanchez:2008:IPE

[RS08] Manuel Rubio-Sánchez. An introduction to problem equivalence with combinatorics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):313, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Ronchetti:2009:TAS

[RS09] Marco Ronchetti and Joseph Sant. Towards automatic syllabi matching. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):379, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Rossling:2000:AAA

[RSF00] Guido Rößling, Markus Schüer, and Bernd Freisleben. The ANIMAL algorithm animation tool. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (3):37–40, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rubio-Sanchez:2007:ERF

[RSHL07] Manuel Rubio-Sánchez and Isidoro Hernán-Losada. Exploring recursion with Fibonacci numbers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):359, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Rauchas:2006:EPP

[RSK06] Sarah Rauchas, Ian Sanders, and Benjamin Kumwenda. The effect of prior programming experience in a Scheme-based breadth-first curriculum at Wits. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (3):326, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rubio-Sanchez:2008:GIM

[RSUFPF08] Manuel Rubio-Sánchez, Jaime Urquiza-Fuentes, and Cristóbal Pareja-Flores. A gentle introduction to mutual recursion. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):235–239, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Rubio-Sanchez:2009:TRU

[RSVI09] Manuel Rubio-Sánchez and J. Ángel Velázquez-Iturbide. Tail recursion by using function generalization. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):394, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Ratcliffe:2003:CCD

[RTET03] Mark Ratcliffe, Lynda Thomas, Wayne Ellis, and Benjy Thomasson. Capturing collaborative designs to assist the pedagogical process. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):79–83, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rui:2008:ITT

[RTH^{+}08] Deng Rui, John T. Thompson, Yang Hong, Zhou Xing-sheng, Liu Ke-jing, and Neil Alexander Macintyre. Imagery training in the teaching of the data structure curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):92–94, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rossling:2004:ECL

[RTM^{+}04] Guido Rößling, Christoph Trompler, Max Mühlhäuser, Susanne Köbler, and Susanne Wolf. Enhancing classroom lectures with digital sliding blackboards. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):218–222, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ruckert:2007:TCC

[Ruc07] Martin Ruckert. Teaching compiler construction and language design: making the case for unusual compiler projects with PostScript as the target language. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):435–439, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Russell:2009:ACC

[Rus09] Ingrid Russell. Analyzing CS competencies using the SOLO taxonomy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):1, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Rutherfoord:2001:UPI

[Rut01] Rebecca H. Rutherfoord. Using personality inventories to help form teams for software engineering class projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):73–76, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Romero:2008:TDC

[RVP08] Miguel Romero, Aurora Vizcaíno, and Mario Piattini. Toward a definition of the competences for global requirements elicitation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):364, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Rountree:2004:PSS

[RVWB04] Nathan Rountree, Tamar Vilner, Brenda Cantwell Wilson, and Roger Boyle. Predictors for success in studying CS. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):145–146, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Richards:2002:INCa

[RW02] Brad Richards and Nate Waisbrot. Illustrating networking concepts with wireless handheld devices. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):29–33, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reid:2005:LDI

[RW05] Karen L. Reid and Gregory V. Wilson. Learning by doing: introducing version control as a way to manage student assignments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):272–276, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Reid:2007:DSP

[RW07] Karen L. Reid and Gregory V. Wilson. DrProject: a software project management portal to meet educational needs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):317–321, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Robbert:2000:DCP

[RWGM00] Mary Ann Robbert, Ming Wang, Mario Guimaraes, and Martha E. Myers. The database course (panel session): what must be taught. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):403–404, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Rodger:2009:IEA

[RWL^{+}09] Susan H. Rodger, Eric Wiebe, Kyung Min Lee, Chris Morgan, Kareem Omar, and Jonathan Su. Increasing engagement in automata theory with JFLAP. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1): 403–407, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Redmond:2003:EIM

[RWP03] James A. Redmond, Cathal Walsh, and Adrian Parkinson. Equilibrating instructional media for cognitive styles. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):55–59, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ryan:2001:HCI

[Rya01] Christopher D. Ryan. The human-computer interface: challenges for educational multimedia and Web designers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(4):51–54, December 2001. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2001.bib.

Saad:2007:VRS

[SA07] Mohamed K. Saad and Syed V. Ahamed. Vulnerabilities of RFID systems in infant abduction protection and patient wander prevention. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):160– 165, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2007.bib.

Saeli:2009:HTP

[Sae09] Mara Saeli. How to teach programming in secondary education: first results of a PhD project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (3):356, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Simon:2004:PET

[SAHS04] Beth Simon, Ruth Anderson, Crystal Hoyer, and Jonathan Su. Preliminary experiences with a tablet PC based system to support active learning in computer science courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):213–217, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Salcedo:2003:FCS

[Sal03a] Michaelangelo Salcedo. Faculty and the 21^{s}t century student in USA higher education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):83– 87, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2003.bib.

Salcedo:2003:FSC

[Sal03b] Michaelangelo Salcedo. Faculty and the 21st century student in USA higher education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):83– 87, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Samaka:2002:CCS

[Sam02] Mohammed Samaka. Changing a computer science curriculum in light of computing curricula 2001. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):32–35, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2002.bib.

Samaka:2005:UFP

[Sam05] M. Samaka. Using a faculty portfolio in the distinction of teaching. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):31–35, June 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2005.bib.

Sturgeon:2009:WEV

[SAM09] Thomas Sturgeon, Colin Allison, and Alan Miller. 802.11 wireless experiments in a virtual world. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):85–89, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Sanders:2002:TEA

[San02] Ian Sanders. Teaching empirical analysis of algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):321–325, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Sanders:2005:DFC

[San05] Alton F. Sanders. A discussion format for computer ethics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):352–355, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sanders:2008:IPB

[San08] Norman Sanders. An industry perspective on the beginnings of CAD. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):128–134, June 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2008.bib.

Sanderson:2009:RAP

[San09a] Donald B. Sanderson. Revising an assessment plan to conform to the new ABET–CAC guidelines. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):352–356, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Sant:2009:MIE

[San09b] Joseph A. Sant. “mailing it in”: email-centric automated assessment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):308–312, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Saravanos:2008:PBC

[Sar08] Antonios Saravanos. Potential benefits of corrected-errors in AVBL. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):374, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Savidis:2005:ITI

[Sav05] Anthony Savidis. Introductory teaching of imperative programming through an anthropomorphic computation model. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):49–53, June 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2005.bib.

Saxon:2003:OOR

[Sax03] Charles S. Saxon. Object-oriented recursive descent parsing in C#. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):82–85, December 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

Sudol:2006:TST

[SB06] Leigh Ann Sudol and Kathryn Bambino. Is there such a thing as too much support?: a discussion from a teacher and a student’s perspective of support groups for girls in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):309, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schweitzer:2007:IVA

[SB07a] Dino Schweitzer and Wayne Brown. Interactive visualization for the active learning classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (1):208–212, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sturm:2007:RCA

[SB07b] D. Sturm and R. S. Beiss. A root cause analysis interface for error reporting. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):166– 170, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2007.bib.

Schlueter:2008:CDC

[SB08] Kirsten Schlueter and Torsten Brinda. Characteristics and dimensions of a competence model of theoretical computer science in secondary education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):367, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Schlueter:2009:ATD

[SB09] Kirsten Schlueter and Torsten Brinda. The attributes of task difficulty in informatics in secondary education: first results of an empirical study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):357, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Sanders:2008:SUO

[SBE^{+}08] Kate Sanders, Jonas Boustedt, Anna Eckerdal, Robert Mc-Cartney, Jan Erik Moström, Lynda Thomas, and Carol Zander. Student understanding of object-oriented programming as expressed in concept maps. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1): 332–336, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Schnabel:2008:ODC

[SBG^{+}08] Robert Schnabel, Duncan Buell, Joanna Goode, J. Strother Moore, and Chris Stephenson. An open dialogue concerning the state of education policy in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):114–115, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Smith:2007:TPC

[SBLH07] Wanda J. Smith, France Bélanger, Tracy L. Lewis, and Kristi Honaker. Training to persist in computing careers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):119–120, December 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Stamouli:2005:EES

[SBM05] Ioanna Stamouli, Marjahan Begum, and Rebecca Mancy. ExploreCSEd: exploring skills and difficulties in programming education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):371, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stallmann:2007:PAE

[SBR^{+}07] Matthias F. Stallmann, Suzanne P. Balik, Robert D. Rodman, Sina Bahram, Michael C. Grace, and Susan D. High. ProofChecker: an accessible environment for automata theory correctness proofs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):48– 52, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Starr:2009:DIC

[SBZ09] Christopher W. Starr, Doug Bergman, and Phil Zaubi. The development and implementation of a context-based curricular framework for computer science education in high schools. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):283–287, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Sarkar:2004:ICH

[SC04] Nurul I. Sarkar and Trevor M. Craig. Illustrating computer hardware concepts using PIC-based projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):270–274, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sarkar:2007:LCP

[SC07] Nurul I. Sarkar and Trevor M. Craig. A low-cost PIC unit for teaching computer hardware fundamentals to undergraduates. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):88–91, June 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Subramanian:2008:CDV

[SC08] K. R. Subramanian and T. Cassen. A cross-domain visual learning engine for interactive generation of instructional materials. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):488–492, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Scanlan:2000:SPM

[Sca00] David A. Scanlan. Student preference for multimedia-based lectures (poster session): a preliminary report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):192, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shackelford:2004:CCO

[SCC^{+}04] Russell Shackelford, Lillian Cassel, James Cross, John Impagliazzo, Eydie Lawson, Richard LeBlanc, Andrew McGettrick, Robert Sloan, and Heikki Topi. Computing curricula 2004: the overview project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1): 501, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shaffer:2007:AVR

[SCE07] Clifford A. Shaffer, Matthew Cooper, and Stephen H. Edwards. Algorithm visualization: a report on the state of the field. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):150–154, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schaub:2000:TJG

[Sch00] Stephen Schaub. Teaching Java with graphics in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(2):71–73, June 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schneider:2002:NMR

[Sch02a] G. Michael Schneider. A new model for a required senior research experience. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):48–51, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2002.bib.

Schreiner:2002:JTT

[Sch02b] Wolfgang Schreiner. A Java toolkit for teaching distributed algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):111–115, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-

3927 (electronic). URL ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Compiler/java.
bib.

Schneider:2004:MTC

[Sch04] G. Michael Schneider. A model for a three course introductory sequence. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(2):40–43, June 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2004.bib.

Schaeffer:2007:GCP

[Sch07] Jonathan Schaeffer. The games computers play…: perfectly. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):581, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schaub:2009:TCW

[Sch09a] Stephen Schaub. Teaching CS1 with Web applications and test-driven development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):113– 117, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schocken:2009:VMA

[Sch09b] Shimon Schocken. Virtual machines: abstraction and implementation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):203–207, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Salter:2009:NMC

[SCHS09] Richard M. Salter, Robert D. Cupper, Stuart Hirshfield, and Alexa M. Sharp. New models for the CS1 course: a fifteen year retrospective. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):335–336, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Sciore:2007:SSJ

[Sci07] Edward Sciore. SimpleDB: a simple Java-based multiuser syst for teaching database internals. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (1):561–565, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sheard:2008:GSA

[SCL^{+}08] Judy Sheard, Angela Carbone, Raymond Lister, Beth Simon, Errol Thompson, and Jacqueline L. Whalley. Going SOLO to assess novice programmers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):209– 213, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Scott:2002:MMI

[Sco02] Kirk Scott. MISC: the minimal instruction set computer. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):223, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Scott:2003:TGI

[Sco03] Kirk Scott. Teaching graphical interface programming in Java with the Game of Wari. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):254, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Compiler/
java.bib.

Satrazemi:2000:TEE

[SD00] M. Satrazemi and V. Dajdiielis. Telemachus an effective electronic marker of students’ programming assignments (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):191–192, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sanders:2003:JTI

[SD03a] Dean Sanders and Brian Dorn. Jeroo: a tool for introducing object-oriented programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1): 201–204, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sheard:2003:ICP

[SD03b] Judy Sheard and Martin Dick. Influences on cheating practice of graduate students in IT courses: what are the factors? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):45–49, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sivilotti:2003:IMS

[SD03c] Paolo A. G. Sivilotti and Murat Demirbas. Introducing middle school girls to fault tolerant computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):327–331, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Salter:2008:UDS

[SD08] Richard M. Salter and John L. Donaldson. Using DLSim 3: a scalable, extensible, multi-level logic simulator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):315, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Salter:2009:AED

[SD09] Richard M. Salter and John L. Donaldson. Abstraction and extensibility in digital logic simulation software. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):418–422, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Satratzemi:2001:SPV

[SDE01] Maria Satratzemi, Vassilios Dagdilelis, and Georgios Evagelidis. A system for program visualization and problem-solving path assessment of novice programmers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):137–140, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Simon:2008:NTN

[SDG^{+}08] Beth Simon, Krista Davis, William G. Griswold, Michael Kelly, and Roshni Malani. Noteblogging: taking note taking public. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):417–421, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Smarkusky:2005:ETK

[SDLdQ05] D. Smarkusky, R. Dempsey, J. Ludka, and F. de Quillettes. Enhancing team knowledge: instruction vs. experience. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):460–464, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sheard:2002:CPP

[SDM^{+}02] Judy Sheard, Martin Dick, Selby Markham, Ian Macdonald, and Meaghan Walsh. Cheating and plagiarism: perceptions and practices of first year IT students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):183–187, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Skupas:2009:DCC

[SDR09] Bronius Skupas, Valentina Dagiene, and Miguel Revilla. Developing classification criteria for programming tasks.

SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):373, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Sweedyk:2005:CGC

[SdSK05] Elizabeth Sweedyk, Marianne deLaet, Michael C. Slattery, and James Kuffner. Computer games and CS education: why and how. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):256–257, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Seeling:2008:L

[See08] Patrick Seeling. Labs@home. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4): 75–77, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Seidman:2009:AFI

[Sei09] Robert H. Seidman. Alice first: 3D interactive game programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):345, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Skevoulis:2002:IFM

[SF02] Sotiris Skevoulis and Maria Falidas. Integrating formal methods tools into undergraduate computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):232, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stuurman:2004:ETD

[SF04] Sylvia Stuurman and Gert Florijn. Experiences with teaching design patterns. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):151–155, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Spiegel:2008:IIT

[SFD08] Daniel S. Spiegel, Lisa M. Frye, and Linda L. Day. Issues in the instantiation of template classes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (2):48–51, June 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2008.bib.

Simon:2007:DAN

[SFM^{+}07] Beth Simon, Sue Fitzgerald, Renée McCauley, Susan M. Haller, John Hamer, Brian Hanks, Michael T. Helmick, Jan Erik Moström, Judy Sheard, and Lynda Thomas. Debugging assistance for novices: a video repository. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):137–151, December 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Sierra:2008:ESA

[SFPFV08] José-Luis Sierra, Ana M. Fernández-Pampillon, and Alfredo Fernández-Valmayor. An environment for supporting active learning in courses on language processing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):128–132, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Sanders:2007:SMM

[SG07] Ian Douglas Sanders and Vashti C. Galpin. Students’ mental models of recursion at Wits. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3): 317, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Stasko:2001:MAC

[SGC^{+}01] John Stasko, Mark Guzdial, Mike Clancy, Nell Dale, and Sally Fincher. Models and areas for CS education research. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):388–389, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sanders:2006:MMR

[SGG06] Ian Sanders, Vashti Galpin, and Tina Götschi. Mental models of recursion revisited. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):138–142, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stewart-Gardiner:2001:CVP

[SGKL^{+}01] Carolee Stewart-Gardiner, David G. Kay, Joyce Currie Little, Joseph D. Chase, John Fendrich, Laurie A. Williams, and Ursula Wolz. Collaboration vs plagiarism in computer science programming courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1): 406–407, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Said:2009:DDA

[SGMJ09] Huwida E. Said, Mario A. Guimaraes, Zakaria Maamar, and Leon Jololian. Database and database application security. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):90–93, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Stolin:2007:SUC

[SH07] Yuila Stolin and Orit Hazzan. Students’ understanding of computer science soft ideas: the case of programming paradigm. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):65–69, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Schreiner:2008:SLL

[SH08] Axel-Tobias Schreiner and James E. Heliotis. Sudoku: a little lesson in OOP. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):44– 47, June 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2008.bib.

Shaffer:2000:IBD

[Sha00] Dale Shaffer. Internet-based distance learning: a multicontinental perspective. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(4):22–23, December 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shay:2002:MMC

[Sha02] William Shay. A multiplatform/multilanguage client/server project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):401–405, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Shannon:2003:ABF

[Sha03] Christine Shannon. Another breadth-first approach to CS I using Python. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):248–251, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shaffer:2004:BPF

[Sha04] Clifford A. Shaffer. Buffer pools and file processing projects for an undergraduate data structures course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):175–178, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shaffer:2005:LOP

[Sha05] Steven C. Shaffer. Ludwig: an online programming tutoring and assessment system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):56– 60, June 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2005.bib.

Shaffer:2006:ETG

[Sha06] Clifford A. Shaffer. Experiences teaching a graduate research methods course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):97– 101, June 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2006.bib.

Shene:2002:TST

[She02] Ching-Kuang Shene. ThreadMentor: a system for teaching multithreaded programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):229, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shelburne:2003:TCO

[She03a] Brian Shelburne. Teaching computer organization using a PDP-8 simulator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):69–73, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sherstov:2003:DVG

[She03b] Alexander A. Sherstov. Distributed visualization of graph algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):376–380, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sheard:2004:ELC

[She04a] Judy Sheard. Electronic learning communities: strategies for establishment and management. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (3):37–41, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sher:2004:SIQ

[She04b] David B. Sher. A simple implementation of a queue with a circularly linked list. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):274, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sher:2006:MDS

[She06] David B. Sher. Motivating data structures with caching Internet stock data. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):344, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sheehan:2007:TOS

[She07] Robert J. Sheehan. Teaching operating systems with Ruby. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):38–42, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Sher:2008:VPA

[She08] David B. Sher. A visual proof for an average case of list searching. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):74–78, June 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2008.bib.

Shiflet:2002:CSS

[Shi02] Angela B. Shiflet. Computer science with the sciences: an emphasis in computational science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (4):40–43, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2002.bib.

Shiau:2006:AVS

[Shi06] LieJune Shiau. An application of vector space theory in data transmission. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):33–36, June 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2006.bib.

Stevens:2000:ULP

[SHL^{+}00] K. Todd Stevens, Joel Henry, Pamela B. Lawhead, John Lewis, Constance Bland, and Mary Jane Peters. Using large projects in a computer science curriculum (panel session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):399–400, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shiflet:2000:UCS

[SHN^{+}00] Angela B. Shiflet, Philip Holmes, Chuck Niederriter, Robert M. Panoff, and Ernest Sibert. Undergraduate computational science education (panel session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):419–420, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Spacco:2006:EMD

[SHP^{+}06] Jaime Spacco, David Hovemeyer, William Pugh, Fawzi Emad, Jeffrey K. Hollingsworth, and Nelson Padua-Perez. Experiences with Marmoset: designing and using an advanced submission and testing system for programming courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):13–17, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sturtevant:2008:MSI

[SHS^{+}08] Nathan R. Sturtevant, H. James Hoover, Jonathan Schaeffer, Sean Gouglas, Michael H. Bowling, Finnegan Southey, Matthew Bouchard, and Ghassan Zabaneh. Multidisciplinary students and instructors: a second-year games course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):383–387, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Shumba:2004:TME

[Shu04] Rose Shumba. Towards a more effective way of teaching a cybersecurity basics course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(4): 108–111, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2004.bib.

Shumba:2005:URR

[Shu05] Rose Shumba. Usability of Rational Rose and Visio in a software engineering course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):107– 110, June 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2005.bib.

Shumba:2006:DHC

[Shu06a] Rose Shumba. The development of a Human Computer Interaction course at a senior synthesis course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):102–104, June 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2006.bib.

Shurn:2006:PFI

[Shu06b] Todd Shurn. Programming fundamentals and innovation taught through windows media player skin creation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):364–372, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Simon-Hurtado:2005:LUF

[SHVP05] Ma Aránzazu Simón-Hurtado and Carlos Vivaracho-Pascual. Learning UNIX in first year of computer engineering. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):392, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sitaraman:2009:ESS

[SHW^{+}09] Murali Sitaraman, Jason O. Hallstrom, Jarred White, Svetlana Drachova-Strang, Heather K. Harton, Dana Leonard, Joan Krone, and Rich Pak. Engaging students in specification and reasoning: “hands-on” experimentation and evaluation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):50–54, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Sicilia:2009:HST

[Sic09] Miguel-Angel Sicilia. How should transversal competence be introduced in computing education? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (4):95–98, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Siegel:2000:IIA

[Sie00] Eric V. Siegel. Iambic IBM AI: the palindrome discovery AI project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):95–99, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Siegfried:2006:VPB

[Sie06] Robert M. Siegfried. Visual programming and the blind: the challenge and the opportunity. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1): 275–278, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sigman:2007:ESF

[Sig07] Scott Sigman. Engaging students in formal language theory and theory of computation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):450– 453, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shukla:2002:CTO

[SJ02] Ranjana Shukla and Donald Joyce. Conferencing technologies and online courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):235, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sitthiworachart:2004:EPA

[SJ04] Jirarat Sitthiworachart and Mike Joy. Effective peer assessment for learning computer programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):122–126, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Standish:2005:UPP

[SJ05] Thomas A. Standish and Norman Jacobson. Using O(n) ProxmapSort and O(1) ProxmapSearch to motivate CS2 students (Part I). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4):41– 44, December 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2005.bib;   ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/SE/alspaugh.bib.

Standish:2006:UPP

[SJ06] Thomas A. Standish and Norman Jacobson. Using O(n) ProxmapSort and O(1) ProxmapSearch to motivate CS2 students, Part II. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):29–32, June 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2006.bib.

Soh:2004:ABC

[SJA04] Leen-Kiat Soh, Hong Jiang, and Charles Ansorge. Agentbased cooperative learning: a proof-of-concept experiment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):368–372, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shanmugasundaram:2006:KBU

[SJH06] Vijayakumar Shanmugasundaram, Paul Juell, and Curt Hill. Knowledge building using visualizations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):23–27, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Subhlok:2007:TPV

[SJS^{+}07] Jaspal Subhlok, Olin Johnson, Venkat Subramaniam, Ricardo Vilalta, and Chang Yun. Tablet PC video based hybrid coursework in computer science: report from a pilot project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):74–78, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shah:2002:TSP

[SK02] Harsh Shah and Amruth N. Kumar. A tutoring system for parameter passing in programming languages. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):170–174, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sweedyk:2005:FGN

[SK05] Elizabeth Sweedyk and Robert M. Keller. Fun and games: a new software engineering course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37 (3):138–142, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Saravanos:2009:PEE

[SK09] Antonios Saravanos and Charles Kinzer. The positive effects of explanation after CE in AVBL. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (3):391, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Skadron:2002:MSC

[Ska02] Kevin Skadron. A microprocessor survey course for learning advanced computer architecture. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (1):152–156, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Stewart:2002:MAF

[SKM02] Avare’ Stewart, Parviz Kermani, and Magda Mourad. MediaMime: after-the-fact authoring annotation system for an e-learning environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):243, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Skoric:2007:SSA

[Sko07] Miroslav S. Skoric. Summer schools on the amateur radio computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):346, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Summet:2009:PCR

[SKO^{+}09] Jay Summet, Deepak Kumar, Keith O’Hara, Daniel Walker, Lijun Ni, Doug Blank, and Tucker Balch. Personalizing CS1 with robots. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):433–437, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Shene:2000:CGU

[SL00a] Ching-Kuang Shene and John Lowther. Computing with geometry as an undergraduate course (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):193, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shene:2000:DPS

[SL00b] Ching-Kuang Shene and John Lowther. DesignMentor (poster session): a pedagogical tool for graphics and computer-aided design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):193, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sobel:2003:ICA

[SL03] Ann E. K. Sobel and Tim Lethbridge. IEEE-CS/ACM computing curricula: software engineering volume. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):217–218, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sattar:2006:DSM

[SL06] Abdul Sattar and Torben Lorenzen. Develop a shopping mart Web application. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):68–70, December 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2006.bib.

Sattar:2007:CAS

[SL07a] Abdul Sattar and Torben Lorenzen. Computer applications: a service course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):105–107, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2007.bib.

Sattar:2007:DCJ

[SL07b] Abdul Sattar and Torben Lorenzen. Develop a compiler in Java for a compiler design course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (2):80–82, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2007.bib.

Sanders:2008:SPP

[SL08a] Ian D. Sanders and Sasha Langford. Students’ perceptions of Python as a first programming language at Wits. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):365, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Sivilotti:2008:SSA

[SL08b] Paolo A. G. Sivilotti and Stacey A. Laugel. Scratching the surface of advanced topics in software engineering: a workshop module for middle school students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):291–295, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Santore:2009:UWC

[SL09a] John Santore and Torben Lorenzen. Use writing class techniques to create software design documents. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):136–137, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sattar:2009:TAP

[SL09b] Abdul Sattar and Torben Lorenzen. Teach Alice programming to non-majors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):118–121, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Santore:2009:SEC

[SLC^{+}09] John Santore, Torben Lorenzen, Robert Creed, David Murphy, and Roger Orcutt. The software engineering class builds a GUI for subversion. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4):82–84, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sherrell:2007:NDG

[SLPG07] Linda B. Sherrell, Chang Liu, William M. Pottenger, and Paul Gross. NSF-DGE GK-12 teaching fellowships: changing student perceptions about computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):474–475, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sloan:2002:NCB

[SLSM02] Joseph D. Sloan, Andy Lopez, Randy K. Smith, and Dick Mowe. Networking CS: beyond the first course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):281–282, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Sanders:2000:FBC

[SM00a] Ian Sanders and Conrad Mueller. A fundamentals-based curriculum for first year computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):227–231, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Siltaneva:2000:NED

[SM00b] Jarmo Siltaneva and Erkki Mäkinen. A note on the expected distribution of degrees in random binary trees. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(4):32–33, December 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schep:2002:EUR

[SM02a] Madeleine Schep and Nieves McNulty. Experiences with using robots in an all-female programming class. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):241, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sheard:2002:CII

[SM02b] Judy Sheard and Selby Markham. Creating an interest in IT: a gender study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):236, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Surakka:2002:WEV

[SM02c] Sami Surakka and Lauri Malmi. Work experience vs. cooperative education program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4): 44–47, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2002.bib.

Sanders:2003:PAT

[SM03] Kathryn E. Sanders and Robert McCartney. Program assessment tools in computer science: a report from the trenches. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):31–35, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stelios:2005:HAP

[SM05a] Xinogalos Stelios and Satratzemi Maya. The hands-on activities of the programming microworld objectKarel. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):384, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stelios:2005:UHA

[SM05b] Xinogalos Stelios and Satratzemi Maya. Using hands-on activities for motivating students with OOP concepts before they are asked to implement them. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37 (3):380, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stone:2007:IRW

[SM07] Jeffrey A. Stone and Elinor M. Madigan. Integrating reflective writing in CS/IS. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):42– 45, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2007.bib.

Stone:2008:IPP

[SM08] Jeffrey A. Stone and Elinor M. Madigan. The impact of providing project choices in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2): 65–68, June 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2008.bib.

Small:2007:DER

[Sma07] Margot Small. Design error and reusability. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):185–187, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2007.bib.

Stone:2009:ECH

[SMF09] Jeffrey A. Stone, Darcy L. Medica, and Leah Ann Fetsko. Experiences with a CS1 for the health sciences. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):122–126, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stern:2005:YCL

[SMH05] Linda Stern, Selby Markham, and Ria Hanewald. You can lead a horse to water: how students really use pedagogical software. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):246–250, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Saikkonen:2001:FAA

[SMK01] Riku Saikkonen, Lauri Malmi, and Ari Korhonen. Fully automatic assessment of programming exercises. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):133–136, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stepp:2009:CIW

[SMK09] Marty Stepp, Jessica Miller, and Victoria Kirst. A “CS 1.5” introduction to Web programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (1):121–125, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Sheard:2003:WBD

[SMR03] Judy Sheard, Jan Miller, and Sita Ramakrishnan. Web-based discussion forums: the staff perspective. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):158–162, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shackelford:2006:CCO

[SMS^{+}06] Russell Shackelford, Andrew McGettrick, Robert Sloan, Heikki Topi, Gordon Davies, Reza Kamali, James Cross, John Impagliazzo, Richard LeBlanc, and Barry Lunt. Computing Curricula 2005: The overview report. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):456–457, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Starr:2008:BTR

[SMS08] Christopher W. Starr, Bill Manaris, and RoxAnn H. Stalvey. Bloom’s taxonomy revisited: specifying assessible learning objectives in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1): 261–265, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Sorkin:2000:CDI

[SMTH00] Sylvia Sorkin, Barbara Mento, Donna Tupper, and Kathleen Harmeyer. Curriculum development in Internet and multimedia technology. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):50–54, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stern:2002:VRR

[SN02] Linda Stern and Lee Naish. Visual representations for recursive algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):196–200, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Sauve:2008:TSD

[SN08] Jacques Philippe Sauvé and Osório Lopes Abath Neto. Teaching software development with ATDD and easyaccept. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):542–546, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Snapp:2006:PFA

[Sna06a] Robert R. Snapp. A Puzzles-First approach to computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):310, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Snapp:2006:TGA

[Sna06b] Robert R. Snapp. Teaching graph algorithms in a corn maze. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):347, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schocken:2009:SCH

[SNA09] Shimon Schocken, Noam Nisan, and Michal Armoni. A synthesis course in hardware architecture, compilers, and software engineering. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):443–447, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Shaban-Nejad:2008:WBD

[SNH08] Arash Shaban-Nejad and Volker Haarslev. Web-based dynamic learning through lexical chaining: a step forward towards knowledge-driven education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (3):375, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Sloan:2007:HSI

[SNS07] David Sloan, Barbara Nelson, and Martin Sloan. How should inclusivity influence teaching of ICT design? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):307–308, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Sobel:2000:ERS

[Sob00] Ann E. Kelley Sobel. Empirical results of a software engineering curriculum incorporating formal methods. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):157–161, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Soh:2004:UGD

[Soh04] Leen-Kiat Soh. Using game days to teach a multiagent system class. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):219–223, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Soh:2006:IJM

[Soh06a] Leen-Kiat Soh. Implementing the jigsaw model in CS1 closed labs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):163–167, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Soh:2006:IIT

[Soh06b] Leen-Kiat Soh. Incorporating an intelligent tutoring system into CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):486–490, March 2006.

CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sojka:2003:AP

[Soj03a] Petr Sojka. Animations in PDF. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35 (3):263, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sojka:2003:ITM

[Soj03b] Petr Sojka. Interactive teaching materials in PDF using JavaScript. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):275, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sojka:2003:REU

[Soj03c] Petr Sojka. Rapid evaluation using multiple choice tests and TeX. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):265, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Soklic:2000:ICP

[Sok00] Milan E. Soklic. Impact of computing platforms on the performance of the asymmetric traveling salesman problem. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(2):79–81, June 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Soklic:2002:SLB

[Sok02] Milan E. Soklic. Simulation of load balancing algorithms: a comparative study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):138–141, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2002.bib.

Solomon:2007:LSA

[Sol07a] Andrew Solomon. Linuxgym: software to automate formative assessment of Unix command-line and scripting skills. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):353, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Solomon:2007:PSC

[Sol07b] Justin Solomon. Putting the science into computer science: treating introductory computer science as the study of algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):46–49, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Sondergaard:2009:LPD

[Son09] Harald Sondergaard. Learning from and with peers: the different roles of student peer reviewing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):31–35, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Sooriamurthi:2000:URT

[Soo00] Raja Sooriamurthi. Using recursion as a tool to reinforce functional abstraction (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):194, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sooriamurthi:2001:PJE

[Soo01a] Raja Sooriamurthi. Prelude to the Java event model (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):193, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sooriamurthi:2001:PCR

[Soo01b] Raja Sooriamurthi. Problems in comprehending recursion and suggested solutions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):25–28, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sooriamurthi:2009:IAD

[Soo09] Raja Sooriamurthi. Introducing abstraction and decomposition to novice programmers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):196– 200, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Sorenson:2005:HCA

[Sor05] Jonathan P. Sorenson. An honors course on Alan M. Turing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4):103–106, December 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2005.bib.

Sadiq:2004:SOS

[SOSL04] Shazia Sadiq, Maria Orlowska, Wasim Sadiq, and Joe Lin. SQLator: an online SQL learning workbench. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):223–227, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stevenson:2003:IOE

[SP03] Daniel E. Stevenson and Andrew T. Phillips. Implementing object equivalence in Java using the template method design pattern. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):278–282, January 2003. CODEN SIGSD3. ISBN 1-58113-648-X. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Compiler/java.bib;   ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/2003.
bib.

Sivilotti:2007:SKL

[SP07] Paolo A. G. Sivilotti and Scott M. Pike. The suitability of kinesthetic learning activities for teaching distributed algorithms. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):362–366, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Spafford:2004:WIS

[Spa04] Gene Spafford. What *is* information security? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):342, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Silvestre:2005:PGV

[SPAC^{+}05] Miguel Silvestre, Maria Pinto-Albuquerque, M. B. Carmo, A. P. Cláudio, J. D. Cunha, and H. Coelho. A platform for the generation of virtual environments inhabited by intelligent virtual humans. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):402, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Spertus:2002:GB

[Spe02] Ellen Spertus. Gender benders. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (2):145–146, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/

DBLP/2002.bib. Women and Computing.

Sheard:2000:SDS

[SPM00] Judy Sheard, Margot Postema, and Selby Markham. Subject differences in student attitudes to paper-based and Webbased resources (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (3):192–193, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Spooner:2000:BSI

[Spo00] David L. Spooner. A Bachelor of Science in Information Technology: an interdisciplinary approach. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):285–289, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sung:2008:GTP

[SPW^{+}08] Kelvin Sung, Michael Panitz, Scott Wallace, Ruth Anderson, and John Nordlinger. Game-themed programming assignments: the faculty perspective. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (1):300–304, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Sheridan-Ross:2007:PTE

[SR07] Jakki Sheridan-Ross. Practical tips for engaging students in team formation for multi-discipline computing projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):360, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Sheridan-Ross:2008:PTC

[SRGF08] Jakki Sheridan-Ross, Andrea Gorra, and Janet Finlay. Practical tips for creating podcasts in higher education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):311, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Springsteel:2000:NDD

[SRR00] Frederick Springsteel, Mary Ann Robbert, and Catherine M. Ricardo. The next decade of the database course: three decades speak to the next. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):41– 45, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Spalter:2000:IIC

[SS00] Anne Morgan Spalter and Rosemary Michelle Simpson. Integrating interactive computer-based learning experiences into established curricula: a case study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (3):116–119, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Scheessele:2006:PGP

[SS06] Michael R. Scheessele and Thomas Schriefer. Poker as a group project for artificial intelligence. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (1):548–552, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stevenson:2007:BAV

[SS07] Daniel E. Stevenson and James D. Schwarzmeier. Building an autonomous vehicle by integrating Lego Mindstorms and a Web cam. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):165–169, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Spradling:2008:ETD

[SSA08] Carol Spradling, Leen-Kiat Soh, and Charles Ansorge. Ethics training and decision-making: do computer science programs need help? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):153–157, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Shang:2000:ATC

[SSC00a] Yi Shang, Hongchi Shi, and Su-Shing Chen. Agent technology in computer science and engineering curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):120–123, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shi:2000:MAS

[SSC00b] Hongchi Shi, Yi Shang, and Su-Shing Chen. A multi-agent system for computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):1–4, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shi:2001:SIC

[SSC01] Hongchi Shi, Yi Shang, and Su-Shing Chen. Smart instructional component based course content organization and delivery (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):186, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Saraiya:2004:EFA

[SSMN04] Purvi Saraiya, Clifford A. Shaffer, D. Scott McCrickard, and Chris North. Effective features of algorithm visualizations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):382–386, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Soh:2005:ARB

[SSP^{+}05a] Leen-Kiat Soh, Ashok Samal, Suzette Person, Gwen Nugent, and Jeff Lang. Analyzing relationships between closed labs and course activities in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):183– 187, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Soh:2005:CLE

[SSP^{+}05b] Leen-Kiat Soh, Ashok Samal, Suzette Person, Gwen Nugent, and Jeff Lang. Closed laboratories with embedded instructional research design for CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1): 297–301, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Soh:2005:DIA

[SSP^{+}05c] Leen-Kiat Soh, Ashok Samal, Suzette Person, Gwen Nugent, and Jeff Lang. Designing, implementing, and analyzing a placement test for introductory CS courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):505–509, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sung:2007:EAG

[SSR07] Kelvin Sung, Peter Shirley, and Becky Reed Rosenberg. Experiencing aspects of games programming in an introductory computer graphics class. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):249– 253, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Smith:2006:IAT

[SSS06] Suzanne Smith, Sara Stoecklin, and Catharina Serino. An innovative approach to teaching refactoring. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):349–353, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stoecklin:2007:TSB

[SSS07] Sara Stoecklin, Suzanne Smith, and Catharina Serino. Teaching students to build well formed object-oriented methods through refactoring. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):145–149, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Soh:2009:RCI

[SSS^{+}09] Leen-Kiat Soh, Ashok Samal, Stephen Scott, Stephen Ramsay, Etsuko Moriyama, George Meyer, Brian Moore, William G. Thomas, and Duane F. Shell. Renaissance computing: an initiative for promoting student participation in computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):59–63, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Spradling:2008:IME

[SSW08] Carol Spradling, Jody Strauch, and Craig Warner. An interdisciplinary major emphasizing multimedia. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):388–391, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Sanders:2007:CGO

[ST07] Kate Sanders and Lynda Thomas. Checklists for grading object-oriented CS1 programs: concepts and misconceptions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):166–170, September 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Sloan:2008:CBA

[ST08] Robert H. Sloan and Patrick Troy. CS 0.5: a better approach to introductory computer science for majors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):271–275, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Stansfield:2005:IVC

[Sta05] Sharon Stansfield. An introductory VR course for undergraduates incorporating foundation, experience and capstone. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):197–200, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Standley:2006:PLP

[Sta06] Hilda M. Standley. ’programming language paradigms’ instruction through designing a new paradigm. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):37–40, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stevens:2001:ETS

[Ste01] K. Todd Stevens. Experiences teaching software engineering for the first time. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):77–80, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Steenkiste:2003:NPC

[Ste03] Peter Steenkiste. A network project course based on network processors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):262–266, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stepney:2004:TPC

[Ste04] Susan Stepney. Teaching post-classical computation: (extended abstract). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):3, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stevenson:2006:PPP

[Ste06] Daniel E. Stevenson. PNG palette permuter. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):143–147, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Steenkiste:2009:UCW

[Ste09a] Peter Steenkiste. The use of a controlled wireless testbed in courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):80–84, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Stephenson:2009:VER

[Ste09b] Ben Stephenson. Visual examples of recursion. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):400, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Stone:2006:UML

[Sto06] Jeffrey A. Stone. Using a machine language simulator to teach CS1 concepts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):43–45, December 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2006.bib.

Stratton:2000:NPS

[Str00] David Stratton. “network protocols and services”: a nonspecialist approach to teaching networking (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):194, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stephenson:2009:QBG

[STS09] Ben Stephenson and Craig Taube-Schock. QuickDraw: bringing graphics into first year. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):211– 215, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Sanchez-Torrubia:2007:NIT

[STTBC07] M. Gloria Sánchez-Torrubia, Carmen Torres-Blanc, and Juan Castellanos. New interactive tools for graph algorithms active learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):337, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Sudol:2008:FCB

[Sud08] Leigh Ann Sudol. Forging connections between life and class using reading assignments: a case study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):357–361, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Suraweera:2001:GMA

[Sur01] Francis Suraweera. Getting the most from an algorithms design course: a personal experience. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33 (4):71–74, December 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2001.bib.

Suraweera:2002:EQL

[Sur02] Francis Suraweera. Enhancing the quality of learning and understanding of first-year mathematics for computer science related majors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):117–120, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2002.bib.

Sutinen:2000:FCP

[Sut00] Erkki Sutinen. Future challenges in program visualization (panel session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):176–177, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sanderson:2000:PAS

[SV00] Pete Sanderson and Ken Vollmar. A primer for applying service learning to computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (1):222–226, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schwarm:2003:MCU

[SV03] Sarah Schwarm and Tammy VanDeGrift. Making connections: using classroom assessment to elicit students’ prior knowledge and construction of concepts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):65–69, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Sivilotti:2004:RTS

[SW04] Paolo A. G. Sivilotti and Bruce W. Weide. Research, teaching, and service: the miniconference as a model for CS graduate seminar courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):487–491, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Schreyer:2006:FAM

[SW06a] Boguslaw Schreyer and Wojciech Wawrzynski. Finite automata models for CS problem with binary semaphore. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):330, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stevenson:2006:DRW

[SW06b] Daniel E. Stevenson and Paul J. Wagner. Developing realworld programming assignments for CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):158–162, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Stevenson:2005:SCI

[SWR05] Daniel E. Stevenson, Michael R. Wick, and Steven J. Ratering. Steganography and cartography: interesting assignments that reinforce machine representation, bit manipulation, and discrete structures concepts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):277– 281, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shin:2000:PTL

[SYLL00] Soo-Bum Shin, In-Hwan Yoo, Chul-Hyun Lee, and Tae-Wuk Lee. Plan of teaching & learning for database software through situated learning (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):193–194, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Taylor:2008:MBS

[TA08] Blair Taylor and Shiva Azadegan. Moving beyond security tracks: integrating security in CS0 and CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):320–324, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Tesser:2000:IMS

[TAHA00] Herbert Tesser, Hisham Al-Haddad, and Gary Anderson. Instrumentation: a multi-science integrated sequence. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):232–236, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Talabis:2006:HLD

[Tal06] Ryan Talabis. Honeynet learning: discovering IT security. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):110–114, June 2006. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2006.bib.

Turner:2007:RMT

[TATL07] Elise H. Turner, Erik Albert, Roy M. Turner, and Laurence Latour. Retaining majors through the introductory sequence. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):24–28, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Taylor:2002:LPL

[Tay02a] R. Gregory Taylor. LL parsing, LR parsing, complexity, and automata. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):71–75, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2002.bib.

Taylor:2002:WCC

[Tay02b] Valerie E. Taylor. Women of color in computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):22–23, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.

math.utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/

Misc/DBLP/2002.bib. Women and Computing.

Taub:2009:ECU

[TBAA09] Rivka Taub, Mordechai Ben-Ari, and Michal Armoni. The effect of CS unplugged on middle-school students’ views of CS. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):99–103, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Tikvati:2004:VTB

[TBAK04] Ahuva Tikvati, Mordechai Ben-Ari, and Yifat Ben-David Kolikant. Virtual trees for the Byzantine Generals algorithm. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):392–396, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tyman:2009:ABS

[TBM09] Damon Tyman, Nirupama Bulusu, and Jens Mache. An activity-based sensor networks course for undergraduates with sun spot devices. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):34–38, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Tirado:2009:ISI

[TBMB09] Felipe Tirado, Alfonso Bustos, Alejandro Miranda, and Ana Elena Del Bosque. Inducing student interaction in a virtual environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):378, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Townsend:2008:GHV

[TBMC08] Gloria Childress Townsend, Lecia Barker, Suzanne Menzel, and J. McGrath Cohoon. Grace Hopper visits the neighborhood. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):513–517, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Truong:2005:LPT

[TBR05] Nghi Truong, Peter Bancroft, and Paul Roe. Learning to program through the Web. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):9–13, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tomer:2000:CCP

[TBS^{+}00] D. S. Tomer, Doug Baldwin, Carl H. Smith, Peter B. Hendersen, and Venu Vadisigi. CS1 and CS2 (panel session): foundations of computer science and discrete mathematics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):397–398, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tabrizi:2009:ICS

[TCK09] M. H. N. Tabrizi, Carol B. Collins, and Vipul Kalamkar. An international collaboration in software engineering. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):306–310, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Tucker:2003:TKC

[TDJ^{+}03] Allen Tucker, Fadi Deek, Jill Jones, Dennis McCowan, Chris Stephenson, and Anita Verno. Toward a K-12 computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):305–306, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Thornton:2008:DTE

[TE08] Matthew Thornton and Stephen H. Edwards. A data type to exploit online data sources. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):114– 118, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Teague:2002:WCW

[Tea02] Joy Teague. Women in computing: what brings them to it, what keeps them in it? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):147– 158, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/

2002.bib. Women and Computing.

Tenenberg:2003:FAT

[Ten03] Josh Tenenberg. A framework approach to teaching data structures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):210–214, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tront:2006:WTC

[TEP06] Joseph G. Tront, Vinod Eligeti, and Jane C. Prey. WriteOn: a tool for classroom presentations on tablet PCs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):352, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Terada:2000:PPS

[Ter00] Minoru Terada. Program paper-slide-show (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):194–195, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Terada:2005:EPT

[Ter05] Minoru Terada. ETV: a program trace player for students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):118–122, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Thornton:2008:SSW

[TETPQ08] Matthew Thornton, Stephen H. Edwards, Roy P. Tan, and Manuel A. Pérez-Quiñones. Supporting student-written tests of GUI programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):537–541, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Talton:2007:TGO

[TF07a] Jerry O. Talton and Darren Fitzpatrick. Teaching graphics with the openGL shading language. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (1):259–263, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tenenberg:2007:ODC

[TF07b] Josh Tenenberg and Sally Fincher. Opening the door of the computer science classroom: the disciplinary commons. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):514–518, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tew:2005:TII

[TFG05] Allison Elliott Tew, Charles Fowler, and Mark Guzdial. Tracking an innovation in introductory CS education from a research university to a two-year college. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):416–420, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tenenberg:2008:PCE

[TFIJ08] Josh Tenenberg, Sally Fincher, John Impagliazzo, and Daniel Joyce. Publishing in computing education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):215–216, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Traynor:2005:SAA

[TG05] Des Traynor and J. Paul Gibson. Synthesis and analysis of automatic assessment methods in CS1: generating intelligent MCQs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):495–499, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Teif:2006:PTO

[TH06] Mariana Teif and Orit Hazzan. Partonomy and taxonomy in object-oriented thinking: junior high school students’ perceptions of object-oriented basic concepts. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):55–60, December 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2006.bib.

Tharp:2007:IIR

[Tha07] Alan L. Tharp. Innovating: the importance of right brain skills for computer science graduates. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):126–130, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Tseng:2009:CUL

[THH09] Judy C. R. Tseng, Sunny Y. Y. Hsu, and Gwo-Jen Hwang. A collaborative ubiquitous learning platform for computer science education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):368, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Thompson:2000:LPM

[Tho00] Errol Thompson. Learning process maturity (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):195, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Thomas:2001:CSS

[Tho01] Pete Thomas. The coach supporting students as they learn to program (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):177, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Thomas:2003:ECI

[Tho03a] P. G. Thomas. Examinations in computing over the Internet. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):22–24, June 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

Thomas:2003:EEM

[Tho03b] Pete Thomas. The evaluation of electronic marking of examinations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):50–54, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Thompson:2006:ARR

[Tho06] Andrew A. Thompson. Approaches to recruiting and retaining in computer-science based student organizations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):85–87, December 2006. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2006.bib.

Tanaka-Ishii:2004:EWB

[TIKT04] Kumiko Tanaka-Ishii, Kazuhiko Kakehi, and Masato Takeichi. EMMA: a Web-based report system for programming course-automated verification and enhanced feedback. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):278, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tucker:2001:OCB

[TKB01] Allen B. Tucker, Charles F. Kelemen, and Kim B. Bruce. Our curriculum has become math-phobic! SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):243–247, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Thomas:2001:OSS

[TL01] Pete Thomas and Kit Logan. Observational studies of student errors in a distance learning environment using a remote recording and replay tool. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):117– 120, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Taylor:2009:PVP

[TLH^{+}09] David Scot Taylor, Andrei F. Lurie, Cay S. Horstmenn, Menko B. Johnson, Sean K. Sharma, and Edward C. Yin. Predictive vs. passive animation learning tools. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):494–498, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Tremblay:2007:EMT

[TLS07] Guy Tremblay, Louise Laforest, and Aziz Salah. Extending a marking tool with simple support for testing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):313, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Tusch:2004:TIS

[TLW^{+}04] Guenter Tusch, Paul Leidig, Greg Wolffe, David Elrod, and Carl Strebel. Technology infrastructure supporting a medical & bioinformatics Masters degree. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (3):264, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tusch:2009:TIS

[TLW^{+}09] Guenter Tusch, Paul Leidig, Gregory Wolffe, David Elrod, and Carl Strebel. Technology infrastructure in support of a medical & bioinformatics Masters degree. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):388, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Traynor:2003:SLM

[TM03] Carol Traynor and Maria McKenna. Service learning models connecting computer science to the community.

SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):43–46, December 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

Thomas:2004:ULD

[TM04] Richard C. Thomas and Rebecca Mancy. Use of large databases for group projects at the nexus of teaching and research. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):161–165, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tucker:2004:ICK

[TMD^{+}04] Allen Tucker, Dennis McCowan, Fadi Deek, Chris Stephenson, Jill Jones, and Anita Verno. Implementation challenges for a K-12 computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):334–335, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Townsend:2007:LCP

[TMS07] Gloria Childress Townsend, Suzanne Menzel, and Katie A. Siek. Leveling the CS1 playing field. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39 (1):331–335, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tremblay:2007:ISP

[TMSZ07] Guy Tremblay, Bruno Malenfant, Aziz Salah, and Pablo Zentilli. Introducing students to professional software construction: a “software construction and maintenance” course and its maintenance corpus. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):176– 180, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Tucker:2002:IFM

[TN02] Allen B. Tucker and Robert E. Noonan. Integrating formal models into the programming languages course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):346–350, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

To:2004:DAO

[To04] L. March To. Determinants affecting organizations’ intent to adopt e-tailing: a study based on innovation theories. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):243, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Topi:2007:RUM

[Top07a] Heikki Topi. Revising the IS undergraduate model curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):10–11, June 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Topi:2007:UMC

[Top07b] Heikki Topi. Update on the IS model curriculum revision project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):15–16, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Topi:2008:ERP

[Top08a] Heikki Topi. IS education: the role of programming in undergraduate IS programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):15–16, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Topi:2008:RIS

[Top08b] Heikki Topi. Role of information systems as a business discipline. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):12–14, June 2008. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2008.bib.

Topi:2009:EAD

[Top09a] Heikki Topi. IS education: accreditation of degree programs in information systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):70–71, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Topi:2009:ERC

[Top09b] Heikki Topi. IS education: The role of IS in computing education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4):12–13, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Townsend:2002:PWM

[Tow02] Gloria Childress Townsend. People who make a difference: mentors and role models. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):57– 61, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/

2002.bib. Women and Computing.

Townsend:2009:UGS

[Tow09] Gloria Childress Townsend. Using a groupware system in CS1 to engage introverted students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (3):16–20, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Taylor:2003:PIP

[TP03] Harriet G. Taylor and Jane C. Prey. Putting ITiCSE into practice through NSF funding programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):234, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Talton:2006:SHC

[TPK^{+}06] Jerry O. Talton, Daniel L. Peterson, Sam Kamin, Deborah Israel, and Jalal Al-Muhtadi. Scavenger hunt: computer science retention through orientation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38 (1):443–447, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Turner:2008:MAO

[TQCPQE08] Scott A. Turner, Ricardo Quintana-Castillo, Manuel A. Pérez-Quiñones, and Stephen H. Edwards. Misunderstandings about object-oriented design: experiences using code reviews. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):97–101, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Trakhtenbrot:2003:ATM

[Tra03] Mark Trakhtenbrot. Analysis of typical misconceptions in a theoretical CS course, and how to address them in e-learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):241, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Traver:2007:CUC

[Tra07] V. Javier Traver. Can user-centered interface design be applied to education? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):57– 61, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2007.bib.

Treu:2002:TUC

[Tre02] Kevin Treu. To teach the unteachable class: an experimental course in Web-based application design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):201–205, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Thomas:2003:CWC

[TRR03] Lynda Thomas, Mark Ratcliffe, and Ann Robertson. Code warriors and code-a-phobes: a study in attitude and pair programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):363–367, January 2003.

CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Thomas:2004:SOD

[TRT04] Lynda Thomas, Mark Ratcliffe, and Benjy Thomasson. Scaffolding with object diagrams in first year programming classes: some unexpected results. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (1):250–254, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Turban:2005:BMB

[TRT05] Georg Turban, Guido Rößling, and Christoph Trompler. Bridging media breaks in presence presentations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):377, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Thomasson:2006:IND

[TRT06] Benjy Thomasson, Mark Ratcliffe, and Lynda Thomas. Identifying novice difficulties in object oriented design. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):28–32, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Truong:2005:ELP

[Tru05] Nghi Truong. The environment for learning to program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):383, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Thomas:2002:LSP

[TRWJ02] Lynda Thomas, Mark Ratcliffe, John Woodbury, and Emma Jarman. Learning styles and performance in the introductory programming sequence. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):33–37, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Trytten:2005:DTP

[Try05] Deborah A. Trytten. A design for team peer code review. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):455–459, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tymann:2000:MSD

[TS00] Paul T. Tymann and G. Michael Schneider. Modern software development concepts: a new philosophy for CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(4):63–65, December 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Treu:2002:TSG

[TS02] Kevin Treu and Alisha Skinner. Ten suggestions for a genderequitable CS classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):165– 167, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/

2002.bib. Women and Computing.

Taylor:2008:SER

[TS08] Carol Taylor and Rose Shumba. Security education: a roadmap to the future. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):459–460, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Turner:2002:UCS

[TSC^{+}02] Peter R. Turner, Angela B. Shiflet, Steve Cunningham, Kris Stewart, Andrew T. Phillips, and Ignatios E. Vakalis. Undergraduate computational science and engineering programs and courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):96–97, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Tegos:2005:LTM

[TSO05] Georgios K. Tegos, Diana V. Stoyanova, and Kolyo Z. Onkov. E-learning of trend modeling in a Web-environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):70–74, June 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2005.bib.

Thomas:2007:CAA

[TSW07] Pete G. Thomas, Neil Smith, and Kevin G. Waugh. Computer assisted assessment of diagrams. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):68–72, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Tjaden:2000:WWB

[TT00] Bunny J. Tjaden and Brett C. Tjaden. A worldwide, Webbased study of the attitudes of college freshmen toward computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):29–32, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Turner:2005:TES

[TT05] Elise H. Turner and Roy M. Turner. Teaching entering students to think like computer scientists. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):307–311, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tjaden:2006:TSA

[TT06] Brian Tjaden and Brett Tjaden. Training students to administer and defend computer networks and systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):245–249, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Turner:2001:IER

[Tur01] Joe Turner. Invited editorial: reflections on curriculum development in computing programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33 (2):4–6, June 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2001.bib.

Turau:2003:HEH

[Tur03] Volker Turau. HTTPExplorer: exploring the hypertext transfer protocol. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):198–201, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Turner:2006:WI

[Tur06] A. Joe Turner. What is IFIP? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4): 17–18, December 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2006.bib.

Turner:2007:IEA

[Tur07a] A. Joe Turner. IFIP educational activities. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):18–19, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2007.bib.

Turner:2007:MIE

[Tur07b] A. Joe Turner. More on IFIP educational activities. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):22, December 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Turner:2008:IVW

[Tur08a] A. Joe Turner. IFIP vibes: WCCE 2009 in Brazil. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):23, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Turner:2008:SIH

[Tur08b] A. Joe Turner. Some IFIP happenings. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):20, June 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2008.bib.

Turner:2009:IVS

[Tur09] A. Joe Turner. IFIP vibes: the Seoul accord. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):78–79, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tuttle:2000:CCC

[Tut00] Sharon M. Tuttle. A capstone course for a computer information systems major. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):265–269, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tomic:2008:FTS

[TV08] Bojan Tomić and Sinisa Vlajić. Functional testing for students: a practical approach. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4): 58–62, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tobin:2005:UWA

[TW05] Donald L. Tobin, Jr. and Michael S. Ware. Using a windows attack intRusion emulator (AWARE) to teach computer security awareness. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):213–217, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Talebi:2009:MMM

[TW09a] Mujtaba Talebi and Thomas Way. Methods, metrics and motivation for a green computer science program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):362–366, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Tymann:2009:FAC

[TW09b] Paul T. Tymann and Laurie White. The future of the AP CS program. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):331–332, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Thomas:2005:EAM

[TWS05] Pete Thomas, Kevin Waugh, and Neil Smith. Experiments in the automatic marking of ER-diagrams. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):158–162, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Thomas:2006:UPA

[TWS06] Pete Thomas, Kevin Waugh, and Neil Smith. Using patterns in the automatic marking of ER-diagrams. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):83–87, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Thomas:2009:GDR

[TWS09] Pete Thomas, Kevin Waugh, and Neil Smith. Generalised diagram revision tools with automatic marking. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):318–322, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Tran:2006:ETT

[TY06] Sam Phu Manh Tran and T. Andrew Yang. Evaluations of target tracking in wireless sensor networks. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):97–101, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tungare:2007:TSR

[TYC^{+}07] Manas Tungare, Xiaoyan Yu, William Cameron, GuoFang Teng, Manuel A. Pérez-Quiñones, Lillian Cassel, Weiguo Fan, and Edward A. Fox. Towards a syllabus repository for computer science courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):55–59, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Tahara:2009:ECM

[TYT^{+}09] Yasuyuki Tahara, Nobukazu Yoshioka, Kenji Taguchi, Toshiaki Aoki, and Shinichi Honiden. Evolution of a course on model checking for practical applications. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):38–44, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Turner:2001:JTV

[TZ01] Joseph A. Turner and Joseph L. Zachary. Javiva: a tool for visualizing and validating student-written Java programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):45–49, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Umphress:2004:BJI

[UCJ^{+}04] David A. Umphress, James H. Cross II, Jhilmil Jain, Nischita Meda, and Larry A. Barowski. Bringing J2ME industry practice into the classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):301– 305, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Urban:2001:ADC

[UD01] Susan D. Urban and Suzanne W. Dietrich. Advanced database concepts for undergraduates: experience with teaching a second course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):357–361, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Urban:2003:UUC

[UD03] Susan D. Urban and Suzanne W. Dietrich. Using UML class diagrams for a comparative analysis of relational, objectoriented, and object-relational database mappings. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):21–25, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Udoh:2006:TDI

[Udo06] Emmanuel Udoh. Teaching database in an integrated Oracle environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):71–74, December 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2006.bib.

Urquiza-Fuentes:2006:VST

[UFGCGBVI06] Jaime Urquiza-Fuentes, Micael Gallego-Carrillo, Francisco Gortázar-Bellas, and J. Ángel Velázquez-Iturbide. Visualizing the symbol table. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):341, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Urquiza-Fuentes:2005:RCW

[UFVI05] Jaime Urquiza-Fuentes and J. Ángel Velázquez-Iturbide. Reusable collections of Web-based program animations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):386, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Urban:2009:PRU

[UHFT09] Joseph E. Urban, Jesse M. Heines, Edward A. Fox, and Harriet G. Taylor. Panel on revitalized undergraduate computing education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):69–70, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Utting:2006:PIT

[Utt06] Ian Utting. Problems in the initial teaching of programming using Java: the case for replacing J2SE with J2ME. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):193–196, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

VanDeGrift:2002:LSI

[VA02] Tammy VanDeGrift and Richard J. Anderson. Learning to support the instructor: classroom assessment tools as discussion frameworks in CS 1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):19–23, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

VanRoy:2003:RLP

[VAFM03] Peter Van Roy, Joe Armstrong, Matthew Flatt, and Boris Magnusson. The role of language paradigms in teaching programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):269–270, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Valentine:2004:CER

[Val04] David W. Valentine. CS educational research: a metaanalysis of SIGCSE technical symposium proceedings. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):255–259, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

VanDeGrift:2004:CPP

[Van04] Tammy VanDeGrift. Coupling pair programming and writing: learning about students’ perceptions and processes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):2–6, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Vandever:2008:TBS

[Van08] Kelly Vandever. Teaching the business of software development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):90–92, June 2008. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2008.bib.

Vasiga:2002:WCA

[Vas02] Troy Vasiga. What comes after CS 1 + 2: a deep breadth before specializing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):28–32, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Vat:2000:TSP

[Vat00] Kam Hou Vat. Teaching software psychology: expanding the perspective. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):392–396, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Vat:2001:THS

[Vat01] Kam Hou Vat. Teaching HCI with scenario-based design: the constructivist’s synthesis. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):9–12, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Vaughn:2000:AST

[Vau00] Rayford B. Vaughn, Jr. Application of security to the computing science classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):90– 94, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Vidal:2002:URT

[VB02] José M. Vidal and Paul Buhler. Using RoboCup to teach multiagent systems and the distributed mindset. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):3–7, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Verno:2005:DRS

[VCC^{+}05] Anita Verno, Debbie Carter, Robb Cutler, Michelle Hutton, and Lenny Pitt. Developing resources to support a national computer science curriculum for K-12. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):388–389, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Verno:2007:DRS

[VCC^{+}07] Anita Verno, Steve Cooper, Thomas J. Cortina, Barb Ericson, and Bill Madden. Developing resources to support a national computer science curriculum for K-12. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):377–378, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Villalobos:2009:DPS

[VCJ09] Jorge A. Villalobos, Nadya A. Calderon, and Camilo H. Jiménez. Developing programming skills by using interactive learning objects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):151–155, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

vanderKuyl:2007:WAC

[vdK07] Chris van der Kuyl. Where have all the computer scientists gone? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):2, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Verma:2005:VIA

[Ver05] Rakesh M. Verma. A visual and interactive automata theory course emphasizing breadth of automata. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):325–329, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Voyles:2007:TRS

[VHF07] M. M. Voyles, S. M. Haller, and T. V. Fossum. Teacher responses to student gender differences. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):226–230, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Velazquez-Iturbide:2000:RGS

[VI00] J. Ángel Velázquez-Iturbide. Recursion in gradual steps (is recursion really that difficult?). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1): 310–314, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Velazquez-Iturbide:2001:FPA

[VI01] J. Á. Velázquez-Iturbide. A first problem for the algorithms course (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):187, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Velazquez-Iturbide:2005:PLC

[VI05] J. Ángel Velázquez-Iturbide. A programming languages course for freshmen. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):271–275, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Vinha:2005:RLO

[Vin05] Antonio Vinha. Reusable learning objects: theory to practice. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):413, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Velazquez-Iturbide:2009:ALG

[VIPC09] J. Ángel Velázquez-Iturbide and Antonio Pérez-Carrasco. Active learning of greedy algorithms by means of interactive experimentation. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):119–123, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Velazquez-Iturbide:2008:SAS

[VIPCUF08] J. Ángel Velázquez-Iturbide, Antonio Pérez-Carrasco, and Jaime Urquiza-Fuentes. SRec: an animation system of recursion for algorithm courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3): 225–229, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Velazquez-Iturbide:2009:IVR

[VIPCUF09] J. Ángel Velázquez-Iturbide, Antonio Pérez-Carrasco, and Jaime Urquiza-Fuentes. Interactive visualization of recursion with SRec. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):339, September 2009.

CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Voracek:2001:IMP

[VKV01] Jan Voracek and Nina Kontro-Vesivalo. International Master’s program in information processing and telecommunications (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):181, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Voracek:2002:IEI

[VKV02] Jan Voracek and Nina Kontro-Vesivalo. International education in information technology. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (3):245, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

vanLangeveld:2009:TMD

[vLK09] Mark Christensen van Langeveld and Robert Kessler. Two in the middle: digital character production and machinima courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):463–467, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Veal:2001:AHS

[VMD01] D. Veal, S. P. Maj, and Rick Duley. Assessing “hands on” skills on CS1 computer & network technology units. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):381–385, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ventura:2004:WCS

[VR04] Phil Ventura and Bina Ramamurthy. Wanted: CS1 students. no experience required. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):240– 244, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Vollmar:2006:MEO

[VS06] Kenneth Vollmar and Pete Sanderson. MARS: an educationoriented MIPS assembly language simulator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):239–243, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Vainio:2007:FNP

[VS07a] Vesa Vainio and Jorma Sajaniemi. Factors in novice programmers’ poor tracing skills. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3): 236–240, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

VanBuren:2007:MGM

[VS07b] Brian G. VanBuren and Muhammad Shaaban. MicroTiger: a graphical microcode simulator with a reconfigurable datapath. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):283–287, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Van:2001:FAQ

[VT01] Huu Le Van and Andrea Trentini. A “frequently asked questions” management system that supports voting, built for student evaluation and optimization purposes (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):184, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

vanTonder:2008:JLD

[vTNC08] Martin van Tonder, Kevin Naude, and Charmain Cilliers. Jenuity: a lightweight development environment for intermediate level programming courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40 (3):58–62, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

vanVeen:2004:WLC

[vVML04] Maarten van Veen, Fred Mulder, and Karel Lemmen. What is lacking in curriculum schemes for computing/informatics? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):186–190, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Vieira:2004:LEH

[VVV04] Luiz Filipe M. Vieira, Marcos Augusto M. Vieira, and Newton J. Vieira. Language emulator, a helpful toolkit in the learning process of computer theory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (1):135–139, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Vandenberg:2000:ICS

[VW00a] Scott Vandenberg and Michael Wollowski. Introducing computer science using a breadth-first approach and functional programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):180–184, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

vonWright:2000:DTS

[vW00b] Joakim von Wright. Distance tutorials in a systems design course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):105–107, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Vilner:2006:OSM

[VZ06] Tamar Vilner and Ela Zur. Once she makes it, she is there: gender differences in computer science study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):227–231, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Vilner:2007:FCC

[VZGE07] Tamar Vilner, Ela Zur, and Judith Gal-Ezer. Fundamental concepts of CS1: procedural vs. object oriented paradigm — a case study. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):171–175, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Wilson:2008:CS

[WAC^{+}08] Greg Wilson, Christine Alvarado, Jennifer Campbell, Rubin Landau, and Robert Sedgewick. CS-1 for scientists. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):36–37, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Wagner:2005:TDM

[Wag05] Paul Wagner. Teaching data modeling: process and patterns. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):168–172, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Waite:2006:CCT

[Wai06] William M. Waite. The compiler course in today’s curriculum: three strategies. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):87–91, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:2000:BFT

[Wal00a] Henry M. Walker. Balancing the forest and the trees in courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(4):17–18, December 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:2000:NG

[Wal00b] Henry M. Walker. Notes on grading. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (2):18–19, June 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:2001:NAJ

[Wal01a] Henry Walker. Notes on the academic job market and hiring strategies. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(4):32–34, December 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2001.bib.

Walker:2001:STR

[Wal01b] Henry Walker. SIGCSE Treasurer’s report for the June 2001 SIGCSE Bulletin. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(2):60–62, June 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2001.bib.

Walker:2001:TSD

[Wal01c] Henry MacKay Walker. Teaching and a sense of the dramatic. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(4):16–17, December 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2001.bib.

Walker:2002:SFI

[Wal02a] Henry Walker. SIGCSE 2002 forum: introducing SIGCSE committees: a new initiative to support computing education through SIGCSE-member involvement. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):368, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Walker:2002:TSD

[Wal02b] Henry MacKay Walker. Teaching and a sense of the dramatic: act II. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):18–19, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2002.bib.

Walker:2003:DCG

[Wal03a] Henry MacKay Walker. Do computer games have a role in the computing classroom? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):18–20, December 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2003.bib.

Walker:2003:LC

[Wal03b] Henry MacKay Walker. Lessons from the CUPM. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):19–21, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.

math.utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2003.bib.

Walker:2004:ECP

[Wal04a] Gary N. Walker. Experimentation in the computer programming lab. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(4):69–72, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2004.bib.

Walker:2004:AHC

[Wal04b] Henry MacKay Walker. Academic honesty in the classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(4):18–19, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2004.bib.

Walker:2004:WTS

[Wal04c] Henry MacKay Walker. What teachers should, can, and cannot do. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(2):20–21, June 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2004.bib.

Walden:2005:RTI

[Wal05a] James Walden. A real-time information warfare exercise on a virtual network. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):86–90, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:2005:MCT

[Wal05b] Henry MacKay Walker. Mathematics and CS topics in the CS classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):15–17, June 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2005.bib.

Walker:2005:WSS

[Wal05c] Henry MacKay Walker. What should be in a syllabus? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4):19–21, December 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2005.bib.

Walls:2005:IVE

[Wal05d] David C. Walls. Integrating views on ethical behavior for computer professionals in an interdisciplinary environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):414, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:2006:TAL

[Wal06a] Henry MacKay Walker. Thoughts about lecturing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):13–15, June 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.

math.utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2006.bib.

Walker:2006:TSF

[Wal06b] Henry MacKay Walker. Thoughts on student feedback to help teaching. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(4):13–14, December 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2006.bib.

Walker:2007:RCW

[Wal07a] Henry MacKay Walker. Reading and class work. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):13–14, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.

math.utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2007.bib.

Walker:2007:WID

[Wal07b] Henry MacKay Walker. What image do CS1/CS2 present to our students? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):18–19, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2007.bib.

Walker:2008:CIS

[Wal08a] Henry M. Walker. Classroom issues: staying connected with the big picture. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):18–20, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:2008:AR

[Wal08b]  Henry MacKay Walker.  Advertising and recruiting.

SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):16–17, June 2008. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2008.bib.

Walker:2009:CIC

[Wal09a] Henry M. Walker. Classroom issues: course descriptions and public relations for computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(2):74–75, June 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:2009:CIG

[Wal09b] Henry M. Walker. Classroom issues: Grading and the allocation of points. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4):14–16, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wang:2001:IOR

[Wan01] Ming Wang. Implementation of object-relational DBMSs in a relational database course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):367– 370, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wang:2006:ITC

[Wan06] Yang Wang. It is time for a computer application degree. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(2):105–109, June 2006. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2006.bib.

Warms:2001:TEC

[War01] Tom M. Warms. Tracing the execution of C++ programs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(4):64–67, December 2001. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2001.bib.

Waraich:2004:UNM

[War04] Atif Waraich. Using narrative as a motivating device to teach binary arithmetic and logic gates. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (3):97–101, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Warms:2005:PNM

[War05] Tom M. Warms. The power of notation: modeling pointer operations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):41–45, June 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2005.bib.

Walters:2002:ILD

[WAS^{+}02] Deborah Walters, Carl Alphonce, Barbara Sherman, Debra Burhans, and Helene Kershner. Increasing learning and decreasing costs in a computer fluency course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):208–212, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Way:2005:CBF

[Way05] Thomas P. Way. A company-based framework for a software engineering course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):132–136, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Way:2006:VLM

[Way06] Thomas P. Way. A virtual laboratory model for encouraging undergraduate research. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):203– 207, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wardle:2002:PEE

[WB02] Caroline E. Wardle and Lawrence Burton. Programmatic efforts encouraging women to enter the information technology workforce. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):27–31, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/

2002.bib. Women and Computing.

Woit:2005:SCC

[WB05] D. Woit and K. Bell. Student communication challenges in distributed software engineering environments. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):286–290, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wirth:2006:NAR

[WB06] Anthony Wirth and Michael Bertolacci. New algorithms research for first year students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):128– 132, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wiesner:2009:HDR

[WB09] Bernhard Wiesner and Torsten Brinda. How do robots foster the learning of basic concepts in informatics? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):403, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Wolz:2009:GDL

[WBBC09] Ursula Wolz, Tiffany Barnes, Jessica Bayliss, and Jamie Cromack. Girls do like playing and creating games. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):199–200, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Wolf:2002:OSS

[WBGM02] Marty J. Wolf, Kevin Bowyer, Don Gotterbarn, and Keith Miller. Open source software: intellectual challenges to the status quo. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):317–318, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Wolz:2006:DGV

[WBPW06] Ursula Wolz, Tiffany Barnes, Ian Parberry, and Michael Wick. Digital gaming as a vehicle for learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):394–395, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wong:2004:DMC

[WBS04] Yue-Ling Wong, Jennifer Burg, and Victoria Strokanova. Digital media in computer science curricula. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):427–431, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wang:2004:EBA

[WC04] Chin-Yeh Wang and Gwo-Dong Chen. Extending e-books with annotation, online support and assessment mechanisms to increase efficiency of learning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36 (3):132–136, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wong:2003:PCT

[WD03] Jacqueline Wong and Timon Du. Project-centered teaching on CBIS to IBBA students in Hong Kong. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):35–38, December 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2003.bib.

Wermelinger:2005:PTF

[WD05] Michel Wermelinger and Artur Miguel Dias. A Prolog toolkit for formal languages and automata. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37 (3):330–334, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Weaver:2004:ECS

[Wea04] Alfred C. Weaver. Electronic commerce software laboratory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):453–456, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Weiss:2005:TDP

[Wei05a] Stephen Weiss. Teaching design patterns by stealth. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):492–494, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Weizenbaum:2005:IIS

[Wei05b] Joseph Weizenbaum. Information in the information society. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):1, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wenderholm:2004:CES

[Wen04] Elaine Wenderholm. Challenges and the elements of success in undergraduate research. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(4): 73–75, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.

edu/pub/mirrors/ftp.ira.uka.de/bibliography/Misc/
DBLP/2004.bib.

Wilkinson:2008:TTA

[WF08] Barry Wilkinson and Clayton Ferner. Towards a topdown approach to teaching an undergraduate grid computing course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):126–130, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Whaley:2007:DSK

[WG07] H. Whaley and S. Grice. Do students know best?: experiences of allowing students to become course designers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):330, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Wilkerson:2005:UPI

[WGS05] Michelle Wilkerson, William G. Griswold, and Beth Simon. Ubiquitous presenter: increasing student access and control in a digital lecturing environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37 (1):116–120, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Hwang:2007:TOS

[wH07] Seung won Hwang. Teaching operating systems with Windows: experiences and contributions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):316, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Hwang:2009:BLT

[wH09a] Seung won Hwang. Blended learning for teaching operating systems with windows. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):380, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Hwang:2009:VCT

[wH09b] Seung won Hwang. Video courseware for teaching operating systems with windows. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):340, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Wang:2009:RCE

[WHCT09] Shu-Ling Wang, Gwo-Haur Hwang, Ju-Chun Chu, and Pei-Shan Tsai. The role of collective efficacy and collaborative learning behavior in learning computer science through CSCL. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):352, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

White:2003:CCI

[Whi03] Kent White. A comprehensive CMPS II semester project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):70–73, June 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

White:2009:HLS

[Whi09] Curt M. White. An historical look at the SIGCSE conference. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):131, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

White:2009:RRT

[WI09] Su White and Alastair Irons. Relating research and teaching: learning from experiences and beliefs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):75–79, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Wick:2001:KUD

[Wic01] Michael R. Wick. Kaleidoscope: using design patterns in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):258–262, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wick:2003:OOR

[Wic03] Michael R. Wick. An object-oriented refactoring of Huffman encoding using the Java collections framework. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):283–287, January 2003. CO-DEN SIGSD3. ISBN 1-58113-648-X. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Compiler/java.bib;   ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/2003.
bib.

Wick:2005:TDP

[Wic05] Michael R. Wick. Teaching design patterns in CS1: a closed laboratory sequence based on the game of life. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):487–491, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wick:2007:BCG

[Wic07] Michael R. Wick. Bridging the conceptual gap: assessing the impact on student attitudes toward programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):509–513, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wick:2009:UPH

[Wic09] Michael R. Wick. Using programming to help students understand the value of diversity. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):367– 371, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Williams:2000:DWT

[Wil00] Michael R. Williams. Do we teach computer science as religion or as history? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(4):4–5, December 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Williams:2003:CHM

[Wil03] Michael R. Williams. The computer history museum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(4):12–13, December 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

Williams:2004:DAR

[Wil04a] Michael R. Williams. Does anyone remember the KDF-9? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(2):13–14, June 2004. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2004.bib.

Williams:2004:SC

[Wil04b] Michael R. Williams. Survey calculations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(4):12–13, December 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2004.bib.

Williams:2005:RS

[Wil05] Michael R. Williams. Reflections on standards. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):11–12, June 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.

math.utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2005.bib.

Williams:2007:LLSb

[Wil07] Laurie Williams. Lessons learned from seven years of pair programming at North Carolina State University. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):79–83, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Misc/DBLP/2007.bib.

Wirth:2002:CSE

[Wir02] Niklaus Wirth. Computing science education: the road not taken. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):1–3, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wirth:2003:NUE

[Wir03] Michael A. Wirth. E-notes: using electronic lecture notes to support active learning in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2):57–60, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.

utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2003.bib.

Wirth:2008:IRP

[Wir08] Michael Wirth. Introducing recursion by parking cars. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):52–55, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wiseman:2005:AND

[Wis05] Yair Wiseman. Advanced non-distributed operating systems course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):65–69, June 2005. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2005.bib.

Wise:2006:GJD

[Wis06] Jon Wise. GoJava: a Java development tool for beginners. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):359, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Waite:2003:CC

[WJD03] William M. Waite, Michele H. Jackson, and Amer Diwan. The conversational classroom. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1): 127–131, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Waite:2004:SCV

[WJDL04] William M. Waite, Michele H. Jackson, Amer Diwan, and Paul M. Leonardi. Student culture vs group work in computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):12–16, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Waite:2006:DIM

[WJJD06] William M. Waite, Assad Jarrahian, Michele H. Jackson, and Amer Diwan. Design and implementation of a modern compiler course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):18–22, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Waegel:2006:TTM

[WK06] Daniel B. Waegel and April Kontostathis. TextMOLE: text mining operations library and environment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):553–557, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Woodley:2007:PSC

[WK07] Michael Woodley and Samuel N. Kamin. Programming studio: a course for improving programming skills in undergraduates. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):531–535, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wyatt:2002:ISI

[WKB02] Jason N. Wyatt, Martha J. Kosa, and Mark A. Boshart. Implementing student ideas in CS2: a simple IDE. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):224, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wein:2009:VGT

[WKC^{+}09] Joel Wein, Kirill Kourtchikov, Yan Cheng, Ron Gutierez, Roman Khmelichek, Matthew Topol, and Chris Sherman. Virtualized games for teaching about distributed systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):246–250, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Wilkins:2000:EIT

[WL00] Dawn E. Wilkins and Pamela B. Lawhead. Evaluating individuals in team projects. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):172– 175, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wang:2008:PIP

[WLCL08] Yanqing Wang, Yijun Li, Michael Collins, and Peijie Liu. Process improvement of peer code review and behavior analysis of its participants. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):107–111, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Wolz:2009:SSC

[WLMM09] Ursula Wolz, Henry H. Leitner, David J. Malan, and John Maloney. Starting with scratch in CS 1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):2–3, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Woit:2000:ESL

[WM00] Denise Woit and Dave Mason. Enhancing student learning through on-line quizzes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(1):367– 371, March 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wagner:2003:IXD

[WM03a] Paul J. Wagner and Thomas K. Moore. Integrating XML into a database systems course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):26– 30, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Woit:2003:EOA

[WM03b] Denise Woit and David Mason. Effectiveness of online assessment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):137–141, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:2001:PST

[WMGR01] Henry MacKay Walker, Renée A. McCauley, Judith L. Gersting, and Ingrid Russell, editors. Proceedings of the 32rd SIGCSE Technical Symposium on Computer Science Education, 2001, Charlotte, North Carolina, USA, 2001, SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education). ACM Press, New York, NY 10036, USA, 2001. CODEN SIGSD3. ISBN 1-58113-329-4. ISSN 0097-8418 (print), 2331-3927 (electronic).

Walker:2002:VRR

[WMM02] Henry M. Walker, Weichao Ma, and Dorene Mboya. Variability of referees’ ratings of conference papers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):178–182, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wolz:2008:SYW

[WMP08] Ursula Wolz, John Maloney, and Sarah Monisha Pulimood. ’Scratch’ your way to introductory CS. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):298–299, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Wicentowski:2005:UIP

[WN05] Richard Wicentowski and Tia Newhall. Using image processing projects to teach CS1 topics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37 (1):287–291, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Warford:2007:PPS

[WO07] J. Stanley Warford and Ryan Okelberry. Pep8CPU: a programmable simulator for a central processing unit. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):288–292, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wolz:2001:TDP

[Wol01] Ursula Wolz. Teaching design and project management with Lego RCX robots. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):95–99, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wolfman:2002:MLE

[Wol02] Steven A. Wolfman. Making lemonade: exploring the bright side of large lecture classes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):257– 261, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Wolf:2007:AII

[Wol07] Tilman Wolf. Assessing the impact of inking technology in a large digital design course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):79– 83, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Woszczynski:2006:COI

[Wos06] Amy B. Woszczynski. CyberTech I: online introduction to computer science course for high school students. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):153–157, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wick:2002:CTM

[WP02] Michael R. Wick and Andrew T. Phillips. Comparing the template method and strategy design patterns in a genetic algorithm application. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(4):76–80, December 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2002.bib.

Winters:2006:CLT

[WP06a] Titus Winters and Tom Payne. Closing the loop on test creation: a question assessment mechanism for instructors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):169–170, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Winters:2006:CAA

[WP06b] Titus Winters and Tom Payne. Computer aided assessment with human oversight. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):320, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wolz:2007:IAP

[WP07] Ursula Wolz and Sarah Monisha Pulimood. An integrated approach to project management through classic CS III and video game development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):322– 326, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wilson:2001:LDI

[WPS01] Laura Wilson, Jon Preston, and Russell Shackelford. Online dynamic interviews (ODIN) (poster session): a means of overcoming distance in student–teacher relations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3):178, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wortman:2007:VTS

[WR07] Dana Wortman and Penny Rheingans. Visualizing trends in student performance across computer science courses. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):430–434, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wick:2001:RAC

[WS01a] Michael R. Wick and Daniel E. Stevenson. A reductionist approach to a course on programming languages. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):253–257, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wilson:2001:CSI

[WS01b] Brenda Cantwell Wilson and Sharon Shrock. Contributing to success in an introductory computer science course: a study of twelve factors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):184–188, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Weinberg:2002:PDH

[WS02] Jerry B. Weinberg and Mary L. Stephen. Participatory design in a human-computer interaction course: teaching ethnography methods to computer scientists. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (1):237–241, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Wick:2006:USI

[WS06] Michael R. Wick and Daniel E. Stevenson. On using scheme to introduce Prolog. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):41–45, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wagner:2003:USD

[WSC03] Paul J. Wagner, Elizabeth Shoop, and John V. Carlis. Using scientific data to teach a database systems course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):224–228, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wick:2002:UEC

[WSP02] Michael R. Wick, Daniel E. Stevenson, and Andrew T. Phillips. Using an environment chain model to teach inheritance in C++. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):297–301, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Wick:2004:SDR

[WSP04] Michael R. Wick, Daniel E. Stevenson, and Andrew T. Phillips. Seven design rules for teaching students sound encapsulation and abstraction of object properties and member data. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):100–104, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

White:2005:HSU

[WSS05] Elizabeth White, Ranjan Sen, and Nina Stewart. Hide and show: using real compiler code for teaching. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):12–16, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wick:2005:UTJ

[WSW05] Michael Wick, Daniel Stevenson, and Paul Wagner. Using testing and JUnit across the curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):236–240, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Williams:2008:IEC

[WTTT^{+}08] Andrew B. Williams, David S. Touretzky, Ethan J. Tira-Thompson, LaVonne Manning, Chutima Boonthum, and Clement S. Allen. Introducing an experimental cognitive robotics curriculum at historically black colleges and universities. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):498–502, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Williams:2001:SSP

[WU01] Laurie Williams and Richard L. Upchurch. In support of student pair-programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):327– 331, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wu:2005:TGA

[Wu05] Mingshen Wu. Teaching graph algorithms using online Java package IAPPGA. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4):64–68, December 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2005.bib.

Weir:2005:DTJ

[WVMN05] George R. S. Weir, Tamar Vilner, António José Mendes, and Marie Nordström. Difficulties teaching Java in CS1 and how we aim to solve them. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):344–345, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Weber-Wulff:2000:CCW

[WW00] Debora Weber-Wulff. Combating the code warrior: a different sort of programming instruction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (3):85–88, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wagner:2004:DIC

[WW04] Paul J. Wagner and Jason M. Wudi. Designing and implementing a cyberwar laboratory exercise for a computer security course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):402–406, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

White:2006:JJF

[WW06a] Timothy M. White and Thomas P. Way. jFAST: a Java finite automata simulator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):384– 388, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wick:2006:UMB

[WW06b] Michael R. Wick and Paul J. Wagner. Using market basket analysis to integrate and motivate topics in discrete structures. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):323–327, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

White:2009:BLT

[WW09] William W. White and Jerry B. Weinberg. Breadth-last technical electives: integrating the CS core via computer games and mobile robotics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):54–58, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Wong:2007:TEV

[WWGT07] Ken Wong, Tilman Wolf, Sergey Gorinsky, and Jonathan Turner. Teaching experiences with a virtual network laboratory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):481–485, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Weinberg:2005:MTR

[WWK^{+}05] Jerry B. Weinberg, William W. White, Cem Karacal, George Engel, and Ai-Ping Hu. Multidisciplinary teamwork in a robotics course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):446–450, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Wiseman:2008:OEV

[WWWG08] Charlie Wiseman, Ken Wong, Tilman Wolf, and Sergey Gorinsky. Operational experience with a virtual networking laboratory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):427–431, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Wolffe:2002:TCO

[WYOH02] Gregory S. Wolffe, William Yurcik, Hugh Osborne, and Mark A. Holliday. Teaching computer organization/ architecture with limited resources using simulators. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):176–180, March 2002. CODEN SIGSD3.

ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Wu:2000:EPN

[WZ00] Congxin Wu and Bokan Zhang. Embedding problem of noncompact fuzzy number space E (II). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32 (3):135–142, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Xinogalos:2003:ODM

[Xin03] S. Xinogalos. objectKarel: a didactic microworld for teaching object-oriented programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35 (3):233, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Xizhe:2001:ETS

[Xiz01] Jin Xizhe. Evaluation technique of software configuration management (poster session). SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(3): 186, September 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Xu:2006:CCT

[XM06] Li Xu and Fred G. Martin. Chirp on crickets: teaching compilers using an embedded robot controller. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):82–86, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Xinogalos:2007:TJB

[XSD07] Stelios Xinogalos, Maya Satratzemi, and Vassilios Dagdilelis. Teaching Java with BlueJ: a two-year experience. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):345, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Xu:2007:RMI

[Xu07] Li Xu. RobotStudio: a modern IDE-based approach to reality computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):440–444, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Xu:2008:LEC

[Xu08] Li Xu. Language engineering in the context of a popular, inexpensive robot platform. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(1):43– 47, March 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE 08.

Yusof:2005:EPC

[YA05] Azwina M. Yusof and Rukaini Abdullah. The evolution of programming courses: course curriculum, students, and their performance. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(4):74–78, December 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2005.bib.

Yair:2008:SFO

[Yai08] Yoav Yair. A step further: opening books and educational resources. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):22–23, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Yang:2007:EOA

[Yan07] Feng-Jen Yang. Eliciting an overlooked aspect of Bayesian reasoning. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):45–48, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2007.bib.

Yang:2008:AOL

[Yan08] Feng-Jen Yang. Another outlook on linear recursion. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):38–41, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Yang:2009:SMA

[Yan09a] Feng-Jen Yang. Stopping a myth in artificial neural networks. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(4):25–29, December 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Yang:2009:TDS

[Yan09b] Li Yang. Teaching database security and auditing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):241–245, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Yurcik:2001:WBL

[YB01] William Yurcik and Larry Brumbaugh. A Web-based Little Man Computer simulator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):204– 208, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Yehezkel:2005:CAM

[YBAD05] Cecile Yehezkel, Mordechai Ben-Ari, and Tommy Dreyfus. Computer architecture and mental models. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):101–105, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Yue:2004:DEU

[YD04] Kwok-Bun Yue and Wei Ding. Design and evolution of an undergraduate course on Web application development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):22–26, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Yehezkel:2002:TCA

[Yeh02] Cecile Yehezkel. A taxonomy of computer architecture visualizations. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):101–105, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Yehezkel:2003:MPE

[Yeh03] Cecile Yehezkel. Making program execution comprehensible one level above the machine language. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):124–128, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Yong:2003:ACP

[YH03] Chye-Foong Yong and Colin Higgins. Automatically creating personalised exercises based on student profiles. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):236, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Yoo:2004:RAI

[YH04] Sung Yoo and Scott Hovis. Remote access internetworking laboratory. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):311–314, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Yehezkel:2005:CSA

[YH05] Cecile Yehezkel and Bruria Haberman. “computer science, academia, and industry” educational project. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):364, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Yarmish:2007:RNP

[YK07] Gavriel Yarmish and Danny Kopec. Revisiting novice programmer errors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(2):131–137, June 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2007.bib.

Yaverbaum:2007:OBC

[YLRZ07] Gayle Yaverbaum, Doris Lidtke, Han Reichgelt, and Stu Zweben. Outcomes-based computing accreditation: program assessment. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):281–282, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Yu:2005:PMM

[YLS05] Tin-Tin Yu, John L. Lowther, and Ching-Kuang Shene. Photon mapping made easy. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37 (1):201–205, March 2005. CODEN SIGSD3. ISBN 1-58113-997-7. ISSN 0097-8418 (print), 2331-3927 (electronic).

URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.ira.
uka.de/bibliography/Graphics/rad.bib;   ftp://ftp.
math.utah.edu/pub/mirrors/ftp.ira.uka.de/bibliography/
Misc/DBLP/2005.bib.

Yu:2006:TWS

[YLYX06] H. Yu, W. Liao, X. Yuan, and J. Xu. Teaching a Web security course to practice information assurance. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):12–16, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Young:2002:ISE

[YM02] Alison Young and Samuel Mann. Innovation in software engineering education. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(3):219, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Young:2006:AFS

[YM06] Alison Young and Logan Muller. Aluminum foil satellite dishes and a millennium of experience: sustainability in the high Andes. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):2, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Yadin:2008:FIL

[YOB08] Aharon Yadin and Rachel Or-Bach. Fostering individual learning: when and how. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):83–86, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Young:2008:SSO

[YPdMA^{+}08] Alison Young, Arnold Pears, Pedro de Miguel Anasagasti, Ralf Romeike, Michael Goldweber, Michael Goldwasser, and Vicki Almstrum. Scrambling for students: our graduates are sexier than yours. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):333–334, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Yoo:2000:PSD

[YSLL00] In-Hwan Yoo, Soo-Bum Shin, Chul-Hyun Lee, and Tae-Wuk Lee. Present status and direction of information curriculum of Korea. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):195, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Yuen:2007:NKC

[Yue07] Timothy T. Yuen. Novices’ knowledge construction of difficult concepts in CS1. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(4):49–53, December 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2007.bib.

Yoo:2006:SPM

[YYLH06] Jungsoon Yoo, Sung Yoo, Chris Lance, and Judy Hankins. Student progress monitoring tool using treeview. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):373–377, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Zhang:2008:GIS

[ZB08] Wendy Zhang and Theresa Beaubouef. Geographic information systems: real world applications for computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):124–127, June 2008. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/
ftp.ira.uka.de/bibliography/Misc/DBLP/2008.bib.

Zimmerman:2001:WWC

[ZE01] Guy W. Zimmerman and Dena E. Eber. When worlds collide!: an interdisciplinary course in virtual-reality art. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 33(1):75–79, March 2001. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Zeller:2000:MSR

[Zel00] Andreas Zeller. Making students read and review code. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(3):89–92, September 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Zelle:2004:SLC

[ZF04] John M. Zelle and Charles Figura. Simple, low-cost stereographics: VR for everyone. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):348– 352, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Zupancic:2002:LRU

[ZH02] Bernd Zupancic and Holger Horz. Lecture recording and its use in a traditional university course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34 (3):24–28, September 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Zur:2005:CWE

[ZIBG05] Ela Zur, Lilly Irani, Lecia Barker, and Mark Guzdial. Contrasting women’s experiences in computer science at different institutions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):63–64, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Zilles:2005:SEP

[Zil05] Craig Zilles. SPIMbot: an engaging, problem-based approach to teaching assembly language programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):106–110, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Zingaro:2008:AAR

[Zin08] Daniel Zingaro. Another approach for resisting student resistance to formal methods. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):56–57, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Zachary:2003:EVA

[ZJ03] Joseph L. Zachary and Peter A. Jensen. Exploiting valueadded content in an online course: introducing programming concepts via HTML and JavaScript. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35 (1):396–400, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Zhang:2007:IAT

[ZLL^{+}07] Mingrui Zhang, Eugene Lundak, Chi-Cheng Lin, Tim Gegg-Harrison, and Joan Francioni. Interdisciplinary application tracks in an undergraduate computer science curriculum. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(1):425–429, March 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Zhang:2007:BTO

[ZLOB07] Mingrui Zhang, Chi-Cheng Lin, Gayle Olsen, and Barbara Beck. A bioinformatics track with outreach components. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):186–190, September 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Zweben:2005:CAN

[ZRY05] Stu Zweben, Han Reichgelt, and Gayle Yaverbaum. Computing accreditation: a new criteria structure and new flexibility. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(1):560–561, March 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Zweben:2006:OBC

[ZRY06] Stu Zweben, Han Reichgelt, and Gayle Yaverbaum. Outcomesbased computing accreditation criteria. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):129–130, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Zhang:2004:COC

[ZSW04] Xuesong Zhang, Ken Surendran, and Ming Wang. A computer organization course project: simulation of a modern traffic signal system. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(2):68– 71, June 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp://ftp.math.utah.edu/

pub/mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/
2004.bib.

Zander:2009:LSN

[ZTS^{+}09] Carol Zander, Lynda Thomas, Beth Simon, Laurie Murphy, Renée McCauley, Brian Hanks, and Sue Fitzgerald. Learning styles: novices decide. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3):223–227, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Zur:2004:TCC

[ZV04] Ela Zur and Tamar Vilner. Teaching certificate in computer science: didactics workshop. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(3):240, September 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).